A Complete Bibliography of Publications in Fisheries Research (2010–2019)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

12 June 2023 Version 1.04

Title word cross-reference

2 [MCMK18]. 3 [KRB17, MML^{+}16]. ^{13} [FTJ^{+}16, MCN16]. ^{15} [FTJ^{+}16, MCN16]. ^{18} [KHM^{+}17, SKY^{+}17]. ^{86} [LRY^{+}15]. ^{87} [LRY^{+}15]. a [RGGdL13]. B_{MSY} [PS12]. · [SPMS17]. δ [FTJ^{+}16, KHM^{+}17, MCN16, SKY^{+}17]. F_{MSY} [PS12]. i [BOBP^{+}17]. L_{opt} [Sve15]. M [SLPT17]. N [PHB^{+}19b].

-dependent [LJJ^{+}13]. -dimensional [PHB^{+}19b]. -length [NPK^{+}10].

1 [HVJP^{+}11a]. 1-116 [Ano15t]. 1-cm [Cra19]. 10-year [PG14, RHOCB^{+}18]. 106 [BMF^{+}11]. 107 [CP12]. 110 [VRMF12]. 120C [WMB16]. 127 [Mun12a]. 133 [ASHH15]. 148 [TV14]. 150706-002900 [Cor17, LRM17]. 163 [HLNH15a]. 164 [HL16]. 166 [LKD18]. 167 [PGP^{+}16]. 16S [TdMWC16]. 170 [Cor17]. 171 [PVKH18]. 177 [TC16a]. 180-degree [BWK^{+}13]. 191 [HAB^{+}18a]. 1920s [Ame12, KCvOE13]. 195C [IIT18]. 1990s [Koo12].

2 [HVJP^{+}11b]. 2000s [KCvOE13]. 2009 [Ano10-27]. 2015 [BW17, SA16]. 2019 [Ano19a, Ano19n]. 208 [SKB^{+}19]. 210 [Tho19a]. 218 [WGC^{+}21]. 219 [Tho21]. 23^{o} [dADSRG10]. 29^{o} [dADSRG10].

34 [HL16]. 3M [NPGT17].

44^{o} [MFRM19]. 48^{o} [MFRM19].

54^{o} [CCM14].

6012 [PGJ10].

84 [KPL^{+}09].

90-degree [BWK^{+}13].

A. [BSSH13]. A1 [Ano15t]. A1-A2 [Ano15t]. A2 [Ano15t]. abalone [HHTT11, JIM^{+}15, May10, PBB11, WKT^{+}11]. abbreviata [ZM15]. ability [GHA13, Leg14, SP12]. Abiotic [KKA10]. Abramis [GDW^{+}15]. Abrolhos [AFR^{+}17]. absolute [FPM17]. Abundance [HdSB^{+}14, TSS11, AFOB^{+}19, ÁCDM^{+}14, BSS^{+}13, BB18, BR15, BNK14, CPG^{+}15, Cam15, CAMW11, CAMFMF18, CN18, CC19, CJG^{+}15, CR18, CK18, CVS11, DBSA18, FHB^{+}19, FO11, GMMW17, GRB14, HKB^{+}18, HNY^{+}19, HdED^{+}13, JSB19a, KJKK15, KKE^{+}12, LMB^{+}11, LBH15, LHSS18, LFG11, LW19b, LSL12, MWSN^{+}13, MBF19a, MFMM16, MQV10, NS16, NBG^{+}17, RDR^{+}18, RCLS18, RMB^{+}21, SWB12, ST13, SPS^{+}19, SSH14, SBJ^{+}19, SLCL15, TOA12, WAP^{+}12, WRR^{+}16, WOPB11, YJC13, ZH10, vPBWA17]. abundances [DPAB10, OALS11]. Acadian [PHRH16a]. acanthias [BSKT13, CPS^{+}12, HCF^{+}17, SPC^{+}13]. Acanthocybium [Zis12]. Acanthopagrus [BBM19, CLS^{+}11]. Acanthuridae [OTST11]. accelerated [LCZ^{+}19]. acceleration [WWG^{+}16]. accelerometer [LBCD18]. acceptance [KOD^{+}15]. access [CNJ11, DFAN12, LMHR19, TD19]. access-point [LMHR19]. accordion [TLLZ19]. accordion-shaped [TLLZ19]. account [MPA^{+}12, POM^{+}18, SWC14, SPC^{+}13]. Accounting [BJZB12, DB12, SG15, TW13, TW14, WM13, WKTD13]. Accumulation [ZMG^{+}14, LCWC17, USH^{+}13]. Accuracy [FWGM19, JSB19a, TFB^{+}12, VB15, BZ12, Cra19, Fey18, Sam11, TW10]. Accurate [SGA17, KKE^{+}12]. achieve [BSF^{+}19]. achieving [BAM16, GCL19]. acid [RBG^{+}19a, SRG17]. acidification [RDM^{+}15, SVPC16]. acids [MCP14]. Acipenser [CGKG^{+}17, MLC^{+}19a]. Acknowledgement [Ano11a]. Acoustic [CMB11a, EFWC^{+}19, ETMK13, FBGB16, KR12, LE14, RBGE^{+}10, SHPSK15, ÁCDM^{+}14, BMCS19, BST^{+}17, BCP^{+}12, Ber13, BNA^{+}19, BTB^{+}19, BM19a, BPV^{+}17, CT16, CLP^{+}19, CBD^{+}14, DMC^{+}12, DOO15, EFW^{+}11, FPC^{+}13, FKM11, GSP17b, GBA^{+}18, GHJ13, HMA18, KBT^{+}19, LR19, MBF19a, MKH16, NSM^{+}18, NS16, ODD16, OCL^{+}18, ÖLTA19, Peñ18, PBC^{+}13, SOY14, SMI17, SHM12, SLCL15, TFB^{+}14, WAB^{+}16, WRR^{+}16, WTS^{+}17, WHP15, WBB^{+}16, YGA^{+}12, ZKSF17, ZBB17, dKMS15]. acoustic-tagging [BNA^{+}19]. Acoustically [BOBP^{+}17]. acoustics [KRTG16]. acronotus [BLHS11]. across [BLW^{+}18, BWN^{+}12, CLP^{+}19, CPD^{+}19, ESF^{+}19, FPM^{+}18, GHE^{+}11, LMB^{+}11, LSM^{+}19, LGH^{+}21, MMJ^{+}12, PBK^{+}19, PSJ^{+}19, RGK19, SO18, SLL^{+}16, SF13, WWLY15, WW14]. Act [PSS14]. acting [TXH^{+}19]. Action [RM19, YRH15]. active [KGH^{+}12, MKHK19]. actively [KW10]. activities [EMOLS^{+}15, KSS^{+}16]. activity [BKM13, DBG19, EP17, GDB19, KS17, LDR^{+}12, MCP^{+}17a, MNCPMG15, NCP15, PG14, PMK^{+}10, PGL^{+}13, RPP^{+}11]. actors [MSD^{+}16]. actual [DHC19]. aculeatus [EE19, RvAGB18]. acusticus [MSV^{+}15]. acute [LBH15]. acutispina [JLLL11]. AdaBoost [LJH11]. Adams [CMM^{+}11]. adaptation [MKH16, NNP17, RCF^{+}17, WWLY15]. adaptations [LS19]. Adapting [Whi10]. adaptive [HC17, HMGH12, IB10, SBLB17, YJSR12]. additional [ZW18]. additive [GDA14]. address [BSSH13, MH14, PCP^{+}18, PHLT16]. Addressing [FL13]. adequate [WM17]. adipose [MMMPP19]. adjacent [DSC19, Jab18, TOA12]. adjusted [BPMGP10]. adjustment [SWC14]. Adopting [BGW^{+}16]. Adriaen [BR15]. Adriatic [CMM^{+}11, PBLT14, GPSF10, GBPC10, MCM12b, PSM^{+}18]. Adult [APB^{+}16, ACRM12, APD^{+}18, CGG21, DCP^{+}16, DDP15, DHP^{+}11, FCNB16, GHE^{+}11, GLB^{+}19, HCWS19, JB15, JTHT14, KKE^{+}12, LWSC11, LSWR17, MK17, RBM13, TRH16, ZCLB11, ZDZY19]. Adult-mediated [APB^{+}16]. advanced [ER18, KMKK10]. advanced-generation [ER18]. Advances [ÁCDM^{+}14, BHP^{+}19, GBB^{+}12, MSRM^{+}14]. Advancing [PM13]. advantages [MDH15]. adverse [FBB^{+}17, SKF^{+}10]. advice [BDC^{+}17, KRTG16, PSE15, RLC^{+}17, SHJ19, WMdS14]. Advisory [Ano10e, Ano10f, Ano10g, Ano10-27]. Aegean [KTL11, DV11, DÖA^{+}10, KDPT18, RDP^{+}17]. aeglefinus [CPOH14, DH11, DRE^{+}17, GIW16, Jaw11, KDSAH^{+}18, KHM^{+}10, KHMF11, LDR^{+}12, LHS^{+}18c, MABA12, PBM^{+}18, RRM16, RRRM19, SHG^{+}16, SBH^{+}16, SKK^{+}13]. Aeolian [BRC^{+}10]. aequipinnis [CPH12]. aerial [GBR17, SBA10]. aestivalis [BSSH13]. affect [BKLB14, DMFR16, FPSS15, FS19, GCDAS^{+}10, HCSM17, HAB^{+}17a, HAB^{+}18a, HPNA16, JCT^{+}16, KLH^{+}15, RSEERRSC16, SHJ19]. affected [MRLD18, WKT^{+}11]. affecting [BKS18, GH15a, HHCP13, MTC^{+}17, Pul17, SMO^{+}18a, SBLL17]. affects [APB^{+}16, BGBCM19, BMEGV^{+}19, MEMdU^{+}17]. affinis [TEGM^{+}11]. AFLP [dJSMROZP^{+}13]. AFORO [PBML^{+}10]. Africa [BFL^{+}14, SNOD^{+}13, BMdS^{+}15, CBS17, HND^{+}17, KEMM12, MCDP18]. African [Mun12a, BCPP17, But16, KMdS^{+}18, LBHP16, MM13, MGR13, MKM^{+}16, Mun12b, Mun18, TWCA15, WRH^{+}15]. after [BDM^{+}18, BBM19, CFPF^{+}15, GHE^{+}11, GSØT15, HV10, LBB^{+}13a, LCL12, Lyn14, MSRB15, MWR18, MMK16, OJT^{+}14, SUR16, SMPB17]. against [BdSFB^{+}18, MCB^{+}18]. agassizii [VRW11]. Age [AVA18, ARMLC19, AFR^{+}17, AVB^{+}16, BPRB^{+}14, BLHS11, CH11, DTC16, FHK17, GCK14, HFPL16, HLA19, HSB^{+}13, KHM^{+}17, KPF^{+}18, LBV13, LCA^{+}19, MSFŠ^{+}11, MMBD13, MLC^{+}19b, NML12, OAPS13, PRBGS^{+}10, PBB^{+}15a, SLTPC^{+}18, Sim15, Smi14, SYL^{+}15, WKT^{+}13, AB12, AENC14, BK12, BNK14, BKS18, CH15, CA10, CBMLB13, Cop13, CTWD15, CAC^{+}14, CV14, DS13, DLCK15, EWB^{+}18, Eth15, EMSC15, FPS11, FBM^{+}19, FdSM^{+}16, Fra16b, GS18, HCHA12, HST10, HBG14, HSB19, HLB^{+}15, Hor11, HSH15, JG12, Jaw11, KMH^{+}12, KA15, KHG^{+}17, KWR15, KVH15, KTT13, KOHF^{+}16, LPTK19, Leg14, LHL^{+}15, LH15, LSJD11, MMF^{+}17, MHK^{+}17, MOA^{+}16, MNGPMdMG10, MWR13b, NPK^{+}10, OTST11, ODdA^{+}10, PHE16, PLM16, PACF17, QQG14, ROCC13, Rod19, RTSV15, SPM^{+}19, SDC^{+}15, SFPB17, SHC14, SPM^{+}10, Ste11, SMBP^{+}12, TT14, TMV16, TRW19, TBSE17]. age [USH^{+}13, VSM^{+}18]. age- [DS13, FBM^{+}19, PACF17]. age-0 [DLCK15, HSB19, ROCC13, USH^{+}13, VSM^{+}18]. Age-based [KPF^{+}18, NML12, SLTPC^{+}18, BNK14, KWR15]. age-class [Ste11]. age-composition [CV14]. age-frequency [BKS18]. Age-related [FHK17]. age-structured [CA10, CTWD15, EWB^{+}18, Fra16b, Hor11, MOA^{+}16, TT14, TRW19]. ageing [BB14, CNWM12, SLL^{+}16, TGM13]. agency [ABV16]. agenda [UARC11]. agents [BHT^{+}19]. ages [IAS^{+}17, KBSK15]. aggregate [BCC15, SP15b]. Aggregating [GSDD11, JPM11, SCV^{+}12, EBB16, EMP^{+}10, GBR17, MDC^{+}16, TGDB16]. Aggregation [RDV11, RBV^{+}13, EPTG19, EAM^{+}14, GMJ^{+}13, GDHA13, HF15a, RSG^{+}11, SMRS15, TCS13]. aggregation-fishery [SMRS15]. Aggregations [ALBŠB^{+}15, ADH^{+}13, CMB11a, CT16, EMMM13, IIS10, MMGEMGS10, SLCL15]. aggregative [KAiT^{+}13]. Aging [CM12, KM15, dZB16, dSMSF15, BBF^{+}18, SVN^{+}13, SGA17]. Agreements [HBC^{+}12, BK10, PLGL18]. aid [HKRH12]. air [CDMC18, FWR10, LLM^{+}19, RSQ18, SLT^{+}19, SOY14]. al [BKS^{+}15, Han16a, Han16b, HL16, HM16a, HM16b, Sve14, Sve15]. al. [HAB^{+}18b, Ogl17, PBB^{+}18, RC17]. alalunga [CDL21, CAA^{+}14, CAA^{+}15, LKL^{+}12, MFC^{+}13, MMM^{+}10, SMdUK16, WKT^{+}13, XTP^{+}16, dZB16]. Alan [Hol11b]. alarms [LE14]. alascanus [Ech17]. Alasdair [Ano10a]. Alaska [ST13, APD10, CL17, CS18, CRW^{+}19, CVS11, DMK^{+}19, Ech17, HPP18, HCK15, MFHK15, MCJD^{+}17, OS11, SZ11, SZPV16, WDNB^{+}19, WTS^{+}17, YRH15]. Alaskan [AENC14, CDB^{+}14, PHE16, SZPP17]. albacares [CHL11, SFB11, SHC14, SRMP^{+}18, VIFLP17, WCC^{+}21, XLCMMV19, ZCZ^{+}13, ZMG^{+}13, ZMG^{+}14]. albacore [ABSI^{+}18, CDL21, CAA^{+}14, CAA^{+}15, DBA18, LKL^{+}12, MFC^{+}13, MMM^{+}10, PBR^{+}16, SMdUK16, WKT^{+}13, XTP^{+}16, dZB16]. albatrosses [GJS^{+}13]. albidus [TAS12, WRP10]. albimarginatus [SCT^{+}17]. Alboran [MGB10, UGGR^{+}17, VSLC^{+}19]. Albula [MCD^{+}13, SSC^{+}12, AD16, WBM17]. albus [BMB^{+}18, MMD^{+}11, RUMBA15]. alcohol [Fey18]. alert [GPT^{+}19]. Alewives [AL13a]. algae [GIK^{+}16]. Algarve [PVMG19]. Algerian [BA12a]. algorithm [Ogl17]. algorithms [Tho19c, Tho21]. allocating [PD18]. allocation [BNAE10, BGB^{+}19, JWM^{+}12]. allometric [UK14]. allometry [IO11]. allow [CFF^{+}16]. alloy [HWS^{+}12]. Almejas [MMdlP16]. alone [SLB^{+}12]. along [BB18, BMdS^{+}15, BK16, BC18, CMMVRF^{+}19, CPG^{+}11, DSC19, GPP^{+}16, HBM11, HKB14, JBKA15, JWZ^{+}19, KMH^{+}12, KMKK10, KLB17, LLH^{+}14, MAH14, MKM^{+}16, MLP^{+}14, SCJ^{+}11, SAH15, SBT^{+}18, TWM^{+}10, WDL17, WJC17, YZW^{+}19, ZQZ^{+}17]. Alopias [HAB^{+}10, SHA^{+}15, SWA19]. Alosa [BSSH13, BNA^{+}19, BPV^{+}17]. alosine [CKC13]. alosines [WCS^{+}12]. alpha [JWZ^{+}19]. alpinus [CLG^{+}14]. alter [JHT^{+}10, RFC^{+}17]. alterations [DBB19]. altered [BDHA11]. Alternative [QEH^{+}11a, BNK14, CAM^{+}14, CDL21, DHP^{+}13, FHK18, KHV^{+}12, LJH11, TC15, WXZ^{+}18, WP17]. alternatives [ASH^{+}17, Gan12]. amalgamation [SB16]. Amazon [CMB11b, CAC15, FCIB17, ICSR16]. amberjack [JDNP18, ŠBMG^{+}16, SMP18]. ambushing [KS12]. Ameghinomya [MREV10]. America [ALR^{+}17, CEW14, HW11, MSRM^{+}14, MCJD^{+}17, PRCGD16, RCHSN^{+}15, SGSFL19, VRW11, VF10]. American [OTST11, Ano13a, BW17, BBT^{+}15, BSSH13, BC18, Bow17, BPG^{+}17, CJG^{+}15, GMMW17, KCM19, KRB17, LCRW19, MBS14, MSR14, MJL^{+}17, PWA10, PQ10, PGL^{+}13, Qui17, SMO^{+}18a, WFHG^{+}17, ZCC11]. americanus [BBT^{+}15, BC18, Bow17, BPG^{+}17, PWA10, PGL^{+}13, SBC^{+}18, TC16a, TC16b, TMM19, WFHG^{+}17, ZCC11]. amino [RBG^{+}19a]. Ammodytes [JvDB^{+}17, WRG^{+}18]. among [BBA^{+}19, BBH^{+}12, FCLH19, HKF^{+}12, KAGO^{+}10, LSH13, OH12, PWA10, PGS^{+}11, RMGK18, SMGH17, TMM19, VvHE^{+}11, WT16, dVA18]. amongst [MVML18]. amphidromy [ALR^{+}17]. amplifiable [LQL^{+}15]. anadromous [BGL^{+}13, CKC13, KBEB13, LSJ17, RPH^{+}18]. analis [GDHA13]. analyses [AdM19, AC17, CK18, DTC16, FSBS11, GRGGCACG11, KHG^{+}17, LJH11, MBvDN14, MKM^{+}16, Mun18, PGW17, PPLK18, PSE15, SFB11, SCT^{+}17, SCF17b, VRS13, VSM^{+}18, VSLC^{+}19]. Analysing [KKP11]. Analysis [ASL^{+}12, BBT^{+}14, CRS^{+}19, DPB^{+}14, DM11, EPO10, HWS10, LRM^{+}10, LBMC15, SH13b, AB11, dSMSF15, ASH^{+}17, AHH^{+}12, APHL13, BAY^{+}12, BCTS11, BHPC14, BGL15, BCMS^{+}18, BSD^{+}13, BJKH12, BHC^{+}16, BGA11, BFL^{+}14, BBH^{+}14, BBLS17, BLPF^{+}15, CCC^{+}14a, DBMR17, DDP15, DBA18, DSB^{+}10, DF16, ER18, FTJ^{+}16, Fit12, FQT17, FdMCB^{+}19, FEGMTÁ17, GAM10, GMNPM^{+}13, GQ16, GBK^{+}13, HFCPSEM15, HBB^{+}12, HVJP^{+}11b, IHFAH17, KNSPK13, KBG^{+}14, KBOM14, KWS^{+}15, KPL^{+}07, KPL^{+}09, LSWR17, LMPM11, LCMTM10, LCMdSM13, LR19, LCST10, MT13, MLBTM16, MSCP17, MP13, MMG10, MBG^{+}10, MBNM10, MFVPC19, NMH^{+}11, OGSNH^{+}17, OPR12, OPH15, ODdA^{+}10, PBM^{+}19, PRBGS^{+}10, PBC^{+}16, PAM^{+}18, PGS^{+}11, RLJ^{+}12, RFPG^{+}16, RCE^{+}19, RM19, SCJ^{+}11, SKY^{+}17, SAH15, SJJ^{+}11, SBLB17, SAK^{+}17a, SMO^{+}18b, Smi14, SKH14, TC15, TDH19, TLCSG^{+}15, TCS^{+}15]. analysis [TÜM^{+}16, UGGR^{+}17, VIFLP17, VRMF10, VRMF13, WLZ^{+}15, WHF^{+}14, WFNC18, WvPLW18, ZDZY19, vPHG11, SUR16]. analytic [PD14]. Anarhichas [GH14, GBEP16, Gun17, GSB^{+}19]. anchoita [DPO^{+}11, SSB^{+}18]. anchor [HHM^{+}13, NFK^{+}17, NSM^{+}19]. anchored [JPM11]. anchovies [KJKK15, RGAP18]. Anchovy [BJD12, BA12a, BGF^{+}15, BKM13, CSP^{+}18a, DPO^{+}11, FPC^{+}13, GGB^{+}18, GPBA13, JBKA15, KBG^{+}14, MSDT16, MSD^{+}16, OÁPP12, PB14, QMAM13, SSS^{+}12, TOTI10, UASM12, XYL^{+}19, YA18, ZKSF17]. ancient [SIBR^{+}19]. and/or [Ano13a, BBM19]. andersoni [CJG^{+}19]. Anegada [CBC^{+}10, LCB17]. anemone [FHH^{+}19]. anemonefish [FHH^{+}19]. anesthetic [OPF15]. angelshark [BdSFB^{+}18]. angle [RSEERRSC15, dKMS15]. angled [BBC11, BBH^{+}12, LBCD18, MWC15, SFL^{+}14, SSC^{+}12]. Angler [BKM19, CMP^{+}14, Cur18, GFJ19, PMG18, SCLS15, BNSS19, BAFP18, CAC19, CB17, HF14, JASP11, KL16, LWCC15, MAP14, PCWM16, SRB19, SJA17]. angler-related [MAP14]. angler-tagging [HF14]. anglerfish [HLB^{+}15, LBV13, OAPS13]. anglers [GRS^{+}15, KLB17, LWSC11, MCB^{+}18, MSCP17, MNCPMG15, PPH^{+}17]. Angling [GBF11b, AA13, BDHA11, BDB^{+}16, BMR^{+}19, BCPP17, CTSD10, CC11, DP18, DHP^{+}11, FL11, GSØT15, JHT^{+}10, KKA10, LLM^{+}19, MCDP18, MWR18, PWB^{+}16, PHB19a, SGA^{+}18, SMGH17, TWCA15, TGE^{+}18, VGE11, WSF^{+}18, vPCWV15]. angling-dependent [CTSD10]. angling-induced [CC11]. Anguilla [BDHA11, CM15, KSAA13, KPF^{+}18, LT18, Sim15, VBR^{+}18b, VBR^{+}18a]. angustirostris [AGB19]. Anhui [ZXH^{+}12]. animal [BM19a]. animals [DMFR16, EFW^{+}11]. Anini [AFA11]. Anini-y [AFA11]. anisakid [HPNA16, TS10]. Anisakidae [CSP^{+}18a, GMCL^{+}18]. Anisakis [BPS^{+}18, CSP^{+}18a, CSP^{+}18b, LCM^{+}18, MGP^{+}18, PMC^{+}18]. Anisakis-free [BPS^{+}18]. Annual [MABA12, TEFAR12, ALJC19, BVN^{+}10, CCC^{+}19, DM16, EMP^{+}10, FBGB16, GGB^{+}18, HSM^{+}12, LGZC17, Lyn14, MSV^{+}19, MRT18, NWS^{+}11, SLCL15, USH^{+}13, WMPR17]. annularis [APAFMN10]. annuli [Cor17, LRML^{+}15, SLL^{+}16]. annulus [UIH^{+}18]. Anomura [TEFAR12, PHE16]. Anoplopoma [HCHA12, HKB14, JC18, KSH19]. Antarctic [CEAP18, Fit12, HKK18, KWR15, KWR19, KK15, MDPH17, NS16, OCL^{+}18, WRR^{+}16, WT16, YZW^{+}19, ZYSZ18, ZDZY19]. Antarctica [YZW^{+}19, FPS11]. antarcticus [FWR10]. antenna [PKWP16]. antennae [BAY^{+}12]. antennas [GBB^{+}12]. Anterior [MBK19]. anthropised [BDM^{+}18]. anthropogenic [BKT10, RMGK18]. anthropogenic-polluted [RMGK18]. anthropogenically [VBR^{+}18b]. anti [CSKB13]. anti-freeze [CSKB13]. antiqua [MREV10]. Antique [AFA11]. aper [Sta16]. aperiodic [LMHR19]. Aphanopus [PGS10, SMCR13]. appendage [EHGG16].

Applicability [NMLZ17, AMGH15]. applicable [Ano17k, Ano17l, Ano17m]. Application [BSL10, BGF^{+}15, BB14, CIS13, CRAVC14, CSMO14, HHH^{+}12, IGS^{+}12, KF18, KBCK16, LTHRU17, MW11, Mau12, PAH^{+}13, RLJ^{+}12, RJ16, RMB^{+}21, Sau14, SGT^{+}14, WMNC15, WTK10, YAH^{+}13, ARF^{+}17, ADH^{+}13, BSB^{+}15, BMOZ13, DPB^{+}14, FMD19, FDMH16, GM15b, HSS18, HE15, HRAK15, IB10, JSYS15, JF13, Jaw11, KZ11, LFG11, MAAW^{+}17, MCVS14, MDH15, MCP^{+}16, MK13, MLF^{+}10, MIÁ^{+}10, PM13, RCF^{+}16, SPB^{+}19, TGMM13, WFJ19, dEWB^{+}15]. Applications [FGS19, KK17, FMB^{+}19, GBB^{+}12, GFK^{+}12, KSAA13, Kle13, MRS15, Yee10, KMKK10]. applied [BNAE10, CAM^{+}14, DRGGVL11, EDG16, FBM^{+}19, MPA^{+}12, SSR18, SWC14, WDS^{+}11, XdA19]. Applying [CCC^{+}19, PACF17]. apportion [YAH^{+}13]. Approach [GGM^{+}18, RBR^{+}15, SPR^{+}19, ALH^{+}17, BHI14, BMG16, BDB^{+}16, BSKT13, CLP^{+}19, CAM^{+}14, CN18, CDL21, CMM^{+}11, CM15, CTS^{+}19, DPAB10, DLCK17, DSB^{+}15, DBSA18, FCOM18, FBY^{+}12, FSA^{+}16, FJWD17, HKAA18, HLS^{+}16, KSL^{+}17, KZ11, KJKK15, KSS^{+}16, KC18, KRC14, KOHF^{+}16, LBF^{+}16, LJH11, MVML18, MKK^{+}16, MMSUAGC14, MMNCR^{+}13, MBF^{+}17, MPA^{+}12, MNW12, MSV^{+}19, MB16, OPH15, PKRC13, PTSC^{+}16, PS13, RADR11, RLC^{+}15, RRW^{+}13, RSGR19, SVPC16, SZ11, SF13, SAWS18, TOTI10, TJJ14, TVCT12, TTMC16, XTP^{+}16, XDL^{+}19, YCQ11]. Approaches [MW11, CC19, CGECSEH10, KEH16, KEP^{+}19, LZX^{+}17, LKR11, Pun19a, SG15, SM14, SM17, TT14, WKA13]. appropriate [WP11]. appropriateness [TBF^{+}16]. approximate [XTP^{+}16]. approximation [PLM16]. April [Ano10t, Ano11r, Ano15n, Ano16n, Ano17o, Ano18m, Ano19q]. Apulian [DSC19]. aquaculture [AHH^{+}12, DUC^{+}19, HBD^{+}16, JNN14, RLC^{+}15, RDM^{+}15, TTT16]. aquafeed [GKOC^{+}16]. aquarium [MFKS17, OKAKO16]. aquatic [EFW^{+}11, HMD^{+}19, SGB^{+}10]. Arabesque [MNT^{+}17]. Arabian [SB15, AMAM^{+}17, EAAA^{+}18, FHB^{+}19, Jab18, RLM^{+}17, RU15]. Arafura [EP17]. arapaima [CAC15, LBCD18, LBCD18]. Arawa [KHQH15]. archaeological [DUD^{+}18]. archipelagic [NMLZ17]. Archipelago [Cro15, HL16, RGK19, SAR^{+}15, WWLY15]. archipelagos [FPM^{+}18]. archival [CAA^{+}14, CAA^{+}15, LMM^{+}17, RTML17, SFB11, SFH^{+}15]. archives [DUD^{+}18]. Archosargus [ALB^{+}18]. Arctic [CLG^{+}14, DMK^{+}19, Han18a, HBC^{+}17, SPM^{+}19, SHG^{+}15, SBH^{+}16, SKK^{+}13, TSHB19]. arctus [AGSF19]. Area [GBPC10, MDPH17, BHG^{+}17, BMF^{+}10, BMF^{+}11, BBH^{+}14, FL11, GBM^{+}18, GLG^{+}11, GHJ13, HGG13, IB10, KML11, LMB^{+}11, LA13, MMGEH15, ML10, MLV13, OGSNH^{+}17, RLC^{+}15, RMB^{+}21, SF13, TNKO19, WMC^{+}16, WCM^{+}17, WMM^{+}15, WJC17]. Areas [Ano13a, HSY16, ABD^{+}19, AA13, BFW13, CMPR^{+}16, CAMFMF18, DR13, DSG^{+}17, FCNB16, HJT13, IAJ^{+}11, JSR10, JDLF11, KRB^{+}15, LAIVA^{+}19, LHLR17, LSH13, LMN^{+}12, MCDP18, MCP14, ORMG11, PD18, PAX^{+}17, PBK^{+}19, PSJ^{+}19, PHLT16, RAGP18, TLCSG^{+}15, UEOA19, WSMS14, WGM^{+}11, WSG16, ZCLB11, HHCP13, RRSB13, WGDB13]. areas-as-fleets [WSMS14]. arel [Ata11]. Argentina [CAGVB14, MFRM19, RRW^{+}13, CBC^{+}10, LCB17, MREV10]. Argentine [TV14, ÁCDM^{+}14, BAL^{+}19, BS10, CCC17, DPO^{+}11, LRM14, MRLD18, RM10b, RLMM18, TV13]. Argentinean [CMB11a]. argentinus [AGB15, CCC17, HdSB^{+}14, LCWC17, RLSPdR14, SP10a, SP10b]. arguinensis [MCP^{+}17a]. argus [Ber13, HTB15, MMBD13, TLCSG^{+}15]. Argyrosomus [DUD^{+}18, FWG11, FWIB14, GQdÁdMGP^{+}11, MKM^{+}16, MNGPMdMG10, MNGPM^{+}12]. Arhynchobatidae [IIT18, IIT17]. ARIS [CMJ^{+}19, SBLB17]. Arraial [BAOP18]. array [EFWC^{+}19, KW10, PKWP16]. arrival [GA19, MHRD16]. arrow [PG14]. art [PHM16b]. article [HL16]. artificial [APDI13, AK11, BNSS19, BST^{+}19, CRCMP^{+}14, FMH13, FTWC18, FGK19, GMCL^{+}18, HLJ15, HFL14, INSN14, JBB^{+}18, KSL^{+}16, KSL^{+}17, LHS^{+}18a, LW19a, MBG^{+}10, ÖLTA19, RBV^{+}13, RPP^{+}11, SBAK17, SAWS18, TS11, WGS17]. artisan [SMO^{+}18a]. Artisanal [NGMA11, SOMO17, AMAM^{+}17, ARVACPMN16, AS10, Bai10, BRC^{+}10, CSMEO^{+}11, EPTG19, GDHA13, GBPC10, HD12, HGG13, HVJP^{+}11a, HVJP^{+}11b, LCB17, MBS14, MMK10, MGC^{+}11, MLV13, PPA^{+}15, PBG^{+}13, PFVGGM16, PNAL16, SBS^{+}11, SBD^{+}12]. Ascaridoid [PBM^{+}18, GBM^{+}18]. Ascaridoidea [PMC^{+}18]. aseasonally [HST10]. asepsis [JBMK13]. Asian [WWLY15]. asp [BRP^{+}16]. aspartate [BKM13]. Aspect [JSB19b, RSEERRSC15]. Aspects [DSJ13, FS19, BRC^{+}10, SAK^{+}17b, TXP^{+}16, VSTS10]. aspera [FPM^{+}18, NKW^{+}19]. assemblage [BST^{+}19, HKB^{+}18, SCV^{+}12]. assemblages [AGS^{+}16, BLP^{+}18, FCIB17, IIT17, IIT18, JSR10, JDBF15, NPGT17, NPRGTSR18, PWSP12, PLMT^{+}11, PMF^{+}16, RJKG12, STCE12, SKB^{+}18, SKB^{+}19, TKF11, WPS19]. assess [BGF^{+}15, CTS^{+}19, HCZ^{+}16, KK12, KSS^{+}16, KHV^{+}12, LCH^{+}18, MTH10, MCEP14, NMH^{+}11, RCE^{+}19, TTT16, ZG14]. assessed [CMM^{+}11, LNC^{+}15a, LNC^{+}15b, LBCD18, LJJ^{+}13, OOM15, VIFLP17]. Assessing [Ano13a, ASB^{+}13, BEM13, BHC10, BLHO18, BBH^{+}12, CBMLB13, CJG^{+}19, DDP15, DPL^{+}16, DLCO^{+}19, FSBS11, GBPC10, GRS^{+}15, GBR17, HSY16, HCK15, HAB^{+}17b, HVJP^{+}11a, HVJP^{+}11b, KGB11, LPHH11, LT16, NSM^{+}18, OOM15, PSL^{+}19, PSS15, SWB12, SCV^{+}12, SGB^{+}10, WDNB^{+}19, WMM^{+}15, CM15, FSA^{+}16, GKOC^{+}16, LG13, PTBT17, SSR18, SDWG18, SAB^{+}11, STCE12, STAN^{+}17]. Assessment [BHG^{+}15, BKSF17, CDB^{+}14, CAC^{+}14, HBT16a, HSS^{+}16, HJT13, KK15, MMJ^{+}12, MS12, RPK11, RBG^{+}19a, SK13, SHGL21, TXP^{+}16, VRMF11, VRMF12, VIP13, WGC^{+}19, WGC^{+}21, WCK18, ALB^{+}18, AWSLS15, APCF17, AAS19, ARF^{+}17, ASMD^{+}18, ABSI^{+}18, ÁCDM^{+}14, APB^{+}16, AF13, BA12b, BFHT16, BNK14, BJZB12, BDC^{+}17, BLR^{+}16, BDB^{+}16, BBHF10, BCPP17, CDR16, CCA16b, CCC^{+}19, CGKG^{+}17, CAM^{+}14, CJWT^{+}19, Coo13, CP11, CP12, CGECSEH10, DM16, DDP^{+}16, DBT15, DSW^{+}17, EWB^{+}18, EAM^{+}14, EL18, FRDC10, FBM^{+}19, Fra12, Fra14, Fra16a, Fra16b, Fra17, FDMH16, FEGMTÁ17, Gan13, GBH^{+}19, GVDMG12, Gri10, GYCL19, HPP18, HTB15, HWS10, HFPL16, HA12, HLSD21, HSS^{+}11, HLP15, IOT14, JSYS15, JCT^{+}16, KKK16, KWR15, KWF12, KSAA13, KH15, KEP^{+}19, KHMF11, KPF^{+}18, KOHF^{+}16, LST^{+}10, LMPM11, LMPM12a, LMPM12b]. assessment [LP18, LSC^{+}18, LTHRU17, LKR11, MBHGN13, MAM^{+}13, MAP14, MTV17, MBRDM18, MDS^{+}18, MD10, Mau11, MH11, MP13, MCVS14, MCP^{+}16, MP17, MT19, MLF^{+}10, MKH16, MW13, MBvDN14, MVMSdS19, MLT17, MMAA^{+}18, Nee15, NB19, OLK^{+}14, OS11, PHE16, PBB11, PVKH15, PVKH18, PM13, PHT15, PHLT16, PDF^{+}18, PC19, Pun19b, RADR11, RLC^{+}17, RJ16, RUMBA15, RU15, RSAB11, Sam14, SLH^{+}14, SPB^{+}19, SEG^{+}15, SHJ19, SRB^{+}18, SHT^{+}13, SM14, SP12, SP15b, Szu16, SCOC16, TSS11, TBF^{+}16, TT14, TJMT17, Tho19c, Tho21, TC12, UGGR^{+}17, VRMF10, VMB18, VS16, WMP^{+}14, WMNC15, Way13, WKT^{+}13, WP15, WWSM15, WDS^{+}11, WA15, WTK10, XLCC19, XYL^{+}19, YSS^{+}11, ZW18, ZSF11, ZLT^{+}16, dEWB^{+}15, dZB16]. assessments [AM12, CZW14, CPC^{+}17, CV14, CMLP19, DJD12, DS13, FMB^{+}19, FdSM^{+}16, HB15a, HB15b, HFBS15, HFO^{+}18, KOD^{+}15, KEH16, LMPM11, Leg14, Lor16, Mac13, May10, MOA^{+}16, NB14, NLPS^{+}13, PL18, PHFW14, PHM16b, Pun17, Rob15, SLPT17, SD13, SM17, TC17, Tho19a, Tho19b, WKTD13, ZHDS16]. assets [TD19]. assigned [FBF^{+}18]. Assigning [RPP^{+}11]. assignment [BCMS^{+}18, CMSG18, JWM^{+}18]. associated [BHT^{+}19, CAC19, CDMC18, CV14, GHL^{+}11, HRB14, LBB^{+}13b, LCH^{+}18, MST14, RR12, SMRS15, SCV^{+}12, TWCA15, WWLY15, XLCMMV19]. Association [GBF^{+}11a, EAAA^{+}18, HHH15]. associations [BLP^{+}18, ZCLB11]. assumed [CGM^{+}19]. assumption [CCC^{+}19]. assumptions [CV14, KH15, LH18, MABA12, OS12, Pun19b, THF12]. Asturias [MCJD^{+}17]. asymmetric [CRAVC14]. asymmetrical [MMSUAGC14]. Asymptotic [AM12, WSMS14]. at-haulback [CIS13]. At-sea [FPSS15, PGL^{+}13, CFBS16, KGR^{+}15, NWO^{+}15]. At-vessel [MSRB15, EBB16]. Atlantic [ABL21, BEM^{+}17, CCM14, FGCG^{+}17, GMUD11, IIT18, LAIVA^{+}19, MHC^{+}18, SSD^{+}19, TEGM^{+}11, TJROFS19, VFFG17, WRP10, AMQ^{+}15, ASL^{+}12, ASB^{+}16, AJOE17, AB11, AB12, ALH^{+}17, AFOB^{+}19, ALBŠB^{+}15, AGB15, ADH^{+}13, AFR^{+}17, AL13b, BB18, BOCÁ^{+}18, BAL^{+}19, Ben18, BHG^{+}17, BKM13, BSD^{+}13, BSSH13, BKS^{+}16, BDC^{+}17, BMR^{+}18, BHC^{+}16, BLHG17, BLHO18, BSKT13, CFB^{+}17, CLS^{+}19, CAM^{+}14, CSKB13, CDMC18, CIS13, CSFCA15, CHCS14, Cor17, CAA^{+}14, CKC13, DR13, DJW^{+}18, DRE^{+}17, DDP15, DBA18, DTC16, EPCB17, EMMM13, ERL15, FPM^{+}18, FBHH14, FCCSA15, FMF^{+}17, FBF^{+}18, FH15, FGS19, FFG^{+}16, dSFdSSS17, GSØT15, GGADSRVL15, GBM^{+}18, GWB^{+}17, GHA13, GWVH^{+}16, Goo16, GH14, GBA^{+}18, GBEP16, Gun17, GSB^{+}19, GHJ13, HRT^{+}21, HBP^{+}18, HKAA18, HDS^{+}18, HA12, HMMP18, HCGS19, HLB^{+}15, HPSG15, HLA19, IIT17]. Atlantic [JHL^{+}19, JB15, JPO^{+}19, Jaw11, JS14, JWM^{+}18, JHT^{+}10, JFT^{+}15, JBMK13, JTP^{+}14, KNSJJ11, KNSPK13, KBOM14, KR12, KTT13, Kri13, LRML^{+}15, LAB13, LHS^{+}18c, LCA^{+}19, LCM^{+}18, LWS19a, LLM12, LRAP19, LMSM14, LW19b, MBHGN13, MSGV11, MSS^{+}16, MCC^{+}17, MCI^{+}13, MD16, MGP^{+}18, MW11, Mau12, MM11, MAH16, MFMM16, MLBGV16, MQV10, MQCS^{+}14, MLC^{+}19b, MFS^{+}18, MFVPC19, MSR14, MEMdU^{+}17, MHC^{+}18, NPK^{+}10, NEJ^{+}18, NAR^{+}17, OLS15, OOTV12, OTA^{+}13, OKPL11, PKH^{+}11, PGLG11, PRP^{+}18, PBM^{+}18, PLSM^{+}16, PTM^{+}18, PGS^{+}11, PRR^{+}19, RGG14, RBV^{+}13, RMMSR19, RSAB11, RBG^{+}19b, Sal18, SMCR13, dSSdAdAA17, SC19, SPV19, SLTPC^{+}18, SAK10, ŠBMG^{+}16, SH13a, SJS^{+}17, SSJ^{+}13, SMO^{+}18b, SP15a, SLL^{+}16, SMNdC^{+}19, SDV^{+}16, SSD^{+}19, SVG^{+}17, Sta16, SRB^{+}18, dJSMROZP^{+}13, SPC^{+}13, SBC^{+}18, SLCL15, SKDO^{+}17]. Atlantic [TVCT12, TOA12, TSS11, TWM^{+}10, TIGCDL14, TAS12, TCBR17, UJLC^{+}19, VPC^{+}14, VdSS^{+}14, VSM^{+}18, VSLC^{+}19, VHS^{+}17, VOM12, VFdS10, WRMM18, WTK10, WAT^{+}13, WW14, WBB^{+}16, XdA19, ZC17, dEWB^{+}15, dVLB^{+}15, dZB16]. atlanticus [AL13b, CR18, GBA^{+}18, Ken14, KSKGR15, PHK13, PHK14, VRS13]. Atlantis [FSFP15, NST^{+}17, SDE^{+}18, TWSA16]. Atractoscion [HLSD21, RCHSN^{+}15]. atresia [Gun17, KFGN10, SKK^{+}13]. atricauda [CH11, CH14, CH15]. attachment [Han13, Han18b]. attacks [SMK^{+}13]. attempt [FHM10, RBG^{+}19b, SHC14]. attempts [MLT18]. attenuation [SOY14]. attitudes [BPS^{+}18, BAOP18, GRS^{+}15]. attract [MBJ^{+}16, WTS^{+}17]. attraction [LDS^{+}19]. attribute [SZ11]. attributes [Bai10, CB17, MRLD18, PHG^{+}18]. Auckland [Rob15, HB15a, HB15b]. auction [SKT18]. Auctions [Mar18, FGRIT^{+}12]. audax [SSP^{+}15]. August [Ano11z, Ano12s, Ano13l, Ano14p, Ano15t, Ano18v, Ano19r, Ano21c]. aurata [ALUFJ^{+}12]. auratus [BBH^{+}12, GCRV10, HSC19, JCW12, WFLP11]. austral [ZDZY19]. Australia [GSP17a, GSP17b, KCO14, PGP^{+}16, BWN^{+}12, BdGM^{+}11, BdGD^{+}16, BRW^{+}13, CMS14, CLB^{+}10, CCT^{+}14, DUD^{+}18, EP17, FPM17, FWG11, FBY^{+}12, FWJ^{+}16, FSA^{+}16, GS12, HHTT11, HSM^{+}12, HMvHN13, JCW12, JHCT11, LSM^{+}19, LGH^{+}21, LFG11, ML11, MBW12, MSWW17, NL11, NML12, NWS^{+}11, PGP^{+}15, PTW^{+}13, PHT15, PHLT16, PDH^{+}16, PDF^{+}18, RDM^{+}15, RKH^{+}17, SWH17, SB12, SB16, Ste11, WFLP11, WBI^{+}16, Way13, WKTD13, Zie12, ZGT12, ZLT^{+}16, ZCW11]. Australian [EHH^{+}15, RKH^{+}17, BGW^{+}16, BMB10, BBK12, BSC12, BSM13a, BBM16, BMH17, BMH18, CBB^{+}16, CY12, DAM10, Gri10, HFO^{+}18, HSC19, KSKGR15, LBB^{+}13b, May10, MLF^{+}10, PKHD16, PML^{+}12, TFB^{+}12, VvHE^{+}11, WDS^{+}11]. australis [ABL21, BKS18, BBM19, CRAVC14, MAS17, MCP14]. Austria [PEU19]. Autocorrelated [WWSM15, JCT^{+}16]. automata [Eid12]. Automated [PBML^{+}10, CLP^{+}19, GRVN10, PHM^{+}16a, QQG14, SBLB17]. automatic [BCN^{+}13]. Autoregressive [Tho19a, Tho19b]. autumn [HBMC12]. autumn-spawning [HBMC12]. AUV [SWB12]. AUV-borne [SWB12]. auxiliary [JF13]. Auxis [MEMdU^{+}17, PGAS15]. Availability [NKW^{+}19, CCC17, DPL^{+}16, SSB^{+}18]. available [TPM15]. average [KWF12]. average-length-based [KWF12]. avoid [CH15]. Avoidance [EPCB17, JRD^{+}13, GBK^{+}18, WAB^{+}16]. Avoiding [KHM^{+}10]. away [WTS^{+}17]. azonus [MNT^{+}17]. Azores [CEJI11, FPC^{+}19, PPBM13].

bachus [KJS12]. BACI [MMSUAGC14]. back [Cor17, FL13, LRML^{+}15, NvZH^{+}17]. back-calculation [Cor17, FL13, LRML^{+}15]. backscatter [DPB^{+}14]. bad [PCG16]. badionotus [HFCPSEM15]. baerii [CGKG^{+}17]. bag [PBD21]. bag-and-size-limit [PBD21]. Bahamas [HKB^{+}18, MCD^{+}13]. Bahia [OGSNH^{+}17]. bairdi [Mur19, YRH15]. Bait [LSS^{+}17, BNSS19, CSFCA15, DHFSS18, EPHF18, FCCSA15, IEL17, JEJS17, SMK^{+}13, SJS^{+}17, WLGS19]. bait-based [DHFSS18]. baited [BM19b, PWB^{+}16, TBS13]. baits [AK11, WW11a]. Baja [OGCBMR15, CSMEO^{+}11, FEGMTÁ17, SFB11]. Balance [TNKO19]. balanced [BGP^{+}16, HPS16, Pla18]. Balancing [DMK^{+}17]. Balaton [FHS^{+}16]. Balearic [CAP^{+}10, MMF^{+}17, MMNCR^{+}13, MMM^{+}10, UJLC^{+}19]. Balistes [BMCS19]. ballot [BAFP18]. Baltic [HNW^{+}18, Han16b, HM16a, HM16b, KPL^{+}09, MBvDN14, OALS11, SL12, BVN^{+}10, BKLB14, BTNA13, CN19, DHH18, EÖFPH17, FF10, Han16a, HTHK11, HHH^{+}12, HRUSR13, HPHR14, HPNA16, HMA^{+}16, JBD^{+}18, JS15, KNSJJ11, KNSSL11, KNRJ^{+}15, KHV^{+}12, KNB^{+}19, KPL^{+}07, LSŠ^{+}16b, MTH10, NNP17, PGVV16, RTSV15, RSV^{+}17, SFL^{+}14, WB16, WLGS19, WHMS11, WHF^{+}14, ZRH^{+}11]. ban [BRH^{+}16, CAR19, SCLS15]. band [HSB^{+}13, KA15, ZYSZ18]. banded [ACRM12, KKH^{+}17]. Bangladesh [TBF^{+}16]. Bank [AFR^{+}17, LBV13, MLT18, AHMS10, HJT13, HHSS17, HCR14, LHSS18, MSR14]. bank-side [MLT18]. banning [CGC13]. bar [MMAH15, SBD^{+}12, BCMS^{+}18]. Bar-HRM [BCMS^{+}18]. barbatus [ABD^{+}19, THG^{+}16]. barbed [BBM19]. barbel [PB11, SRB19]. barbless [BBM19]. Barbus [PB11, SRB19, ALR^{+}17]. barcode [BCMS^{+}18, TC12, VZASM^{+}15, FCOM18]. barcode-based [TC12]. barcode-HRM [FCOM18]. barcoding [BdSFB^{+}18, CGdGT^{+}17, HIH19, KTW^{+}13, KMdS^{+}18, Nij15, PF15a, PMM^{+}15, SPM^{+}15a, SAK^{+}17a, TTMC16, WDL17, ZQZ^{+}17]. Barents [PLA16, BLHG17, DH11, GSL14, GPS15, GSH^{+}15, HBC^{+}17, HSLB19, LHS^{+}18a]. barge [LDS^{+}19]. barnacle [VRMF13]. barndoor [LHSS18]. barotrauma [BBA^{+}19, BRS19, BPC^{+}12, BBH^{+}12, CPB^{+}12, HRB14, HSC19, RHB^{+}17, RSE^{+}14, RLFJ11, SGB^{+}10]. barred [EPHF18, ML10, MJL14]. Barrier [EvHRF10, HSS^{+}16, RMT^{+}18, HDW^{+}15]. barriers [WDM15, vPKF^{+}19]. bartramii [FTJ^{+}16, FCL16, KYO^{+}12, YCYC16]. base [Cor17, Sta16]. Based [DM11, FWJ^{+}16, HN16, AW17, AVA18, BCKG13, BNAE10, BVN^{+}10, BNK14, BLR^{+}16, BFL^{+}14, BZ12, BMS15, BMC10, CAM^{+}19, CFB^{+}17, CCA16b, CHT18, CCC^{+}19, CMP^{+}12, CDF^{+}13, CDL21, CRAV13, CMJ^{+}19, CV14, CB11, DHFSS18, DTC16, Eid12, Eid16, ETMK13, EMSC15, FPS11, FHM10, FJWD17, FF11, FSPWG08, FDMH16, GRGGCACG11, GFCK13, GSP17b, Gri12, GM15b, GMHPV^{+}10, GBK^{+}13, HWS10, HCZ^{+}16, HLS^{+}16, Hil11, HE15, HF15b, HLP15, HAB^{+}17b, HBD^{+}16, JDNP18, JAB15, KMW^{+}11, KSL^{+}16, KHGB14, KZ11, KM10, KWR15, KWF12, KTNA15, KRC14, KHMF11, KPF^{+}18, LMHR19, LBV13, LRML^{+}15, LSR16, LHSS18, LMPM11, LG13, LJB16, LCRW19, LKR11, MZC^{+}18, MVML18, MBRDM18, MD13, MGM12, MMF^{+}16, MSV^{+}19, MM18, MMO^{+}13, Mun18, Nee15, NML12, NBF^{+}19, ODMF17, PBM^{+}19, PL18, PMF^{+}16, PHB^{+}19b, PG14, PPBM13, PVKH15, PVKH18]. based [PHT15, PHLT16, RPK11, RCHSN^{+}15, RPP^{+}11, SCJ^{+}11, STY^{+}16, SSR18, SLTPC^{+}18, SDWG18, SMGH17, SHC14, SHG^{+}15, SHGL21, SZ10, SGC11, TJJ14, TGM13, TGMM13, TSS11, THHG18, TJMT17, TKF11, THG^{+}16, TC12, VRS13, VvPM^{+}18, VS16, VAB^{+}14, WRR^{+}13, WBB^{+}16, YSS^{+}11, ZW18, ZCC11, ZCR16, vPBWA17, GKMGO^{+}17, NL11, PBC^{+}16]. baseline [AC17, BV14, NGvdL^{+}13]. baselines [GWVH^{+}16, UP16]. basic [MBRDM18]. Basin [RLJ^{+}12, WAP^{+}12, DUC^{+}19, FCIB17, IdRH10, PKWP16]. basis [GMS12, HHH15, NCP15]. basket [GPSF10, MM13]. Basque [AAH^{+}16, AJP18, GAM10, PBG^{+}13, PCG16]. bass [AAN^{+}12a, BDB^{+}16, CBB^{+}16, CHA15, GCL19, HKM^{+}17, ML10, MJL14, MWC15, SGA^{+}18, SAK^{+}17b, SW19, WMF13]. batch [GRVN10, HTHK11, Kur12, SMdUK16]. Bathymetric [MMD^{+}11]. bathypelagic [JČK^{+}12]. Bathyraja [AENC14, NBG^{+}17]. batoids [Jab18]. battery [KBCK16]. batwing [MBSM15]. Bay [FBM^{+}19, GMMW17, HBT16a, KCO14, MMdlP16, MBJ^{+}16, MMO^{+}13, PBG^{+}13, SL12, TOTI10, BKM13, BJD12, CBC^{+}10, DTG13, JCW12, LWZ^{+}19, LCB17, MAAW^{+}17, MSD^{+}16, MMK16, MMR^{+}18, MMAA^{+}18, MIÁ^{+}10, ODdA^{+}10, PICK17, PCG16, RBT14, RvRW^{+}14, UASM12]. Bayesian [ARF^{+}17, CN18, CBO^{+}15, DPAB10, DLCK17, DSB^{+}15, FBY^{+}12, KY19a, LBZE17, LWS19a, MM14, MB16, OPR12, ST13, SCF17b, SHT^{+}13, TJJ14, TC17, UPS16, WCK18, WJ17, ZH10, vPCWV15]. bays [CBF^{+}14]. BC [PMG18]. BCS [MMdlP16]. be [Ibá14, JCT^{+}16, KNJK^{+}10, LMPM12a, Pla18, PS12, RHB16, SMRS15]. beach [DHP^{+}11, Gra16]. beaches [GJRR14]. beak [BPRB^{+}14, FTJ^{+}16, LFCC15]. beaks [HBG^{+}19, PRBGS^{+}10]. beam [BHL^{+}16, CMJ^{+}19, DDB^{+}14, GFK^{+}12, RBM13, RSEERRSC15, RSEERR^{+}16, RJKG12, SMPB17, vMWvOKvB14]. beaming [JSB19a, RSEERRSC16]. beams [dKMS15]. bed [AHMS10, PKO10]. beds [SBLM14]. before [HGG13]. Behavior [MST14, BHPC14, BBH^{+}14, CLP^{+}19, CSMO14, DMQ^{+}18, EAM^{+}14, FGRIT^{+}12, FSH^{+}15, HRT^{+}21, HMA18, LW19a, ML17, MFIG16, MNW12, RWRT15, SFB11, SBS17, WWG^{+}16, WWH13, vPBWA17]. Behavioral [KBL^{+}15, BMCS19, HJ12, RHB^{+}17, RFC^{+}17]. behaviors [ALR^{+}17, Sto12]. behaviour [AVU^{+}10, BKT10, CY12, CMS^{+}15, CHS^{+}17, CAA^{+}14, DHC19, DRE^{+}17, DHP^{+}11, FHB^{+}19, JPO^{+}19, JHT^{+}10, LWSC11, LWCC15, LBCD18, MK17, MCB^{+}18, MKHK19, NWL^{+}14, QMGP10, QCZC^{+}19, RMT^{+}18, RTŘ^{+}12, RDV11, RPP^{+}11, SDJ^{+}15, SLB^{+}12, TGE^{+}18, VBR^{+}18a, WKA13]. Behavioural [WWR18, MCB^{+}18, MLC^{+}19a, RHC^{+}12]. behaviours [BDM^{+}18]. Béjaia [BA12a]. Belgian [BHL^{+}16, CKVV16, RDV11, RBV^{+}13, SHGP11a, SHGP11b]. Belgium [LTV^{+}13]. Belhabib [BKS^{+}15]. Belief [UPS16]. Belize [BCKG13, GDHA13]. Bell [LHF16]. Bell-shaped [LHF16]. bellicosus [CMMVRF^{+}19]. belly [BRJ^{+}16]. Benchmarking [SAF^{+}15]. benchmarks [HF15b]. bending [OKB^{+}16]. benefit [BAY^{+}12, DSC19, KNSSL11, KNSPK13, KT18b]. benefit-cost [KT18b]. Benefits [SSB^{+}18, BGW^{+}16, BPPR16, HKAA18, JS15, LBHP16]. Benguela [JKS^{+}15, JKK^{+}16, NGvdL^{+}13, WMPR17]. benign [BSM13a]. Benin [SNOD^{+}13]. benthic [BST^{+}19, BMC10, CMB^{+}12, CDR^{+}11, MCB^{+}16, PSL^{+}19, PMF^{+}16, RRPG18, SGT^{+}14, WDS^{+}11]. bentincki [CJCN21, CCC^{+}14a, CCC14c, FE10]. Bering [APCF17, MRT17, MRT18, Mur19, AENC14, BHI14, NKW^{+}19, SP12, SP15b]. Beringraja [KHG^{+}17]. Bertalanffy [CM12, HHTT11, Ogl17, PLM16, PRL^{+}18, WHNS13, XTP^{+}16]. berteroana [CHT18]. Best [CHW^{+}18, MGR14, Nak17a, STAN^{+}17, DDK^{+}18, PHT15, Pun19a, SVN^{+}13]. beta [JWZ^{+}19]. better [CSMO14, KT18b, MEM15, SLP^{+}15]. Between [HN16, ASHH12, ASHH15, ANEA19, BTB^{+}19, CMPR^{+}16, CBB^{+}16, CNWM12, CH11, DH14, DJD12, dADSRG10, DS18, EMOLS^{+}15, FPC^{+}13, Fit12, FMF^{+}17, FKM11, GMNPM^{+}13, GHS^{+}19, GRB14, HBMC12, HL11, HH14, JG12, KNRJ^{+}15, KS17, KDM15, KCvOE13, LKL^{+}12, LBZE14, LYWZ15, LDM^{+}15, LWZ^{+}19, MCP^{+}17a, MWSN^{+}13, MSD^{+}16, MCP14, MCM12a, Mur11, PKWP16, PJR17, PLPN12, PGJ10, PBB^{+}18, PBK^{+}19, PPdR^{+}10, RDP^{+}17, SD10, SWG18, SMO^{+}18a, SSJ^{+}13, SBLM14, SRB^{+}18, SAWS18, SMPB17, TRC19, UK14, VSLC^{+}19, WWN^{+}15, WCS^{+}12, WMF13, YJC13]. between-reader [CNWM12]. beyond [GN12, MQM^{+}15, SCG^{+}12]. Bias [DM16, FHK18, HDS^{+}18, BV14, BDC^{+}17, Fra16a, FS19, GAR10, HH14, KOHF^{+}16, LNC^{+}15a, LNC^{+}15b, OPR12, PJM14, SM15, SWC14, TZW16, TK16, TFB^{+}14, WOPB11]. biased [KPF^{+}18, PWB^{+}16]. Biases [Szu16, VIP13]. big [KM10, KHG^{+}17]. bigeye [APL^{+}18, dSMSF15, FSH^{+}15, OTH^{+}19, SFH^{+}15, SWA19, SLH^{+}14, SZ10, ZCZ^{+}13, ZCD^{+}12]. Bigger [KT18b]. bighead [DS18]. Bight [RKH^{+}17, SMJB11, EHH^{+}15, Foc14, MHC^{+}18, RAG19, RHW^{+}16, SRB^{+}18]. Bijagós [Cro15]. bilinearis [BRJ^{+}16, CPS^{+}12, LLM12]. Billfish [CY12, WB15, PSS15, WPCO15]. billfishes [GH15b, GM15a]. Bimini [HKB^{+}18]. bin [MOA^{+}16, Szu16]. binoculata [KM10, KHG^{+}17]. Bio [GHP^{+}15, GMD^{+}13]. bio-economic [GMD^{+}13]. Bio-economics [GHP^{+}15]. biobanking [GRO^{+}18]. biochemical [PHE16]. biochemistry [HAB^{+}10]. bioclimate [TC16a, TC16b]. biodegradable [GHS^{+}19, LFS^{+}19, WLBB15]. biodiversity [MP14, WUGG16]. Bioeconomic [EHH^{+}11, HMGH12, Eid12, GD17, KPL^{+}07, KPL^{+}09, MMF^{+}16, PKHD16]. Bioeconomics [SVPC16]. bioenergetic [DPSR10]. bioenergetics [vPW16]. biologgers [LBCD18]. Biological [HMMP18, IAJ^{+}11, CMP^{+}12, CPH12, MBHGN13, MBRDM18, MMM^{+}10, MW13, OMI^{+}14, PBR^{+}16, RFPE^{+}12, SGN^{+}10, TLM11, VHS^{+}17, VSTS10, WRH^{+}15, ZCC11, ZD13]. Biologically [MMGEMGS10]. Biology [HGA^{+}12, OtHH^{+}10, AGSF19, AT10, BdGM^{+}11, BdGD^{+}16, BB13, CBC^{+}10, COC14, MVMdS^{+}17, OAT^{+}16, PGLG11, PEAC11, SSH14, SCL^{+}15, TAS12, VSM^{+}18, Zis12]. biomanipulation [JAJ^{+}16]. biomarkers [BBT^{+}15, LMBF13]. Biomass [PPBM13, WCK18, ARF^{+}17, AW12, ASR^{+}18, BWR^{+}18, CH14, CH15, CMB^{+}12, DCSW^{+}12, DMK^{+}19, FPC^{+}13, GSP17a, GYCL19, GMUD11, JCW12, JAB15, KY19a, KJ17, KWR15, Leg14, MSDT16, MBB13, MB16, NL11, NKW^{+}19, OPR12, PL18, PMP^{+}11, PDP^{+}11, POM^{+}18, RG18, RUMBA15, SZPP17, SGRS^{+}18, Szu16, WBI^{+}16]. biometric [GWVH^{+}16, THHG18]. biometric-based [THHG18]. biophysical [FBHH14]. Biosocioeconomic [TÜM^{+}16]. biosphere [EMOLS^{+}15]. biota [CDR^{+}11]. Biotelemetry [JDNP18, DMFR16]. biphenyl [MS15a]. bird [MGR13]. bird-scaring [MGR13]. Biscay [PBG^{+}13, BKM13, BJD12, DTG13, MSD^{+}16, MMK16, MMR^{+}18, MIÁ^{+}10, ODdA^{+}10, PCG16, RBT14, UASM12]. Bissau [Cro15]. bite [ASHH12, ASHH15]. bite-offs [ASHH12, ASHH15]. biting [KLB17]. bivalve [GJRR14]. bivariate [PLPN12]. Black [SYL^{+}15, CLS^{+}11, CMS^{+}15, CGECSEH10, HLB^{+}15, NVS^{+}18, OMB18, PGS10, RFPE^{+}12, RFPG^{+}16, SCL^{+}15, TLLZ19, WPCO15, WMF13, CJP13, GGB^{+}18]. blackfin [CR18]. blackfoot [SKH14]. blacklip [HHTT11]. blacknose [BLHS11]. blacktip [SCT^{+}17]. blacodes [CAFHG10]. bladders [TdMWC16]. bleeding [OJT^{+}14, SFL^{+}14]. Bleeker [LSJD11]. bleekeri [IIS10]. Bleue [LPC^{+}14]. Bloch [MLBTM16]. Blocking [JTHT14]. blood [HAB^{+}10, RR12]. bloodworm [MS12, SBA10]. bloom [GBK^{+}18]. blooms [GIK^{+}16, PBLT14]. Blue [PFJ19, ARMLC19, BWN^{+}12, BB13, CNS11, CAM^{+}14, CBO^{+}15, CW17, CIS13, CRAVC14, FPS11, FGS19, GFO^{+}16, GIK^{+}16, Goo16, HKRH12, HLA19, HMvHN13, JHCT11, KF18, KBOM14, KWS^{+}15, KRTG16, LPMM14, LBB^{+}13b, MCC^{+}17, MAS17, MQV10, MQCS^{+}14, MFS^{+}18, MFVPC19, ODD16, PBB^{+}15b, RHW^{+}16, STJ11, SHR18b, TOA12, TJROFS19, VdSS^{+}14, VFFG17, VHS^{+}17, WRP10, YSS^{+}11]. blue-green [GIK^{+}16]. Bluefin [ALBŠB^{+}15, RCF^{+}17, ASL^{+}12, ASB^{+}16, ALH^{+}17, Ben18, CFB^{+}17, Cor17, FTN^{+}14, HSD^{+}17, IAS^{+}17, LRML^{+}15, MT13, MMF^{+}17, Nak17a, OSO^{+}18, OAT^{+}16, PAH^{+}13, RBG^{+}19b, SKT18, SLL^{+}16, SVG^{+}17, UIH^{+}18, UJLC^{+}19, VSM^{+}18, VSLC^{+}19, WMB12, WMB16]. bluemouth [RMMSR19, SGN^{+}10, SNP^{+}12, SVN^{+}13]. board [Cra19, DRE^{+}17, LDR^{+}12, PCP^{+}18, TW10, VKM^{+}17, Ano10e, Ano10f, Ano10g, Ano10h, Ano10i, Ano10j, Ano10k, Ano10l, Ano10m, Ano11e, Ano11f, Ano11g, Ano11h, Ano11i, Ano11j, Ano11k, Ano11l, Ano11m, Ano11n, Ano11o, Ano12a, Ano12b, Ano12c, Ano12d, Ano12e, Ano12f, Ano12g, Ano12h, Ano12i, Ano12j, Ano12k, Ano13b, Ano13c, Ano13d, Ano13e, Ano13f, Ano13g, Ano13h, Ano13i, Ano13j, Ano14a, Ano14b, Ano14c, Ano14d, Ano14e, Ano14f, Ano14g, Ano14h, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano15k, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano16i, Ano16j, Ano16k, Ano16l, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano17g, Ano17k, Ano17l, Ano17m, Ano17h]. Board [Ano17i, Ano17j, Ano18a, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano18g, Ano18h, Ano18i, Ano18j, Ano18k, Ano18l, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i, Ano19j, Ano19k, Ano19l, Ano19m, Ano21b]. boards [BSC12, MBSM15, TFH^{+}15]. Boarfish [Sta16]. boat [BKS^{+}15, KSL^{+}16, LMHR19, SYJ^{+}11]. boat-based [KSL^{+}16, LMHR19]. boating [HMD^{+}19]. boats [MAY12, Pun18, YMA12]. bocaccio [HWS10]. bodied [PTTC18, WGC^{+}19, WGC^{+}21]. bodies [KBT^{+}19]. Body [BMF^{+}10, BMF^{+}11, BJKH12, EÖFPH17, GGB^{+}18, HKK18, IGSJBS^{+}15, KSAA13, Kni12, KTT13, KT18b, LSJD11, MM13, NGvdL^{+}13, ŘJP^{+}12, Rod19, RHB16, SSJ^{+}13, VKNK12]. Bohai [LSJD11, SLZG18]. bolts [Cam15]. bomb [DTC16, IAS^{+}17, KHG^{+}17]. bonapartii [IIT17, IIT18]. bonariensis [AMV14]. Bonaterre [MMM^{+}10]. bone [CFB^{+}17]. bonefish [MCD^{+}13, SSC^{+}12]. bongo [MBS^{+}15]. bonito [VSLC^{+}19]. Book [Hol11b]. boom [EGK15]. boosted [FF11]. bootstrap [MPA^{+}12]. boreal [GRB14, SVR16]. borealis [LHS^{+}18b, TMC19]. borne [GRO^{+}18, SWB12]. both [ASN17, FdSM^{+}16, FS19, GBK^{+}18]. Bothnian [SL12]. bottlenecks [GVMSCJ12]. bottom [AHC^{+}11, AAH^{+}16, BK15, BHS16, BLHG17, BLHO18, CMMF13, CPP^{+}10, CMP^{+}12, CHS^{+}17, CGR^{+}15, FPC^{+}13, FPSDB17, GH14, Hen11, KML11, LHS^{+}18c, LW13, LW16, LWH19, LHF16, MMR^{+}18, MC10, MMO^{+}13, Mur15, NWO^{+}15, PDBM12, SNK19, SHG^{+}15, SHG^{+}16, SMI17, SKF^{+}10, SABG12, YRH15, ZTC^{+}18, ZC17]. bottom-set [CGR^{+}15]. bottom-trawl [ZTC^{+}18]. boundaries [LMB^{+}11]. boundary [EP17, SF13, WJC17]. Bounty [ODD16]. Bowdich [VFFG17]. box [BAFP18, QDG17]. Boyle [BPC^{+}12]. Brachyura [RM10a]. brachyuran [FT16]. brackish [KT18b, PGVV16]. brain [LDR^{+}12]. brama [GDW^{+}15]. branch [MGR13]. brand [DDK^{+}18]. Bransfield [CEAP18]. brasiliensis [FdMCB^{+}19, GJSR10, SRMP^{+}18, TTMC16]. Brazil [PG14, RAG19, AHC^{+}11, BAOP18, CRS^{+}19, CH11, CH14, CH15, CMM^{+}14, dADSRG10, FP10, GCK14, HdSB^{+}14, LG13, LLH^{+}14, SAH15, SP10b, dPHS^{+}12]. Brazilian [BLHS11, BAOP18, CGdGT^{+}17, FP10, SP10a, SBS^{+}11, SBD^{+}12, TTF^{+}13]. BRD [PTM^{+}18]. breadth [TTT16]. break [PALC19]. breakage [KYO^{+}12]. bream [ACRM12, BBM19, CMPR^{+}16, GDW^{+}15]. breams [LMB^{+}11]. breeders [CFF^{+}16]. breeding [DHFSS18, RM10a, TEFAR12]. Breton [MM11]. Bridging [HPRB16]. Bristol [ERCG10]. British [TGE^{+}18, ZD13, BRGS19, EB15, FRDC10, GFO^{+}16, GHL^{+}11, HSY16, HHD^{+}11, JC18, WMC^{+}16, ZCLB11]. Broad [BSD^{+}13, LBH15, FCLH19, HFO^{+}18]. Broad-scale [LBH15]. broadbill [AMQ^{+}15]. Broadening [SSHM16]. broadly [HKRH12]. brood [TEFAR12]. brooding [ZQZ^{+}17]. brook [CPB^{+}12]. brown [AVB^{+}16, BGL^{+}13, CMMVRF^{+}19, Cur18, DCP^{+}16, KCO14, PEU19, ROCC13, SBAK17, SGT^{+}18, VLSP19]. Brownie [HE15]. Bruising [KRMW15]. BRUV [TBS13]. BRUVs [HNM^{+}12]. bubble [GLG^{+}11]. bubble-plume [GLG^{+}11]. bubbles [SOY14]. Buccinum [AMH19, BMR^{+}18, EHBK18, LBF^{+}16, MLM15, SPM15b, WJ17]. bucket [GFJ19]. bucket-list [GFJ19]. budegassa [HLB^{+}15]. buffer [BLHO18]. build [EE19]. build-up [EE19]. Building [BNAE13, MEM15, Sta16, UPS16]. Bulkley [TGE^{+}18]. Bull [LAIVA^{+}19, GHL^{+}11]. bullet [MEMdU^{+}17]. buoy [SWA19]. buoyancy [CSKB13, NNP17, OÁPP12]. buoyant [DMC^{+}12, SGB^{+}10]. buoys [LMSM14, OLM^{+}19]. business [JBMK14]. Buyer [FGRIT^{+}12]. by-catch [BBHF10, DV11, VLSP19]. by-product [SJS^{+}17]. Bycatch [ASD^{+}18, CBM^{+}19, FRDC10, HL10, LHS^{+}18b, PGS10, Son18, SZR^{+}11, AMAM^{+}17, ASHH12, ASHH15, ASMD^{+}18, BBM^{+}10, BGBCM19, BRJ^{+}16, BSSH13, BSN17, BBK12, CNS11, CW17, CSFCA15, CRH^{+}13, CL17, CHW^{+}18, CGR^{+}16, CKC13, CCT^{+}14, CSMJB11, DLCO^{+}19, EB15, FCCSA15, FMMW13, FPSDB17, GMS^{+}17, GWWW13, HSD^{+}17, HJLW11, HJ12, HLJ15, HB11, HB13, JHL^{+}19, KLH^{+}15, KHM^{+}10, KGB11, LAB13, LCBD12, LE14, LHS^{+}18a, LTV^{+}13, LJH11, LJB16, LT18, LW12, LW13, LW16, LW19a, LSS^{+}17, MTC^{+}17, MM13, MGR13, MGR14, PF15b, PTM^{+}18, PCWM16, QEH^{+}11a, QMAM13, RRO^{+}19, RRPG18, SHM^{+}16, SMK^{+}13, SGW16, SRB^{+}18, VVL18, WLGS19, WWR18, WDBH16, YRH15, ZCR16]. bycatch-reduction [CSMJB11]. bycatches [QEH^{+}11b]. bycaught [YRH15].

C [FTJ^{+}16, MCN16, Mur19]. C. [JBMK14, MSRB15, PTM^{+}18]. Côte [KAGO^{+}10, LPC^{+}14]. cabezon [MNW12]. Cabo [BAOP18, OGCBMR15]. Cabrera [ARVACPMN16]. Cadiz [SJR^{+}11, TRS12, DSJ13, GQdÁdMGP^{+}11, VSTS10, BJD12]. cage [KDdAA^{+}18, RBG^{+}19b, WHS15]. cainii [BdGM^{+}11, BdGD^{+}16]. Calanus [GLG^{+}11]. calcifier [SVPC16]. Calcium [SLL^{+}16]. calculate [MRT17]. calculated [CPG^{+}15]. Calculating [PDP^{+}11]. calculation [CAMW11, Cla14, Cor17, FL13, LRML^{+}15]. Caledonia [JDLF11, JDBF15]. calibration [KFGN10]. California [CSMEO^{+}11, EMP^{+}10, FEGMTÁ17, MDL19, MFRJ13, MBNM10, OGCBMR15, RHW^{+}16, SFB11, BBA^{+}19, CUL^{+}16, EMOLS^{+}15, FLW^{+}19, GRPEMEVV17, HSD^{+}17, HAB^{+}10, HLSD21, JD14, KMH^{+}12, LJJ^{+}13, MSDT16, ML10, MHB^{+}19, RGGdL13, SRSNC17, SCC^{+}15, SHA^{+}15, TVdLOG^{+}21]. californicus [CUL^{+}16, HHD^{+}11]. call [BAFP18]. callarias [KHV^{+}12]. Callinectes [CMMVRF^{+}19, MCC^{+}17]. Callorhinchus [MWR18]. Cambodia [TBC^{+}19]. Cambridge [Hol11b]. camera [BWR^{+}18, KSL^{+}16, RWRT15, SWB12, TFB^{+}14]. cameras [HPRB16, vPCWV15]. campechanus [CTSD10, FGK19, HPSG15, LKZ^{+}17, TS11, WGS17]. camtschaticus [HNS12, PHE16, WHNS13]. Can [BSF^{+}19, BHS16, CPC^{+}17, FCBS15, GH14, JHL^{+}19, JCT^{+}16, KNJK^{+}10, LMPM12a, MKHK19, PHLT16, RHB16, SMRS15, SP15b, VVL18, ASMP^{+}10, Ara15, BDC^{+}17, CSOV14, FSPWG08, HJ12, MMDH18, MBB13, PBB^{+}15b, PS12, SFBV17, vPW16]. Canada [BW17, CDMC18, ERL15, HHD^{+}11, KC11, LRY^{+}15, MCJD^{+}17, PTM^{+}18, PRR^{+}19, RW19, ZD13]. Canadian [KMJ^{+}12]. canal [VBR^{+}18a]. canary [HRB14, CAGTS^{+}10, PGS10, PGS^{+}11, TEGM^{+}11, TJROFS19]. Cancer [BLW^{+}18, JCA^{+}16, TMC19, ZD13]. candei [CMSG18]. Candidate [PBR^{+}16]. canicula [BMRC14]. cannibalism [Han18a, LLM12, UJLC^{+}19]. cannonball [MDL19]. canosai [CLS^{+}19]. cantharus [NVS^{+}18]. canyon [JČK^{+}12]. canyon-shaped [JČK^{+}12]. Cap [MLV13, NPGT17, NPRGTSR18, PRMKASR13, TCBR17]. capacity [AAN^{+}12b, GMZ^{+}14, PGZ^{+}15, QHSI18, WFHG^{+}17]. Cape [MM11, SAK10]. capelin [DH14, MHRD16, MPR18, SPS^{+}19]. capensis [HND^{+}17, JKS^{+}15, JKK^{+}16, WMPR17]. capriscus [BMCS19]. Capros [Sta16]. captive [CFB^{+}17]. captive-reared [CFB^{+}17]. captivity [CRAP^{+}11, JPSR^{+}14]. Capture [BHL12, CK18, DWH10, MC10, OTA^{+}13, BHT^{+}19, CY12, CFPF^{+}15, CHW^{+}18, Cro15, FL13, FWR10, GHE^{+}11, GVDMG12, GH14, GHL^{+}11, HRB14, HDS^{+}18, HZT^{+}19, HBD^{+}16, IEL17, KCBS13, KYO^{+}12, LS10, LWCC15, MSRB15, MFMM16, MWC15, MMAA^{+}18, OJT^{+}14, OKPL11, PALC19, RMT^{+}18, RHB^{+}17, RHC^{+}12, RPK11, SWA19, SBC^{+}18, VKNK12, VBdB^{+}18, ZG14]. capture-and-release [GHE^{+}11]. capture-based [HBD^{+}16]. capture-induced [RHB^{+}17]. capture-mark-recapture [GVDMG12]. capture-recapture [ZG14]. captured [APDJSU11, CIS13, EBB16, GH14, HAB^{+}10, LWSC11, PPA^{+}15, PGP^{+}15, PGP^{+}16, SHA^{+}15, VRMF11, VRMF12]. Captures [ACSAS^{+}17, BAT12, CAGTS^{+}10, PKWP16, PGS^{+}11]. Carangidae [ŠBMG^{+}16]. carapace [KT18a]. Carassius [FHS^{+}16, MLBTM16]. carbo [HL16, SAR^{+}15, SMCR13]. Carbon [OPF15, VRMF13, JD14, RBG^{+}19a]. carcass [PGP^{+}15, PGP^{+}16]. Carcharhinus [BLHS11, GTDJdL^{+}13, LAIVA^{+}19, LMM^{+}17, MSRB15, RPK11, SCT^{+}17]. Carcharias [KCBS13, OGSNH^{+}17]. Carcharodon [OGSNH^{+}17]. cardinalfish [CRAV13]. CARE [BKSF17, EHGG16]. Caretta [Mur15]. Caribbean [PWSP12, BdGN^{+}18, DSB^{+}10, MMGEMGS10, MMSUAGC14, MMBD13, MAH16, PRCGD16, TOA12, TLCSG^{+}15]. Caridea [CCA16a, CCA16b]. carinatus [LCA^{+}19]. carite [GJSR10]. Carolina [CW17]. carp [BJKH12, FHS^{+}16, KG10, KSAA13, OCN^{+}11, RHC^{+}12, WHB^{+}16]. Carpentaria [MSWW17]. carpio [KSAA13, OCN^{+}11, RHC^{+}12, RSEERR^{+}16, WHB^{+}16]. carponotatus [VvHE^{+}11]. carrying [AAN^{+}12b]. Casal2 [DLD^{+}16]. Case [BHM^{+}18b, BLHG17, ABV16, APCF17, AJP18, AVU^{+}10, BGF^{+}15, BKS^{+}14, BKS^{+}15, BPPR16, BAL^{+}19, BJZB12, CMMF13, CHT18, CAP^{+}10, CFF^{+}15, CBO^{+}15, CPCP^{+}12, CMLP19, DJW^{+}18, DP18, DQMV19, DFAN12, FDN^{+}11, GKOC^{+}16, GCDAS^{+}10, GAM10, GPSA11, GGADSRVL15, GSP17b, GBR^{+}18, GPBA13, GBK^{+}13, GBR17, HHD^{+}11, HFPL16, HF19, HWS^{+}15, HSRT17, HMGH12, HMA^{+}16, Ibá14, INSN14, JPM11, JFT^{+}15, JvDB^{+}17, KK12, LOdH16, LS19, LPC^{+}14, LYWZ15, LvVtB^{+}15, LHF16, MKK^{+}16, MCDP18, MMdlP16, MPB^{+}17, MPF^{+}17, MBvDN14, MCM12a, MCH11, NCP15, NAR^{+}17, OMB18, PBG^{+}13, PKWP16, PBC^{+}16, PEU19, PCG16, PHT15, PHLT16, RFPE^{+}12, RTML17, SMRS15, SB14, SP15b, TLLZ19, TRW19, TIGCDL14, TÜM^{+}16, VFMS16, WT16, WMM^{+}15, WP11, WKTD13, WFJ19, YRH15, ZM15, ZTC^{+}18]. case-study [LS19]. cases [KRB^{+}15]. Caspian [YVJ12, ZKSF17]. Catch [CHL11, FB19, LT16, MBNM10, OKAKO16, PWA10, SBLM14, SFBV17, TCAP11, TC15, YMA12, ZGT12, vMWvOKvB14, AMGH15, AMHR16, AVP^{+}15a, BNSS19, BBF^{+}18, BB10, BKS^{+}14, BKS^{+}15, BDC^{+}17, BJKH12, BDB^{+}16, BMR^{+}19, BK15, BLHG17, BT16, BSM14, BBHF10, BKLB14, BMC10, Cam15, CAMW11, CNS11, CWM12, CPW^{+}14, Car18, CFF^{+}15, CW17, CMP^{+}14, CC19, Coo13, CP11, CP12, Cop13, CSMO14, DV11, DDK^{+}18, DJD12, DM11, DTSG^{+}16, EPHF18, FCBS15, FHB^{+}19, FSG^{+}19, FJWD17, GSØT15, GLMK12, Gri12, GSL14, GCRV10, GWB^{+}19, GRS^{+}15, GRB14, HRT^{+}21, HFD^{+}13, HHH15, HAB^{+}10, HPHF15, Hen11, HSRT17, HVD^{+}18, HSM^{+}12, Hua15, HWS^{+}12, IB10, JIM^{+}15, JHT^{+}10, JASP11, JLF^{+}17, JČK^{+}12, JČK^{+}21, Kai19, KL16, KY17, KVK^{+}17, KLB17, KGB11, KKA10, LMHR19, LLWC15, LTV^{+}13, LA13, LJH11, LYWZ15, LWH19]. catch [LSS^{+}17, LČ18, MMGEH15, MCDP18, MCB^{+}18, MLC^{+}19a, MS18, MWC15, MGC^{+}11, NBS15, NSM^{+}19, OHLBL16, OPH15, OGCBMR15, PGS10, PSU^{+}16, PTSC^{+}16, PLPN12, PQ10, PHB19a, PHBU19, PCWM16, PMK^{+}10, PMP^{+}11, PDP^{+}11, RHC^{+}12, SGA^{+}18, SKFM17, SPMS17, SGT^{+}18, SHG^{+}15, SHG^{+}16, SLB^{+}12, SH10, SABG12, TSNRU21, TWCA15, TW14, TCS13, TTF^{+}13, TBSE17, TGE^{+}18, VLSP19, VAB^{+}14, WDL17, WCS^{+}12, WK10, WSF^{+}18, WLGS19, WP15, WFJ19, WKA13, WKA14, WvPLW18, YvZG^{+}17]. catch-and [PHB19a]. catch-and-release [BBF^{+}18, BDB^{+}16, BMR^{+}19, DDK^{+}18, GRS^{+}15, HRT^{+}21, HAB^{+}10, HVD^{+}18, JHT^{+}10, JASP11, LA13, MCDP18, MCB^{+}18, MLC^{+}19a, RHC^{+}12, SGA^{+}18, SGT^{+}18, SLB^{+}12, TGE^{+}18, WSF^{+}18]. catch-at-age [Cop13, TBSE17]. catch-at-length [WvPLW18]. Catch-at-size [MBNM10, TBSE17]. catch-based [DM11, FJWD17]. Catch-curve [TCAP11]. catch-curves [WK10]. catch-only [WP15]. catch-per-unit-effort [GWB^{+}19, WKA13]. catch-ratio [Car18]. catch-sampling [BMC10]. catchability [BAB^{+}12, CAM^{+}14, DQMV19, FHB^{+}19, Hol11a, KLH^{+}15, KY17, LSS^{+}16a, PGL17, SBLL17, TB10, WRR^{+}13, YRP^{+}19, vPWW16]. catchable [CM18, PMP^{+}11]. catchable-sized [CM18]. catche [SA16]. Catches [CAAB^{+}18, PVMG19, ASL^{+}12, BHG^{+}15, BSC12, BBM17, BM19b, CB11, DBB19, FTM^{+}15, FA19, GTCC15, GBD^{+}15, GK18, HHH15, HL16, LHLR17, LBZE14, LPC^{+}14, LHF16, MFHK10, MMDH18, PBB^{+}15b, SAR^{+}15, Sam11, SBD^{+}12, TD19, VKNK12, WKT^{+}11, ZXH^{+}12]. Catching [BDB^{+}11, BGL^{+}13, BSM12, BSM13b, SB17]. catchment [FHS^{+}16, PSJ^{+}19]. catchments [LST^{+}10]. categories [dMGPR^{+}18]. catface [CJG^{+}19]. catfish [DS18, GMSP19, SPH18, dPHS^{+}12]. catshark [BMRC14]. caught [ASN17, BBLD16, BLHO18, BNM11, BCPP17, CMS14, CFPF^{+}15, DRE^{+}17, FMMW13, FFG^{+}16, FEGMTÁ17, GWB^{+}17, GH15b, Gri10, GPdR11, HFD^{+}13, HBD^{+}16, Kai19, KCBS13, LBB^{+}13b, MC10, OTA^{+}13, RSAB11, RSQ18, SMNdC^{+}19, TTF^{+}13, TSHB19, TC12, VGE11, YAT^{+}19, YA18, ZZA^{+}14]. caught-and-released [BCPP17]. causality [RCE^{+}19, SBL^{+}19]. cause [FT16, LOdH16, MLTD17, SL11b]. caused [DM16, Fra16a, IOT14, MP18, OOTV12]. causes [GGB^{+}18, MSWW17, Wri13, YvZG^{+}17, ZXH^{+}12]. Cautions [HF15b]. CBA [HBD^{+}16]. Celebes [PGAS15]. cellular [Eid12]. Celtic [HBMC12, MDR^{+}19, SFR^{+}13]. censored [BDC^{+}17]. censoring [SG15]. census [GFK^{+}12, KS12, PLMT^{+}11]. Central [CHL11, CEW14, QMGP10, VRW11, AMQ^{+}15, AFA11, BGF^{+}15, BRC^{+}10, CCA16b, CHT18, CJCN21, CCC14c, DSC19, DV11, FCIB17, FE10, FMB^{+}19, FSH^{+}15, HLA19, Kai19, LNL^{+}15, MPB^{+}17, MWSN^{+}13, MST14, MAH16, MBNM10, PGLG11, PBC^{+}16, VdSS^{+}14, BLR^{+}16, CCA16a, CPD^{+}19, DPB^{+}14, FPC^{+}13, HW11, LČ18, MFRM19, NMH^{+}11, SLTPC^{+}18, SFH^{+}15]. central-east [PGLG11]. central-southern [MWSN^{+}13]. Centropristis [WMF13]. century [BR15, LTV^{+}13, VOM12]. cepedianum [JSB19b]. Cephalopod [Ano10-27, GAP10, CAGVB14]. Cephalopoda [LLH^{+}14, PPG^{+}11, STY^{+}16, SP10a, SP10b, TS10]. Cephalopods [GPdR11, OtHH^{+}10, PTF^{+}18]. Cephaloscyllium [WWR18]. certification [BGW^{+}16, BPPR16, BFHT16, BAM16, But16, GQ16, GDB^{+}16, LBHP16, MFKS17, PSU^{+}16, PRCGD16, PLA16, SUR16, SMD^{+}16, TTT16]. certified [ABV16]. Cervimunida [CN18, FQT17]. cessation [Ogl17]. cetaceans [Son18]. cf [KDdAA^{+}18, LBCD18, SGT^{+}18]. Chaboud [BKS^{+}15]. Chaceon [GMUD11, MD16, TEGM^{+}11]. Chaetodipterus [SSD^{+}19]. chain [NAR^{+}17, RFS16, vMWvOKvB14]. chains [ABV16]. chalcogramma [APD10, BHI14, Koo12, PPLK18, WWH13]. challenge [TC17]. challenged [EWB^{+}18]. Challenges [AVU^{+}10, GH15b, IWL^{+}11, JZL13, Via19, BGW^{+}16, BPPR16, BGB^{+}19, EHBK18, FT16, HN16, HSG15, HA12, KEH16, KEP^{+}19, PDH^{+}16, Wri13, vPB18]. chamber [HSC19]. chambered [FKM11]. Chamelea [DSJ13, PSM^{+}18]. change [BK10, CAM^{+}14, CHL11, CMB^{+}12, DHP^{+}13, EP12, GPT^{+}16, GPSA11, HCK15, LL14b, LT16, MAM^{+}13, NLPS^{+}13, OSO^{+}18, PTF^{+}18, PRMKASR13, Pun11, PSS14, Qui17, RRM16, TBC^{+}19, ZC17]. changeable [GPSF10]. changed [MS15b]. Changes [AS10, CLWB15, DPSRM10, FBY^{+}12, LL14b, OÁPP12, TIGCDL14, AJP18, ASB^{+}13, BRW^{+}13, CH11, CH14, CH15, CJCN21, CGLS12, CCC14c, HSRT17, HSH15, JB15, Kai19, KDM15, KJ17, KDdAA^{+}18, KGB11, Lam13, LBH15, MWR18, MMDH18, MSWW17, NNS^{+}10, NGvdL^{+}13, RJ16, SMBP^{+}12, VJH17, WMNC15, WDNB^{+}19, WM13, WHS^{+}18, ZXH^{+}12, vPW16, vPWW16]. Changing [BW17, SLZG18, BLHG17, DTC16, FCL16, GTCC15, GKKK15, Hol11a, LCM^{+}18, PGL17, ST16, UKK^{+}12]. Channa [DUC^{+}19]. Channel [SPM15b, CLP^{+}19, SPH18, ARF^{+}17, ERCG10, PRL^{+}18, RBM13, VKM^{+}17]. Chaohu [ZXH^{+}12]. Char [LRY^{+}15]. Character [Ber19]. Characterisation [GM15b, AMHR16, BEM13, WPCO15]. characterise [AMGH15]. Characterising [SWH17]. Characteristics [CL17, BWN^{+}12, BMR^{+}18, BHCS15, CBC^{+}10, CPH12, Lam13, LCM^{+}18, MBJ^{+}16, SHR18b, TMC19, VvPM^{+}18, WAP^{+}12]. Characterization [BRC^{+}10, FAAM15, JPSR^{+}14, ASMP^{+}10, BMOZ13, CMB11a, LKD15, LKD18]. characterize [AVP^{+}15b, CFF^{+}16]. Characterizing [Ber13, FTM^{+}15]. characters [AC17, FdMCB^{+}19]. charr [CLG^{+}14]. charter [GK18]. Chasing [MVML18]. Chemical [GMNPM^{+}13, KBL^{+}15]. chemistry [ASSL11, ALR^{+}17, CSOV14, DH14, FHK17, MDG17, RBM^{+}16, RR12, RSGR19, SAK10, TRC19, WRP10]. Cherax [BdGM^{+}11, BdGD^{+}16]. Chile [AOC11, CLWB15, CCA16a, CCA16b, CHT18, CJWT^{+}19, CJCN21, CN18, CCC^{+}14a, CCHD19, CCC14c, FE10, KA15, MWSN^{+}13, QEH^{+}11b, QWC10, QWCB11, RBGE^{+}10, RCBPK11, SMO^{+}18a, TSNRU21]. Chilean [MCN16, ASSL11, BLPF^{+}15, GPP^{+}16, LCZ^{+}16, MWSN^{+}13, MCP14, MMD^{+}11, QMGP10, QEH^{+}11a, QCZC^{+}19, RUMBA15, WFNC18, ZZA^{+}14]. chilensis [FDMH16, QWCB11]. chilipepper [HFBS11, LBS^{+}18]. China [FTN^{+}14, XDL^{+}19, YGZ^{+}18, ZzCbC^{+}15, JZL13, LSJD11, LCH^{+}18, LWZ^{+}19, LL14b, PTF^{+}18, RLJ^{+}12, TLLZ19, ZXH^{+}12, ZQZ^{+}17]. chinensis [SLZG18]. Chinese [MZC^{+}18]. Chinook [BHT^{+}19, BDB^{+}11, BWK^{+}13, CQ14, CBD^{+}14, DET^{+}16, HSC^{+}12, HKF^{+}12, LKNL16, LW12, LW19a, MBF19a, MBG^{+}10, RBG^{+}19a, SCC^{+}15, SO18, SCF^{+}17a, WAB^{+}16]. Chionoecetes [DWH10, DMK^{+}19, KC11, MRT17, MRT18, Mur19, NWL^{+}14, SP12, WW11b, YRH15]. chlorophyll [RGGdL13]. choice [WGM^{+}11, WDM15]. choices [MABA12]. Chondrichthyan [DV11, MPB^{+}17]. chondrichthyans [CCM14, HST10, IIT17, IIT18, KMdS^{+}18]. Chondrichthyes [BMRC14]. choosing [HH14, JATBC19, SVN^{+}13]. Choptank [DW19]. chromatophore [VFdS10]. Chromosomal [PRR^{+}19]. chronological [AVB^{+}16]. chronologies [KHM^{+}17]. Chrysaora [RT14]. Chrysoblephus [DJF^{+}15]. Chrysophrys [HSC19]. chub [HNY^{+}19, PKWP15, PKWP16, YGZ^{+}18]. chum [CHW^{+}18, KKE^{+}12, UKK^{+}12]. Chun [SSH14]. cibu [SPMS17]. Cichla [BDB^{+}16]. Cíes [GHUG^{+}16]. Circle [GCRV10, BBM19, CB11, HFD^{+}13, PKH^{+}11]. circulation [CAGVB14]. circumference [SL10, SL11a]. cirrhosa [RGG14]. cisco [FBM^{+}19, HSS18, MYJ^{+}15, SMS17]. citizen [FTWC18, SKB^{+}18, SKB^{+}19, WSF^{+}18]. clam [CMM^{+}11, CLF^{+}17, COC14, DSG^{+}17, EPO10, Gra16, MREV10, MTV17, TXP^{+}16, UGGR^{+}17]. clams [FMD19, MCB^{+}16, RvRW^{+}14]. CLARA [CPP^{+}10]. Clarias [DS18]. class [AD16, Eid16, FCLH19, OALS11, RCBPK11, Ste11, TCAP11]. classes [ER18]. Classic [CGECSEH10]. classification [BCN^{+}13, DPB^{+}14, FWGM19, IO11, MGB10, RCBPK11, SHPSK15, THHG18, YAH^{+}13]. classify [CWM12, POV19]. claw [DSW^{+}17]. claws [PQ10]. clay [SJR^{+}11]. Climate [ML11, TQ17, YGZ^{+}18, BPS^{+}19, BK10, CJWT^{+}19, GTCC15, GPT^{+}16, GKKK15, HCK15, LP18, LL14b, MAM^{+}13, MMGEMGS10, MBF^{+}17, MFIG16, NLPS^{+}13, PTF^{+}18, Pun11, Qui17, Tho19a, Tho19b, TKF11, Way13, WKT^{+}11]. climate-change [NLPS^{+}13]. Climate-coastal [ML11]. climate-driven [LP18]. climate-growth [MBF^{+}17]. Climate-induced [YGZ^{+}18, CJWT^{+}19, Way13]. climate-related [MMGEMGS10, MFIG16]. climatic [CAGTS^{+}10, DHP^{+}13, NEJ^{+}18]. clogging [TNKO19]. Close [BST^{+}19, RSEERRSC16]. closed [BFW13, PHLT16]. Closer [CCC^{+}14b]. closure [MBvDN14, SP15a, SMPB17]. closures [ADH^{+}13, CAC19, COS^{+}19, SKFL15]. cloud [BSM15a, LKL^{+}12]. cloud-free [LKL^{+}12]. Clupea [BHG^{+}17, BTNA13, HBMC12, JS14, LCHL11, LCL12, MGP^{+}18, NNS^{+}10, OALS11, SSJ^{+}13, SP15a, STJ11, SLCL15]. clupeaformis [JWM^{+}12]. Clupeiformes [PGAS15]. Clupeonella [ZKSF17]. cluster [Chi11, Peñ18]. clustering [Peñ18]. clutch [MFRM19]. Clyde [HSH15, HSH16]. cm [Cra19]. Cnidaria [MDL19]. Co [GHA13, DAJMB16, GDHA13, GBPC10, HRC16, PFVGGM16, VPSH^{+}17, WCM^{+}17, WUGG16]. co-exist [WUGG16]. co-managed [VPSH^{+}17, WCM^{+}17]. co-management [DAJMB16, GDHA13, GBPC10, HRC16, PFVGGM16]. Co-occurrence [GHA13]. coarse [CB17, DHC19]. coast [ASB^{+}16, ATG11, BA12a, BB18, BBM^{+}10, BOCÁ^{+}18, BH13, BKS^{+}16, BMdS^{+}15, BK16, BC18, BK15, CSMEO^{+}11, CRB^{+}18, CMMVRF^{+}19, CPG^{+}11, DSG^{+}17, dADSRG10, FCNB16, FE10, dSFdSSS17, FEGMTÁ17, GPT^{+}16, GVSB11, GJS^{+}13, HBM11, HL16, KMH^{+}12, Kri13, LW12, LW13, MAH14, MBF^{+}19b, MKM^{+}16, MLP^{+}14, NBF^{+}19, ORMG11, PVMG19, Pun11, RCHSN^{+}15, RMB^{+}21, SCJ^{+}11, SAR^{+}15, dSSdAdAA17, SAH15, SJJ^{+}11, SAK^{+}17a, SH10, VRMF13, VFMS16, WFLP11, WDL17, WP15, WP17, ZQZ^{+}17, BLHS11, GMS^{+}17, GPP^{+}16, HFPL16, HF19, HKB14, KLB17, LW16, LWH19, TDH19]. Coastal [FSH^{+}13, PKAL13, AdM19, AFA11, AA13, Ata11, BKT10, BK16, CLP^{+}19, CPP^{+}10, CDF^{+}13, CM15, DSB^{+}10, EMOLS^{+}15, FL11, FHK17, GCK14, GBA^{+}18, GPT^{+}19, HFD^{+}13, HKB^{+}18, HKM^{+}17, IGSJBS^{+}15, KNJK^{+}10, KDM15, KR12, KHS12, LSŠ^{+}16b, LHLR17, LPC^{+}14, LLH^{+}14, LMN^{+}12, LL14b, LW19b, MP14, MAM^{+}13, MCC^{+}17, MMGEMGS10, MAY12, MHRD16, MMK10, ML11, MCD^{+}13, NMH^{+}11, NWS^{+}11, PTF^{+}18, RM10a, SHPSK15, SABG12, SKB^{+}18, SKB^{+}19, TC16a, TC16b, TD19, VAB^{+}14, WSG16, WBB^{+}16, YMA12, Zie12]. coasts [CHT18, CCC^{+}14a, HA12]. Coccidiosis [PGG10]. Cod [CSKB13, HNW^{+}18, AJOE17, APCF17, ABL21, ADH^{+}13, BHL12, BVN^{+}10, BKT10, BK16, BJP^{+}13, BGD11, BLHG17, BLHO18, BKLB14, CHW^{+}10, CS18, DR13, DJW^{+}18, DRE^{+}17, DBH^{+}10, EPCB17, FA19, GBM^{+}18, GSL14, GSH^{+}15, GHJ13, Han18a, HWS^{+}15, HLS^{+}16, HSLB19, HPNA16, HBD^{+}16, HMA^{+}16, JTP^{+}14, JLF^{+}17, KHM^{+}17, KJS12, KW10, KR12, KLH^{+}15, KTT13, KHM^{+}10, KHV^{+}12, Kri13, KNB^{+}19, Lam13, LDR^{+}12, LAB13, LHS^{+}18c, LHF16, MFHK10, MTH10, MABA12, MM11, MSR14, NPK^{+}10, OS12, OTA^{+}13, OPE^{+}17, OSMP18, OKPL11, PRMKASR13, PHBU19, PRR^{+}19, RBV^{+}13, RW19, RSAB11, SH13a, SJS^{+}17, SHG^{+}15, SBH^{+}16, SDV^{+}16, SGGH10, SKDO^{+}17, TWM^{+}10, TCBR17, TSHB19, WLGS19, WOPB11, WHMS11, WHF^{+}14, WTK10, WAT^{+}13]. cod-end [BGD11, KW10]. Codend [NFK^{+}17, BHL^{+}16, BHS16, BMB10, CFPF^{+}15, FHMK11, Hen11, HWM^{+}13, HSL^{+}13, HKK18, HSLB19, Kim15, MMJ^{+}12, MKHK19, OKB^{+}16, PHRH16a, SL10, SL11a, SNK19, TSHB19, WHMS11]. codends [CEB^{+}19, DÖA^{+}10, FHM10, HWS^{+}15, HLS^{+}16, KHMF11, ÖEG^{+}15, QEH^{+}11a, THA^{+}14, THG^{+}16, WHF^{+}14]. coefficient [vPWW16]. Coelorinchus [FO11]. Coenen [BR15]. coexist [GIK^{+}16]. coexisting [CAMFMF18, SHPSK15]. Cohen [QDG17]. cohesion [MAS17]. Coho [BRGS19, BDB^{+}11]. cohort [BA12a, CJCN21, CHCS14, FE10, SHR18b, TCAP11, TMV16, WKTD13, YA18, YCYC16]. cohort-specific [WKTD13]. COI [CLF^{+}17, FCOM18]. Coilia [RLJ^{+}12, XYL^{+}19]. coindetii [PPG^{+}11, RLVS14]. Cold [DSC19, KSH19, WAT^{+}13]. Cold-water [DSC19]. Collaborative [FLW^{+}19, HHM^{+}13, MNW12, SAF^{+}15]. collapse [CH15, MSDT16]. collected [FBGB16, Lyn14, MGP^{+}18, NS16]. collection [CDL21, HCM^{+}11, LBB^{+}13a, SPM^{+}10, VRMF13]. Cololabis [HMGH12, MSV^{+}19]. Colombia [MFIG16]. Colombian [PWSP12]. Colonization [AGS^{+}16]. Colorado [PKWP16]. colour [JČK^{+}12, JČK^{+}21, MWC15]. Columbia [HHD^{+}11, ZD13, BRGS19, DET^{+}16, EB15, FRDC10, GHL^{+}11, HSY16, HKF^{+}12, JC18, TGE^{+}18, ZCLB11]. column [RH13]. combination [HB13, YMA12]. Combinations [CFF^{+}16, MAY12]. combine [RUMBA15]. Combined [AMV14, FdMCB^{+}19, MMAH15]. Combining [DHC19, ELRH17, PTG^{+}17, QDG17, VPC^{+}14, RRW^{+}13, RMP^{+}18, SRMP^{+}18, WSF^{+}18]. come [CGM^{+}19]. Comment [CFF^{+}15, HL16, Ken14, Rob15, BKS^{+}15, Han16b, HM16a, HM16b, PHK14]. Comments [AAS19, Han16a, NB19, SA16, Sve14, JBMK14]. Commercial [BBV17, SH10, AHH^{+}12, APHL13, BBT^{+}15, BHM^{+}18a, BSN17, CUL^{+}16, CW17, CMB^{+}12, COS^{+}19, CRH^{+}13, Cro15, DTG13, EMP^{+}10, ETMK13, FPSDB17, GKMGO^{+}17, GHP^{+}15, GSP17b, dMGPR^{+}18, GWWW13, HIH19, HPHF15, HKK18, IIS10, JIM^{+}15, LNC^{+}15a, LNC^{+}15b, LAB13, LSC^{+}18, LPHH11, LMBF13, MCI^{+}13, May10, MKH16, NVS^{+}18, NS16, NFK^{+}17, NSM^{+}19, ÖEG^{+}15, PQ10, Pul17, PS19, RKP^{+}12, RHOCB^{+}18, RAG19, RR12, SGSFL19, SLT^{+}19, SJA17, SB17, SRG17, TPM15, TCS13, WM11, WTS^{+}17, WMPR17, WUGG16, YGA^{+}12]. commercially [BMdS^{+}15, BPMGP10, BB13, DJF^{+}15, FBY^{+}12, HND^{+}17, HRAK15, KRMW15, KA15, MFHK15, SD10, SPM^{+}19, SWD^{+}13, TC12, VRS13, VRW11, ZzCbC^{+}15, dVLB^{+}15]. commerson [EPHF18, NML12, RU15]. Common [BTB^{+}19, Dam15, BSF^{+}19, BMdS^{+}15, BPMGP10, CJCN21, CC11, CCC14c, EHBK18, FE10, GDW^{+}15, GBD^{+}15, GHUG^{+}16, HAB^{+}10, HMGH12, HH14, KG10, Kri13, LBF^{+}16, LXCC19, MLM15, OCN^{+}11, PFVGGM16, PRL^{+}18, RHC^{+}12, SHA^{+}15, SMS^{+}19, SCT^{+}17, TS10, UK14, VSTS10, WHB^{+}16, WJ17]. Communication [CPMNS10]. communities [AWSLS15, AVP^{+}15b, AA13, BLR^{+}16, BRW^{+}13, CKVV16, EMOLS^{+}15, FTM^{+}15, HCK15, JWZ^{+}19, KDM15, OH12, VS16, ZCX^{+}18]. Community [MHB^{+}19, ASB^{+}13, BTBP18, CRW^{+}19, HAB^{+}17b, MDS^{+}18, MNAA^{+}14, NMH^{+}11, NPRGTSR18, RRSB13, SFR^{+}13, ST16, VJH17, YAH^{+}13]. community-based [HAB^{+}17b]. Community-level [MHB^{+}19]. Comoros [HDM13]. Comparative [ARVACPMN16, GMCL^{+}18, MLBTM16, VRMF10, DBSA18, GBK^{+}13, TÜM^{+}16, VFdS10, WHF^{+}14, XLCC19]. compared [BHL12, HWS^{+}15, JLF^{+}17, KOHF^{+}16, PJR17, SO18]. compares [FSA^{+}16]. Comparing [BLPF^{+}15, CEB^{+}19, CRAVC14, GCRV10, LMHR19, MBSM15, PBK^{+}19, SGC11, THF12, BSM13a, IHFAH17, KLWG17, PM11]. Comparison [BSS^{+}13, BAY^{+}12, BB17, BB19, BWK^{+}13, CPG^{+}15, CRB^{+}18, GFK^{+}12, GHS^{+}19, HNM^{+}12, HCM^{+}11, Hor11, JBB^{+}18, Kri13, LWZ^{+}19, MBJ^{+}16, MAA^{+}16, PLMT^{+}11, PLPN12, SMK^{+}13, WKA13, WKT^{+}11, AENC14, BTT11, BAB^{+}12, BJD12, BGA11, CJG^{+}15, CBD^{+}14, DAM10, EHB^{+}21, EMOLS^{+}15, FPC^{+}13, FBM^{+}19, FH15, FSFP15, FdMCB^{+}19, GMMW17, LHS^{+}16, MCP14, PKH^{+}11, PMK^{+}10, RKP^{+}12, RHOCB^{+}18, RCBPK11, SBAK17, SBLM14, SHT^{+}13, SAWS18, SBL^{+}19, TT14, THA^{+}14, YJSR12, YJC13, ZBB17, ZHDS16, vMWvOKvB14, KT18b]. Comparisons [MMF^{+}17, PD14, TMM19, BM19a, CPH12, LNC^{+}15a, LNC^{+}15b, MMMPP19, PRBGS^{+}10, SBT^{+}18, WPS19]. Compartive [AdM19]. compatible [MCDP18]. compensated [KNJK^{+}10]. Compensatory [WHB^{+}16]. competency [RHB^{+}17]. competing [RCBS19, Sal18]. competition [JNN14, MS15b, MM18, PB11, STAN^{+}17, SMPB17]. competitive [BCPP17]. Complementary [FWG11]. Complex [DW19, CMP^{+}14, DBSA18, GKL12, HA12, HZT^{+}19, KBOM14, LPM^{+}13, ODMF17, RDP^{+}17, SB16]. complexities [EHBK18]. complexity [AL13a, JBKA15]. compliance [BAFP18, GGADSRVL15, PW19, SB12]. component [FHM10, HBMC12, LPTK19]. components [GPT^{+}16, MSR14]. Composition [VKNK12, VS16, AVA18, CV14, CDR^{+}11, DPSRM10, FWG11, FMF^{+}17, GLMK12, GK18, Gri12, HNS12, IGSJBS^{+}15, KKL^{+}16, KSAA13, Leg14, LRM14, LCM^{+}18, LL14b, Mau11, MVMdS^{+}17, MLT17, MMO^{+}13, OKAKO16, PTSC^{+}16, QEH^{+}11b, RCLS18, RBG^{+}19a, SRSNC17, SCC^{+}15, SH10, SRG17, WDL17, WM17, WCS^{+}12, ZMdSC16]. compositional [SM17]. compositions [DBG19]. Comprehensive [FBB^{+}17, RPH^{+}18, XYL^{+}19, BKSF17]. compressed [RHB16]. compromise [WBD^{+}13]. compulsory [GSH^{+}15]. computational [TFH^{+}15]. computationally [TM13]. Computer [NWO^{+}15, PM13]. concentrations [RGGdL13]. concept [FHMK11, FA19, WKA14]. concepts [Gan13, Kur12, MFHK10]. concern [CAM^{+}19]. Concerns [MZC^{+}18]. conch [BdGN^{+}18, SMBP^{+}12]. Conciliating [LCB17]. conclusions [TBS13]. concurrent [VJH17]. condition [ASN17, BA12a, BRS19, BTNA13, BBLS17, CCA16a, DPO^{+}11, DOM14, DPSRM10, FAAM15, FAAM16, GGB^{+}18, HPNA16, Kni12, KT18b, LRM14, MEMdU^{+}17, NNS^{+}10, NGvdL^{+}13, OSO^{+}18, RHB^{+}17, Rod19, RLMM18, SAK^{+}17a, SAB^{+}11, SSS^{+}12, SBC^{+}18, WBM17, WMF13]. conditional [BHCJ12, LPTK19, TWM11, XTP^{+}16]. conditioned [MLF^{+}10]. Conditions [Kos15, CJCN21, CP11, CP12, CR18, FBHH14, GHM16, HFBS11, JS15, Kri13, MSS^{+}16, MWSN^{+}13, RU15, SKF^{+}10, UKK^{+}12, WSMS14, YLA15, YCYC16]. confidence [PSE15, WMT11]. configuration [EDG16, SL10]. configurations [BSM13a, GPS15, PHT15]. Configuring [BM18a]. confinement [DHP^{+}11]. confirms [LSWR17]. Conflict [GS18, IOT14, MCP^{+}17b]. conflicting [ANEA19, WHC^{+}19]. conflicts [MP17]. confluentus [GHL^{+}11]. Confounding [BSM16]. confounds [JSR10]. congener [CPH12]. conger [CBS11, LZX^{+}17, MLBGV16, CBS11, LZX^{+}17, MLBGV16]. conical [BM19b]. connected [BST^{+}19]. connection [DPO^{+}11, PBB^{+}18]. Connectivity [DH14, APB^{+}16, ASSL11, AFAJ12, APD^{+}18, BDM^{+}18, CMPR^{+}16, CN19, GPP^{+}16, HBS^{+}16, KKH^{+}17, MSN^{+}12, MHC^{+}18, PLSM^{+}16, RMMSR19, dJSMROZP^{+}13, ZZA^{+}14]. consensus [LG13]. Consequences [GKKK15, LLM^{+}19, OTA^{+}13, TGE^{+}18, AAS19, CAR19, DHP^{+}11, DJF^{+}15, GGB^{+}18, Lam13, Lor16, NB19, OJT^{+}14, RHC^{+}12]. Conservation [HSY16, OLS15, ASMD^{+}18, AAN^{+}12b, ANEA19, BCC15, BMAMG18, CMSG18, EL18, FPM^{+}18, HHM^{+}13, HGA^{+}12, HCH^{+}19, KM15, MCDP18, RMGK18, SMGH17, SNOD^{+}13, TdMWC16, WLZ^{+}15, WUGG16, XYL^{+}19]. conserve [MBB13]. Conserving [BAOP18, EHBK18, WGM^{+}11]. consideration [CHL11]. Considerations [DSG^{+}17, GAB^{+}15, Kni12, ALBŠB^{+}15, CAM^{+}19]. considering [KWCT^{+}18, Kur12, MAP14, MAA^{+}16]. consistent [FSFP15]. conspecific [TQ17]. Constance [RvAGB18]. constant [BB19]. constraint [MM18]. constraint-based [MM18]. constraints [HBC^{+}17, KLB17, RCF^{+}17]. constructed [MBJ^{+}16]. Constructing [Cam15, CEM^{+}11]. Construction [ZD13]. Consumer [BPS^{+}18, MFKS17, WDM15]. consumption [BR15, GPS15, KSH19, SVG^{+}17]. consumptive [PLGL18]. contact [SHG^{+}16]. containment [RBW14]. contemporary [CHA15]. content [BBLS17, PPLK18, SKT18, SRG17, TCS^{+}15, VIFLP17, VSM^{+}18, VSLC^{+}19]. Contents [Ano10b, Ano10c, Ano10d, Ano11b, Ano11c, Ano11d, Ano21a]. context [AFA11, EHBK18, PBM^{+}19, Pun19a, SRSNC17, YCQ11]. continental [BMR^{+}18, HNM^{+}12, MS15b, TEFAR12]. contingent [DHC19]. continual [Mun18]. continuation [BK12, CNWM12]. Continuous [HPRB16, Har19, OS19]. continuum [SCF17b]. contraction [ZH10]. Contrasting [AASF16, BTB^{+}19, SMdUK16, WAT^{+}13]. contribute [CLG^{+}14, FSPWG08, GH14, NPRGTSR18]. contributing [JHCT11]. Contribution [MMC13, BMRC14, CLS^{+}11, TTS11]. contributions [Ben18, SHR^{+}18a]. Control [PKRC15, DB12, FHK18, dSFdSSS17, GSL14, GPT^{+}19, JAB15, KG10, Sau14, WP17, ZCZ^{+}13]. controlled [FWR10, KKE^{+}12, MWR18]. Controlling [PAR^{+}12]. controls [PTSC^{+}16]. controversial [WBI^{+}16]. conventional [BFW13, GHS^{+}19, MAY12, MBSM15, SFH^{+}15]. Convergence [PGJ10]. conversion [Che10, HBP^{+}18]. cooperation [GS18, Han13, MM18]. cooperative [EFWC^{+}19, GBA^{+}18, TW14]. coordinated [SKB^{+}18, SKB^{+}19]. Copepoda [LRAP19]. coproduction [CDB^{+}14]. coral [ASB^{+}13, CT16, CBH18, DSC19, PLMT^{+}11, PVKH15, PVKH18, SOMO17, ZBB17]. coral-reef [CBH18]. corals [WGM^{+}11]. cordyla [SCJ^{+}11]. coregonid [EHB^{+}21]. coregonid-dominated [EHB^{+}21]. coregonines [WCS^{+}12]. Coregonus [AD16, HSS18, JWM^{+}12, LCF15, RSV^{+}17, SMS17, VJH17, WBM17]. cormorant [HL16, SAR^{+}15, SA16]. cormorants [KAiT^{+}13, LHLR17]. Correct [DM16]. correction [SOY14, TK16]. correlated [GBM^{+}18]. correlates [KKA10, PSJ^{+}19]. correlation [PGS^{+}11, dVA18]. corridors [HCWS19]. Corrigendum [ASHH15, HAB^{+}18a, HLNH15a, IIT18, LKD18, LNC^{+}15a, Mun12a, PGP^{+}16, PVKH18, SKB^{+}19, TC16a, TV14, Tho19a, Tho21, VRMF12, WGC^{+}21, WMB16]. corroborated [EHB18]. Cortisol [BV14]. Coryphaena [APDJSU11, CM12, CDF^{+}13, FBHH14, FTN^{+}14, MAH14, MAH16, SMB^{+}17, TVdLOG^{+}21]. Coryphaenidae [SMB^{+}17]. Cost [KNSPK13, KKE^{+}12, RAG19, BAY^{+}12, BCMS^{+}18, Cra19, HPRB16, KNSSL11, KGR^{+}15, KT18b, MVMCC16, OLS15, PFVGGM16, PHBU19, PAM^{+}18, RIB^{+}10]. cost-benefit [BAY^{+}12, KNSSL11]. Cost-effective [KKE^{+}12, HPRB16, PFVGGM16]. Costa [CEW14, VRW11, OPL17]. costs [DTG13, Pun18]. Could [BHL^{+}16, LHS^{+}18a, NPRGTSR18, Ibá14]. Council [Ano10-27, BPPR16, BGW^{+}16, BFHT16, PRCGD16, PLA16]. count [CPG^{+}15, UP16, WMT11]. counterparts [TZW16]. counters [vPB18]. Counting [BKS^{+}15, BPRB^{+}14, GRVN10]. countries [Kos15, KMMV10, RFC^{+}17]. Country [GAM10, PBG^{+}13]. counts [DPAB10, KA15]. coupled [DQMV19]. Coupling [ESF^{+}19, BNU^{+}10]. covariate [MD10]. Covariates [DLCK15, HLNH15a, HLNH15b, KY17]. Cover [Ano17k, Ano17l, Ano17m, Ano17h, Ano17i, Ano17j, Ano18k, Ano18l, BSM15a]. coverage [Har19]. Cox [QDG17, Sau14]. cpue [PLPN12, FBKE16, FCL16, FSG^{+}19, FO11, NWS^{+}11, OTK^{+}16, RBV^{+}13, WB15]. Crab [DWH10, AK11, CW17, CMMVRF^{+}19, CCGD19, DMK^{+}19, FT16, FWL18, FLW^{+}19, GMUD11, HNS12, IAJ^{+}11, JCA^{+}16, JHCT11, KGS18, KC11, LBB^{+}13b, LSH13, LCH^{+}18, MCC^{+}17, MD16, MFRM19, MCMK18, MRT17, MRT18, Mur19, NWL^{+}14, OHSG19, PHE16, RM10a, SSFGL19, SGSFL19, SLT^{+}19, SZPV16, SZPP17, SBLL17, SP12, SP15b, TMC19, TEGM^{+}11, VCS^{+}19, WHNS13, WW11b, WLBB15, YSS^{+}11, YRH15, ZD13, ZC17]. crabs [DSW^{+}17, GMMW17, LBF^{+}16, LWY^{+}11, MMAH15, PTM^{+}18, RMGK18]. Crangon [VLSP19]. crassicaudus [CRAV13]. Crassostrea [BPK^{+}17, CDMC18, MAAW^{+}17, MMAA^{+}18]. crayfish [BdGD^{+}16, KHQH15, ZCW11]. create [CGC13]. creates [VFMS16]. creeks [MCD^{+}13]. creel [BHM^{+}18b, BHM^{+}18a, CM18, HPRB16]. CREELSELECT [BHM^{+}18b]. Creus [MLV13]. crews [BAB^{+}12]. criteria [KY19a, SMBP^{+}12]. Critical [WBI^{+}16, Cor17, LRM17]. Critically [CHA15]. criticism [SA16]. croaker [CMM^{+}14, HD12, LSJD11, LZGL17, dSSdAdAA17, TdMWC16, WLZ^{+}15, ZYW^{+}14, dVLB^{+}15]. cromis [OMB18]. Cross [GMNPM^{+}13, WPS19, BM19a, LQL^{+}15, MH11, SBT^{+}18, dVA18]. cross-ecosystem [BM19a]. Cross-site [WPS19, SBT^{+}18]. cross-species [LQL^{+}15]. crowding [FWR10]. cruising [KS12]. crumenophthalmus [PGAS15]. Crustacea [dADSRG10, LMBF13, SAK^{+}17a]. crustacean [FBB^{+}17, KA15, MBF^{+}19b, QEH^{+}11a, QEH^{+}11b, SGW16]. cryptic [FLW^{+}19, WWN^{+}14]. CSFs [MNAA^{+}14]. Ctenochaetus [OTST11]. cubic [HPB10]. cucumber [CUL^{+}16, HHD^{+}11, RLM10]. cucumbers [GHM16, GWVH^{+}16, HFCPSEM15, MCP^{+}17a, PNAL16]. cues [HFL14, KBL^{+}15]. cultivation [HKAA18]. cultural [AFA11]. culture [JWZ^{+}19]. cultured [DS18]. Cumulative [PKRC15, BKSF17]. Current [AT10, JKS^{+}15, JKK^{+}16, KY19b, KEH16, NBS15, TKF11, BOCÁ^{+}18, DCP^{+}16, GDB^{+}16, HBP^{+}18, JPO^{+}19, JTHT14, KEP^{+}19, LMSM14, LF13, PSS15, Pun19a, Pun19b, Qui17, WTF16, WKT^{+}11, ZXHK15, FT16, JD14, LL14b, STY^{+}16]. currents [TBS13]. curve [MMF^{+}17, TCAP11, TC15, WMP^{+}14]. curves [CRAVC14, MH11, WSMS14, WK10, Whi10, WHNS13]. customary [TRH16]. CUSUM [PKRC15, PKRC13]. cutlassfishes [TC12, WDL17]. cutter [BBLD16]. Cutthroat [LSJ17]. cuttlefish [AAOE13, ARF^{+}17, GHUG^{+}16, LXCC19, PVMG19, SMRS15]. Cuvier [PRBGS^{+}10, HSM^{+}12]. cyanea [RB12]. Cycle [CDB^{+}14, VRMF11, VRMF12, AF13, CCC14c, DBT15, FBGB16, FAAM15, FAAM16, LMN^{+}12, ORMG11, RB12, SP15a, VRMF10]. cycles [AD16, KC11, LS19, WFHG^{+}17]. Cyclic [ALR^{+}17]. Cyclopterus [HBG14, KJ17]. cygnus [BEM13]. Cynoglossidae [Ata11]. Cynoglossus [Ata11]. Cynoscion [EHB18]. cypha [PKWP15, PKWP16]. cyprinid [BCMS^{+}18, MLT17, PKWP15, ŠBB^{+}19]. Cyprinidae [RADR11]. cyprinids [GRB14, MLT18]. Cyprinodon [PTTC18]. Cyprinus [KSAA13, OCN^{+}11, RHC^{+}12, RSEERR^{+}16, WHB^{+}16]. Czech [BJKH12].

d [SPMS17, IdRH10, KRB17, MCMK18, MML^{+}16]. D-loop [IdRH10]. D-Xray-tomography [KRB17]. D. [JBMK14]. dactylopterus [RMMSR19, SGN^{+}10, SNP^{+}12]. dactyloterus [SVN^{+}13]. Daily [FGK19, PKO10, SSS^{+}12, BGF^{+}15, Ber13, HHH^{+}12, JCW12, MIÁ^{+}10, QQG14]. Damage [LBB^{+}13a, CC11, DTSG^{+}16, EHGG16, TSHB19, UGGR^{+}17]. dams [LDM^{+}15]. Danish [BNAE10, BNAE13, FBKE16, LE14, NFK^{+}17, NSM^{+}19, PHBU19, SKFM17]. Danube [SGT^{+}14]. dart [FH15]. Dasyatidae [SAH15]. Dasyatis [SAH15]. Data [Ano17a, IOT14, MCP^{+}17b, PDS^{+}17, dSMSF15, ARF^{+}17, AVA18, AM12, AVB^{+}16, BB10, Ben18, BNK14, BFL^{+}14, BKS18, CPG^{+}15, Cam15, CHT18, CAMW11, CWM12, CPW^{+}14, CMP^{+}12, Cer14, CC19, CEW14, Coo13, Cop13, CSMO14, CV14, CMLP19, CTS^{+}19, DHC19, DM11, DDH^{+}15a, DDH^{+}15b, DPL^{+}16, DMK^{+}17, DBSA18, EHB18, EWB^{+}18, ELRH17, EAM^{+}14, EPO10, EMSC15, FBGB16, Fey18, FWL18, FGS19, FdSM^{+}16, Fra17, FJWD17, FPD^{+}16, GSP17b, GLMK12, GRO^{+}18, GDC^{+}19, HF15a, HWS10, HSS^{+}16, HGD19, HPHF15, HF14, HPB10, HBB^{+}12, HF15b, HLP15, HCR14, JIM^{+}15, JAB15, JTP^{+}14, JSG^{+}15, KMW^{+}11, KS17, KDPT18, KM10, KTNA15, KGB11, KNB^{+}19, LPTK19, LCMTM10, LKR11, MD10, Mau11, MP17, May10, MGM12, MM14, MAS15, MKH16, MAH14, MMG10, MVMdS^{+}17, MJC^{+}14, NBS15, NS16, NMLZ17, OTST11]. data [OLM^{+}19, OPE^{+}17, PKRC13, Peñ18, PBC^{+}16, PHBU19, PCP^{+}18, Pun17, RLC^{+}17, RBT14, RPH^{+}18, RU15, RBR^{+}15, RRW^{+}13, RPC11, RPP^{+}11, RCF^{+}16, RMP^{+}18, SRSNC17, Sam11, SFB11, SBLB17, SLH^{+}14, SDWG18, SFPB17, SZPV16, SZPP17, SEG^{+}15, SZ10, SPM^{+}10, SM17, SKB^{+}18, SKB^{+}19, SAF^{+}15, SH13b, SLCL15, TJJ14, TSS11, TC17, TRW19, TCS13, TAS12, WMNC15, WM17, WRR^{+}16, WP11, WP15, WMB12, WMB16, WWG^{+}16, WFJ19, WKA14, WvPLW18, XGT^{+}17, YCQ11, YvZG^{+}17, ZTC^{+}18, ZSF11, ZMdSC16]. data-deficient [CEW14]. data-limited [ARF^{+}17, CPW^{+}14, CC19, Cop13, FWL18, FPD^{+}16, HSS^{+}16, HF15b, KMW^{+}11, LKR11, MMG10, NBS15, RU15, SAF^{+}15, TC17, WFJ19, YCQ11]. data-moderate [TRW19, WP15]. data-poor [CHT18, DM11, DDH^{+}15a, DDH^{+}15b, EAM^{+}14, FJWD17, HLP15, KTNA15, RLC^{+}17, SRSNC17, SDWG18, TRW19, WP11, ZSF11]. data-rich [FPD^{+}16, WFJ19]. database [PBML^{+}10]. databases [GGS^{+}19]. dataset [MFMM16]. datasets [FSG^{+}19, LR19]. dating [CAC^{+}14, IAS^{+}17]. day [GPP^{+}16, SKF^{+}10, USH^{+}13]. Daylight [ÁCDM^{+}14]. dead [LWS19b, TM13]. Dealing [MD10, MP17, RGSAMV19, SH13b, WM17]. debate [HBC^{+}12]. debris [BM19b]. Decadal [BB18, SOMO17, Lyn14, SHR18b, TKF11]. Decadal-scale [BB18]. decade [JCW12]. decades [BSL10, LEO^{+}11]. Decapoda [dADSRG10, BdGM^{+}11, BdGD^{+}16, CCA16a, CCA16b, FdMCB^{+}19, HW11, KTL11, QWMD13, RM10a, SAK^{+}17a, TEFAR12, VRW11]. Decapterus [SJJ^{+}11]. Deccan [RADR11]. December [Ano10w, Ano11p, Ano12m, Ano15q, Ano17n, Ano18u, Ano19a]. Decision [CJP13, BNAE13, GMS12, GSDD11, NBF^{+}19, Pop19, PML^{+}12, SVPC16, PKRC15]. decisions [MMdlP16, Tho19a, Tho19b]. decisive [KHMF11].

Decline [JASP11, JHCT11, BB18, DAM10, HER13, MP18, NEJ^{+}18, VJH17, ZZA^{+}14]. declines [FLW^{+}19, LGH^{+}21, vPWW16]. Declining [HSM^{+}12, MGC^{+}11]. declivis [SWH17, WBI^{+}16]. decompression [CPB^{+}12]. decoy [WTS^{+}17]. decreased [KVH15]. Decreasing [LJH11, KY19b]. Deep [LHS^{+}18b, LHS^{+}18a, PWSP12, SSH14, CAC^{+}14, EHB^{+}21, GAR19, GMUD11, JHL^{+}19, JKS^{+}15, Ken14, KKH^{+}17, LWCC15, MD16, MJC^{+}14, PHK13, PHK14, SWA19, TEGM^{+}11, VRS13, VFMS16, WM13, WGM^{+}11]. Deep-sea [PWSP12, SSH14, GAR19, MD16, TEGM^{+}11, VRS13, WGM^{+}11]. deep-set [SWA19]. Deep-water [LHS^{+}18b, LHS^{+}18a, CAC^{+}14, JHL^{+}19, JKS^{+}15, Ken14, KKH^{+}17, MJC^{+}14, PHK13, PHK14, VFMS16]. DeepVision [RH13]. deepwater [ASR^{+}18, CCT^{+}14, HKRH12, ODMF17, RMB^{+}21, UEOA19, VRW11, WUGG16, HCSM17]. deficient [CEW14]. define [OPH15]. Defining [CEJI11, DMOL15, JNE19, JSG^{+}15, LCMdSM13, MDR^{+}19, NCP15, VFdS10, MS14]. Definition [DL11, PLMT^{+}11]. deformation [ZXHK15]. deformed [JPSR^{+}14]. Degree [USH^{+}13, BWK^{+}13, LJJ^{+}13]. Degree-day [USH^{+}13]. Delaroche [SGN^{+}10, SNP^{+}12, SVN^{+}13]. Delaware [MAAW^{+}17, MMAA^{+}18]. delay [PDP^{+}11]. delay-difference [PDP^{+}11]. Delayed [CS18, RHB^{+}17, SW19]. delineating [MM11]. delineation [HHB18, KMJ^{+}12]. delivery [WAT^{+}13]. Delta [ZD13, GDA14, Mac13, DS18, LJH11]. Delta-AdaBoost [LJH11]. Demand [KL16, vPKF^{+}19, DP18]. Demand-side [vPKF^{+}19]. Demersal [GSP17a, NPGT17, BSF^{+}19, BH13, BHC10, BK15, CLS^{+}19, CEW14, CVS11, DV11, FA19, GSP17b, GLMK12, GPS15, GSH^{+}15, HNM^{+}12, HND^{+}17, HLS^{+}16, HSH15, HSH16, JRG^{+}18, LCZ^{+}19, MMJ^{+}12, MPB^{+}17, MBOAA12, MQM^{+}15, NPRGTSR18, OS19, OOM15, PMF^{+}16, PGP^{+}15, PGP^{+}16, RBT14, RRW^{+}13, SL11a, SFR^{+}13, ST16, TKF11, UEOA19, VKM^{+}17, WHS^{+}18, dVLB^{+}15]. Demographic [HDW^{+}15, CAFHG10, MSWW17, MRT17, PKWP15, PGAS15, PLSM^{+}16, SCT^{+}17, VB15]. demographics [EHBK18, FHK17, HKB^{+}18]. Demography [VdSS^{+}14, GTDJdL^{+}13, HGD19, NML12, SSG17]. Demonstrating [BM19a]. demonstration [GLG^{+}11]. densities [EHB^{+}21]. Density [Han18a, JB15, LC19, MMAA^{+}18, AGB15, CMB^{+}12, CM15, DJD12, EP12, FMMW13, Goo16, GYCL19, GBF11b, HFCPSEM15, KY17, KMKK10, LRM14, MAA^{+}18, NS16, OTK^{+}16, RWRT15, SSFGL19, SSS^{+}12, TQ17, UJLC^{+}19, WBM17]. Density-dependence [LC19]. Density-dependent [Han18a, MMAA^{+}18, EP12, FMMW13, UJLC^{+}19]. dentatus [HF14, MW11, Mau12, PT19, PBD21, YGA^{+}12]. depauperate [HMvHN13]. dependence [LC19, SKY^{+}17, ZD19]. dependent [APB^{+}16, BAY^{+}12, CTSD10, CBMLB13, EP12, FMMW13, Gra16, Han18a, HFBS15, HSH15, JB15, JCA^{+}16, KSH19, KGB11, Kur12, LJJ^{+}13, MMAA^{+}18, OGSNH^{+}17, SKK^{+}13, UJLC^{+}19]. depensatory [PC19]. depleted [MBOAA12]. Depletion [DM11, CHT18, CNJ11, Car18, FPM17, FL13, LT18, MDS^{+}18, RUMBA15, RU15, SBJ^{+}19, WRR^{+}13].

Depletion-Based [DM11]. deployment [PHM^{+}16a, SLXZ15]. deployments [RBG^{+}19b]. DEPM [NL11]. DEPM-based [NL11]. depredation [CGR^{+}15, HPP18, JRG^{+}18]. deprivation [AJOE17]. Depth [WGS17, CY12, CN19, HRB14, HCR14, JTP19, NPGT17, RHB16, TTT16, WM13]. depth-related [NPGT17]. depths [JBB^{+}18, PSJ^{+}19]. derelict [EP17]. Derivation [HE15]. derive [SLCL15]. derived [BWN^{+}12, CV14, DJD12, MFC^{+}13, SHJ19, ZTC^{+}18]. deriving [Cop13, PD14]. descending [BBA^{+}19]. describe [NPRGTSR18]. described [BR15, HBM11]. Describing [VvPM^{+}18, HHTT11]. Description [SJR^{+}11, CDM^{+}19, HHSS17, PPdR^{+}10]. Descriptive [SAK^{+}17a]. descriptors [ERG^{+}11]. Desiccation [FHS^{+}16]. Design [TFH^{+}15, BSB^{+}15, BGL^{+}13, BLHG17, Chi11, GAB^{+}15, Har19, HB11, HSRT17, KBCK16, LCRW19, MMJ^{+}12, MBSM15, PHM^{+}16a, PHBU19, RMB^{+}21, SNK19, SGW16, SWH17, SEG^{+}15, SAB^{+}11, SHM12, TXZ^{+}17, TSHB19]. designated [Cur18]. designed [KSL^{+}17]. designs [Chi10, EFWC^{+}19, GBH^{+}19, HSLB19, KHKM15, Mun18, RBW14, SBH^{+}16, WW11b, YJSR12]. Desmarest [CMM^{+}14]. despite [JASP11]. destined [ASN17]. Detailed [BNU^{+}10]. detect [CEM^{+}11, GMCL^{+}18, RIB^{+}10]. detected [CHA15, CPCP^{+}12]. Detecting [CMB^{+}12, CSP^{+}18b, FWL18, SBLB17, WCM^{+}17, KS12, MPF^{+}17]. Detection [HPAA19, dPHS^{+}12, BCP^{+}12, DMFR16, KBT^{+}19, NMLZ17, RCF^{+}17, dKMS15]. detections [PKWP16, YAH^{+}13]. Determinants [OH12, BCP^{+}12, HSD^{+}17, KKP11, LOdH16]. Determination [WAB^{+}16, HSB^{+}13, KA15, KHG^{+}17, MWR^{+}13a, PGL^{+}13, WP11, ZYSZ18, dCJG^{+}15]. determine [BMG16, CCC17, CC19, MH11, RBM^{+}16, TVCT12, YRH15]. determined [BMCS19, Kim15, MAH14, MNGPM^{+}12, PMK^{+}10, SFB11]. determines [RFC^{+}17, WHC^{+}19]. Determining [Gan13, MBF^{+}17, SDC^{+}15, BHM^{+}18b, CN19, CDR^{+}11, DM11, GSDD11, MHK^{+}17, WWG^{+}16, dKMS15]. deterrent [GWWW13]. developed [AdM19, AFOB^{+}19, BMH17, LHS^{+}18c]. Developing [BSSH13, BMAMG18, FPD^{+}16, KZ11, RSV^{+}17, RMB^{+}21, SZ11, SZ10, YGA^{+}12, DDH^{+}15b, HHD^{+}11, Kos15, MTC^{+}17, NL11, SMD^{+}16]. Development [BWR^{+}18, CPS^{+}12, DMC^{+}12, FMD19, FHMK11, GJRR14, GSL14, HHM^{+}13, JSYS15, MFHK10, MHFK12, MHFH17, PHE16, SCF^{+}17a, BJD12, CBS17, FDMH16, GN12, GA19, KMSRW10, KGR^{+}15, LXCC19, MQCS^{+}14, OÁPP12, TWSA16, TEGM^{+}11, VRW11, WAT^{+}13]. developmental [KC11, LHL^{+}15]. Developments [AW12]. deviations [OALS11]. device [BHCS15, CPS^{+}12, GMJ^{+}13, HLJ15, LW16, LW19a, PTM^{+}18]. Devices [GSDD11, JPM11, SCV^{+}12, WDBH16, BBA^{+}19, BBK12, CRCMP^{+}14, CW17, CCT^{+}14, CSMJB11, EBB16, EPTG19, GSL14, GBR17, HB15a, HJ12, KLH^{+}15, MDC^{+}16, Rob15, SMRS15, TGDB16, HB15b]. devil [ACSAS^{+}17]. devolving [BBK12]. d’Halieumétrie [GBF^{+}11a]. DI-CUSUM-HCR [PKRC15]. diadromous [PTB^{+}15, SDWG18]. diagnostic [CPC^{+}17, HHB18, WMP^{+}14]. diameter [AHMS10]. Diamond [BMB10, CHW^{+}10, CEB^{+}19, DÖA^{+}10]. Diamond- [BMB10]. diamond-mesh [CEB^{+}19]. diamondback [CW17]. diaries [SJA17]. Dicentrachus [SAK^{+}17b]. dictate [SKF^{+}10]. Did [HCSM17]. DIDSON [PJM14, RTŘ^{+}12, TFB^{+}14]. Diel [BLP^{+}18, JPO^{+}19, BSM15a, GMJ^{+}13]. Diet [SVG^{+}17, BGL15, HL16, KNB^{+}19, PPLK18, SAR^{+}15, SMØ16, SWG18, SL12, TCS^{+}15, UJLC^{+}19, ZDZY19]. Dietary [KDdAA^{+}18, UK14]. difference [Hor11, PDP^{+}11]. Differences [FKM11, KNRJ^{+}15, OTH^{+}19, APAFMN10, BA12a, CSKB13, GFM^{+}12, KVK^{+}17, MGM12, RSEERRSC15, WPS19, YRP^{+}19]. Different [dSSdAdAA17, CDM^{+}19, DDP15, DBLB16, FTM^{+}15, FH17, GPT^{+}16, GBH^{+}19, GSTL15, GPS15, GWB^{+}19, GYCL19, HSLB19, JBB^{+}18, KHKM15, MCP^{+}17a, MREV10, PMK^{+}10, RLJ^{+}12, RBV^{+}13, RJKG12, RPC11, SMdUK16, SCV^{+}12, SMS17, WKT^{+}11, YAT^{+}19, ZXHK15]. Differential [BMdS^{+}15, VOM12, CFB^{+}17]. differentiate [CBB^{+}16]. differentiated [RFPG^{+}16]. differentiating [ER18]. differentiation [GTDJdL^{+}13, IdRH10, LSH13, LZGL17, MDL19, SMO^{+}18b, WB16, ZLT^{+}16]. differing [PMB^{+}11]. differs [SBT^{+}18]. difficult [GMSP19, KTW^{+}13]. difficult-to-tag [GMSP19]. digenean [WRH^{+}15]. digestibility [SAK^{+}17b]. digestion [GMCL^{+}18]. digging [SBA10]. dimensional [KMQ^{+}10, LDM^{+}15, PHB^{+}19b, WGS17]. dimensions [RH13]. diminishes [QMAM13]. dimorphism [Cer14]. dioxide [OPF15]. Diplodus [APAFMN10, ACRM12, CPG^{+}11, HVJP^{+}11a, HVJP^{+}11b, PKAL13]. Dipturus [QWCB11]. Direct [BGA11, Cla14, KA15, QMGP10, TRC19, dSMSF15, DCP^{+}16, JTHT14, MLTD17, WTF16]. directed [HCF^{+}17, LHF16, PTM^{+}18]. direction [WHMS11]. directions [Hil11, PSS15]. Directive [RBR^{+}15]. directories [TZW16]. Dirichlet [TJMT17]. Dirichlet-multinomial [TJMT17]. dirty [JBMK14]. disaggregated [MMdlP16]. Discard [Dam15, ERCG10, UGGR^{+}17, BHC10, BHCJ12, CHS^{+}17, CL17, CS18, DBH^{+}10, FCCSA15, KGS18, Pul17, RRS^{+}19, SH10, SBC^{+}18, YGA^{+}12, vPKF^{+}19]. discarded [CMS14, DDB^{+}14, Hoc12, KRM^{+}18, LBB^{+}13a, MCI^{+}13, MMK16, MMR^{+}18, RBM13, SNK19]. Discarding [BH13, BRH^{+}16, CEM^{+}11, PS19]. discards [CFBS16, CHW^{+}10, CGC13, DV11, FPC^{+}19, dMGPR^{+}18, GK18, HL10, JF13, LBB^{+}13b, MUO^{+}17, PMF^{+}16, PCP^{+}18]. discharge [PAR^{+}12]. disciplinary [TBF^{+}16]. discontinuities [PTBT17]. Discordance [TTMC16]. discrete [MHFK12]. discriminant [QQG14]. discriminate [AB11, CSOV14, KTW^{+}13]. Discriminating [BCMS^{+}18, BCN^{+}13, CFB^{+}17]. Discrimination [IHFAH17, LRY^{+}15, CBS11, FCNB16, FWG11, FCOM18, Fit12, HKRH12, HMA^{+}16, MREV10, PGVV16, RBG^{+}19a, WWN^{+}14, WRH^{+}15]. discriminators [SNP^{+}12]. discus [WKT^{+}11]. Disentangling [CBH18, HBS^{+}16, PRMKASR13]. disincentives [PICK17]. disinfection [WBD^{+}13]. disparate [PSJ^{+}19]. Dispersal [TWCA15, CN19, CAGVB14, HDW^{+}15, HLSD21, MHC^{+}18]. dispersed [FHK17]. dispersion [ALUFJ^{+}12, SFH^{+}15]. displaying [PHB^{+}19b]. disseminating [DDK^{+}18]. Dissostichus [BBHF10, CNWM12, OCL^{+}18, YZW^{+}19]. distance [LBZE17]. distances [LBZE17]. distant [CHL11]. distant-water [CHL11]. distinct [CFF^{+}16, LMN^{+}12, SMB^{+}17, VJH17]. distinguish [DR13, GHA13, KBG^{+}14, LKNL16]. distinguishing [GGS^{+}19]. distortion [WMdS14]. distributed [HKRH12, Pla18, SHM12]. Distribution [BST^{+}19, CAMFMF18, JS14, KCvOE13, MQV10, NS16, NBG^{+}17, OSMP18, PPG^{+}11, RLM10, SC19, YZW^{+}19, APL^{+}18, AS10, BSD^{+}13, BJP^{+}13, CSP^{+}18a, CSP^{+}18b, CR18, CSMO14, CAGVB14, CVS11, CDR^{+}11, CTS^{+}19, DM16, EMMM13, Eid12, FBHH14, FWIB14, GFCK13, GFO^{+}16, GMMW17, Goo16, GDA14, GFM^{+}12, GMUD11, HdSB^{+}14, HCHA12, HGG13, HCR14, JKK^{+}16, Kur12, LHSS18, LCM^{+}18, LWZ^{+}19, MMD^{+}11, MJC^{+}14, MMO^{+}13, OtHH^{+}10, OMI^{+}14, OGCBMR15, OÁPP12, ODMF17, PBC^{+}16, PTBT17, PCP^{+}18, PFJ19, PD14, SCC^{+}15, SWD^{+}13, SSH14, SB12, TOA12, TJMT17, TTF^{+}13, TGDB16, VSTS10, WDL17, WRR^{+}16, WOPB11, WBB^{+}16]. Distributional [VPSH^{+}17, BMG16, JBKA15, LCMTM10]. distributions [CZW14, CQ14, CRAV13, FBY^{+}12, GTS^{+}17, GDC^{+}19, Leg14, LCMdSM13, MGM12, MB16, SO18, SH13a, WMPR17, YCYC16, ZZA^{+}14]. distributive [APHL13]. diurnal [BSM15a]. dive [CUL^{+}16]. diver [ZBB17]. diverged [ER18]. divergence [JBD^{+}18, NC13, PGJ10]. divergent [CSKB13, HRB14]. diverse [MP17, MFMM16, WHK13, ZSF11]. diversification [OH12, WDM15]. Diversity [FHB^{+}19, GK18, CT16, CMMVRF^{+}19, CJG^{+}19, CK18, DS18, DUC^{+}19, GBM^{+}18, HBMC12, IÓS^{+}14, JWZ^{+}19, LFG11, PGAS15, PGM^{+}10, RMGK18, ŠBMG^{+}16, SRMP^{+}18, ZQZ^{+}17, dVA18]. divided [MKHK19]. diving [CMS^{+}15]. division [NPGT17]. d’Ivoire [KAGO^{+}10]. DNA [AHH^{+}12, BSD^{+}13, BdSFB^{+}18, CGdGT^{+}17, EL18, GTDJdL^{+}13, HIH19, HDW^{+}15, IdRH10, KBG^{+}14, KMdS^{+}18, LSR16, LCH^{+}18, Nij15, PF15a, PGAS15, PMM^{+}15, STY^{+}16, SPM^{+}15a, SAK^{+}17a, SRMP^{+}18, SLZG18, TC12, VRS13, WDL17, WCC^{+}21, YA18, ZQZ^{+}17]. Do [FSFP15, GIK^{+}16, JPM11, KJ17, MNAA^{+}14, RSEERRSC16, WDBH16, AAOE13, BAOP18, CGM^{+}19, CLG^{+}14, DCSW^{+}12, GCDAS^{+}10, JRG^{+}18, LBZE17, MRD19, MLTD17, NWO^{+}15, SKF^{+}10, WM17]. dockside [SUR16]. documented [MUO^{+}17]. Documenting [Ste18]. Does [BRB^{+}14, CGC13, DMFR16, HPNA16, JHT^{+}10, JNN14, JBMK13, KHKM15, MWC15, QHSI18, RFPE^{+}12, VKNK12, WBD^{+}13, ZZA^{+}14, HAB^{+}17a, HAB^{+}18a, JCT^{+}16, PL18, PHM16b, SHJ19]. dofleini [BTT11, BN12, CS18]. dogfish [BSKT13, CPS^{+}12, CAC^{+}14, HCF^{+}17, KM15, MS15b, SPC^{+}13, TGM13, TGMM13]. Dolly [LRY^{+}15]. dolphin [LCMF^{+}16, XLCMMV19]. dolphin-associated [XLCMMV19]. Dolphinfish [FBHH14, APDJSU11, CM12, CDF^{+}13, FTN^{+}14, MAH14, MAH16, SMB^{+}17, TVdLOG^{+}21]. domestic [DSB^{+}10, LBZE14, MAH16]. domesticated [KBL^{+}15]. dominated [EHB^{+}21, HVD^{+}18, LA13, WCS^{+}12]. Donax [DSG^{+}17]. Dongsha [XDL^{+}19]. Don’t [LR19]. d’Orbigny [CMSG18]. Dorosoma [JSB19b]. dorsal [Cor17, LRML^{+}15]. Doryteuthis [BHPC14, PG14, RDR^{+}18, RLSPdR14, WA15]. Dosidicus [AT10, HBG^{+}19, LGZC17, LFCC15, MBNM10, NMMBCV^{+}10, PAX^{+}17, RGGdL13, STY^{+}16, YC18]. Double [LBS^{+}18, HSTK16, LHS^{+}16, LHS^{+}18b, LHS^{+}18c, LH18]. double-reads [HSTK16]. double-tagging [LH18]. Dover [SDV^{+}16]. down [FP10, RFS16, WM17, WM13]. down-weighting [WM17]. Downstream [VBR^{+}18b, RIB^{+}10]. DPS [WRMM18]. drag [BSB^{+}15, BSM15b, BBS^{+}17]. drainage [PGVV16]. draughtboard [WWR18]. drawing [TBS13]. Dredge [MAAW^{+}17, KRM^{+}18, LHSS18, MMAA^{+}18, Mur11, WRR^{+}13]. dredges [NGMA11]. dredging [BH11, LBH15, UGGR^{+}17]. dried [TdMWC16]. drift [HSD^{+}17, MHB^{+}19]. drifting [CN19, DHF10, EBB16, JPM11, MST14, TGDB16]. driftnet [BAT12]. drinking [JAJ^{+}16]. drive [SWG18]. driven [LP18, TIGCDL14]. driver [NPRGTSR18]. drivers [CSP^{+}18b, CBH18, DH11, FMB^{+}19, MPR18, NEJ^{+}18, RCE^{+}19, WJC17, YZW^{+}19]. drives [FHS^{+}16]. droebachiensis [JEJS17]. drones [RG18]. drought [FCIB17, GS12]. Drum [HCWS19, BHB^{+}10, CBF^{+}14, OMB18]. drumline [GPT^{+}19]. Dry [ASB^{+}13, HHM^{+}13]. Dual [PJM14, HB12, SHGL21]. Dual-frequency [PJM14]. dual-grid [HB12]. duct [KWM^{+}15]. due [BGBCM19, CHA15, PHLT16, SWC14, WOPB11]. dumerili [JDNP18, ŠBMG^{+}16]. Duncan [Ano10a]. Dungeness [ZD13]. durabilities [BBM16]. duration [BS10, FWR10, FO11, Kur12]. during [ASB^{+}16, BST^{+}17, BHT^{+}19, BK15, BDB^{+}11, BCPP17, CDMC18, DW19, DHP^{+}11, DHFSS18, FKM11, GS12, JTP19, JAJ^{+}16, JČK^{+}12, JČK^{+}21, Koo12, KYO^{+}12, LHS^{+}18a, LWS19a, LL14b, Lyn14, MJL14, NS16, OÁPP12, PGG10, RSE^{+}14, ŘJP^{+}12, RWRT15, SKF^{+}10, TPM15, WOPB11, YRP^{+}19, ZDZY19, vPWW16]. dusky [MSRB15]. Dutch [BRH^{+}16, FGRIT^{+}12]. Dynamic [THKB12, Ber19, HPRB16, MB16, NPRGTSR18, RUMBA15, SLXZ15, TJROFS19]. dynamics [ABSI^{+}18, APB^{+}16, ABL21, AAN^{+}12b, BFM14, BOCÁ^{+}18, BKM13, BC18, BFL^{+}14, BBV17, CAC19, CMLP19, CEAP18, DW19, DL11, DH11, FMA^{+}18, FB19, GOD19, GFM^{+}12, HdSB^{+}14, HNY^{+}19, HRUSR13, HD12, ICSR16, JWC^{+}13, KY19a, KTL11, KCO14, LEO^{+}11, LPMM14, LPTK19, LLH^{+}14, LTHRU17, LKZ^{+}17, MSS^{+}16, MPB^{+}17, MDH15, MAH16, MQGK10, NGvdL^{+}13, NMMBCV^{+}10, OPR12, PT19, PDP^{+}11, PTW^{+}13, RRO^{+}19, RRW^{+}13, SWL^{+}15, SPM^{+}19, SMCR13, SF19, SJA17, SMP18, SDE^{+}18, SP15b, SBL^{+}19, TFH^{+}15, TK16, VOM12, WBB^{+}16, XLCMMV19]. Dyneema [HWS^{+}15].

Early [BKS^{+}16, ABSI^{+}18, AFAJ12, BDB^{+}11, BLPF^{+}15, CRS^{+}19, CMM^{+}14, HBM11, JB15, JČK^{+}12, JČK^{+}21, Koo12, MRS15, WBM17]. earth [GWWW13, WWR18]. East [JB15, KEMM12, LCZ^{+}16, MGP^{+}18, MLBGV16, BB13, CRB^{+}18, FP10, HSS^{+}16, MAH14, PGLG11, SJJ^{+}11, FTN^{+}14, KHGB14, PPLK18, Son18, YZW^{+}19, YGZ^{+}18]. Eastern [DV11, SLTPC^{+}18, AMQ^{+}15, dSMSF15, BHI14, BM18a, BKS18, CDMC18, CCT^{+}14, DW19, DUD^{+}18, DLCO^{+}19, EBB16, FH15, GS12, GPBA13, HLB^{+}15, KSKGR15, LBB^{+}13a, LBB^{+}13b, LCMF^{+}16, LSM^{+}19, LGH^{+}21, MVMSdS19, MBW12, MMAA^{+}18, NL11, NKW^{+}19, PGP^{+}15, PGP^{+}16, SF19, SWH17, SSS^{+}12, SYL^{+}15, SCL^{+}15, SP12, SP15b, TIGCDL14, TCS^{+}15, VvHE^{+}11, WBI^{+}16, Way13, WKTD13, XLCMMV19, ZCD^{+}12, ZGT12, APCF17, AMH19, BPK^{+}17, CY12, FPSDB17, HCGS19, KBG^{+}14, KHV^{+}12, KNB^{+}19, LFCC15, MAAW^{+}17, MRT17, MRT18, Mur19, OS12, PPA^{+}15, PRL^{+}18, THA^{+}14, TÜM^{+}16, Cor17, LRML^{+}15, VHS^{+}17]. EBFM [PW19]. EcAp [RBR^{+}15]. echo [LMSM14, OLM^{+}19, YAH^{+}13]. echo-sounder [LMSM14, OLM^{+}19]. echosounder [GFCK13, MML^{+}16, SMS^{+}19]. eco [BSB16, HH16]. eco-efficiency [BSB16]. eco-labeling [HH16]. ecolabelling [GDB^{+}16]. Ecological [Ber13, HSS^{+}11, ZHDS16, BLR^{+}16, DHF10, HBS^{+}16, HRC16, KBČ^{+}16, LG13, LJB16, LF13, MMDH18, NC13, PFVGGM16, PHG^{+}18, TBF^{+}16, VS16, YCQ11, ZSF11, ZCW11]. Ecologically [VH12]. Ecology [KBČ^{+}16, ABSI^{+}18, AGB19, BMS15, CEW14, DDP15, DPSR10, GAP10, HBG^{+}19, PSL^{+}19, SRB19, VdSS^{+}14]. Ecomorphological [VF10]. Economic [INSN14, MFIG16, PHG^{+}18, Wal13, WGM^{+}11, AJP18, APHL13, BMOZ13, BRC^{+}10, CMP^{+}12, CBH18, GHPJ15, GS12, GMD^{+}13, JNE19, KK12, KWCT^{+}18, LBHP16, LGZC17, MMNCR^{+}13, MQM^{+}15, NLPS^{+}13, OLK^{+}14, PBLT14, PBR^{+}16, Pre17, RM19, Seu21, SHR^{+}18a, TTS11, TBC^{+}19, WMM^{+}15]. economically [FBB^{+}17]. economics [GHP^{+}15, GVG18]. Economy [MCJD^{+}17, GRS^{+}15]. ecopath [AdM19]. Ecosim [FSFP15]. Ecosystem [BW17, FWJ^{+}16, RBR^{+}15, SPR^{+}19, APD10, BCKG13, BKSF17, BKM13, BWR^{+}18, BM19a, ETMK13, FSFP15, FP10, FSPWG08, GAR19, GDA14, GDC^{+}19, GPP^{+}16, HCZ^{+}16, Hil11, KZ11, LG13, LKR11, LMLBC13, MKH16, MMG10, MS15b, MML^{+}16, NMH^{+}11, NGvdL^{+}13, NBF^{+}19, NST^{+}17, PBM^{+}19, PHB^{+}19b, RLSPdR14, SHPSK15, SAB^{+}11, SGGH10, SGC11, SB14, TWSA16, TBC^{+}19, Tho19a, Tho19b, TKF11, VFMS16, YCQ11, YSS^{+}11, ZBB17, ZCR16, JKS^{+}15, JKK^{+}16]. ecosystem-based [ETMK13, FSPWG08, HCZ^{+}16, KZ11, LG13, LKR11, PBM^{+}19, SGC11, TKF11, YSS^{+}11, ZCR16]. Ecosystems [BKSF17, AENC14, BSF^{+}19, CT16, HMD^{+}19, MAM^{+}13, CGLS12]. ecotypes [RLJ^{+}12]. ectenes [RLJ^{+}12]. Ecuadorian [VIFLP17]. edible [JCA^{+}16, MDL19, RG18]. Editor [HL16]. Editorial [Ano10e, Ano10f, Ano10g, Ano10h, Ano10i, Ano10j, Ano10k, Ano10l, Ano10m, Ano11e, Ano11f, Ano11g, Ano11h, Ano11i, Ano11j, Ano11k, Ano11l, Ano11m, Ano11n, Ano11o, Ano12a, Ano12b, Ano12c, Ano12d, Ano12e, Ano12f, Ano12g, Ano12h, Ano12i, Ano12j, Ano12k, Ano13b, Ano13c, Ano13d, Ano13e, Ano13f, Ano13g, Ano13h, Ano13i, Ano13j, Ano14a, Ano14b, Ano14c, Ano14d, Ano14e, Ano14f, Ano14g, Ano14h, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano15k, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano16i, Ano16j, Ano16k, Ano16l, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano17g, Ano17k, Ano17l, Ano17m, Ano17h, Ano17i, Ano17j, Ano18a, Ano18b, Ano18c]. Editorial [Ano18d, Ano18e, Ano18f, Ano18g, Ano18h, Ano18i, Ano18j, Ano18k, Ano18l, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i, Ano19j, Ano19k, Ano19l, Ano19m, Ano21b]. edulis [YAT^{+}19]. edwardsii [BBB^{+}16, CGG21, EHGG16, FPM17, LPHH11, LSM^{+}19, LGH^{+}21, PTW^{+}13]. Eel [DBLB16, BA12b, BDHA11, CM15, CBS11, KY19b, KSAA13, KPF^{+}18, KT18b, LCST10, LT18, MLTD17, PJR17, VBR^{+}18b, VBR^{+}18a, WSF^{+}18]. eels [KPF^{+}18, KT18b, LTHRU17, MJL^{+}17, Sim15, dCJG^{+}15]. EEZ [PAX^{+}17]. Efanov [JG12]. Effect [EPHF18, FO11, GSH^{+}15, HSLB19, KNSJJ11, MWR13b, OHSG19, RWRT15, SL11a, SKFM17, SLT^{+}19, SGW16, SZPP17, SHG^{+}15, SABG12, SRG17, TXH^{+}19, TSHB19, WLGS19, WHMS11, BSB^{+}15, CFPF^{+}15, CPB^{+}12, DCP^{+}16, DTSG^{+}16, Fey18, GWWW13, GBD^{+}15, HRB14, HB12, HPHR14, HL16, HSRT17, HCR14, Ing11, KDSAH^{+}18, LW19a, MOA^{+}16, MS18, NKW^{+}19, Pun18, RR12, SL10, SA16, SHG^{+}16, Szu16, TXZ^{+}17, Han16b, HM16a, HM16b]. effective [BC18, BBH^{+}14, CHA15, Chi10, FWJ^{+}16, GHUG^{+}15, HPRB16, KKE^{+}12, Mau11, OPF15, PFVGGM16, PPVP17, RIB^{+}10, TJMT17, TBSE17, WK10]. effectively [WDBH16]. Effectiveness [BBA^{+}19, CSMJB11, MUO^{+}17, UEOA19, WTF16, AAH^{+}16, BEM13, HJ12, LPHH11, MMSUAGC14, SMK^{+}13, TBF^{+}16, WMC^{+}16]. Effects [BHB^{+}10, BNAE10, BKT10, BRS19, BSC12, BSM12, BSM15a, BBS^{+}17, BM19b, Bro13b, CQ14, CSFCA15, CRH^{+}13, CBMLB13, CCT^{+}14, CKC16, CB11, EDG16, FCCSA15, FWR10, GDW^{+}15, HSS18, HFBS15, HF19, HFL14, Hen11, IGSJBS^{+}15, KNSSL11, KL16, KY17, LDR^{+}12, LvVtB^{+}15, LWH19, MMGEH15, MCMK18, MLV13, OHLBL16, OGCBMR15, PB11, PDBM12, RMT^{+}18, ROCC13, RDM^{+}15, SSFGL19, SGSFL19, SKFL15, SGA^{+}18, SNK19, SBS17, SMS17, SMP18, SFL^{+}14, SKDO^{+}17, TD19, VWHB14, WAB^{+}16, WRR^{+}13, YVJ12, AMHR16, BFM14, BEM^{+}17, BMG16, BBA^{+}19, BBT^{+}15, BDM^{+}14, BDB^{+}16, BPG^{+}17, BSM16, CRAP^{+}11, CCC^{+}19, CGKG^{+}17, CNS11, CADG^{+}18, COS15, CW17, Chi11, CKVV16, EGK15, EPO10, FCIB17, FCL16, FT16, FMMW13, FBB^{+}17, GYCL19, HLJ15, HFPL16, HL16, HSS^{+}11, HWS^{+}12, IOT14, IEL17, JSR10]. effects [JDBF15, KCO14, KCBS13, LSŠ^{+}16b, LMBF13, LMLBC13, LMM^{+}17, MK17, MMMPP19, MD16, MHB^{+}19, MAA^{+}18, MABA12, MDG17, MGB10, ML17, MDPH17, MJL^{+}17, NSM^{+}18, OSO^{+}18, OÁPP12, PSU^{+}16, PTTC18, PHB19a, PACF17, RHB^{+}17, SAR^{+}15, SPM^{+}19, SVPC16, SBD^{+}12, SAK^{+}17b, SDV^{+}16, SD13, TNKO19, TW14, TK16, TMV16, WGC^{+}19, WGC^{+}21, WMM^{+}15, WMvdL10, WKT^{+}11, XGT^{+}17, ZCX^{+}18, ZSF11, ZHDS16]. Efficacy [GCL19, SBJ^{+}19, Gri10, HHM^{+}13, HB15a, HB15b, Rob15, WHC^{+}19]. efficiencies [BBM16, HNM^{+}12, PJM14]. Efficiency [DF16, JEJS17, BSB16, BNAE10, BNAE13, BLHG17, BKLB14, CDR16, CPMNS10, CK18, DMFR16, DWH10, DFAN12, EPHF18, FL13, FPSDB17, GHS^{+}19, HPAA19, HSRT17, IEL17, JČK^{+}12, JČK^{+}21, LBB^{+}13b, LWH19, LOC^{+}11, LCLM15, MAAW^{+}17, MMAA^{+}18, PAH^{+}13, PHBU19, SHG^{+}15, SHG^{+}16, SMI17, WHS^{+}18, WW11b]. efficient [BRH^{+}16, BCMS^{+}18, LR19, NGMA11, SCG^{+}12, TM13]. effort [AVP^{+}15a, AWGP11, BNSS19, BNU^{+}10, BNAE10, BHG^{+}15, BB10, BDHA11, BC18, Cam15, CZW14, CAC19, CMP^{+}14, CCM14, CAR19, CDR^{+}11, DMOL15, DSB^{+}10, EHH^{+}11, ERG^{+}11, ELRH17, Foc14, GCDE^{+}15, Gri12, GWB^{+}19, GRB14, HPHF15, HFFdlT15, JF13, KSL^{+}16, KY17, LMHR19, LHLR17, LPC^{+}14, LvVtB^{+}15, LMLBC13, LT16, Lyn14, MCEP14, MBvDN14, MWC15, OMI^{+}14, OHLBL16, OPH15, PD18, PMG18, PLPN12, PGS^{+}11, PDP^{+}11, RPP^{+}11, TBS13, THF12, THKB12, VAB^{+}14, WCS^{+}12, WKA13, WKA14, YvZG^{+}17, ZGT12, vPCWV15, vPWW16, vPB18]. Efforts [LW12, BB19]. Egg [BSW^{+}12, GN12, GSTL15, MFRM19, NNP17, SSS^{+}12, TEGM^{+}11, AW12, BJD12, BJP^{+}13, CDPJ^{+}13, CSKB13, DCSW^{+}12, HHH^{+}12, HW11, JCW12, KHV^{+}12, MABA12, MDPCSR10, MPA^{+}12, MRT17, MIÁ^{+}10, NNS^{+}10, OÁPP12]. eggs [GHA13, OÁPP12, PUSI10]. Egypt [GKMGO^{+}17]. Egyptian [GKOC^{+}16, SAK^{+}17a]. eightbar [WWN^{+}15]. Elasmobranch [BBM^{+}10, Cro15, AHC^{+}11, CSMEO^{+}11, OPL17, SB15]. Elasmobranchii [SAH15]. elasmobranchs [CEW14, Jab18, KF18, SZR^{+}11, YRP^{+}19]. elasticity [BDHA11, SKH14]. elastomer [BN12, LWY^{+}11, PTTC18]. Elbe [Sim15]. electric [BA12b, PWM^{+}16]. Electrical [EDG16, LDR^{+}12]. electrified [SGT^{+}14]. Electrochemical [MK13]. electrodes [JTHT14]. electrofishing [BAB^{+}12, FMD19, HdED^{+}13, HDS^{+}18, MDS^{+}18, MFMM16, MCB^{+}16, PTBT17, STCE12]. electronic [BFW13, Cra19, PHBU19, SEG^{+}15, TW10, WMB12, WMB16]. electropositive [RPK11]. electrotrawling [SDV^{+}16]. Eledone [RGG14]. eleginoides [BBHF10, CNWM12]. element [FMF^{+}17, FBF^{+}18]. elemental [ADSV17, APD^{+}18, CPG^{+}11, CHCS14, DR13, FWG11, SSD^{+}19]. elements [LRY^{+}15, MFC^{+}13, WW14]. elephant [MWR18]. elephas [ARVACPMN16, BMF^{+}10, BMF^{+}11, GVDMG12]. Eleuthera [MCD^{+}13]. Eleutheronema [BWN^{+}12]. eleven [PVKH15, PVKH18]. elongata [Ata11]. elongatus [BBC11, KMH^{+}12]. elusive [MPF^{+}17]. embayment [HKM^{+}17]. Embracing [Lor16, Mun18]. embryo [BRS19, KC11]. embryonic [MNT^{+}17]. emergency [COS^{+}19]. emerging [DDK^{+}18, GRS^{+}15]. emigration [LTHRU17]. Emmelichthyidae [NL11]. Emmelichthys [NL11]. emphasis [FP10, PF15b]. Empirical [DDH^{+}15a, CZW14, KOHF^{+}16, MCM12b, PML^{+}12, PS12]. enable [FWJ^{+}16]. encircling [LOC^{+}11, LCF15]. enclosed [RRW^{+}13]. enclosure [CM15]. encountering [NWL^{+}14]. encrasicolus [GGB^{+}18, BA12a, BGF^{+}15, CMI^{+}16, CSP^{+}18a, JBKA15, RGAP18, YA18].

End [SDE^{+}18, BGD11, BSM14, KW10]. End-to-end [SDE^{+}18]. Endangered [CHA15, LKNL16, MFRJ13, SPM^{+}15a, SBJ^{+}19]. endemic [CLS^{+}19, CMSG18, MFRJ13, SMB^{+}17]. energetic [KSAA13]. Energy [GPS15, BNAE10, BNAE13, BCPP17, DPSRM10, LRM14, LCWC17, MAY12, PGJ10, SKK^{+}13, WMF13]. energy- [SKK^{+}13]. enforcement [EHH^{+}11, PAH^{+}13]. engine [ERG^{+}11]. engineer [GPP^{+}16]. Engineering [BSM14, SB17, BSM12, BSM13b, NWO^{+}15]. England [EPCB17, HKM^{+}17, KRM^{+}18, LF13, MLM15, SPC^{+}13, WDM15]. English [PRL^{+}18, ARF^{+}17, ELRH17, RBM13, VKM^{+}17]. engrauliformis [ZKSF17]. Engraulis [BA12a, BGF^{+}15, CMI^{+}16, CSP^{+}18a, CCC^{+}14a, CCHD19, DPO^{+}11, GGB^{+}18, JBKA15, KBG^{+}14, RGAP18, SSB^{+}18, TOTI10, YA18]. enhance [GPT^{+}19, ZTC^{+}18]. enhanced [WB16]. enhancement [CBF^{+}14, GCL19, HBT16a, HLSD21, PB11, YVJ12]. Enhancing [Ano13a, May10]. enough [CH15, PD18]. Enriched [WMIDS18]. Enrique [Cor17]. Ensemble [GAR19]. Ensis [COC14, FMD19, MCB^{+}16]. ensuring [GGS^{+}19]. entangling [CGR^{+}15, CGR^{+}16]. enter [OS19]. Enteroctopus [BTT11, BN12, CS18, ORMG11]. entrance [JLF^{+}17]. entrances [JLF^{+}17]. entry [MAH16, Pun18]. envelope [TC16a, TC16b]. environment [AMV14, HA12, KHS12, LKL^{+}12, LLH^{+}14, MJC^{+}14]. Environmental [ALBŠB^{+}15, KCO14, LLWC15, SSP^{+}15, WA15, YZW^{+}19, BSF^{+}19, BCP^{+}12, BSM15c, CJCN21, CMLP19, CBH18, CDR^{+}11, EHBK18, EL18, FSG^{+}19, FF11, GCDAS^{+}10, GFM^{+}12, Han16b, HFBS11, HPHR14, HM16a, HM16b, HFL14, IÓS^{+}14, JSR10, KNSJJ11, KY17, MSS^{+}16, MCB^{+}18, MAP14, MWSN^{+}13, MAA^{+}18, MRT17, OALS11, OGCBMR15, Pit14, PSS14, RFPE^{+}12, SSS^{+}12, SGRS^{+}18, TXZ^{+}17, TXP^{+}16, UKK^{+}12, WHP15, WJC17, YLA15, ZD19]. Environmentally [BPMGP10, BSM13a, HBB^{+}19, LMN^{+}12]. environmentally-informed [HBB^{+}19]. environments [BdGM^{+}11, BB17, BB19, ESF^{+}19, HZT^{+}19]. eperlanus [GFCK13]. epi [BST^{+}19]. epi-benthic [BST^{+}19]. Epigonus [CRAV13]. Epinephelus [CJG^{+}19, GTS^{+}17]. epithelial [CC11]. EPMs [GN12]. equal [Har19, Pun18]. Equations [CIS13, JSB19b, TC15]. equatorial [EBB16, FSH^{+}15, MST14, OCN^{+}11, PKH^{+}11, SFH^{+}15]. equilibrium [PLM16, WSMS14]. equipment [STCE12]. equipped [LCHL11, LCF15, MAY12, PBB^{+}15b]. equivalent [FPD^{+}16]. Erie [BJZB12]. Erratum [BMF^{+}11, CP12, KPL^{+}09]. error [BCP^{+}12, CCC^{+}19, CNWM12, HSTK16, Leg14, MCP^{+}17b, MB16, SM17, WWSM15, XLCC19]. errors [BST^{+}17, CNWM12, CRAVC14, WRR^{+}13]. Escanaba [BNSS19]. Escape [AOC11, BBM17, Kim15, MM13, STJ11, ALUFJ^{+}12, Bri10, BMH17, BM18a, BM18b, BMH18, FBF^{+}18, FA19, JLF^{+}17, LCBD12, LPHH11, MFHK10, PALC19, TLLZ19]. escapement [HKF^{+}12, KW10, LHS^{+}16, LW19a, MLTD17, MBF19a, WTF16]. escaping [DÖA^{+}10, HJ12, KK15]. esculentus [KCO14]. Esox [Cur18, KKA10, LWSC11, PALC19, SMØ16, SLB^{+}12, SFL^{+}14]. Essential [FSH^{+}13, MLP^{+}14, DHF10]. Establishing [NGvdL^{+}13]. estimate [BHB^{+}10, CA10, CN18, CAC15, DPAB10, DHC19, DBLB16, GBF11b, HGD19, HHHWJ18, JCW12, KFGN10, Leg14, LSL12, MUO^{+}17, RMB^{+}21, RMP^{+}18, SZPP17, TB10, XTP^{+}16]. estimated [ASD^{+}18, EHB18, IAS^{+}17, JCT^{+}16, JC18, LMPM12a, NS16, PS12]. Estimates [GSP17a, JLLL11, KHGB14, TPM15, ZRH^{+}11, ALH^{+}17, dSMSF15, BST^{+}17, BGL^{+}13, BWR^{+}18, CGM^{+}19, CRAV13, Coo13, DBG19, DMK^{+}19, EHB^{+}21, ETMK13, EMSC15, FPC^{+}13, Fra12, GFK^{+}12, GDB19, GAB^{+}15, GAR10, HST10, HSTK16, HBG14, HdED^{+}13, HDS^{+}18, HLB^{+}15, JDNP18, JSB19a, KM10, KML11, KKE^{+}12, LMHR19, LMPM12b, LCA^{+}19, MABA12, MBF19a, NKW^{+}19, OCL^{+}18, PKWP15, PL18, POM^{+}18, RG18, RCHSN^{+}15, SMdUK16, SWC14, SGC11, SHT^{+}13, Szu16, TWM11, TJMT17, VB15, WAP^{+}12, WWSM15, WOPB11, WRR^{+}13]. Estimating [AVP^{+}15a, BHCJ12, CRPAMN17, CNWM12, CFH14, CFH15, Cer14, EMSC15, FQT17, Foc14, FdSM^{+}16, GKL12, KSL^{+}17, KDPT18, KWR19, LBHP16, LHS^{+}16, LMPM11, Mau11, MBW12, PTC^{+}18, PPH^{+}17, PHM16b, RSGR19, Seu21, SHR^{+}18a, SBA10, TBSE17, WSF^{+}18, WMB12, WMB16, WFNC18, ZCC11, dZB16, vPB18, AW12, BHI14, BNK14, BAFP18, BZ12, DCSW^{+}12, DHP^{+}13, DOO15, FL13, FJWD17, GH15b, HPHF15, KY19a, LWS19b, MW11, Mur15, PLPN12, PJM14, PR18, RCLS18, RBW14, SBJ^{+}19, TJJ14, TRW19, WM11, YJC13]. Estimation [ATG11, CIS13, DBLB16, GPSF10, HF14, HSRT17, LWS19a, MRT18, NB14, ODdA^{+}10, SZPV16, ZMdSC16, AL13b, ASR^{+}18, BTS^{+}19, CFBS16, COS15, CM12, Chi11, FWGM19, FS19, HFPL16, HF19, HPB10, Hoc12, HKF^{+}12, IOT14, KT18a, KGR^{+}15, KTNA15, Kur12, KOHF^{+}16, LBV13, LBF^{+}16, MCVS14, MCP^{+}16, MLF^{+}10, MFMM16, ML17, MOA^{+}16, Mun18, PRBGS^{+}10, QQG14, QWC10, SPB^{+}19, SP12, TSS11, TEGM^{+}11, UASM12, VIP13, ZH10]. estimations [CMJ^{+}19, FO11]. estimator [Car18, ST13]. Estimators [JF13, AMGH15, XLCC19]. estuaries [CBF^{+}14, FF11, LAIVA^{+}19, MBW12, SWL^{+}15]. Estuarine [CMM^{+}14, AMV14, BAY^{+}12, BMB10, LA13, MSN^{+}12, POV19, RRO^{+}19, RJKG12, TRC19, WW14, WJC17]. estuarine-dependent [BAY^{+}12]. Estuary [GMMW17, CDR^{+}11, DET^{+}16, MBS14, GLA19]. Eteline [WWN^{+}14]. EU [AAH^{+}16, Dam15, NCP15]. eulachon [HJLW11]. Euphausia [HKK18, KWR15, KK15, ZYSZ18, ZDZY19]. Eurasian [PGVV16, VKNK12]. eurocutter [DDB^{+}14]. Europe [CLF^{+}17, CCC^{+}14b, FGRIT^{+}12, GCDE^{+}15, GBF^{+}11a, GBK^{+}13, NEJ^{+}18, UARC11]. European [BGF^{+}15, BCMS^{+}18, BCN^{+}13, BLR^{+}16, BPS^{+}19, Bow17, CRAP^{+}11, CRCMP^{+}14, CPD^{+}19, CMPP11, CMP^{+}12, CSP^{+}18a, CSP^{+}18b, CBS11, CPCP^{+}12, DRGGVL11, DPSRM10, DPSR10, ERG^{+}11, EÖFPH17, Eth15, GPSA11, GMMW17, GPBA13, IGS^{+}12, JBKA15, KMKK10, KDPMSR10, KPF^{+}18, LCM^{+}18, LSC^{+}18, LČ18, MCC^{+}17, MBRDM18, MDPCSR10, MCM12a, MLBGV16, MIÁ^{+}10, MQGK10, NSM^{+}18, PRP^{+}18, PG18, PB11, PHB19a, PLSM^{+}16, PPVP17, RFC^{+}17, RSV^{+}17, SD10, SSR18, SRB19, SMJB11, Sim15, SJA17, SGT^{+}14, TVCT12, VBR^{+}18b, VBR^{+}18a, VMB18, VSTS10, WMC^{+}16, WSF^{+}18, WSG16]. euryhaline [SBT^{+}18]. eutrophication [SVR16]. evaluate [CCA16b, CCC^{+}19, FEGMTÁ17, HHM^{+}13, Mac13, SW19]. Evaluating [AVP^{+}15b, BRJ^{+}16, BCC15, BDB^{+}16, CWM12, CPW^{+}14, CTWD15, DB12, DHP^{+}13, DBH^{+}10, DQMV19, GBR^{+}18, GRB14, HJ12, IIT17, IIT18, LZX^{+}17, LWH19, LFS^{+}19, LMLBC13, MCB^{+}18, Mau12, MCH11, PML^{+}12, RBT14, RM19, SMI17, Sto12, SBC^{+}18, TB10, WP15, WFJ19, WDS^{+}11, WLBB15, XGT^{+}17, YGA^{+}12, YRH15, ZCZ^{+}13, ZCX^{+}18, BSB16, FTM^{+}15, NGvdL^{+}13, YA18]. Evaluation [AMGH15, AVU^{+}10, AK11, BNK14, CC11, CV14, CCGD19, GLB^{+}19, GMSP19, GGS^{+}19, Gra16, GWB^{+}19, GAR10, HSB19, HRC16, HMA^{+}16, Jaw11, JAJ^{+}16, KKK16, KH15, LWSC11, LCST10, LW16, LFG11, OMI^{+}14, PLM16, PR18, PBD21, SPMS17, SJS^{+}17, TCAP11, TBF^{+}16, TFB^{+}14, WMP^{+}14, WXZ^{+}18, WHNS13, XdA19, Zha13, APD10, BA12b, BVN^{+}10, BZ12, CHS^{+}17, DMC^{+}12, EHH^{+}11, FBB^{+}17, HBB^{+}19, HLP15, JCW12, KY19a, KMSRW10, KWF12, LMPM11, LCRW19, MBJ^{+}16, Mau11, MVMSdS19, Nak17b, OSO^{+}18, OTK^{+}16, PSS14, Pun19a, RPH^{+}18, RGAP18, SAK^{+}17b, SCF^{+}17a, UASM12, WBD^{+}13, WWSM15, ZCR16, ZD19]. evaluations [FSFP15]. event [BCPP17, FBF^{+}18, LLM^{+}19]. events [BRS19, BJKH12, RFPE^{+}12]. eversion [BNM11]. Evidence [AW17, CHCS14, EB15, FGRIT^{+}12, JBKA15, JTP^{+}14, LSM^{+}19, LGH^{+}21, MP18, MAH16, PKWP15, RFS16, Ste11, UJLC^{+}19, BTBP18, FCLH19, GSB^{+}19, KOD^{+}15, LNC^{+}15a, LNC^{+}15b, LAIVA^{+}19, NWS^{+}11, LL14a]. Evolution [LMSM14, EFWC^{+}19, GVG18]. evolutionarily [VH12]. Evolutionary [RCF^{+}17]. Evolving [BBK12, Gan13]. ex [GM15b, RSEERRSC15]. ex-vessel [GM15b]. examinations [Fey18]. examine [EAAA^{+}18]. Examining [LLM12, WHP15, Wri13]. example [BSF^{+}19, BK10, Goo16, JS15, LPMM14, LCMTM10, LT18, PML^{+}12, VH12]. examples [GMSP19, TK16]. excessive [QHSI18, SH13b]. exclude [WDBH16]. Excluder [MBK19, WDBH16]. excluding [BHCS15]. Exclusion [HB15b, CKVV16, HB15a, KLH^{+}15, LCBD12, MQGK10, Rob15]. exclusive [KWCT^{+}18, LGZC17]. exemplified [KSAA13]. exhaustive [KDSAH^{+}18, SKDO^{+}17]. exist [WUGG16]. exit [LOdH16, MGC^{+}11, vPWW16]. exogenous [BK10]. exophthalmia [RLFJ11]. expand [MBF19a]. Expansion [ERL15]. expectations [PGL^{+}13]. Expected [Pre17]. expenditure [Gri12]. experience [But16, GCDE^{+}15]. Experiences [PKHD16, PRCGD16]. experiment [SKY^{+}17, TDH19, VIP13, WGM^{+}11, WT16, WSF^{+}18, WMF13, ZXHK15]. Experimental [CHW^{+}18, GHM16, HSC19, MH14, BBHF10, GBH^{+}19, LBH15, LDS^{+}19, MMO^{+}13, RBW14, SBS17, SEG^{+}15]. experimentally [WTK10]. Experiments [MTH10, Eth15, HLNH15a, HLNH15b, Kim15, MFIG16, PMB^{+}11, RSEERRSC15]. expert [DAJMB16, GMHPV^{+}10]. explain [ZZA^{+}14]. Explaining [Fra16a, LBZE17]. explicit [GBH^{+}19, GKL12, GDA14, GDC^{+}19, LTHRU17, SAB^{+}11, Whi10].

Exploitation [dADSRG10, BHI14, BKM19, BHS16, DJF^{+}15, FAAM15, FAAM16, Ken14, MSS^{+}16, OOM15, PHK13, PHK14, QWCB11, SJR^{+}11, SMPB17, VOM12]. exploited [AA13, Ara15, BMB^{+}18, BSF^{+}19, BMdS^{+}15, BPMGP10, BB13, CCA16b, CPH12, COC14, FPM^{+}18, FBY^{+}12, GJRR14, HND^{+}17, JF13, JWM^{+}12, MMDH18, NMH^{+}11, PKAL13, RRM16, SD10, Ste11, TMV16, VRS13, VRW11]. exploiters [Sal18]. exploration [ASH^{+}17]. Exploratory [LCMTM10]. Exploring [APCF17, GBH^{+}19, HFFdlT15, KRC14, PH12, SDV^{+}16, SCF17b, SDE^{+}18, dVA18]. Explosion [LAB13]. exposed [BPC^{+}12, GJRR14, SGB^{+}10]. exposure [CDMC18, FWR10, LSC^{+}18, LLM^{+}19, MS15a, PRP^{+}18, RSQ18, SLT^{+}19]. Expressed [RCF^{+}17]. extended [FPM17]. Extending [PC19]. extension [PMB^{+}11]. extent [GMJ^{+}13, MD13, PBB11, SBA10]. External [TSHB19, BPV^{+}17, DMC^{+}12, NSM^{+}18, RCE^{+}19, ZYW^{+}14]. extinction [MP18]. extract [PGJ10]. Extracting [LP18]. Extractive [BAOP18]. extreme [BJKH12]. extremely [KBT^{+}19]. eye [FPS11]. eyed [JECR^{+}18]. eyes [vPKF^{+}19]. eyestalk [ZYSZ18].

F. [FdMCB^{+}19, TTMC16]. F1 [SBAK17]. F1-hatchery-reared [SBAK17]. faber [SSD^{+}19]. Fabricius [GVDMG12]. face [PTW^{+}13, PHT15]. facilitators [KLB17]. facilities [CPB^{+}12]. facing [HA12]. factor [CN19, NNS^{+}10, RGAP18]. Factors [CM18, HHCP13, MTC^{+}17, PS19, SMO^{+}18a, BHC10, CHS^{+}17, FPSS15, FF11, GH15a, GCDAS^{+}10, GFM^{+}12, GSDD11, JATBC19, JHCT11, KNSJJ11, LBZE17, LvVtB^{+}15, MAP14, MHRD16, Mur15, NPRGTSR18, OGCBMR15, SOY14, SSS^{+}12, WHP15, Pul17]. FAD [EPTG19, GMJ^{+}13, HHH15, LFS^{+}19, MST14, SL11b]. FAD-association [HHH15]. FADs [GSDD11, JPM11, DHF10, EBB16, MDC^{+}16, SMRS15, SCV^{+}12]. fail [MH14]. failure [FL13]. Failures [SK13]. Fair [BBOL16]. falciformis [GTDJdL^{+}13]. Falkland [AFAJ12, BBHF10, LAB10, WA15]. fall [DLCK15, HBMC12, HSC^{+}12, HKF^{+}12]. fallax [BNA^{+}19, BPV^{+}17]. false [BOBP^{+}17]. Family [TTMC16, TC12, MCI^{+}13, ZQZ^{+}17]. Farfante [FdMCB^{+}19]. Farfantepenaeus [FdMCB^{+}19, TTMC16]. farm [ALBŠB^{+}15, FBF^{+}18, KDdAA^{+}18]. farmed [ALUFJ^{+}12, KPF^{+}18, KT18b, PJR17, SAK^{+}17b]. farming [RBG^{+}19b]. farms [BHL12, CKVV16, IGSJBS^{+}15]. Faroe [OPE^{+}17, OSMP18]. Faroese [GWB^{+}17, OAPS13]. fasciatus [PHRH16a]. Fast [WMPR17, BCP^{+}12, KS12]. fast-flowing [BCP^{+}12]. faster [BCMS^{+}18]. fat [SKT18]. Fate [BBM19, BHB^{+}10, BRGS19, Jab18, PALC19, WTK10]. Fatty [MCP14, SRG17]. fauna [CLB^{+}10, MCB^{+}16]. faunas [FTWC18]. Fcube [IGS^{+}12]. Feasibility [RTML17]. features [GPdR11, IAJ^{+}11, ZYW^{+}14]. February [Ano10r, Ano11t, Ano13p, Ano14i, Ano15s, Ano16r, Ano17w, Ano18r, Ano19t, Tho19a]. fecund [MD13]. Fecundity [RRRM19, TWM^{+}10, VFFG17, ZMG^{+}13, AW12, BK16, BTNA13, CBMLB13, DBMR17, DMK^{+}19, FQT17, GRVN10, Gan13, Gun17, HTHK11, HRUSR13, HFBS15, HNS12, Ken14, LXCC19, MQCS^{+}14, PHK13, PHK14, RM10a, Rod19, RLMM18, SMdUK16, SGSFL19, SKK^{+}13, SD13, TEGM^{+}11, VCS^{+}19, WBM17, WAT^{+}13, dEWB^{+}15]. federally [BT16]. feed [AJOE17]. feeders [BAL^{+}19]. Feeding [RLSPdR14, UK14, VIFLP17, Ata11, BA12a, BKT10, DBA18, JPO^{+}19, JPM11, MK17, RDV11, TV13, TV14]. feeds [RBG^{+}19a]. Female [PEAC11, BLW^{+}18, JPSR^{+}14, KJ17, KC11, LCWC17, OAT^{+}16, PGL^{+}13, QWMD13, Rod19, WBM17]. females [BPMGP10, GSTL15, LRM14, MCP14]. Fenneropenaeus [SLZG18]. Fernández [AOC11]. ferruginea [HHSS17, KRMW15]. fidelity [FF10, HGG13, ML10, MNW12, SSP^{+}15]. Field [GLG^{+}11, WTS^{+}17, BST^{+}19, BGD11, CDR16, CFBS16, CMJ^{+}19, DSW^{+}17, GMSP19, MFIG16, YA18]. field-based [CMJ^{+}19]. field-caught [YA18]. fields [BDB^{+}11]. Fifty [FH15]. Fight [RSQ18]. Fighting [BK19]. fillet [DTSG^{+}16, DRE^{+}17, KDSAH^{+}18, PMM^{+}15]. fillets [GMCL^{+}18, VRMF11, VRMF12]. filter [LCZ^{+}19]. filtering [SFB11]. fimbria [HCHA12, HKB14, JC18, KSH19]. fin [CMB11a, CFB^{+}17, Cor17, HdSB^{+}14, LRML^{+}15, LKNL16, MMMPP19, Nij15, PGP^{+}15, PGP^{+}16]. fin-muscle [MMMPP19]. fin-weight [PGP^{+}15, PGP^{+}16]. financial [RAG19]. finding [BBOL16]. Fine [BMCS19, KWCT^{+}18, MJL14, SH13a, BSSH13, CK18, EFW^{+}11, JWC^{+}13, MAS17, ÖLTA19, TRH16]. Fine-scale [BMCS19, KWCT^{+}18, MJL14, SH13a, EFW^{+}11, ÖLTA19]. finfish [HB11, HB13]. finfishing [KMMV10]. fingerprint [ADSV17]. fingerprinting [dJSMROZP^{+}13]. fingerprints [DR13]. finite [CSMO14]. Finland [HL16, KK12, PAM^{+}18, SAR^{+}15]. finned [SP10b, SRB^{+}18]. Finnish [LHLR17]. FINS [GKMGO^{+}17]. finspines [CAC^{+}14]. First [BCTS11, Ech17, GDB19, HBG14, HLB^{+}15, OCL^{+}18, SHC14, DBG19, Fra14, JAJ^{+}16, MMF^{+}17, SPC^{+}13, UIH^{+}18, WMC^{+}16]. Firth [HSH15, HSH16]. fiscal [FDN^{+}11]. Fish [ASHH15, AGS^{+}16, BTBP18, BAB^{+}12, BR15, BMF^{+}11, CLP^{+}19, CP12, DRE^{+}17, FSH^{+}13, GVSB11, GWM18, HAB^{+}18a, HLNH15a, Ibá14, Ibá15, IIT18, JSR10, JBMK14, KGH^{+}12, KPL^{+}09, LKD18, MBK19, MDC^{+}16, Mun12a, PGP^{+}16, PMK^{+}10, SRT13, SDJ^{+}15, SB15, TC16a, TV14, Tho19a, Tho21, VRMF12, VBdB^{+}18, WMB16, APDI13, AWSLS15, ALJC19, AAS19, AVP^{+}15b, AVP^{+}15a, AA13, AM12, Ara15, APB^{+}16, AHH^{+}12, ALBŠB^{+}15, AW12, ASB^{+}13, AL13b, BSS^{+}13, BPS^{+}18, BAY^{+}12, BGL^{+}13, BST^{+}19, BDM^{+}18, BLR^{+}16, BWR^{+}18, BLP^{+}18, BJKH12, BK15, BDB^{+}11, BBLS17, CRS^{+}19, CT16, CTSD10, CY12, CCC^{+}19, CLS^{+}19, CADG^{+}18, CA10, Che10, Chi10, Chi11, CC11, Coo13, CMJ^{+}19, Cra19, CKC16, CBH18, DPB^{+}14, DJD12, DDB^{+}14, DM16, DCP^{+}16, DM11, DCSW^{+}12, DMQ^{+}18, DBSA18, DÖA^{+}10, EBB16, EAAA^{+}18, EPTG19]. fish [EHB^{+}21, EvHRF10, ETMK13, FCIB17, FTM^{+}15, FCLH19, Fit12, FMH13, FBF^{+}18, FTWC18, FGRIT^{+}12, FGCG^{+}17, FdSM^{+}16, FJWD17, FKM11, GKOC^{+}16, GKMGO^{+}17, GLB^{+}19, GBM^{+}18, GMSP19, GMJ^{+}13, dMGPR^{+}18, GFK^{+}12, GIK^{+}16, GBK^{+}18, GMCL^{+}18, GGM^{+}18, GRO^{+}18, Gri12, GM15b, GYCL19, GAR10, HJ12, Han13, HKAA18, HFL14, HND^{+}17, HRAK15, HPNA16, HZT^{+}19, HSH15, HSH16, IO11, IHFAH17, IGSJBS^{+}15, JG12, JFT^{+}15, JWZ^{+}19, JATBC19, JSB19b, JP16, JAJ^{+}16, JRD^{+}13, JČK^{+}21, KCJ19, KWM^{+}15, KSL^{+}17, Ken14, Kim15, KMSRW10, KFGN10, KOD^{+}15, KDdAA^{+}18, KY17, KMKK10, KTW^{+}13, KS12, KBČ^{+}16, KAiT^{+}13, Kur12, LNC^{+}15a, LNC^{+}15b, LAB10, LBZE14, LWCC15, LCM^{+}18, LSC^{+}18, LL14a, LBZE17, LDS^{+}19, LC19, LFG11, LL14b, MS15a, MRD19, MMGEMGS10, MHK^{+}17, MBOAA12, MWR18, MMDH18, MDS^{+}18, MAA^{+}18, MBF^{+}17, MPA^{+}12, MKHK19, MW13]. fish [MCM12a, MWC15, NBS15, NB19, OS19, OPF15, OJT^{+}14, PMC^{+}18, PWSP12, PWM^{+}16, PRP^{+}18, PG18, PKAL13, PSL^{+}19, PLMT^{+}11, PL18, PGJ10, PJM14, PTBT17, PBB^{+}18, PMM^{+}15, PHK13, PHK14, PE14, PSJ^{+}19, PVKH15, PVKH18, PGVV16, Pul17, PS19, PDF^{+}18, RTŘ^{+}12, RSE^{+}14, RFC^{+}17, RBGE^{+}10, RCBPK11, RSEERRSC16, RAGP18, RJKG12, RHB16, RMB^{+}21, SWL^{+}15, SLTPC^{+}18, SMRS15, SBS17, SBLB17, SFR^{+}13, SKT18, SJS^{+}17, SCV^{+}12, ŠBB^{+}19, SAB^{+}11, SSS^{+}12, STCE12, SFL^{+}14, SF13, ST16, SBT^{+}18, Ste11, SKB^{+}18, SKB^{+}19, SRG17, SVR16, STAN^{+}17, SMPB17, SGT^{+}14, SBL^{+}19, TZW16, TPM15, TCAP11, THHG18, TKF11, THG^{+}16, TRS12, TGDB16, TFB^{+}14, TLM11, VJH17, VOM12, VGE11, VS16, WGC^{+}19, WGC^{+}21, WFJ19, WFNC18, WHS15, WHP15, XDL^{+}19, YAH^{+}13, ZW18, ZQZ^{+}17, ZCX^{+}18, dKMS15, dVLB^{+}15]. fish [vPW16, vPBWA17, GSDD11, JSYS15, JPM11]. fish-based [VS16, ZW18]. fish-borne [GRO^{+}18]. fish-removal [JAJ^{+}16]. fish-schools [RCBPK11]. fish-stock [MHK^{+}17]. fished [KSKGR15, LCBD12, MM13, QWC10, SAH15]. fisher [AFA11, BMC10, CHS^{+}17, DAJMB16, GDB19, HKEJ13, HGD12, LG13, MMDH18, MMG10, OLM^{+}19, SKFM17, TD19, ZCW11, BAOP18, DSWB18, QHSI18]. fisher-based [BMC10]. fisherie [GTCC15]. Fisheries [Ano13a, BHT^{+}19, BKS^{+}14, BKS^{+}15, BB17, BB19, CFF^{+}15, Cor17, FWJ^{+}16, GMS12, HNW^{+}18, HN16, HDM13, HBS^{+}16, Kle13, NAR^{+}17, PSS14, SPR^{+}19, SMD^{+}16, SB14, VCS^{+}19, AW17, AMAM^{+}17, AHC^{+}11, AHWY13, AGSF19, ACSAS^{+}17, ASMP^{+}10, ASMD^{+}18, ALUFJ^{+}12, AFR^{+}17, AF13, BSL10, Bai10, BMU14, BMOZ13, BNAE10, BNAE13, BRH^{+}16, BSF^{+}19, BMG16, BGW^{+}16, BBA^{+}19, BAT12, BR15, BSW^{+}12, BMAMG18, BJKH12, BK19, BHS16, BT16, Bro13b, BAM16, BHP^{+}19, BB13, BRB^{+}14, CRPAMN17, CFBS16, CHT18, CNS11, CWM12, CPW^{+}14, CUL^{+}16, CEJI11, CAM^{+}14, CMB11b, CLG^{+}14, CAC15, CAP^{+}10, CHA15, CCC17, CMHP14, CMMVRF^{+}19, CCC^{+}14b, CL17, CGLS12, CGR^{+}15, CGR^{+}16, CBM^{+}19, Cro15, CTS^{+}19, Cur18, DSC19, DJW^{+}18, Dam15, DDK^{+}18, DMOL15, DDH^{+}15a, DDH^{+}15b, dADSRG10, DLCO^{+}19, DUC^{+}19]. fisheries [DF16, EHH^{+}11, EP17, EGK15, EPHF18, EPTG19, ERL15, EMP^{+}10, EL18, ETMK13, FPC^{+}19, FCBS15, FWIB14, FT16, FDN^{+}11, FSA^{+}16, FMMW13, Fra17, FB19, FSPWG08, GCDE^{+}15, GVMSCJ12, GBF^{+}11a, GAR19, GLMK12, GD17, GS12, GBB^{+}12, GWVH^{+}16, GVG18, GMD^{+}13, GQ16, GLMP18, GAP10, GA19, GRS^{+}15, GDB^{+}16, GJS^{+}13, GBK^{+}13, GBS19, HH16, HHD^{+}11, HBC^{+}12, HPRB16, HCZ^{+}16, HFPL16, HF19, HL16, HVD^{+}18, HGD12, Hil11, HBB^{+}12, HRC16, HFO^{+}18, HLP15, HGG13, HMD^{+}19, IEL17, IGS^{+}12, ICSR16, JSYS15, JRG^{+}18, JNN14, JNE19, JZL13, KZ11, KRM^{+}18, Kos15, KRC14, Kri18, KMdS^{+}18, KK17, LRM17, LE14, LMPM11, LHLR17, LBZE14, LG13, LTV^{+}13, LTHRU17, LT18, LKR11, LCB17, Lor16, LSJ17, LF13, MZC^{+}18, MOS^{+}18, MSRM^{+}14, MSGV11, MFHK10, MAM^{+}13, MBS14, MMdlP16, MH11, MP13]. fisheries [MT19, MNAA^{+}14, MRS15, MGR14, MGB10, MAH16, ML11, MBvDN14, MBB13, MDR^{+}19, MNCPMG15, MUO^{+}17, MML^{+}16, MLV13, Mun12a, Mun12b, Mun18, Nak17a, Nak17b, NCP15, NML12, NBF^{+}19, Nij15, NLPS^{+}13, NBG^{+}17, NC13, OLM^{+}19, PBLT14, PTF^{+}18, PPA^{+}15, PKHD16, PBG^{+}13, PKRC13, PKRC15, PBM^{+}19, PB11, PTSC^{+}16, PJML15, PAH^{+}13, PS13, PLGL18, PGZ^{+}15, PFVGGM16, PHG^{+}18, PHG^{+}19, PHB^{+}19b, PPdR^{+}10, PG14, PALC19, PML^{+}12, PHFW14, QEH^{+}11a, RB12, RMT^{+}18, RC17, RLC^{+}17, RLC^{+}15, RDM^{+}15, RFC^{+}17, RDP^{+}17, RvAGB18, RAG19, RCBS19, SRT13, SL11a, SLP^{+}15, Sal18, SA16, SOMO17, Sam11, SSR18, SMCR13, SHM^{+}16, SMK^{+}13, SCC^{+}15, SNK19, SPMS17, SVPC16, SWG18, SPM^{+}15a, SZ11, Seu21, SJS^{+}17, SEG^{+}15, SZ10, Son18, SB15, SHJ19, SGC11, SUR16, SRB^{+}18, SM17, SHGP11a, SH13b, SKR^{+}12, STAN^{+}17]. fisheries [TM13, TTS11, TEFAR12, THF12, TMC15, TKF11, THKB12, TFB^{+}12, TÜM^{+}16, TRH16, UEOA19, UP16, UARC11, UGGR^{+}17, VF13, VRMF10, VBdB^{+}18, VVL18, VAB^{+}14, WDL17, WMT11, WCM^{+}17, WGDB13, WKTD13, WDS^{+}11, WHK13, WDM15, XLCC19, XYL^{+}19, YCQ11, Yee10, YvZG^{+}17, ZRH^{+}11, ZCZ^{+}13, ZCR16, ZCD^{+}12, Zis12, ZG14, ZLT^{+}16, Cor17, Hol11b, Ano10b, Ano10c, Ano10d, Ano11b, Ano11c, Ano11d, Ano21a]. Fisheries-related [VCS^{+}19]. fisheries-seabird [STAN^{+}17]. fishermen [ELRH17, GGADSRVL15, HGD12, MSD^{+}16, PFVGGM16, PBLT14]. Fishers [BBOL16, TZW16, CPMNS10, CJP13, EvHRF10, HHHWJ18, INSN14, KRTG16, MVML18, MFIG16, NvZH^{+}17, PTG^{+}17, vPWW16]. fisherwomen [HDM13]. Fishery [BGBCM19, CY12, GSP17a, GMS^{+}17, HSH15, LW19b, MHB^{+}19, PDH^{+}16, RM19, Sam14, VSTS10, AFOB^{+}19, AJP18, APD10, ARVACPMN16, AVU^{+}10, AOC11, AK11, AS10, APHL13, ADH^{+}13, AFAJ12, ASR^{+}18, ANEA19, BCKG13, BTBP18, BBM^{+}10, BBF^{+}18, BOCÁ^{+}18, BRH^{+}16, BRC^{+}10, BOBP^{+}17, BHL^{+}16, BdGM^{+}11, BH13, Ben18, BHCJ12, BCC15, BSSH13, BC18, BHC^{+}16, BK10, BHM^{+}18b, BHM^{+}18a, BBB^{+}16, BSN17, Bri10, BLHG17, BFW13, BBK12, BMC10, BRW^{+}13, BBV17, CDR16, CMMF13, CT16, CLWB15, CNJ11, CSMEO^{+}11, CAP^{+}10, CAC19, CHL11, CW17, CEB^{+}19, CHS^{+}17, Cla14, CEW14, CDMC18, CKVV16, CIS13, CSFCA15, CBC^{+}10, CRH^{+}13, CGC13, CS18, CHW^{+}18, CKC13, CCT^{+}14, CSMJB11, CEAP18, CCC14c, CB11, DRGGVL11, DJW^{+}18, DV11, DSJ13, DDB^{+}14, DSWB18, DBK^{+}10, EPCB17, EB15, EHBK18, Eng18, EAM^{+}14, FBKE16, FPM17, FE10]. fishery [FCCSA15, FWL18, FLW^{+}19, FPSS15, FPSDB17, FHK17, FHMK11, FA19, FFG^{+}16, GHPJ15, GCL19, GSP17b, GS18, GWB^{+}17, GFJ19, GDHA13, GBD^{+}15, GH15b, Gra16, GK18, Gri12, GSL14, GPS15, GSH^{+}15, GQ16, HSD^{+}17, HKRH12, HB15a, HB15b, HNW^{+}18, HBT16a, HJLW11, HER13, HTB15, HWS10, HSS^{+}16, HCF^{+}17, HB11, HFBS15, HAB^{+}10, HSG15, Hen11, HSRT17, HSLB19, HNS12, Hoc12, Hol11a, HD12, HAB^{+}17b, HCR14, IdRH10, INY11, JHL^{+}19, JFT^{+}15, JWC^{+}13, JHCT11, JSG^{+}15, JECR^{+}18, KPM^{+}19, KKP11, KMW^{+}11, KDM15, KBCK16, KZ11, KL16, KRB^{+}15, KHV^{+}12, KVK^{+}17, KGB11, KPL^{+}07, KPL^{+}09, LOdH16, LBHP16, LSS^{+}16a, LHS^{+}18b, LHS^{+}18c, LMPM12a, LPM^{+}13, LPC^{+}14, LHL^{+}15, LA13, LYWZ15, LvVtB^{+}15, LCST10, LPHH11, LSH13, LW12, LW13, LW16, LFS^{+}19, LHF16, LMLBC13, LL14b, LT16, LJJ^{+}13, MMJ^{+}12, MTH10]. fishery [MHFH17, MTC^{+}17, MCI^{+}13, MMGEMGS10, MTV17, MSCP17, MD10, MCVS14, MCP^{+}16, May10, MMNCR^{+}13, MMK10, MLF^{+}10, MMF^{+}16, MGR13, MMR^{+}18, MQM^{+}15, MW13, MBF^{+}19b, MCH11, MCMK18, MGC^{+}11, Mur11, NGMA11, NL11, OKAKO16, OHSG19, OGSNH^{+}17, PKH^{+}11, PGS10, PHM^{+}16a, PTG^{+}17, PF15b, PICK17, PVMG19, PAH^{+}13, PBC^{+}16, PQ10, PBR^{+}16, Pit14, PTM^{+}18, PHRH16a, PHRH16b, PBB^{+}15b, PR18, Pre17, POV19, Pul17, PS19, PHM16b, Pun18, QMGP10, QCZC^{+}19, QMAM13, RADR11, RLM10, RRO^{+}19, RHOCB^{+}18, RRPG18, RCE^{+}19, RUMBA15, Rob15, RSG^{+}11, RFS16, RBR^{+}15, RRW^{+}13, RMB^{+}21, RHW^{+}16, RRS^{+}19, SL10, SRSNC17, SGSFL19, SKFL15, SMRS15, Sau14, SKFM17, SK13, SHA^{+}15, SMO^{+}18a, SPB^{+}19, SBS^{+}11, SBD^{+}12, SGT^{+}18, SHG^{+}15, SBH^{+}16, SJR^{+}11, Sta16, SB16, SRB^{+}18, SABG12, SAF^{+}15, SBC^{+}18, SLCL15, TLLZ19, TSNRU21]. fishery [TMM19, TFB^{+}12, UPS16, VB15, VZASM^{+}15, VFMS16, VLSP19, VPSH^{+}17, WB15, WXZ^{+}18, WBI^{+}16, WHS^{+}18, WHC^{+}19, WLGS19, WWR18, WDBH16, WW11b, WLBB15, WKA13, WKT^{+}11, WJC17, XLCMMV19, YVJ12, YSS^{+}11, YGA^{+}12, YJSR12, ZCC11, ZXH^{+}12, ZD13, ZSF11, ZZA^{+}14, Zie12, ZGT12, ZC17, vPWW16]. fishery-dependent [Gra16, KGB11]. Fishery-independent [LW19b, ASR^{+}18, HWS10, KMW^{+}11, LJJ^{+}13, RMB^{+}21, RHW^{+}16, WXZ^{+}18, YJSR12]. Fishery-induced [HSH15, CCC14c]. fishery-scale [BHCJ12]. Fishery-specific [GMS^{+}17]. fishes [Ata11, BGF^{+}15, BHC10, BCPP17, CVS11, EMP^{+}10, FCCSA15, FHB^{+}19, Gan13, GBR^{+}18, GMHPV^{+}10, HNM^{+}12, HMA^{+}16, JvDB^{+}17, LCZ^{+}19, LC19, MMC13, MMK10, PBML^{+}10, PWB^{+}16, PUSI10, POM^{+}18, RMT^{+}18, SWD^{+}13, SBJ^{+}19, TMC15, TMV16, ZBB17]. Fishing [AHC^{+}11, CJP13, CSP^{+}18a, DDK^{+}18, FCIB17, FP10, GSDD11, Ken14, MGLT16, PHK13, SMGH17, WM13, Zie12, AMGH15, AMHR16, AWGP11, BA12b, BHI14, BNU^{+}10, BNAE10, BHG^{+}15, BRH^{+}16, BDHA11, BH13, BR18, BC18, BGP^{+}16, BBM17, BKS18, BM18b, BBHF10, BSKT13, BdSFB^{+}18, CZW14, CNJ11, CEJI11, CAAB^{+}18, CADG^{+}18, CMB^{+}12, COS^{+}19, CCM14, CAR19, CGC13, CEAP18, CB17, DRGGVL11, DRGA12, DTG13, DBG19, DHC19, DBH^{+}10, DHFSS18, DSB^{+}10, DAM10, DFAN12, EP17, ERG^{+}11, EHB18, ELRH17, EMOLS^{+}15, FBGB16, FTM^{+}15, GSP17a, GOD19, GDB19, GHS^{+}19, GBPC10, GAB^{+}15, GBR17, GPT^{+}19, HHM^{+}13, Han18a, HPS16, HF19, HAB^{+}17a, HAB^{+}18a, HCK15, HSS^{+}11, HVJP^{+}11a, HVJP^{+}11b, IWL^{+}11, JIM^{+}15, JRG^{+}18, JF13, JATBC19, JSG^{+}15, KS17, KDM15, KK12, KSL^{+}16, KDPT18, KCO14, KSS^{+}16, KTNA15, KKA10, LOdH16]. fishing [LMHR19, LKL^{+}12, LSŠ^{+}16b, LBH15, LLP18, LHLR17, LCM^{+}18, LSC^{+}18, LvVtB^{+}15, LSH13, LMBF13, LMLBC13, LF13, LČ18, LT16, Lyn14, MZC^{+}18, MP18, MKK^{+}16, MAP14, MCEP14, MD16, MGP^{+}18, MKH16, MSV^{+}19, MSR14, MM18, Mur15, NMMBCV^{+}10, NS16, NMLZ17, NST^{+}17, OMI^{+}14, OH12, ORMG11, PD18, PBG^{+}13, PAX^{+}17, PGZ^{+}15, PHK14, PBB11, Pla18, PAM^{+}18, PBK^{+}19, POV19, PHT15, PNAL16, QHSI18, RHC^{+}12, RBT14, RRSB13, RSG^{+}11, RGSAMV19, RFS16, RM19, RR12, RPC11, RPP^{+}11, RCF^{+}16, RMP^{+}18, Sam11, SD10, SCV^{+}12, SJA17, SB12, SMP18, SHR^{+}18a, ST16, SFBV17, SHGP11b, SYJ^{+}11, SCG^{+}12, TWM11, TBC^{+}19, TZW16, THF12, TBF^{+}16, TCS13, TIGCDL14, VKNK12, VH12, VRMF10, WM11, WSMS14, WRR^{+}16, WHS^{+}18, WMM^{+}15, WMB12, WMB16, WTS^{+}17, WDS^{+}11, WUGG16, YVJ12, ZH10, ZCX^{+}18, ZSF11, ZHDS16, vPB18]. fishing-related [CADG^{+}18, KKA10]. Fishman [QDG17]. fishmeal [CBS17]. fitness [AD16]. fitting [MM14, PLM16]. Five [MBJ^{+}16, ERG^{+}11, HSH16, SB17]. Five-year [MBJ^{+}16]. fix [Mac13, WM17]. Fixed [HCR14, VB15]. fjord [KT18b]. fjords [BLPF^{+}15]. FL [BBV17]. flatfish [AVU^{+}10, BRH^{+}16, BB14, BBH^{+}14, LW16, MWR^{+}13a, PF15a, SHM^{+}16, SPV19, SMPB17, WP17, vMWvOKvB14]. flavescens [HBT^{+}16b, MBS14]. fleet [AAH^{+}16, BVN^{+}10, CPP^{+}10, CMPP11, CMP^{+}12, COS^{+}19, CSMO14, DL11, GPT^{+}16, dMGPR^{+}18, GSH^{+}15, HD12, KDPT18, NCP15, SMNdC^{+}19, SHGP11b, TTF^{+}13, VRMF11, VRMF12, Wal13, Zie12]. fleet-based [BVN^{+}10]. fleets [CFH14, CFH15, ERG^{+}11, HL10, Hua15, JF13, WSMS14]. Flemish [NPGT17, NPRGTSR18, PRMKASR13, TCBR17]. flesh [OTA^{+}13, SKDO^{+}17]. flesus [EÖFPH17, NSM^{+}18, NNP17]. flexibility [YRH15]. flexible [LW13, SBH^{+}16]. Floated [JLF^{+}17]. floating [BKLB14, OKPL11]. flood [ICSR16]. floodplain [ICSR16, JWZ^{+}19]. floodplains [CMB11b]. Florida [ASB^{+}13, Ber13, HHM^{+}13, BTB^{+}19, BBV17, CSMJB11, CCGD19, FMA^{+}18, GCL19, KGS18, MMF^{+}17, MMBD13, SWD^{+}13]. floridana [HFCPSEM15]. flounder [CN19, EÖFPH17, FSH^{+}13, FSBS11, HF14, HFL14, HWS^{+}15, HHSS17, HCR14, KRMW15, MW11, Mau12, MWA^{+}15, NSM^{+}18, NNP17, PMB^{+}11, PT19, PBD21, TMM19, YGA^{+}12]. flow [BMdS^{+}15, BGD11, DBB19, KGR^{+}15, RMGK18]. flow-through [KGR^{+}15]. flowing [BCP^{+}12]. fluctuations [BKM13, CAGTS^{+}10, KvZ12]. fluid [TFH^{+}15]. flume [NWO^{+}15]. fluorescein [CC11]. fluviatilis [HL16, JČK^{+}12, PGVV16, SAR^{+}15, SJA17, VKNK12]. Fluvio [ALR^{+}17]. Fluvio-marine [ALR^{+}17]. flying [FTJ^{+}16, KYO^{+}12, PAX^{+}17, YCYC16, YC18]. focus [RDP^{+}17, SAF^{+}15]. follicles [Gan12, KMQ^{+}10, WTK10]. following [BSKT13, CMB^{+}12, COS^{+}19, GSTL15, JCW12, LLM^{+}19, RLFJ11, SWA19, SBC^{+}18, WHS^{+}18]. Food [TV13, TV14, FP10, RFS16, SSB^{+}18, TWSA16]. footprint [VRMF13]. footrope [CPS^{+}12, HLJ15, SDJ^{+}15]. for-hire [AW17]. forage [CAM^{+}19, HAB^{+}17a, HAB^{+}18a, MRD19, PBB^{+}18, PE14, STAN^{+}17, SBL^{+}19]. Foraging [RvAGB18, SLB^{+}12]. forced [Ing11]. forces [TXH^{+}19]. forcing [APD10]. forecast [SCF^{+}17a]. Forecasting [JP16, KJKK15, LKZ^{+}17, SGRS^{+}18, EHH^{+}15, HBB^{+}19, Jaw11]. forecasts [HSC^{+}12, JCT^{+}16]. Forensic [AHH^{+}12, OPL17, CGdGT^{+}17]. forensics [PGVV16]. forest [OTK^{+}16]. Fork [RSQ18, SGA17]. forkbeard [MSFŠ^{+}11]. form [LRY^{+}15]. formal [DDH^{+}15b]. formalin [ZYSZ18]. formation [BMB^{+}18, MCM12b]. former [LHS^{+}16]. forms [FPM17, HH14]. formulation [GBS19]. formulations [AAS19, GKOC^{+}16, NB19]. Forsterygion [TRC19]. forum [SMGH17]. forward [CHW^{+}10, KOF11]. four [BAB^{+}12, BK15, Fit12, Jaw11, JECR^{+}18, JAJ^{+}16, MFHK15, RMT^{+}18, RH13, SAH15, SBS17, SRG17, ZLT^{+}16]. four-eyed [JECR^{+}18]. Fourier [KBG^{+}14]. Foveaux [FDMH16]. FP7 [LSC^{+}18]. fraction [Kur12, UASM12]. Fractional [GAM10]. Frailty [COS15]. frame [AVU^{+}10, BMAMG18, CSMJB11, SGT^{+}14]. Framework [RBR^{+}15, APCF17, BTS^{+}19, FBGB16, MAM^{+}13, MW13, MM18, PLSM^{+}16, RUMBA15, SHT^{+}13, SDE^{+}18, TBF^{+}16, WDS^{+}11]. frameworks [HCZ^{+}16]. Francaise [GBF^{+}11a]. franciscanus [CUL^{+}16]. Fraser [ZD13]. free [BPS^{+}18, GSDD11, LKL^{+}12, TXZ^{+}17]. free-swimming [TXZ^{+}17]. freeze [CSKB13]. freezing [WMvdL10]. French [DTG13, FDN^{+}11, VKM^{+}17]. frequencies [MBRDM18]. frequency [BKS18, CDL21, CRH^{+}13, CMJ^{+}19, FHS^{+}16, Kur12, LCMTM10, LCMdSM13, MRLD18, Peñ18, PJM14, RTŘ^{+}12, RG18, SLH^{+}14, SBA10, WMPR17]. Fresh [vPKF^{+}19, GHE^{+}11, VRMF11, VRMF12]. Freshwater [EL18, AMV14, ALR^{+}17, BdGD^{+}16, CC11, DBB19, KAGO^{+}10, KHQH15, LCBD12, PJR17, PGVV16, RFC^{+}17, ŘJP^{+}12, TJJ14, VBR^{+}18b]. freshwater-flow [DBB19]. friendly [Pop19]. Froese [Sve14, Sve15]. Front [Ano17k, Ano17l, Ano17m, Ano17h, Ano17i, Ano17j, Ano18k, Ano18l, BHS16]. frontal [SSB^{+}18]. frontalis [AOC11]. frozen [CKC16]. fuegensis [CDM^{+}19, MCN16]. Fuel [MAY12, BRH^{+}16, SCG^{+}12]. Fuelling [HBC^{+}12]. Full [Ano14c, Ano14d, Ano14e, Ano14f, Ano14g, Ano14h, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano15k, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano16i, Ano16j, Ano16k, Ano16l, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano17g, NWO^{+}15]. full-scale [NWO^{+}15]. fullness [MFRM19]. fully [MUO^{+}17, PBB11]. function [Ogl17, PLM16, QQG14, SZ11, WMdS14]. functional [TB10]. functions [BKS18, BTS^{+}19, Mau11, MCP^{+}17b, MVMCC16]. furnieri [CMM^{+}14, dSSdAdAA17, dVLB^{+}15]. Further [MBF^{+}19b, HB11, SLL^{+}16].

Future [HBC^{+}17, Hil11, BHP^{+}19, CMSG18, DHP^{+}13, FBGB16, GBF^{+}11a, JHCT11, MZC^{+}18, Mar18, PSS15, PHM16b, SWH17]. fuzzy [RLC^{+}15, RM19]. fyke [LCBD12].

Gadget [BCTS11, EWB^{+}18]. Gadidae [MSFŠ^{+}11]. Gadids [AL13a]. Gadiformes [FO11]. gadoid [MWR13b, SPM^{+}19]. Gadus [AJOE17, BHL12, BKT10, BK16, BJP^{+}13, BLHG17, BLHO18, BKLB14, CSKB13, DRE^{+}17, EPCB17, GBM^{+}18, HLS^{+}16, HPNA16, HBD^{+}16, KHM^{+}17, KR12, KHM^{+}10, KHV^{+}12, Kri13, Lam13, LDR^{+}12, LHS^{+}18c, MABA12, MM11, OTA^{+}13, OPE^{+}17, OKPL11, PRMKASR13, RBV^{+}13, RSAB11, SH13a, SHG^{+}15, SBH^{+}16, SDV^{+}16, SKDO^{+}17, TWM^{+}10, TCBR17, TSHB19, WOPB11, WHMS11, WHF^{+}14, WTK10, WAT^{+}13]. Gag [EP12, GTS^{+}17]. gags [Sau14]. gahi [RLSPdR14, WA15]. gain [ZW18]. galapagensis [RPK11]. Galapagos [RPK11, SK13, SCOC16]. Galatheidae [HW11, TEFAR12]. Galeocerdo [HSM^{+}12]. Galeorhinus [BMdS^{+}15, HDW^{+}15, RKH^{+}17]. galeus [BMdS^{+}15, HDW^{+}15, RKH^{+}17]. Galicia [HFFdlT15, AGSF19, PFVGGM16, PHG^{+}18, VRMF10]. Galician [VRMF12, BOCÁ^{+}18, RGG14, VRMF11, VRMF13]. gallina [DSJ13, PSM^{+}18]. GAM [PR18, YJC13]. Game [GLMP18, JFT^{+}15, AL13b, BSL10, CBM^{+}19, CB17, Eng18, Gri12, Han18b]. gammarus [Bow17, SMJB11, WMC^{+}16]. GAMs [BGF^{+}15, LWZ^{+}19]. gap [HPRB16]. gaps [AWGP11, BBM17, BMH17, BM18a, BM18b, BMH18, LPHH11, MM13]. garfish [BKS18]. gas [AWSLS15, RCLS18]. gases [RR12]. Gasterosteus [EE19, RvAGB18]. gastropod [CMSG18]. Gastropoda [BdGN^{+}18, CAMFMF18, FAAM16]. Gaussian [BSN17]. gayi [MWSN^{+}13, QMGP10, MWSN^{+}13, QMGP10]. Gear [NSM^{+}19, AHC^{+}11, BTBP18, BHB^{+}10, BBLD16, BLHG17, CHW^{+}10, CAAB^{+}18, CA10, CAP^{+}10, ERG^{+}11, EAM^{+}14, FL13, GPT^{+}19, HSRT17, KL16, Mur11, Mur15, OHSG19, PTG^{+}17, SWA19, SB16, SAWS18, TXZ^{+}17, VKM^{+}17, WM11, WHP15, ZSF11]. gear-size [ERG^{+}11]. geared [LHL^{+}15]. gears [BSS^{+}13, BSM15c, CY12, FTM^{+}15, KLWG17, NGMA11, OLK^{+}14, SCV^{+}12, SSHM16]. Gender [AJOE17, BK19]. Gender-specific [AJOE17]. gene [BMdS^{+}15, KTW^{+}13, RMGK18, WWLY15]. gene-associated [WWLY15]. genera [PMC^{+}18]. general [APCF17, PM13, TWM11]. generalization [HGD19]. Generalized [CIS13, BHB^{+}10, BB10, GDA14, LJB16, LT18, RUMBA15, RU15, ZH10]. generated [WMT11]. generates [VvHE^{+}11]. generating [CPG^{+}15]. generation [ER18, KK17]. generic [TK16]. Genetic [BHC^{+}16, CMMVRF^{+}19, DJW^{+}18, DJF^{+}15, DUC^{+}19, GRPEMEVV17, GWB^{+}17, HND^{+}17, JECR^{+}18, KWS^{+}15, LAIVA^{+}19, LSH13, MKM^{+}16, PSM^{+}18, PGAS15, PLSM^{+}16, PGVV16, RMGK18, SMO^{+}18b, SSG17, TLCSG^{+}15, WWLY15, dVLB^{+}15, AC17, CMPR^{+}16, CJG^{+}19, DS18, EvHRF10, dSFdSSS17, GTDJdL^{+}13, GRGGCACG11, GGS^{+}19, HBS^{+}16, IdRH10, IÓS^{+}14, JRA10, KKH^{+}17, LZGL17, LCH^{+}18, LRAP19, MDL19, MGP^{+}18, MAS15, MAS17, MLBGV16, PGM^{+}10, PRMKASR13, PPVP17, RGK19, RPH^{+}18, SCC^{+}15, ŠBMG^{+}16, SRMP^{+}18, SLZG18, dJSMROZP^{+}13, TVdLOG^{+}21, VRS13, VvHE^{+}11, WB16, WMC^{+}16, XYL^{+}19, dPHS^{+}12]. genetically [CFF^{+}16, CSKB13, HMvHN13, SO18]. genetics [CMM^{+}11, LSR16, STY^{+}16]. Geneva [CLG^{+}14]. Genidens [ALR^{+}17]. genome [JWM^{+}18, PRR^{+}19]. genomic [SRMP^{+}18]. genus [DBMR17, Fit12, IAJ^{+}11, KTW^{+}13, LQL^{+}15, SAH15, YRP^{+}19]. Genypterus [CAFHG10, HND^{+}17]. geo [BNU^{+}10, SKB^{+}18, SKB^{+}19]. geo-location [BNU^{+}10]. geo-spatial [SKB^{+}18, SKB^{+}19]. Geochemical [LKNL16, MNGPM^{+}12]. geochemistry [PTB^{+}15, TVCT12]. geoduck [ZM15]. geographic [BSD^{+}13, FMF^{+}17, GBM^{+}18, Ibá14, LFCC15, MSN^{+}12, MCM12b]. Geographical [QWMD13, DBG19, PBK^{+}19, TOTI10]. geographically [LJB16]. Geography [JD14]. geolocation [LCZ^{+}19]. Geologic [PSJ^{+}19]. Geometric [KBG^{+}14, SPV19, MREV10, VPC^{+}14]. geometry [Peñ18, TXH^{+}19]. Geophagus [KDdAA^{+}18]. George [BEM13, GCRV10, MDG17]. Georges [AHMS10, HJT13, HHSS17, HCR14, LHSS18]. Georgia [ZDZY19]. georgii [BSD^{+}13]. Geospatial [MAP14]. Geostatistical [HFCPSEM15, AMGH15, JKK^{+}16, PLPN12]. geostatistics [AHMS10, NS16]. GEREM [DBLB16]. German [SMJB11, Foc14]. Germany [HKAA18, Sim15]. Geryonidae [MD16, TEGM^{+}11]. Get [MVMdS^{+}17]. Getting [PD18]. Geumo [KKL^{+}16]. Ghats [RADR11]. Ghost [SIBR^{+}19, BM18b]. Giang [HRC16]. giant [BTT11, BN12, CHA15, CS18, MMAH15, RMGK18]. gibel [FHS^{+}16]. gibelio [FHS^{+}16, MLBTM16]. Gibraltar [SVG^{+}17]. gigas [AT10, BdGN^{+}18, CHA15, HBG^{+}19, KWCT^{+}18, LGZC17, LFCC15, MBNM10, NMMBCV^{+}10, PAX^{+}17, RGGdL13, STY^{+}16, SMBP^{+}12, YC18]. Gila [PKWP15, PKWP16]. gill [DJD12, KVK^{+}17, MMGEMGS10, MMM^{+}10, AENC14]. gill-net [MMGEMGS10]. Gillnet [ZTC^{+}18, ACSAS^{+}17, BZ12, CGR^{+}15, CGR^{+}16, DQMV19, DFAN12, HSD^{+}17, LE14, LYWZ15, LT16, LJJ^{+}13, MHB^{+}19, OTK^{+}16, PMK^{+}10, PMP^{+}11, SKFM17, SMO^{+}18a, SBC^{+}18, WJC17]. gillnets [GBD^{+}15, GHS^{+}19, PMB^{+}11, RJKG12]. gillnetting [AVP^{+}15b, MWR18]. GIS [PBC^{+}16]. GIS-based [PBC^{+}16]. give [FSFP15]. given [PPH^{+}17]. Gizzard [JSB19a, JSB19b, VWHB14]. glacial [PSJ^{+}19, VWHB14]. Gladden [GDHA13]. gladius [LGZC17, SP10a, AMQ^{+}15, CNS11, IB10, LLWC15, LRAP19, WMNC15, XdA19]. Glass [DBLB16, KT18b, LTHRU17, LT18]. glauca [ARMLC19, BB13, CNS11, CAM^{+}14, KWS^{+}15, MQV10, MQCS^{+}14, PBB^{+}15b, RHW^{+}16, TOA12, VdSS^{+}14]. Glaucosoma [CMS14]. GLM [PR18, YJC13]. GLM/GAM [YJC13]. Global [AWGP11, SBL^{+}19, VRS13, ABV16, BMOZ13, CAAB^{+}18, FWIB14, FH15, LPMM14, NEJ^{+}18, TBC^{+}19]. globalized [GD17]. globally [CDF^{+}13]. Glover [BCKG13]. Glyptocephalus [PMB^{+}11]. go [DCSW^{+}12]. goals [MCDP18, WP17]. goby [GBF11b, JBB^{+}18]. goldband [GSP17a]. golden [SZPV16, SZPP17, HBT16a]. goliath [SCLS15]. Gompertz [BB19, HHTT11]. Gonad [VRW11, WMF13]. gonadosomatic [FWGM19, JKS^{+}15]. Good [CMLP19, BSF^{+}19, MS14, PD18, PCG16]. goodei [LBS^{+}18]. goose [VRMF13]. GOV [RKP^{+}12]. governance [Bai10, MS12, PLA16, WHK13]. governmental [CGdGT^{+}17]. grade [FPSS15, May10]. gradient [SBT^{+}18, TWM^{+}10]. Gradients [LMB^{+}11, CDR^{+}11, WW14]. Grand [MSR14]. Granger [RCE^{+}19]. Granger-causality [RCE^{+}19]. Graphical [NvZH^{+}17]. gray [BMCS19, KFBL13, RRS^{+}19]. grayling [PHB19a]. great [HL16, SAR^{+}15, BKM19, DB12, EvHRF10, EHH^{+}15, GMMW17, HSS^{+}16, HSG15, HBT^{+}16b, ODdA^{+}10, RMT^{+}18, RKH^{+}17]. greater [JDNP18, ŠBMG^{+}16, SMP18]. greatest [MLTD17]. Greece [RDP^{+}17]. Greek [CHS^{+}17, MKK^{+}16]. Green [BKLB14, GIK^{+}16, GMMW17, LHS^{+}18a, PTM^{+}18, RLM^{+}17]. Greenland [CRB^{+}18, DTC16, GLA19, HBG14, HSL^{+}13, KHGB14, PFJ19, YLA15]. greenling [MNT^{+}17]. greenstriped [KMH^{+}12]. grenadier [CJWT^{+}19, FO11, HKRH12, KRTG16, LCA^{+}19, SHR18b, TSNRU21, XYL^{+}19]. grey [CLB^{+}10, LS10, SL12]. Grid [HB11, BHPC14, BHCS15, DBK^{+}10, GSH^{+}15, HB12, HB13, HSL^{+}13, JHL^{+}19, LHS^{+}16, LHS^{+}18b, LHS^{+}18c, LW13, LW16, LCHL11, LCL12, QCZC^{+}19, SBH^{+}16, SHGL21, SBS^{+}11, SBD^{+}12]. grids [LHS^{+}16, LOC^{+}11, LCLM15, MHFH17, MBF^{+}19b, OPH15]. ground [BBOL16, BLHG17, BSM15c, CSP^{+}18a, DPO^{+}11, DOM14, RMP^{+}18, SH13a, SPC^{+}13, TV13, TV14]. Groundfish [EB15, BH13, BK15, CL17, EPCB17, FHK18, GJS^{+}13, HSTK16, KJ17, LW13, LWH19, MFHK15, NPGT17, OS11, PHM^{+}16a, Pun11, WP15]. grounds [CNJ11, GIW16, GIW17, HBM11, JIM^{+}15, KBOM14, LKL^{+}12, LCM^{+}18, LSC^{+}18, MKK^{+}16, MAAW^{+}17, MGP^{+}18, MHRD16, MSV^{+}19, NMLZ17, SAK10]. Group [RM19]. grouper [CJG^{+}19, EP12, GTS^{+}17, ODMF17, PS19, RBM^{+}16, SCLS15, WWN^{+}15]. grouper-tilefish [PS19]. groups [SLB^{+}12, SHGP11b]. grown [YA18]. Growth [AB12, CGG21, CRAV13, Fra16b, JNF^{+}19, LRML^{+}15, MCP^{+}16, MVMCC16, SP10b, SNOD^{+}13, TMC19, UKK^{+}12, ZM15, ALB^{+}18, ALH^{+}17, AENC14, dSMSF15, ALJC19, APAFMN10, AVA18, ARMLC19, AFR^{+}17, ATG11, BA12a, BFM14, BMB^{+}18, BLHS11, BHG^{+}17, BBT^{+}14, BMF^{+}10, BMF^{+}11, CH11, CJCN21, CA10, Cer14, CM12, CDF^{+}13, CRAVC14, DDP15, DSB^{+}15, DTC16, Eid12, EHGG16, EPO10, EMSC15, FE10, FdSM^{+}16, FS19, GCK14, GLA19, GAR10, HSS18, Han18a, HST10, HFBS11, HFPL16, HKB14, HHTT11, HCSM17, HLB^{+}15, HLSD21, HLA19, HSH16, HSB^{+}13, JB15, JNN14, KNSJJ11, KMH^{+}12, KA15, KM10, KEH16, KTT13, KSH19, Kri13, KT18b, KOHF^{+}16, LBV13, LP18, LCA^{+}19, Lor16, MRD19, MK17, MMF^{+}17, MLBTM16, MSFŠ^{+}11, MAA^{+}16, MAA^{+}18, MBF^{+}17, MWA^{+}15, MOA^{+}16, MNGPMdMG10, OAPS13, Ogl17, ODdA^{+}10, PJR17]. growth [PB11, PLM16, PKO10, PRL^{+}18, PHT15, PHM16b, QDG17, RFPG^{+}16, ROCC13, RCHSN^{+}15, RSGR19, SPM^{+}19, SMØ16, SG15, SP10a, SWC14, SFPB17, SPS^{+}19, SSJ^{+}13, Sim15, SMS17, SMNdC^{+}19, SSB^{+}18, SHJ19, SYL^{+}15, SHR18b, TJJ14, TWM^{+}10, UK14, USH^{+}13, VvPM^{+}18, WWN^{+}15, WBM17, WT16, WKT^{+}13, WKTD13, WMPR17, WHNS13, XTP^{+}16, YLA15, ZMdSC16, ZYSZ18, dZB16, vPW16, Cor17]. Guaceto [GBPC10]. Guadeloupe [GBR17]. guaranteed [Har19]. Guaymas [NMMBCV^{+}10]. Guessing [Ibá15]. Guidance [Tho19a, Tho19b, PWM^{+}16]. guide [PM11]. Guidelines [DDH^{+}15b]. guiding [JTHT14]. Guinea [Cro15]. Gulf [CAGVB14, EMOLS^{+}15, FPSDB17, GLA19, GTS^{+}17, KTL11, LKZ^{+}17, PF15b, AMAM^{+}17, APDI13, APDJSU11, APD10, Ame12, AL13a, ADH^{+}13, Ata11, BA12a, BFM14, BJD12, CSKB13, COS^{+}19, CS18, CRW^{+}19, CR18, DSJ13, DBSA18, EAAA^{+}18, EPHF18, EPTG19, EP12, EMP^{+}10, FSH^{+}13, FHB^{+}19, FMMW13, FPSS15, FSBS11, GRPEMEVV17, GVSB11, GQdÁdMGP^{+}11, GHJ13, HFD^{+}13, HPAA19, HB11, HCSM17, HCM^{+}11, HGD12, HPSG15, Ibá14, JDNP18, JS14, Lam13, LSS^{+}16a, LAIVA^{+}19, LCRW19, MMF^{+}17, MDL19, MFRJ13, MFHK15, MM11, MCH11, MSWW17, MBNM10, OALS11, PJML15, PHRH16a, PHRH16b, Pul17, PS19, RLM^{+}17, RCLS18, RJ16, RU15, RGGdL13, SVV^{+}14, SRSNC17, SH13a, SMP18, SJR^{+}11, SAWS18, dJSMROZP^{+}13, TWSA16, TRS12, TVdLOG^{+}21, VSTS10, WPS19, WTS^{+}17, WGS17, ZCC11]. Gulland [HGD19]. gummy [FWR10]. Günther [CH11, CH14, CH15, LCA^{+}19]. gurnard [MSV^{+}15, PGW17].

H. [MCP^{+}17a]. Haan [AK11, RLM^{+}17]. Habitat [FSH^{+}13, LKD15, LKD18, ODMF17, RGG14, SBT^{+}18, SAWS18, AMV14, BEM13, BLP^{+}18, CRW^{+}19, CMM^{+}14, ESF^{+}19, EHH^{+}15, FSH^{+}15, GAR19, GHUG^{+}15, GHUG^{+}16, HCH^{+}19, LCZ^{+}16, LZX^{+}17, MBJ^{+}16, ML10, RKH^{+}17, SFB11, SWB12, SZ10, SABG12, Tho19a, Tho19b, WFLP11, XGT^{+}17, YCYC16, YGZ^{+}18, YC18, ZTC^{+}18]. Habitat-based [ODMF17]. Habitat-specific [SAWS18]. habitats [DSC19, DH14, DMQ^{+}18, EAAA^{+}18, MCP^{+}17a, MLP^{+}14, PKAL13, PBC^{+}13, RLM^{+}17, RBV^{+}13, SAWS18, WDS^{+}11, WKT^{+}11]. habits [JPM11, RvAGB18, RLSPdR14, RLVS14, UK14, VIFLP17]. haddock [CPOH14, DH11, DRE^{+}17, GIW16, Ing11, Jaw11, KDSAH^{+}18, KHM^{+}10, KHMF11, LDR^{+}12, LHS^{+}18c, MABA12, PBM^{+}18, RRM16, RRRM19, SHG^{+}16, SBH^{+}16, SKK^{+}13, SPM^{+}10]. hagfish [ERL15]. Haizhou [LWZ^{+}19]. Hake [DRGA12, GPSA11, VRMF12, ÁCDM^{+}14, Ame12, BCTS11, BRJ^{+}16, BAL^{+}19, BBT^{+}14, BS10, CPD^{+}19, CDPJ^{+}13, CPS^{+}12, CSP^{+}18b, DPSRM10, DPSR10, EMMM13, FCOM18, GVMSCJ12, HA12, HSRT17, JKS^{+}15, JKK^{+}16, JF13, KMJ^{+}12, KMKK10, KDPMSR10, LBHP16, LRM14, LLM12, LW12, MRLD18, MSGV11, MWSN^{+}13, MBRDM18, MDPCSR10, MIÁ^{+}10, MQGK10, PRP^{+}18, PLSM^{+}16, PPVP17, QMGP10, QCZC^{+}19, RFPE^{+}12, RM10b, RLMM18, TVCT12, TV13, TV14, TAS12, VRMF11, VMB18, WSG16, WMPR17]. hakes [BAL^{+}19, BLPF^{+}15, HGA^{+}12, ODdA^{+}10, RFPG^{+}16]. half [VOM12]. Halibut [DTC16, YLA15, GLA19, HSL^{+}13, KHGB14, LW13, SPMS17, SM14]. halide [MAY12, YMA12]. Haliotis [HHTT11, SKH14, WKT^{+}11]. halves [CBM^{+}19]. hammerhead [SC19]. Hampshire [GMMW17, TC16a, TC16b]. hand [LBB^{+}13a, LL14a, ÖEG^{+}15, WKA13]. hand-line [WKA13]. Handled [EHGG16]. handline [CRAP^{+}11, DSWB18, EPHF18]. handling [BNA^{+}19, CHW^{+}18, HBT16a, MWR18, RBG^{+}19b, SBC^{+}18]. Hanjiang [MZC^{+}18]. hannai [WKT^{+}11]. Hansson [HM16a, HM16b]. haplotypes [SMB^{+}17]. harbor [HMA18, RMP^{+}18]. harbour [LE14]. hard [HN16, LR19]. hardware [LCZ^{+}19]. hardware-accelerated [LCZ^{+}19]. harengas [LCL12]. harengus [LCHL11, OALS11, BHG^{+}17, BTNA13, HBMC12, JS14, MGP^{+}18, NNS^{+}10, SSJ^{+}13, SP15a, STJ11, SLCL15]. Harms [JBMK14]. harpoon [PBC^{+}16, RBR^{+}15]. Harvest [JAB15, MBB13, PKRC15, ANEA19, BDHA11, BCC15, BGB^{+}19, CRPAMN17, DB12, DBH^{+}10, DDH^{+}15a, DDH^{+}15b, FPS11, FWJ^{+}16, FHK18, FPD^{+}16, HBC^{+}17, HPS16, HHHWJ18, HF15b, HLP15, KSL^{+}17, KWF12, KL16, MMF^{+}16, MS12, ML17, OLS15, PCWM16, SCLS15, SMBP^{+}12, Sve15, TWM11, WK10, WHB^{+}16, WP11, WP17, ZCZ^{+}13]. harvested [CDMC18, KRMW15, KVH15, PGLG11]. harvester [GLG^{+}11]. Harvesting [GS12, JDBF15, SKB^{+}18, SKB^{+}19, BDM^{+}14, BGP^{+}16, BB17, BAD19, GKKK15, HGA^{+}12, IIS10, JEJS17, JDLF11, NGMA11, Pla18, TXP^{+}16]. hatcheries [FMF^{+}17, dPHS^{+}12]. hatchery [CBF^{+}14, CM18, CSOV14, HKAA18, KNSJJ11, LCH^{+}18, PEU19, RBG^{+}19a, SO18, SMJB11, SBAK17, WMIDS18]. hatchery-reared [CSOV14, LCH^{+}18, PEU19, RBG^{+}19a, SMJB11]. hatchery-released [CBF^{+}14]. hatchlings [JPSR^{+}14]. haulback [CIS13]. Having [PSE15]. Hawai’i [MOS^{+}18, SF13, APHL13, CB11, HL11, WB15]. Hawaii-based [CB11]. Hawaiian [ASR^{+}18, JP16, ODMF17]. Hawai‘ [BOBP^{+}17]. hazard [BKS18]. HCR [PKRC15]. heads [SKR^{+}12]. headwater [PTBT17]. health [BCPP17, GGS^{+}19, Wal13]. heart [LDR^{+}12]. heavily [Ara15, CPH12, KPF^{+}18, MM13, NMH^{+}11]. heavy [CRAVC14, MB16]. heavy-tailed [CRAVC14, MB16]. height [BBS^{+}17, OS19]. Heikinheimo [SA16, Han16a, Han16b, HM16a, HM16b]. Helgoland [SMJB11]. Helicolenus [RMMSR19, SWB12, SGN^{+}10, SNP^{+}12, SVN^{+}13]. helicopter [KKE^{+}12]. helminth [FCNB16]. help [Ara15, CPC^{+}17, FHK17, HJ12, QHSI18, vPW16]. Helpers [HKAA18]. Hemolymph [BBT^{+}15, LMBF13]. Henry [BPC^{+}12]. Herbst [RM10a]. herding [BBH^{+}14]. here [DCSW^{+}12]. Herklots [KAGO^{+}10]. Hermann [RDP^{+}17]. hermaphrodite [MSWW17]. hermaphrodites [MAA^{+}16]. hermaphroditic [WWN^{+}15]. hermit [LBF^{+}16]. Hermite [HPB10]. herring [BCC15, BHG^{+}17, BSSH13, BTNA13, CKC13, HBMC12, JS15, JS14, KPL^{+}07, KPL^{+}09, LOC^{+}11, LCHL11, LCL12, NNS^{+}10, OALS11, OOTV12, RPH^{+}18, SSJ^{+}13, SP15a, STJ11, SLCL15, SGRS^{+}18, WDNB^{+}19, WBB^{+}16]. Heterocarpus [CCA16a, CCA16b]. Heterodontus [FWR10]. heterogeneity [KL16, MVML18, PJML15, PRL^{+}18, PHT15, Sau14, WWLY15]. heterogeneous [CRPAMN17]. Heterologous [JRA10]. heuristic [CGECSEH10]. Hexagrammos [LWZ^{+}19]. hidden [BMCS19]. Hiding [TM13]. Hierarchical [RUMBA15, GBS19, TJJ14, WJ17, vPCWV15]. hierarchy [LPM^{+}13]. High [CMI^{+}16, GLMK12, KKH^{+}17, MDL19, SSS^{+}12, dJSMROZP^{+}13, WBM17, AHH^{+}12, BCMS^{+}18, BCPP17, CMJ^{+}19, GFJ19, HSC19, KHM^{+}17, KJS12, LJH11, Lyn14, PRP^{+}18, PLMT^{+}11, PPdR^{+}10, RTŘ^{+}12, RGSAMV19, SRMP^{+}18, VvHE^{+}11]. high-energy [BCPP17]. high-frequency [RTŘ^{+}12]. High-resolution [GLMK12, BCMS^{+}18, KHM^{+}17]. high-threshold [GFJ19]. highest [MLTD17]. highly [AFOB^{+}19, ASSL11, Goo16, KBL^{+}15, LSL12, MD13, RAGP18, VBR^{+}18a]. highly-valuable [RAGP18]. Hilaire [ACRM12]. Hilborn [PBB^{+}18]. hindcasting [KKK16]. Hippocampus [LQL^{+}15]. Hippoglossoides [KCM19, DTC16, GLA19, HSL^{+}13, KHGB14]. Hippoglossus [LW13]. hippurus [APDJSU11, CM12, CDF^{+}13, FBHH14, FTN^{+}14, MAH14, MAH16, SMB^{+}17, TVdLOG^{+}21]. hire [AW17]. histological [KMQ^{+}10]. histophysiological [KDdAA^{+}18]. historic [GRPEMEVV17, SSG17]. Historical [BOCÁ^{+}18, GTDJdL^{+}13, PPG^{+}11, SRSNC17, BSN17, CDR^{+}11, DUD^{+}18, Foc14, GPP^{+}16, MMGEH15, PKRC15, SM14]. histories [CP11, CP12, Eid16, RFPG^{+}16]. History [GDHA13, JHCT11, BWN^{+}12, BJZB12, BKS^{+}16, BMR^{+}18, CRS^{+}19, CAFHG10, CADG^{+}18, CLF^{+}17, CCGD19, DB12, FWIB14, FT16, FFG^{+}16, GQdÁdMGP^{+}11, GAP10, HBM11, HF19, HDW^{+}15, HNS12, IÓS^{+}14, JLLL11, KNSSL11, KNSPK13, Koo12, KM15, Lam13, LHL^{+}15, LC19, MRD19, MSS^{+}16, MRS15, MSN^{+}12, OMB18, OCN^{+}11, PGAS15, PQ10, RLM^{+}17, RBM^{+}16, RRM16, SLTPC^{+}18, SNP^{+}12, UP16, ZZA^{+}14]. hitting [CMMF13]. Hoenig [JG12]. Hoki [MAS15, DOO15, MBHGN13, SAK10]. holding [GHM16]. holes [SMD^{+}16]. holistic [RRW^{+}13]. Holothuria [HFCPSEM15, MCP^{+}17a, RGK19]. Homarida [QWMD13]. Homarus [BBT^{+}15, BC18, Bow17, BPG^{+}17, PWA10, PGL^{+}13, SMJB11, TC16a, TC16b, WFHG^{+}17, WMC^{+}16, ZCC11]. Home [ML10, ÖLTA19, TS11, MNW12, NC13]. homing [GSB^{+}19, MNW12]. Honduras [HGD12]. honeycomb [RBM^{+}16]. Hong [TdMWC16]. Hook [CGECSEH10, ASHH12, ASHH15, CMS14, CAP^{+}10, CSFCA15, FCCSA15, HWS10, IEL17, KOD^{+}15, LWCC15, MS18, PKH^{+}11, PGL17, RRS^{+}19, Sau14, SPMS17, SSC^{+}12]. hook-and-line [CMS14, RRS^{+}19, Sau14]. hook-set [LWCC15]. hookeri [HB15a, HB15b, Rob15]. Hooking [PPA^{+}15, GCRV10, LWCC15, MWC15, MC10, SFL^{+}14, VGE11, WLGS19]. Hooks [PBB^{+}15b, BBM19, CB11, GCRV10, HFD^{+}13, MGR13]. hoop [BBM16, BM19b, CRH^{+}13, LSS^{+}17]. hoop- [BBM16]. Hoplostethus [Ken14, KSKGR15, PHK13, PHK14, VRS13]. horizon [SSHM16, SCG^{+}12, HCSM17]. Horizontal [JSB19b, RSEERRSC15, RSEERR^{+}16, BST^{+}17, FKM11, GFK^{+}12, JSB19a, KBT^{+}19, MJL14, RSEERRSC16]. Horizontal-Aspect [JSB19b]. horizontally [MKHK19]. Horn [SAK10]. horned [RGG14]. horse [BKS^{+}16, FCNB16, FMB^{+}19, HBP^{+}18, VRMF10]. host [IIT17, IIT18, LCM^{+}18]. hosts [SMB^{+}17]. hot [KSH19]. Hotspot [RADR11]. hotspots [CEAP18, SMPB17, YCYC16]. hound [CBC^{+}10]. households [TBC^{+}19, WMM^{+}15]. HRM [BCMS^{+}18, FCOM18, Fit12]. hub [HF15a]. hubbsi [ÁCDM^{+}14, BAL^{+}19, BBT^{+}14, BS10, DOM14, EMMM13, LRM14, MRLD18, RM10b, RLMM18, TV13, TV14]. Hudson [MS15a, WRMM18]. hull [SOY14]. hull-mounted [SOY14]. Human [LS19]. humans [LAB10]. Humboldt [AT10, FT16, STY^{+}16]. humoral [Bow17]. humpback [PKWP15, PKWP16]. Hungary [FHS^{+}16]. Huron [MBJ^{+}16, LBZE17, SH13b, WCS^{+}12]. hybrid [ER18, HHB18, MT13]. hybridization [HRAK15]. hybrids [dPHS^{+}12]. hydro [BCP^{+}12, RSE^{+}14, SGB^{+}10]. hydro-turbine [RSE^{+}14, SGB^{+}10]. Hydroacoustic [EHB^{+}21, GYCL19, ZzCbC^{+}15, DJD12, DMK^{+}17, DMQ^{+}18, GFK^{+}12, GBK^{+}18, JSB19a, RCBPK11]. Hydroacoustics [EAAA^{+}18, SWL^{+}15, DQMV19, EE19, RSEERRSC16]. hydrodynamic [KWCT^{+}18]. hydroelectric [CPB^{+}12]. hydrokinetic [SBS17]. hydrolysate [SJS^{+}17]. hydroturbine [BPC^{+}12]. hyperbaric [HSC19]. Hyperoglyphe [FPS11]. Hyporhamphus [BKS18]. Hyporthodus [WWN^{+}15]. hypotheses [CCA16b, RCBS19]. hypothesis [DHF10, LXCC19, MDH15]. Hysterothylacium [PMC^{+}18].

Iberian [CAMFMF18, GQdÁdMGP^{+}11, AAN^{+}12a, CMP^{+}12, CHCS14, FAAM15, FAAM16, GHA13, JF13, MCP^{+}17a]. Ibero [MSRM^{+}14, MSS^{+}16]. Ibero-America [MSRM^{+}14]. Ibero-Atlantic [MSS^{+}16]. ice [LLM^{+}19, VKNK12]. ice-angling [LLM^{+}19]. Iceland [BKT10, CEB^{+}19, JSR10, Kri13]. Icelandic [BAD19, EWB^{+}18, GBEP16, Gun17, GA19, IÓS^{+}14, IEL17, KJ17, SDE^{+}18, WJ17]. ICES [RAGP18]. Ichthyophonus [HHSS17]. ichthyoplankton [Fey18, HCM^{+}11, PH12]. iconic [RW19]. ICPMS [RSGR19]. ictalurid [GMSP19]. Idaho [CM18]. ide [RTSV15]. Identification [CPP^{+}10, CMPP11, CMP^{+}12, FGCG^{+}17, HIH19, PBML^{+}10, ADSV17, BMRC14, CRB^{+}18, DPB^{+}14, GKMGO^{+}17, GWB^{+}17, GGS^{+}19, GMHPV^{+}10, HHB18, HRAK15, JBD^{+}18, LRAP19, LQL^{+}15, MRS15, OPL17, PF15a, PMM^{+}15, RBGE^{+}10, SVV^{+}14, SCC^{+}15, SP10b, SGN^{+}10, SRMP^{+}18, TTMC16, TFB^{+}12, TdMWC16, TC12, UIH^{+}18, VHS^{+}17, XDL^{+}19, PJM14]. identifications [BdSFB^{+}18, Ste18]. identify [CPC^{+}17, LFCC15, LCH^{+}18, MKK^{+}16, NPRGTSR18, PMC^{+}18, PMF^{+}16, SPV19, SAH15]. Identifying [BFHT16, CAM^{+}19, HKEJ13, IO11, KBEB13, RRSB13, SFPB17, SP12, WRG^{+}18, CRCMP^{+}14, KWM^{+}15, RLJ^{+}12]. idus [RTSV15]. if [Ano17k, Ano17l, Ano17m]. IFQ [PS19]. II [BW17, FCCSA15]. Ijsselmeer [GFCK13]. Illegal [PBB11, BdSFB^{+}18, CAC15, MZC^{+}18, RGSAMV19]. Illex [AGB15, CCC17, HdSB^{+}14, LCWC17, PPG^{+}11, RLSPdR14, RLVS14, SP10a, SP10b]. illicia [LBV13]. illumination [FMH13, LW19a, VVL18]. illustrated [MDH15, TGMM13]. Illustrating [SBT^{+}18]. illustration [dSMSF15, MVMSdS19, ZCD^{+}12]. Image [MLT17, Ano17k, Ano17l, Ano17m, FQT17]. images [PAX^{+}17]. imaging [BGL15, CLP^{+}19, RTŘ^{+}12, RH13, SBLB17]. Immediate [MLC^{+}19a, CTSD10, GHL^{+}11, Pul17, SH10]. immersion [WMIDS18]. immigration [LTHRU17, MSV^{+}19]. immobilisation [EDG16]. immune [Bow17]. Impact [BK15, CDPJ^{+}13, HRT^{+}21, MBvDN14, OALS11, PICK17, PHRH16a, ZC17, ASMP^{+}10, BJKH12, BH11, BLHO18, BBH^{+}14, CGM^{+}19, CBS17, HKEJ13, HVJP^{+}11a, HVJP^{+}11b, INSN14, JBMK13, JDLF11, Ken14, MSR14, PHK13, PHK14, Pun11, RLC^{+}15, RBT14, RDP^{+}17, SLB^{+}12, SCG^{+}12, TBC^{+}19, TMC15, ZH10, dPHS^{+}12]. Impacts [AA13, FPS11, JWZ^{+}19, MLTD17, MCB^{+}16, TKF11, AGS^{+}16, ASMD^{+}18, APHL13, BBV17, DBH^{+}10, DSW^{+}17, DBB19, EHGG16, GWB^{+}19, HDM13, HMD^{+}19, KK12, LBH15, LT16, MMG10, NLPS^{+}13, PBLT14, Qui17, RRSB13, RvAGB18, SSR18, VOM12, WDS^{+}11, WA15]. impairment [CTSD10, CHW^{+}18, GHE^{+}11, HJ12, LLM^{+}19, PHB19a]. impairs [WBM17]. implant [LWY^{+}11, PTTC18]. implantation [CBD^{+}14, HSB19, MK17]. implanted [GDW^{+}15, WBD^{+}13]. implanting [JBMK13]. implement [BAFP18, RBR^{+}15]. implementation [BCTS11, FWJ^{+}16, HN16, HCZ^{+}16, HGG13, JASP11, Lyn14, PCG16, ST13]. implemented [AAH^{+}16, SHT^{+}13]. Implementing [CHT18, Cop13, SPR^{+}19, TK16, LNL^{+}15, PSS14, ZCR16]. Implication [UEOA19, WKT^{+}11, DCP^{+}16, FE10]. Implications [ALB^{+}18, APDI13, AHWY13, AGSF19, AAH^{+}16, APD^{+}18, BA12a, BdGM^{+}11, BSD^{+}13, BBHF10, CAMFMF18, CJWT^{+}19, CMMVRF^{+}19, CPB^{+}12, CDB^{+}14, CK18, DSJ13, DJD12, DUC^{+}19, EP17, FPM^{+}18, FWIB14, FAAM16, FMB^{+}19, FMMW13, GCDE^{+}15, HCF^{+}17, IdRH10, JBD^{+}18, JWC^{+}13, MSGV11, MSCP17, MFMM16, MS18, MMD^{+}11, NNS^{+}10, NC13, PQ10, PLA16, PRR^{+}19, RMGK18, SMCR13, SPS^{+}19, SLZG18, SBLM14, SM17, TEFAR12, TJROFS19, TRH16, TLM11, VBR^{+}18b, WMT11, WKT^{+}13, WWR18, WWSM15, XYL^{+}19, ZCD^{+}12, dSMSF15, AWGP11, BGL^{+}13, BKS^{+}14, BKS^{+}15, BSM14, CDPJ^{+}13, CFF^{+}15, DPL^{+}16, DMQ^{+}18, FL11, GAM10, GVDMG12, GM15b, Han18a, HPP18, HBC^{+}12, HDM13, HPS16, LSS^{+}16a, LBZE14, LGZC17, MDPCSR10, Mur19, NNP17, NWS^{+}11, OMB18, PTF^{+}18, PKWP15, PHG^{+}19, PGVV16, Qui17, RB12, RHC^{+}12, SPM15b, SNOD^{+}13, SPM^{+}10, TKF11]. implications [TdMWC16, WLZ^{+}15, WB16, WMC^{+}16, Way13, YZW^{+}19, YSS^{+}11, ZMdSC16]. Importance [OGSNH^{+}17, BBLS17, CDR^{+}11, EHB18, GGS^{+}19, HCWS19, IWL^{+}11, MHK^{+}17, Mur15, PGW17, RSG^{+}11, TLM11, WJC17]. important [DJF^{+}15, FBB^{+}17, HRAK15, KA15, MFHK15, RGAP18, SWD^{+}13, ZzCbC^{+}15, dVLB^{+}15]. imports [LBZE14]. imposed [CMHP14]. impossible [Dam15]. improve [ABSI^{+}18, BSB16, BHL^{+}16, BWR^{+}18, BHS16, BSM15b, BM18a, BSM18, CHW^{+}18, FHMK11, MOS^{+}18, MNAA^{+}14, MMF^{+}16, MKHK19, PJML15, PL18, POV19, WAP^{+}12, WHS15, ZCW11]. Improved [dSMSF15, HPB10, FWL18, GMS12, Hoc12, JNN14, POM^{+}18, RFS16, SUR16, WK10]. Improvements [DOO15, PRBGS^{+}10, PHBU19, KL16, MBF^{+}19b]. improves [BJZB12, KS12]. Improving [ALH^{+}17, BLHG17, ERG^{+}11, HST10, HdED^{+}13, HGD12, KML11, MTH10, MQM^{+}15, MWR^{+}13a, RLC^{+}17, TLLZ19, TWM11, VKM^{+}17, FPSDB17, HSTK16, MCB^{+}18, WHF^{+}14]. imputation [CAMW11]. Imputing [vPCWV15]. in-river [DET^{+}16, MBF19a]. in-trawl [RH13]. Inaccurate [CPCP^{+}12]. Inbreeding [FPM^{+}18]. incentive [KRC14]. Incentives [KRB^{+}15, CGC13, GGADSRVL15, MGC^{+}11]. incentivised [OLK^{+}14]. incidences [LWCC15]. Incidental [BAT12, CSFCA15, Hua15, HB15a, HB15b, PGS10, Rob15]. incidentally [PPA^{+}15]. incirrate [SSH14]. Inclination [Bai10]. including [CPH12, CMLP19, JTP19, SFB11, TLM11, Way13]. inclusion [FSG^{+}19, SJS^{+}17]. income [HKEJ13, OH12]. incomes [MM13]. incongruous [AAS19, NB19]. incorporate [MVMCC16]. Incorporating [AC17, APD10, CAM^{+}14, Chi10, JSB19b, QWC10, ZCR16, KRB^{+}15]. Incorporation [SD13]. increase [Hen11, JLF^{+}17, KW10, PBB^{+}15b, WHS^{+}18]. Increased [CDMC18, JNN14, BSKT13, CGG21, JASP11]. increases [ALJC19, ERCG10]. Increasing [BSM15b, HPNA16, KY19b, BBM17, IWL^{+}11, MM13, MGC^{+}11]. increasingly [GD17, RHOCB^{+}18]. increment [EPO10, FdSM^{+}16, QDG17, Smi14, SHR18b]. increments [ALJC19, BPRB^{+}14]. incubation [PUSI10]. independent [ASR^{+}18, HWS10, KMW^{+}11, KHV^{+}12, KVK^{+}17, LW19b, LJJ^{+}13, PR18, RMB^{+}21, RHW^{+}16, WXZ^{+}18, YJSR12]. indeterminate [CBMLB13, Gan13, KMKK10, SMdUK16]. index [ASN17, CN18, FWGM19, JTP19, KWM^{+}15, LP18, LCZ^{+}16, MRT17, OTK^{+}16, PL18, RJ16, SAK^{+}17a, SZ10, SVR16, TW13, WAP^{+}12, ZTC^{+}18]. index-based [PL18]. indexing [BSS^{+}13, TCAP11]. India [CAR19, GVSB11, GRS^{+}15, SJJ^{+}11]. Indian [CFH15, CJG^{+}19, DJF^{+}15, MMC13, RBM^{+}16, SSH14, WCC^{+}21, ZMG^{+}13, ZMG^{+}14, ARMLC19, CFH14, DSB^{+}15, EMSC15, FH15, FFG^{+}16, GH15a, GMZ^{+}14, HF15a, HER13, HE15, HL10, JPM11, LKL^{+}12, LLWC15, LMSM14, MEM15, SCJ^{+}11, SDC^{+}15, SLH^{+}14, SHC14, SZ10, SSG17, WWN^{+}15, WMNC15, ZCZ^{+}13]. indica [CMS^{+}15, SYL^{+}15, SCL^{+}15, WPCO15]. indicate [MRD19, MLT18, SB15]. indicates [ASSL11]. indicator [BHC10, GAR19, LSŠ^{+}16b, ST16]. Indicators [KSKGR15, AJP18, AMV14, BCKG13, BTBP18, Eid16, HKB^{+}18, KZ11, MMdlP16, NPRGTSR18, Pit14, SC19, SZ11, SDWG18, YCQ11]. Indices [DBSA18, AFR^{+}17, CPG^{+}15, Cam15, CAMW11, CEM^{+}11, GBS19, HNY^{+}19, JKS^{+}15, JAB15, KJ17, LRM14, LSL12, MHK^{+}17, MBF19a, MFRM19, PR18, SLCL15, WSG16, YJC13, dVA18]. indigenous [BRW^{+}13, PTM^{+}18]. indirect [CRAP^{+}11, Eth15]. indirectly [BZ12]. indiscriminate [SFBV17]. Individual [APAFMN10, EPO10, KPL^{+}07, KPL^{+}09, BNAE10, BNAE13, LHS^{+}16, MGM12, PKO10, ZCC11, dZB16]. individual-based [BNAE10, ZCC11]. individuals [BPRB^{+}14, MLTD17, PTC^{+}18, WT16]. Indo [PVKH18, PVKH15, SCT^{+}17]. Indonesia [Nij15, NMLZ17, SPM^{+}15a, YvZG^{+}17]. Indonesian [DSWB18, NvZH^{+}17, WWLY15]. indoor [SMS^{+}19]. induced [BST^{+}17, CJWT^{+}19, CC11, CCC14c, HAB^{+}10, HSH15, MMGEMGS10, OLS15, OTA^{+}13, RHB^{+}17, SG15, SOY14, Way13, YGZ^{+}18, YC18]. induces [SM15]. industrial [HD12]. industries [DBT15]. Industry [ODD16, WBB^{+}16, BMH17, CBS17, FCBS15, GMCL^{+}18, MSD^{+}16, SJS^{+}17]. Industry-based [WBB^{+}16]. industry-developed [BMH17]. inermis [PKO10]. inexpensive [TW10]. Infection [GBM^{+}18, TS10, HPNA16, LCM^{+}18, PBM^{+}18]. infectious [BHT^{+}19]. infer [MFVPC19]. inference [CBO^{+}15, MP17, PLSM^{+}16, SCC^{+}15, WB15]. inferences [APB^{+}16, MLBTM16]. Inferred [DMQ^{+}18, SO18, ALR^{+}17, APD^{+}18, BKS18, BLPF^{+}15, CLB^{+}10, DH14, GTDJdL^{+}13, HDW^{+}15, JKS^{+}15, JKK^{+}16, MFS^{+}18, PGAS15, SSD^{+}19, SLZG18, dJSMROZP^{+}13, WMPR17, WCC^{+}21]. Inferring [AFOB^{+}19, BB10, FPM17, KMW^{+}11, RCBS19]. inflated [WMT11]. inflow [TNKO19]. inflowing [SMØ16]. Influence [CR18, FBKE16, GLA19, KRTG16, KC11, LWCC15, PCWM16, SKT18, SSS^{+}12, WMNC15, CHS^{+}17, EAM^{+}14, FDN^{+}11, HNW^{+}18, HWM^{+}13, IO11, JRG^{+}18, JPM11, JD14, JČK^{+}12, JČK^{+}21, KDdAA^{+}18, KvZ12, MLTD17, MHRD16, MVMSdS19, MWC15, MMAA^{+}18, OKB^{+}16, OS19, PTC^{+}18, PS19, SLH^{+}14, SSB^{+}18, SSC^{+}12, WMdS14]. Influences [SRB19, BNSS19, Ber19, KWCT^{+}18, LSS^{+}17, OCN^{+}11, SSP^{+}15, USH^{+}13, WHP15]. influencing [BHC10, CM18, CSP^{+}18b, JATBC19]. inform [BEM^{+}17, BWR^{+}18, CZW14, CMLP19, ER18, GAR19, HCH^{+}19, KHG^{+}17, PRL^{+}18, SMI17, SRB^{+}18]. information [FAAM15, HSG15, IAJ^{+}11, JF13, KY19a, MP17, NvZH^{+}17, OLM^{+}19, PKRC15, Pre17, Seu21, SLPT17]. informational [PLA16]. informative [TC17]. informed [HBB^{+}19]. Informing [CDR16, TCS13]. ingens [RLSPdR14]. ingesting [BBM19]. inhabiting [BdGM^{+}11]. Inherent [HH16]. initial [PHE16]. initialization [Peñ18]. injectable [CBD^{+}14]. injected [BWK^{+}13]. injuries [VBdB^{+}18]. Injury [GHL^{+}11, CRH^{+}13, CHW^{+}18, LBB^{+}13b, LWCC15, MWC15, WLGS19]. inland [CRH^{+}13, FCLH19, Mun12a, Mun12b]. inlets [HCWS19]. inline [BBM19]. inner [DBG19, JCW12, MMD^{+}11]. innovative [BSB^{+}15, GPT^{+}19, KSS^{+}16]. inodorus [MKM^{+}16]. inserted [BHS16]. inserting [CHW^{+}10]. inshore [ADH^{+}13, BHPC14, GS12, Hen11, MCP14, RHOCB^{+}18, SLCL15]. Insight [NVS^{+}18, HCGS19, MSN^{+}12, VSM^{+}18]. Insights [BJKH12, BB13, CPG^{+}11, HAB^{+}10, HBG^{+}19, JSG^{+}15, MFC^{+}13, PWB^{+}16, Ame12, AT10, BFL^{+}14, CAA^{+}14, GVG18, GHUG^{+}15, JECR^{+}18, KM15, Pun17, RCHSN^{+}15, RPC11]. instructa [LRAP19]. instruments [SHGP11a]. insularis [LLH^{+}14]. integer [CNWM12]. integrate [MP17]. Integrated [HHHWJ18, HSC^{+}12, PBM^{+}19, ALH^{+}17, dSMSF15, CZW14, CPC^{+}17, CMM^{+}11, CMLP19, DLD^{+}16, DSB^{+}15, GBH^{+}19, IOT14, JCT^{+}16, KOHF^{+}16, LPTK19, MP13, MOA^{+}16, PTTC18, PBB11, PDS^{+}17, Pun17, PC19, ROCC13, RIB^{+}10, RCE^{+}19, SZ10, SCF17b, SHJ19, SCOC16, VMB18, WMP^{+}14]. integrated-model [RCE^{+}19]. Integrating [ABSI^{+}18, BNAE13, CAMW11, GBS19, MLF^{+}10, MAS15, MJC^{+}14, MRT17, NMH^{+}11, AVU^{+}10]. Integration [MBG^{+}10, GAM10, HSG15, MCM12b]. integrative [Pop19, TTMC16]. intensity [LBH15]. intensive [MP18, PTBT17]. intensively [KVH15]. intention [RFC^{+}17]. Inter [BA12a, CJCN21, FE10, GGB^{+}18, LGZC17, IÓS^{+}14, KKH^{+}17, Lyn14, MMMPP19, MSV^{+}15, PB11, VB15]. inter- [MMMPP19, PB11]. Inter-annual [LGZC17, Lyn14]. Inter-cohort [BA12a, CJCN21, FE10]. inter-ocean [KKH^{+}17]. inter-population [IÓS^{+}14]. inter-regional [VB15]. inter-specific [MSV^{+}15]. interaction [SMO^{+}18a, ZCR16]. Interactions [ASHH12, ASHH15, Mur11, BOBP^{+}17, LWH19, LJJ^{+}13, Mur15, OS12, RDP^{+}17, RLSPdR14, SWG18, TW13, WM11]. Interactive [GMS12]. Interannual [BKM13, TKF11, HTHK11, ICSR16]. Interannual-decadal [TKF11]. interceptory [BHC^{+}16]. interchange [HBS^{+}16]. interdependencies [HD12]. interdisciplinary [PS13]. interest [SGSFL19]. Interim [HC17]. intermittent [CDB^{+}14, OTH^{+}19]. intermixed [MBB13]. internal [BN12, NSM^{+}18, ZYW^{+}14]. internally [GMSP19]. internally-tagging [GMSP19]. International [Ano10-27, PLGL18, Sal18, MAH16, MDR^{+}19, Nak17b, ZZA^{+}14]. Interplay [KDM15, SMPB17]. interpolating [RPC11]. interpolation [HPB10, YJC13]. interpretation [FBGB16, GAB^{+}15, KY17]. interrupta [AENC14]. interruptus [GRPEMEVV17]. interspawning [UASM12]. interspecific [MS15b]. intertidal [CAMFMF18, DSG^{+}17, GOD19]. Interval [PKRC15]. intervals [UASM12, WMT11]. interventions [BMAMG18, DSWB18, UPS16]. Intra [IÓS^{+}14, MSV^{+}15, CCC^{+}19, GGB^{+}18, KKH^{+}17, MMMPP19, PB11, SP10a]. Intra- [IÓS^{+}14]. intra-annual [CCC^{+}19]. intra-ocean [KKH^{+}17]. intra-population [MMMPP19]. Intra-specific [MSV^{+}15, PB11, SP10a]. intracoelomic [HSB19]. Intraspecific [OMB18]. Intrinsic [BCP^{+}12, MFRJ13, MHRD16]. introduced [HNS12, NC13, WHNS13]. Introducing [EPTG19, LSC^{+}18]. Introduction [TKO^{+}15, BSW^{+}12, GAP10, MGLT16]. introductions [VWHB14]. introgression [dPHS^{+}12]. invasion [CLF^{+}17, HNS12]. invasions [FHS^{+}16]. invasive [AAN^{+}12a, BGL15, CMM^{+}11, HPAA19, JP16, JBB^{+}18, OCN^{+}11, RvAGB18, WHB^{+}16, dCJG^{+}15]. Inverse [Che10, CDR^{+}11, HHTT11]. inversions [PRR^{+}19]. Invertebrate [JDLF11, BDB^{+}11, CMB^{+}12, DDB^{+}14, JDBF15]. invertebrates [RJKG12]. investigate [GPBA13, MNW12, MDPH17, UPS16]. Investigating [CT16, DH11, HSH16, JTP19, PRL^{+}18, SLH^{+}14, TZW16, WMT11, ZYW^{+}14, AVB^{+}16, FMB^{+}19, SF13]. Investigation [TOTI10, BBHF10, MCM12b, PDF^{+}18]. investigations [BGD11]. Investment [NPK^{+}10, AT10, DHH18, FDN^{+}11, PICK17, QHSI18, SSJ^{+}13]. Invisible [LTV^{+}13]. involving [GBR^{+}18, WT16]. Ionian [KDPT18]. IPez [GMHPV^{+}10]. Iran [ZKSF17]. Ireland [LBV13, DHC19, GFO^{+}16, HBM11]. iris [SKH14]. Irish [Bri10, CGR^{+}15, CGR^{+}16, COC14, CB17, DL11, GSØT15, GLMK12, MABA12, PBM^{+}19, WMC^{+}16, WGM^{+}11, dEWB^{+}15]. Irminger [JPO^{+}19]. Irregular [MRD19]. Irreversible [YA18]. iSimangaliso [MCDP18]. Island [KKL^{+}16, KHQH15, PG14, RBM^{+}16, TMC19, VFFG17, ABL21, HDM13, MMSUAGC14, SMJB11, HER13, KMMV10, MPB^{+}17, RDP^{+}17, SPM15b]. Islands [BRC^{+}10, CUL^{+}16, CAP^{+}10, GHUG^{+}16, MMNCR^{+}13, Rob15, TEGM^{+}11, TJROFS19, XDL^{+}19, AFAJ12, ASR^{+}18, AVB^{+}16, BBHF10, CAGTS^{+}10, CEAP18, HB15a, HB15b, JP16, LAB10, MBOAA12, PGS10, PGS^{+}11, RFS16, SPM15b, WA15]. Isles [GFO^{+}16]. Isolation [WPCO15]. Isostichopus [HFCPSEM15]. isotope [CBS11, DDP15, ESF^{+}19, IGSJBS^{+}15, KHM^{+}17, LSWR17, LGZC17, MMMPP19, MCN16, NMH^{+}11, PPLK18, RBG^{+}19a, TOTI10, TCS^{+}15, VIFLP17, VSM^{+}18, VSLC^{+}19, WMIDS18, ZDZY19]. isotopes [FTJ^{+}16, JD14, LRY^{+}15, WW14]. isotopic [HBG^{+}19]. Israel [FB19]. issue [vPKF^{+}19]. issues [PW19, Pun19b, RRSB13, SMGH17]. Istiompax [CMS^{+}15, SYL^{+}15, SCL^{+}15, WPCO15]. istiophorid [GH15b, GM15a]. Istiophoridae [BSD^{+}13]. Istiophorus [LBMC15, TCS^{+}15]. Isurus [BNM11, BB13, RHW^{+}16]. Italy [BRC^{+}10, CMMF13, GBPC10, PF15a, PGZ^{+}15, RC17, RM19]. iterative [HLP15, TDH19]. ITQ [GHPJ15]. IUU [YVJ12].

J [Han16b, HL16, BBM19, HFD^{+}13, PKH^{+}11]. JABBA [WCK18]. jack [ASSL11, KKL^{+}16, LCZ^{+}16, MFS^{+}18, MFVPC19, TJROFS19, VFFG17, VHS^{+}17, WBI^{+}16, ZZA^{+}14, SWH17]. jackass [Way13]. Jackson [FWR10]. Jacobian [PBK^{+}19]. Jamaica [KWCT^{+}18]. Janeiro [BAOP18]. January [Ano10q, Ano10u, Ano11u, Ano12q, Ano13m, Ano14q, Ano15p, Ano17q, Ano18n, Ano19v]. Japan [PPLK18, HL11, Koo12, OSO^{+}18, OAT^{+}16, TOTI10, TKF11]. Japanese [HMGH12, Kai19, KY19b, KKE^{+}12, LYWZ15, LCST10, MSV^{+}19, SKY^{+}17, SKT18, TS10, UK14, XYL^{+}19]. japonica [LT18, SMS^{+}19]. japonicus [DUD^{+}18, FWG11, FWIB14, KT18a, KKL^{+}16, TOTI10, YGZ^{+}18]. Jasus [AOC11, BBB^{+}16, CGG21, EHGG16, FPM17, LPHH11, LSM^{+}19, LGH^{+}21, PTW^{+}13]. Jellyfish [PBLT14, QMAM13, MDL19, RG18]. Jenyns [GRGGCACG11, WMvdL10]. Jersey [SPM15b]. jigging [CRAP^{+}11, KYO^{+}12, MAY12, YMA12]. John [SVV^{+}14]. johni [CN18, FQT17]. johnii [SVV^{+}14]. joint [MGR13]. Jonah [TMC19]. jordani [FO11, HJLW11, HLJ15]. Jorge [Cor17]. José [CAGVB14]. Juan [CUL^{+}16]. July [Ano10x, Ano11x, Ano14m, Ano15o, Ano16s, Ano17r, Ano19o]. jumbo [AT10, HBG^{+}19, LGZC17, NMMBCV^{+}10, PAX^{+}17, STY^{+}16, YC18, QCZC^{+}19, RGGdL13]. June [Ano10s, Ano11s, Ano12p, Ano13q, Ano14o, Ano17p, Ano19p]. jurisdictional [BGW^{+}16]. Just [WCK18, KSH19, MVML18]. justice [SB14]. Justification [SGT^{+}18]. Juvenile [Ata11, BFM14, AAN^{+}12a, ÁCDM^{+}14, BA12a, BKS^{+}16, BDB^{+}11, BPC^{+}12, BWK^{+}13, CGKG^{+}17, CPB^{+}12, CBD^{+}14, DMC^{+}12, FCNB16, FMMW13, FF11, GLA19, GCRV10, HSS18, HDS^{+}18, HGD19, HLB^{+}15, HSLB19, HCH^{+}19, JB15, JNF^{+}19, KW10, Kim15, KCBS13, LBB^{+}13a, LDM^{+}15, LKNL16, LWY^{+}11, LCL12, LMM^{+}17, LJJ^{+}13, MK17, MJL^{+}17, ODdA^{+}10, PB11, RSE^{+}14, RHW^{+}16, SKY^{+}17, SMJB11, SGA17, SMS17, SGB^{+}10, SHR18b, TRC19, TTMC16, UK14, VIFLP17, VSLC^{+}19, VOM12, WAB^{+}16, WGC^{+}19, WGC^{+}21, ZCLB11]. juveniles [MNGPMdMG10, YLA15].

kōura [KHQH15]. Kachchh [GVSB11]. Kajikia [SSP^{+}15]. Kalman [SFB11]. kanagurta [NMLZ17, PGAS15, SSG17]. Kao [LCST10]. Kao-Ping [LCST10]. Karachi [RM10a]. Katsuwonus [GMZ^{+}14, MST14, OTH^{+}19, SF19, SMNdC^{+}19]. Kattegat [BJP^{+}13]. Keepemwet [DDK^{+}18]. Keeping [GBA^{+}18]. kelp [CHT18, JWC^{+}13, SMS^{+}19]. kelts [HRT^{+}21]. Kenya [MMO^{+}13, OKAKO16, OH12, SOMO17]. Kenyan [MMK10]. kerathurus [KTL11]. Kerguelen [AVB^{+}16, NBG^{+}17]. keta [KKE^{+}12]. key [AFOB^{+}19, BK12, KJS12, LSC^{+}18, PE14, JATBC19]. Keys [BBV17, MMBD13, HBP^{+}18, SPM^{+}10, SMD^{+}16]. khudree [BMR^{+}19, RADR11]. kHz [GFCK13, RSEERR^{+}16, SOY14]. kilka [ZKSF17]. killer [BOBP^{+}17]. kinematics [LMM^{+}17]. kinetic [ZXHK15]. King [GBA^{+}18, CH11, CH14, CH15, CCT^{+}14, HNS12, MFRM19, MSN^{+}12, MSWW17, PHE16, SSFGL19, SLT^{+}19, SZPV16, SZPP17, VCS^{+}19, WHNS13, GCRV10, MDG17]. kingcroaker [dSFdSSS17]. Kingdom [WCM^{+}17]. kingklip [HND^{+}17]. kitt [Smi14]. km [GPP^{+}16]. knot [BSM16, BBS^{+}17]. knotless [CEB^{+}19]. Knowledge [GRS^{+}15, BAOP18, DAJMB16, ELRH17, HAB^{+}17b, LG13, MMDH18, PTG^{+}17, PFVGGM16, Qui17, Sta16, WBI^{+}16, ZCW11]. known [BPRB^{+}14, HCHA12, HGD19]. known-age [HCHA12]. kob [MKM^{+}16]. Kong [TdMWC16]. Kootenay [GHL^{+}11]. Korea [KJKK15, KKL^{+}16, Son18]. Korean [KZ11, PMB^{+}11, UKK^{+}12, YSS^{+}11]. Krasnoyarsk [SRG17]. krill [CEAP18, HKK18, KWR15, KWR19, KK15, NS16, WRR^{+}16, ZYSZ18, ZDZY19]. Kuroshio [LL14b]. Kuskokwim [SCF^{+}17a]. Kyphosidae [CPH12].

L [Cor17, GBEP16, Han16b, HM16a, HM16b, GFK^{+}12]. L. [AJOE17, ALUFJ^{+}12, ALBŠB^{+}15, AVB^{+}16, BMR^{+}18, BHM^{+}18a, BTNA13, CFPF^{+}15, CPD^{+}19, CBS11, DPSRM10, DPSR10, EHBK18, EÖFPH17, FMF^{+}17, FBF^{+}18, GDW^{+}15, Gun17, HRT^{+}21, HBMC12, HBM11, HPHR14, HPHF15, HV10, KMKK10, KDPMSR10, KHV^{+}12, LBF^{+}16, LRAP19, MLBTM16, MGP^{+}18, NNS^{+}10, OALS11, OPE^{+}17, PGVV16, QWMD13, RHC^{+}12, RSGR19, SPM15b, SSJ^{+}13, SDV^{+}16, TWM^{+}10, VBR^{+}18b, VBR^{+}18a, WB16, WTK10, dEWB^{+}15]. L3 [GMCL^{+}18]. LA-ICPMS [RSGR19]. lab [YA18]. lab-grown [YA18]. labeling [HH16]. labelling [CPCP^{+}12]. laboratories [GMNPM^{+}13]. Laboratory [SKY^{+}17, CGKG^{+}17, SMdUK16, SGB^{+}10, WMF13]. Labrador [DR13, KCM19, TCBR17, WW11b, WLBB15]. labrax [SAK^{+}17b]. Lacepede [NML12, BNA^{+}19, BPV^{+}17]. Lack [RRM16, IdRH10]. lactate [MK17]. lacustrine [AVP^{+}15b]. Lagocephalus [GGS^{+}19]. Lagoon [CMM^{+}14, HRC16, CLP^{+}19, MM13]. lagoons [CRS^{+}19]. Laguna [OMB18]. Lake [BNSS19, JZL13, KLB17, MBJ^{+}16, RvAGB18, AVP^{+}15a, DB12, FHS^{+}16, HSB19, JWZ^{+}19, JWM^{+}12, JRD^{+}13, KBCK16, KvZ12, LSWR17, LBZE17, SGA^{+}18, SBAK17, SJA17, BJZB12, CLG^{+}14, EGK15, ETMK13, FHS^{+}16, HN16, KEMM12, LSWR17, LBZE17, NST^{+}17, NC13, SIBR^{+}19, SH13b, VH12, VJH17, YAH^{+}13, ZXH^{+}12]. lakes [CMP^{+}14, EHB^{+}21, FCLH19, GKKK15, GRB14, JBB^{+}18, KvZ12, KBČ^{+}16, KHQH15, OTK^{+}16, PMG18, SVR16, VWHB14, VS16, WAP^{+}12, WCS^{+}12, DB12, HSG15, HBT^{+}16b]. Lamarck [RGG14]. Lamna [CRG^{+}18]. lamp [KBCK16]. lamprey [CPB^{+}12, JTHT14, LST^{+}10]. lamps [BKLB14, MAY12, YMA12]. land [Gri12, SMGH17]. land-based [Gri12, SMGH17]. Landa [Cor17]. landed [CAAB^{+}18, PGP^{+}15, PGP^{+}16]. landed- [PGP^{+}15, PGP^{+}16]. landing [AAH^{+}16, AJP18, BMOZ13, FH17, HFFdlT15, MLM15, PCG16]. landings [BNU^{+}10, BHG^{+}15, CGLS12, CDB^{+}14, CPCP^{+}12, GPT^{+}16, GD17, LRM^{+}10, MBvDN14, MCM12a, PTG^{+}17, PCP^{+}18, RDR^{+}18, RGGdL13, RCF^{+}16, RMP^{+}18, SRSNC17, SPV19]. landlocked [PTB^{+}15]. landscape [VvPM^{+}18]. lane [AFR^{+}17, SMO^{+}18b]. Lanka [DAJMB16]. lanthanide [HWS^{+}12, MK13]. lapilli [VF10]. lapse [vPCWV15]. Large [JKS^{+}15, JKK^{+}16, MJC^{+}14, AENC14, ASMP^{+}10, ASSL11, BRJ^{+}16, BdGD^{+}16, BDM^{+}18, BBLD16, BMC10, CEJI11, CBD^{+}14, CN19, CK18, CDR^{+}11, DQMV19, EFWC^{+}19, FWIB14, FHS^{+}16, FP10, FGK19, GKKK15, GBF11b, HSD^{+}17, HLNH15a, HLNH15b, HL10, JHL^{+}19, JWZ^{+}19, KBCK16, KZ11, LNL^{+}15, LR19, LGH^{+}21, LCH^{+}18, MMC13, RBG^{+}19b, SWL^{+}15, Sau14, SPH18, ST16, SBT^{+}18, SZR^{+}11, SGT^{+}14, WCS^{+}12, ZW18, CGLS12]. large-mesh [BRJ^{+}16, JHL^{+}19]. large-river [SPH18]. Large-scale [MJC^{+}14, ASSL11, BMC10, CEJI11, CN19, DQMV19, EFWC^{+}19, HLNH15a, HLNH15b, HL10, LNL^{+}15, LGH^{+}21, LCH^{+}18, Sau14]. largemouth [AAN^{+}12a, GCL19, MWC15, SGA^{+}18, SW19]. larger [MBB13, SLP^{+}15]. largest [DJW^{+}18, JSG^{+}15, LSH13]. Larimichthys [LSJD11, LZGL17, WLZ^{+}15, ZYW^{+}14]. larvae [BKS^{+}16, CSP^{+}18b, CR18, DPO^{+}11, Qui17, SSB^{+}18, TV13, TV14, UJLC^{+}19]. Larval [BBT^{+}14, DOM14, BHG^{+}17, BS10, CN19, CAGVB14, GVSB11, LST^{+}10, MMF^{+}17, MHC^{+}18, OALS11, QQG14, SPS^{+}19, SP15a, SD13, TS10, YA18]. last [ASL^{+}12]. Late [GA19]. Latent [BSN17]. lateral [BSM15b]. laterally [RHB16]. laterally-compressed [RHB16]. Lates [KEMM12, NC13]. laticeps [WWR18]. Latin [PRCGD16]. latitude [ALJC19, BLW^{+}18]. latitudinal [SPM^{+}19, SBT^{+}18, TWM^{+}10]. Latreille [FdMCB^{+}19]. lavaretus [RSV^{+}17, VJH17]. law [BPC^{+}12]. Lawrence [GLA19, Lam13, MM11]. layer [JČK^{+}12]. lead [GVMSCJ12]. leader [ASHH12, ASHH15]. leading [WMF13]. Leads [SKR^{+}12, LR19]. leaping [MLT18]. learn [WHK13]. learned [JP16, LNL^{+}15]. learning [CFB^{+}17, GMS12, GMHPV^{+}10]. lease [vPHG11]. least [HSS18, SMS17]. LED [KBCK16, MAY12, YMA12]. legal [BBM17, FPM17]. legend [Ano17k, Ano17l, Ano17m]. legislated [WHF^{+}14]. Lehtonen [SA16]. lemon [Smi14]. lemuru [PGAS15]. Length [BCKG13, HBP^{+}18, HE15, LNC^{+}15a, LNC^{+}15b, LSŠ^{+}16b, MMK10, PVKH15, PVKH18, SDWG18, TRS12, BK12, CCA16b, CCC^{+}19, CZW14, Cer14, CDL21, Che10, CMHP14, CMJ^{+}19, CV14, DOO15, FdSM^{+}16, FDMH16, HVD^{+}18, HLP15, HSH15, JAB15, JSB19b, KT18a, KWF12, LPTK19, LCMTM10, LSJD11, MOA^{+}16, Mun18, NPK^{+}10, PLM16, Rod19, RSEERRSC15, SLH^{+}14, SGGH10, SPM^{+}10, TW10, WMPR17, WvPLW18, XTP^{+}16, YLA15, ZMdSC16, WvPLW18]. length- [CV14]. Length-based [BCKG13, HE15, SDWG18, CCA16b, CCC^{+}19, FDMH16, HLP15, JAB15]. length-composition [ZMdSC16]. length-conditional [XTP^{+}16]. length-frequency [WMPR17]. Length-SRA [WvPLW18]. length-structured [SGGH10]. lengths [Chi11, SGA17]. lens [NBG^{+}17]. lentic [BdGM^{+}11, CM18]. Leone [Bai10]. leopardus [HHB18]. Lepeophtheirus [BGL^{+}13]. less [vPWW16]. Lessonia [CHT18]. Lessons [Kri18, LNL^{+}15, Nak17a, JP16]. Let [Tho21, Tho19c]. lethal [BBF^{+}18, BMR^{+}19, KSAA13]. lethostigma [FSBS11, MWA^{+}15]. Letter [Cor17, HL16, JBMK14, LRM17]. letting [KOD^{+}15]. leucas [LAIVA^{+}19]. Leuciscus [RTSV15]. Leucoraja [SBC^{+}18]. level [AHH^{+}12, BGW^{+}16, JBMK13, KJS12, KvZ12, MHB^{+}19, MM14, RFC^{+}17, SRG17, TWCA15, ZXH^{+}12]. levels [GBM^{+}18, KFGN10, MBB13, RGSAMV19, WP11]. lewini [NWS^{+}11]. lice [BGL^{+}13, OLS15]. licence [SB16]. licenses [DP18]. LiDAR [JIM^{+}15]. Life [AF13, CDB^{+}14, DBT15, FWIB14, FFG^{+}16, GQdÁdMGP^{+}11, IÓS^{+}14, Koo12, OCN^{+}11, SNP^{+}12, VRMF10, VRMF11, VRMF12, ABSI^{+}18, AFAJ12, BWN^{+}12, BJZB12, BKS^{+}16, BMR^{+}18, BLPF^{+}15, CRS^{+}19, CADG^{+}18, CMM^{+}14, CCGD19, DB12, FT16, GAP10, HBM11, JB15, JLLL11, KNSSL11, KNSPK13, KM15, Lam13, LSWR17, LHL^{+}15, LC19, MRS15, MQV10, MSN^{+}12, OMB18, PQ10, RLM^{+}17, RBM^{+}16, RRM16, SLTPC^{+}18, SP15a, ZZA^{+}14]. life-cycle [SP15a]. Life-history [GQdÁdMGP^{+}11, Koo12, BWN^{+}12, DB12, FT16, KNSPK13, SLTPC^{+}18]. lifetime [FGCG^{+}17]. lift [BBM16]. lift-nets [BBM16]. Lifting [SCLS15, GSH^{+}15, SHG^{+}15]. light [BK15, HLJ15, LHS^{+}18a, OS19, OTH^{+}19]. lighting [RWRT15]. Ligurian [GPBA13]. like [Jab18, PHM16b]. likelihood [AVA18, LPMM14, LPTK19, Mau11, MCP^{+}17b, NS16, PUSI10, SHT^{+}13]. likelihood-based [AVA18]. Likely [AJP18, BRH^{+}16]. Limanda [HHSS17, KRMW15, NKW^{+}19]. limbatus [SCT^{+}17]. limit [JNN14, Mun18, PBD21]. Limitations [BAFP18, KBT^{+}19, CRCMP^{+}14]. limited [ARF^{+}17, CPW^{+}14, CC19, Cop13, FWL18, FPD^{+}16, HSS^{+}16, HF15b, JAB15, KMW^{+}11, KKH^{+}17, LKR11, MMG10, NBS15, PKRC15, RU15, SAF^{+}15, TC17, WFJ19, YCQ11]. limiting [CAP^{+}10, RCF^{+}17]. Limits [KTNA15, CPW^{+}14, CMHP14, Cop13, EWB^{+}18, FJWD17, Gri10, LPHH11, NBS15]. limpet [FPM^{+}18]. Lincoln [ST13]. Line [BBLD16, BNM11, CMS14, HWS10, MGR13, PALC19, RPK11, RRS^{+}19, Sau14, SMNdC^{+}19, SABG12, SAWS18, SKR^{+}12, WUGG16, WKA13]. line-based [HWS10, RPK11]. line-caught [BNM11]. lineages [ŠBMG^{+}16]. Linear [CIS13, BHB^{+}10, BB10, LJB16, MDH15, TWM11, THF12, ZH10]. lineatus [ADSV17]. lines [MGR13, OS19, PGP^{+}15, PGP^{+}16, SMK^{+}13]. ling [EWB^{+}18, HPHF15, PHT15, PHLT16]. lingering [TC17]. link [PW19]. linkages [HL11]. linked [RCF^{+}17]. Linking [KTT13, LF13, NLPS^{+}13, THF12]. links [ASN17]. Linnaeus [APAFMN10, BK16, CMPR^{+}16, CPOH14, DSG^{+}17, FCNB16, KEMM12, MGB10, MQCS^{+}14, ODdA^{+}10, PKAL13, SMB^{+}17, SCJ^{+}11]. Lion [HB15b, HB15a, Rob15]. lionfish [HPAA19]. lions [HB15a, HB15b, MBS14, Rob15, SMO^{+}18a]. lipids [ZMG^{+}14]. Lipsi [RDP^{+}17]. List [Ano10n, Ano10o, GFJ19]. listed [TZW16]. literacy [NvZH^{+}17]. literature [DDH^{+}15a]. Lithodes [MFRM19, SSFGL19, SLT^{+}19]. lithodid [GSTL15, SGSFL19]. Lithuania [VS16]. Litopenaeus [BFM14]. Little [HKAA18]. littoral [GRB14, PBC^{+}13, SVR16]. Live [BNSS19, LS10, ASN17, DRE^{+}17, GHM16, GH14, OTA^{+}13]. Live-capture [LS10]. lived [AM12, HMGH12, JG12, KC18, MRD19]. livelihoods [HDM13]. liver [HPNA16, Rod19]. lividus [BDM^{+}14]. living [HBC^{+}17]. liza [FGCG^{+}17, GCK14]. LLSIM [FGS19]. load [GGM^{+}18]. loads [MBS^{+}15]. Lobatus [BdGN^{+}18, KWCT^{+}18]. Lobster [BW17, LSM^{+}19, AGSF19, ARVACPMN16, AOC11, ATG11, BEM13, BBT^{+}15, BMF^{+}10, BMF^{+}11, BC18, Bow17, BHM^{+}18b, BBB^{+}16, BPG^{+}17, BM18a, CFPF^{+}15, CN18, CJG^{+}15, DBT15, dADSRG10, EHGG16, FBKE16, FPM17, GRPEMEVV17, GHP^{+}15, GHPJ15, GVDMG12, HTB15, HW11, Hol11a, KRB17, LBB^{+}13a, LCRW19, LPHH11, LGH^{+}21, LMLBC13, MFHK10, MHFH17, MMBD13, MLF^{+}10, MMF^{+}16, MMR^{+}18, PWA10, PQ10, PGL^{+}13, PML^{+}12, QWMD13, Qui17, SCOC16, TLCSG^{+}15, WMC^{+}16, WWR18, ZCC11, vPHG11]. lobsters [ASN17, BPPR16, CGG21, GMMW17, KT18a, PTW^{+}13, SG15, SMJB11, TEFAR12, WFHG^{+}17]. Local [HKB^{+}18, NEJ^{+}18, RM19, AHWY13, ALUFJ^{+}12, CHT18, CN19, FHS^{+}16, FMB^{+}19, GKL12, MMDH18, RCF^{+}17, RFC^{+}17, SMPB17, WWLY15, ZCW11]. local-level [RFC^{+}17]. localized [CQ14]. Locating [CPOH14]. location [BHI14, BNU^{+}10, Ibá14, LDM^{+}15, Mur15, RSAB11, SFL^{+}14, SSC^{+}12, WLGS19, WRR^{+}13, ZG14]. locations [GCRV10, PPA^{+}15, Sam11]. loci [CAFHG10, RCF^{+}17, TVdLOG^{+}21, WPCO15]. Locks [SMD^{+}16]. log [MDH15]. log-linear [MDH15]. logbook [BFL^{+}14, FGS19, HBB^{+}12, RCF^{+}16, Sam11]. logbooks [BNU^{+}10]. loggerhead [BGBCM19, Mur15]. loggers [WWG^{+}16]. logic [RLC^{+}15]. logistic [BB17, HHTT11, Zha13]. logits [BK12]. lognormal [DM16]. Loligo [BPMGP10, CRAP^{+}11, Hen11, IIS10, SD10, VSTS10]. Long [Ben18, BN12, CH14, CH15, CGLS12, DUD^{+}18, GD17, Lam13, LSJD11, MSS^{+}16, NNS^{+}10, OSO^{+}18, WB16, AM12, BSM12, BSM15a, EMP^{+}10, EFW^{+}11, FSBS11, GPSA11, HC17, HFFdlT15, JG12, JC18, LBZE17, MSGV11, MNT^{+}17, MD16, SRMP^{+}18, SABG12, WDNB^{+}19, ZCR16]. long-line [SABG12]. long-lived [AM12, JG12]. Long-term [Ben18, BN12, CH15, CGLS12, DUD^{+}18, GD17, Lam13, LSJD11, MSS^{+}16, NNS^{+}10, OSO^{+}18, WB16, EMP^{+}10, EFW^{+}11, FSBS11, GPSA11, HC17, HFFdlT15, JC18, MD16, WDNB^{+}19, ZCR16]. longevity [JG12, WWN^{+}15]. longfin [BHPC14, Hen11, MDH15, WA15]. longitudinal [BDM^{+}18, EPO10, PTBT17]. longline [AHC^{+}11, BGBCM19, BB10, BOBP^{+}17, BAT12, BBLD16, BBHF10, CMMF13, CY12, CNS11, CMPP11, CHL11, CDL21, CIS13, CSFCA15, CB11, FCCSA15, FGS19, FSG^{+}19, FPSS15, FPSDB17, FFG^{+}16, GWWW13, HPP18, HSRT17, HL10, Hua15, HBD^{+}16, IEL17, JRG^{+}18, LJB16, MSRB15, MGR14, MS18, MC10, PKH^{+}11, PPA^{+}15, PGL17, PBR^{+}16, PBB^{+}15b, SMK^{+}13, SJS^{+}17, SZ10, SLXZ15, SRB^{+}18, SKR^{+}12, TTF^{+}13, WB15, WHS^{+}18, WTS^{+}17]. longline- [HBD^{+}16]. longline-caught [MC10]. longlines [ASHH12, ASHH15]. longlining [MWR18, RSAB11, RR12]. longnose [KHG^{+}17, KM15]. longshank [GCRV10]. longtail [Gri10]. Lonhurst [Hol11b]. look [PHM16b]. Looking [SCG^{+}12]. loop [IdRH10]. loopholes [HBC^{+}12]. Lophius [HLB^{+}15, LBV13, OAPS13, SBC^{+}18]. losers [GPSA11]. loss [ASD^{+}18, GVDMG12, GSTL15, HLNH15a, HLNH15b, OOTV12, ÖEG^{+}15, SL11b]. lost [KNJK^{+}10]. lot [WHK13]. lotic [BdGM^{+}11]. Low [CHA15, PGM^{+}10, RRS^{+}19, SCG^{+}12, Cra19, KGR^{+}15, Lyn14, MAY12, MLM15, PWM^{+}16]. low-energy [MAY12]. low-voltage [PWM^{+}16]. Lower [Ara15, DET^{+}16, LCST10, LČ18, WFLP11, CMB11b, DUC^{+}19, ICSR16, PKWP16]. lowland [STCE12]. Loxechinus [BMB^{+}18, MMD^{+}11, RUMBA15]. LPUE [HFFdlT15]. Lucas [OGCBMR15].

Luciobarbus [RSEERRSC15]. lucioperca [GKKK15, Han16b, HPHR14, HL16, HM16a, HM16b, JČK^{+}12, KVH15, LSŠ^{+}16b, SAR^{+}15, VH12]. lucius [Cur18, KKA10, PALC19, SMØ16, SLB^{+}12, SFL^{+}14]. lumpfish [KJ17]. lumpsucker [HBG14]. lumpus [HBG14, KJ17]. lupus [GH14, GBEP16, Gun17, GSB^{+}19]. Luque [Cor17]. lure [KS12, MWC15, PALC19, SFL^{+}14, WLGS19]. lure-shedding [PALC19]. luscus [RDV11, RBV^{+}13]. Lutjanidae [KKH^{+}17, PBB^{+}15a, SVV^{+}14, SMO^{+}18b]. Lutjanus [CTSD10, FGK19, GDHA13, HPSG15, LKZ^{+}17, SVV^{+}14, SMO^{+}18b, TS11, VvHE^{+}11, WGS17].

m [WB16]. M. [JKS^{+}15]. Macaronesia [FPM^{+}18]. Macaronesian [AASF16, CMSG18]. Maccullochella [DBH^{+}10]. macdonaldi [MFRJ13]. machine [CFB^{+}17, GMHPV^{+}10, ÖEG^{+}15, RCBPK11]. machines [RBGE^{+}10]. Mackerel [JB15, SWH17, ASSL11, BKS^{+}16, BDC^{+}17, CLB^{+}10, DTSG^{+}16, EPHF18, FCNB16, FMB^{+}19, GAM10, HBP^{+}18, HNY^{+}19, HV10, JPO^{+}19, JFT^{+}15, KKL^{+}16, LCM^{+}18, LYWZ15, LCZ^{+}16, MFS^{+}18, MFVPC19, NML12, RU15, SSG17, TSS11, TPM15, TJROFS19, VFFG17, VHS^{+}17, VRMF10, WBI^{+}16, YGZ^{+}18, ZZA^{+}14, ZLT^{+}16]. macleayi [MBW12]. macro [CDR^{+}11, GVSB11, ODdA^{+}10]. macro-biota [CDR^{+}11]. macro-tidal [GVSB11]. Macrobrachium [KAGO^{+}10]. macrocephalus [DS18, JRG^{+}18, KHM^{+}17]. macrochir [MSN^{+}12, MSWW17]. Macrodon [CH11, CH14, CH15]. macrofaunal [CKVV16]. Macroparasites [SGN^{+}10]. macropterus [Way13]. macroscopic [MWR^{+}13a]. macrostructure [RFPE^{+}12]. Macrouridae [FO11]. Macrourus [Fit12, LCA^{+}19]. Macruronus [CJWT^{+}19, DOO15, HKRH12, MBHGN13, MSGV11, MAS15, SHR18b, TSNRU21]. mactroides [TXP^{+}16]. maculatus [HHB18]. Madagascar [BMOZ13, HAB^{+}17b, RB12]. made [HWS^{+}15, PKAL13]. Madeira [VFFG17]. Madre [OMB18]. maenas [PTM^{+}18]. Magdalena [MMdlP16]. magellanicus [CJWT^{+}19, MBHGN13, MSGV11, MAS15, MHC^{+}18, Mur11, TSNRU21]. magister [ZD13]. magnets [PBB^{+}15b, RRPG18, RPK11, WWR18]. magnitude [ASD^{+}18, TMV16, YvZG^{+}17]. mahseer [BMR^{+}19, RADR11]. main [AJP18, GPdR11, LAB13, LSH13, ŘJP^{+}12]. Maine [TC16a, Ame12, AL13a, ADH^{+}13, BC18, CSKB13, DBT15, FSH^{+}13, GHJ13, HB11, Hol11a, JS14, JWC^{+}13, LCRW19, MCH11, PHRH16a, PHRH16b, RJ16, SH13a, SBA10, TC16b, WBB^{+}16, ZCC11]. mainland [LBZE14, RMGK18]. maintenance [BBT^{+}15, LMBF13]. major [CN19, LCM^{+}18, LSC^{+}18, PG18, VRMF10]. Majorca [MNCPMG15]. Makah [SPMS17]. Makaira [HLA19, WRP10]. Making [DQMV19, UP16, CJP13, GMS12]. Mako [BNM11, BB13, RHW^{+}16]. Malawi [HN16]. Malaysia [IAJ^{+}11, INY11, INSN14, TTS11]. Maldives [SL11b]. Male [KFBL13, HCF^{+}17, KWM^{+}15, KAGO^{+}10, QWMD13, SSFGL19, SLT^{+}19, SZPP17, TMM19, WFHG^{+}17]. male-directed [HCF^{+}17]. maliger [Hoc12]. Malindi [MMO^{+}13]. Malindi-Ungwana [MMO^{+}13]. Mallotus [DH14, MHRD16, SPS^{+}19]. malma [LRY^{+}15, LRY^{+}15]. Malta [GFK^{+}12]. mammata [MCP^{+}17a]. man [PKAL13]. man-made [PKAL13]. managed [BT16, GHPJ15, VPSH^{+}17, WCM^{+}17]. Management [AVU^{+}10, Eid16, FWJ^{+}16, LSS^{+}16a, Nak17b, NWS^{+}11, PHG^{+}19, SPR^{+}19, Via19, WB16, Way13, ZMdSC16, AW17, ALB^{+}18, AGS^{+}16, dSMSF15, AHWY13, AGSF19, AFA11, APD10, ALBŠB^{+}15, APD^{+}18, AAN^{+}12b, BVN^{+}10, BHG^{+}15, BRH^{+}16, BdGM^{+}11, BMG16, BCC15, BSD^{+}13, BMAMG18, BWR^{+}18, BBB^{+}16, BB17, BB19, Bro13b, BHP^{+}19, BRB^{+}14, CDR16, CAMFMF18, CDPJ^{+}13, CMMVRF^{+}19, CP11, CP12, CV14, CCGD19, Cur18, DJW^{+}18, Dam15, DAJMB16, DSJ13, DUC^{+}19, EP17, Eng18, ETMK13, FE10, FWIB14, FT16, FWL18, FL11, FMMW13, FSPWG08, FH17, FDMH16, FEGMTÁ17, GCDE^{+}15, GPSA11, GCL19, GAR19, GLMK12, GDHA13, GHUG^{+}15, GBPC10, GM15b, GYCL19, GPT^{+}19, HBB^{+}19, HNW^{+}18, Han13, HC17, HDM13, HCZ^{+}16, HGA^{+}12, HSG15, HA12, HGD12, Hil11, HRC16, HF15b, Hor11, HMGH12, HVJP^{+}11b, IdRH10, IWL^{+}11, JNN14]. management [JWC^{+}13, JHCT11, JECR^{+}18, KZ11, KRTG16, Kni12, KEP^{+}19, KM15, KRC14, KRB^{+}15, KHQH15, LG13, LPHH11, LSJ17, LT16, MSGV11, MMC13, MMdlP16, MSCP17, MHB^{+}19, Mau12, MDPCSR10, MGB10, MMG10, MW13, MMD^{+}11, Mun12a, Mun12b, MQGK10, Nak17a, NAR^{+}17, NBF^{+}19, NC13, OMB18, PD18, PTF^{+}18, PSU^{+}16, PKHD16, PBM^{+}19, PJML15, PS13, PFVGGM16, PHBU19, PHB^{+}19b, PBD21, POV19, PRR^{+}19, PTW^{+}13, RB12, RLC^{+}17, RSG^{+}11, SLP^{+}15, Sal18, SOMO17, Sam14, SSR18, SMCR13, SPM15b, SAB^{+}11, SLZG18, SNOD^{+}13, Sta16, SHJ19, SRB^{+}18, SHGP11a, SP12, TEFAR12, TMC15, TKF11, TRH16, UEOA19, UPS16, VF13, VBR^{+}18b, WMdS14, WLZ^{+}15, WHC^{+}19, WMT11, WWR18, WP17, WWSM15, XLCC19, XYL^{+}19, YZW^{+}19, YSS^{+}11, ZD13, ZCR16, ZCD^{+}12, ZCW11, HN16]. management-risk [FEGMTÁ17]. managerial [KKP11]. managers [AF13, KOD^{+}15, RFC^{+}17]. Managing [CAP^{+}10, GMD^{+}13, IGS^{+}12, PKRC15, RRSB13, WBI^{+}16]. mangrove [IAJ^{+}11]. mangroves [RMGK18]. Manifesto [GBF^{+}11a]. Manila [CMM^{+}11, CLF^{+}17, RvRW^{+}14]. manners [CPMNS10]. manta [ACSAS^{+}17]. mantis [LMBF13]. map [PCP^{+}18]. Mapping [BJP^{+}13, ELRH17, BNU^{+}10, JIM^{+}15]. Maps [RCF^{+}16, GDA14]. maraena [LCF15]. marble [CFF^{+}16]. March [Ano10v, Ano11y, Ano12l, Ano13n, Ano14n, Ano16o, Ano17t, Ano18s, Ano19w]. margin [DSC19]. marginal [CRS^{+}19, Smi14]. Marin [Cor17]. Marine [Ano13a, BCKG13, CGLS12, GBPC10, HSS^{+}16, HHCP13, Hol11b, JKS^{+}15, JKK^{+}16, LPC^{+}14, MDPH17, SK13, SB12, SCOC16, AMAM^{+}17, AENC14, ASMD^{+}18, AA13, AHH^{+}12, AFAJ12, ASB^{+}13, ALR^{+}17, BBLD16, BSW^{+}12, BMF^{+}10, BMF^{+}11, BWR^{+}18, BM19b, BDB^{+}11, CAAB^{+}18, CQ14, CMB^{+}12, DBG19, DDP15, EP12, FMA^{+}18, FDN^{+}11, FL11, FSFP15, FP10, GBF^{+}11a, GAR19, HHM^{+}13, HJT13, HHHWJ18, HGG13, HMA^{+}16, IWL^{+}11, JS15, JSR10, JP16, KNSSL11, KMSRW10, LMB^{+}11, LA13, Lyn14, MOS^{+}18, MP18, MP14, MCDP18, ML10, MMBD13, MFKS17, MLV13, NMH^{+}11, OPF15, OKAKO16, PGZ^{+}15, PHG^{+}18, RC17, RAG19, RLSPdR14, RMB^{+}21, SHR^{+}18a, SF13, SB16, Ste11, TMC15, TMV16, TLCSG^{+}15, TÜM^{+}16, UEOA19, VGE11, WMC^{+}16, WGM^{+}11, WMM^{+}15, WUGG16, XDL^{+}19, BGW^{+}16, BPPR16, BFHT16, BAOP18, PBG^{+}13, PRCGD16, PLA16, RRSB13]. Marine [RBR^{+}15, WGDB13]. Marini [DOM14]. marinus [BPK^{+}17, WRG^{+}18]. Marjomäki [Han16b]. mark [GVDMG12, HTB15, HF14, HDW^{+}15, HE15, LWS19a, LWS19b, MAH14, Sau14, SMJB11, SW19, VIP13]. mark-recapture [Eth15, HF14, HE15, Sau14, SMJB11, SW19]. mark-recovery [HTB15, LWS19a, LWS19b]. mark-resight [VIP13]. marked [CLS^{+}11, RIB^{+}10]. marker [CCHD19]. markers [BGA11, CFF^{+}16, FdMCB^{+}19, HHB18, KBG^{+}14, KR12, Kur12, LSR16, LZGL17, MBG^{+}10, STY^{+}16, WWLY15]. market [ASN17, FGRIT^{+}12, GAM10, HKEJ13, HL11, Ibá15, MCM12a, MCM12b, RDR^{+}18, SB15, vPHG11]. market-testing [HKEJ13]. markets [CPCP^{+}12, FGRIT^{+}12, HL11, OPL17, PF15a, PRP^{+}18, SKT18]. Marking [MNGPMdMG10, WMIDS18]. Markov [BMCS19]. marlin [BSD^{+}13, BMS15, CBO^{+}15, CMS^{+}15, FGS19, Goo16, HLA19, LKD15, LKD18, LPMM14, OGCBMR15, SSP^{+}15, SYL^{+}15, SCL^{+}15, WRP10, WPCO15]. marmorata [LT18]. marmoratus [CFF^{+}16, MNW12, ZTC^{+}18]. marron [BdGD^{+}16]. Marta [Cor17]. Maryland [DW19]. masquinongy [LWSC11]. mass [KAiT^{+}13]. Matanuska [ST13]. material [CM12]. materials [ASHH12, ASHH15, SB17]. Maternal [RLMM18, MRLD18, OSO^{+}18, SD13]. maternity [MQCS^{+}14]. mating [PGL^{+}13, WFHG^{+}17]. matrices [SZPV16]. matrix [CNWM12, PBK^{+}19, SKH14]. matter [BRB^{+}14, KHKM15]. matters [CM12, FSPWG08, FWG15, MCJD^{+}17, Sve14]. Maturation [SMBP^{+}12, BV14, DH11, HSH15, KVH15, KTT13, MM14, SM15, TMM19, UKK^{+}12, dCJG^{+}15]. mature [AJOE17, SZPP17]. Maturity [HKB14, APDJSU11, AT10, AMH19, ATG11, BLW^{+}18, BdGN^{+}18, BPG^{+}17, BSKT13, CCA16a, CH14, DMK^{+}19, FMB^{+}19, FWGM19, HST10, HSTK16, HGD19, HW11, JG12, KHGB14, LSŠ^{+}16b, LSJD11, MLM15, MWR^{+}13a, MQCS^{+}14, MJC^{+}14, NPK^{+}10, QWMD13, RM10a, Rod19]. maturity-at-age [NPK^{+}10]. maturity-at-size [DMK^{+}19]. Mauritania [BFL^{+}14, CBS17, LXCC19]. maw [TdMWC16]. mawsoni [OCL^{+}18, YZW^{+}19]. maxima [FF10, HBM11]. Maximising [BSM15c]. maximum [GHPJ15, NS16, PUSI10, RTSV15, SHT^{+}13]. May [Ano11w, Ano12o, Ano14k, Ano15l, Ano16p, Ano17x, Ano18o, Ano19s, GVMSCJ12]. maya [BPRB^{+}14, DHFSS18, JRA10, JECR^{+}18]. McIntyre [Ano10a]. me [Bai10]. meagre [GQdÁdMGP^{+}11, MNGPM^{+}12]. meals [AHH^{+}12]. mean [Chi11, GRB14, Jaw11]. Means [HPS16, CPOH14, ODdA^{+}10, PM11, RSV^{+}17]. measure [HJ12, TGM13]. measured [BKM13, SMS^{+}19]. Measurement [MBS^{+}15, AMHR16, CMJ^{+}19, Cra19, HBP^{+}18, HSTK16]. Measurements [SOY14, GBK^{+}18, KEMM12, PB14, RT14, RSEERRSC16]. measures [AAH^{+}16, HVJP^{+}11b, KWR15, KT18b, PHE16, PS12, WMF13]. Measuring [GRVN10, INY11, MMSUAGC14, dKMS15, TW10]. mechanism [KYO^{+}12]. mechanized [UGGR^{+}17]. mediated [APB^{+}16]. Mediterranean [ALUFJ^{+}12, BA12a, CMPR^{+}16, CPD^{+}19, CAP^{+}10, DV11, GPBA13, KBG^{+}14, LRM^{+}10, MMF^{+}17, MMNCR^{+}13, MGB10, MLBGV16, MEMdU^{+}17, MLV13, TÜM^{+}16, UJLC^{+}19, UGGR^{+}17, VSLC^{+}19, ASL^{+}12, ASB^{+}16, ABD^{+}19, AB11, ABSI^{+}18, ALBŠB^{+}15, BMEGV^{+}19, BCTS11, BGF^{+}15, BRC^{+}10, BAB^{+}12, BHCS15, BHS16, BHM^{+}18b, BHM^{+}18a, CLP^{+}19, CSP^{+}18a, CSP^{+}18b, Cor17, DPB^{+}14, DSC19, Dam15, FMB^{+}19, FPC^{+}13, FL11, GGS^{+}19, GWVH^{+}16, GBD^{+}15, GFM^{+}12, GM15b, HVJP^{+}11a, HVJP^{+}11b, KPM^{+}19, KBCK16, KDPT18, KM15, LMB^{+}11, LRML^{+}15, LPC^{+}14, MPB^{+}17, MBRDM18, MMM^{+}10, MQM^{+}15, OOM15, ÖEG^{+}15, PAH^{+}13, PBC^{+}16, RLM10, RCF^{+}17, RDP^{+}17, RMMSR19, RBR^{+}15, RLVS14, SMB^{+}17, SL10, SL11a, SLP^{+}15, SD10, ŠBMG^{+}16, SAK^{+}17a, SSS^{+}12, SABG12, TVCT12, THA^{+}14, THG^{+}16, VSM^{+}18, VVL18]. medium [GPdR11, YvZG^{+}17]. medium-scale [YvZG^{+}17]. meet [SHGP11a, WP17]. meets [GGM^{+}18]. megafauna [AMAM^{+}17]. Megalaspis [SCJ^{+}11]. Megaleporinus [APD^{+}18]. megalocyathus [ORMG11]. Megalops [AL13b, GBA^{+}18, PGLG11]. megrims [CPCP^{+}12]. Mekong [DUC^{+}19, DS18, GBR^{+}18]. melanaster [RT14]. Melanogrammus [CPOH14, DH11, DRE^{+}17, GIW16, Jaw11, KDSAH^{+}18, KHM^{+}10, KHMF11, LDR^{+}12, LHS^{+}18c, MABA12, PBM^{+}18, RRM16, RRRM19, SHG^{+}16, SBH^{+}16, SKK^{+}13]. melanostomus [GBF11b, JBB^{+}18]. melanura [CMPR^{+}16]. Melicertus [CCT^{+}14, KTL11]. melting [BCMS^{+}18]. membras [LCHL11, LCL12, OALS11]. memory [LFG11]. menhaden [LSS^{+}16a, LWS19a, LW19b]. Menippe [DSW^{+}17, KGS18]. mentella [CRB^{+}18]. mercenaria [KGS18]. merguiensis [DBB19]. Merlangius [GIW17, PBM^{+}18, VKM^{+}17]. merlangus [GIW17, PBM^{+}18, VKM^{+}17]. Merluccius [ÁCDM^{+}14, BRJ^{+}16, BAL^{+}19, BBT^{+}14, BS10, CPD^{+}19, CDPJ^{+}13, CPS^{+}12, CSP^{+}18b, DOM14, DPSRM10, DPSR10, EMMM13, HSRT17, JKS^{+}15, JKK^{+}16, JF13, KMKK10, KDPMSR10, LRM14, LLM12, LW12, MRLD18, MWSN^{+}13, MBRDM18, MCP14, MGB10, MQGK10, ODdA^{+}10, PLSM^{+}16, QMGP10, RFPE^{+}12, RFPG^{+}16, RM10b, RLMM18, TV13, TV14, TAS12, WSG16, WMPR17]. merobenthic [LH15]. mesh [BSB^{+}15, BHL^{+}16, BRJ^{+}16, BHS16, BHM^{+}18b, BMB10, BSM15b, CHW^{+}10, CFPF^{+}15, CEB^{+}19, CK18, DBK^{+}10, DÖA^{+}10, HSD^{+}17, Hen11, HCM^{+}11, HLS^{+}16, Ing11, JHL^{+}19, KG10, KHMF11, LCF15, MMJ^{+}12, MMGEH15, PMB^{+}11, PHRH16a, RJKG12, RHB16, SL10, SL11a, SLP^{+}15, TXH^{+}19, THA^{+}14]. meshes [WHMS11]. meso [OtHH^{+}10]. meso-scale [OtHH^{+}10]. Mesocentrotus [CUL^{+}16]. mesocosm [SBS17]. mesopelagic [SLTPC^{+}18]. mesopotamicus [IdRH10]. mesoscale [ABSI^{+}18]. mesotrophic [SJA17]. mesozooplankton [BKM13]. meta [CHCS14, DBMR17, FHK17, PD14, TDH19, WFNC18]. meta-analysis [DBMR17, TDH19, WFNC18]. meta-analytic [PD14]. meta-population [CHCS14, FHK17]. metabarcoding [GKOC^{+}16]. metabolic [LMM^{+}17]. metal [GWWW13, HWS^{+}12, MAY12, YMA12]. metals [MK13, RPK11]. metapenaeid [PTC^{+}18]. Metapenaeus [MBW12]. metapopulation [CMMVRF^{+}19, PLSM^{+}16, ZM15]. Metazoan [MMM^{+}10]. Method [KMSRW10, RCLS18, SSS^{+}12, BAFP18, BHM^{+}18b, CMSG18, CPP^{+}10, CDL21, DM11, EL18, FMD19, FJWD17, Gan12, GS12, GMCL^{+}18, HHH^{+}12, HGD19, JCW12, JvDB^{+}17, KT18a, KWM^{+}15, KFGN10, KWF12, LG13, LKNL16, LQL^{+}15, Mac13, MIÁ^{+}10, PUSI10, PSE15, PMF^{+}16, PCP^{+}18, QQG14, SSR18, SVN^{+}13, SP12, TCAP11, TK16, VS16, WMIDS18, WWG^{+}16, WHS15]. Methodological [Wri13, GVG18, MKK^{+}16]. methodologies [BWR^{+}18, GKMGO^{+}17, ZBB17]. Methodology [HSRT17, BTT11, HLP15, IGS^{+}12, Peñ18]. methods [ÁCDM^{+}14, AW12, BEM13, BCN^{+}13, BNK14, BSW^{+}12, BBT^{+}15, BBH^{+}12, CRS^{+}19, CFBS16, CPG^{+}15, CWM12, CPW^{+}14, CRB^{+}18, CC19, CBD^{+}14, CKC16, DCSW^{+}12, DHP^{+}13, Gan13, GN12, GVDMG12, HBP^{+}18, HPS16, HHHWJ18, JG12, Jaw11, JAJ^{+}16, JBB^{+}18, KSAA13, KHV^{+}12, KGH^{+}12, LYWZ15, LMBF13, LSL12, LJJ^{+}13, MVML18, MPA^{+}12, MWR^{+}13a, MH14, NSM^{+}18, NBS15, PD18, PM11, PD14, PSS15, Pun19b, ROCC13, RCBPK11, SHT^{+}13, Szu16, SBL^{+}19, TGM13, TCAP11, VRMF10, WP15, WFJ19, WRR^{+}13, WHP15, ZW18]. métier [DMOL15, MDR^{+}19]. Métiers [LPC^{+}14, CPP^{+}10, CMPP11, DL11, PTG^{+}17, RPC11, RPP^{+}11, CMP^{+}12]. Metriatyla [RGK19]. metrics [MVMSdS19, OPR12]. Mexican [JECR^{+}18]. Mexico [RCLS18, APDI13, APDJSU11, BFM14, CSMEO^{+}11, CMMVRF^{+}19, COS^{+}19, CR18, DBSA18, EP12, EMP^{+}10, FMMW13, FPSS15, FPSDB17, FSBS11, FEGMTÁ17, GRGGCACG11, GTS^{+}17, HFD^{+}13, HPAA19, HCSM17, HFCPSEM15, HCM^{+}11, HPSG15, Ibá14, JDNP18, LSS^{+}16a, LAIVA^{+}19, LKZ^{+}17, MMF^{+}17, MDL19, MMdlP16, MFIG16, MBNM10, NMMBCV^{+}10, OGCBMR15, PF15b, PJML15, Pul17, PS19, SFB11, SMP18, SAWS18, dJSMROZP^{+}13, TWSA16, TLCSG^{+}15, WPS19, WGS17, MFRJ13]. MEY [KC18]. Michigan [LSWR17, WCS^{+}12]. Michigan-origin [LSWR17]. micro [SKY^{+}17]. micro-scale [SKY^{+}17]. microchemistry [AMV14, CBB^{+}16, CADG^{+}18, GPBA13, MAS15, MFS^{+}18, PSJ^{+}19, WRG^{+}18]. Microcystis [GBK^{+}18]. microlepis [GTS^{+}17, Sau14]. Micromesistius [CRAVC14, MAS17, STJ11]. Micronesia [CBH18, RHOCB^{+}18]. microPIT [WGC^{+}21, WGC^{+}19]. Micropogonias [CMM^{+}14, dSSdAdAA17, dVLB^{+}15]. Micropterus [AAN^{+}12a, SGA^{+}18, TWCA15]. microsatellite [CAFHG10, HDW^{+}15, LSR16, LZGL17, LCH^{+}18, RCF^{+}17, STY^{+}16, SLZG18, TVdLOG^{+}21, VRS13, WPCO15]. microsatellites [JRA10, LQL^{+}15, VPC^{+}14]. Microstomus [Smi14]. microstructure [BS10, BLPF^{+}15, Fey18, HBM11, MNT^{+}17, ODdA^{+}10, RFPG^{+}16, SP10a]. microstructures [SDC^{+}15]. microwave [LKL^{+}12]. Mid [BMR^{+}18, BSSH13, CMMF13, Koo12, LDM^{+}15, MW11, Mau12, NL11, SYJ^{+}11, WCS^{+}12, SRB^{+}18]. Mid-Atlantic [BMR^{+}18, MW11, Mau12, SRB^{+}18]. mid-reach [LDM^{+}15]. mid-water [BSSH13, CMMF13, NL11, SYJ^{+}11, WCS^{+}12]. middle [BBOL16, MHC^{+}18]. midpoints [SM15]. midwater [LW19a, SDJ^{+}15, WWH13, WHP15]. Midwestern [DLCK17, DLCK15]. migrating [BMEGV^{+}19, MEMdU^{+}17, PJM14, RIB^{+}10]. Migration [GSB^{+}19, JKK^{+}16, LSJ17, OPE^{+}17, PTB^{+}15, SMØ16, ASB^{+}16, BV14, BHT^{+}19, DHP^{+}11, FMA^{+}18, KNSJJ11, LCM^{+}18, MLT18, PFJ19, RSV^{+}17, SBAK17, SCF^{+}17a, TOTI10, VBR^{+}18b, VBR^{+}18a]. migrations [AGB15, Ber13, BTB^{+}19, CLS^{+}19, MNGPM^{+}12, OMI^{+}14, WA15]. Migratory [YAT^{+}19, AHWY13, BKT10, Goo16, HCGS19, JHT^{+}10, KNJK^{+}10, KMJ^{+}12, LSL12, SSJ^{+}13]. milii [MWR18]. millimeter [TW10]. millimeter-accuracy [TW10]. Miniature [CGKG^{+}17, MJL^{+}17]. Minimal [EHGG16]. minimally [TC17]. minimise [MFHK10]. Minimizing [FPSDB17]. minimum [MLM15]. Minisampler [MHFK12]. mis [IOT14, PACF17]. mis-specification [IOT14]. mis-specified [PACF17]. misidentification [GVMSCJ12]. misinformed [TC17]. mislabeling [MCJD^{+}17, PF15a]. mislabelling [CGdGT^{+}17]. misleading [BdSFB^{+}18]. Mismanagement [Hol11b]. Mismatch [KS17, MCM12a]. misreporting [BKS^{+}14, BKS^{+}15, CFF^{+}15]. missed [RC17]. Missing [PQ10, BKS^{+}15, HGD19, MD10]. Mission [Dam15]. misspecification [CPC^{+}17, SHJ19, WMNC15, WM17]. misspecified [DS13]. mitigate [BBA^{+}19, ZH10]. Mitigating [LCBD12, BFHT16, WSF^{+}18]. mitigation [CNS11, ERCG10, JHL^{+}19, MGR14]. mitigations [RDM^{+}15]. mitochondrial [BGA11, dSFdSSS17, GTDJdL^{+}13, HDW^{+}15, IdRH10, LSR16, PGAS15, STY^{+}16, WCC^{+}21]. mix [PDP^{+}11]. Mixed [BRH^{+}16, dMGPR^{+}18, SPH18, WLZ^{+}15, AVU^{+}10, BEM^{+}17, BB10, BHC^{+}16, CCC^{+}14b, CBM^{+}19, DJW^{+}18, DMOL15, EPO10, FA19, GCDE^{+}15, GVMSCJ12, GSP17b, GMD^{+}13, HCR14, IGS^{+}12, KRB^{+}15, MVML18, MBG^{+}10, MDR^{+}19, SNK19, SAK10, THF12, THKB12, WHC^{+}19]. mixed-effects [EPO10]. Mixed-fish [dMGPR^{+}18]. mixed-fishery [AVU^{+}10]. Mixed-origins [SPH18]. Mixed-stock [WLZ^{+}15, DJW^{+}18]. mixing [KH15, MFC^{+}13, SFH^{+}15]. mixture [AC17, CSMO14]. mixtures [CRAV13, SFPB17]. mm [CHW^{+}10, DBK^{+}10, KOF11, RIB^{+}10]. mobile [ASSL11, SGB^{+}10, Via19]. mobility [WHC^{+}19]. mobilization [ZMG^{+}14]. modality [JATBC19]. Model [CBO^{+}15, DBLB16, LYWZ15, OPR12, PHFW14, Sau14, TJMT17, WB15, WP11, APCF17, AAS19, ARF^{+}17, AVA18, AL13b, BCTS11, BNAE10, BB17, BB19, CCA16b, CHT18, CCC^{+}19, CZW14, CPC^{+}17, CA10, CIS13, Coo13, Cop13, DRGGVL11, EHH^{+}11, EHB18, FF11, GWM18, GBS19, IOT14, JKK^{+}16, KY19a, KK12, KW10, KMMV10, KNB^{+}19, LMPM11, LMPM12b, LPMM14, LPM^{+}13, LJB16, LWS19a, LWS19b, LTHRU17, LT18, LDS^{+}19, MT13, MLBTM16, MH11, MD13, Mun12a, Mun12b, Nee15, NB19, NST^{+}17, PBB11, Pop19, PMP^{+}11, QDG17, RLC^{+}15, RCE^{+}19, RMP^{+}18, SP10b, SKH14, SGGH10, SCF^{+}17a, SCF17b, SHJ19, SHT^{+}13, SDE^{+}18, SW19, TXH^{+}19, TWSA16, TM13, THHG18, THF12, VB15, VMB18, WMNC15, WM17, WSMS14, WJ17, XdA19, ZTC^{+}18, ZH10, Zha13, ZCX^{+}18, ZXHK15, vPCWV15]. Model-based [TJMT17]. model.Part [HVJP^{+}11a, HVJP^{+}11b]. Modeling [BPK^{+}17, CAC19, GOD19, Goo16, HFBS11, Hoc12, JWC^{+}13, LA13, LCZ^{+}16, LJB16, MT19, MHC^{+}18, RCF^{+}16, SWD^{+}13, TC16a, TC16b, ALH^{+}17, DLCK17, DSB^{+}15, GAR19, HFCPSEM15, KOHF^{+}16, Lor16, MS14, MDH15, MMF^{+}16, MWR^{+}13a, Qui17, RJ16, SHM12, SYJ^{+}11, XGT^{+}17]. modelled [PDP^{+}11]. Modelling [AAN^{+}12b, CN19, DTG13, ERG^{+}11, FMMW13, KEH16, LT18, MFMM16, Pun19a, RRO^{+}19, SFR^{+}13, BK12, CCA16a, DPAB10, DLD^{+}16, FPM17, LZX^{+}17, MABA12, OS12, ODMF17, PHG^{+}19, SG15]. Models [CIS13, AdM19, BHB^{+}10, BMCS19, BEM^{+}17, BGF^{+}15, BNAE13, BB10, BNK14, BSN17, CNWM12, CAM^{+}14, COS15, CBO^{+}15, CTWD15, CSMO14, CGECSEH10, DM16, EPO10, FBM^{+}19, FSFP15, Fra12, Fra14, Fra16a, Fra16b, Fra17, GBH^{+}19, GKL12, GDA14, GWB^{+}19, GDC^{+}19, HHD^{+}11, HSTK16, HHTT11, HE15, Hor11, HCR14, JCT^{+}16, KH15, KOHF^{+}16, LMPM12a, LP18, LPTK19, MAA^{+}16, MD10, Mau11, MH11, MCVS14, MDH15, MCP^{+}16, MP17, MB16, MDPH17, MRT18, NWO^{+}15, NB14, OHLBL16, OPR12, PKHD16, PLPN12, PFVGGM16, PBK^{+}19, PDP^{+}11, PDS^{+}17, PACF17, PC19, QQG14, RUMBA15, RU15, SPB^{+}19, SEG^{+}15, SMI17, SBJ^{+}19, SH13b, TSNRU21, TWM11, TW13, TT14, TK16, TJMT17, TRW19, THA^{+}14, TBSE17, TLM11, WMP^{+}14, WMdS14, WMNC15, WMT11, Whi10, XLCC19, ZCC11, ZMdSC16, vPBWA17]. moderate [JRA10, TRW19, WP15]. modern [ARVACPMN16]. modes [CCM14]. modifications [AHC^{+}11, HJLW11, LCBD12, TXH^{+}19]. modified [LS10, WHF^{+}14, WW11b]. module [BST^{+}19]. Mola [HSD^{+}17]. Molecular [BMRC14, TdMWC16, CMM^{+}11, CFF^{+}16, FdMCB^{+}19, XDL^{+}19]. Molina [MMD^{+}11]. Mollusca [CAMFMF18, FAAM16]. molluscan [BMC10]. molt [SZPV16]. molting [BLW^{+}18, Mur19, TMC19]. Molva [EWB^{+}18, HPHF15]. Monachus [RDP^{+}17]. monetary [MGC^{+}11]. money [GRO^{+}18, PHBU19]. monitor [RTML17, SWL^{+}15, VAB^{+}14, WMIDS18]. monitored [MST14, WTK10]. Monitoring [CY12, KSL^{+}16, BOBP^{+}17, BBT^{+}15, BHP^{+}19, DSWB18, DHP^{+}13, DBSA18, ER18, ELRH17, FTWC18, GSP17b, Gra16, HPRB16, HMA18, HAB^{+}17b, JECR^{+}18, KRTG16, LCMF^{+}16, LMBF13, MOS^{+}18, MAAW^{+}17, MSD^{+}16, MKH16, MCH11, PKRC13, PLMT^{+}11, PBC^{+}16, PBC^{+}13, PHBU19, PG14, QHSI18, RLC^{+}17, SGT^{+}14, WHS^{+}18, ZLT^{+}16, dCJG^{+}15]. monk [RDP^{+}17]. monkfish [SBC^{+}18]. Mono [BM19b]. Mono- [BM19b]. monodon [FQT17, HW11]. monofilament [GBD^{+}15]. monospecific [JSG^{+}15]. Montana [LS19]. Monthly [KKL^{+}16, KJKK15, WMPR17]. moored [GBR17, SMRS15]. moratoria [RGSAMV19]. morbidity [JCA^{+}16]. mordax [KBEB13]. Moreton [KCO14, PICK17]. morhua [KHV^{+}12, AJOE17, BHL12, BKT10, BK16, BJP^{+}13, BLHG17, BLHO18, BKLB14, CSKB13, DRE^{+}17, EPCB17, GBM^{+}18, HLS^{+}16, HPNA16, HBD^{+}16, KR12, KHM^{+}10, Kri13, Lam13, LDR^{+}12, LHS^{+}18c, MABA12, MM11, OTA^{+}13, OPE^{+}17, OKPL11, PRMKASR13, RBV^{+}13, RSAB11, SH13a, SHG^{+}15, SBH^{+}16, SDV^{+}16, SKDO^{+}17, TWM^{+}10, TCBR17, TSHB19, WOPB11, WHMS11, WHF^{+}14, WTK10, WAT^{+}13]. Moridae [ABL21]. morio [GTS^{+}17]. Morocco [BAT12]. Morone [HKM^{+}17]. morphologic [PGJ10]. Morphological [JBD^{+}18, LRAP19, FHM10]. morphologically [KTW^{+}13, VJH17]. morphology [KHMF11, TTMC16, THG^{+}16, VFdS10, ZLT^{+}16]. Morphometric [CDM^{+}19, KAGO^{+}10, GRGGCACG11]. morphometrics [KBG^{+}14, LRM14, MREV10, MSFŠ^{+}11, SPV19, VPC^{+}14, WMvdL10]. morphometry [AMV14, ADSV17, TTMC16, WWN^{+}14]. morphotypes [dSSdAdAA17]. morphs [SFPB17]. Mortalities [DÖA^{+}10, BKS18]. Mortality [HV10, KHS12, RBM13, AHC^{+}11, ASHH12, ASHH15, ASH^{+}17, AM12, BGBCM19, BFM14, BH13, BHC10, BHCJ12, BMF^{+}10, BMF^{+}11, BBC11, CTSD10, CHT18, CH15, CA10, Chi11, CDMC18, CIS13, CL17, CS18, DS13, DTSG^{+}16, DBH^{+}10, EBB16, EHB18, FMMW13, FHK18, Fra12, GCK14, GH15b, GHL^{+}11, GAB^{+}15, HRT^{+}21, HB15a, HB15b, HBT16a, HF14, HPNA16, HLNH15a, HLNH15b, JDNP18, KGR^{+}15, KK15, KGS18, KC11, LBF^{+}16, LMPM11, LMPM12b, LCA^{+}19, LBB^{+}13b, LWS19a, LWS19b, MCI^{+}13, MSRB15, MLTD17, MW11, MBW12, MC10, MSR14, Mun12a, Mun12b, MRT18, OLS15, OOTV12, PAR^{+}12, Pla18, PBK^{+}19, Pul17, PHT15, QWC10, RJ16, Rob15, RBW14, SGW16, SLPT17, SMP18, SNOD^{+}13, SH10, Sto12, SBC^{+}18, SW19, TM13, TWCA15, THF12, TQ17, VGE11, VBdB^{+}18, WWN^{+}15, WSF^{+}18, WMB12, WMB16]. mortality [WWG^{+}16, YGA^{+}12, YRH15]. morwong [Way13]. most [Jab18, dVLB^{+}15]. Motivation [MNCPMG15]. Motivations [KLB17, MVML18]. mount [MWR^{+}13a]. Mountain [GHL^{+}11]. mounted [SOY14]. mounting [TLLZ19]. mouth [JČK^{+}12, JRD^{+}13, JČK^{+}21]. move [LBZE17]. Movement [AGB19, BDM^{+}18, CCC17, HFL14, HKM^{+}17, KM10, MAH16, MCD^{+}13, NC13, TRH16, APDI13, ACRM12, BMCS19, BM19a, CRG^{+}18, DSW^{+}17, ESF^{+}19, FGCG^{+}17, FGK19, HF15a, HZT^{+}19, LBZE17, LWS19a, LWS19b, MFC^{+}13, MLF^{+}10, MJL14, PALC19, RTML17, RKH^{+}17, SEG^{+}15, SSP^{+}15, TS11, WFLP11]. Movements [HCH^{+}19, MAH14, SFB11, SFH^{+}15, AASF16, Ber13, CLP^{+}19, CMS^{+}15, EFW^{+}11, FMA^{+}18, FSH^{+}15, FTN^{+}14, GDW^{+}15, GBA^{+}18, SWA19, TVCT12, VB15, WGS17, WW14]. Moving [ETMK13, MDS^{+}18, WM11]. Mozambique [SSH14]. MPA [AASF16, BHG^{+}15, Eid12]. MPAs [AAOE13, WHC^{+}19]. MSC [BPPR16, ABV16, BAM16, But16, LBHP16, PSU^{+}16, SUR16]. MSY [CDPJ^{+}13, MQM^{+}15, ZCR16]. much [KOD^{+}15]. mud [IAJ^{+}11, LWY^{+}11, MMAH15, RMGK18]. Mugil [FGCG^{+}17, GCK14]. Mugilidae [BBC11]. Mulcahy [JBMK14]. Müller [JEJS17]. mullet [BCN^{+}13, BBC11, GCK14, THG^{+}16]. mulloway [DUD^{+}18, FWG11]. Mullus [ABD^{+}19, BCN^{+}13, THG^{+}16]. Multi [BPG^{+}17, KTW^{+}13, LWS19b, MLBTM16, SHR18b, BTBP18, BVN^{+}10, BST^{+}19, BSM13a, BSKT13, CMPR^{+}16, CBO^{+}15, CMJ^{+}19, DWH10, FBKE16, FWJ^{+}16, FHMK11, GS12, HSS^{+}16, KAiT^{+}13, LHL^{+}15, LWS19a, MAA^{+}16, OTST11, PTG^{+}17, Peñ18, RCBPK11, RJKG12, SKFL15, SZ11, SB16, TSNRU21, TBF^{+}16, TB10, WHK13, ZSF11]. multi-annual [BVN^{+}10]. multi-attribute [SZ11]. multi-beam [CMJ^{+}19]. multi-class [RCBPK11]. Multi-decadal [SHR18b]. multi-disciplinary [TBF^{+}16]. multi-frequency [Peñ18]. multi-gear [BTBP18, PTG^{+}17, SB16, ZSF11]. multi-geared [LHL^{+}15]. Multi-gene [KTW^{+}13]. multi-mesh [RJKG12]. multi-method [GS12]. Multi-model [MLBTM16, CBO^{+}15]. multi-module [BST^{+}19]. multi-net [BSM13a]. multi-parameter [BSKT13]. multi-phasic [MAA^{+}16]. multi-rig [FBKE16]. Multi-scale [BPG^{+}17, KAiT^{+}13, OTST11, WHK13]. multi-scaled [CMPR^{+}16]. multi-sector [FWJ^{+}16, ZSF11]. multi-species [DWH10, FWJ^{+}16, FHMK11, HSS^{+}16, SB16, TSNRU21, TB10]. multi-specific [SKFL15]. Multi-state [LWS19b, LWS19a]. multiannual [MDR^{+}19]. multibeam [MML^{+}16, PGJ10]. multicriteria [RM19]. multifilament [BM19b]. multifleet [EHH^{+}11]. multifrequency [CT16]. Multijurisdictional [FSA^{+}16]. multilevel [DPL^{+}16]. multilocus [SMO^{+}18b]. multimodel [WB15]. multinomial [BGF^{+}15, Fra14, TJMT17, TBSE17]. Multiple [BDM^{+}14, HKRH12, BMH18, HGG13, HFPH14, JAJ^{+}16, KLWG17, KVK^{+}17, Kur12, Lyn14, MMJ^{+}12, MDS^{+}18, MP17, OPF15, PDP^{+}11, RRW^{+}13, SUR16]. multiple-use [HGG13]. Multispecies [SAB^{+}11, CLWB15, Car18, CAC19, EFWC^{+}19, FLW^{+}19, ICSR16, KNB^{+}19, MMGEH15, PTG^{+}17, PBK^{+}19, Wal13, WKA13, WKA14]. multispecific [SPV19]. multistage [BTS^{+}19]. multivariate [PTSC^{+}16, WWN^{+}14]. Murky [HMvHN13]. murphyi [ASSL11, LCZ^{+}16, ZZA^{+}14]. Murray [DBH^{+}10, ZCW11]. muscle [MMMPP19, SAK^{+}17b]. muskellunge [LWSC11]. mussel [NGMA11, TJJ14]. mussels [Kni12]. Mustelus [BMdS^{+}15, CBC^{+}10, FWR10, PGM^{+}10, WJC17]. mutabilis [GPSF10]. Mutton [FMA^{+}18, GDHA13]. mutualistic [FHH^{+}19]. my [Han16b]. Mycteroperca [BB18, GTS^{+}17, Sau14]. mykiss [JNF^{+}19, MK17, TGE^{+}18, WAP^{+}12]. myriaster [LZX^{+}17]. Myxus [BBC11].

N [KTL11, FTJ^{+}16, MCN16]. Nabugabo [NC13]. Nadu [CAR19]. NAFO [NPGT17]. naïve [KBL^{+}15]. namaycush [LSWR17]. NAO [BMEGV^{+}19]. narrow [EPHF18]. narrow-barred [EPHF18]. narrownose [PGM^{+}10, WJC17]. nassa [GPSF10]. Nassarius [GPSF10]. nasus [CRG^{+}18, XYL^{+}19]. Natal [UIH^{+}18, ADSV17, CQ14, LKNL16]. National [ARVACPMN16, HHM^{+}13, GDB19, MFMM16, WGDB13]. nationally [FSA^{+}16]. Nationwide [CGdGT^{+}17]. native [FHS^{+}16]. Natural [BKS18, ASH^{+}17, AM12, AK11, Ata11, CHT18, DS13, EHB18, FHK18, Fra12, IIT17, IIT18, JDLF11, LSWR17, LBF^{+}16, LMPM11, LMPM12b, MAAW^{+}17, MSCP17, MW11, MBG^{+}10, MSV^{+}19, MRT18, QWC10, RJ16, SLPT17, SAWS18, WWN^{+}15, WW11a, WMB12, WMB16, WKT^{+}11]. naturalized [RvRW^{+}14]. Nature [HKAA18]. Nautilus [DAM10]. near [HLJ15, HFL14]. nearshore [DMQ^{+}18, EPTG19, KSL^{+}16, KSL^{+}17, SBLB17, SNOD^{+}13]. nebulifer [MJL14]. nebulosus [EHB18]. necessarily [MLTD17]. necessary [SGT^{+}14]. need [FWL18, HN16, HFO^{+}18, SMBP^{+}12, WM17]. needle [BWK^{+}13]. needs [KM15, SOMO17]. negatively [SGB^{+}10]. neglected [PRP^{+}18]. negotiation [MLT18]. Nemadactylus [Way13]. Nematoda [CSP^{+}18a, PMC^{+}18]. nematode [PG18, PBM^{+}18]. nematodes [GBM^{+}18, HPNA16, LSC^{+}18, RGAP18, TS10]. Neogobius [GBF11b, JBB^{+}18]. neon [FTJ^{+}16, KYO^{+}12, YCYC16]. neonate [SPC^{+}13]. Neotropical [AGS^{+}16]. Nephrops [ASN17, ATG11, BHM^{+}18b, BHM^{+}18a, Bri10, CFPF^{+}15, CBM^{+}19, DBK^{+}10, FBKE16, FHM10, FHMK11, HFFdlT15, Ing11, LHF16, MHFH17, MMK16, MKHK19, QWMD13, SL10]. NEPTUN [PWM^{+}16]. nerka [TQ17]. net [BBM^{+}10, BAY^{+}12, BGD11, BSM13a, CRH^{+}13, CGR^{+}15, CGR^{+}16, DJD12, FB19, HJLW11, HER13, HCM^{+}11, JČK^{+}12, JČK^{+}21, KW10, KVK^{+}17, LL14b, MMGEMGS10, SIBR^{+}19, TNKO19, TPM15, VVL18, YSS^{+}11]. Netherlands [BR15]. nets [ASMD^{+}18, ARVACPMN16, Bri10, BBM16, BM19b, EP17, HMA18, KG10, LCBD12, LSS^{+}17, PMB^{+}11, TXH^{+}19]. netting [BSB^{+}15, CHW^{+}10, HWM^{+}13, HWS^{+}15, KOF11, SB17, WAP^{+}12, WHMS11]. Network [MSCP17, vPHG11, AdM19, FMA^{+}18, KC18, MCEP14, SCJ^{+}11, SJJ^{+}11]. Networks [UPS16, EFWC^{+}19, GBA^{+}18, RBGE^{+}10, RPP^{+}11]. neural [KC18, RBGE^{+}10, RPP^{+}11]. neuston [TNKO19]. neutrally [DMC^{+}12, SGB^{+}10]. Newfoundland [DR13, DH14, DWH10, KR12, KCM19, MHRD16, MPR18, NWL^{+}14, RRRM19, SPS^{+}19, TCBR17, WW11b, WLBB15]. newly [HCWS19, LHS^{+}18c]. newly-opened [HCWS19]. next [KK17, MTH10]. next-generation [KK17]. Ngazidja [HDM13]. Nha [DFAN12]. nicked [JCA^{+}16]. Nigerian [AdM19]. Night [GBK^{+}18, JRD^{+}13, ŘJP^{+}12]. nighttime [MGR13, PAX^{+}17]. nigricans [HLA19, WRP10]. nigripenne [TRC19]. Nile [KEMM12, NC13]. niloticus [KEMM12, NC13]. nimis [FWG15, Sve14]. Ningaloo [SB12]. nitidus [NL11]. No [CSKB13, GWWW13, ABD^{+}19, FBB^{+}17, HHM^{+}13, KLB17, MMC13, MMSUAGC14, MMBD13]. no-anchor [HHM^{+}13]. no-fishing [HHM^{+}13]. no-preference [KLB17]. no-take [MMC13, MMSUAGC14, MMBD13]. no-trawl [ABD^{+}19]. nobilis [HLSD21, RCHSN^{+}15]. nocturnal [FMH13]. nomadic [Ber13]. Non [BBF^{+}18, SM15, AC17, BGL15, Ben18, CMPP11, CMP^{+}12, CM15, DDB^{+}14, FHS^{+}16, GPT^{+}19, KSAA13, MJL14, PLGL18, PCP^{+}18, PTM^{+}18, PTW^{+}13, PC19, SPMS17, STCE12, TBC^{+}19, THF12, TT14, ZSF11, dCJG^{+}15]. non-baseline [AC17]. non-consumptive [PLGL18]. non-depensatory [PC19]. non-fishery [Ben18]. non-fishing [TBC^{+}19]. non-Iberian [CMP^{+}12]. non-indigenous [PTM^{+}18]. non-invasive [BGL15, dCJG^{+}15]. Non-lethal [BBF^{+}18, KSAA13]. non-linear [THF12]. non-native [FHS^{+}16]. non-parametric [TT14]. Non-random [SM15, PCP^{+}18]. non-Spanish [CMPP11]. non-spawning [MJL14]. non-stationary [PTW^{+}13]. non-target [DDB^{+}14, GPT^{+}19, SPMS17, ZSF11]. non-tidal [CM15]. non-wadeable [STCE12]. nonlethal [LKNL16]. Nonparametric [LH18, CZW14, MH11]. Nordic [HBC^{+}17]. Nordmøre [JHL^{+}19, SBS^{+}11, SBD^{+}12]. Nordmøre-grid [SBS^{+}11, SBD^{+}12]. norm [KVH15, SM15]. normal [CRAV13]. norms [MM14]. North [BFL^{+}14, BSKT13, KHQH15, LE14, SMJB11, BDM^{+}14, CHT18, GKKK15, HBC^{+}17, HLB^{+}15, HVJP^{+}11a, HVJP^{+}11b, KWS^{+}15, KHG^{+}17, MREV10, MLC^{+}19b, SGA^{+}18, SB12, TFB^{+}12, AMQ^{+}15, AVU^{+}10, BTT11, Ben18, BN12, BMS15, CW17, CKVV16, Cor17, CAA^{+}14, DDB^{+}14, DCP^{+}16, DHP^{+}13, EMMM13, Eth15, FTJ^{+}16, GIW16, GJS^{+}13, HFFdlT15, IB10, IAS^{+}17, JB15, KF18, Kai19, KCvOE13, KTT13, LRML^{+}15, LAIVA^{+}19, LRY^{+}15, LBMC15, MCI^{+}13, MGP^{+}18, MLBGV16, MEMdU^{+}17, OtHH^{+}10, PGS^{+}11, Pop19, RDV11, RBV^{+}13, RCHSN^{+}15, Sal18, SGGH10, ST16, SPM^{+}10, Ste18, TOA12, VdSS^{+}14, VHS^{+}17, VKM^{+}17, WRP10, WKT^{+}13, WSG16, WOPB11, XTP^{+}16, dZB16]. north-central [CHT18]. north-eastern [HLB^{+}15, Cor17, LRML^{+}15, VHS^{+}17]. North-West [BFL^{+}14]. north-western [HVJP^{+}11a, HVJP^{+}11b, SB12]. northeast [BKM13, CSFCA15, FCCSA15, JHL^{+}19, MS15b, NPK^{+}10, NWL^{+}14, TVCT12, TSHB19, Wal13, AHC^{+}11, AB11, BEM^{+}17, BLHG17, BLHO18, DBA18, Han18a, Jaw11, JFT^{+}15, KBOM14, LHS^{+}18c, NAR^{+}17, PT19, PBM^{+}18, SPM^{+}19, SHG^{+}15, SBH^{+}16, SKK^{+}13, Sta16, TSS11, WFJ19]. northeastern [AB12, BLHS11, GCDAS^{+}10, HCF^{+}17, HPSG15, JLLL11, Koo12, LLH^{+}14, MAH16, NWS^{+}11, BOCÁ^{+}18, OGSNH^{+}17, ÖEG^{+}15, PHM^{+}16a, VFdS10]. Northern [CRH^{+}13, CAGVB14, DRGA12, PRR^{+}19, RM10b, RLMM18, RW19, VRMF11, VRMF12, APDI13, ACSAS^{+}17, BFM14, BWN^{+}12, BKM19, CCHD19, DOM14, EP17, FBY^{+}12, FSBS11, FTN^{+}14, GJSR10, HFD^{+}13, HPAA19, HCM^{+}11, JDNP18, JSR10, KNRJ^{+}15, KK12, Kri13, Lam13, LSŠ^{+}16b, LRY^{+}15, MSDT16, MCH11, PSM^{+}18, PALC19, RCLS18, RJ16, STY^{+}16, SFL^{+}14, SL12, UGGR^{+}17, VMB18, WSG16, WPS19, WMPR17, WGS17, AT10, CLWB15, CPP^{+}10, CMM^{+}11, GPSA11, GSP17a, GSP17b, MCN16, MQGK10, ORMG11, PBLT14, UARC11, VFMS16]. Northwest [BSSH13, GHUG^{+}16, MCJD^{+}17, Ata11, BDC^{+}17, FBHH14, FCL16, LLM12, MMD^{+}11, TAS12, BCN^{+}13, BKT10, BHC^{+}16, CBS17, DTC16, HSC^{+}12, LRY^{+}15, MCJD^{+}17, VPC^{+}14, YCYC16]. northwestern [BAT12, CMS^{+}15, RLM10, RLVS14, CLS^{+}11, SPC^{+}13]. norvegicus [ASN17, ATG11, BHM^{+}18b, BHM^{+}18a, CFPF^{+}15, CRB^{+}18, FBKE16, FHM10, FHMK11, HFFdlT15, MHFH17, MMK16, MKHK19, QWMD13].

Norway [ASN17, ATG11, BK16, BHM^{+}18b, CFPF^{+}15, FBKE16, HKAA18, JEJS17, MFHK10, MHFH17, MMR^{+}18, QWMD13]. Norwegian [BK16, HNS12, LHS^{+}18b, NNS^{+}10, WSG16, WHNS13]. Note [Ano10z, Ano15v, Ano10y, Ano19x]. notes [BOCÁ^{+}18]. nothing [LBS^{+}18]. notialis [GMUD11, MD16]. Notoscopelus [SLTPC^{+}18]. Nova [DBT15, MS12]. novaezealandiae [TRH16]. novaezeladiae [DOO15]. novaezelandiae [HKRH12, HBT16a, SHR18b]. Novel [PGL17, BHI14, Bri10, BSM15c, CM15, CTS^{+}19, FCOM18, GLB^{+}19, GLG^{+}11, KT18a, LBF^{+}16, LQL^{+}15, PTM^{+}18, RSV^{+}17, WAB^{+}16, WWG^{+}16, WKA14]. novemaculeata [CBB^{+}16]. November [Ano11v, Ano12t, Ano13o, Ano14j, Ano16t, Ano17u, Ano18q, Ano19n]. nuclear [HHB18, KBG^{+}14]. nuclear-DNA [KBG^{+}14]. nucleotide [JWM^{+}18, RPH^{+}18, TdMWC16]. number [GBR17, HWM^{+}13, PJM14, PPH^{+}17, TEFAR12, WHMS11]. numbers [FPM17]. numerical [NvZH^{+}17]. nurseries [ADSV17, PKAL13]. Nursery [RLM^{+}17, APB^{+}16, DPO^{+}11, HBM11, LAIVA^{+}19, OGSNH^{+}17, TV13, TV14, WFLP11]. nursery-dependent [APB^{+}16]. nutrient [ESF^{+}19, HNW^{+}18]. Nutritional [DPO^{+}11, SAK^{+}17b]. nuts [Cam15]. NW [AGSF19, CAP^{+}10, GPBA13, HFFdlT15, LRM^{+}10, MLV13, PFVGGM16, UJLC^{+}19, VRMF10, CAMFMF18, FAAM15, FAAM16, RGG14, VRMF13]. NW-Mediterranean [UJLC^{+}19]. nylon [CCA16a, CCA16b, GHS^{+}19].

O [KHM^{+}17, SKY^{+}17]. obesus [dSMSF15, FSH^{+}15, OTH^{+}19, SFH^{+}15, SZ10, ZCZ^{+}13, ZCD^{+}12]. Obituary [Ano16m]. Objections [BAM16]. objective [OPH15, RSGR19]. objectives [ANEA19, GCL19, MQM^{+}15, SHGP11a, WHC^{+}19]. Oblada [CMPR^{+}16]. obligation [AAH^{+}16, AJP18, FH17, PCG16]. obligations [HKEJ13]. oblique [KKH^{+}17]. oblique-banded [KKH^{+}17]. obscurus [MSRB15]. observation [BNM11, QMGP10, XLCC19]. observation-error [XLCC19]. Observations [LL14b, SPM^{+}10, SPC^{+}13, CDR16, CPS^{+}12, FBGB16, LW19b, NWL^{+}14, PLM16, RWRT15, WW11b]. observe [RTŘ^{+}12]. observed [CAA^{+}14]. Observer [PJM14, Ste18, CFBS16, FCBS15, PHM^{+}16a, PCP^{+}18]. observers [TFB^{+}12]. observing [MP14, TFB^{+}14]. obsolete [HVD^{+}18]. Obstacle [MLT18]. obtained [Fey18, JPSR^{+}14, PRBGS^{+}10]. obtaining [PUSI10]. obtusidens [APD^{+}18]. occupancy [CMM^{+}14]. occupying [KHS12]. Occurrence [RDP^{+}17, Ben18, GHA13, HPNA16, SKK^{+}13]. occurring [RM10a]. Ocean [CFH15, LKD18, RBM^{+}16, SWB12, YC18, FBHH14, GJRR14, HSD^{+}17, HJLW11, HLJ15, KKH^{+}17, RDM^{+}15, RRPG18, SCC^{+}15, SO18, SVPC16, APL^{+}18, AB12, dSMSF15, ARMLC19, ASSL11, BEM^{+}17, BB13, CFH14, CAM^{+}14, CHL11, CMS^{+}15, CIS13, CJG^{+}19, DMK^{+}19, DSB^{+}15, DLCO^{+}19, DJF^{+}15, EBB16, EMSC15, FTJ^{+}16, FCL16, FGCG^{+}17, FSH^{+}15, GH15a, GCDAS^{+}10, GTDJdL^{+}13, GAR19, GMZ^{+}14, GMUD11, HF15a, HER13, HBC^{+}17, HE15, HPSG15, HLA19, HL10, Hua15, HCH^{+}19, IAS^{+}17, JPM11, KWS^{+}15, KHG^{+}17, KH15, LKD15, LKL^{+}12, LLWC15, LAIVA^{+}19, LPMM14, LCMF^{+}16, LNL^{+}15, LFCC15, LBMC15, MMC13, MD16, MST14, MVMSdS19, MEM15, PKH^{+}11, PGLG11, PPA^{+}15, SC19, SDC^{+}15, SFB11, SFH^{+}15, SF19, SLH^{+}14, SHC14, SSH14, SSD^{+}19, SZ10, SSP^{+}15, TVCT12, TIGCDL14, WMNC15, WRP10, XLCMMV19, YCYC16, YC18, ZCZ^{+}13, ZCD^{+}12, Zis12, ZC17]. Ocean [ZMG^{+}13, ZMG^{+}14]. Oceania [PNAL16]. oceanic [MMSUAGC14, TTF^{+}13]. Oceanographic [HSD^{+}17, CR18, GPdR11, NMLZ17, SBLL17, YCYC16]. oceanography [ABSI^{+}18, LCWC17]. oceans [MSGV11, FH15, FFG^{+}16, LMSM14, MBHGN13, WWN^{+}15, WCC^{+}21]. ocellaris [BDB^{+}16]. ocellata [SBC^{+}18]. ocellatus [BHB^{+}10, CBF^{+}14, HCWS19]. October [Ano10p, Ano12n, Ano13r, Ano15m, Ano17s, Ano18t, Ano19u, WGC^{+}21]. octofasciatus [WWN^{+}15]. Octopodidae [LLH^{+}14]. Octopus [LHL^{+}15, BOCÁ^{+}18, BTT11, BN12, CAGTS^{+}10, CL17, CS18, JECR^{+}18, LH15, ORMG11, PGG10, PRBGS^{+}10, PFVGGM16, PGS^{+}11, RGG14, SSH14, BPRB^{+}14, DHFSS18, GHUG^{+}15, JPSR^{+}14, JRA10, JECR^{+}18, LLH^{+}14, LMN^{+}12, MPF^{+}17, MLP^{+}14, PRBGS^{+}10, RB12, SJR^{+}11, VFdS10]. Odontesthes [AMV14, PEAC11]. odour [WW11a]. off [ASB^{+}16, AHC^{+}11, AFOB^{+}19, ATG11, BOCÁ^{+}18, BLHS11, BAL^{+}19, BKS^{+}16, BFL^{+}14, CLWB15, CCA16a, CCA16b, CJCN21, CN18, CCC^{+}14a, CCC14c, DMK^{+}19, dADSRG10, FCNB16, FE10, FEGMTÁ17, HCF^{+}17, HA12, HND^{+}17, HFFdlT15, JSR10, KOD^{+}15, LGZC17, LXCC19, LWH19, MSDT16, MBF^{+}19b, NMH^{+}11, NMMBCV^{+}10, NWL^{+}14, OGCBMR15, PGS10, PVMG19, PGP^{+}15, PGP^{+}16, PTW^{+}13, PDF^{+}18, QMAM13, QWCB11, RRRM19, RCHSN^{+}15, RLSPdR14, RMB^{+}21, SFB11, SP10b, SZ11, SWH17, SSH14, SH10, SYL^{+}15, SCL^{+}15, TSNRU21, TCBR17, TEGM^{+}11, VFFG17, WBI^{+}16, WLBB15, YC18, ZGT12]. off-bottom [LWH19]. officinalis [AAOE13, ARF^{+}17, GHUG^{+}16, LXCC19, PVMG19]. offs [ASHH12, ASHH15, FHK18, RBW14, WP17]. offset [BBM19]. offshore [ALBŠB^{+}15, CKVV16, EAAA^{+}18, GJSR10, MCP14, RLC^{+}15, SGT^{+}14, TAS12]. ogive [FWGM19]. ogives [HSTK16, HGD19]. Ohrid [SIBR^{+}19]. oil [AWSLS15, HCSM17, RCLS18]. oils [SAK^{+}17b]. Okamejei [JLLL11]. Okinawan [UEOA19]. Old [RTSV15, vPKF^{+}19]. Oman [SVV^{+}14]. Ommastrephes [FTJ^{+}16, FCL16, KYO^{+}12, YCYC16]. Ommastrephid [RLVS14]. Ommastrephidae [PPG^{+}11, STY^{+}16, SP10b, TS10]. omnidirectionality [GBB^{+}12]. on-board [DRE^{+}17, LDR^{+}12, PCP^{+}18]. onboard [DTSG^{+}16]. Oncorhynchus [CQ14, JNF^{+}19, KKE^{+}12, LKNL16, LW12, MK17, MBG^{+}10, TQ17, TGE^{+}18, WAP^{+}12]. one [FKM11, JLF^{+}17]. one- [FKM11]. online [MCEP14, PBML^{+}10, SMGH17]. only [CWM12, GDC^{+}19, WP15]. ONS [Pop19]. onset [ATG11, QWMD13]. Ontario [KLB17, FBM^{+}19]. Ontogenetic [FdMCB^{+}19, GTS^{+}17, PPLK18, SWG18, ZDZY19, KNB^{+}19]. ontogenic [MNGPM^{+}12]. ontogeny [JD14]. Onykia [RLSPdR14]. Oocyte [SMCR13, RLMM18]. oocytes [KMKK10]. Oogenesis [BTNA13]. opalescens [RDR^{+}18]. Open [DFAN12, HBB^{+}12, CNJ11]. Open-access [DFAN12]. Open-source [HBB^{+}12]. OpenArray [CMI^{+}16]. opened [HCWS19]. opening [Ing11, JČK^{+}12, JRD^{+}13, JČK^{+}21]. openings [BSM15b]. operated [AWSLS15, HSY16, MHFK12]. operating [CMPP11, CMP^{+}12, DJW^{+}18, Hor11, SDV^{+}16]. Operational [FBGB16, LF13, BMU14, Mur15, NCP15, SMO^{+}18a, TXZ^{+}17, YvZG^{+}17]. operations [CEJI11, NS16]. operator [STCE12]. Ophidiidiae [HND^{+}17]. opilio [DWH10, DMK^{+}19, KC11, MRT17, MRT18, Mur19, NWL^{+}14, SP12, WW11b]. opinion [RW19]. opportunistic [BAL^{+}19]. Opportunities [HSG15, WDM15, vPB18, CNS11, CN19, JZL13, SK13, Via19]. opposite [HLJ15]. Optimal [ANEA19, AVB^{+}16, BHM^{+}18b, DRGA12, GHM16, HSB19, PDP^{+}11, RHB16]. optimise [SWH17]. optimization [BMU14]. options [PBD21, Pun19a, RRSB13]. orange [Ken14, KSKGR15, PHK13, PHK14, VRS13]. ORCS [FJWD17]. Order [CAC^{+}14, Jab18, VKNK12]. Orectolobidae [HSB^{+}13]. Oregon [Sam11]. organisms [BBLD16, GBK^{+}18]. organizations [Nak17a, Nak17b]. organized [FGRIT^{+}12]. Organizing [RCF^{+}16]. organs [LCWC17, SD10]. orientalis [FTN^{+}14, HSD^{+}17, IAS^{+}17, OSO^{+}18, OAT^{+}16, SKT18, UIH^{+}18, WMB12, WMB16]. orientation [BSB^{+}15, BSM16, BBS^{+}17, HWM^{+}13, JSB19b]. origin [CQ14, FBF^{+}18, Ibá15, LSWR17, SO18, TOTI10, UIH^{+}18, WRMM18]. originating [Kri13]. origins [SPH18]. Oscillation [PGS^{+}11, MEMdU^{+}17]. osmerids [WCS^{+}12]. Osmerus [GFCK13, KBEB13]. Osteichthyes [CFB^{+}17]. Ostrea [FDMH16]. otakii [LWZ^{+}19]. Otaria [MBS14]. other [AMAM^{+}17, CBS17, CPH12, DHP^{+}13, FRDC10, GH15a, GSDD11, KNSSL11, MRS15, WRR^{+}16]. Otolith [ASSL11, ADSV17, BS10, CSOV14, DR13, DBA18, FWG11, KBOM14, MHK^{+}17, MNT^{+}17, MFVPC19, RFPG^{+}16, RCHSN^{+}15, SAK10, TJROFS19, ZLT^{+}16, AB11, AMV14, ALR^{+}17, APD^{+}18, AVB^{+}16, BHG^{+}17, BB14, BLPF^{+}15, CBB^{+}16, CADG^{+}18, CPG^{+}11, CBS11, CHCS14, DH14, EMSC15, Fey18, FHK17, GPBA13, HKRH12, HBM11, HMA^{+}16, IHFAH17, IÓS^{+}14, KBG^{+}14, KBSK15, LRY^{+}15, MSFŠ^{+}11, MDG17, MAS15, MSV^{+}19, MNGPM^{+}12, MFS^{+}18, PBML^{+}10, PTB^{+}15, PSJ^{+}19, RLJ^{+}12, RFPE^{+}12, RBM^{+}16, SVV^{+}14, SKY^{+}17, dSSdAdAA17, SDC^{+}15, SLL^{+}16, SSD^{+}19, SHR18b, TRC19, TVCT12, WWN^{+}14, WRP10, WRG^{+}18]. otolith-ageing [SLL^{+}16]. Otolith-based [RCHSN^{+}15, EMSC15]. otoliths [ALJC19, CNWM12, Eth15, GMNPM^{+}13, JD14, KHM^{+}17, LP18, MFC^{+}13, MRD19, MSV^{+}15, MNGPMdMG10, ODdA^{+}10, SVN^{+}13, SHC14, ZYW^{+}14]. otter [AAH^{+}16, BHPC14, BSC12, CMP^{+}12, CGC13, DL11, GLMK12, GH14, MCI^{+}13, MBSM15, SNK19, SBLM14, TFH^{+}15]. oualaniensis [ZzCbC^{+}15]. Oulujärvi [VH12, VJH17]. outbreak [PGG10]. outcomes [BDHA11, HRC16, HF15b, PW19]. outer [GPdR11]. Outi [Han16b]. outperforms [KRB^{+}15]. output [LBS^{+}18, MVMSdS19, PL18]. outputs [HFPL16, RCE^{+}19]. Ovalipes [AK11]. ovarian [LXCC19]. over-exploited [RRM16]. over-inflow [TNKO19]. Overcoming [BGB^{+}19]. Overestimated [HL16, SA16]. Overexploitation [MSWW17, CHA15, GVMSCJ12]. overexploited [PTF^{+}18, XYL^{+}19]. overfished [MKM^{+}16, Pun11]. overfishing [Cop13, SMRS15]. Overlap [GJS^{+}13, PPdR^{+}10, SRB^{+}18]. Overlooked [CNS11, TTS11]. Overview [Gan12, FPC^{+}19, Kle13, KRB17, Yee10]. own [MVMdS^{+}17]. oxidative [SLT^{+}19]. oxygen [JD14, KHM^{+}17, LSWR17]. oxygeneios [LSR16]. oxyrinchus [BNM11, BB13, RHW^{+}16]. oyster [BPK^{+}17, MAAW^{+}17]. oysters [CDMC18, DW19, FDMH16, MMAA^{+}18]. OYU [FDMH16].

P. [SNOD^{+}13]. pāua [SKH14]. Pablo [Cor17]. Pacific [ASSL11, CHL11, FTJ^{+}16, GAR19, IAS^{+}17, LKD18, LFCC15, LBMC15, MBHGN13, PPA^{+}15, PVKH18, SFH^{+}15, Ste18, WMB16, YCYC16, YC18, APL^{+}18, dSMSF15, APCF17, ASMD^{+}18, BTT11, BN12, BMS15, BB13, CSMEO^{+}11, CBO^{+}15, CMS^{+}15, CEW14, CPB^{+}12, CS18, DHP^{+}13, DLCO^{+}19, EBB16, FCL16, FH15, FHK18, FSH^{+}15, FTN^{+}14, GCDAS^{+}10, GTDJdL^{+}13, GLB^{+}19, GRGGCACG11, GJS^{+}13, HNY^{+}19, HA12, HDW^{+}15, HF15b, HMGH12, HBG^{+}19, Hua15, HCH^{+}19, HSC^{+}12, IB10, IAS^{+}17, JD14, JDLF11, KF18, Kai19, KHM^{+}17, KMJ^{+}12, KWS^{+}15, KHG^{+}17, KH15, KMMV10, LKD15, LPMM14, LCMF^{+}16, LNL^{+}15, LCZ^{+}16, LW12, LW13, LMSM14, MT13, MFC^{+}13, MSGV11, MST14, MVMSdS19, MSV^{+}19, OSO^{+}18, OAT^{+}16, OGSNH^{+}17, PBR^{+}16, PVKH15, RCHSN^{+}15, SFB11, SF19, SAK10, SPMS17, SKT18, SCT^{+}17, SCF17b, SM14, SSP^{+}15, TEFAR12, UIH^{+}18, VRW11]. Pacific [WWN^{+}15, WDNB^{+}19, WKT^{+}13, WMB12, WPCO15, WCC^{+}21, XTP^{+}16, XLCMMV19, ZCD^{+}12, ZZA^{+}14, Zis12, ZD19]. pacificus [HJLW11, HMGH12, TS10, UK14]. package [Tho19a, Tho19b]. packages [DDP^{+}16]. packing [KMKK10]. pacu [IdRH10]. Page [Ano14c, Ano14d, Ano14e, Ano14f, Ano14g, Ano14h, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano15k, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano16i, Ano16j, Ano16k, Ano16l, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano17g, Ano17k, Ano17l, Ano17m, Ano17h, Ano17i, Ano17j, Ano18k, Ano18l]. Page/Cover [Ano17k, Ano17l, Ano17m]. Pages [Ano10t, Ano10w, Ano10r, Ano10q, Ano10u, Ano10s, Ano10x, Ano10v, Ano10p, Ano11r, Ano11z, Ano11p, Ano11t, Ano11u, Ano11s, Ano11x, Ano11y, Ano11w, Ano11v, Ano11q, Ano12s, Ano12m, Ano12q, Ano12p, Ano12l, Ano12o, Ano12t, Ano12n, Ano12r, Ano13l, Ano13p, Ano13m, Ano13q, Ano13n, Ano13o, Ano13r, Ano13k, Ano14p, Ano14i, Ano14q, Ano14o, Ano14m, Ano14n, Ano14k, Ano14j, Ano14l, Ano15n, Ano15t, Ano15q, Ano15s, Ano15p, Ano15o, Ano15l, Ano15m, Ano15r, Ano16n, Ano16r, Ano16s, Ano16o, Ano16p, Ano16t, Ano16q, Ano17o, Ano17n, Ano17w, Ano17q, Ano17p, Ano17r, Ano17t, Ano17x, Ano17u, Ano17s, Ano17v, Ano18m, Ano18v, Ano18u, Ano18r, Ano18n, Ano18s, Ano18o, Ano18q, Ano18t, Ano18p, Ano19q, Ano19r, Ano19t, Ano19v, Ano19p, Ano19o, Ano19w]. Pages [Ano19s, Ano19u, Ano21c]. Pagrus [BBH^{+}12, GCRV10, JCW12, WFLP11]. pagurus [BLW^{+}18, JCA^{+}16]. pair [CPP^{+}10, ŘJP^{+}12]. paired [BSM13b, SMK^{+}13, SAWS18]. Pakistan [RM10a]. Palau [PVKH18, PVKH15]. Palinurus [ARVACPMN16, BMF^{+}10, BMF^{+}11, GVDMG12]. pandalid [GFM^{+}12]. Pandalus [HJLW11, HLJ15, LHS^{+}18b, Sto12]. panel [BHS16, Bri10, FA19, GSH^{+}15, Ing11, JHL^{+}19, KW10, LCF15]. panels [BSM18, DBK^{+}10, ER18, MAY12, YMA12]. Panjin [LCH^{+}18]. Panopea [ZM15]. panulirid [BBM17]. Panulirus [BEM13, Ber13, GRPEMEVV17, HTB15, KT18a, MMBD13, SCOC16, TLCSG^{+}15]. parable [PW19]. Paracentrotus [BDM^{+}14]. parade [SBLB17]. paradigm [Lor16]. paradoxus [JKS^{+}15]. Paraguay [IdRH10]. Paralabrax [MJL14]. paralarvae [CAGVB14, VFdS10]. Paralichthys [FSBS11, HF14, MW11, Mau12, MWA^{+}15, PT19, PBD21, YGA^{+}12]. Paralithodes [HNS12, PHE16, WHNS13]. paramamosain [LWY^{+}11]. parameter [BSKT13, FS19, GAR10, KY19a, LPMM14, Mac13, PKWP15, PC19, SWC14, Whi10]. Parameterization [HVJP^{+}11a]. parameters [AFR^{+}17, ATG11, BBT^{+}15, CM12, CDF^{+}13, DSG^{+}17, GKL12, JLLL11, LMBF13, MBRDM18, PRMKASR13, PGJ10, PRL^{+}18, SLTPC^{+}18, SNP^{+}12, SSS^{+}12, SNOD^{+}13, SGRS^{+}18, TXZ^{+}17, TB10, XdA19, XTP^{+}16]. parametric [KOHF^{+}16, TT14]. Paranephrops [KHQH15]. Paranoplatense [VF10]. Parasite [IIT17, IIT18, CLB^{+}10, CSP^{+}18a, GGM^{+}18, MP18, MGP^{+}18, PGG10, RAGP18, WRH^{+}15, LSC^{+}18]. Parasites [MBHGN13, CLS^{+}19, FCNB16, GRO^{+}18, HCGS19, MM11, MMM^{+}10, PMC^{+}18, PRP^{+}18, VHS^{+}17]. Parastacidae [BdGM^{+}11, BdGD^{+}16]. Parastichopus [CUL^{+}16, HHD^{+}11]. Park [ARVACPMN16, HHM^{+}13, Lyn14, HSS^{+}16, LPC^{+}14, MCDP18, SB12]. Part [FCCSA15, CKVV16, FWIB14, RDV11, RBV^{+}13, TWM11, CSFCA15]. partial [AA13, SGGH10]. participation [AFA11, HGD12, QHSI18]. Participatory [NBF^{+}19, MUO^{+}17, PFVGGM16, RADR11, RLC^{+}17]. particle [GRVN10, LCZ^{+}19]. particularly [CPH12]. partition [SSJ^{+}13]. partitioning [CPP^{+}10, PGW17, VSLC^{+}19, WFLP11]. partnership [MSD^{+}16]. pass [MDS^{+}18, RIB^{+}10]. pass-through [RIB^{+}10]. passage [BPC^{+}12, CPB^{+}12, DMC^{+}12, GLB^{+}19, KCJ19, RSE^{+}14, SGB^{+}10]. passing [LCL12, MVML18]. Passive [PBC^{+}13, FBB^{+}17, HMA18, JBB^{+}18, PTTC18, ROCC13, RIB^{+}10]. passively [LCBD12]. past [VOM12]. Patagonia [CAGVB14, MREV10, MCP14, MCN16, MMD^{+}11, RRW^{+}13, VCS^{+}19]. Patagonian [MFRM19, TV14, ABL21, BBT^{+}14, BBHF10, BLPF^{+}15, CNWM12, CJWT^{+}19, DOM14, EMMM13, GPdR11, MAS15, MCN16, NGMA11, ORMG11, PPdR^{+}10, RLSPdR14, SAK10, TSNRU21, TV13, WA15].

Patagonotothen [LAB13]. patchily [SHM12]. patchiness [SABG12]. Patella [CMSG18, CAMFMF18, FPM^{+}18, FAAM15, FAAM16]. Pathways [BPC^{+}12, AGB15]. Patos [CMM^{+}14]. Patricia [Cor17]. pattern [BGF^{+}15, BHS16, FKM11, Gan13, GSB^{+}19, LCM^{+}18, LXCC19, RLM10, SKFM17, SJR^{+}11, SLZG18, VFdS10, VSTS10, ZH10]. Patterns [Mur19, PGL^{+}13, AGS^{+}16, AMGH15, ACRM12, AS10, BMCS19, BMB^{+}18, BMdS^{+}15, CMPR^{+}16, CAMFMF18, CCC17, CCHD19, CGLS12, CMM^{+}14, CKC13, dADSRG10, FCL16, FGCG^{+}17, FFG^{+}16, FGK19, GLMK12, GPBA13, GMUD11, HSD^{+}17, HFL14, HCGS19, HKM^{+}17, HMD^{+}19, HFPH14, JKS^{+}15, KRMW15, KM10, LCMTM10, LGZC17, LGH^{+}21, LSJ17, LF13, MRD19, MMGEMGS10, MD16, ML17, MMO^{+}13, MCD^{+}13, OtHH^{+}10, OPE^{+}17, PPG^{+}11, PKO10, RSV^{+}17, SMØ16, SC19, SWH17, SRB^{+}18, SM14, SSP^{+}15, TS11, TRH16, VF10, WFLP11, WSG16, ZCLB11, ZXHK15, vPW16]. paucispinis [HWS10]. paulensis [TTMC16]. Paulik [BTS^{+}19]. Pauly [Ogl17]. pay [BPS^{+}18, SCLS15]. PCR [CMSG18, GKMGO^{+}17, MRS15, PMC^{+}18, SAH15]. PCR-RFLP [MRS15, CMSG18, GKMGO^{+}17]. peacock [BDB^{+}16]. pealeii [BHPC14, Hen11]. pearl [CMS14]. Pecten [HBT16a, TRH16]. peelii [DBH^{+}10, DBH^{+}10]. pegreffii [CSP^{+}18a]. pegs [SMD^{+}16]. pelagic [AHC^{+}11, ASHH12, ASHH15, BB10, BST^{+}19, BKM13, BWR^{+}18, BS10, CLWB15, CCC^{+}19, CNS11, CIS13, CSFCA15, CMLP19, CVS11, CB11, DPB^{+}14, EE19, EBB16, EHB^{+}21, FBGB16, FCCSA15, FGS19, FPSS15, GWWW13, Gri12, Kai19, MTH10, MGR14, MCM12b, NAR^{+}17, PKH^{+}11, POM^{+}18, ŘJP^{+}12, RBGE^{+}10, RFS16, Sal18, SMK^{+}13, STJ11, SLXZ15, SRB^{+}18, SKR^{+}12, WRR^{+}16, YAH^{+}13, ZGT12]. pelagics [CBS17]. pelagicus [BMH17, JHCT11, LBB^{+}13b]. pelamis [GMZ^{+}14, MST14, OTH^{+}19, SF19, SMNdC^{+}19]. pen [JWZ^{+}19]. penaeid [BSB16, BBK12, BSC12, BSM12, BSM13a, BSM13b, BSM14, BSM15a, BSM15b, BSM15c, BSM16, BBS^{+}17, BSM18, IGSJBS^{+}15, MBSM15, MBK19, MMO^{+}13, SBS^{+}11]. penaeid-trawl [BBK12, BSM18, MBSM15, SBS^{+}11]. penaeid-trawling [BSM13a, BSM13b]. Penaeidae [TTMC16, FdMCB^{+}19, KTL11]. Penaeids [SAK^{+}17a]. Penaeus [DBB19, KCO14, RLM^{+}17]. penalize [DM16]. penalized [BB19]. penalty [Fra16a]. penetrating [KG10]. penetration [KHMF11]. penicillatus [SCOC16]. Peninsula [GQdÁdMGP^{+}11, HFCPSEM15, CAMFMF18, FAAM15, FAAM16, MCP^{+}17a]. Peninsular [INY11, INSN14, SSG17]. Pennella [LRAP19]. Pennellidae [LRAP19]. per-recruit [LCST10, Mun18, ZCC11]. Perca [HL16, HBT^{+}16b, JČK^{+}12, PGVV16, SAR^{+}15, SJA17, VKNK12]. Percalates [CBB^{+}16]. Perceived [RLC^{+}15, KDM15]. percentage [LJH11]. percentile [HF15b]. percentile-based [HF15b]. perception [CHT18, DTC16, TXP^{+}16]. Perceptions [GGADSRVL15, HFO^{+}18, GRS^{+}15, HKEJ13, KKP11, LS19, PBLT14]. Perch [HL16, SAR^{+}15, SWB12, CMS14, HBT^{+}16b, JČK^{+}12, KEMM12, LEO^{+}11, LCLM15, NC13, OTK^{+}16, PGVV16, SJA17, VKNK12, VWHB14]. Percidae [HBT^{+}16b]. Perciformes [PGAS15, SMB^{+}17, ŠBMG^{+}16]. percoides [SWB12]. Pérez [FdMCB^{+}19]. perform [PHT15]. Performance [BT16, BMH17, DS13, FHK18, KY19a, LSL12, MMAH15, PD18, YJSR12, YJC13, BSM15c, BTS^{+}19, Eid16, FPS11, FSA^{+}16, FDMH16, HC17, KKP11, KCJ19, KWF12, KT18b, LWS19b, LWY^{+}11, MMJ^{+}12, MQM^{+}15, NWO^{+}15, NSM^{+}19, OPR12, PKH^{+}11, PHM^{+}16a, PDBM12, Pun11, PTW^{+}13, Qui17, RAG19, RLFJ11, RM19, SBH^{+}16, SAB^{+}11, SB17, SAWS18, TXZ^{+}17, TXH^{+}19, UPS16, VPSH^{+}17, WAB^{+}16, WGC^{+}19, WGC^{+}21, WP11, WP15, WP17, WFJ19, WHP15, YMA12, ZTC^{+}18]. performances [BSM12, BSM13b]. period [BHB^{+}10, BSM15a, KWM^{+}15, MNT^{+}17, PMP^{+}11, WWG^{+}16, WMF13]. Periodicity [SLL^{+}16, HSB^{+}13, TEFAR12]. periods [CCC^{+}14a, DRGGVL11, LL14b]. Perkinsus [BPK^{+}17]. Permanent [RRPG18, RPK11]. Persian [Ata11, EPHF18, EPTG19, SVV^{+}14]. persistent [HF15b]. personal [Hil11]. Perspective [Tho19c, Tho21, BRB^{+}14, But16, CAM^{+}19, DPSR10, HBS^{+}16, Hil11, LSJ17, PBG^{+}13, SLP^{+}15, WGDB13]. perspectives [FBGB16, JECR^{+}18, Mar18, MFKS17, SCLS15]. Peru [ACSAS^{+}17, ASMP^{+}10, LGZC17, QMAM13, VZASM^{+}15, YC18]. Perumytilus [GPP^{+}16]. Peruvian [AS10, PAX^{+}17, PB14]. Petersen [ST13]. Phalacrocorax [HL16, SAR^{+}15]. phasic [MAA^{+}16]. phenax [BB18]. phenotypic [IHFAH17, MREV10, PRMKASR13, TJROFS19]. philippinarum [CMM^{+}11, CLF^{+}17, RvRW^{+}14]. Philippine [LvVtB^{+}15, RGK19]. Philippines [AFA11, DAM10, JSYS15]. philopatry [LAIVA^{+}19]. Phocarctos [HB15a, HB15b, Rob15]. Phocoena [HMA18]. Photo [CDL21]. Photo-based [CDL21]. photographic [PMF^{+}16]. photography [KT18a, SBA10]. photovoltaic [KBCK16]. photovoltaic-battery-LED [KBCK16]. Phycis [MSFŠ^{+}11]. phylogenetic [TC12]. phylogeny [CLF^{+}17]. phylogeographic [SLZG18]. Phylogeography [WCC^{+}21, LSR16]. Physeter [JRG^{+}18]. Physical [SBLL17, AD16, CCA16a, MRLD18, MEMdU^{+}17, PKWP16, PDBM12, RHB^{+}17]. Physiological [GHE^{+}11, OOTV12, RHC^{+}12, ASN17, BBT^{+}15, FAAM15, FAAM16, KCBS13, LBB^{+}13a, LWY^{+}11, LMBF13, MWR18, MLC^{+}19a, MMAH15, NLPS^{+}13, PHB19a, WMF13]. physiology [CRH^{+}13, CHW^{+}18, DHP^{+}11, JNF^{+}19, LWSC11, LLM^{+}19, SSFGL19, TGE^{+}18]. Piaractus [IdRH10]. picturatus [GHA13, MFS^{+}18, MFVPC19, TJROFS19, VFFG17, VHS^{+}17]. picture [RJKG12]. Pike [Cur18, CRH^{+}13, KKA10, LEO^{+}11, PALC19, SMØ16, SLB^{+}12, SFL^{+}14]. pikeperch [BRP^{+}16, GKKK15, Han16a, Han16b, HPHR14, HL16, HM16a, HM16b, JČK^{+}12, KVH15, LSŠ^{+}16b, SAR^{+}15, VH12]. Pikitch [HAB^{+}18b]. pilchardus [CHCS14, MSS^{+}16]. Pilot [MOS^{+}18, SIBR^{+}19, SRB^{+}18]. Pimelodidae [dPHS^{+}12]. Ping [LCST10]. pink [HB11, HB12, PHT15, PHLT16, TTMC16]. pinniger [HRB14]. pipeline [BLP^{+}18]. piracy [SB14]. Piroddi [RC17]. pirogues [BKS^{+}15]. piscatorius [LBV13, OAPS13]. Pisces [ABL21, KKH^{+}17, RADR11, SVV^{+}14]. piscivores [DLCK17]. PIT [BAY^{+}12, BWK^{+}13, CBD^{+}14, GBR^{+}18, JvDB^{+}17, MMAH15, PKWP16, ŠBB^{+}19]. PIT-injected [BWK^{+}13]. PIT-tagging [JvDB^{+}17]. pitfalls [PRCGD16]. PIV [BGD11]. place [HN16]. placement [DMFR16, GBR^{+}18]. Placopecten [MHC^{+}18, Mur11]. plaice [Eth15, HWS^{+}15, KCM19, MABA12, MSR14, RBM13, SKFM17, SNK19, WHF^{+}14]. Plains [BKM19]. plan [BVN^{+}10, HHM^{+}13]. plane [FKM11]. planifrons [KHQH15]. plankton [GHA13, HCM^{+}11, MBS^{+}15]. Planning [Hol11a, BMU14, BNAE13, JS15, NBF^{+}19, PBG^{+}13]. plans [CFF^{+}16, Dam15, MGB10]. plasma [MK17]. Plasticity [BdGM^{+}11, BdGD^{+}16, LBS^{+}18, AAN^{+}12a, Lor16]. Plata [MBS14, WJC17]. Plateau [NBG^{+}17, OPE^{+}17, OSMP18]. platessa [Eth15, MABA12, RBM13, SNK19, WHF^{+}14]. platessoides [KCM19]. Platform [ODD16, AWSLS15, GRO^{+}18]. platforms [LCMF^{+}16, MGLT16, MDC^{+}16, RCLS18]. Platichthys [CN19, EÖFPH17, NSM^{+}18, NNP17]. platyceros [Sto12]. platypterus [LBMC15, TCS^{+}15]. platypus [SGW16]. plebejus [CCT^{+}14]. Plectropomus [HHB18]. plei [PG14]. Pleurogrammus [MNT^{+}17]. Pleuroncodes [FQT17, HW11]. Pleuronectes [Eth15, MABA12, RBM13, SNK19, WHF^{+}14]. plumbeus [LMM^{+}17, MSRB15]. plume [GLG^{+}11]. plumes [WW11a]. PNA [GQ16]. Poey [BLHS11]. Pogonias [OMB18]. Pohnpei [RHOCB^{+}18]. point [LMHR19]. points [Ber19, Bro13b, CAM^{+}14, CDPJ^{+}13, HMMP18, JAB15, Mau12, MVMSdS19, MSR14, OMI^{+}14, PBR^{+}16, PSS14, SDWG18, TLM11, ZCC11, ZD13]. pole [SMNdC^{+}19]. pole-and-line [SMNdC^{+}19]. policies [BDHA11, BB17, BB19]. Policy [Dam15, SHGP11a, AW17, BSF^{+}19, FSFP15, UARC11]. politics [Kri18]. Pollachius [BHL12, EDG16, PHRH16b, RR12, SPM^{+}19]. polli [RFPE^{+}12, RFPG^{+}16]. Pollicipes [VRMF13]. pollock [APD10, BHI14, Koo12, PPLK18, PHRH16b, WWH13]. polluted [RMGK18]. polyactis [LSJD11, LZGL17, WLZ^{+}15, ZYW^{+}14]. Polychlorinated [MS15a]. Polydactylus [MSN^{+}12, MSWW17]. polymorphic [WPCO15]. polymorphism [RPH^{+}18]. polymorphisms [JWM^{+}18]. Polyprion [LSR16]. pompilius [DAM10]. ponds [PJR17]. ponticus [GGB^{+}18]. pontoon [LOC^{+}11, LCHL11, LCL12, LCLM15, LCF15]. poor [AM12, CHT18, DM11, DDH^{+}15a, DDH^{+}15b, EAM^{+}14, FJWD17, HLP15, KTNA15, PKRC13, RLC^{+}17, SRSNC17, SDWG18, TRW19, TAS12, WP11, ZSF11]. pop [CAA^{+}14, CAA^{+}15, LMM^{+}17, RTML17]. pop-up [CAA^{+}14, CAA^{+}15, LMM^{+}17, RTML17]. popular [CC11, LYWZ15]. Population [Ame12, CAFHG10, CLS^{+}11, CMM^{+}11, CBC^{+}10, DW19, DSG^{+}17, GJSR10, GM15a, HPSG15, HKF^{+}12, KTL11, LLH^{+}14, MSGV11, MGP^{+}18, MLBGV16, MFS^{+}18, NMMBCV^{+}10, OTST11, PGLG11, STY^{+}16, TOA12, TAS12, TVdLOG^{+}21, WMC^{+}16, ASH^{+}17, APB^{+}16, ASB^{+}13, BFM14, BB10, BdGM^{+}11, Ben18, BJZB12, BKM13, BKM19, BGA11, BN12, BBV17, CGM^{+}19, CCA16b, CHT18, CHA15, CPG^{+}11, CHCS14, DLD^{+}16, DAM10, EE19, Eid12, EHBK18, FPM17, FHK17, FSBS11, GRGGCACG11, GVMSCJ12, GVDMG12, GDB19, GFM^{+}12, Gun17, GMUD11, HSY16, HdSB^{+}14, HND^{+}17, HW11, HFCPSEM15, HVJP^{+}11a, HVJP^{+}11b, IAJ^{+}11, IÓS^{+}14, IIT17, IIT18, JKK^{+}16, JBKA15, JCT^{+}16, JRA10, JECR^{+}18, KJ17, KWCT^{+}18, KBEB13, KGB11, LSR16, LEO^{+}11, LST^{+}10, LPMM14, LPTK19, LCMdSM13, LCMF^{+}16, LH15, LBMC15, MT13, MSS^{+}16, MFRJ13, MMMPP19]. population [MDH15, MP17, MM14, MAS15, MAS17, MSN^{+}12, MFVPC19, MDPH17, MMO^{+}13, MRT18, NNS^{+}10, NGvdL^{+}13, NPRGTSR18, NLPS^{+}13, OCN^{+}11, PKWP16, PGAS15, PPG^{+}11, PGM^{+}10, PL18, PLM16, PPVP17, PR18, RGG14, RPH^{+}18, RTSV15, RSV^{+}17, RCHSN^{+}15, SMB^{+}17, SGA^{+}18, SJA17, SMP18, SSS^{+}12, SSG17, SGRS^{+}18, TVCT12, TWCA15, TK16, TLCSG^{+}15, TJROFS19, TLM11, VPC^{+}14, VRS13, WHP15, WRMM18, XYL^{+}19, ZYW^{+}14, dVLB^{+}15]. population-level [MM14, TWCA15]. Population-specific [HKF^{+}12]. populations [ABD^{+}19, Ara15, APB^{+}16, BMB^{+}18, BHG^{+}17, BDM^{+}14, BGB^{+}19, CAM^{+}19, DUD^{+}18, DLCK15, DS18, dSFdSSS17, FHH^{+}19, HKAA18, HHTT11, HF15b, HMvHN13, IdRH10, JDLF11, JWM^{+}12, JBB^{+}18, KKH^{+}17, KAGO^{+}10, KY17, KDPMSR10, KCM19, LFCC15, MCP^{+}17a, OCN^{+}11, PTB^{+}15, PEU19, PVKH15, PVKH18, PGVV16, QQG14, RG18, RGK19, RMMSR19, RJKG12, SNP^{+}12, SSJ^{+}13, SMO^{+}18b, SP15a, Ste11, VWHB14, VvHE^{+}11, VFdS10]. porbeagle [CRG^{+}18]. Porcupine [LBV13]. porcus [ÖLTA19]. porgy [CLS^{+}11]. porpoises [HMA18, LE14]. port [MCH11, FWR10]. portable [RIB^{+}10]. Portugal [PVMG19, BDM^{+}14, GVG18, LBZE14, MBF^{+}19b, VGE11]. Portuguese [ATG11, BBM^{+}10, BKS^{+}16, CPG^{+}11, GPT^{+}16, JF13, MLP^{+}14, NVS^{+}18, SGN^{+}10, SNP^{+}12]. portunid [BM18b, BM19b]. Portunidae [RM10a]. portunids [BBM16]. Portunus [BMH17, JHCT11, LBB^{+}13b, LSH13, LCH^{+}18, RM10a]. portusjacksoni [FWR10]. posed [MH14]. position [BST^{+}17, HPB10, MCN16, PPLK18, RSEERRSC15]. positioning [BCP^{+}12]. positions [MFHK15]. possibilities [HBC^{+}17]. Possible [CDF^{+}13, BDC^{+}17, MP18, SNP^{+}12]. Post [ALUFJ^{+}12, BBC11, GH14, HLSD21, LBCD18, PLA16, RKH^{+}17, SHA^{+}15, SWA19, BMOZ13, DHP^{+}11, EBB16, GH15b, HRB14, IAS^{+}17, JDNP18, KCBS13, KGR^{+}15, LCWC17, MCI^{+}13, MSRB15, MWR18, RMT^{+}18, RBW14, WSF^{+}18, WWG^{+}16]. post-bomb [IAS^{+}17]. Post-capture [GH14]. Post-escape [ALUFJ^{+}12]. post-landing [BMOZ13]. post-recompression [HRB14]. post-recruit [LCWC17]. Post-release [BBC11, HLSD21, LBCD18, RKH^{+}17, SHA^{+}15, SWA19, DHP^{+}11, EBB16, GH15b, JDNP18, KCBS13, KGR^{+}15, MCI^{+}13, MSRB15, MWR18, RMT^{+}18, RBW14, WSF^{+}18, WWG^{+}16]. Post-Soviet [PLA16]. postovulatory [Gan12, KMQ^{+}10]. posts [MCEP14]. posture [KG10]. pot [AK11, BHL12, CS18, HBD^{+}16, OHSG19, SJS^{+}17, SJR^{+}11]. pot-caught [HBD^{+}16]. potamodromous [BDM^{+}18, PKWP15]. potato [PHB^{+}19b]. Potential [CFPF^{+}15, CADG^{+}18, GGB^{+}18, PKAL13, WOPB11, BMEGV^{+}19, BFHT16, BHC10, BLR^{+}16, Bro13b, CMSG18, CBMLB13, CCGD19, EL18, FGCG^{+}17, FHH^{+}19, GBPC10, HLP15, HMD^{+}19, IOT14, JS15, Ken14, KHGB14, KMSRW10, Lam13, LSŠ^{+}16b, LCH^{+}18, LFS^{+}19, MBHGN13, MFRJ13, MS15b, MDPH17, MML^{+}16, NMLZ17, PHK13, PHK14, RBT14, RM10b, SPC^{+}13, TMC15, TFB^{+}14, VLSP19, WLBB15, Wri13, YC18]. potentially [LFCC15]. potentials [PRCGD16]. pots [BKLB14, JLF^{+}17, KLH^{+}15, OKPL11]. potting [MTC^{+}17]. poutassou [STJ11]. pouting [RDV11, RBV^{+}13]. power [BBM17, ERG^{+}11, TIGCDL14]. practical [LG13, PM11]. practice [BHP^{+}19, DM16, HBP^{+}18, MT19, MGR14]. practices [BTBP18, CHW^{+}18, CMLP19, DDK^{+}18, HRT^{+}21, JRG^{+}18, LT16, MS14, Nak17a, Pun19a, RFC^{+}17, SGSFL19, SMGH17, STAN^{+}17]. pragmatic [KSL^{+}17]. prairie [VWHB14]. prawn [ASD^{+}18, BSB^{+}15, CCT^{+}14, DBB19, FRDC10, IGSJBS^{+}15, KCO14, PICK17, RDM^{+}15, SB17, Sto12]. prawn-trawl [SB17]. pre [EDG16, MWR^{+}13a, MLP^{+}14, RDR^{+}18, TGMM13, TQ17]. pre-recruit [MLP^{+}14, RDR^{+}18, TGMM13]. pre-spawning [MWR^{+}13a, TQ17]. pre-stunning [EDG16]. precautionary [Mun12a, Mun12b]. Precise [HZT^{+}19]. Precision [ZYSZ18, JSB19a, Mac13, OPR12, TWM11, VB15, XdA19]. precocity [KNSSL11]. Predation [CAA^{+}15, MBS14, APD10, HL16, KBL^{+}15, PTC^{+}18, RJ16, SA16, SJA17, TLM11]. predator [CAM^{+}19, KBL^{+}15, SWG18, WAB^{+}16]. predators [ABD^{+}19, GMSP19, HAB^{+}17a, HAB^{+}18a, KJS12, PBB^{+}18, RDP^{+}17]. predatory [KS12]. predict [BSN17, NWO^{+}15, VKNK12, vPW16]. predicted [RHB16]. Predicting [BRP^{+}16, DBB19, HLS^{+}16, HMD^{+}19, KGS18, MSV^{+}19, RDR^{+}18, vPWW16, Jaw11, OOM15, PTG^{+}17, Rod19, Sto12]. Prediction [LLP18, NST^{+}17, Che10, KKK16, THKB12, ZLT^{+}16]. Predictive [JIM^{+}15, THA^{+}14, FF11, MWR^{+}13a]. predictor [CSP^{+}18a, YRH15]. predictors [MSFŠ^{+}11]. Preface [Ano15u, Bro13a]. preference [KLB17, LL14a]. preferences [CRCMP^{+}14, CB17, Cur18, GFJ19, KL16, SWB12, WGS17]. Preliminary [FTJ^{+}16, FRDC10, PPdR^{+}10, EPTG19, GWVH^{+}16]. premature [SL11b]. Preparing [PGG10]. Presence [HMA18, BSS^{+}13, GDC^{+}19, HJT13]. presence-only [GDC^{+}19]. Present [GPP^{+}16]. Present-day [GPP^{+}16]. presented [Ano10-27, SA16]. preserve [GRO^{+}18]. preserved [ZYSZ18]. preserving [Fey18]. Press [Hol11b, GMCL^{+}18]. pressure [BSKT13, FTM^{+}15, KMMV10, LSŠ^{+}16b, PBG^{+}13, RJ16, ST16]. prevailing [WOPB11]. Prevalence [HHSS17]. prevalent [LBH15]. prevent [KLH^{+}15, RCF^{+}17]. preventing [SMK^{+}13]. previously [MBOAA12]. prey [BAL^{+}19, BDB^{+}11, BBLS17, KJS12, SWG18, WPS19]. Price [MCM12b, SKT18]. prices [GM15b, SUR16]. primer [EL18, Mar18]. primers [PMC^{+}18]. primers/probe [PMC^{+}18]. Prince [WDNB^{+}19]. principle [Han18b]. Prionace [ARMLC19, BB13, CNS11, CAM^{+}14, KWS^{+}15, MQV10, MQCS^{+}14, PBB^{+}15b, RHW^{+}16, TOA12, VdSS^{+}14]. prior [CTWD15, WOPB11, ZZA^{+}14]. priorities [ASMD^{+}18, SP12]. priors [CC19, TC17]. prism [ST16]. Pristipomoides [KKH^{+}17]. Private [PLA16, DSWB18]. pro [MCB^{+}18]. pro-environmental [MCB^{+}18]. Probabilistic [KVH15, MM14, SM15]. probabilities [FL13]. Probability [BCP^{+}12, VAB^{+}14, BLW^{+}18, Har19, HDS^{+}18, JATBC19, JLF^{+}17, KS12, MFMM16, PD14, Seu21, SZPV16, ŠBB^{+}19]. Probability-based [VAB^{+}14]. probatocephalus [ALB^{+}18]. probe [PMC^{+}18]. problem [HBB^{+}19, PG18, vPKF^{+}19]. problems [Gan12]. Procedure [BAM16, MQGK10, PTW^{+}13, SBLB17, WKA14]. procedures [DPB^{+}14, GJRR14, GAR10, HSS18, HC17, HMGH12, TB10, UASM12]. process [CCC^{+}19, KK12, KYO^{+}12, MAA^{+}18, MCP^{+}17b, NSM^{+}19, OLM^{+}19, SM17, RBR^{+}15]. processed [PF15a]. processes [DHP^{+}13, SBLL17]. processing [CDB^{+}14, HBB^{+}12]. Prochilodus [ADSV17]. produce [GDA14]. produced [PBG^{+}13]. product [SJS^{+}17]. Production [SSS^{+}12, AW12, BMU14, BSW^{+}12, BPS^{+}19, CDPJ^{+}13, CBO^{+}15, CJP13, CSKB13, DSG^{+}17, DCSW^{+}12, GN12, HHD^{+}11, HHH^{+}12, HBG14, HW11, Hor11, JCW12, KHV^{+}12, LF13, MABA12, MDPCSR10, MPA^{+}12, MFRM19, MRT17, MIÁ^{+}10, NNS^{+}10, OHLBL16, PS12, WMdS14, WJ17, XdA19, XLCC19, Zha13]. productivity [Ber19, BKM13, BJKH12, DHP^{+}13, EHGG16, FMB^{+}19, FEGMTÁ17, Hol11a, HF15b, INY11, KWR19, KOD^{+}15, KvZ12, KTT13, Lam13, LSM^{+}19, MSR14, NML12, PSE15]. products [CDB^{+}14, HIH19, MCJD^{+}17, OPL17, PF15a, PMM^{+}15]. productus [LW12]. professional [SCV^{+}12]. profile [FBF^{+}18, GDB19, KFGN10]. profiles [BV14, FHK17, RCF^{+}16, RSGR19]. profiling [LPMM14, LCH^{+}18, WMP^{+}14]. Profit [BAD19, QMAM13]. profitability [BNAE10, GAM10]. profits [TD19]. profound [MSWW17]. program [BMC10, FTWC18, HF14, HLSD21, MCH11, PM13, Ste18]. Programme [MEM15, CGdGT^{+}17]. programmes [LNL^{+}15]. programs [ER18, GH14, GPT^{+}19]. Progress [LKR11]. Progressing [BSM13a]. Progression [TWSA16]. Progressively [BSB16]. project [EFWC^{+}19]. project-specific [EFWC^{+}19]. promote [BRH^{+}16, Nak17b]. Promoting [GQ16]. Proof [WKA14]. Properties [WW11a, DPB^{+}14, DPO^{+}11, MK13]. property [EB15, Via19]. proportions [SH10]. proposal [POV19]. proposed [LF13, SWC14]. prospects [GDB^{+}16]. protandrous [MSWW17]. protect [AAOE13, ADH^{+}13]. Protected [Ano13a, GBPC10, HHCP13, MDPH17, RRSB13, WGDB13, AA13, BMF^{+}10, BMF^{+}11, CMPR^{+}16, DSG^{+}17, FL11, HJT13, HGG13, JSR10, LMB^{+}11, LA13, MCDP18, ML10, MLV13, RMB^{+}21, SF13, TLCSG^{+}15, UEOA19, WM11, WMC^{+}16, WGM^{+}11, WMM^{+}15]. protection [GHJ13, HVJP^{+}11a, HVJP^{+}11b, JSR10, KRB^{+}15]. protein [CSKB13, SJS^{+}17]. protocol [BNA^{+}19, GMSP19, OLM^{+}19]. protocols [AVP^{+}15b, GSTL15, SMS17]. protogynous [NVS^{+}18, WWN^{+}15]. provenance [RSV^{+}17]. provide [ADSV17, BTBP18, HCGS19, LKNL16, MBB13, WRR^{+}16]. provided [Pre17]. provides [RJKG12]. Providing [PBM^{+}19, Nak17a, WLZ^{+}15]. Province [ZXH^{+}12]. proxies [DF16, PSS14]. Proximate [LRM14, DPSRM10, KKL^{+}16, RBG^{+}19a]. proximus [KDdAA^{+}18]. proxy [HFPH14, LH15]. PSA [ZHDS16]. PSATs [LMM^{+}17, RTML17]. Psetta [FF10, HBM11]. pseudoharengus [BSSH13]. Pseudophycis [KJS12]. Pseudopleuronectes [TMM19]. Pseudoterranova [PMC^{+}18]. public [GGS^{+}19, TZW16]. publication [Cor17, LRM17]. publications [MP14]. Publisher [Ano10y, Ano10z, Ano15v, Ano19x]. Puducherry [CAR19]. Puerto [BDB^{+}16]. pufferfish [GGS^{+}19]. Puget [LSJ17, CQ14, HHCP13]. Pulse [DRGA12, DCP^{+}16, RSEERRSC15, VLSP19, vMWvOKvB14]. pulsed [DCP^{+}16, JTHT14, WTF16]. pulses [ICSR16]. pump [PSL^{+}19]. Pumping [DTSG^{+}16]. punctata [GCRV10]. punctatus [AK11]. puniceus [DJF^{+}15]. pupfish [PTTC18]. pupping [SPC^{+}13]. purple [BDM^{+}14]. purposeful [JP16]. purposes [dEWB^{+}15, dZB16]. purpuratus [GPP^{+}16]. Pursat [TBC^{+}19]. Purse [LCMF^{+}16, CHW^{+}18, DTSG^{+}16, DLCO^{+}19, EBB16, GQ16, HHH15, HV10, KBCK16, KZ11, LOdH16, LLP18, LvVtB^{+}15, LMSM14, TXZ^{+}17, TXH^{+}19, TPM15, TIGCDL14, XLCMMV19, ZXHK15]. Purse-seine [LCMF^{+}16, DLCO^{+}19, XLCMMV19]. pursing [HV10]. pursuit [GHPJ15]. Pushing [EWB^{+}18]. pyrosequencing [GKOC^{+}16].

Qualitative [MMG10]. Quality [OJT^{+}14, RSAB11, ALJC19, BLHO18, DTSG^{+}16, DRE^{+}17, DPL^{+}16, DMK^{+}17, KDSAH^{+}18, KBČ^{+}16, MAP14, OTA^{+}13, RLMM18, SAK^{+}17b, SKDO^{+}17, WBM17]. quantification [BPG^{+}17]. quantify [EFW^{+}11, FHM10, GBA^{+}18, PGL17]. Quantifying [AHMS10, BRW^{+}13, GBB^{+}12, PBB11, RSE^{+}14, TTS11, LWZ^{+}19, MPF^{+}17]. Quantitative [BHPC14, BBH^{+}14, HLSD21, Pit14, ZSF11, BHC10, CCC^{+}14a, SMS^{+}19]. quantities [CV14, JEJS17]. quantity [DMK^{+}17]. queen [BdGN^{+}18, SMBP^{+}12]. Queensland [CMS14, CCT^{+}14, HSS^{+}16, HSM^{+}12, ML11, MSWW17, RDM^{+}15]. Quelle [Cor17]. question [FCLH19]. Questioning [AAH^{+}16]. questions [MH14]. quickly [KFGN10]. quid [FWG15, Sve14]. quillback [Hoc12]. quota [BMG16, BGB^{+}19, Mar18, MMF^{+}16, PHBU19, SSR18, vPHG11]. quota-based [MMF^{+}16, SSR18]. quotas [KPL^{+}07, KPL^{+}09].

race [SMPB17]. race-for-fish [SMPB17]. radio [GBB^{+}12, KKE^{+}12]. radio-controlled [KKE^{+}12]. radiocarbon [DTC16, IAS^{+}17, KHG^{+}17]. radiometric [CAC^{+}14]. RADseq [SRMP^{+}18]. raft [KBCK16]. raiding [KFBL13]. Rainbow [CM18, KBL^{+}15, KBEB13, MK17, PMG18, VvPM^{+}18, WAP^{+}12]. raised [CPS^{+}12]. Raja [KM10, KHG^{+}17]. Rajidae [MCI^{+}13, QWC10]. rajids [KRM^{+}18]. Rajiformes [IIT18, IIT17]. Ramsay [CMS14]. ramsayi [LAB13]. random [BAFP18, BB17, PLM16, PCP^{+}18, SM15, TK16, WXZ^{+}18]. random-at-length [PLM16]. randomly [BB19]. Range [JWM^{+}18, GHE^{+}11, JBKA15, ML10, MNW12, NCP15, NC13, ÖLTA19, PBK^{+}19, RSEERRSC16, TS11, YC18]. range-shifting [YC18]. Range-wide [JWM^{+}18]. Ranina [SBLL17]. Rapid [AWSLS15, ASMD^{+}18, Fit12, RG18, WWN^{+}14, BDB^{+}16, CPB^{+}12, FEGMTÁ17, GJRR14]. rare [GWWW13, WWR18]. rare-earth [GWWW13]. Rastrelliger [NMLZ17, PGAS15, SSG17]. rate [BHT^{+}19, BK15, CAMW11, CJCN21, CGG21, CHL11, CSMO14, GH15a, KGB11, LJH11, LMM^{+}17, MK17, MS18, PBK^{+}19, SVG^{+}17, VOM12]. Rates [EBB16, AMGH15, AMHR16, BNSS19, BHI14, BHCJ12, BJD12, CFH14, CFH15, CGdGT^{+}17, COS15, CLS^{+}11, GN12, Gri12, HFD^{+}13, HF14, Hen11, HSM^{+}12, HWS^{+}12, IB10, JASP11, JLF^{+}17, Kai19, KHS12, KKA10, LLWC15, LYWZ15, LWS19b, MMGEH15, MS18, MBW12, OGCBMR15, PB11, PKO10, PTC^{+}18, SGW16, SABG12, TSNRU21, TW14, TMC15, TTF^{+}13, VIP13, WMB12, WMB16, YRH15]. Rathbun [MCC^{+}17]. ratio [BK12, CCA16a, CNWM12, Car18, Cer14, HLP15, KLWG17, KT18b, SMP18]. ratio-at-length [Cer14]. rationale [NGMA11]. ratios [APDJSU11, CBS11, CCGD19, PGP^{+}15, PGP^{+}16, SHR18b, WDNB^{+}19]. Ray [WDBH16, BGL15, SAH15]. rays [ACSAS^{+}17, LKNL16, WDBH16]. razor [COC14, FMD19, MTV17, MCB^{+}16]. Re [MPR18, RGAP18, ZD19, ER18, MAH16, UASM12]. re-entry [MAH16]. Re-evaluation [RGAP18, ZD19, UASM12]. re-stocking [ER18]. Re-visiting [MPR18]. reach [LDM^{+}15]. reaches [LCST10]. reaction [BHPC14, KVH15, MM14, OTH^{+}19, SM15]. reactions [JRD^{+}13]. read [SRMP^{+}18]. readability [CNWM12]. reader [CNWM12]. readings [HCHA12]. reads [HSTK16]. Real [KRB^{+}15, PMC^{+}18, DJW^{+}18, GPT^{+}19, KRC14]. Real-Time [KRB^{+}15, DJW^{+}18, GPT^{+}19, KRC14]. realised [WAT^{+}13]. reality [CCC^{+}14b]. realize [SLPT17]. reallocation [THKB12]. Really [HH16]. reappraisal [GVG18]. reared [CFB^{+}17, CSOV14, KNRJ^{+}15, Kri13, LCH^{+}18, PEU19, RBG^{+}19a, SMJB11, SBAK17]. rearing [KNSJJ11, LKNL16, SKY^{+}17]. reasoning [BHCJ12]. reassess [BSKT13]. Reassessment [SPM15b, FSH^{+}13]. rebound [MFRJ13]. rebuilding [Pun11]. recapture [dSMSF15, BAY^{+}12, Eth15, EMSC15, GVDMG12, HF14, HE15, KM10, MLF^{+}10, MAH14, Sau14, SMJB11, SZPV16, SW19, TJJ14, TSS11, ZG14]. recapture-conditioned [MLF^{+}10]. recaptured [CLS^{+}11]. recaptures [ALUFJ^{+}12]. recently [ER18]. reception [GBB^{+}12]. recertification [BPPR16]. recognize [FCNB16]. recommendations [CC19, HBP^{+}18, Nak17a]. recompressed [RHB^{+}17]. recompression [BRS19, HRB14]. Reconciling [CP11, CP12, KVK^{+}17, VF13]. Reconsideration [OS12]. reconstruct [FPC^{+}13]. Reconstructed [LBZE14, SP10a]. Reconstructing [CAAB^{+}18, CCC^{+}14b, EE19]. Reconstruction [PGZ^{+}15, KMQ^{+}10, RC17, SRSNC17]. record [RFPE^{+}12]. recorded [BNU^{+}10, RTSV15]. recordings [GFCK13]. records [PTG^{+}17]. recoveries [BHB^{+}10, FH15]. Recovery [FHH^{+}19, KDSAH^{+}18, RGSAMV19, RLFJ11, CBF^{+}14, GBH^{+}19, GH14, HSY16, HTB15, HBD^{+}16, KSKGR15, KCBS13, LWS19a, LWS19b, MBOAA12, SFR^{+}13, VB15, WWG^{+}16]. Recreational [BR18, GDB19, Gri12, AW17, AMGH15, AMHR16, AA13, ANEA19, BDHA11, BdGM^{+}11, BBA^{+}19, BJKH12, BAFP18, BMR^{+}19, BBM16, BBM17, BM18a, BHP^{+}19, CRPAMN17, CPMNS10, CAP^{+}10, CAC19, CMHP14, DDK^{+}18, DBG19, DBH^{+}10, FL11, GFJ19, GH15b, GK18, GCRV10, GRS^{+}15, GHL^{+}11, HPRB16, HAB^{+}10, HF14, HVD^{+}18, HHHWJ18, HMD^{+}19, HVJP^{+}11a, HVJP^{+}11b, IWL^{+}11, JATBC19, KSL^{+}16, KSL^{+}17, KDPT18, LMHR19, LCB17, LČ18, LT16, Lyn14, MOS^{+}18, MVML18, MCB^{+}18, MMG10, MNCPMG15, PHG^{+}18, PAM^{+}18, PPH^{+}17, PALC19, RHC^{+}12, RFC^{+}17, RR12, SCC^{+}15, Sau14, SK13, SPMS17, SHA^{+}15, SB12, SHR^{+}18a, SH13b, TWM11, TZW16, TÜM^{+}16, TXP^{+}16, VGE11, WLGS19, ZGT12, ZG14, vPCWV15]. recreationally [LBB^{+}13b, LBCD18]. recreationally-angled [LBCD18]. recruit [CMLP19, DLCK17, GKL12, LCST10, LCWC17, MLP^{+}14, Mun18, RDR^{+}18, SCF17b, TGMM13, Whi10, ZCC11]. recruited [KMKK10, LC19]. Recruitment [DBLB16, HBT^{+}16b, KEP^{+}19, KCM19, SPB^{+}19, ZCLB11, AD16, BA12a, BRP^{+}16, BMS15, BTS^{+}19, CGM^{+}19, CAM^{+}19, CHT18, CCC^{+}19, CJWT^{+}19, CA10, DLCK15, FPM17, Fra16a, GCDAS^{+}10, HBB^{+}19, Han16a, Han16b, HF19, HPHR14, HM16a, HM16b, JCT^{+}16, KF18, KWR19, KTT13, LSWR17, LMPM12a, LLM12, LGH^{+}21, LMLBC13, MS15a, MSCP17, Mau12, MD13, MT19, Mun12a, Mun12b, MPR18, MYJ^{+}15, PHG^{+}19, PD14, PSS14, PHT15, PC19, Pun19a, RvRW^{+}14, SMRS15, SPS^{+}19, SBL^{+}19, TGMM13, TRW19, TCBR17, WMP^{+}14, Way13, WFNC18, WA15, ZCD^{+}12, ZD19]. recruitment-mortality [Mun12a, Mun12b]. recruitments [DM16]. rectangular [BM18a]. recuperation [SKDO^{+}17]. recurrent [CGLS12]. Red [HCWS19, SB15, ABL21, BHB^{+}10, BCN^{+}13, BBB^{+}16, CTSD10, CBF^{+}14, CUL^{+}16, FMMW13, FGK19, GTS^{+}17, GMUD11, HW11, HNS12, HPSG15, KJS12, KW10, LKZ^{+}17, MD16, MMD^{+}11, ORMG11, PF15b, PHE16, SCOC16, THG^{+}16, TS11, TEGM^{+}11, WGS17, WHNS13, ZCLB11, HCSM17]. redd [DPAB10]. redesign [BFW13]. redfish [CEB^{+}19, CRB^{+}18, LHS^{+}16, PHRH16a, VPC^{+}14]. reduce [AHC^{+}11, BRJ^{+}16, BSM15b, BWK^{+}13, FA19, GGM^{+}18, GAR10, HJLW11, HB11, JLF^{+}17, LE14, LHS^{+}18a, LW12, LW13, MM13, PAR^{+}12, QEH^{+}11a, QHSI18, RRPG18, RPK11, RBG^{+}19b, VVL18]. Reduced [HNS12, LHF16]. reduces [HB13, KOHF^{+}16]. Reducing [BBM17, MGR13, PF15b, SHM^{+}16, SGW16, dMGPR^{+}18, HB15a, HB15b, Mur15, PICK17, Rob15, SPMS17, VLSP19, WTF16]. Reduction [DM11, WDBH16, BBK12, CHW^{+}10, CW17, CPS^{+}12, CBM^{+}19, CCT^{+}14, CSMJB11, EB15, HJ12, HLJ15, LHS^{+}18b, LW16, LW19a, MAY12, PF15b, PTM^{+}18, TC15, WvPLW18]. reedi [CCA16a, CCA16b]. Reef [BCKG13, FTWC18, HSS^{+}16, KMMV10, APDI13, ASB^{+}13, BTBP18, BSS^{+}13, BST^{+}19, BH11, BRW^{+}13, CT16, Chi10, Chi11, CBH18, DBSA18, EMP^{+}10, EvHRF10, FMH13, FGK19, HFL14, JP16, KSL^{+}16, KSL^{+}17, LLH^{+}14, MMC13, MBJ^{+}16, MMK10, MM13, PWB^{+}16, PLMT^{+}11, PVKH15, PVKH18, Pul17, PS19, RMB^{+}21, SWD^{+}13, SAB^{+}11, WGS17, ZBB17, EvHRF10, RMT^{+}18]. reef-field [BST^{+}19]. reef-fish [SAB^{+}11]. reefs [AWSLS15, FTWC18, INSN14, ÖLTA19, RBV^{+}13, SOMO17, TS11]. reel [KCBS13]. reestablishment [BDM^{+}18]. Reeve [CMM^{+}11]. Referees [Ano10n, Ano10o]. Reference [Cor17, LRM17, Ber19, BMAMG18, Bro13b, CDPJ^{+}13, HMMP18, JAB15, Mau12, MVMSdS19, MSR14, OMI^{+}14, PBR^{+}16, SDWG18, TLM11, WA15, ZCC11, ZD13, Zis12]. refined [FJWD17]. Refinement [BNA^{+}19, GWM18, PHE16]. Refinements [FWJ^{+}16]. Refining [MBK19, SBS^{+}11]. reflect [KJ17, SP15b]. reflecting [LCM^{+}18]. reflects [BHG^{+}17, ZQZ^{+}17]. reflex [PHB19a, Sto12, YRH15]. reflexes [KGS18]. reformed [Dam15]. Reframing [WM11]. refrigerated [KGR^{+}15]. regalis [RLM10]. Regan [LAB13]. regard [Cor17]. regarding [HFO^{+}18]. regia [PEAC11]. regime [GVSB11, KOD^{+}15, LSM^{+}19, TKF11]. regimes [FH17, NST^{+}17]. Region [BdGN^{+}18, DWH10, dSFdSSS17, GHL^{+}11, HBT^{+}16b, IdRH10, MDPH17, PLA16, TKF11]. Regional [DS18, KDPMSR10, MAS17, RGK19, WRP10, CBS17, DSB^{+}10, FCLH19, GTCC15, HGD12, JWM^{+}18, KK12, LAIVA^{+}19, MSCP17, Nak17a, Nak17b, SSP^{+}15, TDH19, VB15, WPS19]. regions [DH14, GCK14, GSP17a, MCJD^{+}17]. registration [HPB10]. regius [GQdÁdMGP^{+}11, MNGPMdMG10, MNGPM^{+}12]. Regression [Sau14, FF11, GFK^{+}12, LCMTM10]. regressions [MDH15]. regulated [VBR^{+}18b, VBR^{+}18a]. regulating [Nij15]. regulation [BHB^{+}10, GGADSRVL15, JASP11, MS12, VFFG17, ZMG^{+}13]. regulations [ANEA19, FDN^{+}11, HVD^{+}18, KL16, TBF^{+}16]. regulatory [BDHA11, HKEJ13, WHS^{+}18]. Reinhardtius [DTC16, GLA19, HSL^{+}13, KHGB14]. Reinserting [ABV16]. reintroduction [CFF^{+}16]. related [APAFMN10, CADG^{+}18, FHK17, HNS12, KKA10, MMGEMGS10, MAP14, MWR18, MGR14, MFIG16, NPGT17, RGGdL13, VCS^{+}19]. Relating [CTSD10]. Relation [MSS^{+}16, AW12, BNK14, CCM14, DSG^{+}17, FBHH14, GPdR11, GFM^{+}12, ICSR16, KNSPK13, LRM14, LLM12, MK13, OSMP18, OCN^{+}11, PMP^{+}11, YLA15, YCYC16, ZMG^{+}14]. Relations [WMF13, ERG^{+}11]. Relationship [JG12, LKL^{+}12, MWSN^{+}13, UK14, WCS^{+}12, CHT18, CJCN21, DOO15, GCDAS^{+}10, GKL12, GBEP16, GRB14, Han16b, HPHR14, HM16a, HM16b, LMPM12a, LCWC17, Mau12, MMK10, PD14, TRC19, TGMM13, WFNC18, ZCD^{+}12]. Relationships [DJD12, SD10, BEM13, BKM13, HF19, KF18, LJB16, MBF^{+}17, ML11, PC19, TRS12]. Relative [BSM13b, BBM16, BM18b, CDR^{+}11, EHB18, KvZ12, LBB^{+}13b, Sau14, TMV16, WHC^{+}19, BNK14, CPG^{+}15, CAMW11, Car18, CN18, CC19, Cla14, CTWD15, DBH^{+}10, FTN^{+}14, HNM^{+}12, HFBS15, HHTT11, LWCC15, LHL^{+}15, MMJ^{+}12, MK17, NML12, NS16, RCLS18, RMB^{+}21, SD13, TOA12, YJC13, ZH10]. Release [PEU19, BBF^{+}18, BDB^{+}16, BMR^{+}19, BBC11, BBH^{+}12, CTSD10, CFPF^{+}15, DDK^{+}18, DHP^{+}11, EBB16, GHE^{+}11, GSØT15, GH14, GH15b, GRS^{+}15, GAB^{+}15, GPT^{+}19, HRT^{+}21, HAB^{+}10, HLSD21, HVD^{+}18, HLNH15a, HLNH15b, JDNP18, JHT^{+}10, JASP11, KCBS13, KGR^{+}15, LBCD18, LA13, LCH^{+}18, MCI^{+}13, MCDP18, MCB^{+}18, MSRB15, MWR18, MLC^{+}19a, PHB19a, PTC^{+}18, RMT^{+}18, RHC^{+}12, RBW14, RKH^{+}17, SGA^{+}18, SHA^{+}15, SWA19, SGT^{+}18, SLB^{+}12, SH10, TWCA15, TGE^{+}18, WSF^{+}18, WWG^{+}16, WHS15]. released [BCPP17, CBF^{+}14, LWSC11, RSQ18, Sau14, SFH^{+}15, SMJB11, SLB^{+}12]. releases [JP16, KNSPK13]. releasing [BBLD16]. relevance [BMOZ13, Sam14]. Relevant [FAAM15, BJZB12, CPH12, MNCPMG15, SWL^{+}15]. Reliability [KSAA13, Fra12, GGS^{+}19, KTNA15, LMPM12b, OPR12, Ste18, SDE^{+}18]. reliable [AL13b, HSB^{+}13, WWN^{+}14]. reliably [MRD19]. remaining [DJW^{+}18]. remediation [MBJ^{+}16]. remote [Kle13, NMLZ17, PAH^{+}13, PAM^{+}18, TBS13, vPB18]. remotely [AWSLS15, HSY16, JIM^{+}15, MHFK12, RGGdL13]. remotely-sensed [JIM^{+}15]. removal [DSW^{+}17, HPAA19, HdED^{+}13, JAJ^{+}16, PPBM13, SIBR^{+}19, vPBWA17]. removal-based [vPBWA17]. removals [CCC^{+}14b, Hoc12, LTV^{+}13, PGZ^{+}15, RC17, ZRH^{+}11]. rent [BAD19, DFAN12, GA19]. reoligotrophication [CLG^{+}14]. repellents [MK13]. replace [JHL^{+}19]. Replacing [Fra14, Tho19c, Tho21]. Reply [BKS^{+}15, HB15b, Han16b, LMPM12b, Sve15]. reported [Sam11]. reporting [CFH14, CFH15, FSA^{+}16]. reports [FCBS15]. representations [SMGH17]. Representing [GDC^{+}19]. represents [CSMO14]. Reprint [CFH15]. reproducing [HST10]. reproduction [DSJ13, GBEP16, HFBS11, MVMCC16, MLV13, SRT13, SMNdC^{+}19, ZMG^{+}14]. Reproductive [ASB^{+}16, AGSF19, CEW14, GMZ^{+}14, JWM^{+}12, OAT^{+}16, PG14, SSH14, SCL^{+}15, WFHG^{+}17, ABSI^{+}18, AT10, BMEGV^{+}19, BdGD^{+}16, Ber13, BBT^{+}14, BPMGP10, BdGN^{+}18, BMS15, Bro13b, BB13, CH14, CCM14, CBMLB13, COC14, CCC14c, DPSR10, DHH18, FAAM15, FAAM16, IIS10, KJS12, Ken14, KHGB14, KMSRW10, Kni12, KDPMSR10, Lam13, LBS^{+}18, LLH^{+}14, LCWC17, LMN^{+}12, MRD19, MCP^{+}17a, MQGK10, NVS^{+}18, OSO^{+}18, ORMG11, PRMKASR13, PHK13, PHK14, PEAC11, RB12, RM10b, SM15, SD10, SMCR13, SF19, SSJ^{+}13, TMM19, WBM17, WSG16, Wri13]. required [KOD^{+}15]. Res [ASHH15, BMF^{+}11, CP12, HAB^{+}18a, HLNH15a, IIT18, KPL^{+}09, LKD18, Mun12a, PGP^{+}16, PVKH18, SKB^{+}19, TC16a, TV14, Tho19a, Tho21, VRMF12, WGC^{+}21, WMB16]. rescaling [XGT^{+}17]. Research [Ano10-27, Cor17, HL16, BM19a, DMFR16, FRDC10, FLW^{+}19, GBB^{+}12, HB15a, HB15b, Kai19, KCJ19, KK17, LRM17, MSRM^{+}14, MRS15, MUO^{+}17, MML^{+}16, Pre17, PM13, Rob15, SP12, UARC11, Yee10, Cor17, Ano10b, Ano10c, Ano10d, Ano11b, Ano11c, Ano11d, Ano21a]. Reserve [BCKG13, Ber13, BAOP18, EMOLS^{+}15, FMA^{+}18, HHM^{+}13, MMBD13, TRH16, WUGG16, SK13, SCOC16]. reserves [ASB^{+}13, EP12]. Reservoir [SRG17, BRP^{+}16, BDB^{+}16, DLCK17, GHL^{+}11, JAJ^{+}16, JČK^{+}12, SMØ16, TWCA15, WTF16, GFK^{+}12]. reservoirs [AGS^{+}16, BLR^{+}16, BJKH12, DLCK15, GYCL19, KvZ12, KBČ^{+}16, LSS^{+}17, MLBTM16, PMK^{+}10, ŘJP^{+}12, SDJ^{+}15]. residence [BDB^{+}11]. residency [AASF16, ALR^{+}17, ÖLTA19]. resident [BA12b, KNJK^{+}10, KMJ^{+}12, MNCPMG15]. residing [KLB17]. residuals [TCAP11]. resight [VIP13]. resighting [BAY^{+}12]. Resilience [COS^{+}19, BMS15, HSC19, KvZ12, MAM^{+}13]. resilient [CH15]. resistance [TNKO19]. resolution [BCMS^{+}18, BDC^{+}17, CMI^{+}16, GLMK12, KHM^{+}17, KS17, PGW17, PH12, SBLB17]. resolve [FHK17]. Resolved [BGD11]. RESON [PGJ10]. resorption [CFB^{+}17]. Resource [BR18, CMB11b, PGW17, FMB^{+}19, GHUG^{+}15, GA19, LAB10, MCC^{+}17, MKH16, STAN^{+}17, VSLC^{+}19]. resources [AFA11, EGK15, EHBK18, FL11, GYCL19, HBC^{+}17, MSCP17, MS15b]. resplendens [SLTPC^{+}18]. responds [SVR16, UPS16]. Response [HM16a, HM16b, HAB^{+}18b, LRM17, NB19, PHK14, ASB^{+}13, BDHA11, BAFP18, FWG15, Han16b, JASP11, Kim15, KAiT^{+}13, LT16, MLC^{+}19a, MMG10, MGC^{+}11, NST^{+}17, OOTV12, PMG18, PRMKASR13, PBB^{+}18, QMGP10, QCZC^{+}19, SDJ^{+}15, ST16, WBD^{+}13, WHB^{+}16]. responses [AJOE17, BMR^{+}19, DSWB18, DHP^{+}13, KDdAA^{+}18, KBL^{+}15, LWY^{+}11, MMAH15, WWR18]. responsible [SMRS15]. restocked [PTC^{+}18]. restocking [WMIDS18]. restoration [YVJ12]. restore [CLG^{+}14]. Restoring [Ano13a, VFMS16]. restricted [BHG^{+}17]. restrictions [CAP^{+}10, KL16]. result [LJJ^{+}13]. resulting [GAR10, RLFJ11]. Results [MFIG16, TDH19, WMF13, Ech17, FH15, MVMSdS19, PBK^{+}19, RFS16]. retained [GLMK12, Sam11]. retardation [SG15]. Retention [MJL^{+}17, BN12, CN19, HSS18, MSCP17, PHRH16a, PHRH16b, RHC^{+}12, ROCC13, SGW16, SGT^{+}18, SP15a, ŠBB^{+}19, SMS17, SSC^{+}12, WBD^{+}13]. Retrospective [PDF^{+}18, HA12, ML17, SM14]. return [CM18, EHB18, GSTL15]. return-to-creel [CM18]. return-to-the-water [GSTL15]. Returning [NvZH^{+}17]. returns [HSC^{+}12]. Réunion [RBM^{+}16]. reveal [CHW^{+}18, DBG19, HSC19, JRA10, KVH15, MMDH18, SOMO17, VSLC^{+}19, WDL17, WHS^{+}18, WPS19, WW14]. revealed [BSD^{+}13, CLS^{+}19, CADG^{+}18, CN19, GBK^{+}18, HNY^{+}19, HMA18, IdRH10, LKL^{+}12, LZGL17, OPE^{+}17, ÖLTA19, PPLK18, SMO^{+}18b, WWLY15, ZDZY19]. Revealing [VLSP19]. reveals [AHH^{+}12, CGdGT^{+}17, CAC19, FLW^{+}19, dSFdSSS17, MSN^{+}12, PALC19, RFPG^{+}16, RHOCB^{+}18, RPP^{+}11, SAK10, SPM^{+}15a, TLCSG^{+}15, VZASM^{+}15]. reverse [SBL^{+}19]. Review [HB15a, HB15b, Hol11b, Mau11, Rob15, AF13, BRGS19, BK19, DDP^{+}16, DDH^{+}15a, FT16, GH15b, HBB^{+}19, HBC^{+}17, KRTG16, KEP^{+}19, KK17, MP14, MP13, MT19, Peñ18, PRCGD16, PBB^{+}15a, PSS15, Pun19a, Qui17, SEG^{+}15, SBLL17, SM14, VBdB^{+}18, Zis12]. Reviewers [Ano11a]. reviews [CDF^{+}13]. Revised [Kur12]. revisited [BTS^{+}19]. Revisiting [Fra17, RTSV15]. RFLP [CMSG18, GKMGO^{+}17, MRS15, SAH15]. RFMO [KEH16, KEP^{+}19]. rhina [KHG^{+}17]. Rhinopristiformes [Jab18]. Rhizoprionodon [dJSMROZP^{+}13]. Rhode [TMC19]. Rica [CEW14, VRW11]. Rican [BDB^{+}16, OPL17]. rich [FPD^{+}16, WFJ19]. Richardson [MLC^{+}19a, SLTPC^{+}18, VvHE^{+}11]. richness [CCM14, CK18]. ridge [LCA^{+}19]. ridge-scaled [LCA^{+}19]. rig [FBKE16]. Riga [OALS11]. right [KSH19, MVMdS^{+}17, TBS13]. Rights [AW17, EB15, Via19, VPSH^{+}17, HN16]. Rights-based [AW17]. rigid [LCHL11, QCZC^{+}19]. Rikhter [JG12]. Rikhter/Efanov [JG12]. ring [BMB^{+}18]. ringed [SL12]. ringens [CCC^{+}14a, CCHD19]. rings [MSV^{+}19]. Rio [BAOP18, MBS14, WJC17]. rise [HBMC12]. Risk [BKSF17, FPD^{+}16, FEGMTÁ17, GS12, GH14, HSS^{+}11, KMMV10, MVMdS^{+}17, PRP^{+}18, Pit14, RSE^{+}14, RGAP18, RAGP18, WDS^{+}11, ZSF11, ZHDS16]. risks [BFHT16, BCC15, HH16, LCH^{+}18]. Risso [MEMdU^{+}17, ŠBMG^{+}16]. River [DW19, LS19, LKNL16, MS15a, PKWP16, ST13, TGE^{+}18, ZD13, BDM^{+}18, BCP^{+}12, BSSH13, BTB^{+}19, CK18, DET^{+}16, GBR^{+}18, ICSR16, KKE^{+}12, LST^{+}10, MBF19a, RPH^{+}18, SMØ16, SO18, STCE12, SPH18, SGT^{+}14, DET^{+}16, GBR^{+}18, HKF^{+}12, JWZ^{+}19, LCST10, MZC^{+}18, RSQ18, Sim15, SCF^{+}17a, SGT^{+}14, WB16, WRMM18]. river-floodplain [ICSR16]. river/hatchery [SO18]. riverine [BRP^{+}16]. rivers [BA12b, BTB^{+}19, GSØT15, KNJK^{+}10, KNRJ^{+}15, KAiT^{+}13, MLT17, PJM14, ZW18, KAGO^{+}10]. RNA [YA18]. RNA/DNA [YA18]. road [BAM16]. Robertson [HB15b]. Robust [Peñ18, MQGK10, Nak17b, vPBWA17]. Robustness [KHV^{+}12, CV14, FHK18, FSG^{+}19, HH14]. rochei [MEMdU^{+}17]. Rock [LSM^{+}19, AOC11, BEM13, BBB^{+}16, BM18a, CMMF13, CGG21, EHGG16, FPM17, FLW^{+}19, GHP^{+}15, HN16, LAB13, LBB^{+}13a, LPHH11, LGH^{+}21, LMLBC13, PML^{+}12, PTW^{+}13, WWR18, vPHG11]. Rockfish [HSY16, ASD^{+}18, BRS19, FRDC10, HRB14, HWS10, HFBS11, Hoc12, KMH^{+}12, LBS^{+}18, LW12, PKO10, RHB^{+}17, RLFJ11, RWRT15, TLLZ19]. rockfishes [BBA^{+}19, CRW^{+}19, DBMR17, TDH19, ZTC^{+}18]. Rocky [GHL^{+}11, BH11, PBC^{+}13, SMJB11]. rocky-reef [BH11]. rod [KCBS13]. Rodrigues [HER13]. Rodriguez [Cor17]. Rodriguez-Marin [Cor17]. Role [Ano13a, ABD^{+}19, CAGTS^{+}10, CAGVB14, DPL^{+}16, EB15, KCJ19, LBF^{+}16, LHLR17, MMdlP16, MS15b, Sal18, SP15a, SJA17, STCE12, SL12, WGDB13]. roles [PH12]. roller [CSMJB11]. roller-frame [CSMJB11]. rollerframe [SBLM14]. Rope [HB11, KW10]. rosaceus [RLFJ11]. rose [VFMS16]. Ross [MDPH17]. rostrum [BPRB^{+}14]. rosy [RLFJ11]. rotation [BWK^{+}13, FKM11]. Rotorua [KHQH15]. roughy [Ken14, KSKGR15, PHK13, PHK14, VRS13]. round [GBF11b, JBB^{+}18, SMD^{+}16]. roundfish [SHM^{+}16]. roundscale [BSD^{+}13]. routes [MLT18, YAT^{+}19]. Royal [RLM10]. rRNA [TdMWC16]. RTI [KRC14, KRB^{+}15]. ruberrimus [HRB14, Hoc12, RHB^{+}17]. rubra [HHTT11]. Ruditapes [CMM^{+}11, CLF^{+}17, RvRW^{+}14]. Ruiz [Cor17]. Rules [PKRC15, DB12, FHK18, JAB15, PML^{+}12, WP17, ZCZ^{+}13]. run [SO18]. runoff [LS19]. Ruppell [SJJ^{+}11]. russelli [SJJ^{+}11]. Russia [SRG17]. Russian [PLA16]. Rutilus [MLBTM16].

S [CCM14, MFRM19, MGP^{+}18, dADSRG10]. S. [CRB^{+}18, HRB14]. s.l [RSV^{+}17]. Sabah [TTS11]. sablefish [GMS^{+}17, HCHA12, HPP18, HKB14, KSH19, Rod19, JC18]. Saccharina [SMS^{+}19]. sacks [RHC^{+}12]. saddled [CMPR^{+}16]. SAFE [MBK19, SKR^{+}12, SDV^{+}16, ZHDS16]. safer [SKR^{+}12]. sagax [GCDAS^{+}10, GRGGCACG11, NGvdL^{+}13, WMvdL10, WRH^{+}15]. saggital [AB11]. sagitta [dSSdAdAA17]. sagittal [BPRB^{+}14, MSV^{+}15]. sagittatus [RLVS14]. Sagmariasus [BBM17, BM18a, LBB^{+}13a]. sailfish [LBMC15, TCS^{+}15]. Saint [ACRM12]. Saint-Hilaire [ACRM12]. saira [HMGH12, MSV^{+}19]. saithe [BHL12, EDG16, RR12, SPM^{+}19]. salar [FMF^{+}17, FBF^{+}18, GSØT15, HRT^{+}21, HKAA18, JWM^{+}18, KNRJ^{+}15, MFMM16, dEWB^{+}15]. sale [PTG^{+}17, TdMWC16]. Salilota [ABL21]. salinity [MDG17, NNP17]. Salish [ASD^{+}18]. Salmi [HL16]. Salmo [AVB^{+}16, BGL^{+}13, CFF^{+}16, Cur18, FMF^{+}17, FBF^{+}18, GSØT15, HRT^{+}21, HKAA18, JWM^{+}18, KNJK^{+}10, KNRJ^{+}15, MFMM16, PEU19, RSGR19, SBAK17, SDWG18, SGT^{+}18, WB16, dEWB^{+}15]. salmoides [AAN^{+}12a, SGA^{+}18, TWCA15]. Salmon [LKNL16, WAB^{+}16, BHL12, BV14, BHT^{+}19, BRGS19, BHC^{+}16, BDB^{+}11, BWK^{+}13, CQ14, CBD^{+}14, CHW^{+}18, CDB^{+}14, DMC^{+}12, DET^{+}16, DDP15, DHP^{+}11, DHP^{+}13, ER18, FMF^{+}17, FBF^{+}18, GHE^{+}11, GSØT15, GWB^{+}17, HRT^{+}21, HKAA18, HDS^{+}18, HF15b, HSC^{+}12, HKF^{+}12, JWM^{+}18, JHT^{+}10, JBMK13, KNSJJ11, KNSSL11, KNSPK13, KFBL13, KKE^{+}12, LS10, LDM^{+}15, LW12, LW19a, MMMPP19, MBF19a, MFMM16, MBG^{+}10, MCJD^{+}17, NEJ^{+}18, OLS15, PAM^{+}18, RBG^{+}19a, SCC^{+}15, SO18, ST13, SCF^{+}17a, SCF17b, TQ17, UKK^{+}12, WBD^{+}13, dEWB^{+}15]. salmonids [AAN^{+}12b, BPC^{+}12, DPAB10, GLB^{+}19, RIB^{+}10, RSQ18, SGA17, SGB^{+}10, SL12]. salmonis [BGL^{+}13]. Salvelinus [CLG^{+}14, GHL^{+}11, LSWR17, LRY^{+}15]. SAM [AAS19, APHL13, NB19]. Sambo [Ber13]. same [RSAB11, SHT^{+}13]. Samoa [OTST11]. Sample [BV14, Chi10, HFPL16, Mau11, TJMT17, TBSE17]. sampler [PSL^{+}19]. samplers [MBS^{+}15]. samples [Fey18, GHA13, GRO^{+}18]. Sampling [BDB^{+}11, PH12, RH13, RJKG12, BMC10, CRS^{+}19, CFBS16, Chi10, Chi11, EHB^{+}21, FCBS15, GJRR14, Gra16, GAR10, Har19, HHH15, HNM^{+}12, HdED^{+}13, JČK^{+}12, JBB^{+}18, JČK^{+}21, KGH^{+}12, LNC^{+}15a, LNC^{+}15b, LST^{+}10, LCRW19, MHFK12, PWB^{+}16, PTBT17, PM11, PMP^{+}11, ŘJP^{+}12, SM15, Seu21, SMI17, SBLM14, SGT^{+}14, TBS13, YJSR12, ZW18, ZG14]. San [OGCBMR15, CUL^{+}16, CAGVB14]. sand [AK11, BBC11, KCBS13, ML10, MJL14]. sandbar [LMM^{+}17, MSRB15]. sandeel [JvDB^{+}17, WRG^{+}18]. Sander [BNSS19, BJZB12, DLCK15, GKKK15, Han16b, HPHR14, HL16, HM16a, HM16b, JWM^{+}12, JČK^{+}12, KBSK15, KVH15, LSŠ^{+}16b, LLM^{+}19, SAR^{+}15, VH12]. sandfish [RGK19]. sands [PTTC18]. sanguinolentus [RM10a]. santolla [MFRM19, SSFGL19, SLT^{+}19]. São [PG14]. Sap [TBC^{+}19]. sapidissima [BSSH13]. sapidus [MCC^{+}17]. Sarawak [IAJ^{+}11]. Sardina [CHCS14, MSS^{+}16]. Sardine [BAOP18, CJCN21, CCC14c, DRGGVL11, FE10, FPC^{+}13, GCDAS^{+}10, GRGGCACG11, JSYS15, JD14, MSS^{+}16, MSD^{+}16, NGvdL^{+}13, SKY^{+}17, WMvdL10, WRH^{+}15, ZD19]. sardinella [BFL^{+}14, CBS17, PGAS15, HSS18, SMS17]. Sardinia [ASB^{+}16, MPB^{+}17, SZR^{+}11]. Sardinops [GCDAS^{+}10, GRGGCACG11, NGvdL^{+}13, WMvdL10, WRH^{+}15]. Sargassum [PKO10]. sargus [HVJP^{+}11a, HVJP^{+}11b, HVJP^{+}11a, HVJP^{+}11b, PKAL13]. satellite [AGB19, BNU^{+}10, CAA^{+}14, CAA^{+}15, ELRH17, KDPT18, LMM^{+}17, RTML17]. satellite-recorded [BNU^{+}10]. satellite-tagged [AGB19]. satisfaction [GFJ19]. saturation [PMP^{+}11]. Saudi [AMAM^{+}17, RLM^{+}17, RU15, SB15]. saurus [AB11, AB12, AB11, AB12]. saury [AB11, AB12, HMGH12, MSV^{+}19]. saving [BGL15]. saxatilis [HKM^{+}17]. scabbardfish [PGS10]. scabra [RGK19]. scale [ACSAS^{+}17, ASMP^{+}10, ASMD^{+}18, AFOB^{+}19, ADH^{+}13, ASSL11, ADSV17, BB18, BMCS19, BOCÁ^{+}18, BHI14, BMOZ13, BHCJ12, BSSH13, BPG^{+}17, BMC10, CEJI11, CMB11b, CAC15, CRAV13, CN19, Cro15, DQMV19, EFWC^{+}19, EFW^{+}11, FCLH19, GHP^{+}15, GOD19, dMGPR^{+}18, GBD^{+}15, Gra16, GFM^{+}12, GBK^{+}13, HN16, HLNH15a, HLNH15b, HL10, HAB^{+}17b, IO11, Ibá14, Ibá15, IHFAH17, INSN14, JWC^{+}13, KPM^{+}19, KMW^{+}11, KWCT^{+}18, Kos15, KBEB13, KAiT^{+}13, LOdH16, LBH15, LPC^{+}14, LNL^{+}15, LGH^{+}21, LCH^{+}18, MMNCR^{+}13, MAS17, MJL14, MFRM19, MJC^{+}14, NCP15, NvZH^{+}17, NWO^{+}15, OTST11, OtHH^{+}10, OOTV12, ÖLTA19, PTG^{+}17, PTSC^{+}16, PH12, PVMG19, PJML15, PNAL16, RLC^{+}17, RLC^{+}15, RRO^{+}19, RDP^{+}17, RBG^{+}19b, RSGR19, SKY^{+}17, Sau14, SH13a, SFBV17, TLLZ19, TTS11, TD19, TRH16, VPSH^{+}17, VAB^{+}14, WHK13, WJC17, YvZG^{+}17, Zie12]. scaled [CMPR^{+}16, LCA^{+}19]. Scalefish [GSP17a, GSP17b, SB16, Zie12]. scales [BJZB12, CP11, CP12, DDP15, DBLB16, FCL16, FMF^{+}17, FBF^{+}18, IO11, Ibá14, KS17, VF13, WW14]. scaling [LPMM14]. scallop [AHMS10, BH11, HBT16a, KRM^{+}18, LBH15, LHSS18, MHC^{+}18, Mur11, RDM^{+}15]. scallops [HJT13, TRH16]. scamp [BB18]. scampi [MTC^{+}17]. scapulare [CMS14]. scaring [MGR13]. scenario [GD17]. scenarios [BNAE10, CCGD19, MGC^{+}11]. Schaefer [XdA19]. schedules [HSH15]. scheme [MUO^{+}17, SB16]. schemes [MFKS17, TTT16]. schlegelli [CLS^{+}11]. schmitti [CBC^{+}10, PGM^{+}10, WJC17]. school [HDW^{+}15, RKH^{+}17, dKMS15]. schooling [CLP^{+}19]. schools [CDM^{+}19, GSDD11, PB14, PGJ10, RCBPK11, TXZ^{+}17, TGDB16]. Sciaena [PBC^{+}13]. sciaenid [FWIB14]. Sciaenidae [CLS^{+}19, MNGPMdMG10, SNOD^{+}13]. Sciaenops [BHB^{+}10, CBF^{+}14, HCWS19]. Scieanidae [MKM^{+}16]. Science [BW17, PS13, BHP^{+}19, FTWC18, GGM^{+}18, HHM^{+}13, HSG15, MSD^{+}16, MEM15, SKB^{+}18, SKB^{+}19, UARC11, VF13, WSF^{+}18]. Scientific [MML^{+}16, MGLT16, MDC^{+}16, Nak17a, OLM^{+}19, PSE15]. Scientists [CPMNS10, AF13]. sclerochronological [MBF^{+}17]. Scomber [DTSG^{+}16, HBP^{+}18, HV10, JB15, JPO^{+}19, LCM^{+}18, TPM15, YGZ^{+}18]. Scomberesox [AB11, AB12]. Scomberomorus [CLB^{+}10, EPHF18, GJSR10, NML12, RU15, SRMP^{+}18, ZLT^{+}16]. Scombridae [CFB^{+}17]. scombrus [DTSG^{+}16, HBP^{+}18, HV10, JB15, JPO^{+}19, LCM^{+}18, TPM15]. scores [CNWM12, PSU^{+}16]. Scoring [RAGP18]. Scorpaena [MLV13, ÖLTA19]. Scorpaenichthys [MNW12]. scorpaenids [MLV13]. Scorpaeniformes [MSV^{+}15]. Scorpis [CPH12]. Scotia [DBT15, MS12]. Scotian [AMH19, OS12, KC11]. Scotland [BSF^{+}19, GIW16, HSH16, NBF^{+}19]. Scottish [DBK^{+}10, KOF11, QWMD13]. scrofa [ÖLTA19]. Scyliorhinidae [BMRC14]. Scyliorhinus [BMRC14]. Scylla [BMH18, IAJ^{+}11, LWY^{+}11, MMAH15, RMGK18]. Scyllaridae [dADSRG10]. Scyllarus [AGSF19]. scyphomedusa [RT14]. Scyphozoa [MDL19]. SE [GBPC10, PG14, TEFAR12]. Sea [DV11, GBPC10, GPBA13, Han16b, HL16, HM16a, HM16b, KTL11, KJKK15, KPL^{+}09, LSJD11, LRM^{+}10, MMF^{+}17, MCP^{+}17a, MGB10, OALS11, PPLK18, RDP^{+}17, SAR^{+}15, SMJB11, UJLC^{+}19, UGGR^{+}17, XDL^{+}19, ZKSF17, AHMS10, ACRM12, BGBCM19, BMB^{+}18, BGL^{+}13, BAT12, BDM^{+}14, BR18, BKS18, CFBS16, CMPR^{+}16, CUL^{+}16, CPOH14, CHA15, CSFCA15, FPSS15, GAR19, GHM16, GWVH^{+}16, GMUD11, HHD^{+}11, HB15a, HB15b, HJT13, HFCPSEM15, Hua15, JEJS17, JWC^{+}13, JTHT14, KGR^{+}15, LMB^{+}11, LKL^{+}12, LTV^{+}13, LL14a, LCRW19, LCWC17, MBS14, MD16, MPF^{+}17, MMD^{+}11, MHC^{+}18, Mur11, NWO^{+}15, OLS15, PWSP12, PPA^{+}15, PGS^{+}11, PDBM12, PGL^{+}13, PNAL16, RUMBA15, Rob15, RSGR19, SMO^{+}18a, SSH14, SAK^{+}17b, TNKO19, TEGM^{+}11, VRS13, VVL18, WB16, WGM^{+}11, WMF13, ZCLB11, ASL^{+}12, APCF17, AVU^{+}10, ASD^{+}18, ALUFJ^{+}12, BMEGV^{+}19, BHI14, BGF^{+}15]. Sea [BVN^{+}10, BRC^{+}10, Ben18, BHM^{+}18b, BHM^{+}18a, Bri10, BLHG17, BKLB14, CMPR^{+}16, CPD^{+}19, CMM^{+}11, CJP13, CSP^{+}18a, CKVV16, COC14, DDB^{+}14, DCP^{+}16, DH11, DÖA^{+}10, EÖFPH17, Eth15, FMB^{+}19, FF10, FTN^{+}14, GGS^{+}19, GIW16, GPSF10, GSL14, GPS15, GSH^{+}15, GGB^{+}18, HB15b, HNW^{+}18, Han16a, HBMC12, HPHR14, HSLB19, HPNA16, HVJP^{+}11a, HVJP^{+}11b, HMA^{+}16, Jab18, JBD^{+}18, JS15, KNSJJ11, KNSSL11, KNRJ^{+}15, KBG^{+}14, KDPT18, KCvOE13, Koo12, KM15, KPL^{+}07, LSŠ^{+}16b, LE14, LHS^{+}18a, LSJD11, MPB^{+}17, MABA12, MMM^{+}10, MAH16, MBvDN14, MLBGV16, MDR^{+}19, MDPH17, MEMdU^{+}17, MRT17, MRT18, Mur19, NKW^{+}19, NNP17, OtHH^{+}10, OSO^{+}18, OAT^{+}16, PBLT14, PSM^{+}18, PWSP12, PGAS15, PPG^{+}11, PBC^{+}16, Pop19, PLA16, PGVV16, RDV11, RBV^{+}13, RTSV15, RSV^{+}17, RBR^{+}15, RLVS14, SMB^{+}17, SVV^{+}14, SD10, SFR^{+}13, SCV^{+}12]. Sea [Son18, SLZG18, SB15, SGGH10, SFL^{+}14, ST16, SPM^{+}10, SL12, SP12, SP15b, TVCT12, TOA12, TKF11, VKM^{+}17, WMC^{+}16, WLGS19, WSG16, WOPB11, WHF^{+}14, YVJ12, YSS^{+}11, YGZ^{+}18, ZRH^{+}11, ZzCbC^{+}15]. sea-state [PDBM12]. sea-surface [PGS^{+}11, TNKO19]. seabass [HLSD21, WWLY15, RCHSN^{+}15]. SeaBat [PGJ10]. seabird [BBHF10, GMS^{+}17, LJB16, MGR13, MGR14, PAR^{+}12, STAN^{+}17]. seabirds [Hua15, SMK^{+}13]. seabream [ALUFJ^{+}12, CGECSEH10, HVJP^{+}11a, HVJP^{+}11b, KW10, NVS^{+}18]. seafloor [LWH19]. seafood [BPS^{+}19, CGdGT^{+}17, HCZ^{+}16, Ibá15]. seagrass [SBLM14]. seahorses [LQL^{+}15, SFBV17]. Seal [CGR^{+}15, CGR^{+}16, KLH^{+}15, OS12]. seals [KFBL13, LS10, RDP^{+}17, SL12]. seamounts [WUGG16]. Searching [HMvHN13, SLPT17]. seas [BCN^{+}13, HMMP18, MMD^{+}11, PPdR^{+}10, ŠBMG^{+}16, EP17, HBC^{+}17, KDPT18, PTF^{+}18]. seascape [SF13]. season [BdGN^{+}18, DPSRM10, DHFSS18, RM10a, SD10, SNK19, SRG17]. Seasonal [CMS^{+}15, CRW^{+}19, CVS11, EMP^{+}10, EHH^{+}15, HRUSR13, ICSR16, LMN^{+}12, MMGEMGS10, MNGPM^{+}12, CAR19, DHH18, HHSS17, LHSS18, MAH16, Ogl17, OMI^{+}14, SKFL15, SSP^{+}15, SMPB17, WB16]. Seasonality [HCGS19, APDJSU11, CAC19, TMM19]. seasons [BBT^{+}14, MJL14, NMMBCV^{+}10]. seatrout [CSOV14, EHB18, FF11]. Seattle [Ano13a]. seawater [KGR^{+}15, SMS^{+}19]. Sebastes [BBA^{+}19, CEB^{+}19, CRB^{+}18, DBMR17, HRB14, HWS10, Hoc12, KMH^{+}12, LHS^{+}16, LBS^{+}18, LW12, PKO10, PHRH16a, RHB^{+}17, RTML17, RLFJ11, VPC^{+}14]. Sebastian [OGSNH^{+}17]. Sebastião [PG14]. Sebasticus [ZTC^{+}18]. Sebastolobus [Ech17]. second [HM16b, RW19]. secondary [FdMCB^{+}19]. section [GSH^{+}15]. sectioned [SVN^{+}13, SHC14]. sections [BPRB^{+}14, CHW^{+}10, KMQ^{+}10, KOF11, ZYSZ18]. Sector [MFRM19, BAD19, CAAB^{+}18, FWJ^{+}16, INY11, KKP11, ZSF11]. sectors [JSYS15]. security [RFS16]. sedation [BNA^{+}19]. seed [LCH^{+}18]. Seeking [LG13]. seen [ST16]. segregation [ÁCDM^{+}14, APD^{+}18, HCF^{+}17, NWS^{+}11]. seine [BAY^{+}12, CHW^{+}18, CK18, DLCO^{+}19, GQ16, HER13, HHH15, HV10, KBCK16, KZ11, LOdH16, LLP18, LCMF^{+}16, LvVtB^{+}15, NFK^{+}17, NSM^{+}19, TXZ^{+}17, TXH^{+}19, XLCMMV19, ZXHK15]. seine-net [BAY^{+}12, HER13]. seiners [DTSG^{+}16, LMSM14, TIGCDL14]. seines [EBB16]. seining [DHP^{+}11, HLS^{+}16, TPM15]. seismic [MCMK18]. Selar [PGAS15]. SELECT [PMB^{+}11]. selected [MMK10]. Selecting [HPHF15]. Selection [LPM^{+}13, LOC^{+}11, AVB^{+}16, BAL^{+}19, BHM^{+}18a, BSM15a, BSM15b, BM18a, BSM18, CMI^{+}16, CBO^{+}15, FHM10, GHUG^{+}15, GHUG^{+}16, GBK^{+}13, HWS^{+}15, HLS^{+}16, HKK18, Ing11, KHMF11, LHS^{+}16, LHS^{+}18b, LHS^{+}18c, LYWZ15, LBZE17, LHF16, LCHL11, LCL12, LCLM15, LCF15, MH11, OHSG19, PHFW14, RGG14, Sam14, SHGL21, WB15, WHMS11, WHF^{+}14]. Selective [BHCS15, BGP^{+}16, HPS16, KHM^{+}10, WUGG16, BRH^{+}16, CGC13, GKKK15, GQ16, GAR10, LW16, MHFH17, OLK^{+}14, OKPL11, SLP^{+}15, SBH^{+}16, TM13]. selectivities [Cla14]. Selectivity [KWR15, KLWG17, MCVS14, PHRH16b, SJR^{+}11, WW11b, Ara15, AOC11, BHB^{+}10, BHL^{+}16, BZ12, BMB10, BRB^{+}14, CDR16, CA10, CEB^{+}19, CV14, CGECSEH10, DBK^{+}10, EAM^{+}14, FHMK11, FS19, GPSF10, GSH^{+}15, HFD^{+}13, HB12, Hen11, HWM^{+}13, HSL^{+}13, HSLB19, HH14, HFPH14, IOT14, IEL17, JC18, KPM^{+}19, KHKM15, LNC^{+}15a, LNC^{+}15b, Leg14, MTH10, MS14, MH11, NB14, NFK^{+}17, OKB^{+}16, ÖEG^{+}15, PMB^{+}11, PHRH16a, PHFW14, ŘJP^{+}12, SL10, SL11a, SWC14, SHGL21, SSHM16, SM14, SM17, Sve15, TLLZ19, TT14, THA^{+}14, THG^{+}16, VKM^{+}17, WMP^{+}14, WSMS14, WMdS14]. Self [CMHP14, PKRC13, RCF^{+}16, CLG^{+}14, Nee15, Sam11]. Self-imposed [CMHP14]. Self-Organizing [RCF^{+}16]. self-reported [Sam11]. Self-starting [PKRC13]. self-sustaining [CLG^{+}14]. self-testing [Nee15]. seller [FGRIT^{+}12]. semi [BHC10, QQG14, TT14]. semi-automated [QQG14]. semi-parametric [TT14]. semi-quantitative [BHC10]. semifasciatus [CLB^{+}10]. semisulcatus [RLM^{+}17]. Senegal [BKS^{+}15, CFF^{+}15, BKS^{+}14]. senegalensis [RFPE^{+}12, RFPG^{+}16, SNOD^{+}13]. senescence [BV14, PGG10]. sense [HH14]. sensed [JIM^{+}15, RGGdL13]. sensing [Kle13, NMLZ17, PAH^{+}13]. sensitive [BNA^{+}19]. Sensitivity [ASH^{+}17, HVJP^{+}11b, ER18, OMI^{+}14]. sensitization [RGAP18]. separability [CCC^{+}19]. separate [SAK10]. separation [BHPC14, DBA18, MHK^{+}17, MKHK19]. Sepia [AAOE13, ARF^{+}17, GHUG^{+}16, LXCC19, PVMG19]. September [Ano11q, Ano12r, Ano13a, Ano13k, Ano14l, Ano15r, Ano16q, Ano17v, Ano18p, Ano21c]. Sequence [RCF^{+}17]. sequences [LSR16, PGAS15, TdMWC16]. sequencing [KK17, SRMP^{+}18, VZASM^{+}15]. sequential [MAA^{+}16, Pun18, SCF17b]. Serbian [MLBTM16]. Serial [CNJ11]. series [AVB^{+}16, FSBS11, HPHF15, LCMdSM13, LRM^{+}10, Lyn14, PPG^{+}11]. Seriola [JDNP18, ŠBMG^{+}16]. serranids [AASF16]. serrata [BMH18, MMAH15, RMGK18]. Session [Ano13a]. set [ASMD^{+}18, CMPP11, CGR^{+}15, FB19, HMA18, LWCC15, LSS^{+}17, LL14b, PGP^{+}15, PGP^{+}16, SWA19, VVL18]. set-lines [PGP^{+}15, PGP^{+}16]. set-longline [CMPP11]. set-net [LL14b]. setiferus [BFM14]. sets [GQ16, MP17]. Setting [GWVH^{+}16, UARC11, CPW^{+}14, CC19, FWR10, MWR18, MGR13, NBS15, TXZ^{+}17, WFJ19]. settings [Nak17b]. settlement [PKO10]. settler [GKL12]. Sex [APDJSU11, ŠBB^{+}19, APAFMN10, CCA16a, Cer14, EP12, GLA19, MJC^{+}14, NWS^{+}11, OMI^{+}14, SMP18, SHR18b, WDNB^{+}19, dCJG^{+}15]. sex-related [APAFMN10]. Sex-specific [ŠBB^{+}19, OMI^{+}14, SMP18]. Sexual [HCF^{+}17, HW11, MQCS^{+}14, APDJSU11, AMH19, ATG11, CH14, Cer14, FdMCB^{+}19, KTT13, RM10a]. Seychelles [RSG^{+}11]. Shad [JSB19a, BSSH13, BNA^{+}19, BPV^{+}17, JSB19b, VWHB14]. shaking [Kim15]. shallow [EAAA^{+}18, FHS^{+}16, GKKK15, GFK^{+}12, JKS^{+}15, JKK^{+}16, JRD^{+}13, KBT^{+}19, WMPR17, ZBB17]. shallow-water [JKS^{+}15, JKK^{+}16, WMPR17]. Shandong [TLLZ19]. shape [AB11, BCN^{+}13, DBA18, FWG11, HMA^{+}16, IO11, Ibá14, Ibá15, KY19a, KBG^{+}14, KBOM14, MHK^{+}17, MTV17, MH11, MFVPC19, RLJ^{+}12, SVV^{+}14, SSD^{+}19]. shaped [JČK^{+}12, LHF16, TLLZ19]. shapes [IHFAH17]. share [BT16, PSU^{+}16]. Shared [Han18b, EP17, KHS12, LAB10]. Shark [ASHH12, ASHH15, Jab18, JCW12, LAIVA^{+}19, ARMLC19, BLHS11, BB10, BMdS^{+}15, BHCS15, BB13, CNS11, CAM^{+}14, GTDJdL^{+}13, GWWW13, GPT^{+}19, HFD^{+}13, HKB^{+}18, HSS^{+}16, HIH19, HDW^{+}15, HSM^{+}12, HAB^{+}17b, HWS^{+}12, KF18, KWS^{+}15, KHS12, LMM^{+}17, MK13, MQV10, MQCS^{+}14, Nij15, PGLG11, PGM^{+}10, PJML15, PBB^{+}15b, RKH^{+}17, SRSNC17, SC19, SMGH17, SCT^{+}17, dJSMROZP^{+}13, TOA12, TGM13, TGMM13, TFB^{+}12, TTF^{+}13, VdSS^{+}14, VZASM^{+}15, WWR18, WJC17]. shark-control [GPT^{+}19]. shark-excluding [BHCS15]. Shark-like [Jab18]. sharks [CRG^{+}18, CIS13, EBB16, FBY^{+}12, FWR10, FEGMTÁ17, HAB^{+}10, HCH^{+}19, HSB^{+}13, Kai19, LJJ^{+}13, MSRB15, MC10, MJC^{+}14, OGSNH^{+}17, PGP^{+}15, PGP^{+}16, RRPG18, RPK11, RHW^{+}16, SPM^{+}15a, SHA^{+}15, SWA19, WWR18]. sharpnose [dJSMROZP^{+}13]. sharpspine [JLLL11]. shedding [COS15, GH15a, PALC19, VIP13]. Sheepshead [ALB^{+}18]. Shelf [ABL21, CCM14, KC11, PPdR^{+}10, RLSPdR14, SWD^{+}13, AGB15, BBT^{+}14, BMR^{+}18, GPdR11, HNM^{+}12, HSH16, LAB10, MS15b, PT19, Pit14, TEFAR12, AMH19, MM11, OS12]. shelf/slope [AGB15]. shell [KRB17, LBF^{+}16, MTV17]. shell-structure [KRB17]. shellfishery [GOD19]. sheltering [KT18a]. Shetland [CEAP18, MBOAA12]. shield [GRB14]. shift [CJWT^{+}19, LSM^{+}19, UK14, Way13]. shifting [LS19, UP16, YC18]. shifts [BLP^{+}18, KOD^{+}15, PPLK18, SWG18, TKF11]. shipping [VBR^{+}18a]. shoals [TPM15]. shooting [LLP18, ZXHK15]. shore [FL11, KSL^{+}16, MCDP18, PPBM13]. shore-based [KSL^{+}16, PPBM13]. Short [BSM12, CMS14, CKVV16, DDB^{+}14, DCP^{+}16, MK17, MCI^{+}13, VGE11, AVU^{+}10, BHB^{+}10, BSM15a, CAGTS^{+}10, CMB11a, GOD19, HdSB^{+}14, HMGH12, KC18, MRD19, OLK^{+}14, ÖEG^{+}15, SP10b, SRB^{+}18, UASM12]. short-fin [CMB11a]. short-finned [SP10b, SRB^{+}18]. short-lived [HMGH12, KC18, MRD19]. Short-term [CMS14, CKVV16, DDB^{+}14, DCP^{+}16, MCI^{+}13, VGE11, AVU^{+}10, BHB^{+}10, CAGTS^{+}10, GOD19, OLK^{+}14]. shortbill [AGB19]. shortcomings [EL18]. Shortfin [BNM11, BB13, CCC17, RHW^{+}16]. shortspine [Ech17]. should [Pla18]. Show [Bai10, FBB^{+}17]. shows [Lyn14]. Shrimp [KPM^{+}19, LHS^{+}18b, BFM14, CCA16a, CCA16b, CSMJB11, DCP^{+}16, FMMW13, GJS^{+}13, HJLW11, HLJ15, HB11, HB12, HB13, HSLB19, JHL^{+}19, KAGO^{+}10, LHS^{+}18a, MCH11, NWL^{+}14, PF15b, RLM^{+}17, RRO^{+}19, RJ16, SBD^{+}12, TTMC16, TTT16, VFMS16, VLSP19, VRW11, WDBH16]. shrimps [GFM^{+}12, HB12, MMO^{+}13]. Siberia [SRG17]. Siberian [CGKG^{+}17]. sibling [KTW^{+}13]. sic [MBJ^{+}16]. Sicily [FPC^{+}13]. Side [SDV^{+}16, BSM12, BSM18, MLT18, vPKF^{+}19]. Side-effects [SDV^{+}16]. sides [RC17]. Sierra [Bai10]. sieve [JHL^{+}19]. signal [HZT^{+}19]. Signatures [LZGL17, APD^{+}18, CPG^{+}11, CHCS14, CSOV14, HBG^{+}19, LKNL16, MNGPM^{+}12, SSD^{+}19]. significantly [KS12]. siliqua [COC14]. silky [GTDJdL^{+}13, HCH^{+}19]. Sillaginodes [GCRV10]. Sillago [KTW^{+}13]. Siluriformes [dPHS^{+}12]. Siluriforms [VF10]. Silver [GBA^{+}18, BRJ^{+}16, CPS^{+}12, LLM12, MKM^{+}16, VBR^{+}18a, dCJG^{+}15]. silverside [AMV14, PEAC11]. silvertip [SCT^{+}17]. similar [SO18]. similarly [JPM11, TZW16]. simple [Eng18, MDH15, RJ16, MBK19]. simplex [MGP^{+}18]. simplified [CTWD15]. simplify [Tho19c, Tho21]. SIMRAD [PGJ10]. Simulated [HTB15, BPC^{+}12, FBF^{+}18, FSG^{+}19, GHE^{+}11, KRB^{+}15, Leg14, LLM^{+}19, RMT^{+}18, SGB^{+}10, WTF16]. Simulating [CRAP^{+}11]. Simulation [BZ12, CCC14c, SHM12, DHP^{+}13, EHH^{+}11, FHM10, HLS^{+}16, HVJP^{+}11a, HVJP^{+}11b, LMPM11, LJB16, LCRW19, NWO^{+}15, PLPN12, Pun19a, SLXZ15, SYJ^{+}11, WT16, YJSR12, YJC13]. Simulation-based [BZ12, FHM10, HLS^{+}16, LCRW19]. simulations [HVJP^{+}11a, RBT14, Zha13]. simulator [FGS19]. Simultaneous [LCMdSM13, GBK^{+}18]. simultaneously [KT18b, WHF^{+}14]. since [MPR18]. single [BTBP18, BSM13a, FSPWG08, JWM^{+}18, KRC14, MDS^{+}18, Mun12a, Mun12b, RPH^{+}18, SMK^{+}13, SYJ^{+}11, TB10, YAH^{+}13]. single- [BSM13a, TB10]. single-boat [SYJ^{+}11]. single-nucleotide [JWM^{+}18]. single-species [BTBP18, FSPWG08, KRC14, Mun12a, Mun12b]. Sink [Pul17, SBC^{+}18]. sinking [TXZ^{+}17, TXH^{+}19]. site [CB17, FF10, HGG13, ML10, MNW12, PAM^{+}18, SBT^{+}18, WPS19]. sites [APDI13, ALBŠB^{+}15]. situ [BNM11, KT18a, LNC^{+}15a, LNC^{+}15b, PSL^{+}19, RSEERRSC15, DMK^{+}17, HRAK15, OCL^{+}18, PB14, RT14, ZKSF17]. situation [GDB^{+}16, MMG10]. situations [Cop13, LKR11]. Six [LEO^{+}11, DLCK15]. sixteenth [BR15]. sixteenth-century [BR15]. sixty [MP14]. Size [AT10, AMH19, BLW^{+}18, BHM^{+}18a, FSPWG08, FWG15, HFD^{+}13, HWS^{+}15, HKK18, HVD^{+}18, JC18, KHMF11, LHS^{+}16, LHS^{+}18c, LCHL11, LCF15, MLM15, OKPL11, ÖEG^{+}15, PMB^{+}11, RM10a, SBH^{+}16, SKK^{+}13, SMNdC^{+}19, Sve14, TTF^{+}13, APDJSU11, AM12, ATG11, BdGN^{+}18, BPG^{+}17, BM18a, BWK^{+}13, CLWB15, CA10, CAP^{+}10, Cer14, CHA15, CLS^{+}11, CEB^{+}19, CJG^{+}15, CBMLB13, DMK^{+}19, ERG^{+}11, EPHF18, EWB^{+}18, EÖFPH17, FPM17, FMB^{+}19, FBY^{+}12, FBM^{+}19, FHM10, FHMK11, GBM^{+}18, GKKK15, GDB19, Gri10, GBEP16, GYCL19, GAR10, HB12, HFBS15, HFL14, Hen11, HSL^{+}13, HLS^{+}16, HSLB19, HNS12, Ing11, IEL17, IGSJBS^{+}15, JČK^{+}12, JČK^{+}21, Kni12, KTNA15, KVH15, KTT13, KLWG17, KHKM15, LPHH11, LHF16, MK17, MMGEH15, MHK^{+}17, MBRDM18, Mau11, MMBD13, MVMdS^{+}17, MQCS^{+}14, MWC15, MBNM10, MLT17, MC10, Mun18]. size [NWS^{+}11, OTST11, OHSG19, PPVP17, PBD21, PDP^{+}11, PHM16b, PDS^{+}17, PACF17, QWMD13, RG18, ŘJP^{+}12, RHB16, SL11a, SWB12, SFR^{+}13, SFPB17, SZPV16, SFL^{+}14, SSC^{+}12, SSHM16, Sve15, Szu16, TXH^{+}19, TMV16, TJMT17, THG^{+}16, TBSE17, VKNK12, WAB^{+}16, WLGS19, WHMS11, WHF^{+}14, ZCX^{+}18, FWG15]. Size- [SKK^{+}13, CA10, EWB^{+}18]. size-at-age [OTST11, SFPB17]. size-based [KTNA15]. size-composition [MVMdS^{+}17]. size-dependent [CBMLB13, HFBS15]. size-frequency [RG18]. size-limit [Mun18]. size-selective [GKKK15, GAR10]. Size-selectivity [JC18]. size-sorting [HB12]. size-spectrum [ZCX^{+}18]. size-structure [CLWB15, SFR^{+}13]. size-structured [FBM^{+}19, PDP^{+}11, PHM16b, PDS^{+}17, PACF17, Szu16]. sized [CM18, KHKM15, YAT^{+}19]. sizes [BBM17, Chi10, HFPL16, HCM^{+}11, IEL17, MM13, PMB^{+}11, PHRH16a]. Sizing [ABD^{+}19]. Skagerrak [SNK19]. skate [AENC14, ERCG10, JLLL11, KM10, KHG^{+}17, LHSS18, PSE15, QWCB11, SBC^{+}18]. skates [MCI^{+}13, NBG^{+}17]. skew [CRAV13]. skew-normal [CRAV13]. skewed [MB16]. skill [KKK16]. skilled [vPWW16]. skip [Mur19]. skip-molting [Mur19]. Skipjack [GMZ^{+}14, ABV16, HCGS19, HE15, KH15, MST14, OTH^{+}19, SMNdC^{+}19, SF19]. skipped [PKWP15]. Skipper [GSDD11]. skippers [JTP^{+}14]. Slipper [dADSRG10, AGSF19]. slipping [HV10]. Slope [LRY^{+}15, AGB15, GPdR11]. Slow [MBOAA12]. SM20 [PGJ10]. Small [AAOE13, CEJI11, CMLP19, GFM^{+}12, GBK^{+}13, MMNCR^{+}13, AASF16, ACSAS^{+}17, ASMP^{+}10, ASMD^{+}18, AFOB^{+}19, ADH^{+}13, BMEGV^{+}19, BOCÁ^{+}18, BMRC14, BMOZ13, BMF^{+}10, BMF^{+}11, CCC^{+}19, CMB11b, CAC15, CMP^{+}14, CBS17, Cro15, DPB^{+}14, GOD19, dMGPR^{+}18, GBD^{+}15, Gra16, GRB14, HFD^{+}13, HN16, HB13, Hen11, HAB^{+}17b, INSN14, JPSR^{+}14, JvDB^{+}17, KPM^{+}19, KMW^{+}11, Kos15, LOdH16, LAB13, LST^{+}10, LPC^{+}14, LSJD11, LZGL17, MPA^{+}12, MCM12b, NCP15, NvZH^{+}17, OTK^{+}16, PTG^{+}17, PMG18, PTSC^{+}16, PVMG19, PJML15, PTTC18, POV19, POM^{+}18, PNAL16, RLC^{+}17, RLC^{+}15, RRSB13, RRO^{+}19, RDP^{+}17, RBGE^{+}10, RFS16, SFBV17, TLLZ19, TTS11, TD19, VPSH^{+}17, VAB^{+}14, WLZ^{+}15, WAP^{+}12, WGC^{+}19, WGC^{+}21, WJC17, YvZG^{+}17, ZYW^{+}14, Zie12]. small- [YvZG^{+}17]. small-bodied [PTTC18, WGC^{+}19, WGC^{+}21]. small-mesh [Hen11]. small-pelagic [CCC^{+}19]. Small-scale [GFM^{+}12, MMNCR^{+}13, ACSAS^{+}17, ASMP^{+}10, ASMD^{+}18, AFOB^{+}19, BOCÁ^{+}18, BMOZ13, CMB11b, CAC15, Cro15, GOD19, dMGPR^{+}18, GBD^{+}15, Gra16, HN16, HAB^{+}17b, KMW^{+}11, Kos15, LOdH16, LPC^{+}14, NCP15, NvZH^{+}17, PTG^{+}17, PTSC^{+}16, PVMG19, PJML15, PNAL16, RLC^{+}17, RLC^{+}15, RRO^{+}19, RDP^{+}17, SFBV17, TLLZ19, TTS11, TD19, VPSH^{+}17, VAB^{+}14, WJC17, Zie12]. small-scale-fishery [KPM^{+}19]. small-spotted [BMRC14]. SMART [GPT^{+}19]. smelt [GFCK13, KBEB13, MDH15, PTB^{+}15]. smolt [FMF^{+}17, KNSPK13]. smolts [SBAK17]. Smooth [SPM^{+}10, BdGD^{+}16, CBC^{+}10, SC19]. smooth-hound [CBC^{+}10]. smoothhound [BMdS^{+}15, WJC17]. smoothing [PL18]. Snake [RSQ18]. snakehead [DUC^{+}19]. Snapper [FMA^{+}18, AFR^{+}17, BBH^{+}12, CTSD10, FMMW13, FGK19, GSP17a, GDHA13, GCRV10, HPSG15, HSC19, JCW12, KKH^{+}17, LKZ^{+}17, ODMF17, PF15b, SVV^{+}14, SMO^{+}18b, TS11, WGS17, HCSM17]. snapper-grouper [ODMF17]. snappers [ASR^{+}18, WWN^{+}14]. Snook [BTB^{+}19, ESF^{+}19]. Snow [DWH10, DMK^{+}19, KC11, MCMK18, MRT17, MRT18, Mur19, NWL^{+}14, OHSG19, SP12, WW11b, WLBB15, ZC17, SP15b]. SNP [ER18, GKMGO^{+}17, SRMP^{+}18, WWLY15]. SNP-based [GKMGO^{+}17]. SNPs [CMI^{+}16]. Soak [CNS11, JEJS17, OHSG19, PGL17, PMP^{+}11, SKFM17]. social [CBH18, DDK^{+}18, GDB19, HSG15, HRC16, KDM15, LvVtB^{+}15, MCEP14, PHG^{+}18, RRSB13, TBF^{+}16, UARC11]. Society [Ano13a]. Socio [AFA11, BRC^{+}10, KPL^{+}07, KPL^{+}09, MMNCR^{+}13, NLPS^{+}13, TTS11]. socio-bioeconomic [KPL^{+}07, KPL^{+}09]. Socio-cultural [AFA11]. socio-economic [BRC^{+}10, MMNCR^{+}13, NLPS^{+}13, TTS11]. Socioeconomic [FL11, TXP^{+}16, CHS^{+}17, KZ11, MMdlP16, SSR18, SZ11]. sockeye [BV14, DHP^{+}11, ER18, GHE^{+}11, MMMPP19, TQ17]. software [HBB^{+}12]. solandri [Zis12]. sold [OPL17]. sole [Eth15, GBD^{+}15, NKW^{+}19, PRL^{+}18, RBM13, Smi14, SDV^{+}16, ODdA^{+}10]. Solea [Ata11, Eth15, PRL^{+}18, RBM13, SDV^{+}16]. Soleidae [Ata11]. solemdali [CN19]. Solenocera [VRW11]. Solenoceridae [VRW11]. solid [LFG11]. solidissima [HMMP18]. Solis [DHFSS18, BPRB^{+}14]. Solomon [RFS16]. solution [TC15]. solutions [GMS^{+}17, SAF^{+}15]. Somalia [SB14]. somatic [AT10, LCWC17, SD10]. Some [Cam15, Han18a, Pun17, PSS15, Ste11]. sonar [CT16, CMJ^{+}19, MBF19a, RTŘ^{+}12, SBLB17, PJM14]. sonars [PGJ10]. Sonora [CMMVRF^{+}19]. Sort [GSH^{+}15]. Sort-V [GSH^{+}15]. sorting [GLB^{+}19, HB12, HSL^{+}13, LW13, LW16, MHFH17, MBF^{+}19b, SBH^{+}16, vPWW16]. Sound [CQ14, HHCP13, WDNB^{+}19, CPOH14, GMJ^{+}13, SBS17, LSJ17, TMC19]. sounder [LMSM14, OLM^{+}19]. source [HBB^{+}12]. Sources [YvZG^{+}17, CC11, RRW^{+}13]. South [JDLF11, SGSFL19, Ste11, XDL^{+}19, BdGM^{+}11, BdGD^{+}16, BMdS^{+}15, BB13, CAM^{+}14, FE10, HSS^{+}16, LCA^{+}19, LSM^{+}19, LGH^{+}21, LW19b, NL11, PBR^{+}16, VGE11, Way13, WKTD13, ALR^{+}17, BMdS^{+}15, BCPP17, But16, CEAP18, DUD^{+}18, DJF^{+}15, FPM17, FFG^{+}16, GAR19, HND^{+}17, HDW^{+}15, KJKK15, KMdS^{+}18, LBHP16, LCZ^{+}16, LSJ17, MFC^{+}13, MMC13, MCDP18, MBS14, MBOAA12, MLF^{+}10, MGR13, PML^{+}12, QMGP10, RAG19, RM19, RSQ18, SMO^{+}18a, SB16, TWCA15, VF10, WRH^{+}15, XdA19, ZzCbC^{+}15, ZDZY19]. south-central [FE10]. south-east [BB13, HSS^{+}16]. south-eastern [LSM^{+}19, LGH^{+}21, NL11, Way13, WKTD13]. South-West [MMC13]. south-western [BdGM^{+}11, BdGD^{+}16, LCA^{+}19]. southeast [BSS^{+}13, BB18, LFG11, PDF^{+}18, dJSMROZP^{+}13, ALB^{+}18, MBHGN13, YC18]. Southeastern [HdSB^{+}14, ASMD^{+}18, BMB10, CRS^{+}19, CVS11, dADSRG10, GCK14, GBA^{+}18, HBG^{+}19, RMB^{+}21, RRS^{+}19, SAH15, SH10, ZZA^{+}14, LG13]. Southern [LSM^{+}19, MAS17, MBF^{+}19b, PPVP17, RKH^{+}17, BRGS19, CMMF13, CH11, CH14, CH15, CJCN21, CDPJ^{+}13, CRAVC14, CCC14c, DAJMB16, DDB^{+}14, EHGG16, FPM17, FWG11, FWIB14, dSFdSSS17, FSBS11, FDMH16, GCK14, GHP^{+}15, GBPC10, HN16, HAB^{+}10, Kri13, LGH^{+}21, LJJ^{+}13, MCC^{+}17, MCP^{+}17a, MWSN^{+}13, ML10, MM11, MDPCSR10, MWA^{+}15, MFRM19, MKM^{+}16, NGvdL^{+}13, NEJ^{+}18, ODD16, PF15a, PVMG19, PJML15, PML^{+}12, QWCB11, RAGP18, RRRM19, SSFGL19, SAH15, SLT^{+}19, SP10b, SHA^{+}15, SCV^{+}12, VCS^{+}19, WB16, WWR18, WJC17, ZM15, ASSL11, CCC^{+}14b, FGRIT^{+}12, HdSB^{+}14, PPG^{+}11, RHW^{+}16, SGSFL19, SPC^{+}13, TSNRU21]. southernmost [CDM^{+}19, MTV17, ZC17]. Southwest [ABL21, HL16, SAR^{+}15, DSG^{+}17, DBT15, RB12, SMNdC^{+}19, AGB15, EMMM13, LAB13, MBHGN13, MQV10, MQCS^{+}14, SSD^{+}19, dVLB^{+}15]. Southwestern [IIT18, VFdS10, BAL^{+}19, LCST10, MD16, SCL^{+}15, CLS^{+}19, CCM14, FGCG^{+}17, dSFdSSS17, GMUD11, IIT17, dSSdAdAA17, SPV19]. Soviet [PLA16]. sp [HHSS17, RSEERRSC15]. sp. [GLG^{+}11]. space [AAS19, GIK^{+}16, GBS19, KY19a, MDH15, MB16, MRT18, NB14, NB19, SDV^{+}16, TW13, TD19, XLCC19]. spacing [BST^{+}19, HB12, MS18]. spacings [SBD^{+}12]. spadefish [SSD^{+}19]. Spain [AGSF19, DSJ13, GHUG^{+}16, HFFdlT15, MCJD^{+}17, PFVGGM16, SJR^{+}11, TRS12, VRMF10, ARVACPMN16, BPS^{+}18, GDB19, PHG^{+}18, RGAP18, VRMF13]. Spanish [BBHF10, CPP^{+}10, CMPP11, CMP^{+}12, DBG19, DSG^{+}17, EPHF18, GGADSRVL15, GM15b, HSRT17, LOdH16, LYWZ15, LMSM14, NML12, RU15]. spanner [SBLL17]. spanning [KTT13]. sparid [DJF^{+}15]. Sparidae [NVS^{+}18, VGE11, WFLP11]. Sparse [DMK^{+}17]. Sparus [ALUFJ^{+}12]. spat [HBT16a]. Spatial [ALB^{+}18, ACRM12, APD^{+}18, BK12, BK16, BMR^{+}18, CMPR^{+}16, CBF^{+}14, CKC13, CAGVB14, CEAP18, DHH18, EÖFPH17, FMA^{+}18, FCNB16, FAAM16, GLMK12, GMUD11, HTHK11, Kni12, LFCC15, MD16, MWA^{+}15, MSN^{+}12, MYJ^{+}15, PBG^{+}13, Pun19b, SB12, TCBR17, AMGH15, AMHR16, BJZB12, BJP^{+}13, BFL^{+}14, CGM^{+}19, DBLB16, DBSA18, EPO10, EvHRF10, FPS11, FCL16, GMJ^{+}13, GTS^{+}17, HTB15, HFCPSEM15, HD12, HMD^{+}19, HFPH14, JS15, JSG^{+}15, KBEB13, LCMTM10, LCMdSM13, LGH^{+}21, LWZ^{+}19, Lyn14, MHB^{+}19, MD13, MGM12, MLF^{+}10, ML11, MBvDN14, MDPH17, OPH15, PAX^{+}17, PTSC^{+}16, PGJ10, PBC^{+}16, PCP^{+}18, PRL^{+}18, PHT15, POM^{+}18, Pun19a, SCC^{+}15, SWD^{+}13, SH13a, SKB^{+}18, SKB^{+}19, TK16, TRH16, VF13, VB15, Via19, VMB18, WSMS14, WSG16, YJC13, ZH10]. Spatially [PTBT17, BCC15, BGB^{+}19, CMP^{+}14, CP11, CP12, GBH^{+}19, GKL12, GDA14, GDC^{+}19, KCM19, LSS^{+}16a, LJB16, PD18, PHLT16, SAB^{+}11, SP15b, Whi10]. spatially-explicit [GDC^{+}19]. spatially-structured [BCC15, PHLT16]. spatially-varying [LJB16]. Spatio [CCA16a, EHBK18, FF11, GBEP16, Gun17, HBG^{+}19, Kai19, PT19, RvRW^{+}14, Tho19a, Tho19b, YCYC16, CN18, GWB^{+}19, HGG13, OMI^{+}14, OGCBMR15, PLSM^{+}16, RRO^{+}19]. Spatio-Temporal [Tho19a, Tho19b, CCA16a, EHBK18, FF11, GBEP16, Gun17, HBG^{+}19, Kai19, PT19, RvRW^{+}14, YCYC16, CN18, GWB^{+}19, HGG13, OMI^{+}14, OGCBMR15, PLSM^{+}16, RRO^{+}19]. Spatiotemporal [APL^{+}18, BC18, HNY^{+}19, SF19, SPS^{+}19, SRB^{+}18, TGDB16, XLCMMV19, JTP19, TC16a, TC16b]. spawner [CMLP19, DPAB10, HGD19, SMdUK16, SCF17b, ZCD^{+}12]. spawner-recruit [CMLP19, SCF17b]. spawners [GHA13]. Spawning [BHG^{+}17, BJP^{+}13, CCHD19, FMA^{+}18, FF10, GIW16, GIW17, GHUG^{+}15, GHUG^{+}16, HPHR14, HM16a, HM16b, JKS^{+}15, RM10b, WFLP11, WBI^{+}16, APDJSU11, ABL21, AW12, ADH^{+}13, BV14, BGF^{+}15, BHT^{+}19, CRCMP^{+}14, CPD^{+}19, CPOH14, CCC^{+}14a, CCGD19, DR13, DH14, DOM14, DPSRM10, EMMM13, EAM^{+}14, FSH^{+}13, GDHA13, GPBA13, GHJ13, HCWS19, HBMC12, HRUSR13, HLP15, IIS10, JCW12, JS15, JTP^{+}14, KWM^{+}15, KBOM14, KR12, Kur12, Leg14, LXCC19, MRLD18, MBJ^{+}16, MD13, MHRD16, MJL14, MCP14, MWR^{+}13a, MVMSdS19, MBB13, MLT18, MWR13b, Mun18, NNS^{+}10, NL11, ODD16, OALS11, OSMP18, PSL^{+}19, PKWP15, PUSI10, POM^{+}18, RSG^{+}11, SAK10, SWH17, SH13a, SSJ^{+}13, SLCL15, SMPB17, TMM19, TQ17, TCS13, UASM12, WMF13, WBB^{+}16, Han16b]. spearfish [AGB19, BSD^{+}13]. Special [Ano13a]. specialised [Gri12, ZGT12, ZG14]. specializations [PCWM16]. specialize [KFBL13]. specialized [LWSC11]. Species [CCM14, GVMSCJ12, HHB18, KHKM15, MRS15, MMO^{+}13, PMC^{+}18, PMM^{+}15, QEH^{+}11b, ARF^{+}17, AFOB^{+}19, ASSL11, BTBP18, BBF^{+}18, BHPC14, BEM13, BMdS^{+}15, BNA^{+}19, BK15, BB14, BSM15a, BB13, CAMFMF18, Cer14, CAC19, CBMLB13, CPH12, CMLP19, DPB^{+}14, DMFR16, DWH10, DBG19, DDB^{+}14, DCP^{+}16, DDH^{+}15b, DLCO^{+}19, EFWC^{+}19, EvHRF10, ETMK13, FRDC10, FCLH19, FCOM18, Fit12, FWJ^{+}16, FHK17, FHMK11, FSPWG08, GKMGO^{+}17, GBM^{+}18, GGS^{+}19, GWVH^{+}16, Goo16, GWM18, GSTL15, GH14, GBR^{+}18, GDC^{+}19, GM15b, GPT^{+}19, HKRH12, HFD^{+}13, HSS^{+}16, HIH19, HAB^{+}17a, HAB^{+}18a, HRAK15, HMGH12, HSH15, HSH16, JP16, KBG^{+}14, KJS12, KA15, KHS12, KC18, KRC14, KRB^{+}15, KTW^{+}13, KGB11, LAB13, LCMF^{+}16, LCM^{+}18, LSC^{+}18, LYWZ15, LL14b, LQL^{+}15, LSL12, MP18, MMGEMGS10, MNT^{+}17, MPB^{+}17, MFHK15, MD13, MPF^{+}17, MGR14]. species [MSV^{+}15, MWR13b, MCM12b, Mun12a, Mun12b, NVS^{+}18, OPL17, OPF15, ODMF17, PF15a, PGW17, PKAL13, PTSC^{+}16, PDP^{+}11, Pun11, QWC10, RBT14, RCLS18, RBGE^{+}10, RGSAMV19, RAGP18, RHB16, SRSNC17, Sal18, SGSFL19, SPV19, SNK19, SAH15, SBS17, SPMS17, SVPC16, SMS^{+}19, SHM12, SGC11, SB16, Ste18, SRG17, TSNRU21, THHG18, TB10, TFB^{+}12, TCS13, THA^{+}14, TRS12, TAS12, UASM12, VGE11, Via19, WM11, WRR^{+}16, WM13, WHC^{+}19, WPCO15, YAH^{+}13, ZQZ^{+}17, ZSF11, ZLT^{+}16, dPHS^{+}12]. species- [EFWC^{+}19]. Species-specific [MRS15, PMC^{+}18]. specific [AJOE17, BHI14, CAM^{+}14, EFWC^{+}19, GMS^{+}17, HKF^{+}12, MRS15, MSV^{+}15, OMI^{+}14, PMC^{+}18, PB11, SKFL15, SP10a, ŠBB^{+}19, SMP18, SD13, SAWS18, SHR18b, WKTD13]. specification [IOT14]. specifications [HFBS15]. specified [PACF17]. specimen [BCMS^{+}18]. specimens [BCMS^{+}18]. spectrum [ZCX^{+}18]. speed [MMGEH15, PB14]. speeds [dKMS15]. Sperm [HPP18, JRG^{+}18, KWM^{+}15, WTS^{+}17]. Spermatogenesis [TMM19]. Spermonde [NMLZ17]. Sphyrna [NWS^{+}11, SC19]. spill [HCSM17]. spillover [SF13]. Spine [TGM13, CFB^{+}17, Cor17, LRML^{+}15]. Spine-based [TGM13]. spined [JBD^{+}18, RvAGB18]. spiny [ARVACPMN16, BMF^{+}10, BMF^{+}11, BSKT13, CPS^{+}12, GRPEMEVV17, HTB15, HCF^{+}17, KT18a, KM15, MMBD13, MS15b, SG15, SPC^{+}13, SCOC16, TGM13, TGMM13, TLCSG^{+}15]. Spirinchus [MDH15]. Spisula [HMMP18]. Spit [GDHA13]. spline [HPB10]. split [HJT13, RSEERR^{+}16]. split-beam [RSEERR^{+}16]. Spondyliosoma [NVS^{+}18]. sponge [BBV17]. sponges [BBV17]. sport [BBF^{+}18, CC11, ZGT12]. spot [ASD^{+}18, CGECSEH10, FRDC10, RM10a, Sto12]. spotted [BMRC14, CSOV14, EHB18, FF11]. spp [BBA^{+}19, CEB^{+}19, CSP^{+}18b, DSW^{+}17, FAAM16, FMD19, GGS^{+}19, KBG^{+}14, LHS^{+}16, LW12, MDL19, MLV13, MCB^{+}16, NBG^{+}17, VPC^{+}14, ZLT^{+}16]. spp. [PGS10]. SPR [PVKH18, PVKH15]. sprat [CDM^{+}19, DHH18, HTHK11, HHH^{+}12, HRUSR13, MCN16, MLC^{+}19b]. Sprattus [CDM^{+}19, MCN16]. spread [BSM14]. Spreading [GS12, RGG14]. spring [BKM13, CDMC18, GHJ13, NNS^{+}10, OALS11, OSMP18, SSJ^{+}13, YCYC16]. spring-spawning [GHJ13, NNS^{+}10, SSJ^{+}13]. Squaliformes [CAC^{+}14]. squaloid [MJC^{+}14]. Squalus [BSKT13, CPS^{+}12, HCF^{+}17, PGLG11, SPC^{+}13, TGM13]. square [BHS16, BMB10, DBK^{+}10, DÖA^{+}10, HLS^{+}16, KHMF11, LCF15, MMJ^{+}12, SLP^{+}15, SMD^{+}16]. square-mesh [BHS16, BMB10, MMJ^{+}12, SLP^{+}15]. squat [CN18, HW11, TEFAR12]. Squid [LAB10, AT10, AGB15, BHPC14, BPMGP10, BMB10, CRAP^{+}11, CRCMP^{+}14, CMB11a, CCC17, FTJ^{+}16, HdSB^{+}14, HB15a, HB15b, Hen11, HMGH12, HBG^{+}19, IIS10, KYO^{+}12, LGZC17, MAY12, MRS15, NMMBCV^{+}10, PAX^{+}17, PPdR^{+}10, PG14, QCZC^{+}19, RDR^{+}18, Rob15, RGGdL13, RLVS14, SD10, STY^{+}16, SP10b, TS10, UK14, VSTS10, WA15, YAT^{+}19, YMA12, YCYC16, YC18, ZzCbC^{+}15]. Squilla [LMBF13]. Sr [LRY^{+}15, LRY^{+}15]. Sr/ [LRY^{+}15]. SRA [WvPLW18]. Sri [DAJMB16]. SST [RGGdL13]. St [GLA19, Lam13, MM11]. stability [BK10, DL11, GRPEMEVV17, ZCX^{+}18]. Stabilization [Eng18]. Stable [LSWR17, MCN16, CBS11, DDP15, ESF^{+}19, FTJ^{+}16, JD14, KHM^{+}17, LGZC17, MMMPP19, MBB13, NMH^{+}11, PPLK18, TOTI10, TCS^{+}15, VIFLP17, VSM^{+}18, VSLC^{+}19, WMIDS18, WW14, ZDZY19]. stage [ARF^{+}17, BHG^{+}17, CMM^{+}14, LSWR17, PM11, SKK^{+}13]. stage-dependent [SKK^{+}13]. stages [APD^{+}18, JČK^{+}12, JČK^{+}21, MRS15, MQV10, WBM17]. Stakeholder [BBB^{+}16]. standard [BAM16, CJG^{+}15, HWS^{+}15, THA^{+}14, WAP^{+}12]. standardised [FSA^{+}16]. Standardising [BBLS17]. standardization [CAMW11, CHL11, CSMO14, FSG^{+}19, GWB^{+}19, JTP19, MS18, Peñ18, TW13, WB15]. standardize [IB10, TSNRU21, WKA13, WKA14]. standardized [HNY^{+}19]. standardizing [LYWZ15, TW14]. starting [PKRC13]. starvation [SSFGL19, YA18]. State [BAOP18, CUL^{+}16, ABV16, AAS19, GBS19, KY19a, LWS19a, LWS19b, LMSM14, LFG11, MDH15, MB16, MRT18, NB14, NB19, PDBM12, PBK^{+}19, PHM16b, RW19, XLCC19]. state-space [AAS19, GBS19, KY19a, MDH15, MB16, MRT18, NB14, NB19, XLCC19]. stated [LL14a]. States [PF15b, BMCS19, BSS^{+}13, BB18, BT16, DDP^{+}16, GBA^{+}18, HCF^{+}17, HA12, MS15b, NBS15, PHM^{+}16a, RMB^{+}21, SH10, TAS12]. static [HC17]. stationary [DRGA12, GFCK13, PTW^{+}13, SSJ^{+}13]. Statistical [BTS^{+}19, HCHA12, ML17, TSNRU21, TTT16, Cop13, KOHF^{+}16, MP17, MW13, PM11, RUMBA15, TK16]. statistics [Tho19c, Tho21]. statoliths [SP10b]. Status [CUL^{+}16, ASB^{+}16, AJP18, BCKG13, BSF^{+}19, CWM12, CTWD15, FSA^{+}16, FJWD17, HF15b, IOT14, IIS10, KY19b, KHGB14, KSAA13, KTNA15, LCMF^{+}16, ML17, OOM15, PHB19a, RCBS19, SM15, SLT^{+}19, VS16, Zis12]. steady [PBK^{+}19]. steel [LHS^{+}18c]. steelhead [JNF^{+}19, TGE^{+}18]. Steepness [TDH19, CHT18, LMPM12a, Mun18, PD14, Whi10, WFNC18]. stellate [YVJ12]. stelleri [PMB^{+}11]. stenolepis [LW13]. stenothermy [AGB19]. step [FTM^{+}15, Fra14, MTH10, RMB^{+}21]. Steps [PHB^{+}19b]. stereo [BWR^{+}18, KT18a, LNC^{+}15a, LNC^{+}15b, SWB12, HNM^{+}12]. stereo-BRUVs [HNM^{+}12]. stereo-video [LNC^{+}15a, LNC^{+}15b]. Stereolepis [CHA15]. Stereological [KFGN10, BTNA13, KMKK10]. sterile [PMG18]. Stewardship [BGW^{+}16, BPPR16, BFHT16, PRCGD16, PLA16]. Sthenoteuthis [ZzCbC^{+}15]. Stichopus [RLM10]. stickleback [EE19, JBD^{+}18]. sticklebacks [RvAGB18]. stiffness [OKB^{+}16]. still [FWG15]. Stilt [DAJMB16]. stimulating [KW10]. stimulators [MKHK19]. stimuli [BKLB14]. Stochastic [BMU14, SCT^{+}17]. Stock [BWN^{+}12, BVN^{+}10, BJKH12, BMS15, CBS11, DRGA12, DM11, Gri10, KF18, KMJ^{+}12, MW13, PM13, RU15, SCJ^{+}11, SCC^{+}15, SJJ^{+}11, SSD^{+}19, VRMF11, VRMF12, WRH^{+}15, WvPLW18, ZMdSC16, APCF17, AAS19, ARF^{+}17, AM12, APB^{+}16, AW12, BMRC14, BEM^{+}17, BJZB12, Ber19, BDC^{+}17, BHC^{+}16, BBHF10, Cam15, CCA16b, CHT18, CCC^{+}19, CMSG18, CWM12, Car18, CBF^{+}14, CAM^{+}14, CPC^{+}17, CJWT^{+}19, CDPJ^{+}13, CDF^{+}13, CLB^{+}10, CC19, Coo13, CBMLB13, CP11, CP12, CTWD15, CV14, CMLP19, Cur18, DJW^{+}18, DS13, DM16, DDP^{+}16, DJF^{+}15, DBA18, DF16, EWB^{+}18, FTJ^{+}16, FE10, FWG11, FBM^{+}19, FSA^{+}16, Fra12, Fra14, Fra16a, FdSM^{+}16, Fra16b, Fra17, FJWD17, FDMH16, GCDAS^{+}10, GCL19, GWB^{+}17, GBH^{+}19, HKRH12, Han16b, HWS10, HFBS15, HFPL16, HF19, HPHR14, HM16a, HM16b, HLSD21, Hor11, HMA^{+}16, IOT14, JBD^{+}18]. stock [JKK^{+}16, KY19b, KBOM14, KKK16, KWS^{+}15, KMSRW10, KOD^{+}15, KVH15, KH15, KEH16, KEP^{+}19, KHV^{+}12, KPF^{+}18, KOHF^{+}16, Lam13, LMPM11, LMPM12a, LMPM12b, LP18, Leg14, LTHRU17, Lor16, MT13, MHK^{+}17, MAAW^{+}17, MD10, Mau11, MH11, Mau12, MP13, MD13, MCVS14, MCP^{+}16, MT19, MLF^{+}10, MDPCSR10, MBG^{+}10, ML17, MVMSdS19, MOA^{+}16, MMAA^{+}18, Mun18, NB14, NB19, PB11, PL18, PPVP17, PRL^{+}18, PRR^{+}19, PD14, PSS14, PHFW14, PHLT16, PHM16b, Pun17, POM^{+}18, PC19, Pun19b, RADR11, RLC^{+}17, RJ16, RUMBA15, RM10b, RLMM18, RCBS19, SVV^{+}14, Sam14, SMRS15, SAK10, SGN^{+}10, SLH^{+}14, SPB^{+}19, SEG^{+}15, SLPT17, SSS^{+}12, SD13, SHJ19, SHT^{+}13, SM14, SM17, SLCL15, SP12, SP15b, Szu16, SCOC16, TGMM13, TSS11, TC15, TJMT17, TC17, Tho19a, Tho19b, Tho19c, Tho21, VF13, VHS^{+}17, VOM12]. stock [VMB18, WMP^{+}14, WMNC15, WLZ^{+}15, WBM17, Way13, WKT^{+}13, Whi10, WKTD13, WWSM15, WFNC18, WA15, WTK10, Wri13, WRG^{+}18, YVJ12, Zis12, dEWB^{+}15, dZB16, vPWW16, Cop13]. stock-assessment [Mau11]. Stock-based [BVN^{+}10]. stock-enhancement [CBF^{+}14]. Stock-recruitment [BMS15, KF18, CHT18, Han16b, HF19, HM16a, PC19, WFNC18]. stock-reduction [TC15]. stocked [CBB^{+}16, CM18, DLCK15, KT18b, PJR17]. stocking [CLS^{+}11, ER18, KNSSL11, KPF^{+}18, KT18b, KAiT^{+}13, PMG18, RFC^{+}17, SSFGL19, VvPM^{+}18]. Stocks [CLS^{+}19, AB11, AHWY13, AW12, BVN^{+}10, BCN^{+}13, BFL^{+}14, BS10, CDM^{+}19, CFF^{+}16, CBS17, DM11, ER18, FHK18, FGCG^{+}17, FJWD17, GGM^{+}18, GPP^{+}16, Han13, Han18b, HBS^{+}16, HA12, IHFAH17, JAB15, Jaw11, KNJK^{+}10, KNRJ^{+}15, KTNA15, KTT13, LSŠ^{+}16b, LAB13, LRY^{+}15, MREV10, MM11, MPA^{+}12, MKM^{+}16, MBW12, MLT17, NPK^{+}10, NBS15, OOM15, PG18, PSE15, PSS15, PDF^{+}18, RRM16, SO18, SDWG18, SPM15b, SAB^{+}11, TMM19, TCBR17, VJH17, WB16, WP11, WP17, WFJ19, WFNC18, WAT^{+}13, dPHS^{+}12]. Stomach [TCS^{+}15, VSLC^{+}19, BNM11, BBLS17, KNB^{+}19, PPLK18, VIFLP17, VSM^{+}18]. Stomatopoda [LMBF13]. Stomolophus [MDL19]. stone [CCGD19, DSW^{+}17, KGS18]. storage [CDB^{+}14, CKC16, DRE^{+}17, DPSRM10, JTP^{+}14, LDR^{+}12, OTA^{+}13, OPE^{+}17, SAK^{+}17b]. Strait [BRW^{+}13, CEAP18, FPC^{+}13, FDMH16, SVG^{+}17, YAT^{+}19]. Straits [MMF^{+}17]. Strangomera [CJCN21, CCC^{+}14a, CCC14c, FE10]. strata [OPH15]. Strategic [SHGP11b]. Strategies [AVU^{+}10, SSJ^{+}13, BPMGP10, CMSG18, DDH^{+}15a, DDH^{+}15b, FPS11, FWJ^{+}16, FPD^{+}16, GS12, Gra16, GYCL19, Han18a, HCH^{+}19, JSG^{+}15, MMG10, OMB18, Pun11, TIGCDL14, VvPM^{+}18, YVJ12]. Strategy [RBR^{+}15, APD10, BGBCM19, CPD^{+}19, DAJMB16, dMGPR^{+}18, HBB^{+}19, HLP15, JHCT11, KJS12, KWF12, MCB^{+}18, MMF^{+}16, Nak17b, NVS^{+}18, SMCR13, Sve15, VFFG17, WK10, WWSM15, ZMG^{+}13]. stratification [CZW14, IB10]. stratifications [WXZ^{+}18]. stratified [EHB^{+}21, ST13, WXZ^{+}18]. stream [PTBT17, PSJ^{+}19]. streams [AAN^{+}12a, BAB^{+}12, LKNL16, MDS^{+}18, SBJ^{+}19]. strength [AD16, BST^{+}17, DMK^{+}17, DOO15, FCLH19, FKM11, GFCK13, GSP17a, GBB^{+}12, JSB19b, KEMM12, OCL^{+}18, RT14, RSEERRSC15, RSEERRSC16, RSEERR^{+}16, SWG18, SMS^{+}19, TCAP11, ZKSF17]. strength-total [JSB19b]. Stress [MWR18, RR12, DRE^{+}17, EDG16, HAB^{+}10, KCBS13, LBB^{+}13a, OTA^{+}13, SLT^{+}19]. stress-induced [HAB^{+}10]. Stress-related [MWR18]. striata [DUC^{+}19, WMF13]. striatus [OTST11]. striped [BCN^{+}13, BMS15, CBO^{+}15, DSJ13, DUC^{+}19, HKM^{+}17, LKD15, LKD18, MREV10, OGCBMR15, PSM^{+}18, SSP^{+}15]. Strombidae [BdGN^{+}18]. Strombus [SMBP^{+}12]. Strong [EvHRF10, HLJ15, VvHE^{+}11, CHCS14, PBB^{+}18, WOPB11]. Strongylocentrotus [JEJS17]. strontium [SLL^{+}16]. structuration [PRL^{+}18]. Structure [ASMP^{+}10, AL13a, RMMSR19, APCF17, Ame12, BWN^{+}12, BEM^{+}17, BJZB12, BGA11, CGM^{+}19, CMPR^{+}16, CAFHG10, CLWB15, CCA16b, CH15, CDPJ^{+}13, CDF^{+}13, CLB^{+}10, CMMVRF^{+}19, CJG^{+}15, CRW^{+}19, CPG^{+}11, CHCS14, DJF^{+}15, DUC^{+}19, EÖFPH17, FPS11, FHK17, dSFdSSS17, GRGGCACG11, GJSR10, GM15a, GMUD11, HVD^{+}18, HPSG15, HMvHN13, HFPH14, IIT17, IIT18, JKK^{+}16, JRA10, KBOM14, KWS^{+}15, KWCT^{+}18, KBEB13, KRB17, LPMM14, LLH^{+}14, LFCC15, LBMC15, MSGV11, MGP^{+}18, MMBD13, MAS15, MAS17, MLBGV16, MSN^{+}12, MFS^{+}18, MFVPC19, MMO^{+}13, NMH^{+}11, NNS^{+}10, OTST11, PGAS15, PPG^{+}11, PGM^{+}10, PLM16, POM^{+}18, RGK19, RPH^{+}18, RAG19, RTSV15, RSV^{+}17, RCHSN^{+}15, SCJ^{+}11, SJJ^{+}11, SFR^{+}13, SMNdC^{+}19, SSD^{+}19, SSG17, TVCT12, TOA12, TVdLOG^{+}21, TLCSG^{+}15, VPC^{+}14, VRS13, WMP^{+}14, WMC^{+}16, XYL^{+}19, ZYW^{+}14, Zie12, Zis12, dVLB^{+}15]. structure-based [FPS11]. structured [BCC15, BGB^{+}19, CA10, CTWD15, EWB^{+}18, FBM^{+}19, Fra16b, GS18, Hor11, KCM19, MOA^{+}16, PDP^{+}11, PHLT16, PHM16b, PDS^{+}17, PACF17, ŠBMG^{+}16, SGGH10, SP15b, Szu16, TT14, TRW19]. structures [BGD11, KDM15, VF13]. structuring [HND^{+}17, WRG^{+}18]. studied [PTB^{+}15]. studies [Ben18, GBK^{+}13, HFPL16, HF19, LFG11, MMMPP19, MH14, RBW14, SHA^{+}15, WHS15, ZzCbC^{+}15]. study [AJP18, ARVACPMN16, AVU^{+}10, BMB^{+}18, BGF^{+}15, BPPR16, BHM^{+}18b, BLHG17, BTNA13, CMMF13, CGKG^{+}17, CBO^{+}15, CMLP19, DP18, DHP^{+}13, Eid12, EPTG19, EHBK18, FDN^{+}11, GKOC^{+}16, GRGGCACG11, GPSA11, GMCL^{+}18, GBR^{+}18, GPBA13, GBR17, HHD^{+}11, HWS^{+}15, HSRT17, HMGH12, HMA^{+}16, Ibá14, INSN14, JPM11, JC18, JvDB^{+}17, KK12, KHMF11, LS19, LBH15, LSC^{+}18, LL14a, LYWZ15, LJB16, LDS^{+}19, LHF16, MKK^{+}16, MCDP18, MLBTM16, MPB^{+}17, MPF^{+}17, MGB10, MBvDN14, MCM12a, MCH11, MDR^{+}19, MM18, MCJD^{+}17, NCP15, OMB18, PBG^{+}13, PAX^{+}17, PKWP16, PLPN12, PBC^{+}16, PEU19, PCG16, PHT15, PHLT16, RTML17, SMRS15, Sau14, SMJB11, SP10a, SHG^{+}15, SIBR^{+}19, SB14, SP15b, TLLZ19, THG^{+}16, TÜM^{+}16, VPC^{+}14, VFdS10, WGM^{+}11, WT16, WMM^{+}15, WKTD13, WFJ19, XLCC19, YRH15, YJSR12, YJC13, ZTC^{+}18]. studying [DDP15, SRMP^{+}18]. stunner [EDG16]. stunning [EDG16, LDR^{+}12]. Sture [HM16a, HM16b]. Sturgeon [KCJ19, WPS19, WRMM18, CGKG^{+}17, HSB19, MLC^{+}19a, YVJ12]. Stylet [LH15]. Sub [BMR^{+}19, HND^{+}17]. Sub-lethal [BMR^{+}19]. sub-structuring [HND^{+}17]. subadult [WRMM18]. subdivision [EvHRF10, VvHE^{+}11]. submerged [AWSLS15]. submergence [Hoc12]. submitted [JBMK14]. subpopulations [CSKB13]. Subsampling [LR19, PJM14]. subsea [BLP^{+}18]. subsequent [BRS19, RR12, SKDO^{+}17]. subset [HPHF15]. subsidies [SRB19]. subsidy [HBC^{+}12]. subsistence [BRW^{+}13]. substrata [BH11]. subtidal [Kni12]. subtilis [FdMCB^{+}19]. subtropical [CLS^{+}19, CJG^{+}19]. Success [PDH^{+}16, BA12a, Hoc12, JBMK13, JP16, LWCC15, WMIDS18]. Successful [BPV^{+}17]. suckleyi [TGM13]. suggested [SOY14]. suggestions [Pun19a]. suggests [CMPR^{+}16, KBOM14]. suitability [GAR19, GMCL^{+}18, HHTT11, LCZ^{+}16, LZX^{+}17, XGT^{+}17, YGZ^{+}18, ZTC^{+}18]. suitable [CFF^{+}16]. suite [PD18]. sulcus [MSV^{+}15]. Sulu [PGAS15]. Sum [PKRC15]. summer [HF14, HFL14, MW11, Mau12, PT19, PBD21, YGA^{+}12]. Sundarbans [TBF^{+}16]. sunfish [HSD^{+}17]. Sunk [Pun18]. superba [HKK18, KWR15, KK15, ZYSZ18, ZDZY19]. superciliosus [SWA19]. Superior [MBJ^{+}16, YAH^{+}13]. supplement [JHL^{+}19]. Supplementing [BFW13]. supply [LBZE14]. supplying [AFAJ12]. support [HKAA18, HHCP13, NBF^{+}19, Pop19, RBGE^{+}10, RCBPK11, SAF^{+}15]. supported [MNAA^{+}14]. supporting [GGS^{+}19, TWSA16]. supports [BST^{+}19]. suppression [HVD^{+}18]. SURBAR [Nee15]. surf [BCPP17]. Surface [BST^{+}17, CMMF13, LKL^{+}12, LCWC17, PGS^{+}11, RTŘ^{+}12, TNKO19]. Surface-induced [BST^{+}17]. surfclam [HMMP18]. SURFs [PE14]. surgery [JBMK14]. surgical [HSB19, MK17, MH14]. surgically [GDW^{+}15, JBMK13, WBD^{+}13]. surmuletus [BCN^{+}13]. surplus [CBO^{+}15, OHLBL16, PS12, WMdS14, WJ17, Zha13]. Survey [Ata11, BNK14, BK15, BBH^{+}14, Cla14, Coo13, CTS^{+}19, DWH10, DQMV19, EHB18, FMH13, FMD19, Har19, HHHWJ18, HHSS17, JAB15, KLWG17, LSC^{+}18, LRM^{+}10, MMAA^{+}18, Nee15, NKW^{+}19, PPG^{+}11, PDBM12, RBT14, RKP^{+}12, RHW^{+}16, SWH17, STJ11, SAB^{+}11, SHM12, SKF^{+}10, VAB^{+}14, WXZ^{+}18, WOPB11, WMPR17, WBB^{+}16, ZTC^{+}18]. survey- [WMPR17]. survey-based [Nee15]. surveying [ZBB17]. surveys [BST^{+}17, BJP^{+}13, BAFP18, CT16, DMK^{+}17, DMQ^{+}18, DQMV19, GHJ13, GBR17, HPP18, HPRB16, JCW12, KJ17, KBT^{+}19, KML11, KVK^{+}17, LMHR19, LHSS18, MOS^{+}18, MMGEH15, MMO^{+}13, ODD16, OTST11, PWB^{+}16, PH12, PGL17, PR18, Pre17, RWRT15, SB15, TBS13, TZW16, TW14, YJSR12]. Survivability [MMR^{+}18, ASN17]. Survival [CFPF^{+}15, DET^{+}16, GSØT15, HBD^{+}16, LCL12, MMK16, PJR17, BAY^{+}12, BHT^{+}19, BHC10, BJKH12, BCPP17, CMS14, CRH^{+}13, CHW^{+}18, DDB^{+}14, DHP^{+}11, DSW^{+}17, EBB16, ERCG10, GH14, HSS18, HRB14, HLSD21, KNSJJ11, KNSSL11, KCJ19, KRM^{+}18, LWSC11, LBCD18, MK17, MWR18, MDH15, MRT18, NNP17, PTC^{+}18, Qui17, ROCC13, RKH^{+}17, RBG^{+}19a, RRS^{+}19, Sau14, SNK19, SBAK17, SMS17, SD13, SSC^{+}12, TGMM13, TGE^{+}18]. survivorship [HAB^{+}10, JNF^{+}19, KCBS13, SHA^{+}15, SWA19]. susceptibility [FEGMTÁ17, PSE15]. sustainability [BAM16, CMB11b, HFO^{+}18, KSS^{+}16, Kos15, LBZE14, MNAA^{+}14, MEM15, OKAKO16, PICK17, Pit14]. Sustainable [KHQH15, PSS14, AHWY13, BdGM^{+}11, DM11, DHFSS18, FAAM15, HH16, HER13, HCZ^{+}16, IIS10, Kni12, MBB13, PS13, RADR11, SPM15b, VH12]. sustained [CPB^{+}12]. Sustaining [Ano13a, CLG^{+}14, WGDB13]. suture [WBD^{+}13]. sutures [WBD^{+}13]. Svedäng [FWG15]. SW [AFR^{+}17, BA12a, DSJ13, GQdÁdMGP^{+}11, RBM^{+}16, SJR^{+}11, TRS12, VSLC^{+}19, CPG^{+}11, MBRDM18, SSH14]. swash [GJRR14]. Swedish [DBK^{+}10, NCP15]. sweep [SHG^{+}16]. sweeps [LWH19, SHG^{+}15]. swept [BBH^{+}14, KML11, MMGEH15, TNKO19]. swift [Pop19]. Swim [RMT^{+}18, Pul17, TdMWC16]. swimbladder [FKM11]. swimmer [JHCT11, LBB^{+}13b]. Swimming [KG10, CMMVRF^{+}19, GSDD11, KDSAH^{+}18, KCJ19, LCH^{+}18, LMM^{+}17, RM10a, SKDO^{+}17, TXZ^{+}17, WAB^{+}16, WGC^{+}19, WGC^{+}21, dKMS15]. Swordfish [PBC^{+}16, RBR^{+}15, AMQ^{+}15, BGA11, CMMF13, CNS11, IB10, LLWC15, LRAP19, WMNC15, XdA19]. swordtip [YAT^{+}19]. Sykes [RADR11]. sympatric [HSB^{+}13, RFPG^{+}16]. Symposium [BW17, Ano10-27]. Sympterygia [IIT17, IIT18]. synagris [SMO^{+}18b]. synchronicity [BGF^{+}15, TCBR17]. Synchronous [VJH17, PUSI10]. synchrony [FCLH19, HBT^{+}16b, KCM19, MYJ^{+}15]. Syngnathidae [ZQZ^{+}17]. Synoptic [GHJ13]. syntheses [BM19a]. Synthesis [PM13, ZMdSC16, MT13, MW13, Cop13]. System [KBSK15, BCP^{+}12, BSSH13, BBHF10, CMP^{+}14, DAJMB16, DPL^{+}16, GRO^{+}18, GMHPV^{+}10, HB12, HB13, KGR^{+}15, LHS^{+}16, LHS^{+}18b, LHS^{+}18c, LFG11, PHM^{+}16a, PWM^{+}16, RIB^{+}10, STY^{+}16, SWB12, Sim15, SDE^{+}18, VBR^{+}18b, AT10]. systems [Bow17, BSM13a, BSM13b, DBSA18, FBKE16, GMS12, PMC^{+}18, PDH^{+}16, RSEERR^{+}16, RM19, SHGL21, SB17, SAF^{+}15, SYJ^{+}11, WHS^{+}18, WKT^{+}11]. t [MMAH15, GKMGO^{+}17, Pop19]. t-bar [MMAH15]. T-ONS [Pop19]. T-RFLP [GKMGO^{+}17]. T0 [THA^{+}14]. T45 [BSM18]. T90 [BHL^{+}16, THA^{+}14]. Tachibana [TOTI10]. tactic [SKFM17]. tactics [Zie12]. Tag [GH15a, GVDMG12, LWY^{+}11, RCF^{+}17, dSMSF15, BAY^{+}12, CGKG^{+}17, CFH14, CFH15, COS15, EHB18, EMSC15, FH15, GMSP19, GBH^{+}19, HSS18, HZT^{+}19, HLNH15a, HLNH15b, KM10, KH15, MK17, PKWP16, ROCC13, SG15, SFB11, SZPV16, ŠBB^{+}19, SMS17, TJJ14, TSS11, VB15, VIP13, WRH^{+}15]. tag-induced [SG15]. tag-integrated [GBH^{+}19]. Tag-linked [RCF^{+}17]. tag-recapture [EMSC15, KM10, TJJ14]. tag-recovery [GBH^{+}19, VB15]. tag-return [EHB18]. tagged [AGB19, BHB^{+}10, CAA^{+}15, DMFR16, LWY^{+}11, MMAH15, SFH^{+}15, WT16]. tagger [COS15]. Tagging [MEM15, PTTC18, RBG^{+}19b, BTT11, BNA^{+}19, BPV^{+}17, BFW13, Ech17, FdSM^{+}16, GMSP19, HF15a, HBT16a, HF14, HLNH15a, HLNH15b, JC18, LH18, LNL^{+}15, NSM^{+}18, PDS^{+}17, ROCC13, SMS17, WMB12, WMB16, WHS15, JvDB^{+}17]. tagging-increment [FdSM^{+}16]. tags [BN12, CBD^{+}14, CAA^{+}14, FBB^{+}17, GDW^{+}15, GBR^{+}18, IIT17, IIT18, JBMK13, JTP^{+}14, LMM^{+}17, MBHGN13, MM11, MMM^{+}10, MSV^{+}19, OPE^{+}17, RIB^{+}10, RTML17, SFH^{+}15, SGN^{+}10, SEG^{+}15, VHS^{+}17, WGC^{+}19, WGC^{+}21]. tailed [CRAVC14, MSGV11, MB16]. Taiwan [CHL11, CLS^{+}11, CDF^{+}13, JLLL11, LCST10, SYL^{+}15, SCL^{+}15, TCS^{+}15, WDL17]. Taiwanese [HL10, Hua15, LT18]. take [MMC13, MMSUAGC14, MMBD13]. taken [FTM^{+}15]. taking [POM^{+}18]. Tam [HRC16]. Tamil [CAR19]. Taming [SB16]. tandem [LSS^{+}17]. tangle [BM19b]. tank [Kim15, NWO^{+}15, SMS^{+}19]. Tanner [Mur19, YRH15]. Tanzania [BK19, RMGK18]. Tanzanian [KDM15]. taper [BSM12]. Taqman [CMI^{+}16]. Target [FCCSA15, GFCK13, KEMM12, SMS^{+}19, BST^{+}17, CHL11, DDB^{+}14, DMK^{+}17, DOO15, FKM11, GSP17a, GWVH^{+}16, GPT^{+}19, JSB19b, KBT^{+}19, LYWZ15, OCL^{+}18, PBR^{+}16, RT14, RSEERRSC15, RSEERRSC16, RSEERR^{+}16, SNK19, SPMS17, ZKSF17, ZSF11]. target-Strength [JSB19b]. Targeted [Nij15, SFBV17, EvHRF10, PCWM16, RSG^{+}11, SPM^{+}15a]. targeting [BBM16, HSRT17, MGR14, SVPC16, WKA13]. targets [DMK^{+}17, TMC15, WFJ19]. tariff [KRC14]. tariff-based [KRC14]. tarpon [AL13b, GBA^{+}18, WW14]. Tasmania [HHTT11, Zie12]. Tasmanian [WWR18]. Taupo [DP18]. taurus [KCBS13]. taxa [GAR19, PCWM16, SHPSK15]. Taxon [PBML^{+}10]. taxonomic [GMHPV^{+}10]. Te [KHQH15]. technical [AAH^{+}16, GPSA11]. technique [BGL15, BAFP18, CFB^{+}17, CMJ^{+}19, DHFSS18, EFW^{+}11, LST^{+}10]. Techniques [HSTK16, BAY^{+}12, ESF^{+}19, GMNPM^{+}13, GMHPV^{+}10, HKRH12, HSB19, HKF^{+}12, KMKK10, MP14, MLT17, MH14, NWO^{+}15, SMdUK16, YGA^{+}12]. technologies [TIGCDL14]. technology [GLB^{+}19, GGM^{+}18, LMSM14, SEG^{+}15]. Tehuantepec [APDJSU11]. Telemetry [PALC19, BMCS19, BCP^{+}12, Ber13, BTB^{+}19, BM19a, ESF^{+}19, EFWC^{+}19, EFW^{+}11, GBB^{+}12, GBA^{+}18, MBF19a, YGA^{+}12]. Teleost [PBML^{+}10, LWCC15, MSN^{+}12, VRS13]. Teleostei [BSD^{+}13]. teleosts [FFG^{+}16]. telephone [TZW16]. temperate [CLS^{+}19, GHM16, GKKK15, HGG13, MLTD17, MBF^{+}17, PWB^{+}16, SDJ^{+}15, SGA^{+}18]. Temperature [FT16, JCA^{+}16, KSH19, PHB19a, SKY^{+}17, SPM^{+}19, BPG^{+}17, CNS11, CDMC18, CCM14, CRH^{+}13, Han16b, HPHR14, HM16a, HM16b, KC11, Kur12, LKL^{+}12, LHLR17, MDG17, MWR13b, PUSI10, SLT^{+}19, SJA17, SFL^{+}14, WAP^{+}12, ZC17]. Temperature-dependent [JCA^{+}16, KSH19, Kur12]. temperatures [GHE^{+}11, IÓS^{+}14, OSMP18, PGS^{+}11, Qui17]. Temporal [HBMC12, MPB^{+}17, Tho19a, Tho19b, WSG16, ACRM12, Ber19, CCA16a, CBF^{+}14, CEM^{+}11, CN18, CAC19, CKC13, CEAP18, EHBK18, FF11, GWB^{+}19, GBEP16, Gun17, HPRB16, HGG13, HBG^{+}19, Kai19, KS17, KNB^{+}19, LCMTM10, LZGL17, MD13, MT19, MWA^{+}15, OMI^{+}14, OGCBMR15, PGW17, PAX^{+}17, PT19, PBC^{+}16, PLSM^{+}16, RRO^{+}19, RvRW^{+}14, SH13a, SMPB17, TLCSG^{+}15, WMNC15, YCYC16, ZDZY19]. Temporally [ASH^{+}17, LSS^{+}16a]. ten [MLT17]. ten-year [MLT17]. tenacity [OKB^{+}16]. tending [CRH^{+}13]. Tentacle [KYO^{+}12]. Tentacle-breakage [KYO^{+}12]. tenuis [Ame12]. Terengganu [INSN14]. term [AVU^{+}10, BHB^{+}10, Ben18, BN12, CAGTS^{+}10, CMS14, CH14, CH15, CKVV16, CGLS12, DDB^{+}14, DCP^{+}16, DUD^{+}18, EMP^{+}10, EFW^{+}11, FSBS11, GPSA11, GD17, GOD19, HC17, HFFdlT15, JC18, Lam13, LSJD11, MK17, MSS^{+}16, MCI^{+}13, MD16, NNS^{+}10, OLK^{+}14, OSO^{+}18, ÖEG^{+}15, VGE11, WDNB^{+}19, WB16, ZCR16]. terminal [HER13]. terraenovae [dJSMROZP^{+}13]. terrapin [CW17]. territorial [AAN^{+}12b, VPSH^{+}17]. Territories [LRY^{+}15, KMMV10]. Test [KOF11, CFBS16, GHM16, MHFK12, MHFH17]. tested [SWC14]. Testing [AFAJ12, CC19, EFW^{+}11, FSG^{+}19, HL11, MGM12, RSEERRSC15, TVCT12, DHF10, HKEJ13, MDH15, Nee15, NWO^{+}15]. Tests [HLJ15, CPC^{+}17, LH18]. tetradactylum [BWN^{+}12]. Tetrapturus [AGB19, BSD^{+}13, WRP10]. Teuthida [SP10a]. Texas [OMB18, CBF^{+}14, FF11]. Thai [WMM^{+}15]. Thaleichthys [HJLW11, MDH15]. thazard [PGAS15]. their [AMGH15, BHG^{+}17, BDB^{+}11, CJCN21, GHA13, HFPL16, HAB^{+}17a, HAB^{+}18a, KBOM14, KvZ12, KLH^{+}15, LW19a, MBHGN13, MK13, MMM^{+}10, ML11, ML17, MSR14, NCP15, PJML15, PBB^{+}18, PW19, SMK^{+}13, SNOD^{+}13, TZW16, WPCO15]. theoretical [MFKS17]. Theory [MCVS14, MCP^{+}16, SPB^{+}19, BSL10, DM16, GLMP18, Han18b, JFT^{+}15, KMKK10, MT19, SVPC16]. Theragra [APD10, BHI14, Koo12, PPLK18, WWH13]. there [LNC^{+}15a, LNC^{+}15b, SLPT17, WHK13]. Thermaikos [KTL11]. thermal [LKL^{+}12, OCN^{+}11]. thermocline [FTN^{+}14]. these [NWO^{+}15]. thickness [GBD^{+}15, HWM^{+}13, OKB^{+}16]. thin [HWS^{+}15]. thin-twined [HWS^{+}15]. Think [OPH15]. Thirty [CLG^{+}14, Gan12]. thornyhead [Ech17]. those [CPH12]. threadfin [BWN^{+}12, HMvHN13, MSN^{+}12, MSWW17]. threat [MCC^{+}17, PEU19]. threatened [Jab18]. Three [BHL12, HFPL16, KMQ^{+}10, LDM^{+}15, BSL10, BCN^{+}13, BB14, CRS^{+}19, CAMFMF18, CRB^{+}18, CKC16, GSØT15, GSP17a, GPS15, GYCL19, HSLB19, HSH15, HSB^{+}13, JBD^{+}18, KA15, LNL^{+}15, LSH13, MMGEMGS10, MBSM15, MJC^{+}14, PC19, RM10a, RvAGB18, THA^{+}14, VJH17, VGE11, WGS17, ZCLB11]. Three-dimensional [KMQ^{+}10, LDM^{+}15, WGS17]. three-parameter [PC19]. three-spined [JBD^{+}18, RvAGB18]. three-spot [RM10a]. thresher [HAB^{+}10, SHA^{+}15, SWA19]. Threshold [Qui17, GFJ19, WAB^{+}16]. thresholds [CHW^{+}18]. thrill [MVML18]. throughout [FBGB16, MAH16, WPS19]. throughput [SRMP^{+}18]. Throwing [SKF^{+}10]. Thunder [FBM^{+}19, MBJ^{+}16]. Thunnus [ASB^{+}16, dSMSF15, ALBŠB^{+}15, Ben18, CFB^{+}17, CHL11, CDL21, CR18, Cor17, CAA^{+}14, CAA^{+}15, FSH^{+}15, FTN^{+}14, Gri10, HSD^{+}17, IAS^{+}17, LKL^{+}12, LRML^{+}15, MFC^{+}13, MMF^{+}17, MMM^{+}10, OSO^{+}18, OAT^{+}16, OTH^{+}19, SMdUK16, SFB11, SFH^{+}15, SHC14, SKT18, SRMP^{+}18, SLL^{+}16, SZ10, SVG^{+}17, UIH^{+}18, UJLC^{+}19, VIFLP17, WKT^{+}13, WMB12, WMB16, WCC^{+}21, XTP^{+}16, XLCMMV19, ZCZ^{+}13, ZCD^{+}12, ZMG^{+}13, ZMG^{+}14, dZB16]. Thymallus [PHB19a]. thynnus [ASB^{+}16, ALBŠB^{+}15, Ben18, CFB^{+}17, Cor17, LRML^{+}15, MMF^{+}17, SLL^{+}16, SVG^{+}17, UJLC^{+}19]. tickler [vMWvOKvB14]. Tidal [TBS13, CLP^{+}19, CM15, CAGVB14, GVSB11, HCWS19, HKM^{+}17, MCD^{+}13]. tides [BK19]. tier [DPL^{+}16]. tiger [HSM^{+}12, KCO14, RLM^{+}17]. tigers [KCBS13]. tilapia [KDdAA^{+}18]. tilefish [PS19]. till [PSJ^{+}19]. Time [BGD11, KRB^{+}15, MCC^{+}17, ZG14, AS10, BGL15, BRGS19, BHG^{+}17, CY12, CNS11, CDB^{+}14, DJW^{+}18, DS13, FSBS11, GIK^{+}16, Goo16, GPT^{+}19, JEJS17, KS17, KRC14, Kur12, KOHF^{+}16, LP18, LCMdSM13, LRM^{+}10, Lyn14, MVML18, MS14, MC10, MWR13b, MS15b, MRT18, NB14, OHSG19, PGL17, PMP^{+}11, RSAB11, ST13, SBL^{+}19, TB10, TW13, TMC15, WT16, vPCWV15, PMC^{+}18]. time-lapse [vPCWV15]. Time-location [ZG14]. Time-PCR [PMC^{+}18]. Time-Resolved [BGD11]. time-saving [BGL15]. time-series [Lyn14]. time-variation [SBL^{+}19, TMC15]. time-varying [DS13, Goo16, LP18, MS14, MRT18, NB14, TB10]. timed [MDS^{+}18]. timelines [RGSAMV19]. timers [PGL17, RTSV15]. times [RSQ18, RBG^{+}19b]. timing [CBMLB13, GMZ^{+}14, GBEP16, MHRD16, MSV^{+}19, OSMP18, PTC^{+}18, SO18, SCF^{+}17a]. timing-of-release [PTC^{+}18]. Timo [Han16b]. Timor [EP17]. tissue [LCWC17, MMMPP19, WBD^{+}13]. tissues [MCP14, SD10]. Title [Ano14c, Ano14d, Ano14e, Ano14f, Ano14g, Ano14h, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano15k, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano16i, Ano16j, Ano16k, Ano16l, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano17g]. Tivela [TXP^{+}16]. tobianus [JvDB^{+}17]. Todarodes [HMGH12, RLVS14, TS10, UK14]. togata [SSH14]. tomography [KRB17]. Tonga [WCM^{+}17]. tonggol [Gri10]. Tonle [TBC^{+}19]. Tonnara [ASL^{+}12]. tonnare [SZR^{+}11]. Too [KSH19, LR19, MLM15]. Tool [MAM^{+}13, BKSF17, Cop13, DLD^{+}16, DBA18, EL18, GLMK12, HKEJ13, HMA^{+}16, JSYS15, KLWG17, LG13, MTV17, MFVPC19, MIÁ^{+}10, RCE^{+}19, SWL^{+}15, SPV19, TSS11, dCJG^{+}15]. toolkit [Mun12a, Mun12b]. tools [BHG^{+}15, FWG11, LPHH11, MREV10, PFVGGM16, TTT16]. toothfish [BBHF10, CNWM12, MDPH17, NBG^{+}17, OCL^{+}18, WT16, YZW^{+}19]. tope [BMdS^{+}15]. topless [EPCB17, KHKM15]. Tor [BMR^{+}19, RADR11]. tori [SMK^{+}13]. Torre [GBPC10]. Torres [BRW^{+}13]. Tortugas [ASB^{+}13, HHM^{+}13, FMA^{+}18]. Total [PBG^{+}13, BMOZ13, Chi11, CCC^{+}14b, Hoc12, JSB19b, JASP11, MDPCSR10, NNS^{+}10, RC17, WMF13, ZRH^{+}11]. Totoaba [MFRJ13]. tourism [KK12]. tourist [MNCPMG15]. tournament [SW19, TWCA15]. tow [FWR10, FO11]. towed [FMD19, GLG^{+}11]. towel [SKF^{+}10]. towing [BLHO18, MMGEH15]. Townsend [AENC14]. tows [TNKO19]. Trace [FMF^{+}17, WW14, Ibá14, LRY^{+}15, MFC^{+}13]. traceability [GKOC^{+}16, Ibá15]. traceable [GRO^{+}18]. Trachurus [ASSL11, BKS^{+}16, FCNB16, FMB^{+}19, GHA13, HBP^{+}18, KKL^{+}16, LCZ^{+}16, MFS^{+}18, MFVPC19, SWH17, TJROFS19, VFFG17, VHS^{+}17, WBI^{+}16, ZZA^{+}14]. tracked [SJA17]. Tracking [AMQ^{+}15, CGKG^{+}17, HZT^{+}19, IÓS^{+}14, LDM^{+}15, ÖLTA19]. tracks [HPB10, RPC11]. Trade [BBOL16, EGK15, RBW14, BR15, BdSFB^{+}18, CBH18, FHK18, MFKS17, Nij15, WP17]. Trade-offs [RBW14, FHK18, WP17]. Tradeoffs [HH14, HKF^{+}12, PKWP16, ANEA19, CAC19, SWG18]. Traditional [BSM15c, ARVACPMN16, BOCÁ^{+}18, CEB^{+}19, GBS19, KRB^{+}15, LYWZ15, PFVGGM16, SPMS17, SZR^{+}11, YJSR12]. traffic [vPB18]. training [Kai19]. traits [BLPF^{+}15, DB12, FT16, FFG^{+}16, KNSSL11, Koo12, KDPMSR10, KM15, OCN^{+}11, RLM^{+}17, ZQZ^{+}17]. trajectories [HZT^{+}19, PDP^{+}11]. trajectory [LLP18]. trammel [ARVACPMN16, BBM^{+}10, PMB^{+}11, YSS^{+}11]. Trang [DFAN12]. trans [EP17]. trans-boundary [EP17]. Transatlantic [CRG^{+}18]. transboundary [Pun18]. transcarbamylase [BKM13]. transducers [SOY14]. transect [Har19]. transects [PLMT^{+}11]. transferable [KPL^{+}07, KPL^{+}09]. transformation [QDG17]. transit [DET^{+}16, LTHRU17]. transition [SZPV16]. Translating [HCZ^{+}16]. translocation [CGG21, GHP^{+}15, GDW^{+}15]. transmission [BPK^{+}17]. transmitter [DMFR16, HSS18, HSB19, WAB^{+}16]. transmitters [CBD^{+}14, DMC^{+}12, MST14, MJL^{+}17, WBD^{+}13]. transmontanus [MLC^{+}19a]. transponder [PTTC18, ROCC13, RIB^{+}10]. transport [GVSB11, GHM16, HBT16a]. transportable [Pop19]. transported [DET^{+}16]. Trap [BGL^{+}13, ASD^{+}18, AOC11, BSS^{+}13, BOCÁ^{+}18, DHF10, FHB^{+}19, GPSF10, JEJS17, KPM^{+}19, LS10, LBB^{+}13b, LOC^{+}11, LCHL11, LCL12, LCF15, PVMG19, RRPG18, RSG^{+}11, RHB16, SGW16, TLLZ19, WW11b]. trapping [LBB^{+}13a]. traps [ARVACPMN16, ASD^{+}18, BBM17, BMH17, BM18a, BM18b, BMH18, CJG^{+}15, FRDC10, HNM^{+}12, KFBL13, LNC^{+}15a, LNC^{+}15b, LCLM15, MM13, PWA10, SGW16, SZR^{+}11]. travel [BHT^{+}19, PAM^{+}18]. travelers [ALR^{+}17]. Traveling [SAK10]. traversing [SBAK17]. Trawl [FWR10, HJLW11, ABD^{+}19, ASN17, AAH^{+}16, AJP18, BSB^{+}15, BSB16, BHPC14, BHL^{+}16, BH13, BSSH13, BK15, BHCS15, BHS16, Bri10, BLHG17, BLHO18, BBK12, BSC12, BSM14, BBS^{+}17, BSM18, BBH^{+}14, CHW^{+}10, CFPF^{+}15, CPP^{+}10, CMP^{+}12, CPS^{+}12, CHS^{+}17, CCT^{+}14, DV11, DL11, DWH10, DDB^{+}14, DRE^{+}17, DBK^{+}10, EPCB17, ERG^{+}11, EHB^{+}21, FCBS15, FBKE16, FPC^{+}13, FHM10, GH14, GPS15, GHJ13, HB15a, HB15b, HLJ15, HB12, HB13, Hen11, HKK18, HSLB19, Ing11, JF13, JČK^{+}12, JRD^{+}13, JČK^{+}21, KML11, KK15, KHKM15, KOF11, LBHP16, LHS^{+}18c, LRM^{+}10, LW13, LW16, LW19a, LWH19, LHF16, MMJ^{+}12, MTH10, MHFK12, MCI^{+}13, MBSM15, MKHK19, MMR^{+}18, MBF^{+}19b, MMO^{+}13, Mur15, NL11, NWL^{+}14, NKW^{+}19, OTA^{+}13, PF15b, PICK17, PPG^{+}11, PWA10, Pit14, PHRH16a, PHRH16b, PDBM12, QEH^{+}11a, QEH^{+}11b, QCZC^{+}19]. trawl [RTŘ^{+}12, RKP^{+}12, RRO^{+}19, ŘJP^{+}12, Rob15, RRW^{+}13, RH13, RJKG12, SDJ^{+}15, SL10, SL11a, SLP^{+}15, Sam11, SNK19, SBS^{+}11, SHG^{+}15, STJ11, SMI17, SHM12, SSHM16, SB17, SKF^{+}10, SYJ^{+}11, SGT^{+}14, TFH^{+}15, TSNRU21, TKF11, THG^{+}16, TSHB19, VFMS16, Wal13, WHF^{+}14, WHP15, YGA^{+}12, ZTC^{+}18, vMWvOKvB14]. trawl-acoustic [SHM12]. trawl-caught [ASN17, CFPF^{+}15, TSHB19]. trawler [CGC13, GSH^{+}15, PAR^{+}12, PMF^{+}16]. trawlers [GLMK12, MKK^{+}16, PCG16, RBM13, SMPB17, VKM^{+}17]. trawling [BSM13a, BSM13b, CDR^{+}11, DCP^{+}16, Foc14, HPB10, KHM^{+}10, LHS^{+}18a, MMK16, QMGP10, RSAB11, SHG^{+}15, WCS^{+}12, YRP^{+}19]. trawls [BRJ^{+}16, BMB10, BSM12, BSM15a, BSM15b, BSM15c, BSM16, CSMJB11, MBK19, NWO^{+}15, OS19, SBLM14, VLSP19, WWH13, YRH15, vMWvOKvB14]. treated [VB15]. treatments [GWB^{+}19]. tree [BNAE13, RCBPK11]. trees [FF11, LCMTM10, MGB10, YAH^{+}13]. trend [BGBCM19, LČ18]. Trends [AWGP11, EP17, GPT^{+}16, JDLF11, LSM^{+}19, NWS^{+}11, PPVP17, PNAL16, ZXH^{+}12, AFOB^{+}19, BMOZ13, BB10, BH13, CGdGT^{+}17, CEM^{+}11, CTS^{+}19, EMP^{+}10, FPC^{+}13, FSBS11, GD17, HBMC12, HSM^{+}12, HF15b, HSH16, KMW^{+}11, KVH15, KY17, LBZE14, LMN^{+}12, MCM12a, NPGT17, SOMO17, WCM^{+}17]. trevalla [FPS11]. Trial [FA19, AW17, GWWW13, PHBU19]. Trialing [MVML18]. trials [HSC19, LHS^{+}16, WTS^{+}17]. Tribe [SPMS17]. tributary [SPH18]. Trichiuridae [TC12]. trigger [TQ17]. triggerfish [BMCS19, RRS^{+}19]. Triglidae [MSV^{+}15]. trigonal [TXP^{+}16]. trimmed [PGP^{+}15, PGP^{+}16]. trimmed- [PGP^{+}16]. trimmed-carcass [PGP^{+}15]. Trinidad [GJSR10]. trip [BNAE13]. triplefin [TRC19]. triploid [JNF^{+}19]. Trisopterus [RDV11, RBV^{+}13]. trituberculatus [LSH13, LCH^{+}18]. troll [GHL^{+}11]. Trophic [AAN^{+}12a, RLVS14, ZXH^{+}12, AHH^{+}12, DDP15, HBG^{+}19, KJS12, LGZC17, MFHK15, MCN16, PGW17, PPLK18, RLSPdR14, SRB19, SRG17, VSM^{+}18, ZCR16]. trophic-level [KJS12].

Tropical [SPV19, BBF^{+}18, CFH14, CFH15, CAC15, CSFCA15, EMSC15, FCCSA15, FH15, GH15a, GBR^{+}18, HNM^{+}12, JDLF11, JDBF15, KMW^{+}11, KvZ12, LCMF^{+}16, LMSM14, LFS^{+}19, MSN^{+}12, Mun12a, Mun12b, Mun18, Pit14, PG14, POV19, SDC^{+}15, TIGCDL14, WDBH16]. Trout [CM18, LSJ17, AVB^{+}16, BGL^{+}13, BR18, CFF^{+}16, Cur18, GHL^{+}11, KNJK^{+}10, KBL^{+}15, LSWR17, MK17, MBJ^{+}16, PMG18, PEU19, ROCC13, RSGR19, SBAK17, SDWG18, SGT^{+}18, VvPM^{+}18, WAP^{+}12, WMIDS18, WB16]. truly [BAL^{+}19]. truncation [CBMLB13, Ste11]. trunculus [DSG^{+}17]. truss [SCJ^{+}11, SJJ^{+}11]. trutta [WB16, AVB^{+}16, BGL^{+}13, Cur18, KNJK^{+}10, PEU19, RSGR19, SBAK17, SDWG18, SGT^{+}18, WB16]. TS [GFK^{+}12, RSEERRSC16]. TS/L [GFK^{+}12]. tshawytscha [CQ14, LKNL16, LW12, MBG^{+}10]. Tsushima [TKF11, YAT^{+}19]. tularosa [PTTC18]. Tuna [ALBŠB^{+}15, CY12, GMZ^{+}14, HF15a, MEM15, ASL^{+}12, ASB^{+}16, ABV16, ALH^{+}17, dSMSF15, ABSI^{+}18, BMEGV^{+}19, Ben18, CFB^{+}17, CFH14, CFH15, CHL11, CR18, Cor17, CAA^{+}14, CB11, DSWB18, DSB^{+}15, DLCO^{+}19, DBA18, EBB16, EHH^{+}15, FH15, FPSS15, FFG^{+}16, FSH^{+}15, FTN^{+}14, Gri10, GSDD11, HSD^{+}17, HCGS19, HE15, HLNH15a, HLNH15b, HL10, HL11, IAS^{+}17, JSYS15, JPM11, KH15, KEH16, KEP^{+}19, LRML^{+}15, LLP18, LNL^{+}15, LvVtB^{+}15, LMSM14, LFS^{+}19, MT13, MMF^{+}17, MST14, MGR13, MGR14, MVMSdS19, MEMdU^{+}17, NvZH^{+}17, OSO^{+}18, OAT^{+}16, OTH^{+}19, PAH^{+}13, RCF^{+}17, RBG^{+}19b, SFB11, SFH^{+}15, SLH^{+}14, SHC14, SKT18, SLL^{+}16, SMNdC^{+}19, SZ10, SVG^{+}17, SZR^{+}11, TXZ^{+}17, TTF^{+}13, TIGCDL14, UIH^{+}18, UJLC^{+}19, VIFLP17, VSM^{+}18, VSLC^{+}19, WMB12, WMB16, WCC^{+}21, XLCMMV19, YvZG^{+}17, ZCZ^{+}13, ZXHK15, ZCD^{+}12, ZMG^{+}13, ZMG^{+}14, SF19]. tunas [EMSC15, GH15a, Nak17a, SDC^{+}15]. tuning [Tho19c, Tho21]. Tunisia [FCNB16, VFMS16]. tunnel [KG10]. turbid [JRD^{+}13, MLT17]. turbidity [PMK^{+}10]. turbine [DMC^{+}12, RSE^{+}14, SBS17, SGB^{+}10]. turbines [RDV11]. turbot [FF10, HBM11, IÓS^{+}14]. TURF [MAM^{+}13]. Turkey [TÜM^{+}16]. Turkish [UP16]. turn [MCC^{+}17]. turned [THA^{+}14]. turnover [LBF^{+}16]. Turtle [WDBH16, ASMD^{+}18, BGBCM19, CSFCA15, Mur15, VVL18]. Turtles [AMAM^{+}17, BAT12, Hua15, LCBD12, Mur11, PPA^{+}15]. Tutuila [OTST11]. twaite [BNA^{+}19, BPV^{+}17]. twelve [LMPM11, TDH19]. twelve-year [TDH19]. twin [FBKE16]. twine [HWM^{+}13, OKB^{+}16]. twined [HWS^{+}15]. twines [LFS^{+}19, WLBB15]. Two [HF19, ŠBMG^{+}16, AdM19, AASF16, AENC14, ARF^{+}17, ACRM12, BEM13, BBT^{+}15, BBT^{+}14, BMdS^{+}15, BTB^{+}19, BBM16, BS10, BB13, CPG^{+}15, CDM^{+}19, CBM^{+}19, DLCK17, EHB^{+}21, EMOLS^{+}15, EvHRF10, FHK18, FKM11, GKKK15, GWM18, GSTL15, GBR^{+}18, HFD^{+}13, HCM^{+}11, JLF^{+}17, JBB^{+}18, KNRJ^{+}15, KBG^{+}14, KWR15, KHS12, Leg14, LL14a, LDM^{+}15, LWZ^{+}19, LMN^{+}12, MLBTM16, MSV^{+}15, MWR13b, MCJD^{+}17, PGJ10, PM11, PMK^{+}10, QWC10, RFC^{+}17, RRM16, RBG^{+}19a, RMB^{+}21, SGSFL19, SCV^{+}12, SBH^{+}16, TWM11, TCAP11, VRMF10, WWN^{+}14, WCS^{+}12, WHC^{+}19, WKA13, dKMS15, dPHS^{+}12]. two-banded [ACRM12]. two-chambered [FKM11]. two-part [TWM11]. two-stage [ARF^{+}17, PM11]. two-step [RMB^{+}21]. type [CFPF^{+}15, CAAB^{+}18, EPHF18, JEJS17, LWCC15, LSS^{+}17, STCE12, SFL^{+}14, SSC^{+}12, WLGS19]. types [ASHH12, ASHH15, BHL12]. typical [ALR^{+}17]. typing [SMI17]. typology [MDR^{+}19]. typus [SNOD^{+}13]. Tyrrhenian [PPG^{+}11, LRM^{+}10, SCV^{+}12].

U.S. [BH13, BSSH13, BMR^{+}18, BK15, GMS^{+}17, GTS^{+}17, GJS^{+}13, HFPL16, HF19, HKB14, HPSG15, KMH^{+}12, LLM12, LKZ^{+}17, LW12, LW16, LWH19, MW11, Mau12, MAH14, Mur11, Mur15, Pun11, SRB^{+}18, Wal13, WP15, WP17, WFJ19]. U.S.A [RCLS18]. Uganda [NC13]. UK [PB11]. ultrasonic [HZT^{+}19, MST14]. Ultrasonography [dCJG^{+}15]. ultraviolet [VVL18]. umbra [PBC^{+}13]. umbrella [BBHF10]. Umbrina [CLS^{+}19]. un-assessed [OOM15]. unassociated [GQ16]. unassociated-sets [GQ16]. uncertain [Eng18, FHK18, HF15b, RGSAMV19]. Uncertainties [HA12]. Uncertainty [YvZG^{+}17, AMHR16, APCF17, CRPAMN17, FPS11, LPM^{+}13, LJH11, MPA^{+}12, MFIG16, Mun18, PH12, PHLT16, PDF^{+}18, QWC10, SHT^{+}13, SP12, WCS^{+}12, ZCR16, ZCD^{+}12]. undatum [AMH19, BMR^{+}18, EHBK18, LBF^{+}16, MLM15, SPM15b, WJ17]. underestimating [OHLBL16]. Underlying [ASN17, PSJ^{+}19]. underpin [NL11]. underpinning [PSE15]. underreporting [BDC^{+}17]. undersized [BBM17, SNK19]. understand [HMD^{+}19, MAS15, PTSC^{+}16]. Understanding [HSL^{+}13, KNB^{+}19, MAA^{+}18, PJML15, RRW^{+}13, THG^{+}16, AW12, BEM13, MVML18, MAM^{+}13]. Underwater [MP14, NWL^{+}14, BSS^{+}13, FHB^{+}19, GMJ^{+}13, GBB^{+}12, GBS19, LFG11, RWRT15, TBS13, WW11b]. unequal [Seu21]. unexploited [BKM19, DAM10, HW11]. Ungwana [MMO^{+}13]. unharmed [BGL^{+}13]. Uniform [TC17, ZXHK15]. uninformed [TC17]. Unintended [CAR19]. unit [AVP^{+}15a, DMOL15, GWB^{+}19, GRB14, HPHF15, HFFdlT15, KY17, LHLR17, MWC15, PLPN12, PGS^{+}11, WKA13]. United [PHM^{+}16a, TAS12, BSS^{+}13, BB18, BT16, DDP^{+}16, GBA^{+}18, HCF^{+}17, HA12, MS15b, NBS15, PF15b, RMB^{+}21, SH10]. units [LCMdSM13, LF13]. univariate [KJKK15, PLPN12]. University [Hol11b]. unlikely [WSMS14]. unlisted [TZW16]. unlocking [PHB^{+}19b]. unmarked [KPF^{+}18]. unobservable [WM11]. unprotected [CMPR^{+}16]. unravel [ESF^{+}19]. Unraveling [HBB^{+}19]. unravelling [RSV^{+}17]. unregulated [CNJ11, MZC^{+}18]. unreliable [Coo13]. unreported [APDI13, LTV^{+}13, MZC^{+}18, PBB11]. unscented [SFB11]. unstocked [CBB^{+}16]. unstressed [SKDO^{+}17]. unsustainable [BTBP18, RHOCB^{+}18, SB15]. Untangling [ASMD^{+}18]. unveils [PF15a]. update [KHGB14]. upper [GPdR11, Ing11, MS15a, ST13, EMOLS^{+}15, IdRH10, OMB18, RvAGB18]. Upper-Paraguay [IdRH10]. upriver [DHP^{+}11]. upstream [LTHRU17]. urban [JDLF11]. urchin [BMB^{+}18, BDM^{+}14, CUL^{+}16, JEJS17, JWC^{+}13, MMD^{+}11, RUMBA15]. urchins [ZCLB11]. Urgent [SMBP^{+}12]. Urophycis [Ame12]. Uroteuthis [YAT^{+}19]. Uruguay [HD12]. Uruguayan [GD17]. USA [Ber13, LS19, Ano13a, ASB^{+}13, BC18, BBV17, CUL^{+}16, GMMW17, HHM^{+}13, MS15a, MAAW^{+}17, MMBD13, OMB18, RvRW^{+}14, ST13]. Use [AB11, FEGMTÁ17, HFPH14, IAS^{+}17, JAJ^{+}16, KS12, LPMM14, Mac13, MREV10, MTV17, MDH15, MM11, PAX^{+}17, RTŘ^{+}12, SVV^{+}14, SW19, YAH^{+}13, ZH10, AHH^{+}12, ABL21, AMV14, BBLD16, CAAB^{+}18, CRW^{+}19, DDP15, ESF^{+}19, GN12, GFK^{+}12, GBH^{+}19, GPBA13, HWS10, HGG13, HBD^{+}16, HCH^{+}19, JASP11, LCBD12, LP18, LPTK19, LFS^{+}19, Lyn14, MMMPP19, ML10, MMM^{+}10, MVMdS^{+}17, OTST11, PKHD16, PGL17, PFVGGM16, PTM^{+}18, PBK^{+}19, PM13, PC19, RKH^{+}17, SMI17, SBT^{+}18, VHS^{+}17, VPSH^{+}17, WRR^{+}16, Whi10, WLBB15]. used [Bri10, CSMJB11, CV14, GPS15, GSH^{+}15, LMSM14, LSL12, MHFH17, PB11, PGJ10, SP12]. useful [BHG^{+}15, Ibá14, KLWG17, TSS11]. Usefulness [CRCMP^{+}14, IIT17, IIT18, RBM^{+}16]. user [Pop19]. users [SMGH17]. uses [KWF12]. Using [BHB^{+}10, BMB^{+}18, BSKT13, CRCMP^{+}14, CBB^{+}16, CCA16b, CN18, DPAB10, FMH13, GBA^{+}18, GDA14, HHD^{+}11, IB10, LCH^{+}18, MCEP14, MH11, MBF19a, MMF^{+}16, MGB10, MDPH17, Nee15, SEG^{+}15, SLCL15, WSMS14, XTP^{+}16, YCQ11, ZCW11, AdM19, ALH^{+}17, AWSLS15, APCF17, BEM13, BPPR16, BHC10, BHCJ12, BCN^{+}13, BK12, Ber13, BDC^{+}17, BTB^{+}19, BAFP18, BB14, BKS18, CAFHG10, CNWM12, CLP^{+}19, CWM12, CGdGT^{+}17, CPP^{+}10, CC19, CJG^{+}15, CC11, Coo13, CTWD15, CBS11, CCGD19, DM16, DQMV19, DHFSS18, DBSA18, DF16, EE19, EHH^{+}11, EWB^{+}18, EPO10, Eth15, EL18, EMSC15, FRDC10, FPM17, FL13, FBY^{+}12, Fit12, FQT17, FGS19, FdMCB^{+}19, FdSM^{+}16, FSBS11, FF11, GKOC^{+}16, GRVN10, Gan12, GMSP19, GBS19, HKEJ13, HHB18, HSTK16, HIH19, HF14, HMA18, HPB10, HF15b, Hor11, HLP15]. using [HVJP^{+}11a, HVJP^{+}11b, HMA^{+}16, Ibá15, JIM^{+}15, JKK^{+}16, JF13, JAB15, JWM^{+}18, JCT^{+}16, JSB19a, JvDB^{+}17, KHM^{+}17, KWM^{+}15, KKK16, KSL^{+}17, KDPT18, KJKK15, KGS18, KKE^{+}12, KGB11, KNB^{+}19, LMPM11, LPMM14, LPM^{+}13, LJH11, LJB16, LWS19a, LCST10, LT18, LDS^{+}19, LQL^{+}15, May10, MGR13, NB14, NST^{+}17, OPR12, PBML^{+}10, PMB^{+}11, PGW17, PSL^{+}19, PAH^{+}13, PJM14, PLM16, PMM^{+}15, PBB11, PDP^{+}11, PS12, RG18, RBT14, RBGE^{+}10, RTML17, RSEERR^{+}16, RM19, RPP^{+}11, RSGR19, SG15, SCC^{+}15, SP10b, SWB12, SJJ^{+}11, SGA17, Seu21, SBLB17, SMS^{+}19, SZPV16, SCV^{+}12, SJA17, Smi14, SGB^{+}10, SHT^{+}13, Sto12, SAWS18, SDE^{+}18, SH13b, SGRS^{+}18, TFH^{+}15, TWM11, TTMC16, TBF^{+}16, TK16, TJMT17, Tho19a, Tho19b, TCS13, TVdLOG^{+}21, TdMWC16, VSM^{+}18, WWN^{+}14, WK10, WRH^{+}15, WMB12, WMB16, WWG^{+}16, WFJ19, WFNC18, WMPR17, WHP15]. using [WvPLW18, YMA12, YA18, ZCC11, ZYW^{+}14, ZCX^{+}18, ZCD^{+}12, ZMdSC16, vPCWV15]. utilisation [CMJ^{+}19]. Utility [BEM^{+}17, BMH18, Ber19, BB14, BBH^{+}12, GBH^{+}19, Rod19, SZ11, WPCO15]. utilization [SFB11]. UV [GMCL^{+}18, WBD^{+}13]. UV-press [GMCL^{+}18].

V [GSH^{+}15]. VA [MAM^{+}13]. VALDUVIS [KSS^{+}16]. Valenciennes [FGCG^{+}17]. valid [MIÁ^{+}10]. validate [FCBS15, IAS^{+}17, KHG^{+}17, Nee15]. validates [SLL^{+}16]. Validation [ASN17, CMJ^{+}19, KBSK15, PHE16, AVB^{+}16, BPRB^{+}14, CAC^{+}14, Eth15, GMNPM^{+}13, KHM^{+}17, LPM^{+}13, MSFŠ^{+}11, MH11, MFRM19, MNGPMdMG10, PBB^{+}15a, SRMP^{+}18, Smi14, ZHDS16]. valuable [RAGP18, SPM^{+}19]. value [ABV16, BMOZ13, BR18, BM19a, BAM16, DMOL15, DHC19, JTP19, LA13, MBHGN13, NAR^{+}17, Pre17, RGSAMV19, SBT^{+}18, WGM^{+}11]. values [BV14, CAAB^{+}18, LGZC17, Mac13, PLGL18]. Valuing [PAM^{+}18]. Varden [LRY^{+}15]. Variability [BdGN^{+}18, MFHK15, MDPCSR10, PTF^{+}18, TNKO19, APL^{+}18, ALB^{+}18, APAFMN10, Ber19, CAGTS^{+}10, DBG19, DHH18, EPO10, FCNB16, FE10, FAAM16, GRPEMEVV17, GSP17a, Gun17, HTHK11, HRUSR13, HF19, HFFdlT15, HBG^{+}19, KWR19, KDPMSR10, KNB^{+}19, LGZC17, Lyn14, MMMPP19, MAA^{+}18, MWA^{+}15, MSV^{+}15, PSM^{+}18, PAX^{+}17, PH12, PSJ^{+}19, RvRW^{+}14, SF19, SP10a, SPS^{+}19, SSS^{+}12, TC16a, TC16b, TKF11, TJROFS19, USH^{+}13, VvHE^{+}11, ZM15, dZB16]. variable [BB19, BSM14, Bro13b, DB12, ESF^{+}19, EHBK18, HFBS11, LBZE17, WKTD13]. variables [FSG^{+}19, MRLD18, MRT17, TXZ^{+}17]. Variation [CCC^{+}14a, FMB^{+}19, KMH^{+}12, MCP^{+}17a, PGP^{+}15, PGP^{+}16, WT16, BK16, BMR^{+}18, CBF^{+}14, DS18, FTJ^{+}16, FMF^{+}17, GMJ^{+}13, GWB^{+}19, GGB^{+}18, GBEP16, Ibá14, KNSPK13, KKL^{+}16, Kni12, KAGO^{+}10, LMLBC13, MTV17, MT19, ML11, MSN^{+}12, OMB18, PWA10, QWMD13, RCF^{+}17, SHR18b, SBL^{+}19, TMC15, TRW19, WRP10, Wri13, ZDZY19]. Variations [WWN^{+}15, YLA15, LLWC15, LSJD11, LFCC15, MSS^{+}16, NNS^{+}10, VCS^{+}19, YGZ^{+}18]. varied [NST^{+}17]. variety [ZQZ^{+}17]. various [DCP^{+}16, MAY12, PGP^{+}15, PGP^{+}16]. Varying [MSR14, ASH^{+}17, BB19, CP11, CP12, DS13, Goo16, KOHF^{+}16, LSS^{+}16a, LP18, LJB16, MS14, MRT18, NB14, TB10]. VAST [Tho19a, Tho19b]. Vector [Tho19a, RBGE^{+}10, RCBPK11, Tho19b]. vehicle [HSY16]. vehicles [AWSLS15]. Velodona [SSH14]. vendace [AD16, WBM17]. Veneridae [MREV10]. Venetian [PBC^{+}13]. Venezuela [GJSR10]. ventless [CJG^{+}15]. vents [AOC11, TLLZ19]. venture [MGR13]. venus [DSJ13, PSM^{+}18]. Verification [GWM18, MLC^{+}19b]. Verifying [HHH15]. verreauxi [BBM17, BM18a, LBB^{+}13a]. versus [BNSS19, CEJI11, GMCL^{+}18, GCRV10, OLS15, THF12]. vertebral [UIH^{+}18]. Vertical [EMMM13, FSH^{+}15, FTN^{+}14, APL^{+}18, ÁCDM^{+}14, GFCK13, JPO^{+}19, JTHT14, MJL14, OÁPP12, RTML17, SAWS18]. very [PCG16, SGT^{+}14]. Vessel [HGG13, WHS^{+}18, BNU^{+}10, DBSA18, EBB16, GM15b, MSRB15, TW14]. vessels [CDL21, DTG13, DFAN12, FBGB16, Kai19, LOdH16, LCMF^{+}16, MMJ^{+}12, MKH16, MGLT16, POV19, RCF^{+}16, WRR^{+}16, WTS^{+}17]. VGAMs [Yee10]. VGLMs [Yee10]. via [HZT^{+}19, KT18a, LRY^{+}15, WMIDS18]. viability [GMD^{+}13]. vicinity [BHL12]. Victoria [KEMM12, EGK15, ETMK13, NST^{+}17, PTW^{+}13]. video [BSS^{+}13, BBH^{+}14, CPG^{+}15, FHB^{+}19, FMD19, GBS19, LNC^{+}15a, LNC^{+}15b, LFG11, MP14, PWB^{+}16, PLMT^{+}11, SAWS18, TBS13]. Vietnam [DFAN12, HRC16, NMH^{+}11, SFBV17]. view [FH17]. viewpoint [Pun19b]. vigour [ASN17]. villosus [DH14, MHRD16, SPS^{+}19]. virens [BHL12, EDG16, PHRH16b, RR12, SPM^{+}19]. virgin [WMP^{+}14]. virginica [BPK^{+}17, CDMC18, MAAW^{+}17, MMAA^{+}18]. Virtual [MT13, ASH^{+}17]. visible [LWY^{+}11, PAX^{+}17]. vision [FTWC18]. visiting [MPR18]. visits [LČ18]. Vistula [WB16]. visual [ALJC19, BKLB14, KS12, PLMT^{+}11, PTTC18, RLFJ11]. visualisation [HBB^{+}12]. vital [TMC15]. vitality [Sto12]. vitellogenic [WTK10]. vitreus [BNSS19, BJZB12, DLCK15, JWM^{+}12, KBSK15, LLM^{+}19]. vitro [SAK^{+}17b]. Vizcaino [OGSNH^{+}17]. VMEs [PPdR^{+}10]. VMS [HBB^{+}12, JSG^{+}15, KS17, RPC11, RPP^{+}11, RCF^{+}16, RMP^{+}18, WHS^{+}18]. VMStools [HBB^{+}12]. voice [HGD12]. vollenhovenii [KAGO^{+}10]. voltage [EDG16, PWM^{+}16]. Volume [Ano10b, Ano10c, Ano10d, Ano11b, Ano11c, Ano11d, Ano21a, KWM^{+}15, TPM15]. voluntary [HKAA18]. Vool [JSYS15]. Voss [BPRB^{+}14, DHFSS18]. vs [BVN^{+}10, BMB10, BSM12, BM19b, BPC^{+}12, CN19, CV14, DRGA12, PGL^{+}13, ZBB17]. vs. [BSM15c, GSDD11]. vulgaris [ACRM12, BPMGP10, CRAP^{+}11, CPG^{+}11, GHUG^{+}15, JPSR^{+}14, LMN^{+}12, MPF^{+}17, MLP^{+}14, PRBGS^{+}10, SD10, SJR^{+}11, VFdS10, VSTS10]. vulgata [FAAM15]. Vulnerability [QWCB11, DLCO^{+}19, FEGMTÁ17, GPT^{+}16, HCK15, JSYS15, MAM^{+}13, OS11, OOM15, TCS13]. vulnerable [BPS^{+}19, GAR19, KRB^{+}15, MMC13, RGSAMV19]. vulpes [MCD^{+}13, SSC^{+}12]. vulpinus [HAB^{+}10, SHA^{+}15].

W [MMNCR^{+}13, UGGR^{+}17, BSB16]. wadeable [MDS^{+}18, STCE12]. wahoo [Zis12]. wait [KS12]. wait-ambushing [KS12]. Walbaum [AB11, AB12]. Wales [Ste11, HKAA18]. Walleye [WWH13, APD10, BNSS19, BHI14, BJZB12, BKM19, DLCK15, JWM^{+}12, KBSK15, Koo12, LLM^{+}19, PPLK18, USH^{+}13, VWHB14, WTF16]. wandering [KCvOE13]. Warm [TKF11, WAT^{+}13]. warming [BPS^{+}19, HMMP18, OSMP18, YC18]. warming-induced [YC18]. warp [MBS^{+}15]. Wars [GWM18, JFT^{+}15]. Washington [Ano13a, CUL^{+}16, RvRW^{+}14, ASD^{+}18, CQ14, KMH^{+}12, LKNL16]. waste [PAR^{+}12]. wastes [CDB^{+}14]. Water [AGB15, ABL21, Ata11, BSSH13, BJKH12, CMMF13, CHL11, CRH^{+}13, CAC^{+}14, DSC19, GHE^{+}11, GSTL15, JHL^{+}19, JKS^{+}15, JKK^{+}16, JAJ^{+}16, Ken14, KKH^{+}17, KG10, KBT^{+}19, LHS^{+}18b, LHS^{+}18a, MJC^{+}14, NL11, PHK13, PHK14, RH13, SFL^{+}14, SYJ^{+}11, TRC19, VFMS16, WCS^{+}12, WMPR17, ZBB17, ZC17]. waterbody [PPH^{+}17]. waters [AMAM^{+}17, AdM19, ASD^{+}18, BWR^{+}18, CM18, CMPP11, CMP^{+}12, CQ14, CDF^{+}13, CM15, CHCS14, CGR^{+}15, CGR^{+}16, EMMM13, EWB^{+}18, ELRH17, GJSR10, GHA13, GBA^{+}18, Gri10, GBEP16, Gun17, HLB^{+}15, HNS12, HPSG15, HMvHN13, IÓS^{+}14, Jab18, JLLL11, KMJ^{+}12, KSKGR15, LW19b, MCC^{+}17, MCD^{+}13, NVS^{+}18, NWS^{+}11, NMLZ17, OAPS13, PBM^{+}19, PBM^{+}18, PFJ19, PLA16, QWMD13, RLM^{+}17, RM10a, RGG14, RU15, SP10a, SGN^{+}10, SNP^{+}12, SSH14, SNOD^{+}13, SSB^{+}18, SDE^{+}18, SSG17, TC16a, TC16b, TOA12, TCS^{+}15, VIFLP17, WGM^{+}11, WHNS13, WAT^{+}13, WJ17, WBB^{+}16, ZZA^{+}14, IGS^{+}12]. waved [AMH19, BMR^{+}18]. Wavelet [PGS^{+}11]. waves [LTHRU17]. way [Bai10, BCMS^{+}18, MPF^{+}17]. Weak [FCLH19]. weakfish [CH11, CH14, CH15]. wealth [VFMS16]. weather [SKF^{+}10, TD19]. web [HPRB16, TWSA16]. webs [FP10]. wedge [DSG^{+}17, UGGR^{+}17]. weight [AL13b, CKC16, DMK^{+}19, Jaw11, KOHF^{+}16, LH15, May10, MMK10, PGP^{+}15, PGP^{+}16, SD13, TRS12]. weight-at-age [Jaw11, KOHF^{+}16]. weight-at-size [DMK^{+}19]. weight-grade [May10]. weight-specific [SD13]. weighted [LJB16]. weighting [Ano17a, Fra17, MCP^{+}17b, MGR13, PDS^{+}17, Pun17, SZPP17, SM17, WMNC15, WM17, XGT^{+}17]. weights [BB14, SKR^{+}12]. Welfare [EGK15, JNE19, VBdB^{+}18]. well [NWO^{+}15, PS12]. Wenatchee [LKNL16]. West [BFL^{+}14, DJF^{+}15, GMS^{+}17, HFPL16, HF19, HKB14, LW16, LWH19, SNOD^{+}13, AFA11, BBM^{+}10, BSF^{+}19, BH13, BKS^{+}16, BMdS^{+}15, BK15, FEGMTÁ17, GFO^{+}16, GVSB11, GJS^{+}13, HBM11, KMH^{+}12, LBV13, LW12, LW13, NBF^{+}19, Pun11, SJJ^{+}11, WFLP11, WP15, WP17, GIW16, MMC13, MLBGV16, SWD^{+}13, TDH19]. Western [CPD^{+}19, GSP17a, GSP17b, IGS^{+}12, JCW12, MGB10, ASB^{+}16, BdGM^{+}11, BdGD^{+}16, BEM13, FH15, GMZ^{+}14, GFM^{+}12, HER13, HPSG15, HLA19, HSH16, HVJP^{+}11a, HVJP^{+}11b, JS15, JPM11, Kai19, KH15, LAIVA^{+}19, LCA^{+}19, LNL^{+}15, LMLBC13, MCI^{+}13, MPB^{+}17, MMM^{+}10, MAH16, MCM12a, PKH^{+}11, PPLK18, PTW^{+}13, SD10, SH13a, SB12, SAWS18, VSM^{+}18, VvHE^{+}11, WRP10, WLGS19, WTS^{+}17, AB11, ALH^{+}17, ABSI^{+}18, ALUFJ^{+}12, BGW^{+}16, Ber13, BLR^{+}16, BSKT13, CMPR^{+}16, CHL11, CJG^{+}19, FWJ^{+}16, HNW^{+}18, HMvHN13, JHCT11, MBvDN14, MEMdU^{+}17, NML12, OH12, RADR11, SMO^{+}18b, WCC^{+}21, ZMG^{+}13, ZMG^{+}14]. wet [CKC16]. Wetland [MCDP18]. whale [HPP18, JRG^{+}18]. whales [BOBP^{+}17, SRB^{+}18, WTS^{+}17]. whaling [PPBM13]. whatever [KLB17]. Whelk [SPM15b, AMH19, BMR^{+}18, EHBK18, LBF^{+}16, MLM15, WJ17]. Where [ASMP^{+}10, DCSW^{+}12, CGM^{+}19]. Which [PHT15, GFK^{+}12, OS19, WHK13]. while [BBM17, HGD19, MM13]. White [Ame12, BFM14, BSD^{+}13, HLSD21, HVJP^{+}11a, HVJP^{+}11b, LJJ^{+}13, MLC^{+}19a, OGSNH^{+}17, PTTC18, WRP10, RCHSN^{+}15]. whitefish [DB12, JWM^{+}12, KOF11, LBZE17, LCF15, RKP^{+}12, RSV^{+}17, VJH17]. whitemouth [CMM^{+}14, HD12, dSSdAdAA17, dVLB^{+}15]. whitespotted [LZX^{+}17]. whitetip [TTF^{+}13]. whiting [BEM^{+}17, CRAVC14, GFO^{+}16, GIW17, GCRV10, Ing11, KBOM14, KCvOE13, MDG17, MAS17, ODD16, PBM^{+}18, PFJ19, STJ11, VKM^{+}17]. Whole [SVN^{+}13, AVP^{+}15a, MWR^{+}13a, PGP^{+}15, PGP^{+}16]. whole- [PGP^{+}15, PGP^{+}16]. whole-lake [AVP^{+}15a]. wholesale [FGRIT^{+}12]. whom [BR18]. wide [JWM^{+}18]. widely [FHK17]. widely-dispersed [FHK17]. Wider [BdGN^{+}18, DSB^{+}10, HSH16]. width [MOA^{+}16, Szu16]. wild [ALBŠB^{+}15, BHL12, BN12, CFB^{+}17, CGKG^{+}17, CSOV14, DS18, GSØT15, HBT16a, HKAA18, IGSJBS^{+}15, KNRJ^{+}15, KDdAA^{+}18, LDR^{+}12, MLC^{+}19a, MMM^{+}10, OLS15, PJR17, PEU19, RDM^{+}15, SBAK17, SGT^{+}18, SAK^{+}17b, TGE^{+}18, UJLC^{+}19, WB16, dEWB^{+}15, dPHS^{+}12]. Willapa [RvRW^{+}14]. William [WDNB^{+}19]. Willingness [MGC^{+}11, BPS^{+}18, SCLS15]. wind [CKVV16, RDV11, SOY14]. wind-induced [SOY14]. Windermere [LEO^{+}11]. windmill [RBV^{+}13]. window [BRJ^{+}16]. windows [MFHK10]. winds [WOPB11]. wing [BSM14]. wing-end [BSM14]. Winnebago [KBSK15]. Winners [GPSA11]. winter [FSH^{+}13, LLM^{+}19, SBC^{+}18, TMM19, YCYC16, ZDZY19]. Wisconsin [BNSS19]. Within [YRP^{+}19, AL13a, BDM^{+}18, BHG^{+}17, BMF^{+}10, BMF^{+}11, CMP^{+}14, EHBK18, EMOLS^{+}15, FMA^{+}18, FCLH19, FMB^{+}19, FLW^{+}19, GSP17a, KTW^{+}13, LMPM11, ML10, PHM16b, Sau14, SBS17, TLCSG^{+}15, TRH16, WM13, WWR18, ZMdSC16]. Within-genus [YRP^{+}19]. within-species [WM13]. without [AVA18, BM18b, SZPV16]. wobbegong [HSB^{+}13]. Wolffish [GBEP16, GH14, Gun17, GSB^{+}19]. Wondering [KCvOE13]. work [LR19]. Working [MCB^{+}18]. world [JSG^{+}15, LSH13, SMD^{+}16]. worldwide [ERL15]. worth [LL14a]. wound [BWK^{+}13]. woven [ÖEG^{+}15]. wrong [DM16, Fra16a].

X [BGL15]. X-ray [BGL15]. Xiphias [AMQ^{+}15, CNS11, IB10, LLWC15, LRAP19, WMNC15, XdA19]. Xray [KRB17].

yabby [SGW16]. Yangtze [JWZ^{+}19, RLJ^{+}12]. Year [AD16, ÁCDM^{+}14, Eid16, FCLH19, Ibá14, KRTG16, KSH19, MBJ^{+}16, MLT17, OALS11, PKO10, PG14, RHOCB^{+}18, TCAP11, TMV16, TDH19]. Year-class [AD16, Eid16, FCLH19, OALS11, TCAP11]. yearling [DET^{+}16]. years [AdM19, CLG^{+}14, FBY^{+}12, FH15, Gan12, JAJ^{+}16, MP14, PPG^{+}11, SJA17]. yellow [CN18, CM15, HBT^{+}16b, LSJD11, LZGL17, VWHB14, WLZ^{+}15, ZYW^{+}14, LSJD11, SLZG18, YSS^{+}11]. yelloweye [BRS19, HRB14, Hoc12, RHB^{+}17]. yellowfin [BBM19, CHL11, DSB^{+}15, FPSS15, MVMSdS19, NKW^{+}19, SFB11, SHC14, VIFLP17, WCC^{+}21, XLCMMV19, ZCZ^{+}13, ZMG^{+}13, ZMG^{+}14]. yellownose [QWCB11]. Yellowstone [LS19]. yellowtail [HHSS17, HCR14, KRMW15]. Yield [DLCK17, GHPJ15, Jaw11, LČ18, MGB10, PBK^{+}19, RvAGB18, SWG18, WJC17]. Yield-per-recruit [DLCK17]. yields [DM11, MBB13, SGC11]. York [KLB17, MS15a, LW19b]. young [ÁCDM^{+}14, APD^{+}18, FCNB16, KSH19, PKO10]. young-of-the-year [ÁCDM^{+}14, KSH19, PKO10]. Yucatan [HFCPSEM15]. Yukon [LRY^{+}15].

Zealand [Rob15, BBB^{+}16, DP18, DLD^{+}16, FDMH16, HB15a, HB15b, HBT16a, HHHWJ18, KHQH15, MTC^{+}17, ODD16, PKHD16, PDH^{+}16, SKH14, SHR^{+}18a]. zero [WMT11]. zero-inflated [WMT11]. zeros [LJH11, SH13b]. zigzag [Har19]. Zonal [Han13, Han18b]. zonatus [KKH^{+}17]. Zone [FPM17, AVB^{+}16, GJRR14, KWCT^{+}18, LGZC17, MMC13, MWSN^{+}13, MMSUAGC14, PBC^{+}13]. zones [BCPP17, DDP15, LL14b]. zoning [SB12]. zoonotic [GGM^{+}18, GRO^{+}18, LSC^{+}18, PRP^{+}18, RGAP18]. zooplankton [GLG^{+}11]. zygaena [SC19].

References

Alos:2013:IPM

[AA13] Josep Alós and Robert Arlinghaus. Impacts of partial marine protected areas on coastal fish communities exploited by recreational angling. Fisheries Research, 137(??):88–96, January 2013. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783612002767.

Alzorriz:2016:QET

[AAH^{+}16] Nekane Alzorriz, Luis Arregi, Bent Herrmann, Manu Sistiaga, John Casey, and Jan Jaap Poos. Questioning the effectiveness of technical measures implemented by the Basque bottom otter trawl fleet: Implications under the EU landing obligation. Fisheries Research, 175(??):116–126, March 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783615301454.

Almeida:2012:TPI

[AAN^{+}12a] David Almeida, Ana Almodóvar, Graciela G. Nicola, Benigno Elvira, and Gary D. Grossman. Trophic plasticity of invasive juvenile largemouth bass Micropterus salmoides in Iberian streams. Fisheries Research, 113(1):153–158, January 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783611003298.

Ayllon:2012:MCC

[AAN^{+}12b] Daniel Ayllón, Ana Almodóvar, Graciela G. Nicola, Irene Parra, and Benigno Elvira. Modelling carrying capacity dynamics for the conservation and management of territorial salmonids. Fisheries Research, 134–136(??):95–103, December 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783612002391.

Abecasis:2013:SMD

[AAOE13] David Abecasis, Pedro Afonso, Ron K. O’Dor, and Karim Erzini. Small MPAs do not protect cuttlefish (Sepia officinalis). Fisheries Research, 147(??):196–201, October 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783613001367.

Aldrin:2019:CIF

[AAS19] M. Aldrin, S. Aanes, and S. Subbey. Comments on incongruous formulations in the SAM (state-space assessment model) model and consequences for fish stock assessment. Fisheries Research, 210(??):224–227, February 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618302224.

Afonso:2016:CMR

[AASF16] Pedro Afonso, David Abecasis, Ricardo Serrão Santos, and Jorge Fontes. Contrasting movements and residency of two serranids in a small Macaronesian MPA. Fisheries Research, 177(??):59–70, May 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615301661.

Aguera:2011:USO

[AB11] Antonio Agüera and Deirdre Brophy. Use of saggital otolith shape analysis to discriminate Northeast Atlantic and Western Mediterranean stocks of Atlantic saury, Scomberesox saurus saurus (Walbaum). Fisheries Research, 110(3):465–471, August 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783611002207.

Aguera:2012:GAA

[AB12] Antonio Agüera and Deirdre Brophy. Growth and age of Atlantic saury, Scomberesox saurus saurus (Walbaum), in the northeastern Atlantic Ocean. Fisheries Research, 131–133(??):60–66, November 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783612002202.

Agnetta:2019:SRP

[ABD^{+}19] Davide Agnetta, Fabio Badalamenti, Giovanni D’Anna, Mauro Sinopoli, Franco Andaloro, Salvatrice Vizzini, and Carlo Pipitone. Sizing up the role of predators on Mullus barbatus populations in Mediterranean trawl and no-trawl areas. Fisheries Research, 213(??):196–203, May 2019. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619300244.

Arkhipkin:2021:UIW

[ABL21] Alexander Arkhipkin, Paul Brickle, and Vladimir Laptikhovsky. The use of island water dynamics by spawning red cod, Salilota australis (Pisces: Moridae) on the Patagonian Shelf (Southwest Atlantic). Fisheries Research, 105(3):156– 162, ???? 2021. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578361000086X.

Alvarez-Berastegui:2018:IRE

[ABSI^{+}18] Diego Alvarez-Berastegui, Sámar Saber, G. Walter Ingram, Lara Díaz-Barroso, Patricia Reglero, David Macías, Salvador García-Barcelona, Josetxu Ortiz de Urbina, Joaquín Tintoré, and Francisco Alemany. Integrating reproductive ecology, early life dynamics and mesoscale oceanography to improve albacore tuna assessment in the Western Mediterranean. Fisheries Research, 208(??):329–338, December 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783618302352.

Adolf:2016:RSA

[ABV16] Steven Adolf, Simon R. Bush, and Sietze Vellema. Reinserting state agency in global value chains: the case of MSC certified skipjack tuna. Fisheries Research, 182(??):79–87, October 2016. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783615301430.

Adkison:2017:INB

[AC17] Milo D. Adkison and Keith R. Criddle. Incorporating nonbaseline characters into genetic mixture analyses. Fisheries Research, 193(??):217–222, September 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617301212.

Alvarez-Colombo:2014:DVS

[ÁCDM^{+}14] G. L. Álvarez-Colombo, C. V. Dato, L. Machinandiarena, F. Castro-Machado, and P. Betti. Daylight vertical segregation of young-of-the-year Argentine hake Merluccius hubbsi: Advances in assessment of juvenile abundance with acoustic methods. Fisheries Research, 160(??):85–95, December 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783614000964.

Alos:2012:STP

[ACRM12] Josep Alós, Miguel Cabanellas-Reboredo, and David March. Spatial and temporal patterns in the movement of adult twobanded sea bream Diplodus vulgaris (Saint-Hilaire, 1817). Fisheries Research, 115–116(??):82–88, March 2012. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783611003559.

Alfaro-Cordova:2017:CMD

[ACSAS^{+}17] E. Alfaro-Cordova, A. Del Solar, J. Alfaro-Shigueto, J. C. Mangel, B. Diaz, O. Carrillo, and D. Sarmiento. Captures of manta and devil rays by small-scale gillnet fisheries in northern Peru. Fisheries Research, 195(??):28–36, November 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783617301674.

Axenrot:2016:YCS

[AD16] Thomas Axenrot and Erik Degerman. Year-class strength, physical fitness and recruitment cycles in vendace (Coregonus albula). Fisheries Research, 173(??):61–69, January 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783615001071.

Armstrong:2013:ASS

[ADH^{+}13] Michael P. Armstrong, Micah J. Dean, William S. Hoffman, Douglas R. Zemeckis, Thomas A. Nies, David E. Pierce, Paul J. Diodati, and Daniel J. McKiernan. The application of small scale fishery closures to protect Atlantic cod spawning aggregations in the inshore Gulf of Maine. Fisheries Research, 141(??):62–69, April 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612002780.

Adebola:2019:CNA

[AdM19] Tunde Adebola and Kim de Mutsert. Compartive network analyses for Nigerian coastal waters using two ecopath models developed for the years 1985 and 2000. Fisheries Research, 213(??):33–41, May 2019. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618303849.

Avigliano:2017:OEF

[ADSV17] Esteban Avigliano, Alejandro Domanico, Sebastián Sánchez, and Alejandra V. Volpedo. Otolith elemental fingerprint and scale and otolith morphometry in Prochilodus lineatus provide identification of natal nurseries. Fisheries Research, 186 (part 1)(??):1–10, February 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783616302387.

Ainsley:2014:CAG

[AENC14] Shaara M. Ainsley, David A. Ebert, Lisa J. Natanson, and Gregor M. Cailliet. A comparison of age and growth of the Bering skate, Bathyraja interrupta (Gill and Townsend, 1897), from two Alaskan large marine ecosystems. Fisheries Research, 154(??):17–25, June 2014. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614000411.

Avadi:2013:LCA

[AF13] Angel Avadí and Pierre Fréon. Life cycle assessment of fisheries: a review for fisheries scientists and managers. Fisheries Research, 143(??):21–38, June 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361300009X.

Aldon:2011:SCC

[AFA11] Ma. E. T. Aldon, A. C. Fermin, and R. F. Agbayani. Sociocultural context of fishers’ participation in coastal resources management in Anini-y, Antique in west central Philippines. Fisheries Research, 107(1–3):112–121, January 2011. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783610002730.

Ashford:2012:TEL

[AFAJ12] Julian R. Ashford, Bettina A. Fach, Alexander I. Arkhipkin, and Cynthia M. Jones. Testing early life connectivity supplying a marine fishery around the Falkland Islands. Fisheries Research, 121–122(??):144–152, June 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612000549.

Alonso-Fernandez:2019:IAT

[AFOB^{+}19] Alexandre Alonso-Fernández, Jaime Otero, Rafael Bañón, José Manuel Campelos, Fernando Quintero, Jorge Ribó, Francisco Filgueira, Luisa Juncal, Fabio Lamas, Asunción Gancedo, and José Molares. Inferring abundance trends of key species from a highly developed small-scale fishery off NE Atlantic. Fisheries Research, 209(??):101–116, January 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783618302467.

Aschenbrenner:2017:AGP

[AFR^{+}17] Alexandre Aschenbrenner, Matheus Oliveira Freitas, Gecely Rodrigues Alves Rocha, Rodrigo Leão de Moura, Ronaldo Bastos Francini-Filho, Carolina Minte-Vera, and Beatrice Padovani Ferreira. Age, growth parameters and fisheries indices for the lane snapper in the Abrolhos Bank, SW Atlantic. Fisheries Research, 194(??):155–163, October 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617301510.

Arkhipkin:2015:WDP

[AGB15] Alexander I. Arkhipkin, Michaël Gras, and Alex Blake. Water density pathways for shelf/slope migrations of squid Illex argentinus in the Southwest Atlantic. Fisheries Research, 172(??):234–242, December 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615300321.

Arostegui:2019:MES

[AGB19] M. C. Arostegui, P. Gaube, and C. D. Braun. Movement ecology and stenothermy of satellite-tagged shortbill spearfish (Tetrapturus angustirostris). Fisheries Research, 215(??):21–26, July 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783619300700.

Agostinho:2016:FAN

[AGS^{+}16] Angelo A. Agostinho, Luiz C. Gomes, Natália C. L. Santos, Jean C. G. Ortega, and Fernando M. Pelicice. Fish assemblages in neotropical reservoirs: Colonization patterns, impacts and management. Fisheries Research, 173(??):26–36, January 2016. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783615001290.

Albores:2019:RBS

[AGSF19] Ignacio Alborés, Carlos García-Soler, and Luis Fernández. Reproductive biology of the slipper lobster Scyllarus arctus in Galicia (NW Spain): Implications for fisheries management. Fisheries Research, 212(??):1–11, April 2019. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578361830345X.

Afonso:2011:FGM

[AHC^{+}11] André S. Afonso, Fábio H. V. Hazin, Felipe Carvalho, José C. Pacheco, Humberto Hazin, David W. Kerstetter, Debra Murie, and George H. Burgess. Fishing gear modifications to reduce elasmobranch mortality in pelagic and bottom longline fisheries off Northeast Brazil. Fisheries Research, 108(2–3):336–343, March 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611000117.

Ardura:2012:FDA

[AHH^{+}12] A. Ardura, J. L. Horreo, E. Hernandez, A. Jardon, I. G. Pola, J. L. Martinez, and E. Garcia-Vazquez. Forensic DNA analysis reveals use of high trophic level marine fish in commercial aquaculture fish meals. Fisheries Research, 115–116(??):115–120, March 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611002803.

Adams:2010:QSS

[AHMS10] Charles F. Adams, Bradley P. Harris, Michael C. Marino, and Kevin D. E. Stokesbury. Quantifying sea scallop bed diameter on Georges Bank with geostatistics. Fisheries Research, 106(3):460–467, December 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610002419.

Al-Humaidhi:2013:LMM

[AHWY13] Alia W. Al-Humaidhi, James A. Wilson, and Thomas H. Young. The local management of migratory stocks: Implications for sustainable fisheries management. Fisheries Research, 141(??):13–23, April 2013. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612001713.

Ageeva:2017:GSR

[AJOE17] Tatiana N. Ageeva, Malcolm Jobling, Ragnar L. Olsen, and Margrethe Esaiassen. Gender-specific responses of mature Atlantic cod (Gadus morhua L.) to feed deprivation. Fisheries Research, 188(??):95–99, April 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616304209.

Alzorriz:2018:LSC

[AJP18] Nekane Alzorriz, Ernesto Jardim, and Jan Jaap Poos. Likely status and changes in the main economic and fishery indicators under the landing obligation: a case study of the Basque trawl fishery. Fisheries Research, 205(??):86–95, September 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783618301036.

Archdale:2011:EAN

[AK11] Miguel Vazquez Archdale and Gunzo Kawamura. Evaluation of artificial and natural baits for the pot fishery of the sand crab Ovalipes punctatus (De Haan, 1833). Fisheries Research, 111(3):159–163, November 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611002517.

Ames:2013:GAS

[AL13a] Edward P. Ames and John Lichter. Gadids and alewives: Structure within complexity in the Gulf of Maine. Fisheries Research, 141(??):70–78, April 2013. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612002809.

Ault:2013:RGF

[AL13b] Jerald S. Ault and Jiangang Luo. A reliable game fish weight estimation model for Atlantic tarpon (Megalops atlanticus). Fisheries Research, 139(??):110–117, March 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612003001.

Adams:2018:SVG

[ALB^{+}18] Grant D. Adams, Robert T. Leaf, Joseph C. Ballenger, Stephen A. Arnott, and Christopher J. McDonough. Spatial variability in the growth of sheepshead (Archosargus probatocephalus) in the Southeast US: Implications for assessment and management. Fisheries Research, 206(??):35–43, October 2018. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783618301279.

Arechavala-Lopez:2015:AWA

[ALBŠB^{+}15] Pablo Arechavala-Lopez, Joseph A. Borg, Tanja Šegvić-Bubić, Paolo Tomassetti, Aytaç Özgül, and Pablo Sanchez-Jerez. Aggregations of wild Atlantic bluefin tuna (Thunnus thynnus L.) at Mediterranean offshore fish farm sites: Environmental and management considerations. Fisheries Research, 164(??):178–184, April 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614003373.

Ailloud:2017:IGE

[ALH^{+}17] Lisa E. Ailloud, Matthew V. Lauretta, Alex R. Hanke, Walter J. Golet, Robert J. Allman, Matthew R. Siskey, David H. Secor, and John M. Hoenig. Improving growth estimates for Western Atlantic bluefin tuna using an integrated modeling approach. Fisheries Research, 191(??):17–24, July 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783617300504.

Albuquerque:2019:VQA

[ALJC19] Cristiano Q. Albuquerque, Luiz Cayky S. Lopes, Andres J. Jaureguizar, and Mario V. Condini. The visual quality of annual growth increments in fish otoliths increases with latitude. Fisheries Research, 220(??):Article 105351, December 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783619302061.

Avigliano:2017:FMT

[ALR^{+}17] Esteban Avigliano, Mathieu Leisen, Rurik Romero, Barbara Carvalho, Gonzalo Velasco, Marcelo Vianna, Fernando Barra, and Alejandra Vanina Volpedo. Fluvio-marine travelers from South America: Cyclic amphidromy and freshwater residency, typical behaviors in Genidens barbus inferred by otolith chemistry. Fisheries Research, 193(??):184–194, September 2017. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783617301169.

Arechavala-Lopez:2012:PED

[ALUFJ^{+}12] P. Arechavala-Lopez, I. Uglem, D. Fernandez-Jover, J. T. Bayle-Sempere, and P. Sanchez-Jerez. Post-escape dispersion of farmed seabream (Sparus aurata L.) and recaptures by local fisheries in the Western Mediterranean Sea. Fisheries Research, 121–122(??):126–135, June 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612000586.

Andrews:2012:ASN

[AM12] Kate I. Andrews and Marc Mangel. Asymptotic size and natural mortality of long-lived fish for data poor stock assessments. Fisheries Research, 127–128(??):45–48, September 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783612001336.

Abdulqader:2017:TOM

[AMAM^{+}17] Ebrahim A. A. Abdulqader, Jeff Miller, Ahmed Al-Mansi, Khaled Al-Abdulkader, Nabil Fita, Hussein Al-Nadhiri, and Lotfi Rabaoui. Turtles and other marine megafauna bycatch in artisanal fisheries in the Saudi waters of the Arabian Gulf. Fisheries Research, 196(??):75–84, December 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617302229.

Ames:2012:WHU

[Ame12] Edward P. Ames. White hake (Urophycis tenuis) in the Gulf of Maine: Population structure insights from the 1920s. Fisheries Research, 114(??):56–65, February 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611002761.

Aidoo:2015:EGE

[AMGH15] Eric N. Aidoo, Ute Mueller, Pierre Goovaerts, and Glenn A. Hyndes. Evaluation of geostatistical estimators and their applicability to characterise the spatial patterns of recreational fishing catch rates. Fisheries Research, 168(??):20– 32, August 2015. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783615001034.

Ashfaq:2019:SSM

[AMH19] Umair Ashfaq, Adam Mugridge, and Bruce G. Hatcher. Size at sexual maturity of waved whelk (Buccinum undatum) on the Eastern Scotian Shelf. Fisheries Research, 212(??):12– 20, April 2019. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783618303345.

Aidoo:2016:EMU

[AMHR16] Eric N. Aidoo, Ute Mueller, Glenn A. Hyndes, and Karina L. Ryan. The effects of measurement uncertainty on spatial characterisation of recreational fishing catch rates. Fisheries Research, 181(??):1–13, September 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361630090X.

Abascal:2015:TBS

[AMQ^{+}15] F. J. Abascal, J. Mejuto, M. Quintans, B. García-Cortés, and A. Ramos-Cartelle. Tracking of the broadbill swordfish, Xiphias gladius, in the central and eastern North Atlantic. Fisheries Research, 162(??):20–28, February 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614002793.

Avigliano:2014:CUO

[AMV14] Esteban Avigliano, Carlos Fernando Riaños Martinez, and Alejandra Vanina Volpedo. Combined use of otolith microchemistry and morphometry as indicators of the habitat of the silverside (Odontesthes bonariensis) in a freshwater– estuarine environment. Fisheries Research, 149(??):55–60, January 2014. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783613002233.

Ayllon:2019:OHR

[ANEA19] Daniel Ayllón, Graciela G. Nicola, Benigno Elvira, and Ana Almodóvar. Optimal harvest regulations under conflicting tradeoffs between conservation and recreational fishery objectives. Fisheries Research, 216(??):47–58, August 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619300864.

Anonymous:2010:ADM

[Ano10a] Anonymous. Alasdair Duncan McIntyre (1926–2010). Fisheries Research, 105(1):1, June 2010. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610001074.

Anonymous:2010:CFRa

[Ano10b] Anonymous. Contents of Fisheries Research, volume 101. Fisheries Research, 101(3):207–208, January 15, 2010. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609002951.

Anonymous:2010:CFRb

[Ano10c] Anonymous. Contents of Fisheries Research, volume 102. Fisheries Research, 102(3):340–341, March 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610000354.

Anonymous:2010:CFRc

[Ano10d] Anonymous. Contents of Fisheries Research, volume 106. Fisheries Research, 106(3):579–582, December 2010. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783610002663.

Anonymous:2010:EABa

[Ano10e] Anonymous. Editorial Advisory Board. Fisheries Research, 101(1–2):iii, January 5, 2010. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783609002859.

Anonymous:2010:EABb

[Ano10f] Anonymous. Editorial Advisory Board. Fisheries Research, 101(3):i, January 15, 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783609002938.

Anonymous:2010:EABc

[Ano10g] Anonymous. Editorial Advisory Board. Fisheries Research, 102(1–2):iii, February 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783609003300.

Anonymous:2010:EBa

[Ano10h] Anonymous. Editorial Board. Fisheries Research, 102(3): i, March 2010. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783610000330.

Anonymous:2010:EBb

[Ano10i] Anonymous. Editorial Board. Fisheries Research, 103(1–3): iii, April 2010. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783610000603.

Anonymous:2010:EBc

[Ano10j] Anonymous. Editorial Board. Fisheries Research, 105(1): iii, June 2010. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783610001207.

Anonymous:2010:EBd

[Ano10k] Anonymous. Editorial Board. Fisheries Research, 105(2): i, July 2010. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783610001463.

Anonymous:2010:EBe

[Ano10l] Anonymous. Editorial Board. Fisheries Research, 106(1):iii, October 2010. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578361000189X.

Anonymous:2010:EBf

[Ano10m] Anonymous. Editorial Board. Fisheries Research, 106(3):i, December 2010. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783610002614.

Anonymous:2010:LRa

[Ano10n] Anonymous. List of referees. Fisheries Research, 104(1–3): iv, May 2010. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783610000950.

Anonymous:2010:LRb

[Ano10o] Anonymous. List of referees. Fisheries Research, 106(2):iv, November 2010. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783610002390.

Anonymous:2010:PO

[Ano10p] Anonymous. Pages 1–116 (October 2010). Fisheries Research, 106(1):??, October 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2010:PJa

[Ano10q] Anonymous. Pages 1–132 (5 January 2010). Fisheries Research, 101(1–2):??, January 5, 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2010:PF

[Ano10r] Anonymous. Pages 1–228 (February 2010). Fisheries Research, 102(1–2):??, February 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2010:PJc

[Ano10s] Anonymous. Pages 1–62 (June 2010). Fisheries Research, 105(1):??, June 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2010:PA

[Ano10t] Anonymous. Pages 1–72 (April 2010). Fisheries Research, 103(1–3):??, April 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2010:PJb

[Ano10u] Anonymous. Pages 133–208 (15 January 2010). Fisheries Research, 101(3):??, January 15, 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2010:PM

[Ano10v] Anonymous. Pages 229–342 (March 2010). Fisheries Research, 102(3):??, March 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2010:PD

[Ano10w] Anonymous. Pages 249–582 (December 2010). Fisheries Research, 106(3):??, December 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2010:PJd

[Ano10x] Anonymous. Pages 63–124 (July 2010). Fisheries Research, 105(2):??, July 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2010:PNa

[Ano10y] Anonymous. Publisher’s note. Fisheries Research, 102(3): iii, March 2010. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783610000470.

Anonymous:2010:PNb

[Ano10z] Anonymous. Publisher’s note. Fisheries Research, 106(1):1, October 2010. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783610002018.

Anonymous:2010:RPC

[Ano10-27] Anonymous. Research presented at the 2009 Cephalopod International Advisory Council Symposium. Fisheries Research, 106(2):??, November 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2011:AR

[Ano11a] Anonymous. Acknowledgement of reviewers. Fisheries Research, 108(1):232–234, February 2011. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611000257.

Anonymous:2011:CFRa

[Ano11b] Anonymous. Contents of Fisheries Research, volume 107. Fisheries Research, 107(1–3):301–302, January 2011. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783610003322.

Anonymous:2011:CFRb

[Ano11c] Anonymous. Contents of Fisheries Research, volume 108. Fisheries Research, 108(2–3):405–407, March 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611000695.

Anonymous:2011:CFRc

[Ano11d] Anonymous. Contents of Fisheries Research, volume 109. Fisheries Research, 109(2–3):384–385, May 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611001457.

Anonymous:2011:EBa

[Ano11e] Anonymous. Editorial Board. Fisheries Research, 107(1–3): iii, January 2011. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783610003280.

Anonymous:2011:EBb

[Ano11f] Anonymous. Editorial Board. Fisheries Research, 108(1):iii, February 2011. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783611000191.

Anonymous:2011:EBc

[Ano11g] Anonymous. Editorial Board. Fisheries Research, 108(2–3): i, March 2011. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783611000622.

Anonymous:2011:EBd

[Ano11h] Anonymous. Editorial Board. Fisheries Research, 109(1): iii, April 2011. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783611000993.

Anonymous:2011:EBe

[Ano11i] Anonymous. Editorial Board. Fisheries Research, 109(2–3): i, May 2011. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783611001408.

Anonymous:2011:EBf

[Ano11j] Anonymous. Editorial Board. Fisheries Research, 110(1): iii, June 2011. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783611001780.

Anonymous:2011:EBg

[Ano11k] Anonymous. Editorial Board. Fisheries Research, 110(2): i, July 2011. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783611002086.

Anonymous:2011:EBh

[Ano11l] Anonymous. Editorial Board. Fisheries Research, 110(3): i, August 2011. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783611002438.

Anonymous:2011:EBi

[Ano11m] Anonymous. Editorial Board. Fisheries Research, 111(1– 2):iii, September 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783611002633.

Anonymous:2011:EBj

[Ano11n] Anonymous. Editorial Board. Fisheries Research, 111(3):i, November 2011. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783611002888.

Anonymous:2011:EBk

[Ano11o] Anonymous. Editorial Board. Fisheries Research, 112 (1–2):iii, December 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783611003043.

Anonymous:2011:PD

[Ano11p] Anonymous. Pages 1–104 (December 2011). Fisheries Research, 112(1–2):??, December 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2011:PS

[Ano11q] Anonymous. Pages 1–130 (September 2011). Fisheries Research, 111(1–2):??, September 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2011:PAa

[Ano11r] Anonymous. Pages 1–216 (April 2011). Fisheries Research, 109(1):??, April 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2011:PJb

[Ano11s] Anonymous. Pages 1–224 (June 2011). Fisheries Research, 110(1):??, June 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2011:PF

[Ano11t] Anonymous. Pages 1–234 (February 2011). Fisheries Research, 108(1):??, February 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2011:PJa

[Ano11u] Anonymous. Pages 1–302 (January 2011). Fisheries Research, 107(1–3):??, January 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2011:PN

[Ano11v] Anonymous. Pages 131–188 (November 2011). Fisheries Research, 111(3):??, November 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2011:PMb

[Ano11w] Anonymous. Pages 217–386 (May 2011). Fisheries Research, 109(2–3):??, May 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2011:PJc

[Ano11x] Anonymous. Pages 225–386 (July 2011). Fisheries Research, 110(2):??, July 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2011:PMa

[Ano11y] Anonymous. Pages 235–408 (March 2011). Fisheries Research, 108(2–3):??, March 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2011:PAb

[Ano11z] Anonymous. Pages 387–484 (August 2011). Fisheries Research, 110(3):??, August 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2012:EBa

[Ano12a] Anonymous. Editorial Board. Fisheries Research, 113(1):iii, January 2012. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783611003614.

Anonymous:2012:EBb

[Ano12b] Anonymous. Editorial Board. Fisheries Research, 115– 116(??):i, March 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783612000100.

Anonymous:2012:EBc

[Ano12c] Anonymous. Editorial Board. Fisheries Research, 117– 118(??):ii, April 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783612000197.

Anonymous:2012:EBd

[Ano12d] Anonymous. Editorial Board. Fisheries Research, 119– 120(??):ii, May 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783612000628.

Anonymous:2012:EBe

[Ano12e] Anonymous. Editorial Board. Fisheries Research, 121– 122(??):ii, June 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783612001129.

Anonymous:2012:EBf

[Ano12f] Anonymous. Editorial Board. Fisheries Research, 123–124 (??):ii, July 2012. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783612000707.

Anonymous:2012:EBg

[Ano12g] Anonymous. Editorial Board. Fisheries Research, 125– 126(??):iii, August 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016578361200149X.

Anonymous:2012:EBh

[Ano12h] Anonymous. Editorial Board. Fisheries Research, 127–128 (??):i, September 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016578361200183X.

Anonymous:2012:EBi

[Ano12i] Anonymous. Editorial Board. Fisheries Research, 129– 130(??):i, October 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783612002263.

Anonymous:2012:EBj

[Ano12j] Anonymous. Editorial Board. Fisheries Research, 131–133 (??):i, November 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783612002421.

Anonymous:2012:EBk

[Ano12k] Anonymous. Editorial Board. Fisheries Research, 134–136 (??):i, December 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783612002901.

Anonymous:2012:PMa

[Ano12l] Anonymous. Pages 1–120 (March 2012). Fisheries Research, 115–116(??):1–120, March 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2012:PD

[Ano12m] Anonymous. Pages 1–134 (December 2012). Fisheries Research, 134–136(??):1–134, December 2012. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2012:PO

[Ano12n] Anonymous. Pages 1–152 (October 2012). Fisheries Research, 129–130(??):1–152, October 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2012:PMb

[Ano12o] Anonymous. Pages 1–156 (May 2012). Fisheries Research, 119–120(??):1–156, May 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2012:PJb

[Ano12p] Anonymous. Pages 1–162 (June 2012). Fisheries Research, 121–122(??):1–162, June 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2012:PJa

[Ano12q] Anonymous. Pages 1–190 (January 2012). Fisheries Research, 113(1):??, January 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2012:PS

[Ano12r] Anonymous. Pages 1–194 (September 2012). Fisheries Research, 127–128(??):1–194, September 2012. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2012:PA

[Ano12s] Anonymous. Pages 1–330 (August 2012). Fisheries Research, 125–126(??):1–330, August 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2012:PN

[Ano12t] Anonymous. Pages 1–84 (November 2012). Fisheries Research, 131–133(??):1–84, November 2012. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2013:AFS

[Ano13a] Anonymous. American Fisheries Society Special Session: Assessing the Role of Marine Protected Areas in Restoring, Sustaining, and/or Enhancing Fisheries. September 6, 2011. Seattle, Washington, USA. Fisheries Research, 144(??):111– 116, July 2013. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783613000751.

Anonymous:2013:EBa

[Ano13b] Anonymous. Editorial Board. Fisheries Research, 137(??): i, January 2013. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783612003098.

Anonymous:2013:EBb

[Ano13c] Anonymous. Editorial Board. Fisheries Research, 139(??): i, March 2013. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783612003608.

Anonymous:2013:EBc

[Ano13d] Anonymous. Editorial Board. Fisheries Research, 140(??):i, February 2013. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783613000271.

Anonymous:2013:EBd

[Ano13e] Anonymous. Editorial Board. Fisheries Research, 143(??): i, June 2013. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783613000957.

Anonymous:2013:EBe

[Ano13f] Anonymous. Editorial Board. Fisheries Research, 144(??): ii, July 2013. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783613001124.

Anonymous:2013:EBf

[Ano13g] Anonymous. Editorial Board. Fisheries Research, 145(??): ii, August 2013. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783613001264.

Anonymous:2013:EBg

[Ano13h] Anonymous. Editorial Board. Fisheries Research, 146(??):ii, September 2013. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783613001409.

Anonymous:2013:EBh

[Ano13i] Anonymous. Editorial Board. Fisheries Research, 147(??): ii, October 2013. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783613001938.

Anonymous:2013:EBi

[Ano13j] Anonymous. Editorial Board. Fisheries Research, 148(??):i, November 2013. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783613002579.

Anonymous:2013:PS

[Ano13k] Anonymous. Pages 1–120 (September 2013). Fisheries Research, 146(??):1–120, September 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2013:PA

[Ano13l] Anonymous. Pages 1–136 (August 2013). Fisheries Research, 145(??):1–136, August 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2013:PJa

[Ano13m] Anonymous. Pages 1–140 (January 2013). Fisheries Research, 137(??):1–140, January 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2013:PM

[Ano13n] Anonymous. Pages 1–144 (March 2013). Fisheries Research, 139(??):1–144, March 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2013:PN

[Ano13o] Anonymous. Pages 1–146 (November 2013). Fisheries Research, 148(??):1–146, November 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2013:PF

[Ano13p] Anonymous. Pages 1–154 (February 2013). Fisheries Research, 140(??):1–154, February 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2013:PJb

[Ano13q] Anonymous. Pages 1–204 (June 2013). Fisheries Research, 143(??):1–204, June 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2013:PO

[Ano13r] Anonymous. Pages 1–454 (October 2013). Fisheries Research, 147(??):1–454, October 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2014:EBa

[Ano14a] Anonymous. Editorial Board. Fisheries Research, 149(??): i, January 2014. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578361300266X.

Anonymous:2014:EBb

[Ano14b] Anonymous. Editorial Board. Fisheries Research, 150(??):i, February 2014. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783613002774.

Anonymous:2014:FTPa

[Ano14c] Anonymous. Full title page (Editorial Board). Fisheries Research, 151(??):i, March 2014. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614000137.

Anonymous:2014:FTPb

[Ano14d] Anonymous. Full title page (Editorial Board). Fisheries Research, 154(??):ii, June 2014. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614001027.

Anonymous:2014:FTPc

[Ano14e] Anonymous. Full title page (Editorial Board). Fisheries Research, 155(??):ii, July 2014. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614001337.

Anonymous:2014:FTPd

[Ano14f] Anonymous. Full title page (Editorial Board). Fisheries Research, 158(??):ii, October 2014. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614001192.

Anonymous:2014:FTPe

[Ano14g] Anonymous. Full title page (Editorial Board). Fisheries Research, 159(??):ii, November 2014. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614002033.

Anonymous:2014:FTPf

[Ano14h] Anonymous. Full title page (Editorial Board). Fisheries Research, 160(??):i, December 2014. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614002318.

Anonymous:2014:PF

[Ano14i] Anonymous. Pages 1–108 (February 2014). Fisheries Research, 150(??):1–108, February 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2014:PN

[Ano14j] Anonymous. Pages 1–114 (November 2014). Fisheries Research, 159(??):1–114, November 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2014:PMb

[Ano14k] Anonymous. Pages 1–122 (May 2014). Fisheries Research, 153(??):1–122, May 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2014:PS

[Ano14l] Anonymous. Pages 1–192 (September 2014). Fisheries Research, 157(??):1–192, September 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2014:PJc

[Ano14m] Anonymous. Pages 1–198 (July 2014). Fisheries Research, 155(??):1–198, July 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2014:PMa

[Ano14n] Anonymous. Pages 1–212 (March 2014). Fisheries Research, 151(??):1–212, March 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2014:PJb

[Ano14o] Anonymous. Pages 1–224 (June 2014). Fisheries Research, 154(??):1–224, June 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2014:PA

[Ano14p] Anonymous. Pages 1–38 (August 2014). Fisheries Research, 156(??):1–38, August 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2014:PJa

[Ano14q] Anonymous. Pages 1–94 (January 2014). Fisheries Research, 149(??):1–94, January 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2015:FTPa

[Ano15a] Anonymous. Full title page (Editorial Board). Fisheries Research, 162(??):iii, February 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614003233.

Anonymous:2015:FTPb

[Ano15b] Anonymous. Full title page (Editorial Board). Fisheries Research, 163(??):ii, March 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614003506.

Anonymous:2015:FTPc

[Ano15c] Anonymous. Full title page (Editorial Board). Fisheries Research, 164(??):i, April 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361500020X.

Anonymous:2015:FTPd

[Ano15d] Anonymous. Full title page (Editorial Board). Fisheries Research, 165(??):i, May 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783615000478.

Anonymous:2015:FTPe

[Ano15e] Anonymous. Full title page (Editorial Board). Fisheries Research, 166(??):ii, June 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361500065X.

Anonymous:2015:FTPf

[Ano15f] Anonymous. Full title page (Editorial Board). Fisheries Research, 167(??):i, July 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783615001150.

Anonymous:2015:FTPg

[Ano15g] Anonymous. Full title page (Editorial Board). Fisheries Research, 168(??):i, August 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361500137X.

Anonymous:2015:FTPh

[Ano15h] Anonymous. Full title page (Editorial Board). Fisheries Research, 169(??):i, September 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615001599.

Anonymous:2015:FTPi

[Ano15i] Anonymous. Full title page (Editorial Board). Fisheries Research, 170(??):i, October 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615002040.

Anonymous:2015:FTPj

[Ano15j] Anonymous. Full title page (Editorial Board). Fisheries Research, 171(??):ii, November 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615002143.

Anonymous:2015:FTPk

[Ano15k] Anonymous. Full title page (Editorial Board). Fisheries Research, 172(??):i, December 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615002337.

Anonymous:2015:PM

[Ano15l] Anonymous. Pages 1–132 (May 2015). Fisheries Research, 165(??):1–132, May 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2015:PO

[Ano15m] Anonymous. Pages 1–236 (October 2015). Fisheries Research, 170(??):1–236, October 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2015:PA

[Ano15n] Anonymous. Pages 1–332 (April 2015). Fisheries Research, 164(??):1–332, April 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2015:PJb

[Ano15o] Anonymous. Pages 1–338 (July 2015). Fisheries Research, 167(??):1–338, July 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2015:PJa

[Ano15p] Anonymous. Pages 1–408 (January 2015). Fisheries Research, 161(??):1–408, January 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2015:PD

[Ano15q] Anonymous. Pages 1–452 (December 2015). Fisheries Research, 172(??):1–452, December 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2015:PS

[Ano15r] Anonymous. Pages 1–60 (September 2015). Fisheries Research, 169(??):1–60, September 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2015:PF

[Ano15s] Anonymous. Pages 1–74 (February 2015). Fisheries Research, 162(??):1–74, February 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2015:PAA

[Ano15t] Anonymous. Pages A1-A2, 1-116 (August 2015). Fisheries Research, 168(??):A1–, August 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2015:P

[Ano15u] Anonymous. Preface. Fisheries Research, 166(??):1–3, June 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783615000430.

Anonymous:2015:PN

[Ano15v] Anonymous. Publisher’s note. Fisheries Research, 168(??): A1, August 2015. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783615001411.

Anonymous:2016:FTPa

[Ano16a] Anonymous. Full title page (Editorial Board). Fisheries Research, 173(??):ii, January 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615002428.

Anonymous:2016:FTPb

[Ano16b] Anonymous. Full title page (Editorial Board). Fisheries Research, 174(??):i, February 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615002659.

Anonymous:2016:FTPc

[Ano16c] Anonymous. Full title page (Editorial Board). Fisheries Research, 175(??):i, March 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615003021.

Anonymous:2016:FTPd

[Ano16d] Anonymous. Full title page (Editorial Board). Fisheries Research, 176(??):i, April 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616000047.

Anonymous:2016:FTPe

[Ano16e] Anonymous. Full title page (Editorial Board). Fisheries Research, 177(??):i, May 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783616300340.

Anonymous:2016:FTPf

[Ano16f] Anonymous. Full title page (Editorial Board). Fisheries Research, 178(??):ii, June 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616300753.

Anonymous:2016:FTPg

[Ano16g] Anonymous. Full title page (Editorial Board). Fisheries Research, 179(??):i, July 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783616301047.

Anonymous:2016:FTPh

[Ano16h] Anonymous. Full title page (Editorial Board). Fisheries Research, 180(??):ii, August 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616301175.

Anonymous:2016:FTPi

[Ano16i] Anonymous. Full title page (Editorial Board). Fisheries Research, 181(??):i, September 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616301618.

Anonymous:2016:FTPj

[Ano16j] Anonymous. Full title page (Editorial Board). Fisheries Research, 182(??):ii, October 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616301813.

Anonymous:2016:FTPk

[Ano16k] Anonymous. Full title page (Editorial Board). Fisheries Research, 183(??):i, November 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616302442.

Anonymous:2016:FTPl

[Ano16l] Anonymous. Full title page (Editorial Board). Fisheries Research, 184(??):ii, December 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616302880.

Anonymous:2016:O

[Ano16m] Anonymous. Obituary. Fisheries Research, 173(??):viii, January 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783615002477.

Anonymous:2016:PA

[Ano16n] Anonymous. Pages 1–106 (April 2016). Fisheries Research, 176(??):1–106, April 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2016:PMa

[Ano16o] Anonymous. Pages 1–148 (March 2016). Fisheries Research, 175(??):1–148, March 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2016:PMb

[Ano16p] Anonymous. Pages 1–164 (May 2016). Fisheries Research, 177(??):1–164, May 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2016:PS

[Ano16q] Anonymous. Pages 1–254 (September 2016). Fisheries Research, 181(??):1–254, September 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2016:PF

[Ano16r] Anonymous. Pages 1–308 (February 2016). Fisheries Research, 174(??):1–308, February 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2016:PJ

[Ano16s] Anonymous. Pages 1–362 (July 2016). Fisheries Research, 179(??):1–362, July 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2016:PN

[Ano16t] Anonymous. Pages 1–634 (November 2016). Fisheries Research, 183(??):1–634, November 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2017:DW

[Ano17a] Anonymous. Data weighting. Fisheries Research, 192(??): ??, August 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2017:FTPa

[Ano17b] Anonymous. Full title page (Editorial Board). Fisheries Research, 185(??):i, January 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616303496.

Anonymous:2017:FTPb

[Ano17c] Anonymous. Full title page (Editorial Board). Fisheries Research, 186 (part 1)(??):i, February 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616303630.

Anonymous:2017:FTPc

[Ano17d] Anonymous. Full title page (Editorial Board). Fisheries Research, 187(??):i, March 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616304246.

Anonymous:2017:FTPd

[Ano17e] Anonymous. Full title page (Editorial Board). Fisheries Research, 188(??):i, April 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617300103.

Anonymous:2017:FTPe

[Ano17f] Anonymous. Full title page (Editorial Board). Fisheries Research, 189(??):i, May 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783617300371.

Anonymous:2017:FTPf

[Ano17g] Anonymous. Full title page (Editorial Board). Fisheries Research, 190(??):i, June 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617300681.

Anonymous:2017:IFCd

[Ano17h] Anonymous. Inside front cover — Editorial Board page. Fisheries Research, 194(??):ifc, October 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617301856.

Anonymous:2017:IFCe

[Ano17i] Anonymous. Inside front cover — Editorial Board page. Fisheries Research, 195(??):ifc, November 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617302357.

Anonymous:2017:IFCf

[Ano17j] Anonymous. Inside front cover — Editorial Board page. Fisheries Research, 196(??):ifc, December 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617302473.

Anonymous:2017:IFCa

[Ano17k] Anonymous. Inside front cover — Editorial Board page/ cover image legend if applicable. Fisheries Research, 191(??):ifc, July 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783617300954.

Anonymous:2017:IFCb

[Ano17l] Anonymous. Inside front cover — Editorial Board page/cover image legend if applicable. Fisheries Research, 192(??):ifc, August 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783617301273.

Anonymous:2017:IFCc

[Ano17m] Anonymous. Inside front cover — Editorial Board page/cover image legend if applicable. Fisheries Research, 193(??):ifc, September 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S016578361730156X.

Anonymous:2017:PD

[Ano17n] Anonymous. Pages 1–170 (December 2017). Fisheries Research, 196(??):1–170, December 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2017:PA

[Ano17o] Anonymous. Pages 1–182 (April 2017). Fisheries Research, 188(??):1–182, April 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2017:PJb

[Ano17p] Anonymous. Pages 1–196 (June 2017). Fisheries Research, 190(??):1–196, June 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2017:PJa

[Ano17q] Anonymous. Pages 1–200 (January 2017). Fisheries Research, 185(??):1–200, January 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2017:PJc

[Ano17r] Anonymous. Pages 1–222 (July 2017). Fisheries Research, 191(??):1–222, July 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2017:PO

[Ano17s] Anonymous. Pages 1–222 (October 2017). Fisheries Research, 194(??):1–222, October 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2017:PMa

[Ano17t] Anonymous. Pages 1–226 (March 2017). Fisheries Research, 187(??):1–226, March 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2017:PN

[Ano17u] Anonymous. Pages 1–246 (November 2017). Fisheries Research, 195(??):1–246, November 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2017:PS

[Ano17v] Anonymous. Pages 1–262 (September 2017). Fisheries Research, 193(??):1–262, September 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2017:PF

[Ano17w] Anonymous. Pages 1–406 (February 2017). Fisheries Research, 186 (part 1)(??):1–406, February 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2017:PMb

[Ano17x] Anonymous. Pages 1–94 (May 2017). Fisheries Research, 189(??):1–94, May 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2018:EBa

[Ano18a] Anonymous. Editorial Board. Fisheries Research, 199(??): ii, March 2018. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783617303478.

Anonymous:2018:EBb

[Ano18b] Anonymous. Editorial Board. Fisheries Research, 200(??): ii, April 2018. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783618300080.

Anonymous:2018:EBc

[Ano18c] Anonymous. Editorial Board. Fisheries Research, 201(??): ii, May 2018. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783618300432.

Anonymous:2018:EBd

[Ano18d] Anonymous. Editorial Board. Fisheries Research, 202(??): ii, June 2018. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783618300614.

Anonymous:2018:EBe

[Ano18e] Anonymous. Editorial Board. Fisheries Research, 203(??): ii, July 2018. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783618300675.

Anonymous:2018:EBf

[Ano18f] Anonymous. Editorial Board. Fisheries Research, 204(??): ii, August 2018. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783618301188.

Anonymous:2018:EBg

[Ano18g] Anonymous. Editorial Board. Fisheries Research, 205(??):ii, September 2018. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783618301358.

Anonymous:2018:EBh

[Ano18h] Anonymous. Editorial Board. Fisheries Research, 206(??): ii, October 2018. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783618301693.

Anonymous:2018:EBi

[Ano18i] Anonymous. Editorial Board. Fisheries Research, 207(??):ii, November 2018. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783618302030.

Anonymous:2018:EBj

[Ano18j] Anonymous. Editorial Board. Fisheries Research, 208(??):ii, December 2018. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783618302595.

Anonymous:2018:IFCa

[Ano18k] Anonymous. Inside front cover — Editorial Board page. Fisheries Research, 197(??):ifc, January 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617302771.

Anonymous:2018:IFCb

[Ano18l] Anonymous. Inside front cover — Editorial Board page. Fisheries Research, 198(??):ifc, February 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617303028.

Anonymous:2018:PAa

[Ano18m] Anonymous. Pages 1–116 (April 2018). Fisheries Research, 200(??):1–116, April 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2018:PJ

[Ano18n] Anonymous. Pages 1–130 (January 2018). Fisheries Research, 197(??):1–130, January 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2018:PMb

[Ano18o] Anonymous. Pages 1–152 (May 2018). Fisheries Research, 201(??):1–152, May 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2018:PS

[Ano18p] Anonymous. Pages 1–158 (September 2018). Fisheries Research, 205(??):1–158, September 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2018:PN

[Ano18q] Anonymous. Pages 1–204 (November 2018). Fisheries Research, 207(??):1–204, November 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2018:PF

[Ano18r] Anonymous. Pages 1–224 (February 2018). Fisheries Research, 198(??):1–224, February 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2018:PMa

[Ano18s] Anonymous. Pages 1–272 (March 2018). Fisheries Research, 199(??):1–272, March 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2018:PO

[Ano18t] Anonymous. Pages 1–310 (October 2018). Fisheries Research, 206(??):1–310, October 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2018:PD

[Ano18u] Anonymous. Pages 1–346 (December 2018). Fisheries Research, 208(??):1–346, December 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2018:PAb

[Ano18v] Anonymous. Pages 1–458 (August 2018). Fisheries Research, 204(??):1–458, August 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2019:D

[Ano19a] Anonymous. December 2019. Fisheries Research, 220(??): ??, December 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2019:EBa

[Ano19b] Anonymous. Editorial Board. Fisheries Research, 209(??): ii, January 2019. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783618303096.

Anonymous:2019:EBb

[Ano19c] Anonymous. Editorial Board. Fisheries Research, 210(??):ii, February 2019. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783618303151.

Anonymous:2019:EBc

[Ano19d] Anonymous. Editorial Board. Fisheries Research, 211(??): ii, March 2019. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783618303655.

Anonymous:2019:EBd

[Ano19e] Anonymous. Editorial Board. Fisheries Research, 212(??): ii, April 2019. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783619300311.

Anonymous:2019:EBe

[Ano19f] Anonymous. Editorial Board. Fisheries Research, 213(??): ii, May 2019. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783619300475.

Anonymous:2019:EBf

[Ano19g] Anonymous. Editorial Board. Fisheries Research, 214(??): ii, June 2019. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783619300670.

Anonymous:2019:EBg

[Ano19h] Anonymous. Editorial Board. Fisheries Research, 215(??): ii, July 2019. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783619301018.

Anonymous:2019:EBh

[Ano19i] Anonymous. Editorial Board. Fisheries Research, 216(??): ii, August 2019. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783619301341.

Anonymous:2019:EBi

[Ano19j] Anonymous. Editorial Board. Fisheries Research, 217(??):ii, September 2019. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783619301407.

Anonymous:2019:EBj

[Ano19k] Anonymous. Editorial Board. Fisheries Research, 218(??): ii, October 2019. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783619301766.

Anonymous:2019:EBk

[Ano19l] Anonymous. Editorial Board. Fisheries Research, 219 (??):Article 105360, November 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783619302152.

Anonymous:2019:EBl

[Ano19m] Anonymous. Editorial Board. Fisheries Research, 220 (??):Article 105391, December 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783619302462.

Anonymous:2019:N

[Ano19n] Anonymous. November 2019. Fisheries Research, 219(??): ??, November 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2019:PJc

[Ano19o] Anonymous. Pages 1–150 (July 2019). Fisheries Research, 215(??):1–150, July 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2019:PJb

[Ano19p] Anonymous. Pages 1–182 (June 2019). Fisheries Research, 214(??):1–182, June 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2019:PAa

[Ano19q] Anonymous. Pages 1–208 (April 2019). Fisheries Research, 212(??):1–208, April 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2019:PAb

[Ano19r] Anonymous. Pages 1–212 (August 2019). Fisheries Research, 216(??):1–212, August 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2019:PMb

[Ano19s] Anonymous. Pages 1–226 (May 2019). Fisheries Research, 213(??):1–226, May 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2019:PF

[Ano19t] Anonymous. Pages 1–230 (February 2019). Fisheries Research, 210(??):1–230, February 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2019:PO

[Ano19u] Anonymous. Pages 1–268 (October 2019). Fisheries Research, 218(??):1–268, October 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2019:PJa

[Ano19v] Anonymous. Pages 1–270 (January 2019). Fisheries Research, 209(??):1–270, January 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2019:PMa

[Ano19w] Anonymous. Pages 1–350 (March 2019). Fisheries Research, 211(??):1–350, March 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2019:PN

[Ano19x] Anonymous. Publisher’s note. Fisheries Research, 219 (??):Article 105304, November 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783619301511.

Anonymous:2021:CFR

[Ano21a] Anonymous. Contents of Fisheries Research, volume 105. Fisheries Research, 105(3):261–262, ???? 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610001694.

Anonymous:2021:EB

[Ano21b] Anonymous. Editorial Board. Fisheries Research, 105(3): i, ???? 2021. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783610001670.

Anonymous:2021:PAS

[Ano21c] Anonymous. Pages 125–262 (August–September 2010). Fisheries Research, 105(3):??, ???? 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Arana:2011:EVT

[AOC11] Patricio M. Arana, Juan Carlos Orellana, and Álvaro De Caso. Escape vents and trap selectivity in the fishery for the juan Fernández rock lobster (Jasus frontalis), Chile. Fisheries Research, 110(1):1–9, June 2011. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611001305.

Alos:2010:IVS

[APAFMN10] Josep Alós, Miquel Palmer, Alexandre Alonso-Fernández, and Beatriz Morales-Nin. Individual variability and sexrelated differences in the growth of Diplodus annularis (Linnaeus, 1758). Fisheries Research, 101(1–2):60–69, January 5, 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783609002513.

Archambault:2016:AMC

[APB^{+}16] Benoit Archambault, Olivier Le Pape, Loic Baulier, Youen Vermard, Matthieu Véron, and Etienne Rivot. Adultmediated connectivity affects inferences on population dynamics and stock assessment of nursery-dependent fish populations. Fisheries Research, 181(??):198–213, September 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783616300911.

Akselrud:2017:EMS

[APCF17] Caitlin I. Allen Akselrud, André E. Punt, and Lee Cronin-Fine. Exploring model structure uncertainty using a general stock assessment framework: the case of Pacific cod in the Eastern Bering Sea. Fisheries Research, 193(??):104–120, September 2017. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783617300814.

Amar:2010:IEF

[APD10] Z. Teresa A’mar, André E. Punt, and Martin W. Dorn. Incorporating ecosystem forcing through predation into a management strategy evaluation for the Gulf of Alaska walleye pollock (Theragra chalcogramma) fishery. Fisheries Research, 102(1–2):98–114, February 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783609002781.

Avigliano:2018:SSC

[APD^{+}18] Esteban Avigliano, Jorge Pisonero, Alejandro Dománico, Natalia Silva, Sebastián Sánchez, and Alejandra Vanina Volpedo. Spatial segregation and connectivity in young and adult stages of Megaleporinus obtusidens inferred from otolith elemental signatures: Implications for management. Fisheries Research, 204(??):239–244, August 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361830081X.

Addis:2013:IRF

[APDI13] Dustin T. Addis, William F. Patterson, Michael A. Dance, and G. Walter Ingram. Implications of reef fish movement from unreported artificial reef sites in the northern Gulf of Mexico. Fisheries Research, 147(??):349–358, October 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783613001756.

Alejo-Plata:2011:SRS

[APDJSU11] Carmen Alejo-Plata, Píndaro Díaz-Jaimes, and Isaías H. Salgado-Ugarte. Sex ratios, size at sexual maturity, and spawning seasonality of dolphinfish (Coryphaena hippurus) captured in the Gulf of Tehuantepec, Mexico. Fisheries Research, 110(1):207–216, June 2011. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611001573.

Arita:2013:DEI

[APHL13] Shawn Arita, Minling Pan, Justin Hospital, and PingSun Leung. The distributive economic impacts of Hawaii’s commercial fishery: a SAM analysis. Fisheries Research, 145(??):82– 89, August 2013. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783613000489.

Abascal:2018:SVB

[APL^{+}18] F. J. Abascal, T. Peatman, B. Leroy, S. Nicol, K. Schaefer, D. W. Fuller, and J. Hampton. Spatiotemporal variability in bigeye vertical distribution in the Pacific Ocean. Fisheries Research, 204(??):371–379, August 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618300870.

Aramayo:2015:LSC

[Ara15] Víctor Aramayo. Lower selectivity can help heavily exploited fish populations. Fisheries Research, 172(??):261–264, December 2015. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783615300370.

Alemany:2017:BTS

[ARF^{+}17] Juliette Alemany, Etienne Rivot, Eric Foucher, Joël Vigneau, and Jean-Paul Robin. A Bayesian two-stage biomass model for stock assessment of data-limited species: an application to cuttlefish (Sepia officinalis) in the English Channel. Fisheries Research, 191(??):131–143, July 2017. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783617300772.

Andrade:2019:AGB

[ARMLC19] I. Andrade, D. Rosa, R. Muñoz-Lechuga, and R. Coelho. Age and growth of the blue shark (Prionace glauca) in the Indian Ocean. Fisheries Research, 211(??):238–246, March 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578361830331X.

Amengual-Ramis:2016:AFS

[ARVACPMN16] J. F. Amengual-Ramis, M. Vázquez-Archdale, C. Cánovas-Pérez, and B. Morales-Nin. The artisanal fishery of the spiny lobster Palinurus elephas in Cabrera National Park, Spain: Comparative study on traditional and modern traps with trammel nets. Fisheries Research, 179(??):23–32, July 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783616300224.

Arellano:2010:PAF

[AS10] Carlota Estrella Arellano and Gordon Swartzman. The Peruvian artisanal fishery: Changes in patterns and distribution over time. Fisheries Research, 101(3):133–145, January 15, 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783609002288.

Ault:2013:ACR

[ASB^{+}13] Jerald S. Ault, Steven G. Smith, James A. Bohnsack, Jiangang Luo, Natalia Zurcher, David B. McClellan, Tracy A.

Ziegler, David E. Hallac, Matt Patterson, Michael W. Feeley, Benjamin I. Ruttenberg, John Hunt, Dan Kimball, and Billy Causey. Assessing coral reef fish population and community changes in response to marine reserves in the Dry Tortugas, Florida, USA. Fisheries Research, 144(??):28– 37, July 2013. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783612003037.

Addis:2016:RSA

[ASB^{+}16] Piero Addis, Marco Secci, Cecilia Biancacci, Daniela Loddo, Danila Cuccu, Francesco Palmas, and Andrea Sabatini. Reproductive status of Atlantic bluefin tuna, Thunnus thynnus, during migration off the coast of Sardinia (western Mediterranean). Fisheries Research, 181(??):137–147, September 2016. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783616301138.

Antonelis:2018:BRS

[ASD^{+}18] Kyle Antonelis, James Selleck, Joan Drinkwin, Annika Saltman, Dan Tonnes, and Jeffrey June. Bycatch of rockfish in spot prawn traps and estimated magnitude of trap loss in Washington waters of the Salish Sea. Fisheries Research, 208(??):105–115, December 2018. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361830184X.

Allen:2017:TVN

[ASH^{+}17] Shanae D. Allen, William H. Satterthwaite, David G. Hankin, Diana J. Cole, and Michael S. Mohr. Temporally varying natural mortality: Sensitivity of a virtual population analysis and an exploration of alternatives. Fisheries Research, 185(??):185–197, January 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616302958.

Afonso:2012:SBM

[ASHH12] André S. Afonso, Renato Santiago, Humberto Hazin, and Fábio H. V. Hazin. Shark bycatch and mortality and hook bite-offs in pelagic longlines: Interactions between hook types and leader materials. Fisheries Research, 131–133(??):9–14, November 2012. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0165783612002056. See corrigendum [ASHH15].

Afonso:2015:CSB

[ASHH15] André S. Afonso, Renato Santiago, Humberto Hazin, and Fábio H. V. Hazin. Corrigendum to “Shark bycatch and mortality and hook bite-offs in pelagic longlines: Interactions between hook types and leader materials” [Fish. Res. 131–133 (2015) 9–14]. Fisheries Research, 167(??):278–279, July 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615000971. See [ASHH12].

Addis:2012:AAB

[ASL^{+}12] Piero Addis, Marco Secci, Ivan Locci, Angelo Cau, and Andrea Sabatini. Analysis of Atlantic bluefin tuna catches from the last Tonnara in the Mediterranean Sea: 1993–2010. Fisheries Research, 127–128(??):133–141, September 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578361200166X.

Alfaro-Shigueto:2018:UIN

[ASMD^{+}18] Joanna Alfaro-Shigueto, Jeffrey C. Mangel, Jodie Darquea, Miguel Donoso, Andres Baquero, Phil D. Doherty, and Brendan J. Godley. Untangling the impacts of nets in the southeastern Pacific: Rapid assessment of marine turtle bycatch to set conservation priorities in small-scale fisheries. Fisheries Research, 206(??):185–192, October 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618301103.

Alfaro-Shigueto:2010:WSC

[ASMP^{+}10] Joanna Alfaro-Shigueto, Jeffrey C. Mangel, Mariela Pajuelo, Peter H. Dutton, Jeffrey A. Seminoff, and Brendan J. Godley. Where small can have a large impact: Structure and characterization of small-scale fisheries in Peru.

Fisheries Research, 106(1):8–17, October 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610001347.

Albalat:2017:VVI

[ASN17] Amaya Albalat, Simon Sinclair, and Douglas Neil. Validation of a vigour index for trawl-caught Norway lobsters (Nephrops norvegicus) destined for the live market: Underlying links to both physiological condition and survivability. Fisheries Research, 191(??):25–29, July 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361730053X.

Ault:2018:TFI

[ASR^{+}18] Jerald S. Ault, Steven G. Smith, Benjamin L. Richards, Annie J. Yau, Brian J. Langseth, Joseph M. O’Malley, Christofer H. Boggs, Michael P. Seki, and Gerard T. Di-Nardo. Towards fishery-independent biomass estimation for Hawaiian Islands deepwater snappers. Fisheries Research, 208(??):321–328, December 2018. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618302339.

Ashford:2011:OCI

[ASSL11] Julian Ashford, Rodolfo Serra, Juan Carlos Saavedra, and Jaime Letelier. Otolith chemistry indicates large-scale connectivity in Chilean jack mackerel (Trachurus murphyi), a highly mobile species in the Southern Pacific Ocean. Fisheries Research, 107(1–3):291–299, January 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610002973.

Arguelles:2010:NIB

[AT10] Juan Argüelles and Ricardo Tafur. New insights on the biology of the jumbo squid Dosidicus gigas in the Northern Humboldt Current System: Size at maturity, somatic and reproductive investment. Fisheries Research, 106(2):185–192, November 2010. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783610001359.

Atabak:2011:SNF

[Ata11] N. Atabak. Survey on natural feeding of juvenile Cynoglossus arel and Solea elongata fishes (Cynoglossidae and Soleidae) in the northwest of Persian Gulf coastal water. Fisheries Research, 108(1):9–14, February 2011. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610002985.

Ayza:2011:ESO

[ATG11] Olga Ayza, Victor M. Tuset, and José A. González. Estimation of size at onset of sexual maturity and growth parameters in Norway lobster (Nephrops norvegicus) off the Portuguese coast. Fisheries Research, 108(1):205–208, February 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783610003000.

Alves:2018:ACG

[AVA18] Diego Corrêa Alves, Lilian Paula Vasconcelos, and Angelo Antonio Agostinho. Age composition and growth without age data: a likelihood-based model. Fisheries Research, 204(??):361–370, August 2018. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618300493.

Aymes:2016:AVK

[AVB^{+}16] J. C. Aymes, M. Vignon, E. Beall, F. Guéraud, and P. Gaudin. Age validation of the Kerguelen Islands brown trout, Salmo trutta L., and selection of the otolith optimal zone for investigating chronological data series. Fisheries Research, 176(??):22–29, April 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361530148X.

Alexander:2015:EWL

[AVP^{+}15a] T. J. Alexander, P. Vonlanthen, G. Periat, F. Degiorgi, J. C. Raymond, and O. Seehausen. Estimating whole-lake fish catch per unit effort. Fisheries Research, 172(??):287–302, December 2015. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783615300333.

Alexander:2015:EGP

[AVP^{+}15b] Timothy J. Alexander, Pascal Vonlanthen, Guy Periat, François Degiorgi, Jean-Claude Raymond, and Ole Seehausen. Evaluating gillnetting protocols to characterize lacustrine fish communities. Fisheries Research, 161(??):320–329, January 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783614002549.

Andersen:2010:CIS

[AVU^{+}10] Bo Sølgaard Andersen, Youen Vermard, Clara Ulrich, Trevor Hutton, and Jan-Jaap Poos. Challenges in integrating shortterm behaviour in a mixed-fishery management strategies evaluation frame: a case study of the North Sea flatfish fishery. Fisheries Research, 102(1–2):26–40, February 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609002689.

Armstrong:2012:DUF

[AW12] Michael J. Armstrong and Peter R. Witthames. Developments in understanding of fecundity of fish stocks in relation to egg production methods for estimating spawning stock biomass. Fisheries Research, 117–118(??):35–47, April 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783611000026.

Abbott:2017:RBM

[AW17] Joshua K. Abbott and Daniel Willard. Rights-based management for recreational for-hire fisheries: Evidence from a policy trial. Fisheries Research, 196(??):106–116, December 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578361730228X.

Anticamara:2011:GFE

[AWGP11] J. A. Anticamara, R. Watson, A. Gelchu, and D. Pauly. Global fishing effort (1950–2010): Trends, gaps, and implications. Fisheries Research, 107(1–3):131–136, January 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783610002754.

Ajemian:2015:RAF

[AWSLS15] Matthew J. Ajemian, Jennifer Jarrell Wetz, Brooke Shipley-Lozano, and Gregory W. Stunz. Rapid assessment of fish communities on submerged oil and gas platform reefs using remotely operated vehicles. Fisheries Research, 167(??):143– 155, July 2015. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783615000569.

Bacha:2012:ICD

[BA12a] Mahmoud Bacha and Rachid Amara. Inter-cohort differences in growth, condition and feeding of juvenile anchovy (Engraulis encrasicolus) in the Gulf of Béjaia (Algerian coast, SW Mediterranean): Implications for recruitment success. Fisheries Research, 129–130(??):73–81, October 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783612001968.

Baldwin:2012:EEF

[BA12b] Liz Baldwin and Miran Aprahamian. An evaluation of electric fishing for assessment of resident eel in rivers. Fisheries Research, 123–124(??):4–8, July 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611003407.

Benejam:2012:FCC

[BAB^{+}12] Lluís Benejam, Carles Alcaraz, Josep Benito, Nuno Caiola, Frederic Casals, Alberto Maceda-Veiga, Adolf de Sostoa, and Emili García-Berthou. Fish catchability and comparison of four electrofishing crews in Mediterranean streams. Fisheries Research, 123–124(??):9–15, July 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611003511.

Byrne:2019:PRI

[BAD19] Conor Byrne, Sveinn Agnarsson, and Brynhildur Davidsdottir. Profit and rent in the Icelandic harvesting sector. Fisheries Research, 220(??):Article 105349, December 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619302048.

Bova:2018:LRR

[BAFP18] Christopher S. Bova, Shankar Aswani, Matthew W. Farthing, and Warren M. Potts. Limitations of the random response technique and a call to implement the ballot box method for estimating recreational angler compliance using surveys. Fisheries Research, 208(??):34–41, December 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783618301875.

Baio:2010:SMW

[Bai10] Andrew Baio. Show me the way: Inclination towards governance attributes in the artisanal fisheries of Sierra Leone. Fisheries Research, 102(3):311–322, March 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361000007X.

Belleggia:2019:HTO

[BAL^{+}19] Mauro Belleggia, Nadia Marina Alves, Melisa Mariana Leyton, Gustavo Álvarez-Colombo, Brenda Temperoni, Diego Giberto, and Claudia Bremec. Are hakes truly opportunistic feeders? A case of prey selection by the Argentine hake Merluccius hubbsi off southwestern Atlantic. Fisheries Research, 214(??):166–174, June 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619300529.

Brown:2016:RFC

[BAM16] Simon Brown, David J. Agnew, and Will Martin. On the road to fisheries certification: the value of the objections procedure in achieving the MSC sustainability standard. Fisheries Research, 182(??):136–148, October 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615301120.

Braga:2018:CBS

[BAOP18] Heitor O. Braga, Ulisses M. Azeiteiro, Henrique M. F. Oliveira, and Miguel A. Pardal. Conserving Brazilian sardine: Fisher’s attitudes and knowledge in the Marine Extractive Reserve of Arraial do Cabo, Rio de Janeiro State, Brazil. Fisheries Research, 204(??):402–411, August 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783618300936.

Benhardouze:2012:ICS

[BAT12] Wafae Benhardouze, Mustapha Aksissou, and Manjula Tiwari. Incidental captures of sea turtles in the driftnet and longline fisheries in northwestern Morocco. Fisheries Research, 127–128(??):125–132, September 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612001300.

Barbour:2012:CCB

[BAY^{+}12] Andrew B. Barbour, Aaron J. Adams, Tanner Yess, Donald C. Behringer, and R. Kirby Wolfe. Comparison and costbenefit analysis of PIT tag antennae resighting and seine-net recapture techniques for survival analysis of an estuarinedependent fish. Fisheries Research, 121–122(??):153–160, June 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783612000446.

Baum:2010:ISP

[BB10] Julia K. Baum and Wade Blanchard. Inferring shark population trends from generalized linear mixed models of pelagic longline catch and effort data. Fisheries Research, 102(3):229–239, March 2010. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783609003087.

Bustamante:2013:IRB

[BB13] Carlos Bustamante and Michael B. Bennett. Insights into the reproductive biology and fisheries of two commercially exploited species, shortfin mako (Isurus oxyrinchus) and blue shark (Prionace glauca), in the south-east Pacific Ocean. Fisheries Research, 143(??):174–183, June 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783613000507.

Britton:2014:AUU

[BB14] J. Robert Britton and Rachel Blackburn. Application and utility of using otolith weights in the ageing of three flatfish species. Fisheries Research, 154(??):147–151, June 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783614000459.

Brites:2017:FMR

[BB17] Nuno M. Brites and Carlos A. Braumann. Fisheries management in random environments: Comparison of harvesting policies for the logistic model. Fisheries Research, 195(??):238–246, November 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617301996.

Bacheler:2018:DSD

[BB18] Nathan M. Bacheler and Joseph C. Ballenger. Decadalscale decline of scamp (Mycteroperca phenax) abundance along the southeast United States Atlantic coast. Fisheries Research, 204(??):74–87, August 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618300341.

Brites:2019:FMR

[BB19] Nuno M. Brites and Carlos A. Braumann. Fisheries management in randomly varying environments: Comparison of constant, variable and penalized efforts policies for the Gompertz model. Fisheries Research, 216(??):196–203, August 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783619300803.

Bellquist:2019:EDD

[BBA^{+}19] Lyall Bellquist, Sabrina Beyer, Morgan Arrington, Jordan Maeding, Alayna Siddall, Paul Fischer, John Hyde, and Nicholas C. Wegner. Effectiveness of descending devices to mitigate the effects of barotrauma among rockfishes (Sebastes spp.) in California recreational fisheries. Fisheries Research, 215(??):44–52, July 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619300621.

Breen:2016:SMN

[BBB^{+}16] Paul A. Breen, Andrew R. Branson, Nokome Bentley, Vivian Haist, Malcolm Lawson, Paul J. Starr, Daryl R. Sykes, and D’Arcy N. Webber. Stakeholder management of the New Zealand red rock lobster (Jasus edwardsii) fishery. Fisheries Research, 183(??):530–538, November 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615301569.

Broadhurst:2011:PRM

[BBC11] Matt K. Broadhurst, Paul A. Butcher, and Brian R. Cullis. Post-release mortality of angled sand mullet (Myxus elongatus: Mugilidae). Fisheries Research, 107(1–3):272–275, January 2011. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783610002948.

Baker:2018:NLA

[BBF^{+}18] Ronald Baker, Michael Bradley, Stefano Freddi, Katya Abrantes, Adam Barnett, and Marcus Sheaves. Non-lethal aging of tropical catch-and-release sport fishery species. Fisheries Research, 207(??):110–117, November 2018. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783618301814.

Butcher:2012:ABA

[BBH^{+}12] Paul A. Butcher, Matt K. Broadhurst, Karina C. Hall, Brian R. Cullis, and Shane R. Raidal. Assessing barotrauma among angled snapper (Pagrus auratus) and the utility of release methods. Fisheries Research, 127–128(??):49–55, September 2012. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783612001427.

Bryan:2014:QVA

[BBH^{+}14] David R. Bryan, Keith L. Bosley, Allan C. Hicks, Melissa A. Haltuch, and W. Waldo Wakefield. Quantitative video analysis of flatfish herding behavior and impact on effective area swept of a survey trawl. Fisheries Research, 154(??):120– 126, June 2014. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783614000472.

Brown:2010:EIU

[BBHF10] Judith Brown, Paul Brickle, Sarah Hearne, and Gemma French. An experimental investigation of the ‘umbrella’ and ‘Spanish’ system of longline fishing for the Patagonian toothfish (Dissostichus eleginoides) in the Falkland Islands: Implications for stock assessment and seabird by-catch. Fisheries Research, 106(3):404–412, December 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610002225.

Broadhurst:2012:EDB

[BBK12] Matt K. Broadhurst, Craig P. Brand, and Steven J. Kennelly. Evolving and devolving bycatch reduction devices in an Australian penaeid-trawl fishery. Fisheries Research, 113(1):68–75, January 2012. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611002980.

Bergmann:2016:LCU

[BBLD16] Charles Bergmann, James Barbour, Lionel LaForce, and William B. Driggers. Line cutter for use when releasing large marine organisms caught on longline gear. Fisheries Research, 177(??):124–127, May 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616300194.

Buckland:2017:SFS

[BBLS17] Amanda Buckland, Ronald Baker, Neil Loneragan, and Marcus Sheaves. Standardising fish stomach content analysis: the importance of prey condition. Fisheries Research, 196(??):126–140, December 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617302175.

Baeta:2010:EBT

[BBM^{+}10] Filipa Baeta, Marisa Batista, Anabela Maia, Maria José Costa, and Henrique Cabral. Elasmobranch bycatch in a trammel net fishery in the Portuguese west coast. Fisheries Research, 102(1–2):123–129, February 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578360900280X.

Broadhurst:2016:RED

[BBM16] Matt K. Broadhurst, Paul B. Butcher, and Russell B. Millar. Relative efficiencies and durabilities of recreational hoop- and lift-nets targeting two Australian portunids. Fisheries Research, 179(??):115–123, July 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616300418.

Broadhurst:2017:EGR

[BBM17] Matt K. Broadhurst, Paul A. Butcher, and Russell B. Millar. Escape gaps in recreational panulirid traps: Reducing catches of undersized Sagmariasus verreauxi while increasing fishing power for legal sizes. Fisheries Research, 189(??): 55–61, May 2017. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783617300061.

Broadhurst:2019:FYB

[BBM19] Matt K. Broadhurst, Paul A. Butcher, and Russell B. Millar. Fate of yellowfin bream, Acanthopagrus australis after ingesting offset and/or inline barbed and barbless J and circle hooks. Fisheries Research, 211(??):183–190, March 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578361830328X.

Bailey:2016:FFT

[BBOL16] Megan Bailey, Simon Bush, Peter Oosterveer, and Laksmi Larastiti. Fishers, fair trade, and finding middle ground. Fisheries Research, 182(??):59–68, October 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615301508.

Broadhurst:2017:EKO

[BBS^{+}17] Matt K. Broadhurst, Cheslav Balash, David J. Sterling, Russell B. Millar, and Shinsuke Matsubara. Effects of knot orientation on the height and drag of a penaeid trawl. Fisheries Research, 186 (part 1)(??):48–54, February 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616302557.

Betti:2014:LGH

[BBT^{+}14] Paola Betti, Daniel R. Brown, Brenda Temperoni, Laura Machinandiarena, and Martín D. Ehrlich. Larval growth of hake (Merluccius hubbsi) in the Patagonian shelf: Analysis of two reproductive seasons. Fisheries Research, 160(??):69–76, December 2014. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783614000587.

Bernardi:2015:HPP

[BBT^{+}15] Cristian Bernardi, Luciana Baggiani, Erica Tirloni, Simone Stella, Fabio Colombo, Vittorio Maria Moretti, and Patrizia Cattaneo. Hemolymph parameters as physiological biomarkers in American lobster (Homarus americanus) for monitoring the effects of two commercial maintenance methods.

Fisheries Research, 161(??):280–284, January 2015. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783614002586.

Butler:2017:CSF

[BBV17] Mark J. Butler, Donald C. Behringer, and Marla M. Valentine. Commercial sponge fishery impacts on the population dynamics of sponges in the Florida keys, FL (USA). Fisheries Research, 190(??):113–121, June 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617300449.

Boenish:2018:SDE

[BC18] Robert Boenish and Yong Chen. Spatiotemporal dynamics of effective fishing effort in the American lobster (Homarus americanus) fishery along the coast of Maine, USA. Fisheries Research, 199(??):231–241, March 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617303090.

Benson:2015:ECR

[BCC15] Ashleen J. Benson, Sean P. Cox, and Jaclyn S. Cleary. Evaluating the conservation risks of aggregate harvest management in a spatially-structured herring fishery. Fisheries Research, 167(??):101–113, July 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615000405.

Babcock:2013:LBI

[BCKG13] Elizabeth A. Babcock, Robin Coleman, Mandy Karnauskas, and Janet Gibson. Length-based indicators of fishery and ecosystem status: Glover’s Reef Marine Reserve, Belize. Fisheries Research, 147(??):434–445, October 2013. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783613000799.

Behrens-Chapuis:2018:DEC

[BCMS^{+}18] Simone Behrens-Chapuis, Tadeusz Malewski, Ewa Suchecka, Matthias F. Geiger, Fabian Herder, and Wiesław Bogdanowicz. Discriminating European cyprinid specimens by barcode high-resolution melting analysis (Bar-HRM) — a cost efficient and faster way for specimen assignment? Fisheries Research, 204(??):61–73, August 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618300353.

Benzinou:2013:DSS

[BCN^{+}13] Abdesslam Benzinou, Sébastien Carbini, Kamal Nasreddine, Romain Elleboode, and Kélig Mahé. Discriminating stocks of striped red mullet (Mullus surmuletus) in the Northwest European seas using three automatic shape classification methods. Fisheries Research, 143(??):153–160, June 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783613000192.

Berge:2012:PDP

[BCP^{+}12] Julien Bergé, Hervé Capra, Hervé Pella, Tracey Steig, Michaël Ovidio, Elise Bultel, and Nicolas Lamouroux. Probability of detection and positioning error of a hydro acoustic telemetry system in a fast-flowing river: Intrinsic and environmental determinants. Fisheries Research, 125–126(??):1– 13, August 2012. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783612000835.

Butler:2017:AHS

[BCPP17] Edward C. Butler, Amber-Robyn Childs, Matthew C. Parkinson, and Warren M. Potts. An assessment of the health and survival of fishes caught-and-released in highenergy surf zones during a South African competitive angling event. Fisheries Research, 195(??):152–168, November 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783617301765.

Bartolino:2011:FIG

[BCTS11] Valerio Bartolino, Francesco Colloca, Lorna Taylor, and Gunnar Stefansson. First implementation of a gadget model for the analysis of hake in the Mediterranean. Fisheries Research, 107(1–3):75–83, January 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610002584.

Brodeur:2011:CPS

[BDB^{+}11] Richard D. Brodeur, Elizabeth A. Daly, Cassandra E. Benkwitt, Cheryl A. Morgan, and Robert L. Emmett. Catching the prey: Sampling juvenile fish and invertebrate prey fields of juvenile Coho and Chinook salmon during their early marine residence. Fisheries Research, 108(1):65–73, February 2011. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783610003085.

Bower:2016:EEC

[BDB^{+}16] Shannon D. Bower, Andy J. Danylchuk, Jacob W. Brownscombe, Jason D. Thiem, and Steven J. Cooke. Evaluating effects of catch-and-release angling on peacock bass (Cichla ocellaris) in a Puerto Rican reservoir: a rapid assessment approach. Fisheries Research, 175(??):95–102, March 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783615301375.

Beveren:2017:HCU

[BDC^{+}17] Elisabeth Van Beveren, Daniel Duplisea, Martin Castonguay, Thomas Doniol-Valcroze, Stéphane Plourde, and Noel Cadigan. How catch underreporting can bias stock assessment of and advice for northwest Atlantic mackerel and a possible resolution using censored catch. Fisheries Research, 194(??):146–154, October 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361730142X.

Beatty:2016:PRB

[BdGD^{+}16] Stephen Beatty, Martin de Graaf, Rodney Duffy, Vinh Nguyen, and Brett Molony. Plasticity in the reproductive biology of the smooth marron Cherax cainii (Decapoda: Parastacidae): a large freshwater crayfish of south-western Australia. Fisheries Research, 177(??):128–136, May 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578361630011X.

Beatty:2011:PPB

[BdGM^{+}11] Stephen Beatty, Martin de Graaf, Brett Molony, Vinh Nguyen, and Kenneth Pollock. Plasticity in population biology of Cherax cainii (Decapoda: Parastacidae) inhabiting lentic and lotic environments in south-western Australia: Implications for the sustainable management of the recreational fishery. Fisheries Research, 110(2):312–324, July 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783611001706.

Boman:2018:VSM

[BdGN^{+}18] Erik Maitz Boman, Martin de Graaf, Leopold A. J. Nagelkerke, Allan W. Stoner, Caroline E. Bissada, Omar Hernando Avila-Poveda, Erick Raul Baqueiro-Cardenas, and Aad C. Smaal. Variability in size at maturity and reproductive season of queen conch Lobatus gigas (Gastropoda: Strombidae) in the wider Caribbean region. Fisheries Research, 201(??): 18–25, May 2018. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783617303582.

Beardmore:2011:EFE

[BDHA11] Ben Beardmore, Malte Dorow, Wolfgang Haider, and Robert Arlinghaus. The elasticity of fishing effort response and harvest outcomes to altered regulatory policies in eel (Anguilla anguilla) recreational angling. Fisheries Research, 110(1):136–148, June 2011. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611001329.

Bertocci:2014:MEH

[BDM^{+}14] Iacopo Bertocci, R. Dominguez, I. Machado, C. Freitas, J. Domínguez Godino, I. Sousa-Pinto, M. Gonçalves, and M. B. Gaspar. Multiple effects of harvesting on populations of the purple sea urchin Paracentrotus lividus in north Portugal. Fisheries Research, 150(??):60–65, February 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783613002373.

Benitez:2018:MBP

[BDM^{+}18] Jean-Philippe Benitez, Arnaud Dierckx, Billy Nzau Matondo, Xavier Rollin, and Michaël Ovidio. Movement behaviours of potamodromous fish within a large anthropised river after the reestablishment of the longitudinal connectivity. Fisheries Research, 207(??):140–149, November 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783618301784.

Bunholi:2018:FIT

[BdSFB^{+}18] Ingrid Vasconcellos Bunholi, Bruno Lopes da Silva Ferrette, Juliana Beltramin De Biasi, Carolina de Oliveira Magalhães, Matheus Marcos Rotundo, Claudio Oliveira, Fausto Foresti, and Fernando Fernandes Mendonça. The fishing and illegal trade of the angelshark: DNA barcoding against misleading identifications. Fisheries Research, 206(??):193–197, October 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783618301541.

Bellchambers:2013:AET

[BEM13] L. M. Bellchambers, S. N. Evans, and J. J. Meeuwig. Assessing the effectiveness of two methods of habitat characterisation for understanding species habitat relationships, using the western rock lobster (Panulirus cygnus George). Fisheries Research, 139(??):5–10, March 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612003384.

Barrios:2017:UME

[BEM^{+}17] Alexander Barrios, Bruno Ernande, Kelig Mahé, Verena Trenkel, and Marie-Joëlle Rochet. Utility of mixed effects models to inform the stock structure of whiting in the Northeast Atlantic Ocean. Fisheries Research, 190(??):132–139, June 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783617300334.

Bennema:2018:LTO

[Ben18] Floris P. Bennema. Long-term occurrence of Atlantic bluefin tuna Thunnus thynnus in the North Sea: contributions of non-fishery data to population studies. Fisheries Research, 199(??):177–185, March 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617303272.

Bertelsen:2013:CDM

[Ber13] Rodney D. Bertelsen. Characterizing daily movements, nomadic movements, and reproductive migrations of Panulirus argus around the Western Sambo Ecological Reserve (Florida, USA) using acoustic telemetry. Fisheries Research, 144(??):91–102, July 2013. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612003529.

Berger:2019:CTV

[Ber19] Aaron M. Berger. Character of temporal variability in stock productivity influences the utility of dynamic reference points. Fisheries Research, 217(??):185–197, September 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783618303424.

Bellchambers:2016:IMP

[BFHT16] Lynda M. Bellchambers, Emily A. Fisher, Alastair V. Harry, and Kendra L. Travaille. Identifying and mitigating potential risks for Marine Stewardship Council assessment and certification. Fisheries Research, 182(??):7–17, October 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783616300625.

Braham:2014:NIS

[BFL^{+}14] Cheikh-Baye Braham, Pierre Fréon, Alain Laurec, Hervé Demarcq, and Nicolas Bez. New insights in the spatial dynamics of sardinella stocks off Mauritania (North-West Africa) based on logbook data analysis. Fisheries Research, 154(??):195–204, June 2014. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614000605.

Baker:2014:JGM

[BFM14] Ronald Baker, Masami Fujiwara, and Thomas J. Minello. Juvenile growth and mortality effects on white shrimp Litopenaeus setiferus population dynamics in the northern Gulf of Mexico. Fisheries Research, 155(??):74–82, July 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783614000721.

Bris:2013:SET

[BFW13] Arnault Le Bris, Alain Fréchet, and Joseph S. Wroblewski. Supplementing electronic tagging with conventional tagging to redesign fishery closed areas. Fisheries Research, 148(??):106–116, November 2013. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613002063.

Bradman:2011:DCM

[BGA11] H. Bradman, P. Grewe, and B. Appleton. Direct comparison of mitochondrial markers for the analysis of swordfish population structure. Fisheries Research, 109(1):95–99, April 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783611000415.

Bosley:2019:OCH

[BGB^{+}19] Katelyn M. Bosley, Daniel R. Goethel, Aaron M. Berger, Jonathan J. Deroba, Kari H. Fenske, Dana H. Hanselman, Brian J. Langseth, and Amy M. Schueller. Overcoming challenges of harvest quota allocation in spatially structured populations. Fisheries Research, 220(??):Article 105344, December 2019. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783619301997.

Baez:2019:FSA

[BGBCM19] José Carlos Báez, Salvador García-Barcelona, Juan Antonio Camiñas, and David Macías. Fishery strategy affects the loggerhead sea turtle mortality trend due to the longline bycatch. Fisheries Research, 212(??):21–28, April 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783618303369.

Bouhoubeiny:2011:TRP

[BGD11] Elkhadim Bouhoubeiny, Grégory Germain, and Philippe Druault. Time-resolved PIV investigations of the flow field around cod-end net structures. Fisheries Research, 108(2–3):344–355, March 2011. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783611000142.

Basilone:2015:AGM

[BGF^{+}15] G. Basilone, K. Ganias, R. Ferreri, M. D’Elia, E. M. Quinci, S. Mazzola, and A. Bonanno. Application of GAMs and multinomial models to assess the spawning pattern of fishes with daily spawning synchronicity: a case study in the European anchovy (Engraulis encrasicolus) in the central Mediterranean Sea. Fisheries Research, 167(??):92–100, July 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783615000399.

Barlaup:2013:TDC

[BGL^{+}13] Bjørn Torgeir Barlaup, Sven-Erik Gabrielsen, Jon Løyland, Marie-Lise Schläppy, Tore Wiers, Knut Wiik Vollset, and Ulrich Pulg. Trap design for catching fish unharmed and the implications for estimates of sea lice (Lepeophtheirus salmonis) on anadromous brown trout (Salmo trutta). Fisheries Research, 139(??):43–46, March 2013. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612000550.

Beckmann:2015:XRI

[BGL15] Melanie C. Beckmann, James F. Gilliam, and R. Brian Langerhans. X-ray imaging as a time-saving, non-invasive technique for diet analysis. Fisheries Research, 161(??):1–7, January 2015. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783614001878.

Breen:2016:SFB

[BGP^{+}16] Mike Breen, Norman Graham, Michael Pol, Pingguo He, Dave Reid, and Petri Suuronen. Selective fishing and balanced harvesting. Fisheries Research, 184(??):2–8, December 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783616300704.

Bellchambers:2016:AMS

[BGW^{+}16] L. M. Bellchambers, D. J. Gaughan, B. S. Wise, G. Jackson, and W. J. Fletcher. Adopting Marine Stewardship Council certification of Western Australian fisheries at a jurisdictional level: the benefits and challenges. Fisheries Research, 183(??):609–616, November 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616302260.

Boulcott:2011:ISD

[BH11] Philip Boulcott and Trevor R. W. Howell. The impact of scallop dredging on rocky-reef substrata. Fisheries Research, 110(3):415–420, August 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611001950.

Bellman:2013:DFM

[BH13] Marlene A. Bellman and Eliza Heery. Discarding and fishing mortality trends in the U.S. west coast groundfish demersal trawl fishery. Fisheries Research, 147(??):115–126, October 2013. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783613001057.

Bacheler:2010:UGL

[BHB^{+}10] Nathan M. Bacheler, Joseph E. Hightower, Summer M. Burdick, Lee M. Paramore, Jeffrey A. Buckel, and Kenneth H. Pollock. Using generalized linear models to estimate selectivity from short-term recoveries of tagged red drum Sciaenops ocellatus: Effects of gear, fate, and regulation period. Fisheries Research, 102(3):266–275, March 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783609003233.

Benoit:2010:AFI

[BHC10] Hugues P. Benoît, Thomas Hurlbut, and Joël Chassé. Assessing the factors influencing discard mortality of demersal fishes using a semi-quantitative indicator of survival potential. Fisheries Research, 106(3):436–447, December 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783610002274.

Bradbury:2016:GMS

[BHC^{+}16] Ian R. Bradbury, Lorraine C. Hamilton, Gerald Chaput, Martha J. Robertson, Herlé Goraguer, Arthur Walsh, Vicki Morris, David Reddin, J. Brian Dempson, Timothy F. Sheehan, Timothy King, and Louis Bernatchez. Genetic mixed stock analysis of an interceptory Atlantic salmon fishery in the Northwest Atlantic. Fisheries Research, 174(??):234–244, February 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615301065.

Benoit:2012:EFS

[BHCJ12] Hugues P. Benoît, Thomas Hurlbut, Joël Chassé, and Ian D. Jonsen. Estimating fishery-scale rates of discard mortality using conditional reasoning. Fisheries Research, 125–126(??):318–330, August 2012. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611003705.

Brcic:2015:SCS

[BHCS15] Jure Brčić, Bent Herrmann, Francesco De Carlo, and Antonello Sala. Selective characteristics of a shark-excluding grid device in a Mediterranean trawl. Fisheries Research, 172(??):352–360, December 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615300448.

Batista:2015:ACL

[BHG^{+}15] M. I. Batista, B. Horta e Costa, L. Gonçalves, M. Henriques, K. Erzini, J. E. Caselle, E. J. Gonçalves, and H. N. Cabral. Assessment of catches, landings and fishing effort as useful tools for MPA management. Fisheries Research, 172(??):197–208, December 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615300291.

Berg:2017:STA

[BHG^{+}17] Florian Berg, Åse Husebø, Jane Aanestad Godiksen, Aril Slotte, and Arild Folkvord. Spawning time of Atlantic herring (Clupea harengus) populations within a restricted area reflects their otolith growth at the larval stage. Fisheries Research, 194(??):68–75, October 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617301364.

Barbeaux:2014:NAE

[BHI14] Steven J. Barbeaux, John K. Horne, and James N. Ianelli. A novel approach for estimating location and scale specific fishing exploitation rates of eastern Bering Sea walleye pollock (Theragra chalcogramma). Fisheries Research, 153(??): 69–82, May 2014. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578361400006X.

Bagdonas:2012:CWS

[BHL12] Kasparas Bagdonas, Odd-Børre Humborstad, and Svein Løkkeborg. Capture of wild saithe (Pollachius virens) and cod (Gadus morhua) in the vicinity of salmon farms: Three pot types compared. Fisheries Research, 134–136(??):1–5, December 2012. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783612002044.

Bayse:2016:CTM

[BHL^{+}16] Shannon M. Bayse, Bent Herrmann, Heleen Lenoir, Jochen Depestele, Hans Polet, Els Vanderperren, and Bart Verschueren. Could a T90 mesh codend improve selectivity in the Belgian beam trawl fishery? Fisheries Research, 174(??):201–209, February 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615301089.

Brcic:2018:SSN

[BHM^{+}18a] Jure Brčić, Bent Herrmann, Marina Mašanović, Mateja Baranović, Svjetlana Krstulović Šifner, and Frane Škeljo. Size selection of Nephrops norvegicus (L.) in commercial creel fishery in the Mediterranean Sea. Fisheries Research, 200(??):25–32, April 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783617303417.

Brcic:2018:CMD

[BHM^{+}18b] Jure Brčić, Bent Herrmann, Marina Mašanović, Svjetlana Krstulović Šifner, and Frane Škeljo. CREELSE-LECT — a method for determining the optimal creel mesh: Case study on Norway lobster (Nephrops norvegicus) fishery in the Mediterranean Sea. Fisheries Research, 204(??):433–440, August 2018. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618300948.

Brownscombe:2019:FRF

[BHP^{+}19] Jacob W. Brownscombe, Kieran Hyder, Warren Potts, Kyle L. Wilson, Kevin L. Pope, Andy J. Danylchuk, Steven J. Cooke, Adrian Clarke, Robert Arlinghaus, and John R. Post. The future of recreational fisheries: Advances in science, monitoring, management, and practice. Fisheries Research, 211(??):247–255, March 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618302881.

Bayse:2014:QAB

[BHPC14] Shannon M. Bayse, Pingguo He, Michael V. Pol, and David M. Chosid. Quantitative analysis of the behavior of longfin inshore squid (Doryteuthis pealeii) in reaction to a species separation grid of an otter trawl. Fisheries Research, 152(??):55–61, April 2014. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613000866.

Brcic:2016:CSM

[BHS16] Jure Brčić, Bent Herrmann, and Antonello Sala. Can a square-mesh panel inserted in front of the codend improve the exploitation pattern in Mediterranean bottom trawl fisheries? Fisheries Research, 183(??):13–18, November 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783616301448.

Bass:2019:FCI

[BHT^{+}19] Arthur L. Bass, Scott G. Hinch, Amy K. Teffer, David A. Patterson, and Kristina M. Miller. Fisheries capture and infectious agents are associated with travel rate and survival of Chinook salmon during spawning migration. Fisheries Research, 209(??):156–166, January 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618302443.

Bernal:2012:AED

[BJD12] M. Bernal, M. P. Jiménez, and J. Duarte. Anchovy egg development in the Gulf of Cádiz and its comparison with development rates in the Bay of Biscay. Fisheries Research, 117–118(??):112–120, April 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611001597.

Boukal:2012:SCA

[BJKH12] David S. Boukal, Martin Jankovský, Jan Kubečka, and Mikko Heino. Stock–catch analysis of carp recreational fisheries in Czech reservoirs: Insights into fish survival, water body productivity and impact of extreme events. Fisheries Research, 119–120(??):23–32, May 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611003687.

Borjesson:2013:SKC

[BJP^{+}13] Patrik Börjesson, Patrik Jonsson, Selma Pacariz, Göran Björk, Martin I. Taylor, and Henrik Svedäng. Spawning of Kattegat cod (Gadus morhua) — mapping spatial distribution by egg surveys. Fisheries Research, 147(??):63–71, October 2013. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578361300129X.

Berger:2012:ASP

[BJZB12] Aaron M. Berger, Michael L. Jones, Yingming Zhao, and James R. Bence. Accounting for spatial population structure at scales relevant to life history improves stock assessment: the case for Lake Erie walleye Sander vitreus. Fisheries Research, 115–116(??):44–59, March 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611003353.

Brandt:2010:SFA

[BK10] Urs Steiner Brandt and Lone Grønbæk Kronbak. On the stability of fishery agreements under exogenous change: an example of agreements under climate change. Fisheries Research, 101(1–2):11–19, January 5, 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783609002331.

Berg:2012:SAL

[BK12] Casper W. Berg and Kasper Kristensen. Spatial age–length key modelling using continuation ratio logits. Fisheries Research, 129–130(??):119–126, October 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612002007.

Bradburn:2015:ILC

[BK15] Mark J. Bradburn and Aimee A. Keller. Impact of light on catch rate of four demersal fish species during the 2009–2010 U.S. west coast groundfish bottom trawl survey. Fisheries Research, 164(??):193–200, April 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361400335X.

Blom:2016:SVF

[BK16] M. Blom and J. Kennedy. Spatial variation in fecundity of Norwegian coastal cod, Gadus morhua (Linnaeus), along the coast of Norway. Fisheries Research, 183(??):401–403, November 2016. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783616301709.

Bradford:2019:FTR

[BK19] Kirsten Bradford and Robert Eliakim Katikiro. Fighting the tides: a review of gender and fisheries in Tanzania. Fisheries Research, 216(??):79–88, August 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619300943.

Bryhn:2014:GLV

[BKLB14] Andreas C. Bryhn, Sara J. Königson, Sven-Gunnar Lunneryd, and Mikaela A. J. Bergenius. Green lamps as visual stimuli affect the catch efficiency of floating cod (Gadus morhua) pots in the Baltic Sea. Fisheries Research, 157(??):187–192, September 2014. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614001477.

Bergeron:2013:IFS

[BKM13] Jean-Pierre Bergeron, Noussithé Koueta, and Jacques Massé. Interannual fluctuations in spring pelagic ecosystem productivity in the Bay of Biscay (northeast Atlantic) measured by mesozooplankton aspartate transcarbamylase activity and relationships with anchovy population dynamics. Fisheries Research, 143(??):184–190, June 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613000490.

Blackwell:2019:AEU

[BKM19] Brian G. Blackwell, Todd M. Kaufman, and Tyrel S. Moos. Angler exploitation of an unexploited walleye population in the northern Great Plains. Fisheries Research, 216(??):59– 64, August 2019. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783619300918.

Belhabib:2014:FCM

[BKS^{+}14] Dyhia Belhabib, Viviane Koutob, Aliou Sall, Vicky W. Y. Lam, and Daniel Pauly. Fisheries catch misreporting and its implications: the case of Senegal. Fisheries Research, 151(??):1–11, March 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783613003007.

Belhabib:2015:CPM

[BKS^{+}15] Dyhia Belhabib, Viviane Koutob, Aliou Sall, Vicky W. Y. Lam, Dirk Zeller, and Daniel Pauly. Counting pirogues and missing the boat: Reply to Chaboud et al.’s comment on Belhabib et al. “Fisheries catch misreporting and its implications: the case of Senegal”. Fisheries Research, 164(??):325– 328, April 2015. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783614002847.

Beveren:2016:ELH

[BKS^{+}16] Elisabeth Van Beveren, Maria Klein, Ester A. Serrão, Emanuel J. Gonçalves, and Rita Borges. Early life history of larvae and early juvenile Atlantic horse mackerel Trachurus trachurus off the Portuguese west coast. Fisheries Research, 183(??):111–118, November 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616301552.

Broadhurst:2018:NFM

[BKS18] Matt K. Broadhurst, Marco Kienzle, and John Stewart. Natural and fishing mortalities affecting eastern sea garfish, Hyporhamphus australis inferred from age-frequency data using hazard functions. Fisheries Research, 198(??):43–49, February 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783617302916.

Battista:2017:CAR

[BKSF17] Willow Battista, Kendra Karr, Nicole Sarto, and Rod Fujita. Comprehensive Assessment of Risk to Ecosystems (CARE): a cumulative ecosystem risk assessment tool. Fisheries Research, 185(??):115–129, January 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616303101.

Bjornsson:2010:EAF

[BKT10] Björn Björnsson, Hjalti Karlsson, and Vilhjálmur Thorsteinsson. Effects of anthropogenic feeding on the migratory behaviour of coastal cod (Gadus morhua) in Northwest Iceland. Fisheries Research, 106(1):81–92, October 2010. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783610001761.

Brinkhof:2017:ICE

[BLHG17] Jesse Brinkhof, Roger B. Larsen, Bent Herrmann, and Eduardo Grimaldo. Improving catch efficiency by changing ground gear design: Case study of Northeast Atlantic cod (Gadus morhua) in the Barents Sea bottom trawl fishery. Fisheries Research, 186 (part 1)(??):269–282, February 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783616303368.

Brinkhof:2018:AIB

[BLHO18] Jesse Brinkhof, Roger B. Larsen, Bent Herrmann, and Stein H. Olsen. Assessing the impact of buffer towing on the quality of Northeast Atlantic cod (Gadus morhua) caught with a bottom trawl. Fisheries Research, 206(??):209–219, October 2018. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783618301577.

Barreto:2011:AGB

[BLHS11] Rodrigo R. Barreto, Rosangela P. Lessa, Fabio H. Hazin, and Francisco M. Santana. Age and growth of the blacknose shark, Carcharhinus acronotus (Poey, 1860) off the northeastern Brazilian Coast. Fisheries Research, 110(1):170–176, June 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783611001354.

Bond:2018:DSH

[BLP^{+}18] T. Bond, T. J. Langlois, J. C. Partridge, M. J. Birt, B. E. Malseed, L. Smith, and D. L. McLean. Diel shifts and habitat associations of fish assemblages on a subsea pipeline. Fisheries Research, 206(??):220–234, October 2018. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783618301474.

Bustos:2015:CEL

[BLPF^{+}15] Claudia A. Bustos, Mauricio F. Landaeta, Pámela Palacios-Fuentes, Nicole Jahnsen-Guzmán, and Fernando Balbontín. Comparing early life traits of hakes from Chilean Patagonian fjords inferred by otolith microstructure analysis. Fisheries Research, 164(??):35–44, April 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614003099.

Blabolil:2016:AEP

[BLR^{+}16] Petr Blabolil, Maxime Logez, Daniel Ricard, Marie Prchalová, Milan Říha, Alban Sagouis, Jiří Peterka, Jan Kubečka, and Christine Argillier. An assessment of the ecological potential of Central and Western European reservoirs based on fish communities. Fisheries Research, 173(??):80–87, January 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361500171X.

Bakke:2018:SMM

[BLW^{+}18] Snorre Bakke, Wenche E. Larssen, Astrid K. Woll, Guldborg Søvik, Agnes C. Gundersen, Carsten Hvingel, and Einar M. Nilssen. Size at maturity and molting probability across latitude in female Cancer pagurus. Fisheries Research, 205(??):43–51, September 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618300985.

Broadhurst:2018:CEG

[BM18a] Matt K. Broadhurst and Russell B. Millar. Configuring escape gaps in recreational rectangular traps to improve size selection for eastern rock lobster, Sagmariasus verreauxi. Fisheries Research, 207(??):182–186, November 2018. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783618301644.

Broadhurst:2018:RGF

[BM18b] Matt K. Broadhurst and Russell B. Millar. Relative ghost fishing of portunid traps with and without escape gaps. Fisheries Research, 208(??):202–209, December 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618302169.

Boucek:2019:DVC

[BM19a] R. E. Boucek and D. Morley. Demonstrating the value of cross-ecosystem syntheses and comparisons in animal movement and acoustic telemetry research. Fisheries Research, 216(??):74–78, August 2019. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619300840.

Broadhurst:2019:MVM

[BM19b] Matt K. Broadhurst and Russell B. Millar. Mono- vs multifilament baited conical hoop (tangle) nets: Effects on portunid catches and marine debris. Fisheries Research, 220(??):Article 105356, December 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619302115.

Bladon:2018:DFR

[BMAMG18] Annabelle Jade Bladon, Essam Yassin Mohammed, Liaquat Ali, and E. J. Milner-Gulland. Developing a frame of reference for fisheries management and conservation interventions. Fisheries Research, 208(??):296–308, December 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783618302315.

Broadhurst:2010:DVS

[BMB10] Matt K. Broadhurst, Russell B. Millar, and Craig P. Brand. Diamond- vs. square-mesh codend selectivity in southeastern Australian estuarine squid trawls. Fisheries Research, 102(3):276–285, March 2010. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783609003257.

Balboa:2018:UGR

[BMB^{+}18] Cecilia Balboa, Carlos Molinet, Nancy Barahona, Patricio A. Díaz, Dagoberto Subiabre, and Paulina Gebauer. Using growth ring formation to study growth patterns of exploited sea urchin (Loxechinus albus) populations. Fisheries Research, 201(??):88–97, May 2018. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618300183.

Burch:2010:FBC

[BMC10] Paul Burch, Stephen Mayfield, and Rowan C. Chick. A fisher-based, catch-sampling program for a large-scale, benthic molluscan fishery. Fisheries Research, 106(3):535–542, December 2010. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783610002535.

Bacheler:2019:FSM

[BMCS19] Nathan M. Bacheler, Théo Michelot, Robin T. Cheshire, and Kyle W. Shertzer. Fine-scale movement patterns and behavioral states of gray triggerfish Balistes capriscus determined from acoustic telemetry and hidden Markov models. Fisheries Research, 215(??):76–89, July 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619300578.

Bitalo:2015:DGF

[BMdS^{+}15] Daphne N. Bitalo, Simo N. Maduna, Charlene da Silva, Rouvay Roodt-Wilding, and Aletta E. Bester van der Merwe. Differential gene flow patterns for two commercially exploited shark species, tope (Galeorhinus galeus) and common smoothhound (Mustelus mustelus) along the south–west coast of South Africa. Fisheries Research, 172(??):190–196, December 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783615300126.

Baez:2019:NAR

[BMEGV^{+}19] José Carlos Báez, Pedro Muñoz-Exposito, M. J. Gómez-Vives, D. Godoy-Garrido, and David Macías. The NAO affects the reproductive potential of small tuna migrating from the Mediterranean Sea. Fisheries Research, 216(??):41–46, August 2019. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783619300888.

Bevacqua:2010:BGM

[BMF^{+}10] Daniele Bevacqua, Paco Melià, Maria C. Follesa, Giulio A. De Leo, Marino Gatto, and Angelo Cau. Body growth and mortality of the spiny lobster Palinurus elephas within and outside a small marine protected area. Fisheries Research, 106(3):543–549, December 2010. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610002560. See erratum [KPL^{+}07].

Bevacqua:2011:EBG

[BMF^{+}11] Daniele Bevacqua, Paco Melià, Maria C. Follesa, Giulio A. De Leo, Marino Gatto, and Angelo Cau. Erratum to “Body growth and mortality of the spiny lobster Palinurus elephas within and outside a small marine protected area” [Fish. Res. 106 (2010) 543–549]. Fisheries Research, 108(2–3):404, March 2011. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611000099. See [BMF^{+}10].

Bellanger:2016:NAD

[BMG16] Manuel Bellanger, Claire Macher, and Olivier Guyader. A new approach to determine the distributional effects of quota management in fisheries. Fisheries Research, 181(??):116–126, September 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783616300972.

Broadhurst:2017:PID

[BMH17] Matt K. Broadhurst, Russell B. Millar, and Brian Hughes. Performance of industry-developed escape gaps in Australian Portunus pelagicus traps. Fisheries Research, 187(??):120–126, March 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783616303988.

Broadhurst:2018:UME

[BMH18] Matt K. Broadhurst, Russell B. Millar, and Brian Hughes. Utility of multiple escape gaps in Australian Scylla serrata traps. Fisheries Research, 204(??):88–94, August 2018. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783618300249.

Barnes-Mauthe:2013:TEV

[BMOZ13] Michele Barnes-Mauthe, Kirsten L. L. Oleson, and Bienvenue Zafindrasilivonona. The total economic value of smallscale fisheries with a characterization of post-landing trends: an application in Madagascar with global relevance. Fisheries Research, 147(??):175–185, October 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613001537.

Borsetti:2018:SVL

[BMR^{+}18] S. Borsetti, D. Munroe, D. B. Rudders, C. Dobson, and E. A. Bochenek. Spatial variation in life history characteristics of waved whelk (Buccinum undatum L.) on the U.S. mid-Atlantic continental shelf. Fisheries Research, 198(??):129–137, February 2018. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617302746.

Bower:2019:SLR

[BMR^{+}19] Shannon D. Bower, Neethi Mahesh, Rajeev Raghavan, Andy J. Danylchuk, and Steven J. Cooke. Sub-lethal responses of mahseer (Tor khudree) to catch-and-release recreational angling. Fisheries Research, 211(??):231–237, March 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783618303047.

Barbieri:2014:MCS

[BMRC14] M. Barbieri, F. Maltagliati, M. I. Roldán, and A. Castelli. Molecular contribution to stock identification in the smallspotted catshark, Scyliorhinus canicula (Chondrichthyes, Scyliorhinidae). Fisheries Research, 154(??):11–16, June 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578361400037X.

Brodziak:2015:SRR

[BMS15] Jon Brodziak, Marc Mangel, and Chi-Lu Sun. Stockrecruitment resilience of North Pacific striped marlin based on reproductive ecology. Fisheries Research, 166(??):140– 150, June 2015. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783614002537.

Bakhrankova:2014:SOO

[BMU14] Krystsina Bakhrankova, Kjetil T. Midthun, and Kristin T. Uggen. Stochastic optimization of operational production planning for fisheries. Fisheries Research, 157(??):147–153, September 2014. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783614001118.

Brewer:2012:LTR

[BN12] Reid S. Brewer and Brenda L. Norcross. Long-term retention of internal elastomer tags in a wild population of North Pacific giant octopus (Enteroctopus dofleini). Fisheries Research, 134–136(??):17–20, December 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612002330.

Bolland:2019:RAT

[BNA^{+}19] Jonathan D. Bolland, Andrew D. Nunn, Natalie V. Angelopoulos, Jamie R. Dodd, Peter Davies, Catherine Gutmann Roberts, J. Robert Britton, and Ian G. Cowx. Refinement of acoustic-tagging protocol for twaite shad Alosa fallax (Lacépède), a species sensitive to handling and sedation. Fisheries Research, 212(??):183–187, April 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618303503.

Bastardie:2010:EFE

[BNAE10] Francois Bastardie, J. Rasmus Nielsen, Bo Sølgaard Andersen, and Ole Ritzau Eigaard. Effects of fishing effort allocation scenarios on energy efficiency and profitability: an individual-based model applied to Danish fisheries. Fisheries Research, 106(3):501–516, December 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610002456.

Bastardie:2013:IIT

[BNAE13] Francois Bastardie, J. Rasmus Nielsen, Bo Sølgaard Andersen, and Ole Ritzau Eigaard. Integrating individual trip planning in energy efficiency — building decision tree models for Danish fisheries. Fisheries Research, 143(??):119–130, June 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783613000222.

Berg:2014:EAA

[BNK14] Casper W. Berg, Anders Nielsen, and Kasper Kristensen. Evaluation of alternative age-based methods for estimating relative abundance from survey data in relation to assessment models. Fisheries Research, 151(??):91–99, March 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783613002324.

Brunnschweiler:2011:SOS

[BNM11] Juerg M. Brunnschweiler, Frank Nielsen, and Philip Motta. In situ observation of stomach eversion in a linecaught shortfin mako (Isurus oxyrinchus). Fisheries Research, 109(1):212–216, April 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783611000737.

Bailey:2019:LVA

[BNSS19] Christopher T. Bailey, Austin M. Noring, Stephanie L. Shaw, and Greg G. Sass. Live versus artificial bait influences on walleye (Sander vitreus) angler effort and catch rates on Escanaba Lake, Wisconsin, 1993–2015. Fisheries Research, 219(??):Article 105330, November 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619301857.

Bastardie:2010:DMF

[BNU^{+}10] Francois Bastardie, J. Rasmus Nielsen, Clara Ulrich, Josefine Egekvist, and Henrik Degel. Detailed mapping of fishing effort and landings by coupling fishing logbooks with satellite-recorded vessel geo-location. Fisheries Research, 106(1):41–53, October 2010. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610001554.

Bayless:2017:AMH

[BOBP^{+}17] Alexandra R. Bayless, Erin M. Oleson, Simone Baumann-Pickering, Anne E. Simonis, Jamie Marchetti, Sean Martin, and Sean M. Wiggins. Acoustically monitoring the Hawai‘ i longline fishery for interactions with false killer whales. Fisheries Research, 190(??):122–131, June 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617300346.

Banon:2018:TSS

[BOCÁ^{+}18] Rafael Bañón, Jaime Otero, José Manuel Campelos-Álvarez, Alberto Garazo, and Alexandre Alonso-Fernández. The traditional small-scale octopus trap fishery off the Galician coast (Northeastern Atlantic): Historical notes and current fishery dynamics. Fisheries Research, 206(??):115–128, October 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783618301413.

Bowden:2017:HIS

[Bow17] Timothy J. Bowden. The humoral immune systems of the American lobster (Homarus americanus) and the European lobster (Homarus gammarus). Fisheries Research, 186 (part 1)(??):367–371, February 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616302351.

Brown:2012:PBJ

[BPC^{+}12] Richard S. Brown, Brett D. Pflugrath, Alison H. Colotelo, Colin J. Brauner, Thomas J. Carlson, Z. Daniel Deng, and Adam G. Seaburg. Pathways of barotrauma in juvenile salmonids exposed to simulated hydroturbine passage: Boyle’s law vs. Henry’s law. Fisheries Research, 121–122(??):43–50, June 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783612000379.

Bris:2017:MSQ

[BPG^{+}17] Arnault Le Bris, Andrew J. Pershing, Julien Gaudette, Tracy L. Pugh, and Kathleen M. Reardon. Multi-scale quantification of the effects of temperature on size at maturity in the American lobster (Homarus americanus). Fisheries Research, 186 (part 1)(??):397–406, February 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616303010.

Bidegain:2017:MTP

[BPK^{+}17] G. Bidegain, E. N. Powell, J. M. Klinck, E. E. Hofmann, T. Ben-Horin, D. Bushek, S. E. Ford, D. M. Munroe, and X. Guo. Modeling the transmission of Perkinsus marinus in the Eastern oyster Crassostrea virginica. Fisheries Research, 186 (part 1)(??):82–93, February 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616302582.

Boavida-Portugal:2010:EAR

[BPMGP10] J. Boavida-Portugal, A. Moreno, L. Gordo, and J. Pereira. Environmentally adjusted reproductive strategies in females of the commercially exploited common squid Loligo vulgaris. Fisheries Research, 106(2):193–198, November 2010. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783610001323.

Bellchambers:2016:CRB

[BPPR16] Lynda M. Bellchambers, Bruce F. Phillips, and Mónica Pérez-Ramírez. From certification to recertification the benefits and challenges of the Marine Stewardship Council (MSC): a case study using lobsters. Fisheries Research, 182(??):88–97, October 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783615300758.

Barcenas:2014:AVO

[BPRB^{+}14] Guadalupe Villegas Bárcenas, Catalina Perales-Raya, Aurora Bartolomé, Eduardo Almansa, and Carlos Rosas. Age validation in Octopus maya (Voss and Solís, 1966) by counting increments in the beak rostrum sagittal sections of known age individuals. Fisheries Research, 152(??):93–97, April 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783613001884.

Bao:2018:CAW

[BPS^{+}18] Miguel Bao, Graham J. Pierce, Norval J. C. Strachan, Cristina Martínez, Rosa Fernández, and Ioannis Theodossiou. Consumers’ attitudes and willingness to pay for Anisakis-free fish in Spain. Fisheries Research, 202(??):149– 160, June 2018. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578361730173X.

Blanchet:2019:HVE

[BPS^{+}19] Marie-Anne Blanchet, Raul Primicerio, Aslak Smalås, Juliana Arias-Hansen, and Michaela Aschan. How vulnerable is the European seafood production to climate warming? Fisheries Research, 209(??):251–258, January 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578361830239X.

Breine:2017:SEA

[BPV^{+}17] Jan Breine, Ine S. Pauwels, Pieterjan Verhelst, Lore Vandamme, Raf Baeyens, Jan Reubens, and Johan Coeck. Successful external acoustic tagging of twaite shad Alosa fallax (Lacépède 1803). Fisheries Research, 191(??):36–40, July 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783617300632.

Bennema:2015:FAF

[BR15] Floris P. Bennema and Adriaan D. Rijnsdorp. Fish abundance, fisheries, fish trade and consumption in sixteenthcentury Netherlands as described by adriaen coenen. Fisheries Research, 161(??):384–399, January 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614002677.

Blicharska:2018:RFS

[BR18] Malgorzata Blicharska and Patrik Rönnbäck. Recreational fishing for sea trout — resource for whom and to what value? Fisheries Research, 204(??):380–389, August 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361830078X.

Butterworth:2014:DSM

[BRB^{+}14] Doug S. Butterworth, Rebecca A. Rademeyer, Anabela Brandão, Helena F. Geromont, and Susan J. Johnston. Does selectivity matter? A fisheries management perspective. Fisheries Research, 158(??):194–204, October 2014. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783614000447.

Battaglia:2010:CAF

[BRC^{+}10] Pietro Battaglia, Teresa Romeo, Pierpaolo Consoli, Gianfranco Scotti, and Franco Andaloro. Characterization of the artisanal fishery and its socio-economic aspects in the central Mediterranean Sea (Aeolian Islands, Italy). Fisheries Research, 102(1–2):87–97, February 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578360900277X.

Bendriem:2019:RFS

[BRGS19] Nathan Bendriem, Raphael Roman, Darah Gibson, and U. Rashid Sumaila. A review of the fate of southern British Columbia Coho salmon over time. Fisheries Research, 218(??):10–21, October 2019. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619300931.

Batsleer:2016:MFM

[BRH^{+}16] J. Batsleer, A. D. Rijnsdorp, K. G. Hamon, H. M. J. van Overzee, and J. J. Poos. Mixed fisheries management: Is the ban on discarding likely to promote more selective and fuel efficient fishing in the Dutch flatfish fishery? Fisheries Research, 174(??):118–128, February 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615300825.

Briggs:2010:NEP

[Bri10] R. P. Briggs. A novel escape panel for trawl nets used in the Irish Sea Nephrops fishery. Fisheries Research, 105(2):118– 124, July 2010. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783610000767.

Bayse:2016:ELM

[BRJ^{+}16] Shannon M. Bayse, Chris B. Rillahan, Natalie F. Jones, Vincent Balzano, and Pingguo He. Evaluating a large-mesh belly window to reduce bycatch in silver hake (Merluccius bilinearis) trawls. Fisheries Research, 174(??):1–9, February 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783615300680.

Brock:2013:P

[Bro13a] Robert J. Brock. Preface. Fisheries Research, 144(??):1, July 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578361300074X.

Brooks:2013:EVR

[Bro13b] Elizabeth N. Brooks. Effects of variable reproductive potential on reference points for fisheries management. Fisheries Research, 138(??):152–158, February 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612001750.

Blabolil:2016:PAP

[BRP^{+}16] Petr Blabolil, Daniel Ricard, Jiří Peterka, Milan Říha, Tomáš Jůza, Mojmír Vašek, Marie Prchalová, Martin Čech, Milan Muška, Jaromír Seďa, Tomáš Mrkvička, David S. Boukal, and Jan Kubečka. Predicting asp and pikeperch recruitment in a riverine reservoir. Fisheries Research, 173(??):45–52, January 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615300503.

Blain-Roth:2019:EBR

[BRS19] Brittany J. Blain-Roth and Trent M. Sutton. Effects of barotrauma and recompression events on subsequent embryo condition of yelloweye rockfish. Fisheries Research, 211(??):212–216, March 2019. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618303357.

Busilacchi:2013:QCS

[BRW^{+}13] Sara Busilacchi, Garry R. Russ, Ashley J. Williams, Gavin A. Begg, and Stephen G. Sutton. Quantifying changes in the subsistence reef fishery of indigenous communities in Torres Strait, Australia. Fisheries Research, 137(??):50–58, January 2013. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783612002652.

Buratti:2010:OMP

[BS10] C. C. Buratti and B. A. Santos. Otolith microstructure and pelagic larval duration in two stocks of the Argentine hake, Merluccius hubbsi. Fisheries Research, 106(1):2–7, October 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783610001104.

Balash:2015:EMO

[BSB^{+}15] Cheslav Balash, David Sterling, Jonathan Binns, Giles Thomas, and Neil Bose. The effect of mesh orientation on netting drag and its application to innovative prawn trawl design. Fisheries Research, 164(??):206–213, April 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783614003440.

Balash:2016:PEP

[BSB16] Cheslav Balash, David Sterling, and Matt K. Broadhurst. Progressively evaluating a penaeid W trawl to improve ecoefficiency. Fisheries Research, 181(??):148–154, September 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783616301242.

Broadhurst:2012:EOB

[BSC12] Matt K. Broadhurst, David J. Sterling, and Brian R. Cullis. Effects of otter boards on catches of an Australian penaeid trawl. Fisheries Research, 131–133(??):67–75, November 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783612002214.

Bernard:2013:BGD

[BSD^{+}13] Andrea M. Bernard, Mahmood S. Shivji, Rodrigo Rodrigues Domingues, Fabio Hissa Vieira Hazin, Alberto Ferreira de Amorim, Andres Domingo, Freddy Arocha, Eric D. Prince, John P. Hoolihan, and Alexandre Wagner Silva Hilsdorf. Broad geographic distribution of roundscale spearfish (Tetrapturus georgii) (Teleostei, Istiophoridae) in the Atlantic revealed by DNA analysis: Implications for white marlin and roundscale spearfish management. Fisheries Research, 139(??):93–97, March 2013. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612003050.

Baudron:2019:CCF

[BSF^{+}19] Alan R. Baudron, Natalia Serpetti, Niall G. Fallon, Johanna J. Heymans, and Paul G. Fernandes. Can the common fisheries policy achieve good environmental status in exploited ecosystems: the west of Scotland demersal fisheries example. Fisheries Research, 211(??):217–230, March 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783618302935.

Bubley:2013:UMP

[BSKT13] Walter J. Bubley, James A. Sulikowski, David M. Koester, and Paul C. W. Tsang. Using a multi-parameter approach to reassess maturity of spiny dogfish, Squalus acanthias, following increased fishing pressure in the Western North Atlantic. Fisheries Research, 147(??):202–212, October 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783613001562.

Bailey:2010:AGT

[BSL10] Megan Bailey, U. Rashid Sumaila, and Marko Lindroos. Application of game theory to fisheries over three decades. Fisheries Research, 102(1–2):1–8, February 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783609003051.

Broadhurst:2012:SVL

[BSM12] Matt K. Broadhurst, David J. Sterling, and Russell B. Millar. Short vs long penaeid trawls: Effects of side taper on engineering and catching performances. Fisheries Research, 134–136(??):73–81, December 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361200255X.

Broadhurst:2013:PME

[BSM13a] Matt K. Broadhurst, David J. Sterling, and Russell B. Millar. Progressing more environmentally benign penaeidtrawling systems by comparing Australian single- and multi-net configurations. Fisheries Research, 146(??):7–17, September 2013. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578361300088X.

Broadhurst:2013:REC

[BSM13b] Matt K. Broadhurst, David J. Sterling, and Russell B. Millar. Relative engineering and catching performances of paired penaeid-trawling systems. Fisheries Research, 143 (??):143–152, June 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783613000234.

Broadhurst:2014:ECI

[BSM14] Matt K. Broadhurst, David J. Sterling, and Russell B. Millar. Engineering and catch implications of variable wing-end spread on a penaeid trawl. Fisheries Research, 153(??):24– 30, May 2014. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783614000022.

Broadhurst:2015:EDP

[BSM15a] Matt K. Broadhurst, David J. Sterling, and Russell B. Millar. Effects of diel period and diurnal cloud cover on the species selection of short and long penaeid trawls. Fisheries Research, 170(??):144–151, October 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615001769.

Broadhurst:2015:ILM

[BSM15b] Matt K. Broadhurst, David J. Sterling, and Russell B. Millar. Increasing lateral mesh openings in penaeid trawls to improve selection and reduce drag. Fisheries Research, 170(??):68–75, October 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615001551.

Broadhurst:2015:TVN

[BSM15c] Matt K. Broadhurst, David J. Sterling, and Russell B. Millar. Traditional vs. novel ground gears: Maximising the environmental performance of penaeid trawls. Fisheries Research, 167(??):199–206, July 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615000594.

Broadhurst:2016:CEK

[BSM16] Matt K. Broadhurst, David J. Sterling, and Russell B. Millar. Confounding effects of knot orientation in penaeid trawls. Fisheries Research, 179(??):124–130, July 2016. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783616300509.

Broadhurst:2018:TSP

[BSM18] Matt K. Broadhurst, David J. Sterling, and Russell B. Millar. T45 side panels improve penaeid-trawl selection. Fisheries Research, 204(??):8–15, August 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618300225.

Breivik:2017:LGM

[BSN17] Olav Nikolai Breivik, Geir Storvik, and Kjell Nedreaas. Latent Gaussian models to predict historical bycatch in commercial fishery. Fisheries Research, 185(??):62–72, January 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783616303265.

Bacheler:2013:CTU

[BSS^{+}13] Nathan M. Bacheler, Christina M. Schobernd, Zeb H. Schobernd, Warren A. Mitchell, David J. Berrane, G. Todd Kellison, and Marcel J. M. Reichert. Comparison of trap and underwater video gears for indexing reef fish presence and abundance in the southeast United States. Fisheries Research, 143(??):81–88, June 2013. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613000167.

Bethoney:2013:DFS

[BSSH13] N. David Bethoney, Bradley P. Schondelmeier, Kevin D. E. Stokesbury, and William S. Hoffman. Developing a fine scale system to address river herring (Alosa pseudoharengus, A. aestivalis) and American shad (A. sapidissima) bycatch in the U.S. Northwest Atlantic mid-water trawl fishery. Fisheries Research, 141(??):79–87, April 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361200272X.

Balk:2017:SIE

[BST^{+}17] Helge Balk, Bendik S. Søvegjarto, Michal Tušer, Jaroslava Frouzová, Milan Muška, Vladislav Draštík, Roman Baran, and Jan Kubečka. Surface-induced errors in target strength and position estimates during horizontal acoustic surveys. Fisheries Research, 188(??):149–156, April 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616304362.

Becker:2019:DPE

[BST^{+}19] Alistair Becker, James A. Smith, Matthew D. Taylor, James McLeod, and Michael B. Lowry. Distribution of pelagic and epi-benthic fish around a multi-module artificial reeffield: Close module spacing supports a connected assemblage. Fisheries Research, 209(??):75–85, January 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783618302546.

Bernal:2012:EPM

[BSW^{+}12] M. Bernal, S. Somarakis, P. R. Witthames, C. J. G. van Damme, A. Uriarte, N. C. H. Lo, and M. Dickey-Collas. Egg production methods in marine fisheries: an introduction. Fisheries Research, 117–118(??):1–5, April 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612000148.

Brinson:2016:PFM

[BT16] Ayeisha A. Brinson and Eric M. Thunberg. Performance of federally managed catch share fisheries in the United States. Fisheries Research, 179(??):213–223, July 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616300649.

Boucek:2019:CRM

[BTB^{+}19] R. E. Boucek, A. A. Trotter, D. A. Blewett, J. L. Ritch, R. Santos, P. W. Stevens, J. A. Massie, and J. Rehage. Contrasting river migrations of common snook between two Florida rivers using acoustic telemetry. Fisheries Research, 213(??):219–225, May 2019. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618303618.

Babcock:2018:FCS

[BTBP18] Elizabeth A. Babcock, Alexander Tewfik, and Virginia Burns-Perez. Fish community and single-species indicators provide evidence of unsustainable practices in a multi-gear reef fishery. Fisheries Research, 208(??):70–85, December 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783618301930.

Bucholtz:2013:OFC

[BTNA13] R. Hagstrøm Bucholtz, J. Tomkiewicz, J. Randel Nyengaard, and J. Bremholm Andersen. Oogenesis, fecundity and condition of Baltic herring (Clupea harengus L.): a stereological study. Fisheries Research, 145(??):100–113, August 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783613000362.

Brooks:2019:PRS

[BTS^{+}19] Elizabeth N. Brooks, James T. Thorson, Kyle W. Shertzer, Richard D. M. Nash, Jon K. T. Brodziak, Kelli F. Johnson, Nikolai Klibansky, Brian K. Wells, and Jonathan White. Paulik revisited: Statistical framework and estimation performance of multistage recruitment functions. Fisheries Research, 217(??):58–70, September 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618301887.

Barry:2011:CTM

[BTT11] Patrick D. Barry, Sherry L. Tamone, and David A. Tallmon. A comparison of tagging methodology for North Pacific giant octopus Enteroctopus dofleini. Fisheries Research, 109(2–3):370–372, May 2011. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611000828.

Butterworth:2016:SAE

[But16] Doug S. Butterworth. The South African experience with MSC certification: a perspective. Fisheries Research, 182(??):124–127, October 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616300510.

Baker:2014:CPS

[BV14] M. R. Baker and C. H. Vynne. Cortisol profiles in sockeye salmon: Sample bias and baseline values at migration, maturation, spawning, and senescence. Fisheries Research, 154(??):38–43, June 2014. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361400023X.

Bastardie:2010:SBV

[BVN^{+}10] Francois Bastardie, Morten Vinther, J. Rasmus Nielsen, Clara Ulrich, and Marie Storr Paulsen. Stock-based vs. fleet-based evaluation of the multi-annual management plan for the cod stocks in the Baltic Sea. Fisheries Research, 101(3):188–202, January 15, 2010. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783609002732.

Battison:2017:ALC

[BW17] Andrea Battison and Carl Wilson. The American lobster in a changing ecosystem II: a US–Canada Science Symposium 2015. Fisheries Research, 186 (part 1)(??):356–357, February 2017. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783616303423.

Bryson:2013:CDD

[BWK^{+}13] Amanda J. Bryson, Christa M. Woodley, Rhonda K. Karls, Kathleen D. Hall, Mark A. Weiland, Z. Daniel Deng, Thomas J. Carlson, and M. Brad Eppard. Comparison of 180-degree and 90-degree needle rotation to reduce wound size in PIT-injected juvenile Chinook salmon. Fisheries Research, 143(??):201–204, June 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613000635.

Ballagh:2012:SSB

[BWN^{+}12] Aaron C. Ballagh, David J. Welch, Stephen J. Newman, Quentin Allsop, and Jason M. Stapley. Stock structure of the blue threadfin (Eleutheronema tetradactylum) across northern Australia derived from life-history characteristics. Fisheries Research, 121–122(??):63–72, June 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612000422.

Boldt:2018:DSC

[BWR^{+}18] Jennifer L. Boldt, Kresimir Williams, Christopher N. Rooper, Richard H. Towler, and Stéphane Gauthier. Development of stereo camera methodologies to improve pelagic fish biomass estimates and inform ecosystem management in marine waters. Fisheries Research, 198(??):66–77, February 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783617302886.

Brenden:2012:SBE

[BZ12] Travis O. Brenden and Yingming Zhao. Simulation-based evaluation of the accuracy in indirectly estimating gillnet selectivity. Fisheries Research, 134–136(??):64–72, December 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783612002500.

Catalano:2010:SAS

[CA10] Matthew J. Catalano and Micheal S. Allen. A size- and agestructured model to estimate fish recruitment, growth, mortality, and gear selectivity. Fisheries Research, 105(1):38– 45, June 2010. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783610000664.

Cosgrove:2014:NIB

[CAA^{+}14] Ronan Cosgrove, Igor Arregui, Haritz Arrizabalaga, Nicolas Goni, and Michael Sheridan. New insights to behaviour of North Atlantic albacore tuna (Thunnus alalunga) observed with pop-up satellite archival tags. Fisheries Research, 150(??):89–99, February 2014. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613002403.

Cosgrove:2015:PPS

[CAA^{+}15] Ronan Cosgrove, Igor Arregui, Haritz Arrizabalaga, Nicolas Goni, and John D. Neilson. Predation of pop-up satellite archival tagged albacore (Thunnus alalunga). Fisheries Research, 162(??):48–52, February 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614002707.

Cashion:2018:RGM

[CAAB^{+}18] Tim Cashion, Dalal Al-Abdulrazzak, Dyhia Belhabib, Brittany Derrick, Esther Divovich, Dimitrios K. Moutopoulos, Simon-Luc Noël, Maria L. Deng Palomares, Lydia C. L. Teh, Dirk Zeller, and Daniel Pauly. Reconstructing global marine fishing gear use: Catches and landed values by gear type and sector. Fisheries Research, 206(??):57–64, October 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783618301097.

Cotton:2014:ARD

[CAC^{+}14] Charles F. Cotton, Allen H. Andrews, Gregor M. Cailliet, R. Dean Grubbs, Sarah B. Irvine, and John A. Musick. Assessment of radiometric dating for age validation of deep-water dogfish (order: Squaliformes) finspines. Fisheries Research, 151(??):107–113, March 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613002415.

Cavole:2015:HIT

[CAC15] Letícia Maria Cavole, Caroline Chaves Arantes, and Leandro Castello. How illegal are tropical small-scale fisheries? An estimate for arapaima in the Amazon. Fisheries Research, 168(??):1–5, August 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615001022.

Chagaris:2019:MTC

[CAC19] David Chagaris, Micheal Allen, and Edward Camp. Modeling temporal closures in a multispecies recreational fishery reveals tradeoffs associated with species seasonality and angler effort dynamics. Fisheries Research, 210(??):106–120, February 2019. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578361830287X.

Catalan:2018:PFR

[CADG^{+}18] Ignacio A. Catalán, Josep Alós, Carlos Díaz-Gil, Silvia Pérez-Mayol, Gotzon Basterretxea, Beatriz Morales-Nin, and Miquel Palmer. Potential fishing-related effects on fish life history revealed by otolith microchemistry. Fisheries Research, 199(??):186–195, March 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617303168.

Canales-Aguirre:2010:PSD

[CAFHG10] Cristian B. Canales-Aguirre, Sandra Ferrada, Cristián E. Hernández, and Ricardo Galleguillos. Population structure and demographic history of Genypterus blacodes using microsatellite loci. Fisheries Research, 106(1):102–106, October 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783610001402.

Caballero-Alfonso:2010:RCV

[CAGTS^{+}10] A. M. Caballero-Alfonso, U. Ganzedo, A. Trujillo-Santana, J. Polanco, A. Santana del Pino, G. Ibarra-Berastegi, and J. J. Castro-Hernández. The role of climatic variability on the short-term fluctuations of octopus captures at the Canary Islands. Fisheries Research, 102(3):258–265, March 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783609003221.

Crespi-Abril:2014:SDC

[CAGVB14] A. C. Crespi-Abril, G. L. Villanueva Gomila, L. A. Venerus, and P. J. Barón. Spatial distribution of cephalopod paralarvae in San José Gulf (Northern Patagonia, Argentina): the role of tidal circulation in larval dispersal. Fisheries Research, 152(??):13–20, April 2014. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613001665.

Carvalho:2014:ISC

[CAM^{+}14] Felipe Carvalho, Robert Ahrens, Debra Murie, José M. Ponciano, Alexandre Aires da Silva, Mark N. Maunder, and Fábio Hazin. Incorporating specific change points in catchability in fisheries stock assessment models: an alternative approach applied to the blue shark (Prionace glauca) stock in the south Atlantic Ocean. Fisheries Research, 154(??):135– 146, June 2014. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783614000381.

Campbell:2015:CSA

[Cam15] Robert A. Campbell. Constructing stock abundance indices from catch and effort data: Some nuts and bolts. Fisheries Research, 161(??):109–130, January 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614002185.

Camp:2019:IFP

[CAM^{+}19] Edward V. Camp, Robert N. M. Ahrens, Timothy C. Mac-Donald, Kevin A. Thompson, and Kai Lorenzen. Identifying forage populations of concern: a new perspective based on predator recruitment considerations. Fisheries Research, 219(??):Article 105319, November 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619301663.

Casal:2018:DAP

[CAMFMF18] Gema Casal, Sara Aceña-Matarranz, Diana Fernández-Márquez, and Nuria Fernández. Distribution and abundance patterns of three coexisting species of Patella (Mollusca Gastropoda) in the intertidal areas of the NW Iberian Peninsula: Implications for management. Fisheries Research, 198(??):86–98, February 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617302862.

Carruthers:2011:IIS

[CAMW11] Thomas R. Carruthers, Robert N. M. Ahrens, Murdoch K. McAllister, and Carl J. Walters. Integrating imputation and standardization of catch rate data in the calculation of relative abundance indices. Fisheries Research, 109(1):157–167, April 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783611000531.

Cerda:2010:MRF

[CAP^{+}10] Margalida Cerdà, Josep Alós, Miquel Palmer, Antoni Maria Grau, and Francisco Riera. Managing recreational fisheries through gear restrictions: the case of limiting hook size in the recreational fishery from the Balearic Islands (NW Mediterranean). Fisheries Research, 101(3):146–155, January 15, 2010. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783609002628.

Carruthers:2018:MCR

[Car18]  Thomas R. Carruthers.  A multispecies catch-ratio

estimator of relative stock depletion. Fisheries Research, 197(??):25–33, January 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783617302655.

Colwell:2019:UCS

[CAR19] Julia M. Novak Colwell, Mark Axelrod, and Brian Roth. Unintended consequences of a seasonal ban on fishing effort in Tamil Nadu & Puducherry, India. Fisheries Research, 212(??):72–80, April 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783618303473.

Curran:2011:ECH

[CB11] Daniel Curran and Keith Bigelow. Effects of circle hooks on pelagic catches in the Hawaii-based tuna longline fishery. Fisheries Research, 109(2–3):265–275, May 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611000841.

Curtis:2017:ICG

[CB17] John Curtis and Benjamin Breen. Irish coarse and game anglers’ preferences for fishing site attributes. Fisheries Research, 190(??):103–112, June 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617300255.

Cameron:2016:UOM

[CBB^{+}16] Leo Cameron, Lee Baumgartner, Gavin Butler, Daniel Bucher, and Dong-Chun Lou. Using otolith microchemistry to differentiate between stocked and unstocked Australian bass (Percalates novemaculeata). Fisheries Research, 183(??):86–91, November 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616301515.

Colautti:2010:PBF

[CBC^{+}10] Dario Colautti, Claudio Baigun, Andrea Lopez Cazorla, Facundo Llompart, Juan M. Molina, Patricia Suquele, and Sebastián Calvo. Population biology and fishery characteristics of the smooth-hound Mustelus schmitti in anegada Bay, Argentina. Fisheries Research, 106(3):351–357, December 2010. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783610002134.

Cook:2014:CIM

[CBD^{+}14] Katrina V. Cook, Richard S. Brown, Z. Daniel Deng, Ryan S. Klett, Huidong Li, Adam G. Seaburg, and M. Brad Eppard. A comparison of implantation methods for large PIT tags or injectable acoustic transmitters in juvenile Chinook salmon. Fisheries Research, 154(??):213–223, June 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783613002841.

Carson:2014:STV

[CBF^{+}14] Evan W. Carson, Britt W. Bumguardner, Mark Fisher, Eric Saillant, and John R. Gold. Spatial and temporal variation in recovery of hatchery-released red drum (Sciaenops ocellatus) in stock-enhancement of Texas bays and estuaries. Fisheries Research, 151(??):191–198, March 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613002907.

Cuetos-Bueno:2018:DES

[CBH18] Javier Cuetos-Bueno and Peter Houk. Disentangling economic, social, and environmental drivers of coral-reef fish trade in Micronesia. Fisheries Research, 199(??):263–270, March 2018. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783617302850.

Cosgrove:2019:GTH

[CBM^{+}19] Ronán Cosgrove, Daragh Browne, Cóilín Minto, Peter Tyndall, Martin Oliver, Mike Montgomerie, and Matthew McHugh. A game of two halves: Bycatch reduction in Nephrops mixed fisheries. Fisheries Research, 210(??):31–40, February 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783618302558.

Cooper:2013:ASR

[CBMLB13] Wade T. Cooper, Luiz R. Barbieri, Michael D. Murphy, and Susan K. Lowerre-Barbieri. Assessing stock reproductive potential in species with indeterminate fecundity: Effects of age truncation and size-dependent reproductive timing. Fisheries Research, 138(??):31–41, February 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612001725.

Chang:2015:MSM

[CBO^{+}15] Yi-Jay Chang, Jon Brodziak, Joseph O’Malley, Hui-Hua Lee, Gerard DiNardo, and Chi-Lu Sun. Model selection and multi-model inference for Bayesian surplus production models: a case study for Pacific blue and striped marlin. Fisheries Research, 166(??):129–139, June 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614002781.

Correia:2011:SDE

[CBS11] A. T. Correia, F. Barros, and A. N. Sial. Stock discrimination of European conger eel (Conger conger L.) using otolith stable isotope ratios. Fisheries Research, 108(1):88–94, February 2011. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783610003140.

Corten:2017:DFI

[CBS17] Ad Corten, Cheikh-Baye Braham, and Ahmed Sidi Sadegh. The development of a fishmeal industry in Mauritania and its impact on the regional stocks of sardinella and other small pelagics in Northwest Africa. Fisheries Research, 186 (part 1)(??):328–336, February 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361630337X.

Colotelo:2011:ECA

[CC11] Alison H. Colotelo and Steven J. Cooke. Evaluation of common angling-induced sources of epithelial damage for popular freshwater sport fish using fluorescein. Fisheries Research, 109(2–3):217–224, May 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610003176.

Chrysafi:2019:TAD

[CC19] Anna Chrysafi and Jason M. Cope. Testing approaches to determine relative stock abundance priors when setting catch recommendations using data-limited methods. Fisheries Research, 219(??):Article 105343, November 2019. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619301985.

Canales:2016:STM

[CCA16a] Cristian M. Canales, Joan B. Company, and Patricio M. Arana. Spatio-temporal modelling of the maturity, sex ratio, and physical condition of nylon shrimp Heterocarpus reedi (Decapoda, Caridea), off Central Chile. Fisheries Research, 179(??):1–9, July 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616300236.

Canales:2016:ULB

[CCA16b] Cristian M. Canales, Joan B. Company, and Patricio M. Arana. Using a length-based stock assessment model to evaluate population structure hypotheses of nylon shrimp Heterocarpus reedi (Decapoda, Caridea) exploited off central Chile. Fisheries Research, 183(??):360–370, November 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783616302090.

Claramunt:2014:VSP

[CCC^{+}14a] Gabriel Claramunt, Luis A. Cubillos, Leonardo Castro, Carola Hernández, and Marcos Arteaga. Variation in the spawning periods of Engraulis ringens and Strangomera bentincki off the coasts of Chile: a quantitative analysis. Fisheries Research, 160(??):96–102, December 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613002208.

Coll:2014:CRR

[CCC^{+}14b] M. Coll, M. Carreras, M. J. Cornax, E. Massutí, E. Morote, X. Pastor, A. Quetglas, R. Sáez, L. Silva, I. Sobrino, M. A. Torres, S. Tudela, S. Harper, D. Zeller, and D. Pauly. Closer to reality: Reconstructing total removals in mixed fisheries from Southern Europe. Fisheries Research, 154(??):179–194, June 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783614000216.

Cubillos:2014:SFI

[CCC14c] Luis A. Cubillos, Gabriel Claramunt, and Leonardo R. Castro. Simulation of fishery-induced changes on the reproductive cycle of common sardine, Strangomera bentincki, off central southern Chile. Fisheries Research, 160(??):103–111, December 2014. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578361300297X.

Chiu:2017:MPD

[CCC17] Tsan-Yu Chiu, Tai-Sheng Chiu, and Chih-Shin Chen. Movement patterns determine the availability of Argentine shortfin squid Illex argentinus to fisheries. Fisheries Research, 193(??):71–80, September 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617300905.

Canales:2019:ASA

[CCC^{+}19] Cristian M. Canales, Luis A. Cubillos, María José Cuevas, Nicolás Adasme, and Nazareth Sánchez. Applying a separability assumption in a length-based stock assessment model to evaluate intra-annual effects of recruitment process error of small-pelagic fish. Fisheries Research, 217(??):108–121, September 2019. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783618302911.

Crowley:2019:EFS

[CCGD19] Claire Crowley, David Chagaris, Ryan Gandy, and Kendra Daly. Evaluation of Florida stone crab life history and management scenarios using spawning potential ratios. Fisheries Research, 209(??):196–207, January 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618302522.

Claramunt:2019:SMP

[CCHD19] Gabriel Claramunt, Luis Cubillos, Gustavo Herrera, and Eduardo Díaz. Spawning marker patterns of Engraulis ringens of northern Chile. Fisheries Research, 219 (??):Article 105306, November 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783619301535.

Colonello:2014:SRR

[CCM14] Jorge H. Colonello, Federico Cortés, and Ana M. Massa. Species richness and reproductive modes of chondrichthyans in relation to temperature and fishing effort in the Southwestern Atlantic Shelf (34–54^{o}S). Fisheries Research, 160(??):8–17, December 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783614001519.

Courtney:2014:EBR

[CCT^{+}14] A. J. Courtney, M. J. Campbell, M. L. Tonks, D. P. Roy, S. W. Gaddes, J. A. Haddy, P. M. Kyne, D. G. Mayer, and K. E. Chilcott. Effects of bycatch reduction devices in Queensland’s (Australia) deepwater eastern king prawn (Melicertus plebejus) trawl fishery. Fisheries Research, 157(??):113–123, September 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614001143.

Cooper:2014:LCA

[CDB^{+}14] Joyce Cooper, Steven Diesburg, Alaric Babej, Michael Noon, Ezra Kahn, Maureen Puettmann, and John Colt. Life cycle assessment of products from Alaskan salmon processing wastes: Implications of coproduction, intermittent landings, and storage time. Fisheries Research, 151(??):26–38, March 2014. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783613002932.

Chang:2013:PSS

[CDF^{+}13] Shui-Kai Chang, Gerard DiNardo, Jessica Farley, Jon Brodziak, and Zih-Lun Yuan. Possible stock structure of dolphinfish (Coryphaena hippurus) in Taiwan coastal waters and globally based on reviews of growth parameters. Fisheries Research, 147(??):127–136, October 2013. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783613001355.

Chang:2021:PBA

[CDL21] Shui-Kai Chang, Gerard DiNardo, and Ta-Te Lin. Photobased approach as an alternative method for collection of albacore (Thunnus alalunga) length frequency from longline vessels. Fisheries Research, 105(3):148–155, ???? 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783610000858.

Casarsa:2019:MDS

[CDM^{+}19] Leandro Casarsa, Mariano J. Diez, Adrián Madirolas, Ariel G. Cabreira, and Claudio C. Buratti. Morphometric description of schools from two different stocks of the southernmost sprat Sprattus fuegensis. Fisheries Research, 212(??):29–34, April 2019. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618303485.

Clements:2018:IMH

[CDMC18] Jeff C. Clements, John D. P. Davidson, Jarrod Gunn Mc-Quillan, and Luc A. Comeau. Increased mortality of harvested eastern oysters (Crassostrea virginica) is associated with air exposure and temperature during a spring fishery in Atlantic Canada. Fisheries Research, 206(??):27–34, October 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783618301280.

Cervino:2013:IEP

[CDPJ^{+}13] Santiago Cerviño, Rosario Domínguez-Petit, Ernesto Jardim, Sonia Mehault, Carmen Piñeiro, and Fran Saborido-Rey. Impact of egg production and stock structure on MSY reference points and its management implications for southern hake (Merluccius merluccius). Fisheries Research, 138(??):168–178, February 2013. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783612002226.

Currie:2011:RIE

[CDR^{+}11] David R. Currie, Cameron D. Dixon, Shane D. Roberts, Graham E. Hooper, Shirley J. Sorokin, and Timothy M. Ward. Relative importance of environmental gradients and historical trawling effort in determining the composition and distribution of benthic macro-biota in a large inverse estuary. Fisheries Research, 107(1–3):184–195, January 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783610002845.

Cadrin:2016:IFA

[CDR16] Steven X. Cadrin, Gregory R. DeCelles, and David Reid. Informing fishery assessment and management with field observations of selectivity and efficiency. Fisheries Research, 184(??):9–17, December 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615300734.

Cruz:2018:STD

[CEAP18] Francisco Santa Cruz, Billy Ernst, Javier A. Arata, and Carolina Parada. Spatial and temporal dynamics of the Antarctic krill fishery in fishing hotspots in the Bransfield Strait and South Shetland Islands. Fisheries Research, 208(??):157–166, December 2018. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618302194.

Cheng:2019:CSS

[CEB^{+}19] Zhaohai Cheng, Haraldur Arnar Einarsson, Shannon Bayse, Bent Herrmann, and Paul Winger. Comparing size selectivity of traditional and knotless diamond-mesh codends in the Iceland redfish (Sebastes spp.) fishery. Fisheries Research, 216(??):138–144, August 2019. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619301067.

Carvalho:2011:DSF

[CEJI11] Natacha Carvalho, Gareth Edwards-Jones, and Eduardo Isidro. Defining scale in fisheries: Small versus largescale fishing operations in the Azores. Fisheries Research, 109(2–3):360–369, May 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783611001159.

Catchpole:2011:CID

[CEM^{+}11] T. L. Catchpole, R. Enever, D. L. Maxwell, M. J. Armstrong, A. Reese, and A. S. Revill. Constructing indices to detect temporal trends in discarding. Fisheries Research, 107(1–3):94–99, January 2011. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610002717.

Cervino:2014:EGS

[Cer14] Santiago Cerviño. Estimating growth from sex ratio-at-length data in species with sexual size dimorphism. Fisheries Research, 160(??):112–119, December 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613002889.

Clarke:2014:RED

[CEW14] Tayler M. Clarke, Mario Espinoza, and Ingo S. Wehrtmann. Reproductive ecology of demersal elasmobranchs from a data-deficient fishery, Pacific of Costa Rica, Central America. Fisheries Research, 157(??):96–105, September 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783614001295.

Campobasso:2017:MLT

[CFB^{+}17] Francesco Campobasso, Annarita Fanizzi, Giambattista Bello, Nicoletta Santamaria, and Aldo Corriero. A ‘machine learning’ technique for discriminating captive-reared from wild Atlantic bluefin tuna, Thunnus thynnus (Osteichthyes: Scombridae), based on differential fin spine bone resorption. Fisheries Research, 194(??):42–49, October 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617301352.

Cahalan:2016:FTF

[CFBS16] Jennifer Cahalan, Craig Faunce, Julie Bonney, and Robert Swanson. A field test of fisheries observer sampling methods for estimation of at-sea discards. Fisheries Research, 174(??):219–233, February 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615301016.

Chaboud:2015:CFC

[CFF^{+}15] Christian Chaboud, Massal Fall, Jocelyne Ferraris, André Fontana, Alain Fonteneau, Francis Laloë, Alassane Samba, and Djiga Thiao. Comment on “Fisheries catch misreporting and its implications: the case of Senegal”. Fisheries Research, 164(??):322–324, April 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614003051.

Chiesa:2016:CDM

[CFF^{+}16] Stefania Chiesa, Laura Filonzi, Claudio Ferrari, Marina Vaghi, Fabiana Bilò, Armando Piccinini, Giulia Zuccon, Robert C. Wilson, Jørn Ulheim, and Francesco Nonnis Marzano. Combinations of distinct molecular markers allow to genetically characterize marble trout (Salmo marmoratus) breeders and stocks suitable for reintroduction plans. Fisheries Research, 176(??):55–64, April 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615301612.

Carruthers:2014:ETR

[CFH14] T. Carruthers, A. Fonteneau, and J. P. Hallier. Estimating tag reporting rates for tropical tuna fleets of the Indian Ocean. Fisheries Research, 155(??):20–32, July 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783614000514.

Carruthers:2015:RET

[CFH15] T. Carruthers, A. Fonteneau, and J. P. Hallier. Reprint of “Estimating tag reporting rates for tropical tuna fleets of the Indian Ocean”. Fisheries Research, 163(??):85–97, March 2015. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783614002148.

Campos:2015:STC

[CFPF^{+}15] Aida Campos, Paulo Fonseca, Tereza Pilar-Fonseca, Ana Maria Leocádio, and Margarida Castro. Survival of trawl-caught Norway lobster (Nephrops norvegicus L.) after capture and release — potential effect of codend mesh type on survival. Fisheries Research, 172(??):415–422, December 2015. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783615300473.

Condie:2013:DBD

[CGC13] Harriet M. Condie, Alastair Grant, and Thomas L. Catchpole. Does banning discards in an otter trawler fishery create incentives for more selective fishing? Fisheries Research, 148(??):137–146, November 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361300221X.

Carvalho:2017:NBG

[CGdGT^{+}17] Daniel Cardoso Carvalho, Danusa Guedes, Maria da Gloria Trindade, Regina Melo Sartori Coelho, and Paulo Humberto de Lima Araujo. Nationwide Brazilian governmental forensic programme reveals seafood mislabelling trends and rates using DNA barcoding. Fisheries Research, 191(??):30– 35, July 2017. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783617300589.

Czerwinski:2010:HSM

[CGECSEH10] Ivone A. Czerwinski, Juan C. Gutiérrez-Estrada, Mila Casimiro-Soriguer-Escofet, and José A. Hernando. Hook selectivity models assessment for black spot seabream. Classic and heuristic approaches. Fisheries Research, 102(1– 2):41–49, February 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783609002690.

Chandrapavan:2021:GRA

[CGG21] Arani Chandrapavan, Caleb Gardner, and Bridget S. Green. Growth rate of adult rock lobsters Jasus edwardsii increased through translocation. Fisheries Research, 105(3):244–247, ???? 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783610000780.

Carrera-Garcia:2017:TJS

[CGKG^{+}17] Erika Carrera-García, Juliette Kordek, Charline Gesset, Louis Jacobs, and Marie-Laure Acolas. Tracking juvenile sturgeon in the wild: Miniature tag effects assessment in a laboratory study on Siberian sturgeon (Acipenser baerii). Fisheries Research, 186 (part 1)(??):337–344, February 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783616303447.

Conti:2012:LTC

[CGLS12] Lorenza Conti, Gaël Grenouillet, Sovan Lek, and Michele Scardi. Long-term changes and recurrent patterns in fisheries landings from Large Marine Ecosystems (1950–2004). Fisheries Research, 119–120(??):1–12, May 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611003560.

Cadrin:2019:WDY

[CGM^{+}19] Steven X. Cadrin, Daniel R. Goethel, Molly R. Morse, Gavin Fay, and Lisa A. Kerr. “So, where do you come from?” the impact of assumed spatial population structure on estimates of recruitment. Fisheries Research, 217(??):156–168, September 2019. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783618303448.

Cosgrove:2015:SDB

[CGR^{+}15] Ronan Cosgrove, Martha Gosch, David Reid, Michael Sheridan, Nicholas Chopin, Mark Jessopp, and Michelle Cronin. Seal depredation in bottom-set gillnet and entangling net fisheries in Irish waters. Fisheries Research, 172(??):335–344, December 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615300497.

Cosgrove:2016:SBG

[CGR^{+}16] Ronan Cosgrove, Martha Gosch, David Reid, Michael Sheridan, Nicolas Chopin, Mark Jessopp, and Michelle Cronin. Seal bycatch in gillnet and entangling net fisheries in Irish waters. Fisheries Research, 183(??):192–199, November 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783616301965.

Cardoso:2011:ACG

[CH11] Luis Gustavo Cardoso and Manuel Haimovici. Age and changes in growth of the king weakfish Macrodon atricauda (Günther, 1880) between 1977 and 2009 in southern Brazil. Fisheries Research, 111(3):177–187, November 2011. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783611002384.

Cardoso:2014:LTC

[CH14] Luis Gustavo Cardoso and Manuel Haimovici. Long term changes in the sexual maturity and in the reproductive biomass of the southern king weakfish Macrodon atricauda (Günther, 1880) in southern Brazil. Fisheries Research, 160(??):120–128, December 2014. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614001659.

Cardoso:2015:LTC

[CH15] Luís Gustavo Cardoso and Manuel Haimovici. Long-term changes in the age structure, mortality and biomass of the king weakfish Macrodon atricauda (Günther, 1880) in southern Brazil: Is it resilient enough to avoid collapse? Fisheries Research, 167(??):174–179, July 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615000570.

Chabot:2015:LCE

[CHA15] Chris L. Chabot, Holly A. Hawk, and Larry G. Allen. Low contemporary effective population size detected in the critically endangered giant sea bass, Stereolepis gigas, due to fisheries overexploitation. Fisheries Research, 172(??):71–78, December 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615300011.

Correia:2014:EMP

[CHCS14] A. T. Correia, P. Hamer, B. Carocinho, and A. Silva. Evidence for meta-population structure of Sardina pilchardus in the Atlantic Iberian waters from otolith elemental signatures of a strong cohort. Fisheries Research, 149(??):76–85, January 2014. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783613002269.

Cheng:2010:IPF

[Che10] Yuk W. Cheng. Inverse prediction for fish length–length conversion. Fisheries Research, 106(1):112–114, October 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578361000158X.

Chih:2010:IES

[Chi10] Ching-Ping Chih. Incorporating effective sample sizes into sampling designs for reef fish. Fisheries Research, 105(2):102–110, July 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S016578361000072X.

Chih:2011:DEC

[Chi11] Ching-Ping Chih. The design effects of cluster sampling on the estimation of mean lengths and total mortality of reef fish. Fisheries Research, 109(2–3):295–302, May 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783611000877.

Chang:2011:CRS

[CHL11] Shui-Kai Chang, Simon Hoyle, and Hung-I. Liu. Catch rate standardization for yellowfin tuna (Thunnus albacares) in Taiwan’s distant-water longline fishery in the Western and Central Pacific Ocean, with consideration of target change. Fisheries Research, 107(1–3):210–220, January 2011. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783610002870.

Christou:2017:ESF

[CHS^{+}17] Maria Christou, John Haralabous, Konstantinos I. Stergiou, Dimitrios Damalas, and Christos D. Maravelias. An evaluation of socioeconomic factors that influence fishers’ discard behaviour in the Greek bottom trawl fishery. Fisheries Research, 195(??):105–115, November 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617301777.

Canales:2018:IMD

[CHT18] Cristian M. Canales, Claudia Hurtado, and Carlos Techeira. Implementing a model for data-poor fisheries based on steepness of the stock-recruitment relationship, natural mortality and local perception of population depletion. the case of the kelp Lessonia berteroana on coasts of north-central Chile. Fisheries Research, 198(??):31–42, February 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617302928.

Campbell:2010:RCD

[CHW^{+}10] R. Campbell, T. Harcus, D. Weirman, R. J. Fryer, R. J. Kynoch, and F. G. O’Neill. The reduction of cod discards by inserting 300 mm diamond mesh netting in the forward sections of a trawl gear. Fisheries Research, 102(1–2):221–226, February 2010. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783609003178.

Cook:2018:ECH

[CHW^{+}18] Katrina V. Cook, Scott G. Hinch, Maryann S. Watson, David A. Patterson, Andrea J. Reid, and Steve J. Cooke. Experimental capture and handling of chum salmon reveal thresholds in injury, impairment, and physiology: Best practices to improve bycatch survival in a purse seine fishery. Fisheries Research, 206(??):96–108, October 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618301243.

Coelho:2013:AGL

[CIS13] Rui Coelho, Paulo Infante, and Miguel N. Santos. Application of generalized linear models and generalized estimation equations to model at-haulback mortality of blue sharks captured in a pelagic longline fishery in the Atlantic Ocean. Fisheries Research, 145(??):66–75, August 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613000532.

Castillo-Jordan:2021:ICG

[CJCN21] Claudio Castillo-Jordán, Luis A. Cubillos, and Eduardo Navarro. Inter-cohort growth rate changes of common sardine (Strangomera bentincki) and their relationship with environmental conditions off central southern Chile. Fisheries Research, 105(3):228–236, ???? 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610001293.

Clark:2015:CAL

[CJG^{+}15] Abigail S. Clark, Steven H. Jury, Jason S. Goldstein, Thomas G. Langley, and Winsor H. Watson. A comparison of American lobster size structure and abundance using standard and ventless traps. Fisheries Research, 167(??):243–251, July 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783615000612.

Coppinger:2019:AGD

[CJG^{+}19] C. R. Coppinger, N. James, G. Gouws, B. Q. Mann, A. Guissamulo, and M. Mwale. Assessing the genetic diversity of catface grouper Epinephelus andersoni in the subtropical Western Indian Ocean. Fisheries Research, 218(??):186–197, October 2019. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783619301249.

Cinar:2013:DMF

[CJP13] E. Mine Cinar, Joseph Johnson, and Andrew Palmer. Decision making: Fishing production and fishers in the Black Sea. Fisheries Research, 147(??):296–303, October 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578361300163X.

Castillo-Jordan:2019:ICI

[CJWT^{+}19] Claudio Castillo-Jordán, Sally E. Wayte, Geoffrey N. Tuck, Sean Tracey, Stewart D. Frusher, and André E. Punt. Implications of a climate-induced recruitment shift in the stock assessment of Patagonian grenadier (Macruronus magellanicus) in Chile. Fisheries Research, 212(??):114–122, April 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783618303692.

Crane:2018:CEF

[CK18] Derek P. Crane and Kevin L. Kapuscinski. Capture efficiency of a fine mesh seine in a large river: Implications for abundance, richness, and diversity analyses. Fisheries Research, 205(??):149–157, September 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618301231.

Cournane:2013:STP

[CKC13] Jamie M. Cournane, Jacob P. Kritzer, and Steven J. Correia. Spatial and temporal patterns of anadromous alosine bycatch in the US Atlantic herring fishery. Fisheries Research, 141(??):88–94, April 2013. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612002342.

Crane:2016:ETF

[CKC16] Derek P. Crane, Christina C. Killourhy, and Mark D. Clapsadl. Effects of three frozen storage methods on wet weight of fish. Fisheries Research, 175(??):142–147, March 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783615301466.

Coates:2016:STE

[CKVV16] Delphine A. Coates, Danae-Athena Kapasakali, Magda Vincx, and Jan Vanaverbeke. Short-term effects of fishery exclusion in offshore wind farms on macrofaunal communities in the Belgian part of the North Sea. Fisheries Research, 179(??):131–138, July 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616300492.

Conners:2017:CDM

[CL17] M. Elizabeth Conners and Michael Levine. Characteristics and discard mortality of octopus bycatch in Alaska groundfish fisheries. Fisheries Research, 185(??):169–175, January 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783616303034.

Clark:2014:DCR

[Cla14] William G. Clark. Direct calculation of relative fishery and survey selectivities. Fisheries Research, 158(??):135–137, October 2014. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783613002452.

Charters:2010:SSG

[CLB^{+}10] Robbie A. Charters, Robert J. G. Lester, Rik C. Buckworth, Stephen J. Newman, Jennifer R. Ovenden, Damien Broderick, Olena Kravchuk, Aaron Ballagh, and David J. Welch. The stock structure of grey mackerel Scomberomorus semifasciatus in Australia as inferred from its parasite fauna. Fisheries Research, 101(1–2):94–99, January 5, 2010. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609002574.

Chiesa:2017:HIG

[CLF^{+}17] Stefania Chiesa, Livia Lucentini, Rosa Freitas, Francesco Nonnis Marzano, Silvia Breda, Etelvina Figueira, Nathalie Caill-Milly, Roger J. H. Herbert, Amadeu M. V. M. Soares, and Emanuele Argese. A history of invasion: COI phylogeny of Manila clam Ruditapes philippinarum in Europe. Fisheries Research, 186 (part 1)(??):25–35, February 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616302363.

Caudron:2014:TYR

[CLG^{+}14] Arnaud Caudron, Emilien Lasne, Christian Gillet, Jean Guillard, and Alexis Champigneulle. Thirty years of reoligotrophication do not contribute to restore self-sustaining fisheries of Arctic charr, Salvelinus alpinus, in Lake Geneva. Fisheries Research, 154(??):165–171, June 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614000393.

Capoccioni:2019:FMS

[CLP^{+}19] Fabrizio Capoccioni, Chiara Leone, Domitilla Pulcini, Massimo Cecchetti, Alessandro Rossi, and Eleonora Ciccotti. Fish movements and schooling behavior across the tidal channel in a Mediterranean coastal lagoon: an automated approach using acoustic imaging. Fisheries Research, 219(??):Article 105318, November 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619301651.

Chang:2011:PSS

[CLS^{+}11] Wei-Che Chang, Ying-Chou Lee, Chun-Han Shih, Ta-Jen Chu, and Ping-Hung Chang. Population size and stocking contribution rates for marked and recaptured black porgy Acanthopagrus schlegelli, in Northwestern Taiwan, 2005– 2008. Fisheries Research, 109(2–3):252–256, May 2011. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783611000804.

Canel:2019:SMD

[CLS^{+}19] Delfina Canel, Eugenia Levy, Iris A. Soares, Paola E. Braicovich, Manuel Haimovici, José L. Luque, and Juan T. Timi. Stocks and migrations of the demersal fish Umbrina canosai (Sciaenidae) endemic from the subtropical and temperate Southwestern Atlantic revealed by its parasites. Fisheries Research, 214(??):10–18, June 2019. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619300384.

Canales:2015:CSS

[CLWB15] T. Mariella Canales, Richard Law, Rodrigo Wiff, and Julia L. Blanchard. Changes in the size-structure of a multispecies pelagic fishery off Northern Chile. Fisheries Research, 161(??):261–268, January 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614002501.

Chang:2012:AMM

[CM12] Shui-Kai Chang and Mark N. Maunder. Aging material matters in the estimation of von Bertalanffy growth parameters for dolphinfish (Coryphaena hippurus). Fisheries Research, 119–120(??):147–153, May 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612000392.

Claus:2015:NEA

[CM15] Ubl Claus and Dorow Malte. A novel enclosure approach to assessing yellow eel (Anguilla anguilla) density in nontidal coastal waters. Fisheries Research, 161(??):57–63, January 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783614001970.

Cassinelli:2018:FIR

[CM18] John D. Cassinelli and Kevin A. Meyer. Factors influencing return-to-creel of hatchery catchable-sized rainbow trout stocked in Idaho lentic waters. Fisheries Research, 204(??):316–323, August 2018. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618300791.

Cabreira:2011:ACA

[CMB11a] A. G. Cabreira, A. Madirolas, and N. E. Brunetti. Acoustic characterization of the Argentinean short-fin squid aggregations. Fisheries Research, 108(1):95–99, February 2011.

CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783610003152.

Castello:2011:RSS

[CMB11b] Leandro Castello, David G. McGrath, and Pieter S. A. Beck. Resource sustainability in small-scale fisheries in the Lower Amazon floodplains. Fisheries Research, 110(2):356–364, July 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783611001913.

Chick:2012:DCD

[CMB^{+}12] R. C. Chick, S. Mayfield, P. Burch, S. N. Turich, and R. Mc-Garvey. Detecting change in density and biomass of a benthic marine invertebrate following commercial fishing. Fisheries Research, 129–130(??):94–105, October 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612001932.

Chizinski:2014:SIL

[CMHP14] Christopher J. Chizinski, Dustin R. Martin, Keith L. Hurley, and Kevin L. Pope. Self-imposed length limits in recreational fisheries. Fisheries Research, 155(??):83–89, July 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578361400068X.

Catanese:2016:HRS

[CMI^{+}16] Gaetano Catanese, Iratxe Montes, Mikel Iriondo, Andone Estonba, Daniele Iudicone, and Gabriele Procaccini. High resolution SNPs selection in Engraulis encrasicolus through Taqman OpenArray. Fisheries Research, 177(??):31–38, May 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783616300145.

Cook:2019:VFL

[CMJ^{+}19] Denham Cook, Karen Middlemiss, Peter Jaksons, William Davison, and Alistair Jerrett. Validation of fish length estimations from a high frequency multi-beam sonar (ARIS) and its utilisation as a field-based measurement technique.

Fisheries Research, 218(??):59–68, October 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619301213.

Crone:2019:GPI

[CMLP19] Paul R. Crone, Mark N. Maunder, HuiHua Lee, and Kevin R. Piner. Good practices for including environmental data to inform spawner-recruit dynamics in integrated stock assessments: Small pelagic species case study. Fisheries Research, 217(??):122–132, September 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618303825.

Chiesa:2011:IMC

[CMM^{+}11] Stefania Chiesa, Francesco Nonnis Marzano, Giovanni Minervini, Davide De Lucrezia, Gianluca Baccarani, Guido Bordignon, Irene Poli, Giampietro Ravagnan, and Emanuele Argese. The invasive Manila clam Ruditapes philippinarum (Adams and Reeve, 1850) in Northern Adriatic Sea: Population genetics assessed by an integrated molecular approach. Fisheries Research, 110(2):259–267, July 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611001627.

Costa:2014:EEL

[CMM^{+}14] M. D. P. Costa, J. H. Muelbert, L. E. Moraes, J. P. Vieira, and J. P. Castello. Estuarine early life stage habitat occupancy patterns of whitemouth croaker Micropogonias furnieri (Desmarest, 1830) from the Patos Lagoon, Brazil. Fisheries Research, 160(??):77–84, December 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613002543.

Cambie:2013:SMW

[CMMF13] Giulia Cambiè, Ramón Muiño, Toni Mingozzi, and Juan Freire. From surface to mid-water: Is the swordfish longline fishery “hitting rock bottom”? A case study in southern Italy. Fisheries Research, 140(??):114–122, February 2013.

CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783613000027.

Cisneros-Mata:2019:GDM

[CMMVRF^{+}19] Miguel Ángel Cisneros-Mata, Adrián Munguía-Vega, Demetrio Rodríguez-Félix, Eugenio Alberto Aragón-Noriega, José Manuel Grijalva-Chon, José Alfredo Arreola-Lizárraga, and Luis A. Hurtado. Genetic diversity and metapopulation structure of the brown swimming crab (Callinectes bellicosus) along the coast of Sonora, Mexico: Implications for fisheries management. Fisheries Research, 212(??):97–106, April 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783618303333.

Castro:2012:IMB

[CMP^{+}12] J. Castro, M. Marín, N. Pérez, G. J. Pierce, and A. Punzón. Identification of métiers based on economic and biological data: the Spanish bottom otter trawl fleet operating in non-Iberian European waters. Fisheries Research, 125–126(??):77–86, August 2012. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612000902.

Chizinski:2014:AEC

[CMP^{+}14] Christopher J. Chizinski, Dustin R. Martin, Kevin L. Pope, Tony J. Barada, and Jeffrey J. Schuckman. Angler effort and catch within a spatially complex system of small lakes. Fisheries Research, 154(??):172–178, June 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614000538.

Castro:2011:IMS

[CMPP11] J. Castro, M. Marín, G. J. Pierce, and A. Punzón. Identification of métiers of the Spanish set-longline fleet operating in non-Spanish European waters. Fisheries Research, 107(1–3):100–111, January 2011. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610002729.

Calo:2016:SGS

[CMPR^{+}16] Antonio Calò, Irene Muñoz, Ángel Pérez-Ruzafa, Carlos Vergara-Chen, and José Antonio García-Charton. Spatial genetic structure in the saddled sea bream (Oblada melanura [Linnaeus, 1758]) suggests multi-scaled patterns of connectivity between protected and unprotected areas in the Western Mediterranean Sea. Fisheries Research, 176(??):30–38, April 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783615301533.

Campbell:2014:STS

[CMS14] M. J. Campbell, M. F. McLennan, and W. D. Sumpton. Short-term survival of discarded pearl perch (Glaucosoma scapulare Ramsay, 1881) caught by hook-and-line in Queensland, Australia. Fisheries Research, 151(??):206–212, March 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783613002993.

Chiang:2015:SMD

[CMS^{+}15] Wei-Chuan Chiang, Michael K. Musyl, Chi-Lu Sun, Gerard DiNardo, Hsiao-Min Hung, Hsien-Chung Lin, Shin-Chi Chen, Su-Zan Yeh, Wen-Yie Chen, and Chin-Lau Kuo. Seasonal movements and diving behaviour of black marlin (Istiompax indica) in the northwestern Pacific Ocean. Fisheries Research, 166(??):92–102, June 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614003178.

Carreira:2018:PRM

[CMSG18] Gilberto P. Carreira, Niall J. McKeown, Paul W. Shaw, and João M. A. Gonçalves. A PCR-RFLP method for stock assignment in the endemic Macaronesian gastropod, Patella candei d’Orbigny, and its potential for future conservation strategies. Fisheries Research, 199(??):8–11, March 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783617303193.

Cavieres:2018:UST

[CN18] Joaquin Cavieres and Orietta Nicolis. Using a spatiotemporal Bayesian approach to estimate the relative abundance index of yellow squat lobster (Cervimunida johni) off Chile. Fisheries Research, 208(??):97–104, December 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783618301929.

Corell:2019:MLD

[CN19] Hanna Corell and Anders Nissling. Modelling of larval dispersal of Baltic flounder (Platichthys solemdali) revealed drifting depth as a major factor determining opportunities for local retention vs large-scale connectivity. Fisheries Research, 218(??):127–137, October 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619301262.

Cardinale:2011:SDF

[CNJ11] Massimiliano Cardinale, Duto Nugroho, and Patrik Jonson. Serial depletion of fishing grounds in an unregulated, open access fishery. Fisheries Research, 108(1):106–111, February 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578361000319X.

Carruthers:2011:OBM

[CNS11] Erin H. Carruthers, John D. Neilson, and Sean C. Smith. Overlooked bycatch mitigation opportunities in pelagic longline fisheries: Soak time and temperature effects on swordfish (Xiphias gladius) and blue shark (Prionace glauca) catch. Fisheries Research, 108(1):112–120, February 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783610003206.

Candy:2012:EAE

[CNWM12] Steven G. Candy, Gabrielle B. Nowara, Dirk C. Welsford, and John P. McKinlay. Estimating an ageing error matrix for Patagonian toothfish (Dissostichus eleginoides) otoliths using between-reader integer errors, readability scores, and continuation ratio models. Fisheries Research, 115–116(??):14–23, March 2012. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611003341.

Cross:2014:RBE

[COC14] M. E. Cross, R. M. O’Riordan, and S. C. Culloty. The reproductive biology of the exploited razor clam, Ensis siliqua, in the Irish Sea. Fisheries Research, 150(??):11–17, February 2014. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783613002300.

Cook:2013:FSA

[Coo13] R. M. Cook. A fish stock assessment model using survey data when estimates of catch are unreliable. Fisheries Research, 143(??):1–11, June 2013. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613000064.

Cope:2013:ISC

[Cop13] Jason M. Cope. Implementing a statistical catch-at-age model (Stock Synthesis) as a tool for deriving overfishing limits in data-limited situations. Fisheries Research, 142(??):3–14, May 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783612001257.

Cort:2017:LCR

[Cor17] José L. Cort. Letter critical of a recent Fisheries Research publication [150706-002900] [reference: 150706-002900]: With regard to your recent publication: Growth of bluefin tuna (Thunnus thynnus) in the North-eastern Atlantic and Mediterranean base on back-calculation of dorsal fin spine annuli by Jorge Landa, Enrique Rodriguez-Marin, Patricia L. Luque, Marta Ruiz, Pablo Quelle [Fisheries Research 170 (2015): 190–198]. Fisheries Research, 185(??):198–199, January 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0165783616302314. See [LRML^{+}15] and response [LRM17].

Chambers:2015:FME

[COS15] Mark S. Chambers, Ben O’Neill, and Leesa A. Sidhu. Frailty models for the estimation of tag shedding rates with tagger effects. Fisheries Research, 161(??):303–311, January 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783614002641.

Cockrell:2019:RCF

[COS^{+}19] Marcy L. Cockrell, Shay O’Farrell, James Sanchirico, Steven A. Murawski, Larry Perruso, and Andrew Strelcheck. Resilience of a commercial fishing fleet following emergency closures in the Gulf of Mexico. Fisheries Research, 218(??):69–82, October 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783619301146.

Cope:2011:RSA

[CP11] Jason M. Cope and André E. Punt. Reconciling stock assessment and management scales under conditions of spatially varying catch histories. Fisheries Research, 107(1–3):22–38, January 2011. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578361000250X. See erra-

tum [CP12].

Cope:2012:ERS

[CP12] Jason M. Cope and André E. Punt. Erratum to “Reconciling stock assessment and management scales under conditions of spatially varying catch histories” [Fish. Res. 107 (1–3) (2011) 22–38]. Fisheries Research, 134–136(??):133, December 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783612002792. See [CP11].

Colotelo:2012:ERS

[CPB^{+}12] Alison H. Colotelo, Brett D. Pflugrath, Richard S. Brown, Colin J. Brauner, Robert P. Mueller, Thomas J. Carlson, Z. Daniel Deng, Martin L. Ahmann, and Bradly A. Trumbo. The effect of rapid and sustained decompression on barotrauma in juvenile brook lamprey and Pacific lamprey:

Implications for passage at hydroelectric facilities. Fisheries Research, 129–130(??):17–20, October 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612001737.

Carvalho:2017:CDT

[CPC^{+}17] Felipe Carvalho, André E. Punt, Yi-Jay Chang, Mark N. Maunder, and Kevin R. Piner. Can diagnostic tests help identify model misspecification in integrated stock assessments? Fisheries Research, 192(??):28–40, August 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783616303113.

Crego-Prieto:2012:ILD

[CPCP^{+}12] Victor Crego-Prieto, Daniel Campo, Juliana Perez, Jose L. Martinez, Eva Garcia-Vazquez, and Agustin Roca. Inaccurate labelling detected at landings and markets: the case of European megrims. Fisheries Research, 129–130(??):106–109, October 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783612002019.

Carbonara:2019:SSE

[CPD^{+}19] Pierluigi Carbonara, Cristina Porcu, Marilena Donnaloia, Paola Pesci, Letizia Sion, Maria Teresa Spedicato, Walter Zupa, Francesca Vitale, and Maria Cristina Follesa. The spawning strategy of European hake (Merluccius merluccius, L. 1758) across the western and Central Mediterranean Sea. Fisheries Research, 219(??):Article 105333, November 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783619301882.

Correia:2011:IPS

[CPG^{+}11] A. T. Correia, T. Pipa, J. M. S. Gonçalves, K. Erzini, and P. A. Hamer. Insights into population structure of Diplodus vulgaris along the SW Portuguese coast from otolith elemental signatures. Fisheries Research, 111(1–2):82–91, September 2011. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783611002311.

Campbell:2015:CRA

[CPG^{+}15] Matthew D. Campbell, Adam G. Pollack, Christopher T. Gledhill, Theodore S. Switzer, and Douglas A. DeVries. Comparison of relative abundance indices calculated from two methods of generating video count data. Fisheries Research, 170(??):125–133, October 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615001526.

Coulson:2012:BCS

[CPH12] Peter G. Coulson, Ian C. Potter, and Norman G. Hall. The biological characteristics of Scorpis aequipinnis (Kyphosidae), including relevant comparisons with those of other species and particularly of a heavily exploited congener. Fisheries Research, 125–126(??):272–282, August 2012. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783612001233.

Cardona-Pons:2010:SRF

[CPMNS10] Federico Cardona-Pons, Beatriz Morales-Nin, and Stephen G. Sutton. Scientists and recreational fishers: Communication manners and its efficiency. Fisheries Research, 106(3):575–578, December 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783610002468.

Casaretto:2014:LSH

[CPOH14] Licia Casaretto, Marta Picciulin, Kjell Olsen, and Anthony D. Hawkins. Locating spawning haddock (Melanogrammus aeglefinus, Linnaeus, 1758) at sea by means of sound. Fisheries Research, 154(??):127–134, June 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614000502.

Castro:2010:IMN

[CPP^{+}10] José Castro, Antonio Punzón, Graham J. Pierce, Manuel Marín, and Esther Abad. Identification of métiers of the Northern Spanish coastal bottom pair trawl fleet by using the partitioning method CLARA. Fisheries Research, 102(1–2):184–190, February 2010. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783609003130.

Chosid:2012:DOS

[CPS^{+}12] David M. Chosid, Michael Pol, Mark Szymanski, Frank Mirarchi, and Andrew Mirarchi. Development and observations of a spiny dogfish Squalus acanthias reduction device in a raised footrope silver hake Merluccius bilinearis trawl. Fisheries Research, 114(??):66–75, February 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611001160.

Carruthers:2014:EMS

[CPW^{+}14] Thomas R. Carruthers, André E. Punt, Carl J. Walters, Alec MacCall, Murdoch K. McAllister, Edward J. Dick, and Jason Cope. Evaluating methods for setting catch limits in data-limited fisheries. Fisheries Research, 153(??):48–68, May 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783613003081.

Chamberlin:2014:ENO

[CQ14] Joshua W. Chamberlin and Thomas P. Quinn. Effects of natal origin on localized distributions of Chinook salmon, Oncorhynchus tshawytscha, in the marine waters of Puget Sound, Washington. Fisheries Research, 153(??):113–122, May 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783614000101.

Cornic:2018:IOC

[CR18] Maëlle Cornic and Jay R. Rooker. Influence of oceanographic conditions on the distribution and abundance of blackfin tuna (Thunnus atlanticus) larvae in the Gulf of Mexico. Fisheries Research, 201(??):1–10, May 2018. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783617303570.

Craig:2019:LCC

[Cra19] James R. Craig. A low cost 1-cm accuracy electronic fish measurement board. Fisheries Research, 210(??):198–203, February 2019. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783618302340.

Cabanellas-Reboredo:2011:SIH

[CRAP^{+}11] Miguel Cabanellas-Reboredo, Josep Alós, Miquel Palmer, Roman Grädel, and Beatriz Morales-Nin. Simulating the indirect handline jigging effects on the European squid Loligo vulgaris in captivity. Fisheries Research, 110(3):435–440, August 2011. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783611002025.

Contreras-Reyes:2013:GEC

[CRAV13] Javier E. Contreras-Reyes and Reinaldo B. Arellano-Valle. Growth estimates of cardinalfish (Epigonus crassicaudus) based on scale mixtures of skew-normal distributions. Fisheries Research, 147(??):137–144, October 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613001343.

Contreras-Reyes:2014:CGC

[CRAVC14] Javier E. Contreras-Reyes, Reinaldo B. Arellano-Valle, and T. Mariella Canales. Comparing growth curves with asymmetric heavy-tailed errors: Application to the southern blue whiting (Micromesistius australis). Fisheries Research, 159(??):88–94, November 2014. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614001593.

Christensen:2018:CTM

[CRB^{+}18] Helle Torp Christensen, Frank Rigét, Marie Balslev Backe, Atal Saha, Torild Johansen, and Rasmus B. Hedeholm.

Comparison of three methods for identification of redfish (Sebastes mentella and S. norvegicus) from the Greenland east coast. Fisheries Research, 201(??):11–17, May 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783618300031.

Cabanellas-Reboredo:2014:UAD

[CRCMP^{+}14] Miguel Cabanellas-Reboredo, Matías Calvo-Manazza, Miquel Palmer, Jorge Hernández-Urcera, Manuel E. Garci, Ángel F. González, Ángel Guerra, and Beatriz Morales-Nin. Using artificial devices for identifying spawning preferences of the European squid: Usefulness and limitations. Fisheries Research, 157(??):70–77, September 2014. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783614001131.

Cameron:2018:TMP

[CRG^{+}18] Luke W. J. Cameron, William Roche, Peter Green, Jonathan D. R. Houghton, and Paul J. Mensink. Transatlantic movement in porbeagle sharks, Lamna nasus. Fisheries Research, 207(??):25–27, November 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618301516.

Colotelo:2013:NPB

[CRH^{+}13] Alison H. Colotelo, Graham D. Raby, Caleb T. Hasler, Tim J. Haxton, Karen E. Smokorowski, Gabriel Blouin-Demers, and Steven J. Cooke. Northern pike bycatch in an inland commercial hoop net fishery: Effects of water temperature and net tending frequency on injury, physiology, and survival. Fisheries Research, 137(??):41–49, January 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783612002676.

Cross:2015:ECC

[Cro15] Helen Cross. Elasmobranch capture by commercial smallscale fisheries in the Bijagós Archipelago, Guinea Bissau. Fisheries Research, 168(??):105–108, August 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615001083.

Cabanellas-Reboredo:2017:EHU

[CRPAMN17] Miguel Cabanellas-Reboredo, Miquel Palmer, Josep Alós, and Beatriz Morales-Nin. Estimating harvest and its uncertainty in heterogeneous recreational fisheries. Fisheries Research, 188(??):100–111, April 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616304180.

Cabacinha:2019:ATS

[CRS^{+}19] Fernanda K. R. Cabacinha, Gustavo R. Rosa, Guilherme M. Santos, Gilberto N. Salvador, Andréa Bialetzki, and Gilmar B. Santos. Analysis of three sampling methods of early life history of fish in marginal lagoons of southeastern Brazil. Fisheries Research, 219(??):Article 105327, November 2019. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783619301821.

Conrath:2019:SHU

[CRW^{+}19] Christina L. Conrath, Christopher N. Rooper, Rachel E. Wilborn, Brian A. Knoth, and Darin T. Jones. Seasonal habitat use and community structure of rockfishes in the Gulf of Alaska. Fisheries Research, 219(??):Article 105331, November 2019. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783619301869.

Conrath:2018:DDM

[CS18] Christina L. Conrath and Nicholas B. Sisson. Delayed discard mortality of the giant Pacific octopus, Enteroctopus dofleini, in the Gulf of Alaska cod pot fishery. Fisheries Research, 197(??):10–14, January 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617302710.

Coelho:2015:EHB

[CSFCA15] Rui Coelho, Miguel N. Santos, Joana Fernandez-Carvalho, and Sérgio Amorim. Effects of hook and bait in a tropical northeast Atlantic pelagic longline fishery: Part I — incidental sea turtle bycatch. Fisheries Research, 164(??):302– 311, April 2015. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783614003336.

Clapp:2013:NDE

[CSKB13] Amanda R. Clapp, Todd R. Smith, Adrienne I. Kovach, and David L. Berlinsky. No differences in egg buoyancy and anti-freeze protein production in genetically divergent subpopulations of Gulf of Maine Atlantic cod (Gadus morhua). Fisheries Research, 141(??):130–135, April 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612001610.

Cartamil:2011:AEF

[CSMEO^{+}11] Daniel Cartamil, Omar Santana-Morales, Miguel Escobedo-Olvera, Dovi Kacev, Leonardo Castillo-Geniz, Jeffrey B. Graham, Robert D. Rubin, and Oscar Sosa-Nishizaki. The artisanal elasmobranch fishery of the Pacific coast of Baja California, Mexico. Fisheries Research, 108(2–3):393–403, March 2011. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783611000397.

Crawford:2011:EBR

[CSMJB11] Charles R. Crawford, Philip Steele, Anne L. McMillen-Jackson, and Theresa M. Bert. Effectiveness of bycatchreduction devices in roller-frame trawls used in the Florida shrimp fishery. Fisheries Research, 108(2–3):248–257, March 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783610003164.

Cosgrove:2014:AFM

[CSMO14] Ronan Cosgrove, Michael Sheridan, Cóilín Minto, and Rick Officer. Application of finite mixture models to catch rate standardization better represents data distribution and fleet behavior. Fisheries Research, 153(??):83–88, May 2014. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783614000071.

Curtis:2014:OCC

[CSOV14] Judson M. Curtis, Gregory W. Stunz, R. Deborah Overath, and Robert R. Vega. Otolith chemistry can discriminate signatures of hatchery-reared and wild spotted seatrout. Fisheries Research, 153(??):31–40, May 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361300307X.

Cipriani:2018:APN

[CSP^{+}18a] Paolo Cipriani, Gian Luca Sbaraglia, Marialetizia Palomba, Lucilla Giulietti, Bruno Bellisario, Ivana Bušelić, Ivona Mladineo, Roberto Cheleschi, Giuseppe Nascetti, and Simonetta Mattiucci. Anisakis pegreffii (Nematoda: Anisakidae) in European anchovy Engraulis encrasicolus from the Mediterranean Sea: Fishing ground as a predictor of parasite distribution. Fisheries Research, 202(??):59–68, June 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783617300875.

Cipriani:2018:MEH

[CSP^{+}18b] Paolo Cipriani, Gian Luca Sbaraglia, Michela Paoletti, Lucilla Giulietti, Bruno Bellisario, Marialetizia Palomba, Ivana Bušelić, Ivona Mladineo, Giuseppe Nascetti, and Simonetta Mattiucci. The Mediterranean European hake, Merluccius merluccius: Detecting drivers influencing the Anisakis spp. larvae distribution. Fisheries Research, 202(??):79–89, June 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783617301935.

Campanella:2016:IAD

[CT16] Fabio Campanella and J. Christopher Taylor. Investigating acoustic diversity of fish aggregations in coral reef ecosystems from multifrequency fishery sonar surveys. Fisheries Research, 181(??):63–76, September 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616300959.

Currie:2019:NAA

[CTS^{+}19] Jock C. Currie, James T. Thorson, Kerry J. Sink, Lara J. Atkinson, Tracey P. Fairweather, and Henning Winker. A novel approach to assess distribution trends from fisheries survey data. Fisheries Research, 214(??):98–109, June 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619300414.

Campbell:2010:RAD

[CTSD10] Matthew D. Campbell, James Tolan, Richard Strauss, and Sandra L. Diamond. Relating angling-dependent fish impairment to immediate release mortality of red snapper (Lutjanus campechanus). Fisheries Research, 106(1):64–70, October 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783610001591.

Cope:2015:EPR

[CTWD15] Jason M. Cope, James T. Thorson, Chantell R. Wetzel, and John DeVore. Evaluating a prior on relative stock status using simplified age-structured models. Fisheries Research, 171(??):101–109, November 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614002513.

Carson:2016:SCS

[CUL^{+}16] Henry S. Carson, Michael Ulrich, Dayv Lowry, Robert E. Pacunski, and Robert Sizemore. Status of the California sea cucumber (Parastichopus californicus) and red sea urchin (Mesocentrotus franciscanus) commercial dive fisheries in the San Juan Islands, Washington State, USA. Fisheries Research, 179(??):179–190, July 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616300571.

Curtis:2018:PEL

[Cur18] John Curtis. Pike (Esox lucius) stock management in designated brown trout (Salmo trutta) fisheries: Anglers’ preferences. Fisheries Research, 207(??):37–48, November 2018.

CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783618301565.

Crone:2014:ELV

[CV14] Paul R. Crone and Juan L. Valero. Evaluation of lengthvs. age-composition data and associated selectivity assumptions used in stock assessments based on robustness of derived management quantities. Fisheries Research, 158(??):165–171, October 2014. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614000800.

Csepp:2011:SAD

[CVS11] David J. Csepp, Johanna J. Vollenweider, and Michael F. Sigler. Seasonal abundance and distribution of pelagic and demersal fishes in southeastern Alaska. Fisheries Research, 108(2–3):307–320, March 2011. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611000051.

Chavez:2017:EBR

[CW17] Stephanie Chavez and Amanda Southwood Williard. The effects of bycatch reduction devices on diamondback terrapin and blue crab catch in the North Carolina commercial crab fishery. Fisheries Research, 186 (part 1)(??):94–101, February 2017. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783616302636.

Carruthers:2012:EMC

[CWM12] Thomas R. Carruthers, Carl J. Walters, and Murdoch K. McAllister. Evaluating methods that classify fisheries stock status using only fisheries catch data. Fisheries Research, 119–120(??):66–79, May 2012. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611003894.

Campbell:2012:MBL

[CY12] Robert A. Campbell and Jock W. Young. Monitoring the behaviour of longline gears and the depth and time of fish capture in the Australian Eastern tuna and billfish fishery. Fisheries Research, 119–120(??):48–65, May 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611003729.

Candy:2014:NME

[CZW14] Steven Candy, Philippe Ziegler, and Dirk Welsford. A nonparametric model of empirical length distributions to inform stratification of fishing effort for integrated assessments. Fisheries Research, 159(??):34–44, November 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783614001556.

Duarte:2010:SLC

[dADSRG10] Luis Felipe de Almeida Duarte, Evandro Severino-Rodrigues, and Maria A. Gasalla. Slipper lobster (Crustacea, Decapoda, Scyllaridae) fisheries off the southeastern coast of Brazil: I. exploitation patterns between 23^{o}00’ and 29^{o}65’ S. Fisheries Research, 102(1–2):141–151, February 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609003063.

Deepananda:2016:SFK

[DAJMB16] K. H. M. Ashoka Deepananda, Upali S. Amarasinghe, Udith K. Jayasinghe-Mudalige, and Fikret Berkes. Stilt fisher knowledge in southern Sri Lanka as an expert system: a strategy towards co-management. Fisheries Research, 174(??):288–297, February 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615301247.

Dunstan:2010:NPF

[DAM10] A. Dunstan, O. Alanis, and J. Marshall. Nautilus pompilius fishing and population decline in the Philippines: a comparison with an unexploited Australian Nautilus population. Fisheries Research, 106(2):239–247, November 2010.

CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783610001542.

Damalas:2015:MID

[Dam15] Dimitrios Damalas. Mission impossible: Discard management plans for the EU Mediterranean fisheries under the reformed common fisheries policy. Fisheries Research, 165(??): 96–99, May 2015. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783615000259.

Deroba:2012:EHC

[DB12] Jonathan J. Deroba and James R. Bence. Evaluating harvest control rules for lake whitefish in the Great Lakes: Accounting for variable life-history traits. Fisheries Research, 121–122(??):88–103, June 2012. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361200046X.

Duncan:2018:OSA

[DBA18] Roxanne Duncan, Deirdre Brophy, and Haritz Arrizabalaga. Otolith shape analysis as a tool for stock separation of albacore tuna feeding in the Northeast Atlantic. Fisheries Research, 200(??):68–74, April 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617303533.

Duggan:2019:PIF

[DBB19] Melissa Duggan, Peter Bayliss, and Michele A. Burford. Predicting the impacts of freshwater-flow alterations on prawn (Penaeus merguiensis) catches. Fisheries Research, 215(??): 27–37, July 2019. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783619300566.

Dedeu:2019:FES

[DBG19] Arnau L. Dedeu, Jordi Boada, and Ana Gordoa. The first estimates of species compositions of Spanish marine recreational fishing reveal the activity’s inner and geographical variability. Fisheries Research, 216(??):65–73, August 2019.

CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619300906.

Douglas:2010:ERI

[DBH^{+}10] John Douglas, Paul Brown, Taylor Hunt, Mark Rogers, and Micheal Allen. Evaluating relative impacts of recreational fishing harvest and discard mortality on Murray cod (Maccullochella peelii peelii). Fisheries Research, 106(1):18–21, October 2010. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783610001360.

Drewery:2010:SSG

[DBK^{+}10] J. Drewery, D. Bova, R. J. Kynoch, A. Edridge, R. J. Fryer, and F. G. O’Neill. The selectivity of the Swedish grid and 120 mm square mesh panels in the Scottish Nephrops trawl fishery. Fisheries Research, 106(3):454–459, December 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783610002407.

Drouineau:2016:GGE

[DBLB16] H. Drouineau, C. Briand, P. Lambert, and L. Beaulaton. GEREM (Glass Eel Recruitment Estimation Model): a model to estimate glass eel recruitment at different spatial scales. Fisheries Research, 174(??):68–80, February 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783615300795.

Dick:2017:MAF

[DBMR17] E. J. Dick, Sabrina Beyer, Marc Mangel, and Stephen Ralston. A meta-analysis of fecundity in rockfishes (genus Sebastes). Fisheries Research, 187(??):73–85, March 2017. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783616303745.

Ducharme-Barth:2018:IAG

[DBSA18] Nicholas D. Ducharme-Barth, Kyle W. Shertzer, and Robert N. M. Ahrens. Indices of abundance in the Gulf of Mexico reef fish complex: a comparative approach using spatial data from vessel monitoring systems. Fisheries Research, 198(??):1–13, February 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617302965.

Driscoll:2015:LCA

[DBT15] John Driscoll, Catherine Boyd, and Peter Tyedmers. Life cycle assessment of the Maine and southwest Nova Scotia lobster industries. Fisheries Research, 172(??):385–400, December 2015. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783615300540.

duColombier:2015:UNI

[dCJG^{+}15] Sarah Bureau du Colombier, Louis Jacobs, Charline Gesset, Pierre Elie, and Patrick Lambert. Ultrasonography as a non-invasive tool for sex determination and maturation monitoring in silver eels. Fisheries Research, 164(??):50– 58, April 2015. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783614003087.

Desender:2016:STE

[DCP^{+}16] Marieke Desender, Koen Chiers, Hans Polet, Bart Verschueren, Jimmy H. Saunders, Bart Ampe, Atle Mortensen, Velmurugu Puvanendran, and Annemie Decostere. Shortterm effect of pulsed direct current on various species of adult fish and its implication in pulse trawling for brown shrimp in the North Sea. Fisheries Research, 179(??):90– 97, July 2016. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783616300480.

Dickey-Collas:2012:WDE

[DCSW^{+}12] Mark Dickey-Collas, Stylianos Somarakis, Peter R. Witthames, Cindy J. G. van Damme, Andres Uriarte, Nancy C. H. Lo, and Miguel Bernal. Where do egg production methods for estimating fish biomass go from here? Fisheries Research, 117–118(??):6–11, April 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361200015X.

Depestele:2014:STS

[DDB^{+}14] Jochen Depestele, Marieke Desender, Hugues P. Benoît, Hans Polet, and Magda Vincx. Short-term survival of discarded target fish and non-target invertebrate species in the “eurocutter” beam trawl fishery of the southern North Sea. Fisheries Research, 154(??):82–92, June 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614000344.

Dowling:2015:EHS

[DDH^{+}15a] N. A. Dowling, C. M. Dichmont, M. Haddon, D. C. Smith, A. D. M. Smith, and K. Sainsbury. Empirical harvest strategies for data-poor fisheries: a review of the literature. Fisheries Research, 171(??):141–153, November 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614003300.

Dowling:2015:GDF

[DDH^{+}15b] N. A. Dowling, C. M. Dichmont, M. Haddon, D. C. Smith, A. D. M. Smith, and K. Sainsbury. Guidelines for developing formal harvest strategies for data-poor species and fisheries. Fisheries Research, 171(??):130–140, November 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783614002811.

Danylchuk:2018:KFE

[DDK^{+}18] Andy J. Danylchuk, Sascha Clark Danylchuk, Andrew Kosiarski, Steven J. Cooke, and Bryan Huskey. Keepemwet fishing — an emerging social brand for disseminating best practices for catch-and-release in recreational fisheries. Fisheries Research, 205(??):52–56, September 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618301048.

Dixon:2015:AUD

[DDP15] Heather J. Dixon, J. Brian Dempson, and Michael Power. Assessing the use of different marine growth zones of adult Atlantic salmon scales for studying marine trophic ecology with stable isotope analysis. Fisheries Research, 164(??):112–119, April 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783614003026.

Dichmont:2016:RSA

[DDP^{+}16] Catherine M. Dichmont, Roy A. Deng, André E. Punt, Jon Brodziak, Yi-Jay Chang, Jason M. Cope, James N. Ianelli, Christopher M. Legault, Richard D. Methot, Jr., Clay E. Porch, Michael H. Prager, and Kyle W. Shertzer. A review of stock assessment packages in the United States. Fisheries Research, 183(??):447–460, November 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616302132.

Dietrich:2016:STR

[DET^{+}16] Joseph Dietrich, Kai Eder, Donald Thompson, Rebecca Buchanan, John Skalski, Geoffrey McMichael, Derek Fryer, and Frank Loge. Survival and transit of in-river and transported yearling Chinook salmon in the lower Columbia River and estuary. Fisheries Research, 183(??):435–446, November 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783616302181.

deEyto:2015:FWI

[dEWB^{+}15] E. de Eyto, J. White, P. Boylan, B. Clarke, D. Cotter, D. Doherty, P. Gargan, R. Kennedy, P. McGinnity, N. O’Maoiléidigh, and K. O’Higgins. The fecundity of wild Irish Atlantic salmon Salmo salar L. and its application for stock assessment purposes. Fisheries Research, 164(??):159– 169, April 2015. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783614003439.

Duy:2016:EAF

[DF16] Nguyen Ngoc Duy and Ola Flaaten. Efficiency analysis of fisheries using stock proxies. Fisheries Research, 181(??):102–113, September 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783616301011.

Duy:2012:OAF

[DFAN12] Nguyen Ngoc Duy, Ola Flaaten, Nguyen Thi Kim Anh, and Quach Thi Khanh Ngoc. Open-access fishing rent and efficiency — the case of gillnet vessels in nha trang, Vietnam. Fisheries Research, 127–128(??):98–108, September 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783612001373.

Devine:2011:IDM

[DH11] Jennifer A. Devine and Mikko Heino. Investigating the drivers of maturation dynamics in Barents Sea haddock (Melanogrammus aeglefinus). Fisheries Research, 110(3):441–449, August 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783611002050.

Davoren:2014:CCM

[DH14] Gail K. Davoren and Norman M. Halden. Connectivity of capelin (Mallotus villosus) between regions and spawning habitats in Newfoundland inferred from otolith chemistry. Fisheries Research, 159(??):95–104, November 2014. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783614001635.

Deely:2019:CAC

[DHC19] John Deely, Stephen Hynes, and John Curtis. Combining actual and contingent behaviour data to estimate the value of coarse fishing in Ireland. Fisheries Research, 215(??):53– 61, July 2019. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783619300736.

Dagorn:2010:DFE

[DHF10] Laurent Dagorn, Kim N. Holland, and John Filmalter. Are drifting FADs essential for testing the ecological trap hypothesis? Fisheries Research, 106(1):60–63, October 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783610001578.

Duarte:2018:ISF

[DHFSS18] José A. Duarte, Alvaro Hernández-Flores, Silvia Salas, and Juan Carlos Seijo. Is it sustainable fishing for Octopus maya Voss and Solis, 1966, during the breeding season using a baitbased fishing technique? Fisheries Research, 199(??):119– 126, March 2018. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783617303284.

Doring:2018:SSV

[DHH18] Julian Döring, Helena Hauss, and Holger Haslob. Spatial and seasonal variability in reproductive investment of Baltic sprat. Fisheries Research, 204(??):49–60, August 2018. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783618300304.

Donaldson:2011:CAB

[DHP^{+}11] Michael R. Donaldson, Scott G. Hinch, David A. Patterson, Jayme Hills, Jim O. Thomas, Steven J. Cooke, Graham D. Raby, Lisa A. Thompson, David Robichaud, Karl K. English, and Anthony P. Farrell. The consequences of angling, beach seining, and confinement on the physiology, postrelease behaviour and survival of adult sockeye salmon during upriver migration. Fisheries Research, 108(1):133–141, February 2011. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783610003231.

Dorner:2013:EAM

[DHP^{+}13] Brigitte Dorner, Kendra R. Holt, Randall M. Peterman, Chris Jordan, David P. Larsen, Anthony R. Olsen, and Omar I. Abdul-Aziz. Evaluating alternative methods for monitoring and estimating responses of salmon productivity in the North Pacific to future climatic change and other processes: a simulation study. Fisheries Research, 147(??):10–23, October 2013. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613000854.

Dennerline:2012:RBH

[DJD12] Donald E. Dennerline, Cecil A. Jennings, and Donald J. Degan. Relationships between hydroacoustic derived density and gill net catch: Implications for fish assessments. Fisheries Research, 123–124(??):78–89, July 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611003419.

Duncan:2015:GSC

[DJF^{+}15] Murray Duncan, Nicola James, Sean T. Fennessy, Rui Jorge Mutombene, and Monica Mwale. Genetic structure and consequences of stock exploitation of Chrysoblephus puniceus, a commercially important sparid in the South West Indian Ocean. Fisheries Research, 164(??):64–72, April 2015. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783614003129.

Suarez-Moo:2013:HGC

[dJSMROZP^{+}13] Pablo de Jesús Suárez-Moo, Axayácatl Rocha-Olivares, Omar Zapata-Pérez, Adriana Quiroz-Moreno, and Lorenzo Felipe Sánchez-Teyer. High genetic connectivity in the Atlantic sharpnose shark, Rhizoprionodon terraenovae, from the southeast Gulf of Mexico inferred from AFLP fingerprinting. Fisheries Research, 147(??):338–343, October 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783613001677.

Dahle:2018:GMM

[DJW^{+}18] Geir Dahle, Torild Johansen, Jon-Ivar Westgaard, Asgeir Aglen, and Kevin A. Glover. Genetic management of mixedstock fisheries “real-time”: the case of the largest remaining cod fishery operating in the Atlantic in 2007–2017. Fisheries Research, 205(??):77–85, September 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361830105X.

deKerckhove:2015:MFS

[dKMS15] Derrick T. de Kerckhove, Scott Milne, and Brian J. Shuter. Measuring fish school swimming speeds with two acoustic beams and determining the angle of the school detection. Fisheries Research, 172(??):432–439, December 2015. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783615300485.

Davie:2011:DDS

[DL11] Sarah Davie and Colm Lordan. Definition, dynamics and stability of métiers in the Irish otter trawl fleet. Fisheries Research, 111(3):145–158, November 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611002505.

Doll:2015:CAW

[DLCK15] Jason C. Doll, Thomas E. Lauer, and Sandra Clark-Kolaks. Covariates of age-0 walleye Sander vitreus fall recruitment from stocked populations in six midwestern reservoirs. Fisheries Research, 172(??):274–286, December 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615300436.

Doll:2017:YPR

[DLCK17] Jason C. Doll, Thomas E. Lauer, and Sandra Clark-Kolaks. Yield-per-recruit modeling of two piscivores in a midwestern reservoir: a Bayesian approach. Fisheries Research, 191(??):200–210, July 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617300796.

Duffy:2019:AVB

[DLCO^{+}19] Leanne M. Duffy, Cleridy E. Lennert-Cody, Robert J. Olson, Carolina V. Minte-Vera, and Shane P. Griffiths. Assessing vulnerability of bycatch species in the tuna purseseine fisheries of the eastern Pacific Ocean. Fisheries Research, 219(??):Article 105316, November 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619301638.

Doonan:2016:CNZ

[DLD^{+}16] Ian Doonan, Kath Large, Alistair Dunn, Scott Rasmussen, Craig Marsh, and Sophie Mormede. Casal2: New Zealand’s integrated population modelling tool. Fisheries Research, 183(??):498–505, November 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616301412.

Dick:2011:DBS

[DM11] E. J. Dick and Alec D. MacCall. Depletion-based stock reduction analysis: a catch-based method for determining sustainable yields for data-poor fish stocks. Fisheries Research, 110(2):331–341, July 2011. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611001962.

Deroba:2016:CTW

[DM16] Jonathan J. Deroba and Timothy J. Miller. Correct in theory but wrong in practice: Bias caused by using a lognormal distribution to penalize annual recruitments in fish stock assessment models. Fisheries Research, 176(??):86–93, April 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783615301545.

Deng:2012:DEN

[DMC^{+}12] Z. Daniel Deng, Jayson J. Martinez, Alison H. Colotelo, Tylor K. Abel, Andrea P. LeBarge, Richard S. Brown, Brett D. Pflugrath, Robert P. Mueller, Thomas J. Carlson, Adam G. Seaburg, Robert L. Johnson, and Martin L.

Ahmann. Development of external and neutrally buoyant acoustic transmitters for juvenile salmon turbine passage evaluation. Fisheries Research, 113(1):94–105, January 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783611003092.

Dance:2016:DTP

[DMFR16] Michael A. Dance, David L. Moulton, Nathan B. Furey, and Jay R. Rooker. Does transmitter placement or species affect detection efficiency of tagged animals in biotelemetry research? Fisheries Research, 183(??):80–85, November 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783616301461.

Gil:2018:MFC

[dMGPR^{+}18] Maria del Mar Gil, Miquel Palmer, Rosario Rosselló, Josep Alós, Miguel Cabanellas-Reboredo, Antoni Maria Grau, and Beatriz Morales-Nin. Mixed-fish commercial categories as a strategy for reducing discards from the small-scale fleet. Fisheries Research, 205(??):105–114, September 2018. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783618301115.

DuFour:2017:STH

[DMK^{+}17] Mark R. DuFour, Christine M. Mayer, Patrick M. Kocovsky, Song S. Qian, Dave M. Warner, Richard T. Kraus, and Christopher S. Vandergoot. Sparse targets in hydroacoustic surveys: Balancing quantity and quality of in situ target strength data. Fisheries Research, 188(??):173–182, April 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783616304398.

Divine:2019:NEW

[DMK^{+}19] Lauren M. Divine, Franz J. Mueter, Gordon H. Kruse, Bodil A. Bluhm, Stephen C. Jewett, and Katrin Iken. New estimates of weight-at-size, maturity-at-size, fecundity, and biomass of snow crab, Chionoecetes opilio, in the Arctic Ocean off Alaska. Fisheries Research, 218(??):246–258, October 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783619301183.

Davie:2015:DVP

[DMOL15] Sarah Davie, Cóilín Minto, Rick Officer, and Colm Lordan. Defining value per unit effort in mixed métier fisheries. Fisheries Research, 165(??):1–10, May 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614003634.

DuFour:2018:IFB

[DMQ^{+}18] Mark R. DuFour, Christine M. Mayer, Song S. Qian, Christopher S. Vandergoot, Richard T. Kraus, Patrick M. Kocovsky, and David M. Warner. Inferred fish behavior its implications for hydroacoustic surveys in nearshore habitats. Fisheries Research, 199(??):63–75, March 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617303260.

Duzbastilar:2010:MFE

[DÖA^{+}10] F. Ozan Düzbastılar, Hüseyin Özbilgin, Celalettin Aydın, Gülnur Metin, Ali Ulaş, Altan Lök, and Cengiz Metin. Mortalities of fish escaping from square and diamond mesh codends in the Aegean Sea. Fisheries Research, 106(3):386–392, December 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783610002171.

Diaz:2014:LCM

[DOM14] Marina V. Diaz, M. Pilar Olivar, and Gustavo J. Macchi. Larval condition of Merluccius hubbsi (Marini, 1933) in the northern Patagonian spawning ground. Fisheries Research, 160(??):60–68, December 2014. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613002877.

Dunford:2015:IEA

[DOO15] Adam J. Dunford, Richard L. O’Driscoll, and Johannes Oeffner. Improvements in estimating an acoustic target strength–length relationship for hoki (Macruronus novaezeladiae). Fisheries Research, 162(??):12–19, February 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783614002756.

Dedual:2018:DAL

[DP18] Michel Dedual and Michael Pickford. The demand for angling licenses: a case study from Taupo, New Zealand. Fisheries Research, 204(??):289–296, August 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618300730.

Dauphin:2010:URC

[DPAB10] Guillaume Dauphin, Etienne Prévost, Colin E. Adams, and Patrick Boylan. Using redd counts to estimate salmonids spawner abundances: a Bayesian modelling approach. Fisheries Research, 106(1):32–40, October 2010. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610001530.

DElia:2014:ABP

[DPB^{+}14] M. D’Elia, B. Patti, A. Bonanno, I. Fontana, G. Giacalone, G. Basilone, and P. G. Fernandes. Analysis of backscatter properties and application of classification procedures for the identification of small pelagic fish species in the Central Mediterranean. Fisheries Research, 149(??):33–42, January 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783613001872.

doPrado:2012:DHG

[dPHS^{+}12] Fernanda Dotti do Prado, Diogo Teruo Hashimoto, José Augusto Senhorini, Fausto Foresti, and Fabio Porto-Foresti. Detection of hybrids and genetic introgression in wild stocks of two catfish species (Siluriformes: Pimelodidae): the impact of hatcheries in Brazil. Fisheries Research, 125–126(??):300–305, August 2012. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612001221.

Dowling:2016:AMT

[DPL^{+}16] Natalie A. Dowling, André E. Punt, L. Richard Little, Catherine M. Dichmont, David C. Smith, Malcolm Haddon, Miriana Sporcic, Elizabeth A. Fulton, and Rebecca J. Gorton. Assessing a multilevel tier system: the role and implications of data quality and availability. Fisheries Research, 183(??):588–593, November 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616301357.

Diaz:2011:NCA

[DPO^{+}11] Marina V. Diaz, Marcelo Pájaro, M. Pilar Olivar, Patricia Martos, and Gustavo J. Macchi. Nutritional condition of Argentine anchovy Engraulis anchoita larvae in connection with nursery ground properties. Fisheries Research, 109(2–3):330–341, May 2011. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611000919.

Dominguez-Petit:2010:NBP

[DPSR10] R. Domínguez-Petit and F. Saborido-Rey. New bioenergetic perspective of European hake (Merluccius merluccius L.) reproductive ecology. Fisheries Research, 104(1–3):83–88, May 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783609002458.

Dominguez-Petit:2010:CPC

[DPSRM10] R. Domínguez-Petit, F. Saborido-Rey, and I. Medina. Changes of proximate composition, energy storage and condition of European hake (Merluccius merluccius, L. 1758) through the spawning season. Fisheries Research, 104(1–3): 73–82, May 2010. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783609001416.

DuFour:2019:ECL

[DQMV19] Mark R. DuFour, Song. S. Qian, Christine M. Mayer, and Christopher S. Vandergoot. Evaluating catchability in a large-scale gillnet survey using hydroacoustics: Making the case for coupled surveys. Fisheries Research, 211(??):309– 318, March 2019. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783618303217.

DAvignon:2013:OEF

[DR13] Geneviève D’Avignon and George A. Rose. Otolith elemental fingerprints distinguish Atlantic cod spawning areas in Newfoundland and Labrador. Fisheries Research, 147(??):1– 9, October 2013. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783613001045.

Digre:2017:BLS

[DRE^{+}17] Hanne Digre, Carolyn Rosten, Ulf Erikson, John R. Mathiassen, and Ida G. Aursand. The on-board live storage of Atlantic cod (Gadus morhua) and haddock (Melanogrammus aeglefinus) caught by trawl: Fish behaviour, stress and fillet quality. Fisheries Research, 189(??):42–54, May 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783617300048.

Da-Rocha:2012:PVO

[DRGA12] José-María Da-Rocha, María-José Gutiérrez, and Luis T. Antelo. Pulse vs. optimal stationary fishing: the northern stock of hake. Fisheries Research, 121–122(??):51–62, June 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783612000409.

Da-Rocha:2011:MFP

[DRGGVL11] José M. Da-Rocha, M. Dolores Garza-Gil, and Manuel M. Varela-Lafuente. A model of fishing periods applied to the European sardine fishery. Fisheries Research, 109(1):16–24, April 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783611000129.

Deroba:2013:PSA

[DS13] Jonathan J. Deroba and Amy M. Schueller. Performance of stock assessments with misspecified age- and time-varying natural mortality. Fisheries Research, 146(??):27–40, September 2013. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783613000830.

Duong:2018:RVG

[DS18] Thuy-Yen Duong and Kim T. Scribner. Regional variation in genetic diversity between wild and cultured populations of bighead catfish (Clarias macrocephalus) in the Mekong Delta. Fisheries Research, 207(??):118–125, November 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783618301826.

Dunn:2010:RAC

[DSB^{+}10] D. C. Dunn, K. Stewart, R. H. Bjorkland, M. Haughton, S. Singh-Renton, R. Lewison, L. Thorne, and P. N. Halpin. A regional analysis of coastal and domestic fishing effort in the wider Caribbean. Fisheries Research, 102(1–2):60–68, February 2010. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783609002744.

Dortel:2015:IBM

[DSB^{+}15] E. Dortel, F. Sardenne, N. Bousquet, E. Rivot, J. Million, G. Le Croizier, and E. Chassot. An integrated Bayesian modeling approach for the growth of Indian Ocean yellowfin tuna. Fisheries Research, 163(??):69–84, March 2015. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783614002203.

DOnghia:2019:CWC

[DSC19] Gianfranco D’Onghia, Letizia Sion, and Francesca Capezzuto. Cold-water coral habitats benefit adjacent fisheries along the Apulian margin (central Mediterranean). Fisheries Research, 213(??):172–179, May 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619300219.

Freitas:2017:MCR

[dSFdSSS17] Adrianne dos Santos Freitas, Raimundo da Silva, Iracilda Sampaio, and Horacio Schneider. The mitochondrial control region reveals genetic structure in southern kingcroaker populations on the coast of the Southwestern Atlantic. Fisheries Research, 191(??):87–94, July 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617300759.

Delgado:2017:PPP

[DSG^{+}17] Marina Delgado, Luis Silva, Santiago Gómez, Esther Masferrer, Miguel Cojan, and Miguel B. Gaspar. Population and production parameters of the wedge clam Donax trunculus (Linnaeus, 1758) in intertidal areas on the southwest Spanish coast: Considerations in relation to protected areas. Fisheries Research, 193(??):232–241, September 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783617301170.

Delgado:2013:ARS

[DSJ13] Marina Delgado, Luis Silva, and Ana Juárez. Aspects of reproduction of striped venus Chamelea gallina in the Gulf of Cádiz (SW Spain): Implications for fishery management. Fisheries Research, 146(??):86–95, September 2013. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783613001033.

Aires-da-Silva:2015:IGE

[dSMSF15] Alexandre M. Aires da Silva, Mark N. Maunder, Kurt M. Schaefer, and Daniel W. Fuller. Improved growth estimates from integrated analysis of direct aging and tag–recapture data: an illustration with bigeye tuna (Thunnus obesus) of the eastern Pacific Ocean with implications for management. Fisheries Research, 163(??):119–126, March 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361400126X.

Santos:2017:DSO

[dSSdAdAA17] Rosa da Silva Santos, Márcia Cristina Costa de Azevedo, Cristiano Queiroz de Albuquerque, and Francisco Gerson Araújo. Different sagitta otolith morphotypes for the whitemouth croaker Micropogonias furnieri in the Southwestern Atlantic coast. Fisheries Research, 195(??):222–229, November 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783617302102.

Duermit:2017:FAC

[DSW^{+}17] Elizabeth Duermit, Virginia Shervette, J. David Whitaker, Peter R. Kingsley-Smith, and Dara Wilber. A field assessment of claw removal impacts on the movement and survival of stone crabs Menippe spp. Fisheries Research, 193(??):43–50, September 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617300863.

Doddema:2018:FRP

[DSWB18] Mandy Doddema, Gert Spaargaren, Budy Wiryawan, and Simon R. Bush. Fisher responses to private monitoring interventions in an Indonesian tuna handline fishery. Fisheries Research, 208(??):49–57, December 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618301991.

Dwyer:2016:AGG

[DTC16] K. S. Dwyer, M. A. Treble, and S. E. Campana. Age and growth of Greenland halibut (Reinhardtius hippoglossoides) in the Northwest Atlantic: a changing perception based on bomb radiocarbon analyses. Fisheries Research, 179(??):342–350, July 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616300169.

Daures:2013:MFC

[DTG13] Fabienne Daurès, Verena M. Trenkel, and Olivier Guyader. Modelling the fishing costs of French commercial vessels in the Bay of Biscay. Fisheries Research, 146(??):74–85, September 2013. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783613001069.

Digre:2016:PMS

[DTSG^{+}16] Hanne Digre, Guro M. Tveit, Torfinn Solvang-Garten, Aleksander Eilertsen, and Ida G. Aursand. Pumping of mackerel (Scomber scombrus) onboard purse seiners, the effect on mortality, catch damage and fillet quality. Fisheries Research, 176(??):65–75, April 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615301624.

Duong:2019:GDS

[DUC^{+}19] Thuy-Yen Duong, Sophorn Uy, Phen Chheng, Nam So, Thanh-Hien Thi Tran, Ngoc-Tran Thi Nguyen, Robert Pomeroy, and Hillary Egna. Genetic diversity and structure of striped snakehead (Channa striata) in the Lower Mekong Basin: Implications for aquaculture and fisheries management. Fisheries Research, 218(??):166–173, October 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619301304.

Disspain:2018:LTA

[DUD^{+}18] Morgan C. F. Disspain, Sean Ulm, Neale Draper, Jeffery Newchurch, Stewart Fallon, and Bronwyn M. Gillanders. Long-term archaeological and historical archives for mulloway, Argyrosomus japonicus, populations in eastern South Australia. Fisheries Research, 205(??):1–10, September 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783618300833.

Damalas:2011:CCD

[DV11] Dimitrios Damalas and Vassiliki Vassilopoulou. Chondrichthyan by-catch and discards in the demersal trawl fishery of the central Aegean Sea (Eastern Mediterranean). Fisheries Research, 108(1):142–152, February 2011. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783610003371.

delValle:2018:ECC

[dVA18] I. del Valle and K. Astorkiza. Exploring cross correlation among diversity indices. Fisheries Research, 204(??):

103–115, August 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783618300365.

deVasconcellos:2015:GPS

[dVLB^{+}15] Anderson Vilasboa de Vasconcellos, Daíza Lima, François Bonhomme, Marcelo Vianna, and Antonio Mateo Solé-Cava. Genetic population structure of the commercially most important demersal fish in the Southwest Atlantic: the whitemouth croaker (Micropogonias furnieri). Fisheries Research, 167(??):333–337, July 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615000983.

Damiano:2019:PDE

[DW19] Matthew D. Damiano and Michael J. Wilberg. Population dynamics of eastern oysters in the Choptank River Complex, Maryland during 1989–2015. Fisheries Research, 212(??):196–207, April 2019. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618303758.

Dawe:2010:CEM

[DWH10] Earl G. Dawe, Stephen J. Walsh, and Elaine M. Hynick. Capture efficiency of a multi-species survey trawl for snow crab (Chionoecetes opilio) in the Newfoundland region. Fisheries Research, 101(1–2):70–79, January 5, 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609002525.

deZarate:2016:EIG

[dZB16] V. Ortiz de Zárate and E. A. Babcock. Estimating individual growth variability in albacore (Thunnus alalunga) from the North Atlantic stock: Aging for assessment purposes. Fisheries Research, 180(??):54–66, August 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615300394.

Egerton:2018:HEF

[EAAA^{+}18] Jack P. Egerton, Mohsin Al-Ansi, Mohamed Abdallah, Mark Walton, Jamie Hayes, John Turner, Brad Erisman, Ibrahim Al-Maslamani, Mohammed Mohannadi, and Lewis Le Vay. Hydroacoustics to examine fish association with shallow offshore habitats in the Arabian Gulf. Fisheries Research, 199(??):127–136, March 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617303375.

Erisman:2014:IGS

[EAM^{+}14] Brad E. Erisman, Ashley M. Apel, Alec D. MacCall, Martha J. Román, and Rod Fujita. The influence of gear selectivity and spawning behavior on a data-poor assessment of a spawning aggregation fishery. Fisheries Research, 159(??):75–87, November 2014. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614001829.

Edinger:2015:RPR

[EB15] Tonya Edinger and Jungho Baek. The role of property rights in bycatch reduction: Evidence from the British Columbia groundfish fishery. Fisheries Research, 168(??):100–104, August 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783615001241.

Eddy:2016:RVM

[EBB16] Corey Eddy, Richard Brill, and Diego Bernal. Rates of atvessel mortality and post-release survival of pelagic sharks captured with tuna purse seines around drifting fish aggregating devices (FADs) in the equatorial eastern Pacific Ocean. Fisheries Research, 174(??):109–117, February 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783615300849.

Echave:2017:FRT

[Ech17] Katy B. Echave. First results of the tagging of shortspine thornyhead, Sebastolobus alascanus, in Alaska. Fisheries Research, 195(??):46–51, November 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617301704.

Erikson:2016:EIS

[EDG16] Ulf Erikson, Hanne Digre, and Leif Grimsmo. Electrical immobilisation of saithe (Pollachius virens): Effects of prestunning stress, applied voltage, and stunner configuration. Fisheries Research, 179(??):148–155, July 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616300479.

Eckmann:2019:RBP

[EE19] Reiner Eckmann and Brigitte Engesser. Reconstructing the build-up of a pelagic stickleback (Gasterosteus aculeatus) population using hydroacoustics. Fisheries Research, 210(??):189–192, February 2019. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618302236.

Espinoza:2011:TNA

[EFW^{+}11] Mario Espinoza, Thomas J. Farrugia, Dale M. Webber, Frank Smith, and Christopher G. Lowe. Testing a new acoustic telemetry technique to quantify long-term, fine-scale movements of aquatic animals. Fisheries Research, 108(2–3):364–371, March 2011. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611000300.

Ellis:2019:ATA

[EFWC^{+}19] Robert D. Ellis, Kerry E. Flaherty-Walia, Angela B. Collins, Joel W. Bickford, Ross Boucek, Sarah L. Walters Burnsed, and Susan K. Lowerre-Barbieri. Acoustic telemetry array evolution: From species- and project-specific designs to large-scale, multispecies, cooperative networks. Fisheries Research, 209(??):186–195, January 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618302509.

Eggert:2015:TRW

[EGK15] Håkan Eggert, Mads Greaker, and Asmerom Kidane. Trade and resources: Welfare effects of the Lake Victoria fisheries boom. Fisheries Research, 167(??):156–163, July 2015. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783615000600.

Ellis:2018:RIF

[EHB18] Timothy A. Ellis, Joseph E. Hightower, and Jeffrey A. Buckel. Relative importance of fishing and natural mortality for spotted seatrout (Cynoscion nebulosus) estimated from a tag-return model and corroborated with survey data. Fisheries Research, 199(??):81–93, March 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617303120.

Emmrich:2021:HEF

[EHB^{+}21] M. Emmrich, I. P. Helland, S. Busch, S. Schiller, and T. Mehner. Hydroacoustic estimates of fish densities in comparison with stratified pelagic trawl sampling in two deep, coregonid-dominated lakes. Fisheries Research, 105(3):178– 186, ???? 2021. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783610001049.

Emmerson:2018:CCC

[EHBK18] J. A. Emmerson, J. A. Haig, I. S. M. Bloor, and M. J. Kaiser. The complexities and challenges of conserving common whelk (Buccinum undatum L.) fishery resources: Spatio-temporal study of variable population demographics within an environmental context. Fisheries Research, 204(??):125–136, August 2018. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618300511.

Emery:2016:HCM

[EHGG16] Timothy J. Emery, Klaas Hartmann, Bridget S. Green, and Caleb Gardner. Handled with care: Minimal impacts of appendage damage on the growth and productivity of the southern rock lobster (Jasus edwardsii). Fisheries Research, 175(??):75–86, March 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615301387.

Edwards:2011:BEF

[EHH^{+}11] C. T. T. Edwards, R. Hillary, E. Hoshino, J. Pearce, and D. J. Agnew. Bioeconomic evaluation of fisheries enforcement effort using a multifleet simulation model. Fisheries Research, 107(1–3):253–260, January 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610002924.

Eveson:2015:SFT

[EHH^{+}15] J. Paige Eveson, Alistair J. Hobday, Jason R. Hartog, Claire M. Spillman, and Kirsten M. Rough. Seasonal forecasting of tuna habitat in the Great Australian Bight. Fisheries Research, 170(??):39–49, October 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615001496.

Eide:2012:BMS

[Eid12] Arne Eide. A bioeconomic MPA study based on cellular automata population growth and distribution. Fisheries Research, 113(1):118–132, January 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611003122.

Eide:2016:MPI

[Eid16] Arne Eide. Management performance indicators based on year-class histories. Fisheries Research, 174(??):280–287, February 2016. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783615301223.

Evans:2018:FFA

[EL18] Nathan T. Evans and Gary A. Lamberti. Freshwater fisheries assessment using environmental DNA: a primer on the method, its potential, and shortcomings as a conservation tool. Fisheries Research, 197(??):60–66, January 2018. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783617302618.

Enever:2017:MFE

[ELRH17] R. Enever, S. Lewin, A. Reese, and T. Hooper. Mapping fishing effort: Combining fishermen’s knowledge with satellite monitoring data in English waters. Fisheries Research, 189(??):67–76, May 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783617300188.

Ehrlich:2013:VDH

[EMMM13] Martín Ehrlich, Gustavo Macchi, Adrián Madirolas, and Laura Machinandiarena. Vertical distribution of hake Merluccius hubbsi in spawning aggregations in North Patagonian waters of the Southwest Atlantic. Fisheries Research, 138(??):89–98, February 2013. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612002135.

Erisman:2015:CFA

[EMOLS^{+}15] Brad Erisman, Ismael Mascareñas-Osorio, Catalina López-Sagástegui, Marcia Moreno-Báez, Victoria Jiménez-Esquivel, and Octavio Aburto-Oropeza. A comparison of fishing activities between two coastal communities within a biosphere reserve in the Upper Gulf of California. Fisheries Research, 164(??):254–265, April 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783614003671.

Erisman:2010:SAL

[EMP^{+}10] Brad Erisman, Ismael Mascarenas, Gustavo Paredes, Yvonne Sadovy de Mitcheson, Octavio Aburto-Oropeza, and Philip Hastings. Seasonal, annual, and long-term trends in commercial fisheries for aggregating reef fishes in the Gulf of California, Mexico. Fisheries Research, 106(3):279–288, December 2010. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783610001840.

Eveson:2015:EGT

[EMSC15] J. Paige Eveson, Julien Million, Fany Sardenne, and Gaël Le Croizier. Estimating growth of tropical tunas in the Indian Ocean using tag-recapture data and otolith-based age estimates. Fisheries Research, 163(??):58–68, March 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578361400188X.

Engwerda:2018:SUS

[Eng18] Jacob Engwerda. Stabilization of an uncertain simple fishery management game. Fisheries Research, 203(??):63–73, July 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783617302011.

Erlandsson:2017:SSB

[EÖFPH17] Johan Erlandsson, Örjan Östman, Ann-Britt Florin, and Zeynep Pekcan-Hekim. Spatial structure of body size of European flounder (Platichthys flesus L.) in the Baltic Sea. Fisheries Research, 189(??):1–9, May 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617300012.

Ellis:2012:GGM

[EP12] Robert D. Ellis and Joseph E. Powers. Gag grouper, marine reserves, and density-dependent sex change in the Gulf of Mexico. Fisheries Research, 115–116(??):89–98, March 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783611003535.

Edyvane:2017:TDF

[EP17] Karen S. Edyvane and Shane S. Penny. Trends in derelict fishing nets and fishing activity in northern Australia: Implications for trans-boundary fisheries management in the shared Arafura and Timor Seas. Fisheries Research, 188(??):23–37, April 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616304064.

Eayrs:2017:AAC

[EPCB17] Stephen Eayrs, Michael Pol, Shelly Tallack Caporossi, and Carl Bouchard. Avoidance of Atlantic cod (Gadus morhua) with a topless trawl in the New England groundfish fishery. Fisheries Research, 185(??):145–152, January 2017. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783616303071.

Eighani:2018:EBT

[EPHF18] Morteza Eighani, Seyed Yousef Paighambari, Bent Herrmann, and Jordan Feekings. Effect of bait type and size on catch efficiency of narrow-barred Spanish mackerel (Scomberomorus commerson) in the Persian Gulf handline fisheries. Fisheries Research, 199(??):32–35, March 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783617303314.

Escati-Penaloza:2010:ALG

[EPO10] Gabriela Escati-Peñaloza, Ana M. Parma, and J. M. (Lobo) Orensanz. Analysis of longitudinal growth increment data using mixed-effects models: Individual and spatial variability in a clam. Fisheries Research, 105(2):91–101, July 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783610000718.

Eighani:2019:INF

[EPTG19] Morteza Eighani, Seyed Yousef Paighambari, Marc Taquet, and Jean-Claude Gaertner. Introducing nearshore fish aggregation devices (FAD) to artisanal Persian Gulf fisheries: a preliminary study. Fisheries Research, 212(??):35–39, April 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783618303497.

Elliott:2018:SPD

[ER18] Lucas Elliott and Michael A. Russello. SNP panels for differentiating advanced-generation hybrid classes in recently diverged stocks: a sensitivity analysis to inform monitoring of sockeye salmon re-stocking programs. Fisheries Research, 208(??):339–345, December 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618302364.

Enever:2010:DMI

[ERCG10] R. Enever, A. S. Revill, R. Caslake, and A. Grant. Discard mitigation increases skate survival in the Bristol Channel. Fisheries Research, 102(1–2):9–15, February 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783609002598.

Eigaard:2011:IFE

[ERG^{+}11] Ole Ritzau Eigaard, Dominic Rihan, Norman Graham, Antonello Sala, and Kristian Zachariassen. Improving fishing effort descriptors: Modelling engine power and gear-size relations of five European trawl fleets. Fisheries Research, 110(1):39–46, June 2011. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611001196.

Ellis:2015:EHF

[ERL15] Jessica E. Ellis, Sherrylynn Rowe, and Heike K. Lotze. Expansion of hagfish fisheries in Atlantic Canada and worldwide. Fisheries Research, 161(??):24–33, January 2015. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783614001994.

Eggenberger:2019:CTS

[ESF^{+}19] C. W. Eggenberger, R. O. Santos, T. A. Frankovich, W. R. James, C. J. Madden, J. A. Nelson, and J. S. Rehage. Coupling telemetry and stable isotope techniques to unravel movement: Snook habitat use across variable nutrient environments. Fisheries Research, 218(??):35–47, October 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619301055.

Etherton:2015:EPP

[Eth15] Mark Etherton. European plaice (Pleuronectes platessa) and sole (Solea solea) indirect age validation using otoliths from mark-recapture experiments from the North Sea. Fisheries Research, 170(??):76–81, October 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615001502.

Everson:2013:AEC

[ETMK13] Inigo Everson, Anthony Taabu-Munyaho, and Robert Kayanda. Acoustic estimates of commercial fish species in Lake Victoria: Moving towards ecosystem-based fisheries management. Fisheries Research, 139(??):65–75, March 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783612002962.

Evans:2010:SGS

[EvHRF10] R. D. Evans, L. van Herwerden, G. R. Russ, and A. J. Frisch. Strong genetic but not spatial subdivision of two reef fish species targeted by fishers on the Great Barrier Reef. Fisheries Research, 102(1–2):16–25, February 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609002665.

Elvarsson:2018:PLD

[EWB^{+}18] Bjarki Tór Elvarsson, Pamela J. Woods, Höskuldur Björnsson, Jamie Lentin, and Gudmundur Thordarson. Pushing the limits of a data challenged stock: a size- and age-structured assessment of ling (Molva molva) in Icelandic waters using gadget. Fisheries Research, 207(??):95–109, November 2018. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783618301759.

Fraser:2019:TNE

[FA19] Shaun Fraser and Chevonne H. Angus. Trial of a new escape panel concept to reduce cod catches in a mixed demersal fishery. Fisheries Research, 213(??):212–218, May 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619300256.

Fernandez:2015:CRC

[FAAM15] Nuria Fernández, Ignacio Alborés, and Sara Aceña-Matarranz. Characterization of the reproductive cycle and physiological condition of Patella vulgata in the NW of the Iberian Peninsula: Relevant information for a sustainable exploitation. Fisheries Research, 164(??):293–301, April 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783614003683.

Fernandez:2016:SVR

[FAAM16] Nuria Fernández, Ignacio Alborés, and Sara Aceña-Matarranz. Spatial variability of the reproductive cycle and physiological condition of Patella spp. (Mollusca Gastropoda) in the NW of the Iberian Peninsula: Implications for exploitation. Fisheries Research, 179(??):76–85, July 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783616300406.

Frid:2019:CDS

[FB19] Ori Frid and Jonathan Belmaker. Catch dynamics of set net fisheries in Israel. Fisheries Research, 213(??):1–11, May 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783618303710.

Fullbrook:2017:CEP

[FBB^{+}17] Liam R. Fullbrook, Dylan Barrett, Patricia Browne, Susanne P. Eriksson, Conor T. Graham, Lee Hancox, Valentina Lauria, Anette Ungfors, Mark P. Johnson, and Anne Marie Power. Comprehensive evaluation of passive tags show no adverse effects in an economically important crustacean. Fisheries Research, 187(??):209–217, March 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616304106.

Flem:2018:FFO

[FBF^{+}18] Belinda Flem, Tonje Fagertun Benden, Tor Erik Finne, Vidar Moen, Thor Mikkel Nordahl, Ketil Skår, Øystein Nordgulen, and Arne Solli. The fish farm of origin is assigned by the element profile of Atlantic salmon (Salmo salar L.) scales in a simulated escape event. Fisheries Research, 206(??):1–13, October 2018. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618301309.

Fassler:2016:ADC

[FBGB16] Sascha M. M. Fässler, Thomas Brunel, Sven Gastauer, and Dirk Burggraaf. Acoustic data collected on pelagic fishing vessels throughout an annual cycle: Operational framework, interpretation of observations, and future perspectives. Fisheries Research, 178(??):39–46, June 2016. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578361530117X.

Farrell:2014:DCH

[FBHH14] Edward R. Farrell, André M. Boustany, Patrick N. Halpin, and Donald L. Hammond. Dolphinfish (Coryphaena hippurus) distribution in relation to biophysical ocean conditions in the northwest Atlantic. Fisheries Research, 151(??):177– 190, March 2014. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783613002920.

Feekings:2016:ITM

[FBKE16] Jordan Feekings, Casper Willestofte Berg, Ludvig Ahm Krag, and Ole Ritzau Eigaard. Influence of twin and multi-rig trawl systems on CPUE in the Danish Norway lobster (Nephrops norvegicus) fishery. Fisheries Research, 175(??):51–56, March 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615301405.

Fisch:2019:CAS

[FBM^{+}19] Nicholas C. Fisch, James R. Bence, Jared T. Myers, Eric K. Berglund, and Daniel L. Yule. A comparison of age- and sizestructured assessment models applied to a stock of cisco in Thunder Bay, Ontario. Fisheries Research, 209(??):86–100, January 2019. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783618302492.

Field:2012:CSD

[FBY^{+}12] Iain C. Field, Rik C. Buckworth, Guo-Jing Yang, Mark G. Meekan, Grant Johnson, John D. Stevens, Richard D. Pillans, Clive R. McMahon, and Corey J. A. Bradshaw. Changes in size distributions of commercially exploited sharks over 25 years in northern Australia using a Bayesian approach. Fisheries Research, 125–126(??):262–271, August 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783612001245.

Faunce:2015:COS

[FCBS15] Craig H. Faunce, Jennifer Cahalan, Julie Bonney, and Robert Swanson. Can observer sampling validate industry catch reports from trawl fisheries? Fisheries Research, 172(??):34–43, December 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615001873.

Fernandez-Carvalho:2015:EHB

[FCCSA15] Joana Fernandez-Carvalho, Rui Coelho, Miguel N. Santos, and Sérgio Amorim. Effects of hook and bait in a tropical northeast Atlantic pelagic longline fishery: Part II — target, bycatch and discard fishes. Fisheries Research, 164(??):312– 321, April 2015. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783614003348.

Fabre:2017:FDE

[FCIB17] Nidia N. Fabré, Leandro Castello, Victoria J. Isaac, and Vandick S. Batista. Fishing and drought effects on fish assemblages of the central Amazon Basin. Fisheries Research, 188(??):157–165, April 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616304349.

Feng:2016:ECS

[FCL16] Yongjiu Feng, Xinjun Chen, and Yan Liu. The effects of changing spatial scales on spatial patterns of CPUE for Ommastrephes bartramii in the northwest Pacific Ocean. Fisheries Research, 183(??):1–12, November 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616301436.

Feiner:2019:QSW

[FCLH19] Zachary S. Feiner, David P. Coulter, Matthew D. Linn, and Tomas O. Höök. A question of scale: Weak evidence for broad regional synchrony in fish year-class strength within or among species in inland lakes. Fisheries Research, 214(??):45–55, June 2019. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619300268.

Feki:2016:SVH

[FCNB16] M. Feki, M. Châari, L. Neifar, and L. Boudaya. Spatial variability of helminth parasites to recognize the discrimination of juvenile and young adult areas of horse mackerel, Trachurus trachurus (Linnaeus, 1758) off the coast of Tunisia. Fisheries Research, 183(??):318–325, November 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616301710.

Fernandes:2018:CBH

[FCOM18] Telmo J. R. Fernandes, Joana Costa, M. Beatriz P. P. Oliveira, and Isabel Mafra. COI barcode-HRM as a novel approach for the discrimination of hake species. Fisheries Research, 197(??):50–59, January 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361730262X.

Franca:2019:FSP

[FdMCB^{+}19] Nielson Felix Caetano França, Alex Barbosa de Moraes, Abner Carvalho-Batista, Marília Carla Ramos Barreto de Melo, Laura López-Greco, Fernando Luis Mantelatto, and Fúlvio Aurélio de Morais Freire. Farfantepenaeus subtilis (Pérez–Farfante, 1967) and F. brasiliensis (Latreille, 1817) (Decapoda, Penaeidae): Ontogenetic comparison using the combined analysis of secondary sexual characters and molecular markers. Fisheries Research, 216(??):89–95, August 2019. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578361930089X.

Fu:2016:DPL

[FDMH16] Dan Fu, Alistair Dunn, Keith P. Michael, and Julie Hills. The development and performance of a length-based stock assessment of Foveaux Strait oysters (Ostrea chilensis, OYU 5) in southern New Zealand, and application to management. Fisheries Research, 183(??):506–517, November 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783616301370.

Floch:2011:IFR

[FDN^{+}11] Pascal Le Floc’h, Fabienne Daurès, Myriam Nourry, Olivier Thébaud, Muriel Travers, and Sylvie Van Iseghem. The influence of fiscal regulations on investment in marine fisheries: a French case study. Fisheries Research, 109(2–3):257–264, May 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578361100083X.

Francis:2016:EFG

[FdSM^{+}16] R. I. C. Chris Francis, Alexandre M. Aires da Silva, Mark N. Maunder, Kurt M. Schaefer, and Daniel W. Fuller. Estimating fish growth for stock assessments using both age–length and tagging-increment data. Fisheries Research, 180(??):113–118, August 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615001915.

Feltrim:2010:ICG

[FE10] Marcelo Feltrim and Billy Ernst. Inter-cohort growth variability and its implication for fishery management of the common sardine (Strangomera bentincki) stock off the coast of south-central Chile. Fisheries Research, 106(3):368–377, December 2010. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783610002158.

Furlong-Estrada:2017:UPS

[FEGMTÁ17] Emmanuel Furlong-Estrada, Felipe Galván-Magaña, and Javier Tovar-Ávila. Use of the productivity and susceptibility analysis and a rapid management-risk assessment to evaluate the vulnerability of sharks caught off the west coast of Baja California Sur, Mexico. Fisheries Research, 194(??):197–208, October 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617301637.

Fey:2018:EPI

[Fey18] Dariusz P. Fey. The effect of preserving ichthyoplankton samples in alcohol on the accuracy of data obtained from otolith microstructure examinations. Fisheries Research, 206(??):198–201, October 2018. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618301528.

Florin:2010:SSF

[FF10] Ann-Britt Florin and Fredrik Franzén. Spawning site fidelity in Baltic Sea turbot (Psetta maxima). Fisheries Research, 102(1–2):207–213, February 2010. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578360900318X.

Froeschke:2011:STP

[FF11] John T. Froeschke and Bridgette F. Froeschke. Spatiotemporal predictive model based on environmental factors for juvenile spotted seatrout in Texas estuaries using boosted regression trees. Fisheries Research, 111(3):131–138, November 2011. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783611002530.

Fredou:2016:LHT

[FFG^{+}16] Flávia Lucena Frédou, Thierry Frédou, Daniel Gaertner, Laurie Kell, Michel Potier, Pascal Bach, Paulo Travassos, Fábio Hazin, and Frédéric Ménard. Life history traits and fishery patterns of teleosts caught by the tuna longline fishery in the South Atlantic and Indian Oceans. Fisheries Research, 179(??):308–321, July 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616300698.

Fortunato:2017:IPF

[FGCG^{+}17] Roberta Callicó Fortunato, Mariano González-Castro, Aida Reguera Galán, Ignacio García Alonso, Cecilia Kunert, Vicent Benedito Durà, and Alejandra Volpedo. Identification of potential fish stocks and lifetime movement patterns of Mugil liza Valenciennes 1836 in the Southwestern Atlantic Ocean. Fisheries Research, 193(??):164–172, September 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783617301108.

Froehlich:2019:DMP

[FGK19] Catheline Y. M. Froehlich, Andres Garcia, and Richard J. Kline. Daily movement patterns of red snapper (Lutjanus campechanus) on a large artificial reef. Fisheries Research, 209(??):49–57, January 2019. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618302418.

Fluvia:2012:BSB

[FGRIT^{+}12] Modest Fluvià, Anna Garriga, Ricard Rigall-I-Torrent, Ernesto Rodríguez-Carámbula, and Albert Saló. Buyer and seller behavior in fish markets organized as Dutch auctions: Evidence from a wholesale fish market in Southern Europe. Fisheries Research, 127–128(??):18–25, September 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783612001397.

Forrestal:2019:ALS

[FGS19] Francesca C. Forrestal, C. Phillip Goodyear, and Michael Schirripa. Applications of the longline simulator (LLSIM) using US pelagic longline logbook data and Atlantic blue marlin. Fisheries Research, 211(??):331–337, March 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783618303436.

Fonteneau:2015:FYD

[FH15] Alain Fonteneau and Jean-Pierre Hallier. Fifty years of dart tag recoveries for tropical tuna: a global comparison of results for the western Pacific, eastern Pacific, Atlantic, and Indian Oceans. Fisheries Research, 163(??):7–22, March 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783614001155.

Frost:2017:LOV

[FH17] Hans S. Frost and Ayoe Hoff. The landing obligation in view of different management regimes. Fisheries Research, 195(??):202–213, November 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617301972.

Finucci:2019:DAB

[FHB^{+}19] B. Finucci, R. J. Hurst, N. W. Bagley, S. S. Al Dhaheri, and E. M. Grandcourt. Diversity, abundance, behaviour, and catchability of fishes from trap catch and underwater video in the Arabian Gulf. Fisheries Research, 220(??):Article 105342, December 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619301973.

Frisch:2019:RPM

[FHH^{+}19] Ashley J. Frisch, Jean-Paul A. Hobbs, Scott T. Hansen, David H. Williamson, Mary C. Bonin, Geoffrey P. Jones, and Justin R. Rizzari. Recovery potential of mutualistic anemone and anemonefish populations. Fisheries Research, 218(??):1–9, October 2019. CODEN FISRDJ.

ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783619301158.

Fowler:2017:ARO

[FHK17] Anthony John Fowler, Paul Andrew Hamer, and Jodie Kemp. Age-related otolith chemistry profiles help resolve demographics and meta-population structure of a widely-dispersed, coastal fishery species. Fisheries Research, 189(??):77–94, May 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S016578361730019X.

Forrest:2018:PAH

[FHK18] Robyn E. Forrest, Kendra R. Holt, and A. Robert Kronlund. Performance of alternative harvest control rules for two Pacific groundfish stocks with uncertain natural mortality: Bias, robustness and trade-offs. Fisheries Research, 206(??):259–286, October 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618301061.

Frandsen:2010:SBA

[FHM10] Rikke Petri Frandsen, Bent Herrmann, and Niels Madsen. A simulation-based attempt to quantify the morphological component of size selection of Nephrops norvegicus in trawl codends. Fisheries Research, 101(3):156–167, January 15, 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578360900263X.

Frandsen:2011:DCC

[FHMK11] R. P. Frandsen, B. Herrmann, N. Madsen, and L. A. Krag. Development of a codend concept to improve size selectivity of Nephrops (Nephrops norvegicus) in a multi-species fishery. Fisheries Research, 111(1–2):116–126, September 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783611002360.

Ferincz:2016:DFD

[FHS^{+}16] Árpád Ferincz, Zsófia Horváth, Ádám Staszny, András Ács, Nora Kováts, Csaba F. Vad, Judit Csaba, Szandra Sütő, and Gábor Paulovits. Desiccation frequency drives local invasions of non-native gibel carp (Carassius gibelio) in the catchment of a large, shallow lake (Lake Balaton, Hungary). Fisheries Research, 173(??):37–44, January 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361500154X.

Fitzcharles:2012:RDB

[Fit12] E. M. Fitzcharles. Rapid discrimination between four Antarctic fish species, genus Macrourus, using HRM analysis. Fisheries Research, 127–128(??):166–170, September 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783612000574.

Free:2017:ROA

[FJWD17] Christopher M. Free, Olaf P. Jensen, John Wiedenmann, and Jonathan J. Deroba. The refined ORCS approach: a catch-based method for estimating stock status and catch limits for data-poor fish stocks. Fisheries Research, 193(??):60–70, September 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361730084X.

Frouzova:2011:DAT

[FKM11] Jaroslava Frouzova, Jan Kubecka, and Tomas Mrkvicka. Differences in acoustic target strength pattern between fish with one- and two-chambered swimbladder during rotation in the horizontal plane. Fisheries Research, 109(1):114–118, April 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783611000452.

Font:2011:SIR

[FL11] Toni Font and Josep Lloret. Socioeconomic implications of recreational shore angling for the management of coastal resources in a Mediterranean marine protected area. Fisheries Research, 108(1):214–217, February 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610003073.

Fetherman:2013:ADF

[FL13] Eric R. Fetherman and Jesse M. Lepak. Addressing depletion failure and estimating gear efficiency using back-calculation of capture probabilities. Fisheries Research, 147(??):284–289, October 2013. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783613001574.

Fitzgerald:2019:CRR

[FLW^{+}19] Sean P. Fitzgerald, Hunter S. Lenihan, Jono R. Wilson, Carolynn S. Culver, and Matthew Potoski. Collaborative research reveals cryptic declines within the multispecies California rock crab fishery. Fisheries Research, 220(??):Article 105340, December 2019. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361930195X.

Feeley:2018:SMM

[FMA^{+}18] Michael W. Feeley, Danielle Morley, Alejandro Acosta, Paul Barbera, John Hunt, Ted Switzer, and Michael Burton. Spawning migration movements of mutton snapper in Tortugas, Florida: Spatial dynamics within a marine reserve network. Fisheries Research, 204(??):209–223, August 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783618300560.

Ferreri:2019:VSM

[FMB^{+}19] Rosalia Ferreri, Richard S. McBride, Marco Barra, Antonella Gargano, Salvatore Mangano, Maurizio Pulizzi, Salvatore Aronica, Angelo Bonanno, and Gualtiero Basilone. Variation in size at maturity by horse mackerel (Trachurus trachurus) within the central Mediterranean Sea: Implications for investigating drivers of local productivity and applications for resource assessments. Fisheries Research, 211(??):291–299, March 2019. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618303394.

Fox:2019:DAE

[FMD19] Clive J. Fox, Anne McLay, and Stephanie Dickens. Development and application of electrofishing with towed video as a new survey method for razor clams (Ensis spp.). Fisheries Research, 214(??):76–84, June 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361930027X.

Flem:2017:TEC

[FMF^{+}17] Belinda Flem, Vidar Moen, Tor Erik Finne, Hildegunn Viljugrein, and Anja Bråthen Kristoffersen. Trace element composition of smolt scales from Atlantic salmon (Salmo salar L.), geographic variation between hatcheries. Fisheries Research, 190(??):183–196, June 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617300474.

Fitzpatrick:2013:UAI

[FMH13] Conor Fitzpatrick, Dianne McLean, and Euan S. Harvey. Using artificial illumination to survey nocturnal reef fish. Fisheries Research, 146(??):41–50, September 2013. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783613000842.

Forrest:2013:MED

[FMMW13] Robyn E. Forrest, Murdoch K. McAllister, Steven J. D. Martell, and Carl J. Walters. Modelling the effects of density-dependent mortality in juvenile red snapper caught as bycatch in Gulf of Mexico shrimp fisheries: Implications for management. Fisheries Research, 146(??):102–120, September 2013. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783613000891.

Fulanda:2011:ETD

[FO11] Bernerd Fulanda and Jun Ohtomi. Effect of tow duration on estimations of CPUE and abundance of the grenadier Coelorinchus jordani (Gadiformes, Macrouridae). Fisheries Research, 110(2):298–304, July 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611001688.

Fock:2014:EHT

[Foc14] H. O. Fock. Estimating historical trawling effort in the German Bight from 1924 to 1938. Fisheries Research, 154(??):26–37, June 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S016578361400040X.

Freire:2010:FBM

[FP10] Kátia M. F. Freire and Daniel Pauly. Fishing down Brazilian marine food webs, with emphasis on the east Brazil large marine ecosystem. Fisheries Research, 105(1):57–62, June 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578361000055X.

Fiorentino:2013:CBA

[FPC^{+}13] Fabio Fiorentino, Bernardo Patti, Francesco Colloca, Angelo Bonanno, Gualtiero Basilone, Vita Gancitano, Germana Garofalo, Sergey Goncharov, Michele Gristina, Giuseppe Sinacori, and Salvatore Mazzola. A comparison between acoustic and bottom trawl estimates to reconstruct the biomass trends of sardine and anchovy in the Strait of Sicily (Central Mediterranean). Fisheries Research, 147(??):290–295, October 2013. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613001525.

Fauconnet:2019:OFD

[FPC^{+}19] L. Fauconnet, C. K. Pham, A. Canha, P. Afonso, H. Diogo, M. Machete, H. M. Silva, F. Vandeperre, and T. Morato. An overview of fisheries discards in the Azores. Fisheries Research, 209(??):230–241, January 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618302704.

Fulton:2016:DRE

[FPD^{+}16] Elizabeth A. Fulton, André E. Punt, Catherine M. Dichmont, Rebecca Gorton, Miriana Sporcic, Natalie Dowling, L. Richard Little, Malcolm Haddon, Neil Klaer, and David C. Smith. Developing risk equivalent datarich and data-limited harvest strategies. Fisheries Research, 183(??):574–587, November 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616302168.

Feenstra:2017:IAR

[FPM17] John Feenstra, André E. Punt, and Richard McGarvey. Inferring absolute recruitment and legal size population numbers of southern rock lobster (Jasus edwardsii) in South Australia’s southern zone fishery using extended forms of depletion modelling. Fisheries Research, 191(??):164–178, July 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783617300565.

Faria:2018:IEL

[FPM^{+}18] João Faria, Alfonso Pita, Gustavo M. Martins, Pedro A. Ribeiro, Stephen J. Hawkins, Pablo Presa, and Ana I. Neto. Inbreeding in the exploited limpet Patella aspera across the Macaronesia archipelagos (NE Atlantic): Implications for conservation. Fisheries Research, 198(??):180–188, February 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578361730245X.

Fay:2011:ISU

[FPS11] Gavin Fay, André E. Punt, and Anthony D. M. Smith. Impacts of spatial uncertainty on performance of age structurebased harvest strategies for blue eye trevalla (Hyperoglyphe antarctica). Fisheries Research, 110(3):391–407, August 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783611001640.

Foster:2017:MBI

[FPSDB17] Daniel G. Foster, Jeffrey R. Pulver, Elizabeth Scott-Denton, and Charles Bergmann. Minimizing bycatch and improving efficiency in the commercial bottom longline fishery in the Eastern Gulf of Mexico. Fisheries Research, 196(??):117–125, December 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617302217.

Foster:2015:SFA

[FPSS15] Daniel G. Foster, Glenn R. Parsons, Derke Snodgrass, and Arvind Shah. At-sea factors that affect yellowfin tuna grade in the Gulf of Mexico pelagic longline tuna fishery. Fisheries Research, 164(??):59–63, April 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614003063.

Flores:2017:EFU

[FQT17] Andrés Flores, Dante Queirolo, and Marcia Tomasoni. Estimating fecundity using image analysis in Cervimunida johni and Pleuroncodes monodon. Fisheries Research, 190(??):34– 42, June 2017. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783617300279.

Francis:2012:REN

[Fra12] R. I. C. Chris Francis. The reliability of estimates of natural mortality from stock assessment models. Fisheries Research, 119–120(??):133–134, May 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611003717.

Francis:2014:RMS

[Fra14] R. I. C. Chris Francis. Replacing the multinomial in stock assessment models: a first step. Fisheries Research, 151(??):70–84, March 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783613003093.

Francis:2016:EBC

[Fra16a] R. I. C. Chris Francis. Explaining bias caused by the wrong recruitment penalty in stock assessment models. Fisheries Research, 181(??):114–115, September 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616300960.

Francis:2016:GAS

[Fra16b] R. I. C. Chris Francis. Growth in age-structured stock assessment models. Fisheries Research, 180(??):77–86, August 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783615000909.

Francis:2017:RDW

[Fra17] R. I. C. Chris Francis. Revisiting data weighting in fisheries stock assessment models. Fisheries Research, 192(??):5–15, August 2017. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783616301953.

Favaro:2010:BRO

[FRDC10] Brett Favaro, Dennis T. Rutherford, Stefanie D. Duff, and Isabelle M. Côté. Bycatch of rockfish and other species in British Columbia spot prawn traps: Preliminary assessment using research traps. Fisheries Research, 102(1–2):199–206, February 2010. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783609003154.

Frater:2019:ABG

[FS19] Paul N. Frater and Gunnar Stefansson. Aspects of both growth and selectivity affect growth parameter estimation bias. Fisheries Research, 212(??):154–161, April 2019. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783618303722.

Flood:2016:MFP

[FSA^{+}16] Matthew J. Flood, Ilona Stobutzki, James Andrews, Crispian Ashby, Gavin A. Begg, Rick Fletcher, Caleb Gardner, Lee Georgeson, Scott Hansen, Klaas Hartmann, Patrick Hone, James Larcombe, Luke Maloney, Anthony Moore, Justin Roach, Anthony Roelofs, Keith Sainsbury, Thor Saunders, Sean Sloan, Anthony D. M. Smith, John Stewart, Carolyn Stewardson, and Brent S. Wise. Multijurisdictional fisheries performance reporting: How Australia’s nationally standardised approach to assessing stock status compares. Fisheries Research, 183(??):559–573, November 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783616300248.

Froeschke:2011:ASF

[FSBS11] Bridgette F. Froeschke, Blair Sterba-Boatwright, and Gregory W. Stunz. Assessing southern flounder (Paralichthys lethostigma) long-term population trends in the northern Gulf of Mexico using time series analyses. Fisheries Research, 108(2–3):291–298, March 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610003486.

Forrest:2015:DME

[FSFP15] Robyn E. Forrest, Marie Savina, Elizabeth A. Fulton, and Tony J. Pitcher. Do marine ecosystem models give consistent policy evaluations? A comparison of Atlantis and Ecosim. Fisheries Research, 167(??):293–312, July 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615001009.

Forrestal:2019:TRC

[FSG^{+}19] Francesca C. Forrestal, Michael Schirripa, C. Phillip Goodyear, Haritz Arrizabalaga, Elizabeth A. Babcock, Rui Coelho, Walter Ingram, Matthew Lauretta, Mauricio Ortiz, Rishi Sharma, and John Walter. Testing robustness of CPUE standardization and inclusion of environmental variables with simulated longline catch datasets. Fisheries Research, 210(??):1–13, February 2019. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783618302662.

Fairchild:2013:CSW

[FSH^{+}13] Elizabeth A. Fairchild, Laughlin Siceloff, W. Huntting Howell, Bill Hoffman, and Michael P. Armstrong. Coastal spawning by winter flounder and a reassessment of essential fish habitat in the Gulf of Maine. Fisheries Research, 141(??):118–129, April 2013. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612001634.

Fuller:2015:VMB

[FSH^{+}15] Daniel W. Fuller, Kurt M. Schaefer, John Hampton, Sylvain Caillot, Bruno M. Leroy, and David G. Itano. Vertical movements, behavior, and habitat of bigeye tuna (Thunnus obesus) in the equatorial central Pacific Ocean. Fisheries Research, 172(??):57–70, December 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615300047.

Froese:2008:SMH

[FSPWG08] Rainer Froese, Amanda Stern-Pirlot, Henning Winker, and Didier Gascuel. Size matters: How single-species management can contribute to ecosystem-based fisheries management. Fisheries Research, 92(2–3):231–241, August 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360800043X. See comments [Sve14,

Sve15] and response [FWG15].

Fischer:2016:TEL

[FT16] Sönke Fischer and Sven Thatje. Temperature effects on life-history traits cause challenges to the management of brachyuran crab fisheries in the Humboldt Current: a review. Fisheries Research, 183(??):461–468, November 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578361630220X.

Fang:2016:PAB

[FTJ^{+}16] Zhou Fang, Katherine Thompson, Yue Jin, Xinjun Chen, and Yong Chen. Preliminary analysis of beak stable isotopes (δ ^{13}C and δ ^{15}N) stock variation of neon flying squid, Ommastrephes bartramii, in the North Pacific Ocean. Fisheries Research, 177(??):153–163, May 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616300121.

Fauconnet:2015:CCT

[FTM^{+}15] Laurence Fauconnet, Verena M. Trenkel, Gilles Morandeau, Nathalie Caill-Milly, and Marie-Joëlle Rochet. Characterizing catches taken by different gears as a step towards evaluating fishing pressure on fish communities. Fisheries Research, 164(??):238–248, April 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614003452.

Furukawa:2014:VMP

[FTN^{+}14] Seishiro Furukawa, Yuichi Tsuda, Gregory N. Nishihara, Ko Fujioka, Seiji Ohshimo, Seitaro Tomoe, Naoyuki Nakatsuka, Hideaki Kimura, Takashi Aoshima, Hisao Kanehara, Takashi Kitagawa, Wei-Chuan Chiang, Hideaki Nakata, and Ryo Kawabe. Vertical movements of Pacific bluefin tuna (Thunnus orientalis) and dolphinfish (Coryphaena hippurus) relative to the thermocline in the northern East China Sea. Fisheries Research, 149(??):86–91, January 2014. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783613002142.

Florisson:2018:RVC

[FTWC18] James H. Florisson, James R. Tweedley, Tim H. E. Walker, and Jennifer A. Chaplin. Reef vision: a citizen science program for monitoring the fish faunas of artificial reefs. Fisheries Research, 206(??):296–308, October 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618301425.

Ferguson:2011:OSE

[FWG11] Greg J. Ferguson, Tim M. Ward, and Bronwyn M. Gillanders. Otolith shape and elemental composition: Complementary tools for stock discrimination of mulloway (Argyrosomus japonicus) in southern Australia. Fisheries Research, 110(1):75–83, June 2011. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611001238.

Froese:2015:SSM

[FWG15] Rainer Froese, Henning Winker, and Didier Gascuel. Size still matters. a response to Svedäng (2013): Size matters: Ne quid nimis. Fisheries Research, 164(??):329–330, April 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783614003014. See [FSPWG08, Sve14].

Flores:2019:AGI

[FWGM19] Andrés Flores, Rodrigo Wiff, Konstantinos Ganias, and C. Tara Marshall. Accuracy of gonadosomatic index in maturity classification and estimation of maturity ogive. Fisheries Research, 210(??):50–62, February 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618302789.

Ferguson:2014:LHA

[FWIB14] Greg J. Ferguson, Tim M. Ward, Alex Ivey, and Thomas Barnes. Life history of Argyrosomus japonicus, a large sciaenid at the southern part of its global distribution: Implications for fisheries management. Fisheries Research, 151(??):148–157, March 2014. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613002749.

Fletcher:2016:RHS

[FWJ^{+}16] W. J. Fletcher, B. S. Wise, L. M. Joll, N. G. Hall, E. A. Fisher, A. V. Harry, D. V. Fairclough, D. J. Gaughan, K. Travaille, B. W. Molony, and M. Kangas. Refinements to harvest strategies to enable effective implementation of ecosystem based fisheries management for the multisector, multi-species fisheries of Western Australia. Fisheries Research, 183(??):594–608, November 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616301278.

Fitzgerald:2018:DNI

[FWL18] Sean P. Fitzgerald, Jono R. Wilson, and Hunter S. Lenihan. Detecting a need for improved management in a datalimited crab fishery. Fisheries Research, 208(??):133–144, December 2018. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578361830208X.

Frick:2010:TCP

[FWR10] Lorenz H. Frick, Terence I. Walker, and Richard D. Reina. Trawl capture of Port Jackson sharks, Heterodontus portusjacksoni, and gummy sharks, Mustelus antarcticus, in a controlled setting: Effects of tow duration, air exposure and crowding. Fisheries Research, 106(3):344–350, December 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783610002109.

Gunnlaugsson:2019:LAD

[GA19] Stefan B. Gunnlaugsson and Sveinn Agnarsson. Late arrival: the development of resource rent in Icelandic fisheries. Fisheries Research, 214(??):126–135, June 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619300505.

Gutowsky:2015:CDI

[GAB^{+}15] Lee F. G. Gutowsky, Waleed Aslam, Reza Banisaeed, Lauren R. Bell, Kenzie L. Bove, Jacob W. Brownscombe, Graham J. J. Burrows, Elaine Chu, Jennifer M. T. Magel, Andrew M. Rous, and Steven J. Cooke. Considerations for the design and interpretation of fishing release mortality estimates. Fisheries Research, 167(??):64–70, July 2015. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783615000375.

Garcia:2010:FIA

[GAM10] J. García, J. Arteche, and A. Murillas. Fractional integration analysis and its implications on profitability: the case of the mackerel market in the Basque country. Fisheries Research, 106(3):420–429, December 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610002249.

Ganias:2012:TYU

[Gan12] Konstantinos Ganias. Thirty years of using the postovulatory follicles method: Overview, problems and alternatives. Fisheries Research, 117–118(??):63–74, April 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611000130.

Ganias:2013:DIE

[Gan13] Konstantinos Ganias. Determining the indeterminate: Evolving concepts and methods on the assessment of the fecundity pattern of fishes. Fisheries Research, 138(??):23–30, February 2013. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783612001622.

Guerra:2010:CLH

[GAP10] Ángel Guerra, Louise Allcock, and João Pereira. Cephalopod life history, ecology and fisheries: an introduction. Fisheries Research, 106(2):117–124, November 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610002110.

Gwinn:2010:EPR

[GAR10] Daniel C. Gwinn, Micheal S. Allen, and Mark W. Rogers. Evaluation of procedures to reduce bias in fish growth parameter estimates resulting from size-selective sampling. Fisheries Research, 105(2):75–79, July 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361000069X.

Georgian:2019:EHS

[GAR19] Samuel E. Georgian, Owen F. Anderson, and Ashley A. Rowden. Ensemble habitat suitability modeling of vulnerable marine ecosystem indicator taxa to inform deep-sea fisheries management in the South Pacific Ocean. Fisheries Research, 211(??):256–274, March 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618303321.

Griffin:2018:KSK

[GBA^{+}18] Lucas P. Griffin, Jacob W. Brownscombe, Aaron J. Adams, Ross E. Boucek, John T. Finn, Michael R. Heithaus, Jennifer S. Rehage, Steven J. Cooke, and Andy J. Danylchuk. Keeping up with the silver king: Using cooperative acoustic telemetry networks to quantify the movements of Atlantic tarpon (Megalops atlanticus) in the coastal waters of the southeastern United States. Fisheries Research, 205(??):65–76, September 2018. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618301073.

Gingerich:2012:QRS

[GBB^{+}12] Andrew J. Gingerich, Brian J. Bellgraph, Richard S. Brown, Noel T. Tavan, Z. Daniel Deng, and Joseph R. Brown. Quantifying reception strength and omnidirectionality of underwater radio telemetry antennas: Advances and applications for fisheries research. Fisheries Research, 121–122(??):1–8, June 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783612000343.

Grati:2015:EMT

[GBD^{+}15] Fabio Grati, Luca Bolognini, Filippo Domenichetti, Gianna Fabi, Piero Polidori, Angela Santelli, Giuseppe Scarcella, and Alessandra Spagnolo. The effect of monofilament thickness on the catches of gillnets for common sole in the Mediterranean small-scale fishery. Fisheries Research, 164(??):170–177, April 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614003403.

Gunnarsson:2016:STV

[GBEP16] Ásgeir Gunnarsson, Höskuldur Björnsson, Bjarki Elvarsson, and Christophe Pampoulie. Spatio-temporal variation in the reproduction timing of Atlantic wolffish (Anarhichas lupus l) in Icelandic waters and its relationship with size. Fisheries Research, 183(??):404–409, November 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616302144.

Gascuel:2011:FMF

[GBF^{+}11a] D. Gascuel, N. Bez, A. Forest, P. Guillotreau, F. Laloë, J. Lobry, S. Mahévas, B. Mesnil, E. Rivot, S. Rochette, and V. Trenkel. A future for marine fisheries in Europe (Manifesto of the Association Française d’Halieumétrie). Fisheries Research, 109(1):1–6, April 2011. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611000579.

Gutowsky:2011:AED

[GBF11b] Lee F. G. Gutowsky, Jacob W. Brownscombe, and Michael G. Fox. Angling to estimate the density of large round goby (Neogobius melanostomus). Fisheries Research, 108(1):228–231, February 2011. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783610003395.

Goethel:2019:EUD

[GBH^{+}19] Daniel R. Goethel, Katelyn M. Bosley, Dana H. Hanselman, Aaron M. Berger, Jonathan J. Deroba, Brian J. Langseth, and Amy M. Schueller. Exploring the utility of different tag-recovery experimental designs for use in spatially explicit, tag-integrated stock assessment models. Fisheries Research, 219(??):Article 105320, November 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619301675.

Guyader:2013:SSF

[GBK^{+}13] O. Guyader, P. Berthou, C. Koutsikopoulos, F. Alban, S. Demanèche, M. B. Gaspar, R. Eschbaum, E. Fahy, O. Tully, L. Reynal, O. Curtil, K. Frangoudes, and F. Maynou. Small scale fisheries in Europe: a comparative analysis based on a selection of case studies. Fisheries Research, 140(??):1–13, February 2013. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612003438.

Godlewska:2018:NFA

[GBK^{+}18] Małgorzata Godlewska, Helge Balk, Zbigniew Kaczkowski, Tomasz Jurczak, Katarzyna Izydorczyk, Bronisław Długoszewski, Aleksandra Jaskulska, Ilona Gagała-Borowska, and Joanna Mankiewicz-Boczek. Night fish avoidance of Microcystis bloom revealed by simultaneous hydroacoustic measurements of both organisms. Fisheries Research, 207(??):74–84, November 2018. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783618301619.

Gay:2018:ILS

[GBM^{+}18] M. Gay, M. Bao, K. MacKenzie, S. Pascual, K. Buchmann, O. Bourgau, C. Couvreur, S. Mattiucci, M. Paoletti, L. C. Hastie, A. Levsen, and G. J. Pierce. Infection levels and species diversity of ascaridoid nematodes in Atlantic cod, Gadus morhua, are correlated with geographic area and fish size. Fisheries Research, 202(??):90–102, June 2018. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783617301534.

Guidetti:2010:APA

[GBPC10] Paolo Guidetti, Simona Bussotti, Fausto Pizzolante, and Alessandro Ciccolella. Assessing the potential of an artisanal fishing co-management in the marine protected area of Torre Guaceto (southern Adriatic Sea, SE Italy). Fisheries Research, 101(3):180–187, January 15, 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783609002707.

Guyader:2017:ANM

[GBR17] Olivier Guyader, Robert Bauer, and Lionel Reynal. Assessing the number of moored fishing aggregating devices through aerial surveys: a case study from Guadeloupe. Fisheries Research, 185(??):73–82, January 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616303319.

Grieve:2018:EPP

[GBR^{+}18] Bettina Grieve, Lee J. Baumgartner, Wayne Robinson, Luiz G. M. Silva, Karl Pomorin, Garry Thorncraft, and Nathan Ning. Evaluating the placement of PIT tags in tropical river fishes: a case study involving two Mekong River species. Fisheries Research, 200(??):43–48, April 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617303508.

Gwinn:2019:IUV

[GBS19] Daniel C. Gwinn, Nathan M. Bacheler, and Kyle W. Shertzer. Integrating underwater video into traditional fisheries indices using a hierarchical formulation of a state-space model. Fisheries Research, 219(??):Article 105309, November 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783619301560.

Galindo-Cortes:2010:HDE

[GCDAS^{+}10] Gabriela Galindo-Cortes, Juan Antonio De Anda-Montañez, Francisco Arreguín-Sánchez, Silvia Salas, and Eduardo Francisco Balart. How do environmental factors affect the stock– recruitment relationship? The case of the Pacific sardine (Sardinops sagax) of the northeastern Pacific Ocean. Fisheries Research, 102(1–2):173–183, February 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783609003129.

Garcia-Carreras:2015:REE

[GCDE^{+}15] Bernardo García-Carreras, Paul Dolder, Georg H. Engelhard, Christopher P. Lynam, Georgia A. Bayliss-Brown, and Steven Mackinson. Recent experience with effort management in Europe: Implications for mixed fisheries. Fisheries Research, 169(??):52–59, September 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615001332.

Garbin:2014:AGM

[GCK14] Thaís Garbin, Jorge Pablo Castello, and Paul Gerhard Kinas. Age, growth, and mortality of the mullet Mugil liza in Brazil’s southern and southeastern coastal regions. Fisheries Research, 149(??):61–68, January 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361300218X.

Garlock:2019:ELB

[GCL19] T. M. Garlock, E. V. Camp, and K. Lorenzen. Efficacy of largemouth bass stock enhancement in achieving fishery management objectives in Florida. Fisheries Research, 213(??):180–189, May 2019. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619300104.

Grixti:2010:CVL

[GCRV10] Daniel Grixti, Simon D. Conron, Karina Ryan, and Vincent L. Versace. Circle versus longshank hooks: Comparing hooking locations and recreational catch for juvenile snapper Pagrus auratus and King George whiting Sillaginodes punctata. Fisheries Research, 106(1):27–31, October 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783610001529.

Gianelli:2017:UFU

[GD17] Ignacio Gianelli and Omar Defeo. Uruguayan fisheries under an increasingly globalized scenario: Long-term landings and bioeconomic trends. Fisheries Research, 190(??):53– 60, June 2017. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783617300309.

Gruss:2014:UDG

[GDA14] Arnaud Grüss, Michael Drexler, and Cameron H. Ainsworth. Using delta generalized additive models to produce distribution maps for spatially explicit ecosystem models. Fisheries Research, 159(??):11–24, November 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614001581.

Gutierrez:2016:CSP

[GDB^{+}16] Nicolas L. Gutierrez, Omar Defeo, Simon R. Bush, Doug S. Butterworth, Cathy A. Roheim, and Andre E. Punt. The current situation and prospects of fisheries certification and ecolabelling. Fisheries Research, 182(??):1–6, October 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783616301382.

Gordoa:2019:RFS

[GDB19] Ana Gordoa, Arnau L. Dedeu, and Jordi Boada. Recreational fishing in Spain: First national estimates of fisher population size, fishing activity and fisher social profile. Fisheries Research, 211(??):1–12, March 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618302947.

Gruss:2019:RSD

[GDC^{+}19] Arnaud Grüss, Michael D. Drexler, Emily Chancellor, Cameron H. Ainsworth, Jeffrey S. Gleason, John M. Tirpak, Matthew S. Love, and Elizabeth A. Babcock. Representing species distributions in spatially-explicit ecosystem models from presence-only data. Fisheries Research, 210(??):89–105, February 2019. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783618302807.

Granados-Dieseldorff:2013:HCM

[GDHA13] Pablo Granados-Dieseldorff, William D. Heyman, and James Azueta. History and co-management of the artisanal mutton snapper (Lutjanus analis) spawning aggregation fishery at Gladden Spit, Belize, 1950–2011. Fisheries Research, 147(??):213–221, October 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613001598.

Gardner:2015:ESI

[GDW^{+}15] Chris J. Gardner, D. Charles Deeming, Ian Wellby, Carl D. Soulsbury, and Paul E. Eady. Effects of surgically implanted tags and translocation on the movements of common bream Abramis brama (L.). Fisheries Research, 167(??):252–259, July 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783615000934.

Gastauer:2013:TSV

[GFCK13] Sven Gastauer, Sascha M. M. Fässler, Bram Couperus, and Ada Marieke Keller. Target strength and vertical distribution of smelt (Osmerus eperlanus) in the Ijsselmeer based on stationary 200 kHz echosounder recordings. Fisheries Research, 148(??):100–105, November 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613002087.

Golden:2019:APS

[GFJ19] Abigail S. Golden, Christopher M. Free, and Olaf P. Jensen. Angler preferences and satisfaction in a highthreshold bucket-list recreational fishery. Fisheries Research, 220(??):Article 105364, December 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361930219X.

Godlewska:2012:CHE

[GFK^{+}12] M. Godlewska, J. Frouzova, J. Kubecka, W. Wiśniewolski, and J. Szlakowski. Comparison of hydroacoustic estimates with fish census in shallow Malta Reservoir — which TS/L regression to use in horizontal beam applications? Fisheries Research, 123–124(??):90–97, July 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611003523.

Guijarro:2012:SSD

[GFM^{+}12] Beatriz Guijarro, Emanuela Fanelli, Joan Moranta, Joan E. Cartes, and Enric Massutí. Small-scale differences in the distribution and population dynamics of pandalid shrimps in the western Mediterranean in relation to environmental factors. Fisheries Research, 119–120(??):33–47, May 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783611003547.

Gastauer:2016:DBW

[GFO^{+}16] Sven Gastauer, Sascha M. M. Fässler, Ciaran O’Donnell, Åge Høines, Jan Arge Jakobsen, Alexander I. Krysov, Leon Smith, Øyvind Tangen, Valantine Anthonypillai, Ebba Mortensen, Eric Armstrong, Matthias Schaber, and Ben Scoulding. The distribution of blue whiting west of the British Isles and Ireland. Fisheries Research, 183(??):32–43, November 2016. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783616301497.

Garza-Gil:2015:PIC

[GGADSRVL15] M. Dolores Garza-Gil, Lucy Amigo-Dobaño, Juan C. Surís-Regueiro, and Manuel Varela-Lafuente. Perceptions on incentives for compliance with regulation. The case of Spanish fishermen in the Atlantic. Fisheries Research, 170(??):30–38, October 2015. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783615001538.

Gucu:2018:IIA

[GGB^{+}18] Ali Cemal Gücü, Yaşar Genç, Nimet Selda Başçınar, Murat Dağtekin, Elvan Atılgan, Murat Erbay, İlkay Özcan Akpınar, and Sebahattin Kutlu. Inter and intra annual variation in body condition of the Black Sea anchovy, Engraulis encrasicolus ponticus — potential causes and consequences. Fisheries Research, 205(??):21–31, September 2018. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783618300894.

Gonzalez:2018:ARZ

[GGM^{+}18] A. F. González, J. Gracia, I. Miniño, J. Romón, C. Larsson, J. Maroto, M. Regueira, and S. Pascual. Approach to reduce the zoonotic parasite load in fish stocks: When science meets technology. Fisheries Research, 202(??):140– 148, June 2018. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783617302308.

Giusti:2019:IDP

[GGS^{+}19] Alice Giusti, Marcella Guarducci, Nir Stern, Nadav Davidovich, Daniel Golani, and Andrea Armani. The importance of distinguishing pufferfish species (Lagocephalus spp.) in the Mediterranean Sea for ensuring public health: Evaluation of the genetic databases reliability in supporting species identification. Fisheries Research, 210(??):14–21, February 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783618302728.

Grant:2014:PCS

[GH14] Scott M. Grant and Wade Hiscock. Post-capture survival of Atlantic wolffish (Anarhichas lupus) captured by bottom otter trawl: Can live release programs contribute to the recovery of species at risk? Fisheries Research, 151(??):169–176, March 2014. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783613002816.

Gaertner:2015:TST

[GH15a] Daniel Gaertner and Jean Pierre Hallier. Tag shedding by tropical tunas in the Indian Ocean and other factors affecting the shedding rate. Fisheries Research, 163(??):98–105, March 2015. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578361400071X.

Graves:2015:CEP

[GH15b] John E. Graves and Andrij Z. Horodysky. Challenges of estimating post-release mortality of istiophorid billfishes caught in the recreational fishery: a review. Fisheries Research, 166(??):163–168, June 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614003075.

Goncalves:2013:COT

[GHA13] Patrícia Gonçalves, Elisabete Henriques, and Maria Manuel Angélico. Co-occurrence of Trachurus trachurus and Trachurus picturatus spawners in Atlantic Iberian waters and the ability to distinguish their eggs in plankton samples. Fisheries Research, 138(??):139–145, February 2013. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783612002354.

Gale:2011:PIA

[GHE^{+}11] Marika Kirstin Gale, Scott G. Hinch, Erika J. Eliason, Steven J. Cooke, and David A. Patterson. Physiological impairment of adult sockeye salmon in fresh water after simulated capture-and-release across a range of temperatures. Fisheries Research, 112(1–2):85–95, December 2011. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783611002839.

Gurshin:2013:SAT

[GHJ13] Christopher W. D. Gurshin, W. Huntting Howell, and J. Michael Jech. Synoptic acoustic and trawl surveys of spring-spawning Atlantic cod in the Gulf of Maine cod spawning protection area. Fisheries Research, 141(??):44– 61, April 2013. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783612002871.

Gutowsky:2011:IIM

[GHL^{+}11] Lee F. G. Gutowsky, Phil M. Harrison, Sean J. Landsman, Michael Power, and Steven J. Cooke. Injury and immediate mortality associated with recreational troll capture of bull trout (Salvelinus confluentus) in a reservoir in the Kootenay–Rocky Mountain region of British Columbia. Fisheries Research, 109(2–3):379–383, May 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611000944.

Gianasi:2016:ETO

[GHM16] Bruno L. Gianasi, Jean-François Hamel, and Annie Mercier. Experimental test of optimal holding conditions for live transport of temperate sea cucumbers. Fisheries Research, 174(??):298–308, February 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615301272.

Gardner:2015:BEC

[GHP^{+}15] Caleb Gardner, Klaas Hartmann, André E. Punt, Sarah Jennings, and Bridget Green. Bio-economics of commercial scale translocation of southern rock lobster. Fisheries Research, 162(??):29–36, February 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614002835.

Gardner:2015:PME

[GHPJ15] Caleb Gardner, Klaas Hartmann, André E. Punt, and Sarah Jennings. In pursuit of maximum economic yield in an ITQ managed lobster fishery. Fisheries Research, 161(??):285–292, January 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783614002604.

Grimaldo:2019:CFE

[GHS^{+}19] Eduardo Grimaldo, Bent Herrmann, Biao Su, Heidi Moe Føre, Jørgen Vollstad, Leonore Olsen, Roger B. Larsen, and Ivan Tatone. Comparison of fishing efficiency between biodegradable gillnets and conventional nylon gillnets. Fisheries Research, 213(??):67–74, May 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619300037.

Guerra:2015:SHS

[GHUG^{+}15] Ángel Guerra, Jorge Hernández-Urcera, Manuel E. Garci, Marta Sestelo, Marcos Regueira, Ángel F. González, Miguel Cabanellas-Reboredo, Matías Calvo-Manazza, and Beatriz Morales-Nin. Spawning habitat selection by Octopus vulgaris: New insights for a more effective management of this resource. Fisheries Research, 167(??):313–322, July 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783615001010.

Guerra:2016:SHS

[GHUG^{+}16] Ángel Guerra, Jorge Hernández-Urcera, Manuel E. Garci, Marta Sestelo, Marcos Regueira, Miguel Gilcoto, and Ángel F. González. Spawning habitat selection by the common cuttlefish Sepia officinalis in the Cíes Islands (Northwest Spain). Fisheries Research, 183(??):44–54, November 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783616301400.

Godlewska:2016:DFB

[GIK^{+}16] Małgorzata Godlewska, Katarzyna Izydorczyk, Zbigniew Kaczkowski, Adam Jóźwik, Bronisław Długoszewski, Shaowen Ye, Yuxi Lian, and Jean Guillard. Do fish and bluegreen algae blooms coexist in space and time? Fisheries Research, 173(??):93–100, January 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783615001952.

Gonzalez-Irusta:2016:SGH

[GIW16] J. M. González-Irusta and P. J. Wright. Spawning grounds of haddock (Melanogrammus aeglefinus) in the North Sea and West of Scotland. Fisheries Research, 183(??):180–191, November 2016. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783616301771.

Gonzalez-Irusta:2017:SGW

[GIW17] J. M. González-Irusta and P. J. Wright. Spawning grounds of whiting (Merlangius merlangus). Fisheries Research, 195(??):141–151, November 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617301790.

Gray:2014:DRS

[GJRR14] Charles A. Gray, Daniel D. Johnson, Darren Reynolds, and Douglas Rotherham. Development of rapid sampling procedures for an exploited bivalve in the swash zone on exposed ocean beaches. Fisheries Research, 154(??):205–212, June 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783614000733.

Guy:2013:ONP

[GJS^{+}13] T. J. Guy, S. L. Jennings, R. M. Suryan, E. F. Melvin, M. A. Bellman, L. T. Ballance, B. A. Blackie, D. A. Croll, T. Deguchi, T. O. Geernaert, R. W. Henry, M. Hester, K. D. Hyrenbach, J. Jahncke, M. A. Kappes, K. Ozaki, J. Roletto, F. Sato, W. J. Sydeman, and J. E. Zamon. Overlap of North Pacific albatrosses with the U.S. west coast groundfish and shrimp fisheries. Fisheries Research, 147(??):222–234, October 2013. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783613001616.

Gold:2010:PSC

[GJSR10] J. R. Gold, A. M. C. Jobity, E. Saillant, and M. A. Renshaw. Population structure of carite (Scomberomorus brasiliensis) in waters offshore of Trinidad and northern Venezuela. Fisheries Research, 103(1–3):30–39, April 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610000135.

Gray:2018:DCC

[GK18] Charles A. Gray and Steven J. Kennelly. Diversity and composition of catches and discards in a recreational charter fishery. Fisheries Research, 199(??):44–52, March 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783617303296.

Ginter:2015:CSS

[GKKK15] Kai Ginter, Andu Kangur, Peeter Kangur, and Külli Kangur. Consequences of size-selective harvesting and changing climate on the pikeperch Sander lucioperca in two large shallow north temperate lakes. Fisheries Research, 165(??):63– 70, May 2015. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783615000041.

Gruss:2012:ELS

[GKL12] Arnaud Grüss, David M. Kaplan, and Christophe Lett. Estimating local settler–recruit relationship parameters for complex spatially explicit models. Fisheries Research, 127–128(??):34–39, September 2012. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612001543.

Galal-Khallaf:2017:SBP

[GKMGO^{+}17] Asmaa Galal-Khallaf, Khaled Mohammed-Geba, Alaa G. M. Osman, Khaled Y. AbouelFadl, Yaisel J. Borrell, and Eva Garcia-Vazquez. SNP-based PCR-RFLP, T-RFLP and FINS methodologies for the identification of commercial fish species in Egypt. Fisheries Research, 185(??):34–42, January 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783616303241.

Galal-Khallaf:2016:CSA

[GKOC^{+}16] Asmaa Galal-Khallaf, Alaa G. M. Osman, Carlos E. Carleos, Eva Garcia-Vazquez, and Yaisel J. Borrell. A case study for assessing fish traceability in Egyptian aquafeed formulations using pyrosequencing and metabarcoding. Fisheries Research, 174(??):143–150, February 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615300850.

Ghinter:2019:ISG

[GLA19] Léopold Ghinter, Yvan Lambert, and Céline Audet. Influence of sex on growth of juvenile Greenland halibut (Reinhardtius hippoglossoides) in the Estuary and Gulf of St. Lawrence. Fisheries Research, 219(??):Article 105332, November 2019. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783619301870.

Garavelli:2019:EPS

[GLB^{+}19] Lysel Garavelli, Timothy J. Linley, Brian J. Bellgraph, Briana M. Rhode, Jill M. Janak, and Alison H. Colotelo. Evaluation of passage and sorting of adult Pacific salmonids through a novel fish passage technology. Fisheries Research, 212(??):40–47, April 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783618303539.

Grimaldo:2011:FDN

[GLG^{+}11] Eduardo Grimaldo, Ira Leifer, Svein Helge Gjøsund, Roger B. Larsen, Henrik Jeuthe, and Sünnje Basedow. Field demonstration of a novel towed, area bubbleplume zooplankton (Calanus sp.) harvester. Fisheries Research, 107(1–3):147–158, January 2011. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783610002778.

Gerritsen:2012:SPR

[GLMK12] H. D. Gerritsen, C. Lordan, C. Minto, and S. B. M. Kraak. Spatial patterns in the retained catch composition of Irish demersal otter trawlers: High-resolution fisheries data as a management tool. Fisheries Research, 129–130(??):127–136, October 2012. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783612002032.

Gronbaek:2018:GTF

[GLMP18] Lone Grønbæk, Marko Lindroos, Gordon Munro, and Pedro Pintassilgo. Game theory and fisheries. Fisheries Research, 203(??):1–5, July 2018. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617303351.

Graves:2015:PSI

[GM15a] John E. Graves and Jan R. McDowell. Population structure of istiophorid billfishes. Fisheries Research, 166(??):21–28, June 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783614002616.

Guillen:2015:CFS

[GM15b] Jordi Guillen and Francesc Maynou. Characterisation of fish species based on ex-vessel prices and its management implications: an application to the Spanish Mediterranean. Fisheries Research, 167(??):22–29, July 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615000302.

Gomez-Morales:2018:UPM

[GMCL^{+}18] Maria Angeles Gómez-Morales, Cristina Martínez Castro, Marco Lalle, Rosa Fernández, Patrizio Pezzotti, Elvira Abollo, and Edoardo Pozio. UV-press method versus artificial digestion method to detect Anisakidae L3 in fish fillets: Comparative study and suitability for the industry. Fisheries Research, 202(??):22–28, June 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616304210.

Gourguet:2013:MMF

[GMD^{+}13] S. Gourguet, C. Macher, L. Doyen, O. Thébaud, M. Bertignac, and O. Guyader. Managing mixed fisheries for bio-economic viability. Fisheries Research, 140(??):46–62, February 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783612003499.

Guisande:2010:PIE

[GMHPV^{+}10] C. Guisande, A. Manjarrés-Hernández, P. Pelayo-Villamil, C. Granado-Lorencio, I. Riveiro, A. Acuña, E. Prieto-Piraquive, E. Janeiro, J. M. Matías, C. Patti, B. Patti, S. Mazzola, S. Jiménez, V. Duque, and F. Salmerón. IPez: an expert system for the taxonomic identification of fishes based on machine learning techniques. Fisheries Research, 102(3):240–247, March 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783609003191.

Ghazali:2013:DVS

[GMJ^{+}13] S. M. Ghazali, J. C. Montgomery, A. G. Jeffs, Z. Ibrahim, and C. A. Radford. The diel variation and spatial extent of the underwater sound around a fish aggregation device (FAD). Fisheries Research, 148(??):9–17, November 2013.

CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783613001793.

Goldstein:2017:CDA

[GMMW17] Jason S. Goldstein, Elizabeth M. Morrissey, Erika D. Moretti, and Winsor H. Watson. A comparison of the distribution and abundance of European green crabs and American lobsters in the Great Bay Estuary, New Hampshire, USA. Fisheries Research, 189(??):10–17, May 2017. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783617300024.

Geffen:2013:CAO

[GMNPM^{+}13] Audrey J. Geffen, Beatriz Morales-Nin, Sílvia Pérez-Mayol, Alicia M. Cantarero-Roldán, Julie Skadal, and Antonio Tovar-Sánchez. Chemical analysis of otoliths: Cross validation between techniques and laboratories. Fisheries Research, 143(??):67–80, June 2013. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613000088.

Garrett:2012:FLS

[GMS12] Angus Garrett, Phil MacMullen, and David Symes. Fisheries as learning systems: Interactive learning as the basis for improved decision making. Fisheries Research, 127–128(??):182–187, September 2012. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612000458.

Gladics:2017:FSS

[GMS^{+}17] Amanda J. Gladics, Edward F. Melvin, Robert M. Suryan, Thomas P. Good, Jason E. Jannot, and Troy J. Guy. Fishery-specific solutions to seabird bycatch in the U.S. West Coast sablefish fishery. Fisheries Research, 196(??):85–95, December 2017. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783617302291.

Gerber:2019:EFP

[GMSP19] Kayla M. Gerber, Martha E. Mather, Joseph M. Smith, and Zachary J. Peterson. Evaluation of a field protocol for internally-tagging fish predators using difficultto-tag ictalurid catfish as examples. Fisheries Research, 209(??):58–66, January 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783618302388.

Gutierrez:2011:SDP

[GMUD11] Nicolás L. Gutiérrez, Arianna Masello, Gabriela Uscudun, and Omar Defeo. Spatial distribution patterns in biomass and population structure of the deep sea red crab Chaceon notialis in the Southwestern Atlantic Ocean. Fisheries Research, 110(1):59–66, June 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783611001214.

Grande:2014:RTR

[GMZ^{+}14] Maitane Grande, Hilario Murua, Iker Zudaire, Nicolas Goñi, and Nathalie Bodin. Reproductive timing and reproductive capacity of the skipjack tuna (Katsuwonus pelamis) in the western Indian Ocean. Fisheries Research, 156(??):14–22, August 2014. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783614001453.

Geffen:2012:EDR

[GN12] Audrey J. Geffen and Richard D. M. Nash. Egg development rates for use in egg production methods (EPMs) and beyond. Fisheries Research, 117–118(??):48–62, April 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361100004X.

Gianelli:2019:MST

[GOD19] Ignacio Gianelli, Leonardo Ortega, and Omar Defeo. Modeling short-term fishing dynamics in a small-scale intertidal shellfishery. Fisheries Research, 209(??):242–250, January 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783618302698.

Goodyear:2016:MTV

[Goo16] C. Phillip Goodyear. Modeling the time-varying density distribution of highly migratory species: Atlantic blue marlin as an example. Fisheries Research, 183(??):469–481, November 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783616302284.

Guidetti:2013:UOM

[GPBA13] Paolo Guidetti, Mario Petrillo, Giuseppe De Benedetto, and Giancarlo Albertelli. The use of otolith microchemistry to investigate spawning patterns of European anchovy: a case study in the eastern Ligurian Sea (NW Mediterranean). Fisheries Research, 139(??):1–4, March 2013. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783612003372.

Guerra:2011:CCO

[GPdR11] Ángel Guerra, Julio M. Portela, and José Luis del Río. Cephalopods caught in the outer Patagonian shelf and its upper and medium slope in relation to the main oceanographic features. Fisheries Research, 109(1):179–186, April 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783611000580.

Guinez:2016:PDC

[GPP^{+}16] R. Guiñez, A. Pita, M. Pérez, C. Briones, S. A. Navarrete, J. Toro, and P. Presa. Present-day connectivity of historical stocks of the ecosystem engineer Perumytilus purpuratus along 4500 km of the Chilean Coast. Fisheries Research, 179(??):322–332, July 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616300431.

Grimaldo:2015:ECT

[GPS15] Eduardo Grimaldo, Roar Pedersen, and Manu Sistiaga. Energy consumption of three different trawl configurations used in the Barents Sea demersal trawl fishery. Fisheries Research, 165(??):71–73, May 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615000090.

Garcia:2011:WLT

[GPSA11] Dorleta Garcia, Raúl Prellezo, Marina Santurtun, and Luis Arregi. Winners and losers of a technical change: a case study of long-term management of the Northern European hake. Fisheries Research, 110(1):98–110, June 2011. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783611001275.

Grati:2010:EBT

[GPSF10] Fabio Grati, Piero Polidori, Giuseppe Scarcella, and Gianna Fabi. Estimation of basket trap selectivity for changeable nassa (Nassarius mutabilis) in the Adriatic Sea. Fisheries Research, 101(1–2):100–107, January 5, 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783609002586.

Gamito:2016:TLV

[GPT^{+}16] Rita Gamito, Cristina Pita, Célia Teixeira, Maria J. Costa, and Henrique N. Cabral. Trends in landings and vulnerability to climate change in different fleet components in the Portuguese coast. Fisheries Research, 181(??):93–101, September 2016. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783616301126.

Guyomard:2019:IFG

[GPT^{+}19] David Guyomard, Christophe Perry, Pierre Ugo Tournoux, Geremy Cliff, Victor Peddemors, and Sébastien Jaquemet. An innovative fishing gear to enhance the release of non-target species in coastal shark-control programs: the SMART (shark management alert in real-time) drumline. Fisheries Research, 216(??):6–17, August 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619300761.

Groeneveld:2016:PSF

[GQ16] Rolf A. Groeneveld and Martin F. Quaas. Promoting selective fisheries through certification? An analysis of the PNA unassociated-sets purse seine fishery. Fisheries Research, 182(??):69–78, October 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615301119.

Gonzalez-Quiros:2011:LHM

[GQdÁdMGP^{+}11] Rafael González-Quirós, Juan del Árbol, María del Mar García-Pacheco, Alfonso J. Silva-García, José María Naranjo, and Beatriz Morales-Nin. Life-history of the meagre Argyrosomus regius in the Gulf of Cádiz (SW Iberian Peninsula). Fisheries Research, 109(1):140–149, April 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783611000518.

Gray:2016:EFD

[Gra16] Charles A. Gray. Evaluation of fishery-dependent sampling strategies for monitoring a small-scale beach clam fishery. Fisheries Research, 177(??):24–30, May 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616300078.

Guzzo:2014:ERB

[GRB14] Matthew M. Guzzo, Michael D. Rennie, and Paul J. Blanchfield. Evaluating the relationship between mean catch per unit effort and abundance for littoral cyprinids in small boreal shield lakes. Fisheries Research, 150(??):100–108, February 2014. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783613002476.

Garcia-Rodriguez:2011:SPS

[GRGGCACG11] Francisco Javier García-Rodríguez, Silvia Alejandra García-Gasca, José De La Cruz-Agüero, and Víctor Manuel Cota-Gómez. A study of the population structure of the Pacific sardine Sardinops sagax (Jenyns, 1842) in Mexico based on morphometric and genetic analyses. Fisheries Research, 107(1–3):169–176, January 2011. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610002821.

Griffiths:2010:SAE

[Gri10] S. P. Griffiths. Stock assessment and efficacy of size limits on longtail tuna (Thunnus tonggol) caught in Australian waters. Fisheries Research, 102(3):248–257, March 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609003208.

Griffiths:2012:RCC

[Gri12] Shane P. Griffiths. Recreational catch composition, catch rates, effort and expenditure in a specialised landbased pelagic game fish fishery. Fisheries Research, 127–128(??):40–44, September 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783612001385.

Gonzalez:2018:BPF

[GRO^{+}18] Ángel F. González, Helena Rodríguez, Luis Outeiriño, Carlos Vello, Christian Larsson, and Santiago Pascual. A biobanking platform for fish-borne zoonotic parasites: a traceable system to preserve samples, data and money. Fisheries Research, 202(??):29–37, June 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617300838.

Garcia-Rodriguez:2017:GVH

[GRPEMEVV17] Francisco J. García-Rodríguez, Ricardo Perez-Enriquez, Armando Medina-Espinoza, and Armando Vega-Velázquez. Genetic variability and historic stability of the California spiny lobster Panulirus interruptus in the Gulf of California. Fisheries Research, 185(??):130–136, January 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783616303095.

Gupta:2015:ARF

[GRS^{+}15] Nishikant Gupta, Rajeev Raghavan, Kuppusamy Sivakumar, Vinod Mathur, and Adrian C. Pinder. Assessing recreational fisheries in an emerging economy: Knowledge, perceptions and attitudes of catch-and-release anglers in India. Fisheries Research, 165(??):79–84, May 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615000168.

Ganias:2010:MBF

[GRVN10] Konstantinos Ganias, Maria Rakka, Theocharis Vavalidis, and Cristina Nunes. Measuring batch fecundity using automated particle counting. Fisheries Research, 106(3):570–574, December 2010. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783610002250.

Gillson:2012:SER

[GS12] Jonathan Gillson and James Scandol. Spreading the economic risk: Harvesting strategies of multi-method inshore fisheries during drought in eastern Australia. Fisheries Research, 113(1):106–117, January 2012. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611003109.

Gauteplass:2018:CCA

[GS18] Asle Gauteplass and Anders Skonhoft. Conflict and cooperation in an age structured fishery. Fisheries Research, 203(??):35–45, July 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783617302412.

Gunnarsson:2019:MPE

[GSB^{+}19] Ásgeir Gunnarsson, Jón Sólmundsson, Höskuldur Björnsson, Gudjón Sigurdsson, and Christophe Pampoulie. Migration pattern and evidence of homing in Atlantic wolffish (Anarhichas lupus). Fisheries Research, 215(??):69–75, July 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783619300608.

Guillotreau:2011:FTA

[GSDD11] Patrice Guillotreau, Frédéric Salladarré, Patrice Dewals, and Laurent Dagorn. Fishing tuna around Fish Aggregating Devices (FADs) vs. free swimming schools: Skipper decision and other determining factors. Fisheries Research, 109(2–3):234–242, May 2011. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611000786.

Grimaldo:2015:ELP

[GSH^{+}15] Eduardo Grimaldo, Manu Sistiaga, Bent Herrmann, Svein H. Gjøsund, and Terje Jørgensen. Effect of the lifting panel on selectivity of a compulsory grid section (Sort-V) used by the demersal trawler fleet in the Barents Sea cod fishery. Fisheries Research, 170(??):158–165, October 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783615001770.

Grimaldo:2014:DCC

[GSL14] Eduardo Grimaldo, Manu Sistiaga, and Roger B. Larsen. Development of catch control devices in the Barents Sea cod fishery. Fisheries Research, 155(??):122–126, July 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783614000812.

Gargan:2015:SWA

[GSØT15] Patrick G. Gargan, Trevor Stafford, Finn Økland, and Eva B. Thorstad. Survival of wild Atlantic salmon (Salmo salar) after catch and release angling in three Irish rivers. Fisheries Research, 161(??):252–260, January 2015. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783614002495.

Gastauer:2017:EVG

[GSP17a] Sven Gastauer, Ben Scoulding, and Miles Parsons. Estimates of variability of goldband snapper target strength and biomass in three fishing regions within the Northern Demersal Scalefish Fishery (Western Australia). Fisheries Research, 193(??):250–262, September 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617301224.

Gastauer:2017:TAM

[GSP17b] Sven Gastauer, Ben Scoulding, and Miles Parsons. Towards acoustic monitoring of a mixed demersal fishery based on commercial data: the case of the Northern Demersal Scalefish Fishery (Western Australia). Fisheries Research, 195(??):91–104, November 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617301820.

Gowland-Sainz:2015:ELF

[GSTL15] María Gowland-Sainz, Federico Tapella, and Gustavo A. Lovrich. Egg loss in females of two lithodid species following different return-to-the-water protocols. Fisheries Research, 161(??):77–85, January 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614002124.

Gamito:2015:RFC

[GTCC15] Rita Gamito, Célia M. Teixeira, Maria J. Costa, and Henrique N. Cabral. Are regional fisheries’ catches changing with climate? Fisheries Research, 161(??):207–216, January 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783614002288.

Galvan-Tirado:2013:HDG

[GTDJdL^{+}13] Carolina Galván-Tirado, Píndaro Díaz-Jaimes, Francisco J. García de León, Felipe Galván-Magaña, and Manuel Uribe-Alcocer. Historical demography and genetic differentiation inferred from the mitochondrial DNA of the silky shark (Carcharhinus falciformis) in the Pacific Ocean. Fisheries Research, 147(??):36–46, October 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613000921.

Gruss:2017:OSD

[GTS^{+}17] Arnaud Grüss, James T. Thorson, Skyler R. Sagarese, Elizabeth A. Babcock, Mandy Karnauskas, John F. Walter, and Michael Drexler. Ontogenetic spatial distributions of red grouper (Epinephelus morio) and gag grouper (Mycteroperca microlepis) in the U.S. Gulf of Mexico. Fisheries Research, 193(??):129–142, September 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361730111X.

Gunnarsson:2017:STV

[Gun17] Á. Gunnarsson. Spatio-temporal variability in fecundity and atresia of Atlantic wolffish (Anarhichas lupus L.) population in Icelandic waters. Fisheries Research, 195(??):214–221, November 2017. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783617302060.

Gonzalez-Vicente:2012:TLL

[GVDMG12] Lucía González-Vicente, David Díaz, Sandra Mallol, and Raquel Goñi. Tag loss in the lobster Palinurus elephas (Fabricius, 1787) and implications for population assessment with capture-mark-recapture methods. Fisheries Research, 129–130(??):1–7, October 2012. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612001701.

Goulart:2018:EFP

[GVG18] Pedro Goulart, Francisco J. Veiga, and Catarina Grilo. The evolution of fisheries in Portugal: a methodological reappraisal with insights from economics. Fisheries Research, 199(??):76–80, March 2018. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617303302.

Garcia-Vazquez:2012:SMM

[GVMSCJ12] Eva Garcia-Vazquez, Gonzalo Machado-Schiaffino, Daniel Campo, and Francis Juanes. Species misidentification in mixed hake fisheries may lead to overexploitation and population bottlenecks. Fisheries Research, 114(??):52–55, February 2012. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783611002013.

George:2011:FLT

[GVSB11] Grinson George, Ponnumony Vethamony, Kotteppad Sudheesh, and Madavana Thomas Babu. Fish larval transport in a macro-tidal regime: Gulf of Kachchh, west coast of India. Fisheries Research, 110(1):160–169, June 2011. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783611001342.

Gilbey:2017:GSI

[GWB^{+}17] John Gilbey, Vidar Wennevik, Ian R. Bradbury, Peder Fiske, Lars Petter Hansen, Jan Arge Jacobsen, and Ted Potter. Genetic stock identification of Atlantic salmon caught in the Faroese fishery. Fisheries Research, 187(??):110–119, March 2017. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783616304052.

Gruss:2019:EID

[GWB^{+}19] Arnaud Grüss, John F. Walter, Elizabeth A. Babcock, Francesca C. Forrestal, James T. Thorson, Matthew V. Lauretta, and Michael J. Schirripa. Evaluation of the impacts of different treatments of spatio-temporal variation in catch-per-unit-effort standardization models. Fisheries Research, 213(??):75–93, May 2019. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619300086.

Gorniewicz:2018:VRT

[GWM18] Oskar Górniewicz and Agnieszka Wiszniewska-Matyszkiel. Verification and refinement of a two species fish wars model. Fisheries Research, 203(??):22–34, July 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617302977.

Gonzalez-Wanguemert:2016:SPB

[GWVH^{+}16] Mercedes González-Wangüemert, Sara Valente, Filipe Henriques, Jorge A. Domínguez-Godino, and Ester A. Serrão. Setting preliminary biometric baselines for new target sea cucumbers species of the NE Atlantic and Mediterranean fisheries. Fisheries Research, 179(??):57–66, July 2016. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783616300303.

Godin:2013:NER

[GWWW13] Aurelie Cosandey Godin, Tonya Wimmer, John H. Wang, and Boris Worm. No effect from rare-earth metal deterrent on shark bycatch in a commercial pelagic longline trial. Fisheries Research, 143(??):131–135, June 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613000246.

Guo:2019:HAF

[GYCL19] Aihuan Guo, Julin Yuan, Tianjiang Chu, and Qingping Lian. Hydroacoustic assessment of fish resources in three reservoirs: the effects of different management strategies on fish density, biomass and size. Fisheries Research, 215(??): 90–96, July 2019. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578361930061X.

Helser:2012:RHS

[HA12] Thomas E. Helser and Larry Alade. A retrospective of the hake stocks off the Atlantic and Pacific coasts of the United States: Uncertainties and challenges facing assessment and management in a complex environment. Fisheries Research, 114(??):2–18, February 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611003080.

Heberer:2010:ICR

[HAB^{+}10] C. Heberer, S. A. Aalbers, D. Bernal, S. Kohin, B. Di-Fiore, and C. A. Sepulveda. Insights into catch-and-release survivorship and stress-induced blood biochemistry of common thresher sharks (Alopias vulpinus) captured in the southern California recreational fishery. Fisheries Research, 106(3):495–500, December 2010. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610002444.

Hilborn:2017:WDF

[HAB^{+}17a] Ray Hilborn, Ricardo O. Amoroso, Eugenia Bogazzi, Olaf P. Jensen, Ana M. Parma, Cody Szuwalski, and Carl J. Walters. When does fishing forage species affect their predators? Fisheries Research, 191(??):211–221, July 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617300176. See [HAB^{+}18a, PBB^{+}18].

Humber:2017:ASS

[HAB^{+}17b] Frances Humber, Emmanuel Trabonjy Andriamahaino, Thomas Beriziny, Radonirina Botosoamananto, Brendan John Godley, Charlotte Gough, Stephanie Pedron, Volanirina Ramahery, and Annette Cameron Broderick. Assessing the small-scale shark fishery of Madagascar through community-based monitoring and knowledge. Fisheries Research, 186 (part 1)(??):131–143, February 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361630265X.

Hilborn:2018:CWD

[HAB^{+}18a] Ray Hilborn, Ricardo O. Amoroso, Eugenia Bogazzi, Olaf P. Jensen, Ana M. Parma, Cody Szuwalski, and Carl J. Walters. Corrigendum to “When does fishing forage species affect their predators?” [Fish. Res. 191 (2017) 211–221]. Fisheries Research, 206(??):309, October 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618301310. See [HAB^{+}17a].

Hilborn:2018:RPA

[HAB^{+}18b] Ray Hilborn, Ricardo O. Amoroso, Eugenia Bogazzi, Olaf P. Jensen, Ana M. Parma, Cody Szuwalski, and Carl J. Walters. Response to Pikitch et al. Fisheries Research, 198(??):224, February 2018. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617302084.

Hannesson:2013:ZAF

[Han13] Rögnvaldur Hannesson. Zonal attachment of fish stocks and management cooperation. Fisheries Research, 140(??): 149–154, February 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783613000040.

Hansson:2016:CHA

[Han16a] Sture Hansson. Comments to Heikinheimo et al. (2014) — on recruitment in Baltic Sea pikeperch. Fisheries Research, 179(??):351, July 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/

S0165783616300170. See [HPHR14, HM16a, Han16b].

Hansson:2016:RRO

[Han16b] Sture Hansson. Reply to response from Outi Heikinheimo and Timo J. Marjomäki on my comment on Heikinheimo et al., 2014: Spawning stock-recruitment relationship in pikeperch Sander lucioperca (L.) in the Baltic Sea, with temperature as an environmental effect. Fisheries Research, 183(??):482, November 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0165783616302193. See [HPHR14, HM16a, HM16b].

Hannesson:2018:DDG

[Han18a] Rögnvaldur Hannesson. Density-dependent growth and cannibalism in Northeast Arctic cod: Some implications for fishing strategies. Fisheries Research, 204(??):1–7, August 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783618300328.

Hannesson:2018:SSG

[Han18b] Rögnvaldur Hannesson. Shared stocks, game theory and the zonal attachment principle. Fisheries Research, 203(??):6– 11, July 2018. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783617302096.

Harbitz:2019:ZSD

[Har19] Alf Harbitz. A zigzag survey design for continuous transect sampling with guaranteed equal coverage probability. Fisheries Research, 213(??):151–159, May 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619300165.

He:2011:RGN

[HB11] Pingguo He and Vincent Balzano. Rope grid: a new grid design to further reduce finfish bycatch in the Gulf of Maine pink shrimp fishery. Fisheries Research, 111(1–2):100–107, September 2011. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783611002347.

He:2012:EGS

[HB12] Pingguo He and Vincent Balzano. The effect of grid spacing on size selectivity of shrimps in a pink shrimp trawl with a dual-grid size-sorting system. Fisheries Research, 121–122(??):81–87, June 2012. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612000434.

He:2013:NST

[HB13] Pingguo He and Vincent Balzano. A new shrimp trawl combination grid system that reduces small shrimp and finfish bycatch. Fisheries Research, 140(??):20–27, February 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578361200344X.

Hamilton:2015:RRAa

[HB15a] Sheryl Hamilton and G. Barry Baker. Review of research and assessments on the efficacy of sea lion exclusion devices in reducing the incidental mortality of New Zealand sea lions Phocarctos hookeri in the Auckland Islands squid trawl fishery. Fisheries Research, 161(??):200–206, January 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783614002240.

Hamilton:2015:RRAb

[HB15b] Sheryl Hamilton and G. Barry Baker. Review of research and assessments on the efficacy of Sea Lion Exclusion Devices in reducing the incidental mortality of New Zealand sea lions Phocarctos hookeri in the Auckland Islands squid trawl fishery: Reply to Robertson (2015). Fisheries Research, 165(??):130–132, May 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614003579.

Hintzen:2012:VOS

[HBB^{+}12] Niels T. Hintzen, Francois Bastardie, Doug Beare, Gerjan J. Piet, Clara Ulrich, Nicolas Deporte, Josefine Egekvist, and Henrik Degel. VMStools: Open-source software for the processing, analysis and visualisation of fisheries logbook and VMS data. Fisheries Research, 115–116(??):31–43, March 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783611003365.

Haltuch:2019:URP

[HBB^{+}19] M. A. Haltuch, E. N. Brooks, J. Brodziak, J. A. Devine, K. F. Johnson, N. Klibansky, R. D. M. Nash, M. R. Payne, K. W. Shertzer, S. Subbey, and B. K. Wells. Unraveling the recruitment problem: a review of environmentallyinformed forecasting and management strategy evaluation. Fisheries Research, 217(??):198–216, September 2019. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783618303606.

Harper:2012:FFS

[HBC^{+}12] Sarah Harper, Daniele Bevacqua, Rachel Chudnow, Sabrina Giorgi, Victoire Guillonneau, Frédéric Le Manach, Tim Sutor, and Ussif Rashid Sumaila. Fuelling the fisheries subsidy debate: Agreements, loopholes and implications. Fisheries Research, 113(1):143–146, January 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611003250.

Haug:2017:FHL

[HBC^{+}17] Tore Haug, Bjarte Bogstad, Melissa Chierici, Harald Gjøsæter, Elvar H. Hallfredsson, Åge S. Høines, Alf Håkon Hoel, Randi B. Ingvaldsen, Lis Lindal Jørgensen, Tor Knutsen, Harald Loeng, Lars-Johan Naustvoll, Ingolf Røttingen, and Knut Sunnanå. Future harvest of living resources in the Arctic Ocean north of the Nordic and Barents Seas: a review of possibilities and constraints. Fisheries Research, 188(??):38–57, April 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616304131.

Humborstad:2016:SRL

[HBD^{+}16] Odd-Børre Humborstad, Michael Breen, Michael W. Davis, Svein Løkkeborg, Anders Mangor-Jensen, Kjell Ø. Midling, and Rolf E. Olsen. Survival and recovery of longline- and pot-caught cod (Gadus morhua) for use in capture-based aquaculture (CBA). Fisheries Research, 174(??):103–108, February 2016. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783615300771.

Hedeholm:2014:FEA

[HBG14] R. Hedeholm, M. E. Blicher, and P. Grønkjær. First estimates of age and production of lumpsucker (Cyclopterus lumpus) in Greenland. Fisheries Research, 149(??):1–4, January 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783613002099.

Hu:2019:STV

[HBG^{+}19] Guanyu Hu, Robert Boenish, Chunxia Gao, Bai Li, Xinjun Chen, Yong Chen, and André E. Punt. Spatio-temporal variability in trophic ecology of jumbo squid (Dosidicus gigas) in the southeastern Pacific: Insights from isotopic signatures in beaks. Fisheries Research, 212(??):56–62, April 2019.

CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783618303540.

Haynes:2011:ELH

[HBM11] Paula S. Haynes, Deirdre Brophy, and David McGrath. The early life history of turbot (Psetta maxima L.) on nursery grounds along the west coast of Ireland: 2007– 2009, as described by otolith microstructure. Fisheries Research, 110(3):478–482, August 2011. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611002037.

Harma:2012:RFA

[HBMC12] Clémentine Harma, Deirdre Brophy, Cóilín Minto, and Maurice Clarke. The rise and fall of autumn-spawning herring (Clupea harengus L.) in the Celtic Sea between 1959 and 2009: Temporal trends in spawning component diversity. Fisheries Research, 121–122(??):31–42, June 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612000367.

Hansen:2018:LMM

[HBP^{+}18] Flemming Thorbjørn Hansen, Finlay Burns, Søren Post, Uffe Høgsbro Thygesen, and Teunis Jansen. Length measurement methods of Atlantic mackerel (Scomber scombrus) and Atlantic horse mackerel (Trachurus trachurus) — current practice, conversion keys and recommendations. Fisheries Research, 205(??):57–64, September 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618301012.

Hawkins:2016:FSE

[HBS^{+}16] S. J. Hawkins, K. Bohn, D. W. Sims, P. Ribeiro, J. Faria, P. Presa, A. Pita, G. M. Martins, A. I. Neto, M. T. Burrows, and M. J. Genner. Fisheries stocks from an ecological perspective: Disentangling ecological connectivity from genetic interchange. Fisheries Research, 179(??):333–341, July 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783616300157.

Handley:2016:ASS

[HBT16a] Sean J. Handley, Stephen N. Brown, and Ian Tuck. Assessment of scallop spat (Pecten novaezelandiae) transport, handling and tagging mortality for wild fishery enhancement, Golden Bay, New Zealand. Fisheries Research, 179(??):86– 89, July 2016. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783616300467.

Honsey:2016:RSY

[HBT^{+}16b] Andrew E. Honsey, David B. Bunnell, Cary D. Troy, David G. Fielder, Michael V. Thomas, Carey T. Knight, Stephen C. Chong, and Tomas O. Höök. Recruitment synchrony of yellow perch (Perca flavescens, Percidae) in the Great Lakes region, 1966–2008. Fisheries Research, 181(??):214–221, September 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616301345.

Harford:2017:ILT

[HC17] William J. Harford and Thomas R. Carruthers. Interim and long-term performance of static and adaptive management procedures. Fisheries Research, 190(??):84–94, June 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783617300310.

Haugen:2017:SSS

[HCF^{+}17] Janne B. Haugen, Tobey H. Curtis, Paul G. Fernandes, Katherine A. Sosebee, and Paul J. Rago. Sexual segregation of spiny dogfish (Squalus acanthias) off the northeastern United States: Implications for a male-directed fishery. Fisheries Research, 193(??):121–128, September 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617301121.

Hermida:2019:SST

[HCGS19] M. Hermida, B. Cavaleiro, L. Gouveia, and A. Saraiva. Seasonality of skipjack tuna parasites in the Eastern Atlantic provide an insight into its migratory patterns. Fisheries Research, 216(??):167–173, August 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619301079.

Hutchinson:2019:MHU

[HCH^{+}19] Melanie Hutchinson, Daniel M. Coffey, Kim Holland, David Itano, Bruno Leroy, Suzanne Kohin, Russell Vetter, Ashley J. Williams, and Johanna Wren. Movements and habitat use of juvenile silky sharks in the Pacific Ocean inform conservation strategies. Fisheries Research, 210(??):131–142, February 2019. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783618302856.

Hanselman:2012:SDA

[HCHA12] Dana H. Hanselman, William G. Clark, Jonathan Heifetz, and Delsa M. Anderl. Statistical distribution of age readings of known-age sablefish (Anoplopoma fimbria). Fisheries Research, 131–133(??):1–8, November 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612002093.

Himes-Cornell:2015:ACC

[HCK15] Amber Himes-Cornell and Stephen Kasperski. Assessing climate change vulnerability in Alaska’s fishing communities. Fisheries Research, 162(??):1–11, February 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361400277X.

Hernandez:2011:CTP

[HCM^{+}11] Frank J. Hernandez, Laure Carassou, Sarah Muffelman, Sean P. Powers, and William M. Graham. Comparison of two plankton net mesh sizes for ichthyoplankton collection in the northern Gulf of Mexico. Fisheries Research, 108(2–3):327–335, March 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611000075.

Hyun:2014:FME

[HCR14] Saang-Yoon Hyun, Steven X. Cadrin, and Sally Roman. Fixed and mixed effect models for fishery data on depth distribution of Georges Bank yellowtail flounder. Fisheries Research, 157(??):180–186, September 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614001441.

Herdter:2017:DDH

[HCSM17] Elizabeth S. Herdter, Don P. Chambers, Christopher D. Stallings, and Steven A. Murawski. Did the Deepwater Horizon oil spill affect growth of Red Snapper in the Gulf of Mexico? Fisheries Research, 191(??):60–68, July 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617300656.

Hall:2019:INO

[HCWS19] Quentin A. Hall, Judson M. Curtis, Jason Williams, and Gregory W. Stunz. The importance of newly-opened tidal inlets as spawning corridors for adult red drum (Sciaenops ocellatus). Fisheries Research, 212(??):48–55, April 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783618303461.

Hazen:2016:TSS

[HCZ^{+}16] Lucie Hazen, Elodie Le Cornu, Amanda Zerbe, Rebecca Martone, Ashley L. Erickson, and Larry B. Crowder. Translating sustainable seafood frameworks to assess the implementation of ecosystem-based fisheries management. Fisheries Research, 182(??):149–157, October 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615301429.

Horta:2012:SDW

[HD12] Sebastián Horta and Omar Defeo. The spatial dynamics of the whitemouth croaker artisanal fishery in Uruguay and interdependencies with the industrial fleet. Fisheries Research, 125–126(??):121–128, August 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612000823.

Hedger:2013:IAE

[HdED^{+}13] Richard D. Hedger, Elvira de Eyto, Mary Dillane, Ola H. Diserud, Kjetil Hindar, Philip McGinnity, Russell Poole, and Ger Rogan. Improving abundance estimates from electrofishing removal sampling. Fisheries Research, 137(??):104–115, January 2013. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783612002846.

Hauzer:2013:FNI

[HDM13] Melissa Hauzer, Philip Dearden, and Grant Murray. The fisherwomen of Ngazidja island, Comoros: Fisheries livelihoods, impacts, and implications for management. Fisheries Research, 140(??):28–35, February 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612003451.

Hedger:2018:BEE

[HDS^{+}18] Richard D. Hedger, Ola H. Diserud, Odd T. Sandlund, Laila Saksgård, Ola Ugedal, and Gunnbjørn Bremset. Bias in estimates of electrofishing capture probability of juvenile Atlantic salmon. Fisheries Research, 208(??):286–295, December 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783618302261.

Haimovici:2014:ADP

[HdSB^{+}14] Manuel Haimovici, Roberta Aguiar dos Santos, Mara C. R. S. Bainy, Luciano Gomes Fischer, and Luis Gustavo Cardoso. Abundance, distribution and population dynamics of the short fin squid Illex argentinus in southeastern and Southern Brazil. Fisheries Research, 152(??):1–12, April 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783613002178.

Hernandez:2015:DHS

[HDW^{+}15] Sebastián Hernández, Ross Daley, Terry Walker, Matias Braccini, Andrea Varela, Malcolm P. Francis, and Peter A. Ritchie. Demographic history and the South Pacific dispersal barrier for school shark (Galeorhinus galeus) inferred by mitochondrial DNA and microsatellite DNA mark. Fisheries Research, 167(??):132–142, July 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615000557.

Hillary:2015:LBB

[HE15] Richard M. Hillary and J. Paige Eveson. Length-based Brownie mark-recapture models: Derivation and application to Indian Ocean skipjack tuna. Fisheries Research, 163(??):141–151, March 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614001921.

Hendrickson:2011:ECM

[Hen11] Lisa C. Hendrickson. Effects of a codend mesh size increase on size selectivity and catch rates in a small-mesh bottom trawl fishery for longfin inshore squid, Loligo pealeii. Fisheries Research, 108(1):42–51, February 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610003048.

Hardman:2013:SNF

[HER13] Emily Ruth Hardman, Alasdair James Edwards, and Jean Stephen Jovani Raffin. The seine-net fishery of Rodrigues Island, western Indian Ocean: Is it sustainable or in terminal decline? Fisheries Research, 139(??):35–42, March 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578361200330X.

Henderson:2014:ESF

[HF14] M. J. Henderson and M. C. Fabrizio. Estimation of summer flounder (Paralichthys dentatus) mortality rates using markrecapture data from a recreational angler-tagging program. Fisheries Research, 159(??):1–10, November 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614001568.

Hallier:2015:TAM

[HF15a] J. P. Hallier and A. Fonteneau. Tuna aggregation and movement from tagging data: a tuna “hub” in the Indian Ocean. Fisheries Research, 163(??):34–43, March 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361400191X.

Holt:2015:CUP

[HF15b] Carrie A. Holt and Michael J. P. Folkes. Cautions on using percentile-based benchmarks of status for data-limited populations of Pacific salmon under persistent trends in productivity and uncertain outcomes from harvest management. Fisheries Research, 171(??):188–200, November 2015. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783615000144.

He:2019:ERV

[HF19] Xi He and John C. Field. Effects of recruitment variability and fishing history on estimation of stock-recruitment relationships: Two case studies from U.S. West Coast fisheries. Fisheries Research, 217(??):21–34, September 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783618301656.

Harvey:2011:MGR

[HFBS11] Chris J. Harvey, John C. Field, Sabrina G. Beyer, and Susan M. Sogard. Modeling growth and reproduction of chilipepper rockfish under variable environmental conditions. Fisheries Research, 109(1):187–200, April 2011. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783611000592.

He:2015:ESD

[HFBS15] Xi He, John C. Field, Sabrina G. Beyer, and Susan M. Sogard. Effects of size-dependent relative fecundity specifications in fishery stock assessments. Fisheries Research, 165(??):54–62, May 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783615000119.

Hernandez-Flores:2015:GAS

[HFCPSEM15] A. Hernández-Flores, A. Condal, A. Poot-Salazar, and J. C. Espinoza-Méndez. Geostatistical analysis and spatial modeling of population density for the sea cucumbers Isostichopus badionotus and Holothuria floridana on the Yucatan Peninsula, Mexico. Fisheries Research, 172(??):114–124, December 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578361530014X.

Hannan:2013:SSC

[HFD^{+}13] Kristin M. Hannan, Alex Q. Fogg, William B. Driggers, Eric R. Hoffmayer, G. Walter Ingram, and Mark A. Grace. Size selectivity and catch rates of two small coastal shark species caught on circle and J hooks in the northern Gulf of Mexico. Fisheries Research, 147(??):145–149, October 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578361300146X.

Herraiz:2015:ELT

[HFFdlT15] Isabel González Herraiz, A. Celso Fariña, Juan Freire, and José Ramón Cancelo de la Torre. Exploring longterm variability of Nephrops norvegicus landing per unit effort (LPUE) off North Galicia (NW Spain). Fisheries Research, 165(??):121–126, May 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615000089.

Henderson:2014:MPS

[HFL14] M. J. Henderson, M. C. Fabrizio, and J. A. Lucy. Movement patterns of summer flounder near an artificial reef: Effects of fish size and environmental cues. Fisheries Research, 153(??):1–8, May 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783614000034.

Hobday:2018:PRN

[HFO^{+}18] Alistair J. Hobday, Aysha Fleming, Emily M. Ogier, Linda Thomas, Jason R. Hartog, Sara Hornborg, and Robert L. Stephenson. Perceptions regarding the need for broad sustainability assessments of Australian fisheries. Fisheries Research, 208(??):247–257, December 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618302273.

Hurtado-Ferro:2014:UMS

[HFPH14] Felipe Hurtado-Ferro, André E. Punt, and Kevin T. Hill. Use of multiple selectivity patterns as a proxy for spatial structure. Fisheries Research, 158(??):102–115, October 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783613002282.

He:2016:ASS

[HFPL16] Xi He, John C. Field, Donald E. Pearson, and Lyndsey S. Lefebvre. Age sample sizes and their effects on growth estimation and stock assessment outputs: Three case studies from U.S. West Coast fisheries. Fisheries Research, 180(??):92–102, August 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615300643.

He:2012:BHM

[HGA^{+}12] Pingguo He, Christopher Glass, John Annala, Steve Cadrin, Kenneth La Valley, and Michael Pol. Biology, harvesting, management and conservation of hakes. Fisheries Research, 114(??):1, February 2012. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611003821.

Heyman:2012:VFG

[HGD12] William D. Heyman and Pablo Granados-Dieseldorff. The voice of the fishermen of the Gulf of Honduras: Improving regional fisheries management through fisher participation. Fisheries Research, 125–126(??):129–148, August 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783612000963.

Heino:2019:GGM

[HGD19] Mikko Heino, Olav Rune Godø, and Ulf Dieckmann. The generalization of Gulland’s method: How to estimate maturity ogives when juvenile data are missing while spawner demography is known. Fisheries Research, 219(??):Article 105265, November 2019. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361930092X.

HortaeCosta:2013:VSF

[HGG13] Bárbara Horta e Costa, Leonel Gonçalves, and Emanuel J. Gonçalves. Vessels’ site fidelity and spatio-temporal distribution of artisanal fisheries before the implementation of a temperate multiple-use marine protected area. Fisheries Research, 148(??):27–37, November 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613001823.

Hulson:2014:TBB

[HH14] Peter-John F. Hulson and Dana H. Hanselman. Tradeoffs between bias, robustness, and common sense when choosing selectivity forms. Fisheries Research, 158(??):63–73, October 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578361300310X.

Hadjimichael:2016:RSI

[HH16] Maria Hadjimichael and Troels J. Hegland. Really sustainable? Inherent risks of eco-labeling in fisheries. Fisheries Research, 174(??):129–135, February 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615300886.

He:2018:SDH

[HHB18] Song He, Hugo B. Harrison, and Michael L. Berumen. Species delineation and hybrid identification using diagnostic nuclear markers for Plectropomus leopardus and Plectropomus maculatus. Fisheries Research, 206(??):287–291, October 2018. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783618301589.

Hoelting:2013:FAS

[HHCP13] Kristin R. Hoelting, Clara H. Hard, Patrick Christie, and Richard B. Pollnac. Factors affecting support for Puget Sound Marine Protected Areas. Fisheries Research, 144(??): 48–59, July 2013. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783612003025.

Hajas:2011:UPM

[HHD^{+}11] Wayne Hajas, Claudia Hand, Nicholas Duprey, Janet Lochead, and Julie Deault. Using production models with new and developing fisheries: a case study using the sea cucumber Parastichopus californicus in British Columbia, Canada. Fisheries Research, 110(3):421–434, August 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611001986.

Haslob:2012:ADE

[HHH^{+}12] H. Haslob, H. Hauss, H. H. Hinrichsen, R. Voss, U. Böttcher, and G. Kraus. Application of the daily egg production method to Baltic sprat. Fisheries Research, 127–128(??):73–82, September 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783612001270.

Hare:2015:VFA

[HHH15] Steven R. Hare, Shelton J. Harley, and W. John Hampton. Verifying FAD-association in purse seine catches on the basis of catch sampling. Fisheries Research, 172(??):361–372, December 2015. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783615300515.

Holdsworth:2018:ISM

[HHHWJ18] John C. Holdsworth, Bruce W. Hartill, Andy Heinemann, and Jeremy Wynne-Jones. Integrated survey methods to estimate harvest by marine recreational fishers in New Zealand. Fisheries Research, 204(??):424–432, August 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783618300924.

Hallac:2013:DCS

[HHM^{+}13] David E. Hallac, John H. Hunt, Douglas Morrison, Alice Clarke, Tracy A. Ziegler, William C. Sharp, and Robert Johnson. Development of a collaborative science plan to evaluate the conservation efficacy of a no-fishing, no-anchor marine reserve in Dry Tortugas National Park, Florida, USA. Fisheries Research, 144(??):15–22, July 2013. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783612002834.

Huntsberger:2017:PDI

[HHSS17] Carl J. Huntsberger, Jillian R. Hamlin, Ronald J. Smolowitz, and Roxanna M. Smolowitz. Prevalence and description of Ichthyophonus sp. in yellowtail flounder (Limanda ferruginea) from a seasonal survey on Georges Bank. Fisheries Research, 194(??):60–67, October 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S016578361730139X.

Helidoniotis:2011:RSB

[HHTT11] Fay Helidoniotis, Malcolm Haddon, Geoff Tuck, and David Tarbath. The relative suitability of the von Bertalanffy, Gompertz and inverse logistic models for describing growth in blacklip abalone populations (Haliotis rubra) in Tasmania, Australia. Fisheries Research, 112(1–2):13–21, December 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783611002748.

Hellberg:2019:ISS

[HIH19] Rosalee S. Hellberg, Rachel B. Isaacs, and Eduardo L. Hernandez. Identification of shark species in commercial products using DNA barcoding. Fisheries Research, 210(??):81–88, February 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783618302790.

Hilborn:2011:FDE

[Hil11] Ray Hilborn. Future directions in ecosystem based fisheries management: a personal perspective. Fisheries Research, 108(2–3):235–239, March 2011. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611000087.

Hannah:2012:EBI

[HJ12] Robert W. Hannah and Stephen A. Jones. Evaluating the behavioral impairment of escaping fish can help measure the effectiveness of bycatch reduction devices. Fisheries Research, 131–133(??):39–44, November 2012. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783612002160.

Hannah:2011:TNM

[HJLW11] Robert W. Hannah, Stephen A. Jones, Mark J. M. Lomeli, and W. Waldo Wakefield. Trawl net modifications to reduce the bycatch of eulachon (Thaleichthys pacificus) in the ocean shrimp (Pandalus jordani) fishery. Fisheries Research, 110(2):277–282, July 2011. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611001652.

Hart:2013:SSA

[HJT13] Deborah R. Hart, Larry D. Jacobson, and Jiashen Tang. To split or not to split: Assessment of Georges Bank sea scallops in the presence of marine protected areas. Fisheries Research, 144(??):74–83, July 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612003396.

Harrison:2018:NLH

[HKAA18] Hannah L. Harrison, Sophia Kochalski, Robert Arlinghaus, and Øystein Aas. “Nature’s little helpers”: a benefits approach to voluntary cultivation of hatchery fish to support wild Atlantic salmon (Salmo salar) populations in Norway, Wales, and Germany. Fisheries Research, 204(??):348–360, August 2018. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783618300729.

Head:2014:MGS

[HKB14] Melissa A. Head, Aimee A. Keller, and Mark Bradburn. Maturity and growth of sablefish, Anoplopoma fimbria, along the U.S. West Coast. Fisheries Research, 159(??):56–67, November 2014. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578361400160X.

Hansell:2018:LIA

[HKB^{+}18] Alexander C. Hansell, Steven T. Kessel, Lauran R. Brewster, Steven X. Cadrin, Samuel H. Gruber, Gregory B. Skomal, and Tristan L. Guttridge. Local indicators of abundance and demographics for the coastal shark assemblage of bimini, Bahamas. Fisheries Research, 197(??):34–44, January 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783617302643.

Hadjimichael:2013:IRO

[HKEJ13] Maria Hadjimichael, Michel J. Kaiser, and Gareth Edwards-Jones. The impact of regulatory obligations on fishers’ income: Identifying perceptions using a market-testing tool.

Fisheries Research, 137(??):129–140, January 2013. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783612002822.

Hyun:2012:PSE

[HKF^{+}12] Saang-Yoon Hyun, Matthew L. Keefer, Jeff K. Fryer, Michael A. Jepson, Rishi Sharma, Christopher C. Caudill, John M. Whiteaker, and George P. Naughton. Populationspecific escapement of Columbia River fall Chinook salmon: Tradeoffs among estimation techniques. Fisheries Research, 129–130(??):82–93, October 2012. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361200197X.

Herrmann:2018:SSA

[HKK18] Bent Herrmann, Ludvig A. Krag, and Bjørn A. Krafft. Size selection of Antarctic krill (Euphausia superba) in a commercial codend and trawl body. Fisheries Research, 207(??):49–54, November 2018. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618301632.

Hollema:2017:MPS

[HKM^{+}17] Heather M. Hollema, Jeff Kneebone, Stephen D. Mc-Cormick, Greg B. Skomal, and Andy J. Danylchuk. Movement patterns of striped bass (Morone saxatilis) in a tidal coastal embayment in New England. Fisheries Research, 187(??):168–177, March 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361630371X.

Hamer:2012:MOT

[HKRH12] Paul A. Hamer, Jodie Kemp, Simon Robertson, and Jeremy S. Hindell. Multiple otolith techniques aid stock discrimination of a broadly distributed deepwater fishery species, blue grenadier, Macruronus novaezelandiae. Fisheries Research, 113(1):21–34, January 2012. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611002852.

Huang:2010:BDT

[HL10] Hsiang-Wen Huang and Kwang-Ming Liu. Bycatch and discards by Taiwanese large-scale tuna longline fleets in the Indian Ocean. Fisheries Research, 106(3):261–270, December 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783610001827.

Huang:2011:TML

[HL11] Hui Huang and PingSun Leung. Testing for market linkages between Hawaii and Japan’s tuna markets. Fisheries Research, 109(2–3):351–359, May 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611001135.

Heikinheimo:2016:LEO

[HL16] Outi Heikinheimo and Hannu Lehtonen. Letter to the Editor: Overestimated effect of cormorant predation on fisheries catches: Comment to the article by Salmi, J. A. et al., 2015: Perch (Perca fluviatilis) and pikeperch (Sander lucioperca) in the diet of the great cormorant (Phalacrocorax carbo) and effects on catches in the Archipelago Sea, Southwest coast of Finland. Fisheries Research 164, 26–34. Fisheries Research, 179(??):354–357, July 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616300182. See [SAR^{+}15].

Hoolihan:2019:AGB

[HLA19] J. P. Hoolihan, J. Luo, and F. Arocha. Age and growth of blue marlin Makaira nigricans from the central western Atlantic Ocean. Fisheries Research, 220(??):Article 105346, December 2019. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783619302012.

Hernandez:2015:FEA

[HLB^{+}15] C. Hernández, J. Landa, J. Barrado, A. Antolínez, and M. B. Santos. First estimates of age and growth of juvenile black anglerfish (Lophius budegassa), in north-eastern Atlantic waters. Fisheries Research, 161(??):269–272, January 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783614002392.

Hannah:2015:TAL

[HLJ15] Robert W. Hannah, Mark J. M. Lomeli, and Stephen A. Jones. Tests of artificial light for bycatch reduction in an ocean shrimp (Pandalus jordani) trawl: Strong but opposite effects at the footrope and near the bycatch reduction device. Fisheries Research, 170(??):60–67, October 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615001514.

Hoyle:2015:CCR

[HLNH15a] Simon D. Hoyle, Bruno M. Leroy, Simon J. Nicol, and W. John Hampton. Corrigendum to “Covariates of release mortality and tag loss in large-scale tuna tagging experiments” [Fish. Res. 163 (2015) 106–118]. Fisheries Research, 168(??):47–48, August 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361500096X. See [HLNH15b].

Hoyle:2015:CRM

[HLNH15b] Simon D. Hoyle, Bruno M. Leroy, Simon J. Nicol, and W. John Hampton. Covariates of release mortality and tag loss in large-scale tuna tagging experiments. Fisheries Research, 163(??):106–118, March 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0165783614000691. See corrigendum [HLNH15a].

Hordyk:2015:EIH

[HLP15] Adrian R. Hordyk, Neil R. Loneragan, and Jeremy D. Prince. An evaluation of an iterative harvest strategy for data-poor fisheries using the length-based spawning potential ratio assessment methodology. Fisheries Research, 171(??):20–32, November 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615000065.

Herrmann:2016:PSS

[HLS^{+}16] Bent Herrmann, Roger B. Larsen, Manu Sistiaga, Nina A. H. Madsen, Karl G. Aarsæther, Eduardo Grimaldo, and Olafur A. Ingolfsson. Predicting size selection of cod (Gadus morhua) in square mesh codends for demersal seining: a simulation-based approach. Fisheries Research, 184(??):36–46, December 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615300242.

Hervas:2021:QAW

[HLSD21] Susana Hervas, Kai Lorenzen, Michael A. Shane, and Mark A. Drawbridge. Quantitative assessment of a white seabass (Atractoscion nobilis) stock enhancement program in California: Post-release dispersal, growth and survival. Fisheries Research, 105(3):237–243, ???? 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610001311.

Heikinheimo:2016:RSHa

[HM16a] Outi Heikinheimo and Timo J. Marjomäki. Response to Sture Hansson’s comment on Heikinheimo et al., 2014: Spawning stock-recruitment relationship in pikeperch Sander lucioperca (L.) in the Baltic Sea, with temperature as an environmental effect. Fisheries Research, 179(??):352–353, July 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/

S0165783616300686. See [HPHR14, Han16a, Han16b].

Heikinheimo:2016:RSHb

[HM16b] Outi Heikinheimo and Timo J. Marjomäki. Response to Sture Hansson’s second comment on Heikinheimo et al. (2014): Spawning stock–recruitment relationship in pikeperch Sander lucioperca (L.) in the Baltic Sea, with temperature as an environmental effect. Fisheries Research, 183(??):483–484, November 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0165783616302302. See [HPHR14, HM16a, Han16b].

Hussy:2016:EOS

[HMA^{+}16] Karin Hüssy, Henrik Mosegaard, Christoffer Moesgaard Albertsen, Einar Eg Nielsen, Jakob Hemmer-Hansen, and Margit Eero. Evaluation of otolith shape as a tool for stock discrimination in marine fishes using Baltic Sea cod as a case study. Fisheries Research, 174(??):210–218, February 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783615301107.

Higashisaka:2018:PBH

[HMA18] Hiroki Higashisaka, Takashi Matsuishi, and Tomonari Akamatsu. Presence and behavior of harbor porpoises (Phocoena phocoena) around set nets revealed using passive acoustic monitoring. Fisheries Research, 204(??):269–274, August 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783618300778.

Hunt:2019:PSP

[HMD^{+}19] Len M. Hunt, David M. Morris, D. Andrew R. Drake, Jeffrey D. Buckley, and Timothy B. Johnson. Predicting spatial patterns of recreational boating to understand potential impacts to fisheries and aquatic ecosystems. Fisheries Research, 211(??):111–120, March 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618303199.

Hoshino:2012:BAM

[HMGH12] Eriko Hoshino, E. J. Milner-Gulland, and Richard M. Hillary. Bioeconomic adaptive management procedures for short-lived species: a case study of Pacific saury (Cololabis saira) and Japanese common squid (Todarodes pacificus). Fisheries Research, 121–122(??):17–30, June 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612000380.

Hennen:2018:BRP

[HMMP18] Daniel R. Hennen, Roger Mann, Daphne M. Munroe, and Eric N. Powell. Biological reference points for Atlantic surfclam (Spisula solidissima) in warming seas. Fisheries Research, 207(??):126–139, November 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618301838.

Horne:2013:MWS

[HMvHN13] John B. Horne, Paolo Momigliano, Lynne van Herwerden, and Stephen J. Newman. Murky waters: Searching for structure in genetically depauperate blue threadfin populations of Western Australia. Fisheries Research, 146(??):1–6, September 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783613000817.

Hara:2016:BRH

[HN16] Mafaniso Hara and Friday Njaya. Between a rock and a hard place: the need for and challenges to implementation of Rights Based Fisheries Management in small-scale fisheries of southern Lake Malawi. Fisheries Research, 174(??):10–18, February 2016. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783615300527.

Henriques:2017:GPS

[HND^{+}17] Romina Henriques, Erica S. Nielsen, Deon Durholtz, David Japp, and Sophie von der Heyden. Genetic population substructuring of kingklip (Genypterus capensis — ophidiidiae), a commercially exploited demersal fish off South Africa. Fisheries Research, 187(??):86–95, March 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616303721.

Harvey:2012:CRE

[HNM^{+}12] Euan S. Harvey, Stephen J. Newman, Dianne L. McLean, Mike Cappo, Jessica J. Meeuwig, and Craig L. Skepper. Comparison of the relative efficiencies of stereo-BRUVs and traps for sampling tropical continental shelf demersal fishes.

Fisheries Research, 125–126(??):108–120, August 2012. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783612000859.

Hjelset:2012:RSC

[HNS12] Ann Merete Hjelset, Einar M. Nilssen, and Jan H. Sundet. Reduced size composition and fecundity related to fishery and invasion history in the introduced red king crab (Paralithodes camtschaticus) in Norwegian waters. Fisheries Research, 121–122(??):73–80, June 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612000410.

Hammarlund:2018:FMU

[HNW^{+}18] Cecilia Hammarlund, Max Nielsen, Staffan Waldo, Rasmus Nielsen, Ayoe Hoff, and Valerio Bartolino. Fisheries management under nutrient influence: Cod fishery in the Western Baltic Sea. Fisheries Research, 201(??):109–119, May 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783618300195.

Hashimoto:2019:SDP

[HNY^{+}19] Midori Hashimoto, Shota Nishijima, Ryuji Yukami, Chikako Watanabe, Yasuhiro Kamimura, Sho Furuichi, Momoko Ichinokawa, and Hiroshi Okamura. Spatiotemporal dynamics of the Pacific chub mackerel revealed by standardized abundance indices. Fisheries Research, 219 (??):Article 105315, November 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783619301626.

Hochhalter:2012:MSS

[Hoc12] Samuel J. Hochhalter. Modeling submergence success of discarded yelloweye rockfish (Sebastes ruberrimus) and quillback rockfish (Sebastes maliger): Towards improved estimation of total fishery removals. Fisheries Research, 127–128(??):142–147, September 2012. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612001282.

Holland:2011:PCP

[Hol11a] Daniel S. Holland. Planning for changing productivity and catchability in the Maine lobster fishery. Fisheries Research, 110(1):47–58, June 2011. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611001202.

Holt:2011:BRM

[Hol11b] Sidney Holt. Book review: Mismanagement of Marine Fisheries, Alan Lonhurst. Cambridge University Press (2010). Fisheries Research, 110(3):483, August 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611001974.

Horbowy:2011:CSM

[Hor11] Jan Horbowy. Comparison of stock management with production, difference, and age-structured models using operating models. Fisheries Research, 108(1):153–162, February 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783610003401.

Harris:2019:DRE

[HPAA19] Holden E. Harris, William F. Patterson, Robert N. M. Ahrens, and Micheal S. Allen. Detection and removal efficiency of invasive lionfish in the northern Gulf of Mexico. Fisheries Research, 213(??):22–32, May 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619300025.

Hintzen:2010:IET

[HPB10] Niels T. Hintzen, Gerjan J. Piet, and Thomas Brunel. Improved estimation of trawling tracks using cubic Hermite spline interpolation of position registration data. Fisheries Research, 101(1–2):108–115, January 5, 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783609002604.

Helle:2015:SSC

[HPHF15] Kristin Helle, Michael Pennington, Nils-Roar Hareide, and Inge Fossen. Selecting a subset of the commercial catch data for estimating catch per unit effort series for ling (Molva molva L.). Fisheries Research, 165(??):115–120, May 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578361500003X.

Heikinheimo:2014:SSR

[HPHR14] Outi Heikinheimo, Zeynep Pekcan-Hekim, and Jari Raitaniemi. Spawning stock–recruitment relationship in pikeperch Sander lucioperca (L.) in the Baltic Sea, with temperature as an environmental effect. Fisheries Research, 155(??):1–9, July 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/

S0165783614000551. See comments [Han16a] response [HM16a, Han16b, HM16b].

Horbowy:2016:IOA

[HPNA16] Jan Horbowy, Magdalena Podolska, and Katarzyna Nadolna-Ałtyn. Increasing occurrence of anisakid nematodes in the liver of cod (Gadus morhua) from the Baltic Sea: Does infection affect the condition and mortality of fish? Fisheries Research, 179(??):98–103, July 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S016578361630042X.

Hanselman:2018:SWD

[HPP18] Dana H. Hanselman, Brian J. Pyper, and Megan J. Peterson. Sperm whale depredation on longline surveys and implications for the assessment of Alaska sablefish. Fisheries Research, 200(??):75–83, April 2018. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617303594.

Hartill:2016:BTG

[HPRB16] Bruce W. Hartill, George W. Payne, Nicola Rush, and Richard Bian. Bridging the temporal gap: Continuous and cost-effective monitoring of dynamic recreational fisheries by web cameras and creel surveys. Fisheries Research, 183(??):488–497, November 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616301904.

He:2016:SFB

[HPS16] Pingguo He, Michael Pol, and Petri Suuronen. Selective fishing and balanced harvest: Means, methods and implications. Fisheries Research, 184(??):1, December 2016. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783616302685.

Hollenbeck:2015:PSR

[HPSG15] Christopher M. Hollenbeck, David S. Portnoy, Eric Saillant, and John R. Gold. Population structure of red snapper (Lutjanus campechanus) in U.S. waters of the western Atlantic Ocean and the northeastern Gulf of Mexico. Fisheries Research, 172(??):17–25, December 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615001976.

Hofmann:2015:ASH

[HRAK15] Thorben Hofmann, Michael J. Raupach, Pedro Martinez Arbízu, and Thomas Knebelsberger. An application of in situ hybridization for the identification of commercially important fish species. Fisheries Research, 170(??):1–8, October 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783615001435.

Hannah:2014:DEC

[HRB14] Robert W. Hannah, Polly S. Rankin, and Matthew T. O. Blume. The divergent effect of capture depth and associated barotrauma on post-recompression survival of canary (Sebastes pinniger) and yelloweye rockfish (S. ruberrimus). Fisheries Research, 157(??):106–112, September 2014. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783614001416.

Ho:2016:ESE

[HRC16] Nga Thi Thanh Ho, Helen Ross, and Jeffrey Coutts. Evaluation of social and ecological outcomes of fisheries comanagement in Tam giang lagoon, Vietnam. Fisheries Research, 174(??):151–159, February 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615300898.

Halttunen:2021:ICR

[HRT^{+}21] Elina Halttunen, Audun H. Rikardsen, Eva B. Thorstad, Tor F. Næsje, Jenny L. A. Jensen, and Øystein Aas. Impact of catch-and-release practices on behavior and mortality of Atlantic salmon (Salmo salar L.) kelts. Fisheries Research, 105(3):141–147, ???? 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610000810.

Haslob:2013:SVF

[HRUSR13] Holger Haslob, Sonia Rabade-Uberos, and Fran Saborido-Rey. Seasonal variability of fecundity and spawning dynamics of Baltic sprat. Fisheries Research, 138(??):99–109, February 2013. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783612002378.

Huveneers:2013:AGD

[HSB^{+}13] Charlie Huveneers, Joanna Stead, Michael B. Bennett, Kate A. Lee, and Robert G. Harcourt. Age and growth determination of three sympatric wobbegong sharks: How reliable is growth band periodicity in orectolobidae? Fisheries Research, 147(??):413–425, October 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613000829.

Hegna:2019:EOS

[HSB19] Jonathan Hegna, Kim Scribner, and Edward Baker. Evaluation of optimal surgical techniques for intracoelomic transmitter implantation in age-0 lake sturgeon. Fisheries Research, 218(??):198–208, October 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361930147X.

Hyun:2012:IFF

[HSC^{+}12] Saang-Yoon Hyun, Rishi Sharma, John K. Carlile, James G. Norris, Gayle Brown, Ryan J. Briscoe, and Diana Dobson. Integrated forecasts of fall Chinook salmon returns to the Pacific Northwest. Fisheries Research, 125–126(??):306–317, August 2012. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783612001178.

Hughes:2019:EHC

[HSC19] Julian M. Hughes, John Stewart, and Damian Collins. Experimental hyperbaric chamber trials reveal high resilience to barotrauma in Australian snapper (Chrysophrys auratus). Fisheries Research, 212(??):172–182, April 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361830376X.

Hahlbeck:2017:ODO

[HSD^{+}17] Nick Hahlbeck, Kylie L. Scales, Heidi Dewar, Sara M. Maxwell, Steven J. Bograd, and Elliott L. Hazen. Oceanographic determinants of ocean sunfish (Mola mola) and bluefin tuna (Thunnus orientalis) bycatch patterns in the California large mesh drift gillnet fishery. Fisheries Research, 191(??):154–163, July 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617300784.

Heck:2015:ISS

[HSG15] Nadine Heck, Richard C. Stedman, and Marc Gaden. The integration of social science information into Great Lakes fishery management: Opportunities and challenges. Fisheries Research, 167(??):30–37, July 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615000272.

Hunter:2015:FIC

[HSH15] Aidan Hunter, Douglas C. Speirs, and Michael R. Heath. Fishery-induced changes to age and length dependent maturation schedules of three demersal fish species in the Firth of Clyde. Fisheries Research, 170(??):14–23, October 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783615001459.

Hunter:2016:ITG

[HSH16] Aidan Hunter, Douglas C. Speirs, and Michael R. Heath. Investigating trends in the growth of five demersal fish species from the Firth of Clyde and the wider western shelf of Scotland. Fisheries Research, 177(??):71–81, May 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616300054.

Herrmann:2013:USG

[HSL^{+}13] Bent Herrmann, Manu Sistiaga, Roger B. Larsen, Kåre N. Nielsen, and Eduardo Grimaldo. Understanding sorting grid and codend size selectivity of Greenland halibut (Reinhardtius hippoglossoides). Fisheries Research, 146(??): 59–73, September 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016578361300091X.

Herrmann:2019:ETD

[HSLB19] Bent Herrmann, Manu Sistiaga, Roger B. Larsen, and Jesse Brinkhof. Effect of three different codend designs on the size selectivity of juvenile cod in the Barents Sea shrimp trawl fishery. Fisheries Research, 219(??):Article 105337, November 2019. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783619301924.

Holmes:2012:DTA

[HSM^{+}12] Bonnie J. Holmes, Wayne D. Sumpton, David G. Mayer, Ian R. Tibbetts, David T. Neil, and Mike B. Bennett. Declining trends in annual catch rates of the tiger shark (Galeocerdo cuvier) in Queensland, Australia. Fisheries Research, 129–130(??):38–45, October 2012. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612001786.

Herrmann:2017:EEG

[HSRT17] Bent Herrmann, Manu Sistiaga, Lasse Rindahl, and Ivan Tatone. Estimation of the effect of gear design changes on catch efficiency: Methodology and a case study for a Spanish longline fishery targeting hake (Merluccius merluccius). Fisheries Research, 185(??):153–160, January 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578361630306X.

Hobday:2011:ERA

[HSS^{+}11] A. J. Hobday, A. D. M. Smith, I. C. Stobutzki, C. Bulman, R. Daley, J. M. Dambacher, R. A. Deng, J. Dowdney, M. Fuller, D. Furlani, S. P. Griffiths, D. Johnson, R. Kenyon, I. A. Knuckey, S. D. Ling, R. Pitcher, K. J. Sainsbury, M. Sporcic, T. Smith, C. Turnbull, T. I. Walker, S. E. Wayte, H. Webb, A. Williams, B. S. Wise, and S. Zhou. Ecological risk assessment for the effects of fishing. Fisheries Research, 108(2–3):372–384, March 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611000324.

Harry:2016:ADL

[HSS^{+}16] Alastair V. Harry, Richard J. Saunders, Jonathan J. Smart, Peter M. Yates, Colin A. Simpfendorfer, and Andrew J. Tobin. Assessment of a data-limited, multi-species shark fishery in the Great Barrier Reef Marine Park and southeast Queensland. Fisheries Research, 177(??):104–115, May 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783615301600.

Hadden:2018:ETA

[HSS18] Jeremy T. Hadden, Nicholas. J. Smith, and Trent M. Sutton. Effects of transmitter application procedures on growth, survival, and tag retention of juvenile least cisco Coregonus sardinella. Fisheries Research, 199(??):196–201, March 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783617303156.

Harry:2010:IAG

[HST10] Alastair V. Harry, Colin A. Simpfendorfer, and Andrew J. Tobin. Improving age, growth, and maturity estimates for aseasonally reproducing chondrichthyans. Fisheries Research, 106(3):393–403, December 2010. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610002195.

Head:2016:TIE

[HSTK16] Melissa A. Head, Gretchen L. Stokes, James T. Thorson, and Aimee A. Keller. Techniques for improving estimates of maturity ogives in groundfish using double-reads and measurement error models. Fisheries Research, 179(??):251–258, July 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783616300601.

Haggarty:2016:APR

[HSY16] Dana R. Haggarty, Jonathan B. Shurin, and K. Lynne Yamanaka. Assessing population recovery inside British Columbia’s rockfish conservation areas with a remotely operated vehicle. Fisheries Research, 183(??):165–179, November 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783616301898.

Harford:2015:SMR

[HTB15] William J. Harford, Caroline Ton, and Elizabeth A. Babcock. Simulated mark-recovery for spatial assessment of a spiny lobster (Panulirus argus) fishery. Fisheries Research, 165(??):42–53, May 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615000120.

Haslob:2011:SIV

[HTHK11] H. Haslob, J. Tomkiewicz, H. H. Hinrichsen, and G. Kraus. Spatial and interannual variability in Baltic sprat batch fecundity. Fisheries Research, 110(2):289–297, July 2011. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783611001676.

Huang:2015:ICS

[Hua15] Hsiang-Wen Huang. Incidental catch of seabirds and sea turtles by Taiwanese longline fleets in the Pacific Ocean. Fisheries Research, 170(??):179–189, October 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615001848.

Huse:2010:MMS

[HV10] Irene Huse and Aud Vold. Mortality of mackerel (Scomber scombrus L.) after pursing and slipping from a purse seine. Fisheries Research, 106(1):54–59, October 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610001566.

Hessenauer:2018:SSS

[HVD^{+}18] Jan-Michael Hessenauer, Jason Vokoun, Justin Davis, Robert Jacobs, and Eileen O’Donnell. Size structure suppression and obsolete length regulations in recreational fisheries dominated by catch-and-release. Fisheries Research, 200(??):33–42, April 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783617303429.

Hussein:2011:AIAa

[HVJP^{+}11a] C. Hussein, M. Verdoit-Jarraya, J. Pastor, A. Ibrahim, G. Saragoni, D. Pelletier, S. Mahévas, and P. Lenfant. Assessing the impact of artisanal and recreational fishing and protection on a white seabream (Diplodus sargus sargus) population in the north-western Mediterranean Sea using a simulation model.Part 1: Parameterization and simulations. Fisheries Research, 108(1):163–173, February 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783610003425.

Hussein:2011:AIAb

[HVJP^{+}11b] C. Hussein, M. Verdoit-Jarraya, J. Pastor, A. Ibrahim, G. Saragoni, D. Pelletier, S. Mahévas, and P. Lenfant. Assessing the impact of artisanal and recreational fishing and protection on a white seabream (Diplodus sargus sargus) population in the north-western Mediterranean Sea, using a simulation model.Part 2: Sensitivity analysis and management measures. Fisheries Research, 108(1):174–183, February 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783610003437.

Hernaez:2011:SME

[HW11] Patricio Hernáez and Ingo S. Wehrtmann. Sexual maturity and egg production in an unexploited population of the red squat lobster Pleuroncodes monodon (Decapoda, Galatheidae) from Central America. Fisheries Research, 107(1–3):276–282, January 2011. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361000295X.

Herrmann:2013:ITT

[HWM^{+}13] Bent Herrmann, Harald Wienbeck, Waldemar Moderhak, Daniel Stepputtis, and Ludvig Ahm Krag. The influence of twine thickness, twine number and netting orientation on codend selectivity. Fisheries Research, 145(??):22–36, August 2013. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783613000684.

Harms:2010:AFI

[HWS10] John H. Harms, John R. Wallace, and Ian J. Stewart. Analysis of fishery-independent hook and line-based data for use in the stock assessment of bocaccio rockfish (Sebastes paucispinis). Fisheries Research, 106(3):298–309, December 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783610002031.

Hutchinson:2012:ELM

[HWS^{+}12] Melanie Hutchinson, John H. Wang, Yonat Swimmer, Kim Holland, Suzanne Kohin, Heidi Dewar, James Wraith, Russ Vetter, Craig Heberer, and Jimmy Martinez. The effects of a lanthanide metal alloy on shark catch rates. Fisheries Research, 131–133(??):45–51, November 2012. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783612002123.

Herrmann:2015:SSC

[HWS^{+}15] Bent Herrmann, Harald Wienbeck, Daniel Stepputtis, Ludvig Ahm Krag, Jordan Feekings, and Waldemar Moderhak. Size selection in codends made of thin-twined dyneema netting compared to standard codends: a case study with cod, plaice and flounder. Fisheries Research, 167(??):82–91, July 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783615000351.

Hou:2019:PCF

[HZT^{+}19] Yiqun Hou, Xuan Zou, Weiming Tang, Wei Jiang, Jiazhi Zhu, Chenlong Deng, and Yongfeng Zhang. Precise capture of fish movement trajectories in complex environments via ultrasonic signal tag tracking. Fisheries Research, 219(??):Article 105307, November 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619301547.

Ikhwanuddin:2011:BIP

[IAJ^{+}11] M. Ikhwanuddin, G. Azmie, H. M. Juariah, M. Z. Zakaria, and M. A. Ambak. Biological information and population features of mud crab, genus Scylla from mangrove areas of sarawak, Malaysia. Fisheries Research, 108(2–3):299–306, March 2011. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783611000038.

Ishihara:2017:UPB

[IAS^{+}17] Taiki Ishihara, Osamu Abe, Tamaki Shimose, Yukio Takeuchi, and Alexandre Aires da Silva. Use of postbomb radiocarbon dating to validate estimated ages of Pacific bluefin tuna, Thunnus orientalis, of the North Pacific Ocean. Fisheries Research, 189(??):35–41, May 2017. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783616304350.

Ichinokawa:2010:UAA

[IB10] Momoko Ichinokawa and Jon Brodziak. Using adaptive area stratification to standardize catch rates with application to North Pacific swordfish (Xiphias gladius). Fisheries Research, 106(3):249–260, December 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610001785.

Ibanez:2014:FSS

[Ibá14] Ana L. Ibáñez. Fish scale shape variation by year and by geographic location, could scales be useful to trace fish? A case study on the Gulf of Mexico. Fisheries Research, 156(??):34– 38, August 2014. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783614001428.

Ibanez:2015:FTG

[Ibá15] Ana L. Ibáñez. Fish traceability: Guessing the origin of fish from a seafood market using fish scale shape. Fisheries Research, 170(??):82–88, October 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615001654.

Isaac:2016:SID

[ICSR16] Victoria J. Isaac, Leandro Castello, Paulo R. Brasil Santos, and Mauro L. Ruffino. Seasonal and interannual dynamics of river-floodplain multispecies fisheries in relation to flood pulses in the Lower Amazon. Fisheries Research, 183(??):352–359, November 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616302065.

Iervolino:2010:LGD

[IdRH10] Fabiana Iervolino, Emiko Kawakami de Resende, and Alexandre Wagner Silva Hilsdorf. The lack of genetic differentiation of pacu (Piaractus mesopotamicus) populations in the Upper-Paraguay Basin revealed by the mitochondrial DNA D-loop region: Implications for fishery management. Fisheries Research, 101(1–2):27–31, January 5, 2010. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360900246X.

Ingolfsson:2017:EHB

[IEL17] Ólafur Arnar Ingólfsson, Haraldur Arnar Einarsson, and Svein Løkkeborg. The effects of hook and bait sizes on size selectivity and capture efficiency in Icelandic longline fisheries. Fisheries Research, 191(??):10–16, July 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617300541.

Iriondo:2012:MMF

[IGS^{+}12] Ane Iriondo, Dorleta García, Marina Santurtún, José Castro, Iñaki Quincoces, Sigrid Lehuta, Stephanie Mahévas, Paul Marchal, Alex Tidd, and Clara Ulrich. Managing mixed fisheries in the European Western Waters: Application of Fcube methodology. Fisheries Research, 134–136(??):6–16, December 2012. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783612002329.

Izquierdo-Gomez:2015:ECF

[IGSJBS^{+}15] D. Izquierdo-Gomez, P. Sanchez-Jerez, J. T. Bayle-Sempere, N. J. Loader, and C. Garcia de Leaniz. Effects of coastal fish farms on body size and isotope composition of wild penaeid prawn. Fisheries Research, 172(??):50–56, December 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783615300059.

Ibanez:2017:DAP

[IHFAH17] Ana L. Ibáñez, Karla Hernández-Fraga, and Sergio Alvarez-Hernández. Discrimination analysis of phenotypic stocks comparing fish otolith and scale shapes. Fisheries Research, 185(??):6–13, January 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783616303186.

Iwata:2010:CHS

[IIS10] Yoko Iwata, Kingo Ito, and Yasunori Sakurai. Is commercial harvesting of spawning aggregations sustainable?

— the reproductive status of the squid Loligo bleekeri. Fisheries Research, 102(3):286–290, March 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610000020.

Irigoitia:2017:EUN

[IIT17] Manuel Marcial Irigoitia, Inés Silvia Incorvaia, and Juan Tomás Timi. Evaluating the usefulness of natural tags for host population structure in chondrichthyans: Parasite assemblages of Sympterygia bonapartii (Rajiformes: Arhynchobatidae) in the Southwestern Atlantic. Fisheries Research, 195(??):80–90, November 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/

S0165783617301807. See corrigendum [IIT18].

Irigoitia:2018:CEU

[IIT18] Manuel Marcial Irigoitia, Inés Silvia Incorvaia, and Juan Tomás Timi. Corrigendum to “Evaluating the usefulness of natural tags for host population structure in chondrichthyans: Parasite assemblages of Sympterygia bonapartii (Rajiformes: Arhynchobatidae) in the Southwestern Atlantic” [Fish. Res. 195C (2017) 80–90]. Fisheries Research, 199(??):272, March 2018. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783617302990. See [IIT17].

Ingolfsson:2011:EFM

[Ing11] Ólafur Arnar Ingólfsson. The effect of forced mesh opening in the upper panel of a Nephrops trawl on size selection of Nephrops, haddock and whiting. Fisheries Research, 108(1):218–222, February 2011. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610003097.

Islam:2014:EIA

[INSN14] Gazi Md. Nurul Islam, Kusairi Mohd Noh, Shaufique F. Sidique, and Aswani Farhana Mohd Noh. Economic impact of artificial reefs: a case study of small scale fishers in terengganu, peninsular Malaysia. Fisheries Research, 151(??):122– 129, March 2014. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783613002464.

Islam:2011:MPF

[INY11] Gazi Md. Nurul Islam, Kusairi Mohd Noh, and Tai Shzee Yew. Measuring productivity in fishery sector of peninsular Malaysia. Fisheries Research, 108(1):52–57, February 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578361000305X.

Ibanez:2011:IFS

[IO11] Ana L. Ibáñez and Paul O’Higgins. Identifying fish scales: the influence of allometry on scale shape and classification. Fisheries Research, 109(1):54–60, April 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361100035X.

Imsland:2014:LHT

[IÓS^{+}14] Albert K. Imsland, Kristinn Ólafsson, Sigurlaug Skírnisdóttir, Snorri Gunnarsson, Matthías Oddgeirsson, Sara Vandamme, Sarah J. Helyar, Julie Skadal, and Arild Folkvord. Life history of turbot in Icelandic waters: Intra- and interpopulation genetic diversity and otolith tracking of environmental temperatures. Fisheries Research, 155(??):185– 193, July 2014. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783614000861.

Ichinokawa:2014:DCC

[IOT14] Momoko Ichinokawa, Hiroshi Okamura, and Yukio Takeuchi. Data conflict caused by model mis-specification of selectivity in an integrated stock assessment model and its potential effects on stock status estimation. Fisheries Research, 158(??):147–157, October 2014. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614000423.

Ihde:2011:IIM

[IWL^{+}11] Thomas F. Ihde, Michael J. Wilberg, David A. Loewensteiner, David H. Secor, and Thomas J. Miller. The increasing importance of marine recreational fishing in the US: Challenges for management. Fisheries Research, 108(2–3):268–276, March 2011. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783610003413.

Jardim:2015:HCR

[JAB15] Ernesto Jardim, Manuela Azevedo, and Nuno M. Brites. Harvest control rules for data limited stocks using lengthbased reference points and survey biomass indices. Fisheries Research, 171(??):12–19, November 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614003397.

Jabado:2018:FMT

[Jab18] R. W. Jabado. The fate of the most threatened order of elasmobranchs: Shark-like batoids (Rhinopristiformes) in the Arabian Sea and adjacent waters. Fisheries Research, 204(??):448–457, August 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618300961.

Jurajda:2016:UMF

[JAJ^{+}16] P. Jurajda, Z. Adámek, M. Janáč, K. Roche, L. Mikl, L. Rederer, T. Zapletal, V. Koza, and J. Špaček. Use of multiple fish-removal methods during biomanipulation of a drinking water reservoir — evaluation of the first four years. Fisheries Research, 173(??):101–108, January 2016. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783615001423.

Johnston:2011:DAU

[JASP11] Fiona D. Johnston, Robert Arlinghaus, Jim Stelfox, and John R. Post. Decline in angler use despite increased catch rates: Anglers’ response to the implementation of a total catch-and-release regulation. Fisheries Research, 110(1):189–197, June 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S016578361100155X.

Jimenez-Alvarado:2019:HFK

[JATBC19] D. Jiménez-Alvarado, B. Tovar, J. F. Baños, and J. J. Castro. How to fish? Key factors influencing the probability of choosing a recreational fishing modality. Fisheries Research, 212(??):87–96, April 2019. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618303527.

Jaworski:2011:EMP

[Jaw11] Andrzej Jaworski. Evaluation of methods for predicting mean weight-at-age: an application in forecasting yield of four haddock (Melanogrammus aeglefinus) stocks in the Northeast Atlantic. Fisheries Research, 109(1):61–73, April 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783611000361.

Jansen:2015:DDG

[JB15] Teunis Jansen and Finlay Burns. Density dependent growth changes through juvenile and early adult life of North East Atlantic mackerel (Scomber scombrus). Fisheries Research, 169(??):37–44, September 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615001344.

Juza:2018:CTP

[JBB^{+}18] Tomáš Jůza, Petr Blabolil, Roman Baran, Vladislav Draštík, Michaela Holubová, Luboš Kočvara, Milan Muška, Milan Říha, Zuzana Sajdlová, Marek Šmejkal, Michal Tušer, Mojmír Vašek, Lukáš Vejřík, Ivana Vejříková, Arco J. Wagenvoort, Jakub Žák, and Henk A. M. Ketelaars. Comparison of two passive methods for sampling invasive round goby (Neogobius melanostomus) populations at different depths in artificial lakes. Fisheries Research, 207(??):175–181, November 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783618301723.

Jakubaviciute:2018:MDT

[JBD^{+}18] Eglė Jakubavičiūtė, Yves De Blick, Justas Dainys, Linas Ložys, and Jens Olsson. Morphological divergence of threespined stickleback in the Baltic Sea — implications for stock identification. Fisheries Research, 204(??):305–315, August 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783618300869.

Jemaa:2015:EPC

[JBKA15] Sharif Jemaa, Mahmoud Bacha, Gaby Khalaf, and Rachid Amara. Evidence for population complexity of the European anchovy (Engraulis encrasicolus) along its distributional range. Fisheries Research, 168(??):109–116, August 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783615001277.

Jepsen:2013:DLA

[JBMK13] Niels Jepsen, Torsten S. Boutrup, Jonathan D. Midwood, and Anders Koed. Does the level of asepsis impact the success of surgically implanting tags in Atlantic salmon? Fisheries Research, 147(??):344–348, October 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613001811.

Jepsen:2014:FSD

[JBMK14] Niels Jepsen, Torsten S. Boutrup, Jonathan D. Midwood, and Anders Koed. Fish surgery — a dirty business? Comments to a letter submitted by D. Mulcahy and C. A. Harms. Fisheries Research, 156(??):6–8, August 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614000757.

Jones:2018:SSB

[JC18] Michelle K. Jones and Sean Cox. Size-selectivity for British Columbia Sablefish (Anoplopoma fimbria) estimated from a long-term tagging study. Fisheries Research, 199(??):94– 106, March 2018. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783617303223.

Johnson:2016:TDM

[JCA^{+}16] Laura Johnson, Christopher J. Coates, Amaya Albalat, Keith Todd, and Douglas Neil. Temperature-dependent morbidity of ‘nicked’ edible crab, Cancer pagurus. Fisheries Research, 175(??):127–131, March 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615301478.

Juza:2012:ITM

[JČK^{+}12] Tomáš Jůza, Martin Čech, Jan Kubečka, Mojmír Vašek, Jiří Peterka, Michal Kratochvíl, Jaroslava Frouzová, and Josef Matěna. The influence of the trawl mouth opening size and net colour on catch efficiency during sampling of early stages of perch (Perca fluviatilis) and pikeperch (Sander lucioperca) in the bathypelagic layer of a canyon-shaped reservoir. Fisheries Research, 123–124(??):21–25, July 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611003432.

Juza:2021:ITM

[JČK^{+}21] Tomáš Jůza, Martin Čech, Jan Kubečka, Mojmír Vašek, Jiří Peterka, and Josef Matěna. The influence of the trawl mouth opening size and net colour on catch efficiency during sampling of early fish stages. Fisheries Research, 105(3):125– 133, ???? 2021. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783610000743.

Johnson:2016:CAR

[JCT^{+}16] Kelli F. Johnson, Elizabeth Councill, James T. Thorson, Elizabeth Brooks, Richard D. Methot, Jr., and André E. Punt. Can autocorrelated recruitment be estimated using integrated assessment models and how does it affect population forecasts? Fisheries Research, 183(??):222–232, November 2016. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783616301928.

Jackson:2012:EDE

[JCW12] G. Jackson, Y. W. Cheng, and C. B. Wakefield. An evaluation of the daily egg production method to estimate spawning biomass of snapper (Pagrus auratus) in inner shark Bay, Western Australia, following more than a decade of surveys 1997–2007. Fisheries Research, 117–118(??):22–34, April 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783610003218.

Javor:2014:GOI

[JD14] Barbara Javor and Emmanis Dorval. Geography and ontogeny influence the stable oxygen and carbon isotopes of otoliths of Pacific sardine in the California Current. Fisheries Research, 154(??):1–10, June 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614000241.

Jimenez:2015:HET

[JDBF15] H. Jimenez, P. Dumas, L. Bigot, and J. Ferraris. Harvesting effects on tropical invertebrate assemblages in New Caledonia. Fisheries Research, 167(??):75–81, July 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615000338.

Jimenez:2011:IHT

[JDLF11] Haizea Jimenez, Pascal Dumas, Marc Léopold, and Jocelyne Ferraris. Invertebrate harvesting on tropical urban areas: Trends and impact on natural populations (New Caledonia, South Pacific). Fisheries Research, 108(1):195–204, February 2011. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783610003462.

Jackson:2018:BBE

[JDNP18] Laura Stone Jackson, J. Marcus Drymon, T. Reid Nelson, and Sean P. Powers. Biotelemetry based estimates of greater amberjack (Seriola dumerili) post-release mortality in the northern Gulf of Mexico. Fisheries Research, 208(??):239–246, December 2018. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618302157.

Juarez:2018:GMM

[JECR^{+}18] Oscar E. Juárez, Luis Enríquez, Faustino Camarena-Rosales, Leticia Arena, Clara E. Galindo-Sánchez, Fabiola Lafarga-De la Cruz, Laura López-Galindo, Karina Nambo, and Carlos Rosas. Genetic monitoring of the Mexican four-eyed octopus Octopus maya population: New insights and perspectives for the fishery management. Fisheries Research, 206(??):109–114, October 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618301383.

James:2017:ETT

[JEJS17] P. James, T. Evensen, R. Jacobsen, and S. Siikavuopio. Efficiency of trap type, soak time and bait type and quantities for harvesting the sea urchin Strongylocentrotus droebachiensis (Müller) in Norway. Fisheries Research, 193(??):15–20, September 2017. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783617300887.

Jardim:2013:EDU

[JF13] Ernesto Jardim and Ana Cláudia Fernandes. Estimators of discards using fishing effort as auxiliary information with an application to Iberian hake (Merluccius merluccius) exploited by the Portuguese trawl fleets. Fisheries Research, 140(??):105–113, February 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612003505.

Jensen:2015:GTF

[JFT^{+}15] Frank Jensen, Hans Frost, Thomas Thøgersen, Peder Andersen, and Jesper Levring Andersen. Game theory and fish wars: the case of the Northeast Atlantic mackerel fishery. Fisheries Research, 172(??):7–16, December 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361500199X.

Jaric:2012:RBL

[JG12] Ivan Jarić and Zoran Gačić. Relationship between the longevity and the age at maturity in long-lived fish: Rikhter/Efanov’s and Hoenig’s methods. Fisheries Research, 129–130(??):61–63, October 2012. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612001944.

Johnston:2011:DBS

[JHCT11] Danielle Johnston, David Harris, Nick Caputi, and Adrian Thomson. Decline of a blue swimmer crab (Portunus pelagicus) fishery in Western Australia — history, contributing factors and future management strategy. Fisheries Research, 109(1):119–130, April 2011. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611000476.

Jacques:2019:CLM

[JHL^{+}19] Nadine Jacques, Bent Herrmann, Roger B. Larsen, Manu Sistiaga, Jure Brčić, Gökhan Gökçe, and Jesse Brinkhof. Can a large-mesh sieve panel replace or supplement the Nordmøre grid for bycatch mitigation in the northeast Atlantic deep-water shrimp fishery? Fisheries Research, 219(??):Article 105324, November 2019. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619301717.

Jensen:2010:DCR

[JHT^{+}10] Jenny L. A. Jensen, Elina Halttunen, Eva B. Thorstad, Tor F. Næsje, and Audun H. Rikardsen. Does catch-and-release angling alter the migratory behaviour of Atlantic salmon? Fisheries Research, 106(3):550–554, December 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783610002067.

Jalali:2015:PMA

[JIM^{+}15] M. Ali Jalali, Daniel Ierodiaconou, Jacquomo Monk, Harry Gorfine, and Alex Rattray. Predictive mapping of abalone fishing grounds using remotely-sensed LiDAR and commercial catch data. Fisheries Research, 169(??):26–36, September 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783615001320.

Jansen:2016:MDP

[JKK^{+}16] Teunis Jansen, Kasper Kristensen, Paulus Kainge, Deon Durholtz, Tore Strømme, Uffe Høgsbro Thygesen, Margit R. Wilhelm, John Kathena, Tracey P. Fairweather, Sarah Paulus, Henrik Degel, Marek R. Lipinski, and Jan E. Beyer. Migration, distribution and population (stock) structure of shallow-water hake (Merluccius capensis) in the Benguela Current Large Marine Ecosystem inferred using a geostatistical population model. Fisheries Research, 179(??):156– 167, July 2016. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578361630056X.

Jansen:2015:SPS

[JKS^{+}15] Teunis Jansen, Paulus Kainge, Larvika Singh, Margit Wilhelm, Deon Durholtz, Tore Strømme, John Kathena, and Victoria Erasmus. Spawning patterns of shallow-water hake (Merluccius capensis) and deep-water hake (M. paradoxus) in the Benguela Current Large Marine Ecosystem inferred from gonadosomatic indices. Fisheries Research, 172(??):168–180, December 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615300187.

Jorgensen:2017:FCP

[JLF^{+}17] Terje Jørgensen, Svein Løkkeborg, Dag Furevik, Odd-Børre Humborstad, and Francesco De Carlo. Floated cod pots with one entrance reduce probability of escape and increase catch rates compared with pots with two entrances. Fisheries Research, 187(??):41–46, March 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616303435.

Joung:2011:ELH

[JLLL11] Shoou-Jeng Joung, Pei-Hsuan Lee, Kwang-Ming Liu, and Yih-Yia Liao. Estimates of life history parameters of the sharpspine skate, Okamejei acutispina, in the northeastern waters of Taiwan. Fisheries Research, 108(2–3):258–267, March 2011. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783610003383.

Jensen:2019:DEW

[JNE19] Frank Jensen, Max Nielsen, and Hans Ellefsen. Defining economic welfare in fisheries. Fisheries Research, 218(??): 138–154, October 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783619301274.

Johnson:2019:GSJ

[JNF^{+}19] Marc A. Johnson, David L. G. Noakes, Thomas A. Friesen, Andrew H. Dittman, Ryan B. Couture, Carl B. Schreck, Craig Banner, Darran May, and Thomas P. Quinn. Growth, survivorship, and juvenile physiology of triploid steelhead (Oncorhynchus mykiss). Fisheries Research, 220(??):Article 105350, December 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361930205X.

Jensen:2014:ICA

[JNN14] Frank Jensen, Max Nielsen, and Rasmus Nielsen. Increased competition for aquaculture from fisheries: Does improved fisheries management limit aquaculture growth? Fisheries Research, 159(??):25–33, November 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361400157X.

Johnston:2016:FSI

[JP16] Matthew W. Johnston and Sam J. Purkis. Forecasting the success of invasive marine species; lessons learned from purposeful reef fish releases in the Hawaiian Islands. Fisheries Research, 174(??):190–200, February 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615301090.

Jaquemet:2011:DDA

[JPM11] S. Jaquemet, M. Potier, and F. Ménard. Do drifting and anchored Fish Aggregating Devices (FADs) similarly influence tuna feeding habits? A case study from the western Indian Ocean. Fisheries Research, 107(1–3):283–290, January 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783610002961.

Jansen:2019:DVF

[JPO^{+}19] Teunis Jansen, Søren Post, Anna Heida Olafsdottir, Páll Reynisson, Gudmundur J. Óskarsson, and Kristine Engel Arendt. Diel vertical feeding behaviour of Atlantic mackerel (Scomber scombrus) in the Irminger current. Fisheries Research, 214(??):25–34, June 2019. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619300207.

Jimenez-Prada:2014:CDH

[JPSR^{+}14] Pablo Jiménez-Prada, Anastasia Scherbakova, Rodrigo Riera, Beatriz C. Felipe, António V. Sykes, Rui A. Gonçalves, José P. Andrade, Catalina Perales-Raya, Covadonga Rodríguez, and Eduardo Almansa. Characterization of deformed hatchlings of Octopus vulgaris obtained under captivity from a small female. Fisheries Research, 152(??):62–65, April 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783613002440.

Juarez:2010:HMR

[JRA10] Oscar E. Juárez, Carlos Rosas, and Leticia Arena. Heterologous microsatellites reveal moderate genetic structure in the Octopus maya population. Fisheries Research, 106(2):209–213, November 2010. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610002043.

Juza:2013:ARF

[JRD^{+}13] T. Jůza, G. Rakowitz, V. Draštík, P. Blabolil, A. Herzig, M. Kratochvíl, M. Muška, M. Říha, Z. Sajdlová, and J. Kubečka. Avoidance reactions of fish in the trawl mouth opening in a shallow and turbid lake at night. Fisheries Research, 147(??):154–160, October 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613001495.

Janc:2018:HDF

[JRG^{+}18] Anaïs Janc, Gaëtan Richard, Christophe Guinet, John P. Y. Arnould, Maria Ching Villanueva, Guy Duhamel, Nicolas Gasco, and Paul Tixier. How do fishing practices influence sperm whale (Physeter macrocephalus) depredation on demersal longline fisheries? Fisheries Research, 206(??):14–26, October 2018. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783618301255.

Jech:2014:DAH

[JS14] J. Michael Jech and Patrick J. Sullivan. Distribution of Atlantic herring (Clupea harengus) in the Gulf of Maine from 1998 to 2012. Fisheries Research, 156(??):26–33, August 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783614001520.

Janssen:2015:PBM

[JS15] Holger Janßen and Franziska Schwarz. On the potential benefits of marine spatial planning for herring spawning conditions — an example from the western Baltic Sea. Fisheries Research, 170(??):106–115, October 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615001721.

Johnson:2019:APH

[JSB19a] Garrett R. Johnson, Daniel E. Shoup, and Kevin M. Boswell. Accuracy and precision of hydroacoustic estimates of gizzard shad abundance using horizontal beaming. Fisheries Research, 212(??):81–86, April 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618303576.

Johnson:2019:IFO

[JSB19b] Garrett R. Johnson, Daniel E. Shoup, and Kevin M. Boswell. Incorporating fish orientation into target strengthtotal length equations: Horizontal-aspect target-strength equations for gizzard shad Dorosoma cepedianum. Fisheries Research, 218(??):155–165, October 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619301250.

Joo:2015:DFS

[JSG^{+}15] Rocio Joo, Omar Salcedo, Mariano Gutierrez, Ronan Fablet, and Sophie Bertrand. Defining fishing spatial strategies from VMS data: Insights from the world’s largest monospecific fishery. Fisheries Research, 164(??):223–230, April 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783614003592.

Jaworski:2010:FAI

[JSR10] Andrzej Jaworski, Jon Solmundsson, and Stefan Aki Ragnarsson. Fish assemblages inside and outside marine protected areas off northern Iceland: protection effects or environmental confounds? Fisheries Research, 102(1–2):50–59, February 2010. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783609002720.

Jacinto:2015:DAF

[JSYS15] Melchor R. Jacinto, Al Jayson G. Songcuan, Graceous Von Yip, and Mudjekeewis D. Santos. Development and application of the fisheries vulnerability assessment tool (Fish Vool) to tuna and sardine sectors in the Philippines. Fisheries Research, 161(??):174–181, January 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614002215.

Johnson:2014:BGA

[JTHT14] Nicholas S. Johnson, Henry T. Thompson, Christopher Holbrook, and John A. Tix. Blocking and guiding adult sea lamprey with pulsed direct current from vertical electrodes. Fisheries Research, 150(??):38–48, February 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613002336.

Jonsdottir:2014:ESS

[JTP^{+}14] Ingibjörg G. Jónsdóttir, Vilhjalmur Thorsteinsson, Ólafur K. Pálsson, Gunnar G. Tómasson, and Christoph Pampoulie. Evidence of spawning skippers in Atlantic cod from data storage tags. Fisheries Research, 156(??):23–25, August 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783614001544.

Johnson:2019:IVI

[JTP19] Kelli F. Johnson, James T. Thorson, and André E. Punt. Investigating the value of including depth during spatiotemporal index standardization. Fisheries Research, 216(??):126–137, August 2019. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619300955.

Jorgensen:2017:PTM

[JvDB^{+}17] Michelle Grace Pinto Jørgensen, Mikael van Deurs, Ian Anthony Ernest Butts, Kasper Jørgensen, and Jane W. Behrens. PIT-tagging method for small fishes: a case study using sandeel (Ammodytes tobianus). Fisheries Research, 193(??):95–103, September 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617300930.

Johnson:2013:MFS

[JWC^{+}13] Teresa R. Johnson, James A. Wilson, Caitlin Cleaver, Graham Morehead, and Robert Vadas. Modeling fine scale urchin and kelp dynamics: Implications for management of the Maine sea urchin fishery. Fisheries Research, 141(??):107–117, April 2013. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612001646.

Johnston:2012:RAE

[JWM^{+}12] T. A. Johnston, D. M.-M. Wong, M. D. Moles, M. D. Wiegand, J. M. Casselman, and W. C. Leggett. Reproductive allocation in exploited lake whitefish (Coregonus clupeaformis) and walleye (Sander vitreus) populations. Fisheries Research, 125–126(??):225–234, August 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612000860.

Jeffery:2018:RWR

[JWM^{+}18] Nicholas W. Jeffery, Brendan F. Wringe, Meghan C. McBride, Lorraine C. Hamilton, Ryan R. E. Stanley, Louis Bernatchez, Matthew Kent, Marie Clément, John Gilbey, Timothy F. Sheehan, Paul Bentzen, and Ian R. Bradbury. Range-wide regional assignment of Atlantic salmon (Salmo salar) using genome wide single-nucleotide polymorphisms. Fisheries Research, 206(??):163–175, October 2018. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578361830153X.

Jiang:2019:IPC

[JWZ^{+}19] Zhongguan Jiang, Chao Wang, Lizhi Zhou, Wen Xiong, and Chunlong Liu. Impacts of pen culture on alpha and beta diversity of fish communities in a large floodplain lake along the Yangtze River. Fisheries Research, 210(??):41–49, February 2019. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783618302765.

Jia:2013:LFC

[JZL13] Peiqiao Jia, Wenbo Zhang, and Qigen Liu. Lake fisheries in China: Challenges and opportunities. Fisheries Research, 140(??):66–72, February 2013. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612003517.

Kilada:2015:DAD

[KA15] Raouf Kilada and Enzo Acuña. Direct age determination by growth band counts of three commercially important crustacean species in Chile. Fisheries Research, 170(??):134–143, October 2015. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783615001757.

Konan:2010:MVA

[KAGO^{+}10] Koffi Mexmin Konan, Abouo Béatrice Adépo-Gourène, Allassane Ouattara, Wanja Dorothy Nyingy, and Germain Gourène. Morphometric variation among male populations of freshwater shrimp Macrobrachium vollenhovenii Herklots, 1851 from Côte d’Ivoire Rivers. Fisheries Research, 103(1–3):1–8, April 2010. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610000093.

Kai:2019:STC

[Kai19] Mikihiko Kai. Spatio-temporal changes in catch rates of pelagic sharks caught by Japanese research and training vessels in the western and central North Pacific. Fisheries Research, 216(??):177–195, August 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361930058X.

Kumada:2013:MSA

[KAiT^{+}13] Nao Kumada, Tomoko Arima, Jun ichi Tsuboi, Akihiko Ashizawa, and Masahiro Fujioka. The multi-scale aggregative response of cormorants to the mass stocking of fish in rivers. Fisheries Research, 137(??):81–87, January 2013.

CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783612002743.

Kubecka:2016:EEQ

[KBČ^{+}16] Jan Kubečka, David S. Boukal, Martin Čech, Phil Hickley, James F. Kitchell, Daniel Ricard, Lars Rudstam, Kateřina Soukalová, and Robin Welcomme. Ecology and ecological quality of fish in lakes and reservoirs. Fisheries Research, 173(??):1–3, January 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615002386.

Kehayias:2016:PBL

[KBCK16] George Kehayias, Dionysis Bouliopoulos, Nikos Chiotis, and Pavlos Koutra. A photovoltaic-battery-LED lamp raft design for purse seine fishery: Application in a large Mediterranean lake. Fisheries Research, 177(??):18–23, May 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783616300030.

Kovach:2013:ISS

[KBEB13] Adrienne I. Kovach, Timothy S. Breton, Claire Enterline, and David L. Berlinsky. Identifying the spatial scale of population structure in anadromous rainbow smelt (Osmerus mordax). Fisheries Research, 141(??):95–106, April 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783612002147.

Karahan:2014:GMF

[KBG^{+}14] Arzu Karahan, Philippe Borsa, Ali Cemal Gucu, Irfan Kandemir, Emel Ozkan, Yesim Ak Orek, Sinan Can Acan, Evren Koban, and Inci Togan. Geometric morphometrics, Fourier analysis of otolith shape, and nuclear-DNA markers distinguish two anchovy species (Engraulis spp.) in the Eastern Mediterranean Sea. Fisheries Research, 159(??):45–55, November 2014. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783614001623.

Kopack:2015:BRH

[KBL^{+}15] Christopher J. Kopack, E. Dale Broder, Jesse M. Lepak, Eric R. Fetherman, and Lisa M. Angeloni. Behavioral responses of a highly domesticated, predator naïve rainbow trout to chemical cues of predation. Fisheries Research, 169(??):1–7, September 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615001289.

Keating:2014:OSA

[KBOM14] James P. Keating, Deirdre Brophy, Rick A. Officer, and Eugene Mullins. Otolith shape analysis of blue whiting suggests a complex stock structure at their spawning grounds in the Northeast Atlantic. Fisheries Research, 157(??):1–6, September 2014. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783614000915.

Koenigs:2015:VOA

[KBSK15] Ryan P. Koenigs, Ronald M. Bruch, Robert S. Stelzer, and Kendall K. Kamke. Validation of otolith ages for walleye (Sander vitreus) in the winnebago system. Fisheries Research, 167(??):13–21, July 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615000326.

Koliada:2019:LTD

[KBT^{+}19] Ievgen Koliada, Helge Balk, Michal Tušer, Ladislav Ptáček, and Jan Kubečka. Limitations of target detection in horizontal acoustic surveys of extremely shallow water bodies. Fisheries Research, 218(??):94–104, October 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619301201.

Kuhn:2011:ITE

[KC11] Penny S. Kuhn and Jae S. Choi. Influence of temperature on embryo developmental cycles and mortality of female Chionoecetes opilio (snow crab) on the Scotian Shelf, Canada.

Fisheries Research, 107(1–3):245–252, January 2011. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783610002912.

Kompas:2018:MSL

[KC18] Tom Kompas and Long Chu. MEY for a short-lived species: a neural network approach. Fisheries Research, 201(??):138– 146, May 2018. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783618300201.

Kneebone:2013:PEC

[KCBS13] Jeff Kneebone, John Chisholm, Diego Bernal, and Gregory Skomal. The physiological effects of capture stress, recovery, and post-release survivorship of juvenile sand tigers (Carcharias taurus) caught on rod and reel. Fisheries Research, 147(??):103–114, October 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613001082.

Katopodis:2019:SSR

[KCJ19] Christos Katopodis, Lu Cai, and David Johnson. Sturgeon survival: the role of swimming performance and fish passage research. Fisheries Research, 212(??):162–171, April 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783618303837.

Kumar:2019:RSS

[KCM19] Rajeev Kumar, Noel G. Cadigan, and M. Joanne Morgan. Recruitment synchrony in spatially structured Newfoundland and Labrador populations of American plaice (Hippoglossoides platessoides). Fisheries Research, 211(??):91– 99, March 2019. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783618302960.

Kienzle:2014:EFE

[KCO14] Marco Kienzle, Anthony J. Courtney, and Michael F. O’Neill. Environmental and fishing effects on the dynamics of brown tiger prawn (Penaeus esculentus) in Moreton Bay (Australia). Fisheries Research, 155(??):138–148, July 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783614000769.

Kerby:2013:WAW

[KCvOE13] Tina K. Kerby, William W. L. Cheung, Cock van Oosterhout, and Georg H. Engelhard. Wondering about wandering whiting: Distribution of North Sea whiting between the 1920s and 2000s. Fisheries Research, 145(??):54–65, August 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783613000647.

Kliemann:2018:DCH

[KDdAA^{+}18] Bruna Caroline Kotz Kliemann, Rosilene Luciana Delariva, João Paulo de Arruda Amorim, Cristiéle da Silva Ribeiro, Bruno da Silva, Rosicleire Veríssimo Silveira, and Igor Paiva Ramos. Dietary changes and histophysiological responses of a wild fish (Geophagus cf. proximus) under the influence of tilapia cage farm. Fisheries Research, 204(??):337–347, August 2018. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783618300857.

Katikiro:2015:IBP

[KDM15] Robert Katikiro, K. H. M. Ashoka Deepananda, and Edison Macusi. Interplay between perceived changes in fishery and social structures in Tanzanian coastal fishing communities. Fisheries Research, 164(??):249–253, April 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614003658.

Korta:2010:RVR

[KDPMSR10] Maria Korta, Rosario Domínguez-Petit, Hilario Murua, and Francisco Saborido-Rey. Regional variability in reproductive traits of European hake Merluccius merluccius L. populations. Fisheries Research, 104(1–3):64–72, May 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609000733.

Keramidas:2018:ERF

[KDPT18] Ioannis Keramidas, Donna Dimarchopoulou, Androniki Pardalou, and Athanassios C. Tsikliras. Estimating recreational fishing fleet using satellite data in the Aegean and Ionian Seas (Mediterranean Sea). Fisheries Research, 208(??):1–6, December 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783618301917.

Karlsson-Drangsholt:2018:RES

[KDSAH^{+}18] Anders Karlsson-Drangsholt, Ragnhild Aven Svalheim, Øyvind Aas-Hansen, Stein-Harris Olsen, Kjell Midling, Michael Breen, Endre Grimsbø, and Helge Kreutzer Johnsen. Recovery from exhaustive swimming and its effect on fillet quality in haddock (Melanogrammus aeglefinus). Fisheries Research, 197(??):96–104, January 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783617302540.

Kolody:2016:MGT

[KEH16] Dale S. Kolody, J. Paige Eveson, and Richard M. Hillary. Modelling growth in tuna RFMO stock assessments: Current approaches and challenges. Fisheries Research, 180(??):177–193, August 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783615300035.

Kayanda:2012:TSM

[KEMM12] Robert Kayanda, Inigo Everson, Taabu Munyaho, and Yunus Mgaya. Target strength measurements of Nile perch (Lates niloticus: Linnaeus, 1758) in Lake Victoria, East Africa. Fisheries Research, 113(1):76–83, January 2012. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783611003110.

Kennedy:2014:CFF

[Ken14] James Kennedy. Comment on “Fishing and fecundity: the impact of exploitation on the reproductive potential of a deep-water fish, orange roughy (Hoplostethus atlanticus)”.

Fisheries Research, 155(??):194–195, July 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614000885.

Kolody:2019:RTR

[KEP^{+}19] Dale S. Kolody, J. Paige Eveson, Ann L. Preece, Campbell R. Davies, and Richard M. Hillary. Recruitment in tuna RFMO stock assessment and management: a review of current approaches and challenges. Fisheries Research, 217(??):217–234, September 2019. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618303412.

Kai:2018:SRR

[KF18] Mikihiko Kai and Yuki Fujinami. Stock-recruitment relationships in elasmobranchs: Application to the North Pacific blue shark. Fisheries Research, 200(??):104–115, April 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783617303004.

Konigson:2013:MGS

[KFBL13] Sara Königson, Arne Fjälling, Malin Berglind, and Sven-Gunnar Lunneryd. Male gray seals specialize in raiding salmon traps. Fisheries Research, 148(??):117–123, November 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783613001781.

Kjesbu:2010:SCP

[KFGN10] Olav Sigurd Kjesbu, Merete Fonn, Barbara Dunia Gonzáles, and Trygve Nilsen. Stereological calibration of the profile method to quickly estimate atresia levels in fish. Fisheries Research, 104(1–3):8–18, May 2010. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610000779.

Kim:2010:SPC

[KG10] Yonghae Kim and Malcolm S. Gordon. Swimming and posture control of common carp when penetrating mesh nets in a water tunnel. Fisheries Research, 102(1–2):166–172, February 2010. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783609003117.

Kuhnert:2011:APC

[KGB11] Petra M. Kuhnert, Shane Griffiths, and David Brewer. Assessing population changes in bycatch species using fishery-dependent catch rate data. Fisheries Research, 108(1):15–21, February 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783610002997.

Kubecka:2012:FSA

[KGH^{+}12] Jan Kubečka, Olav Rune Godø, Phil Hickley, Marie Prchalová, Milan Říha, Lars Rudstam, and Robin Welcomme. Fish sampling with active methods. Fisheries Research, 123–124(??):1–3, July 2012. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611003420.

Knotek:2015:DLC

[KGR^{+}15] Ryan J. Knotek, Shaun M. Gill, David B. Rudders, John W. Mandelman, Hugues P. Benoît, and James A. Sulikowski. The development of a low cost refrigerated flow-through seawater system for at-sea estimation of post-release mortality. Fisheries Research, 170(??):152–157, October 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783615001691.

Kronstadt:2018:PDM

[KGS18] Stephanie M. Kronstadt, Ryan Gandy, and Colin Shea. Predicting discard mortality in Florida stone crab, Menippe mercenaria, using reflexes. Fisheries Research, 197(??): 88–95, January 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783617302552.

Kolody:2015:ETM

[KH15] Dale Kolody and Simon Hoyle. Evaluation of tag mixing assumptions in western Pacific Ocean skipjack tuna stock assessment models. Fisheries Research, 163(??):127–140, March 2015. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783614001611.

King:2017:BRA

[KHG^{+}17] J. R. King, T. Helser, C. Gburski, D. A. Ebert, G. Cailliet, and C. R. Kastelle. Bomb radiocarbon analyses validate and inform age determination of longnose skate (Raja rhina) and big skate (Beringraja binoculata) in the north Pacific Ocean. Fisheries Research, 193(??):195–206, September 2017. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783617301091.

Kennedy:2014:ERP

[KHGB14] James Kennedy, Rasmus B. Hedeholm, Agnes C. Gundersen, and Jesper Boje. Estimates of reproductive potential of Greenland halibut (Reinhardtius hippoglossoides) in East Greenland based on an update of maturity status. Fisheries Research, 154(??):73–81, June 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614000496.

Krag:2015:SSD

[KHKM15] Ludvig A. Krag, Bent Herrmann, Junita D. Karlsen, and Bernd Mieske. Species selectivity in different sized topless trawl designs: Does size matter? Fisheries Research, 172(??):243–249, December 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615300199.

Krag:2010:SHM

[KHM^{+}10] Ludvig A. Krag, Rene Holst, Niels Madsen, Kurt Hansen, and Rikke P. Frandsen. Selective haddock (Melanogrammus aeglefinus) trawling: Avoiding cod (Gadus morhua) bycatch. Fisheries Research, 101(1–2):20–26, January 5, 2010.

CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609002446.

Kastelle:2017:AVP

[KHM^{+}17] Craig R. Kastelle, Thomas E. Helser, Jennifer L. McKay, Chris G. Johnston, Delsa M. Anderl, Mary E. Matta, and Daniel G. Nichol. Age validation of Pacific cod (Gadus macrocephalus) using high-resolution stable oxygen isotope (δ ^{18}O) chronologies in otoliths. Fisheries Research, 185(??):43–53, January 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616303174.

Krag:2011:SSH

[KHMF11] Ludvig A. Krag, Bent Herrmann, Niels Madsen, and Rikke P. Frandsen. Size selection of haddock (Melanogrammus aeglefinus) in square mesh codends: a study based on assessment of decisive morphology for mesh penetration. Fisheries Research, 110(2):225–235, July 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611001184.

Kusabs:2015:SMF

[KHQH15] Ian A. Kusabs, Brendan J. Hicks, John M. Quinn, and David P. Hamilton. Sustainable management of freshwater crayfish (kōura, Paranephrops planifrons) in Te Arawa (Rotorua) lakes, North Island, New Zealand. Fisheries Research, 168(??):35–46, August 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615001058.

Knip:2012:MRT

[KHS12] Danielle M. Knip, Michelle R. Heupel, and Colin A. Simpfendorfer. Mortality rates for two shark species occupying a shared coastal environment. Fisheries Research, 125–126(??):184–189, August 2012. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612001038.

Kraus:2012:REP

[KHV^{+}12] Gerd Kraus, Hans-Harald Hinrichsen, Rüdiger Voss, Eske Teschner, Jonna Tomkiewicz, and Friedrich W. Köster. Robustness of egg production methods as a fishery independent alternative to assess the Eastern Baltic cod stock (Gadus morhua callarias L.). Fisheries Research, 117–118(??):75– 85, April 2012. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783611000439.

Kim:2015:ERJ

[Kim15] Yonghae Kim. Escape response of juvenile fish in a shaking codend determined from tank experiments. Fisheries Research, 161(??):273–279, January 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614002574.

Kennedy:2017:DBI

[KJ17] James Kennedy and Sigurdur Tór Jónsson. Do biomass indices from Icelandic groundfish surveys reflect changes in the population of female lumpfish (Cyclopterus lumpus)? Fisheries Research, 194(??):22–30, October 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617301339.

Kim:2015:FMA

[KJKK15] Jin Yeong Kim, Hyeong Chul Jeong, Heeyong Kim, and Sukyung Kang. Forecasting the monthly abundance of anchovies in the South Sea of Korea using a univariate approach. Fisheries Research, 161(??):293–302, January 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783614002628.

Kemp:2012:RSR

[KJS12] Jodie Kemp, Gregory P. Jenkins, and Stephen E. Swearer. The reproductive strategy of red cod, Pseudophycis bachus, a key prey species for high trophic-level predators. Fisheries Research, 125–126(??):161–172, August 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612001014.

Kauppila:2012:PMA

[KK12] Pekka Kauppila and Timo P. Karjalainen. A process model to assess the regional economic impacts of fishing tourism: a case study in northern Finland. Fisheries Research, 127–128(??):88–97, September 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612001609.

Krafft:2015:AMA

[KK15] Bjørn A. Krafft and Ludvig A. Krag. Assessment of mortality of Antarctic krill (Euphausia superba) escaping from a trawl. Fisheries Research, 170(??):102–105, October 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578361500168X.

Kumar:2017:ANG

[KK17] Girish Kumar and Martin Kocour. Applications of nextgeneration sequencing in fisheries research: a review. Fisheries Research, 186 (part 1)(??):11–22, February 2017. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783616302338.

Kuparinen:2010:AFR

[KKA10] Anna Kuparinen, Thomas Klefoth, and Robert Arlinghaus. Abiotic and fishing-related correlates of angling catch rates in pike (Esox lucius). Fisheries Research, 105(2):111–117, July 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783610000755.

Kudo:2012:CEA

[KKE^{+}12] Hideaki Kudo, Yousuke Koshino, Akihiro Eto, Masaki Ichimura, and Masahide Kaeriyama. Cost-effective accurate estimates of adult chum salmon, Oncorhynchus keta, abundance in a Japanese river using a radio-controlled helicopter. Fisheries Research, 119–120(??):94–98, May 2012.

CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783611003882.

Kennington:2017:HIO

[KKH^{+}17] W. Jason Kennington, Peter W. Keron, Euan S. Harvey, Corey B. Wakefield, Ashley J. Williams, Tuikolongahau Halafihi, and Stephen J. Newman. High intra-ocean, but limited inter-ocean genetic connectivity in populations of the deep-water oblique-banded snapper Pristipomoides zonatus (Pisces: Lutjanidae). Fisheries Research, 193(??):242–249, September 2017. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783617301200.

Kell:2016:EPS

[KKK16] Laurence T. Kell, Ai Kimoto, and Toshihide Kitakado. Evaluation of the prediction skill of stock assessment using hindcasting. Fisheries Research, 183(??):119–127, November 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783616301540.

Kim:2016:MVP

[KKL^{+}16] Su Min Kim, Heeyong Kim, Won Chan Lee, Hyung Chul Kim, Howon Lee, Da Som Lee, Seok Nam Kwak, Naeun Jo, and Sang Heon Lee. Monthly variation in the proximate composition of jack mackerel (Trachurus japonicus) from Geumo Island, Korea. Fisheries Research, 183(??):371–378, November 2016. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783616302211.

Karelakis:2011:AMP

[KKP11] Christos Karelakis, Eleni Kaimakoudi, and Konstantinos Polymeros. Analysing managerial perceptions towards performance determinants in the fishery sector. Fisheries Research, 110(2):244–251, July 2011. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611001603.

Knoche:2016:DFR

[KL16] Scott Knoche and Frank Lupi. Demand for fishery regulations: Effects of angler heterogeneity and catch improvements on preferences for gear and harvest restrictions. Fisheries Research, 181(??):163–171, September 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616301230.

Kuehn:2017:CWB

[KLB17] Diane Kuehn, Valerie Luzadis, and Matthew Brincka. “I catch whatever’s biting!”: Motivations, constraints, and facilitators of no-preference anglers residing along New York’s Lake Ontario Coast. Fisheries Research, 194(??):188–196, October 2017. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783617301509.

Klemas:2013:FAR

[Kle13] Victor Klemas. Fisheries applications of remote sensing: an overview. Fisheries Research, 148(??):124–136, November 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783612001075.

Konigson:2015:SED

[KLH^{+}15] Sara Königson, Johan Lövgren, Joakim Hjelm, Mikael Ovegård, Fredrik Ljunghager, and Sven-Gunnar Lunneryd. Seal exclusion devices in cod pots prevent seal bycatch and affect their catchability of cod. Fisheries Research, 167(??):114–122, July 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S016578361500034X.

Kotwicki:2017:SRU

[KLWG17] Stan Kotwicki, Robert R. Lauth, Kresimir Williams, and Scott E. Goodman. Selectivity ratio: a useful tool for comparing size selectivity of multiple survey gears. Fisheries Research, 191(??):76–86, July 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617300498.

King:2010:MPG

[KM10] Jacquelynne R. King and Gordon A. McFarlane. Movement patterns and growth estimates of big skate (Raja binoculata) based on tag-recapture data. Fisheries Research, 101(1–2):50–59, January 5, 2010. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783609002495.

Kousteni:2015:ALH

[KM15] Vasiliki Kousteni and Persefoni Megalofonou. Aging and life history traits of the longnose spiny dogfish in the Mediterranean Sea: New insights into conservation and management needs. Fisheries Research, 168(??):6–19, August 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783615001046.

Kuguru:2018:DBC

[KMdS^{+}18] Gibbs Kuguru, Simo Maduna, Charlene da Silva, Enrico Gennari, Clint Rhode, and Aletta Bester van der Merwe. DNA barcoding of chondrichthyans in South African fisheries. Fisheries Research, 206(??):292–295, October 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783618301590.

Keller:2012:VAG

[KMH^{+}12] Aimee A. Keller, Kyle J. Molton, Allan C. Hicks, Melissa Haltuch, and Chantell Wetzel. Variation in age and growth of greenstriped rockfish (Sebastes elongatus) along the U.S. west coast (Washington to California). Fisheries Research, 119–120(??):80–88, May 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611003900.

King:2012:SDM

[KMJ^{+}12] Jacquelynne R. King, Gordon A. McFarlane, Simon R. M. Jones, Scott R. Gilmore, and Cathryn L. Abbott. Stock delineation of migratory and resident Pacific hake in Canadian waters. Fisheries Research, 114(??):19–30, February 2012.

CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783610003498.

Korta:2010:HOR

[KMKK10] Maria Korta, Hilario Murua, Yutaka Kurita, and Olav S. Kjesbu. How are the oocytes recruited in an indeterminate fish? Applications of stereological techniques along with advanced packing density theory on European hake (Merluccius merluccius L.). Fisheries Research, 104(1–3):56–63, May 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783610000147.

Kotwicki:2011:IAS

[KML11] Stan Kotwicki, Michael H. Martin, and Edward A. Laman. Improving area swept estimates from bottom trawl surveys. Fisheries Research, 110(1):198–206, June 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611001561.

Kronen:2010:RFP

[KMMV10] M. Kronen, F. Magron, B. McArdle, and A. Vunisea. Reef finfishing pressure risk model for Pacific Island countries and territories. Fisheries Research, 101(1–2):1–10, January 5, 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578360900232X.

Korta:2010:TDR

[KMQ^{+}10] Maria Korta, Hilario Murua, Iñaki Quincoces, Anders Thorsen, and Peter Witthames. Three-dimensional reconstruction of postovulatory follicles from histological sections. Fisheries Research, 104(1–3):38–44, May 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783609000988.

Kjesbu:2010:MDE

[KMSRW10] Olav Sigurd Kjesbu, Hilario Murua, Fran Saborido-Rey, and Peter R. Witthames. Method development and evaluation of stock reproductive potential of marine fish. Fisheries Research, 104(1–3):1–7, May 2010. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610000731.

Karnauskas:2011:ITS

[KMW^{+}11] Mandy Karnauskas, David B. McClellan, Jean W. Wiener, Margaret W. Miller, and Elizabeth A. Babcock. Inferring trends in a small-scale, data-limited tropical fishery based on fishery-independent data. Fisheries Research, 111(1–2):40–52, September 2011. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361100227X.

Kulatska:2019:UOT

[KNB^{+}19] Nataliia Kulatska, Stefan Neuenfeldt, Ulrika Beier, Bjarki Tór Elvarsson, Håkan Wennhage, Gunnar Stefansson, and Valerio Bartolino. Understanding ontogenetic and temporal variability of Eastern Baltic cod diet using a multispecies model and stomach data. Fisheries Research, 211(??):338– 349, March 2019. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783618303370.

Knights:2012:SVB

[Kni12] Antony M. Knights. Spatial variation in body size and reproductive condition of subtidal mussels: Considerations for sustainable management. Fisheries Research, 113(1):45–54, January 2012. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783611002931.

Kallio-Nyberg:2010:CLM

[KNJK^{+}10] Irma Kallio-Nyberg, Eero Jutila, Marja-Liisa Koljonen, Jarmo Koskiniemi, and Irma Saloniemi. Can the lost migratory Salmo trutta stocks be compensated with resident trout stocks in coastal rivers? Fisheries Research, 102(1–2):69–79, February 2010. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783609002756.

Kallio-Nyberg:2015:DBW

[KNRJ^{+}15] Irma Kallio-Nyberg, Atso Romakkaniemi, Erkki Jokikokko, Irma Saloniemi, and Eero Jutila. Differences between wild and reared Salmo salar stocks of two northern Baltic Sea rivers. Fisheries Research, 165(??):85–95, May 2015. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783615000107.

Kallio-Nyberg:2011:EHR

[KNSJJ11] Irma Kallio-Nyberg, Irma Saloniemi, Eero Jutila, and Erkki Jokikokko. Effect of hatchery rearing and environmental factors on the survival, growth and migration of Atlantic salmon in the Baltic Sea. Fisheries Research, 109(2–3):285– 294, May 2011. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783611000865.

Kallio-Nyberg:2013:CBA

[KNSPK13] I. Kallio-Nyberg, M. Salminen, T. Pakarinen, and M.- L. Koljonen. Cost–benefit analysis of Atlantic salmon smolt releases in relation to life-history variation. Fisheries Research, 145(??):6–14, August 2013. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613000726.

Kallio-Nyberg:2011:EMS

[KNSSL11] I. Kallio-Nyberg, M. Salminen, I. Saloniemi, and M. Lindroos. Effects of marine survival, precocity and other life history traits on the cost-benefit of stocking salmon in the Baltic Sea. Fisheries Research, 110(1):111–119, June 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783611001287.

Klaer:2015:HME

[KOD^{+}15] Neil L. Klaer, Robert N. O’Boyle, Jonathan J. Deroba, Sally E. Wayte, L. Richard Little, Larry A. Alade, and Paul J. Rago. How much evidence is required for acceptance of productivity regime shifts in fish stock assessments: Are we letting managers off the hook? Fisheries Research, 168(??):49–55, August 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615001113.

Kynoch:2011:TMN

[KOF11] R. J. Kynoch, F. G. O’Neill, and R. J. Fryer. Test of 300 and 600 mm netting in the forward sections of a Scottish whitefish trawl. Fisheries Research, 108(2–3):277–282, March 2011. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783610003449.

Kuriyama:2016:EWA

[KOHF^{+}16] Peter T. Kuriyama, Kotaro Ono, Felipe Hurtado-Ferro, Allan C. Hicks, Ian G. Taylor, Roberto R. Licandeo, Kelli F. Johnson, Sean C. Anderson, Cole C. Monnahan, Merrill B. Rudd, Christine C. Stawitz, and Juan L. Valero. An empirical weight-at-age approach reduces estimation bias compared to modeling parametric growth in integrated, statistical stock assessment models when growth is time varying. Fisheries Research, 180(??):119–127, August 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615300837.

Kooka:2012:LHT

[Koo12] Kouji Kooka. Life-history traits of walleye pollock, Theragra chalcogramma, in the northeastern Japan Sea during early to mid 1990s. Fisheries Research, 113(1):35–44, January 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578361100292X.

Kosamu:2015:CSS

[Kos15] Ishmael B. M. Kosamu. Conditions for sustainability of small-scale fisheries in developing countries. Fisheries Research, 161(??):365–373, January 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614002690.

Kullmann:2018:ABS

[KPF^{+}18] Björn Kullmann, Jan-Dag Pohlmann, Marko Freese, Alexander Keth, Laura Wichmann, Rüdiger Neukamm, and Ralf Thiel. Age-based stock assessment of the European eel (Anguilla anguilla) is heavily biased by stocking of unmarked farmed eels. Fisheries Research, 208(??):258–266, December 2018. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783618302303.

Kulmala:2007:ITQ

[KPL^{+}07] S. Kulmala, H. Peltomäki, M. Lindroos, P. Söderkultalahti, and S. Kuikka. Individual transferable quotas in the Baltic Sea herring fishery: a socio-bioeconomic analysis. Fisheries Research, 84(3):368–377, May 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783606004267. See erratum [KPL^{+}09].

Kulmala:2009:EIT

[KPL^{+}09] S. Kulmala, H. Peltomäki, M. Lindroos, P. Söderkultalahtir, and S. Kuikka. Erratum to “Individual transferable quotas in the Baltic Sea herring fishery: a socio-bioeconomic analysis” [Fish. Res. 84 (2007) 368–377]. Fisheries Research, 100(2):183, October 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783609001866. See [KPL^{+}07].

Kalogirou:2019:STS

[KPM^{+}19] S. Kalogirou, L. Pihl, C. D. Maravelias, B. Herrmann, C. J. Smith, N. Papadopoulou, E. Notti, and A. Sala. Shrimp trap selectivity in a Mediterranean small-scale-fishery. Fisheries Research, 211(??):131–140, March 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618303187.

Knickle:2012:AMA

[KR12] D. Craig Knickle and George A. Rose. Acoustic markers of Atlantic cod (Gadus morhua) spawning in coastal Newfoundland. Fisheries Research, 129–130(??):8–16, October 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783612001695.

Kraak:2015:RRT

[KRB^{+}15] Sarah B. M. Kraak, David G. Reid, Guillaume Bal, Amos Barkai, Edward A. Codling, Ciarán J. Kelly, and Emer Rogan. RTI (“Real-Time Incentives”) outperforms traditional management in a simulated mixed fishery and cases incorporating protection of vulnerable species and areas. Fisheries Research, 172(??):209–224, December 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615300230.

Kunkel:2017:XTA

[KRB17] Joseph G. Kunkel, Melissa Rosa, and Ali N. Bahadur. 3Dxray-tomography of American lobster shell-structure. An overview. Fisheries Research, 186 (part 1)(??):372–382, February 2017. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783616303216.

Kraak:2014:ERR

[KRC14] Sarah B. M. Kraak, Dave G. Reid, and Edward A. Codling. Exploring the RTI (real-time incentive) tariffbased approach to single-species fisheries management. Fisheries Research, 155(??):90–102, July 2014. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361400054X.

Kristjansson:2013:CGA

[Kri13] Theódór Kristjánsson. Comparison of growth in Atlantic cod (Gadus morhua) originating from the northern and southern coast of Iceland reared under common conditions. Fisheries Research, 139(??):105–109, March 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612003013.

Kristianson:2018:LFP

[Kri18] Gerry Kristianson. Lessons in fisheries politics. Fisheries Research, 205(??):32–34, September 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618300997.

Knotek:2018:SRD

[KRM^{+}18] Ryan J. Knotek, David B. Rudders, John W. Mandelman, Hugues P. Benoît, and James A. Sulikowski. The survival of rajids discarded in the New England scallop dredge fisheries. Fisheries Research, 198(??):50–62, February 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617302904.

Kenney:2015:BPC

[KRMW15] Jessica L. Kenney, Taufiqur Rahman, Heather Manuel, and Paul D. Winger. Bruising patterns in commercially harvested yellowtail flounder (Limanda ferruginea). Fisheries Research, 172(??):79–84, December 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615300072.

Kloser:2016:IMA

[KRTG16] Rudy J. Kloser, Tim E. Ryan, Geoffrey N. Tuck, and Gerry Geen. Influence on management advice of fishers acoustics — 10 year review of blue grenadier monitoring. Fisheries Research, 178(??):82–92, June 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615300916.

Kruschel:2012:ULV

[KS12] Claudia Kruschel and Stewart T. Schultz. Use of a lure in visual census significantly improves probability of detecting wait-ambushing and fast cruising predatory fish. Fisheries Research, 123–124(??):70–77, July 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611003444.

Katara:2017:MBV

[KS17] I. Katara and A. Silva. Mismatch between VMS data temporal resolution and fishing activity time scales. Fisheries Research, 188(??):1–5, April 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616304088.

Klefoth:2013:RNL

[KSAA13] Thomas Klefoth, Christian Skov, Kim Aarestrup, and Robert Arlinghaus. Reliability of non-lethal assessment methods of body composition and energetic status exemplified by applications to eel (Anguilla anguilla) and carp (Cyprinus carpio). Fisheries Research, 146(??):18–26, September 2013. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783613000787.

Krieger:2019:TDG

[KSH19] Joseph R. Krieger, Ashwin Sreenivasan, and Ron Heintz. Temperature-dependent growth and consumption of youngof-the-year sablefish Anoplopoma fimbria: Too hot, too cold or just right? Fisheries Research, 209(??):32–39, January 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783618302406.

Kloser:2015:IRO

[KSKGR15] Rudy J. Kloser, Caroline Sutton, Kyne Krusic-Golub, and Tim E. Ryan. Indicators of recovery for orange roughy (Hoplostethus atlanticus) in eastern Australian waters fished from 1987. Fisheries Research, 167(??):225–235, July 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783615000880.

Keller:2016:MBB

[KSL^{+}16] Krystle Keller, Aldo S. Steffe, Michael Lowry, Jeffrey J. Murphy, and Iain M. Suthers. Monitoring boat-based recreational fishing effort at a nearshore artificial reef with a shore-based camera. Fisheries Research, 181(??):84–92, September 2016. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783616300935.

Keller:2017:ERH

[KSL^{+}17] Krystle Keller, Aldo S. Steffe, Michael B. Lowry, Jeffrey J. Murphy, James A. Smith, and Iain M. Suthers. Estimating the recreational harvest of fish from a nearshore designed artificial reef using a pragmatic approach. Fisheries Research, 187(??):158–167, March 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616304076.

Kinds:2016:VIA

[KSS^{+}16] Arne Kinds, Kim Sys, Laura Schotte, Koen Mondelaers, and Hans Polet. VALDUVIS: an innovative approach to assess the sustainability of fishing activities. Fisheries Research, 182(??):158–171, October 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615301235.

Kawamata:2018:NSM

[KT18a] Shigeru Kawamata and Seiya Taino. A novel in situ method for estimation of the carapace length of sheltering spiny lobsters, Panulirus japonicus, via stereo photography. Fisheries Research, 201(??):56–67, May 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618300158.

Kullmann:2018:BBE

[KT18b] Björn Kullmann and Ralf Thiel. Bigger is better in eel stocking measures? Comparison of growth performance, body condition, and benefit-cost ratio of simultaneously stocked glass and farmed eels in a brackish fjord. Fisheries Research, 205(??):132–140, September 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618301085.

Kevrekidis:2011:PDM

[KTL11] Kosmas Kevrekidis and Maria Thessalou-Legaki. Population dynamics of Melicertus kerathurus (Decapoda: Penaeidae) in Thermaikos Gulf (N. Aegean Sea). Fisheries Research, 107(1–3):46–58, January 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610002547.

Kokkalis:2015:LRS

[KTNA15] Alexandros Kokkalis, Uffe H. Thygesen, Anders Nielsen, and Ken H. Andersen. Limits to the reliability of sizebased fishing status estimation for data-poor stocks. Fisheries Research, 171(??):4–11, November 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614003002.

Koster:2013:LSA

[KTT13] Friedrich W. Köster, Edward A. Trippel, and Jonna Tomkiewicz. Linking size and age at sexual maturation to body growth, productivity and recruitment of Atlantic cod stocks spanning the North Atlantic. Fisheries Research, 138(??):52–61, February 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361200207X.

Kruck:2013:MGB

[KTW^{+}13] Nils C. Krück, Ian R. Tibbetts, Robert D. Ward, Jeffrey W. Johnson, William K. W. Loh, and Jennifer R. Ovenden. Multi-gene barcoding to discriminate sibling species within a morphologically difficult fish genus (Sillago). Fisheries Research, 143(??):39–46, June 2013. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783613000106.

Kurita:2012:RCE

[Kur12] Yutaka Kurita. Revised concepts for estimation of spawning fraction in multiple batch spawning fish considering temperature-dependent duration of spawning markers and spawning time frequency distribution. Fisheries Research, 117–118(??):121–129, April 2012. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611001998.

Kokkonen:2015:PMR

[KVH15] Eevi Kokkonen, Anssi Vainikka, and Outi Heikinheimo. Probabilistic maturation reaction norm trends reveal decreased size and age at maturation in an intensively harvested stock of pikeperch Sander lucioperca. Fisheries Research, 167(??):1–12, July 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615000284.

Kraus:2017:RCD

[KVK^{+}17] Richard T. Kraus, Christopher S. Vandergoot, Patrick M. Kocovsky, Mark W. Rogers, H. Andrew Cook, and Travis O. Brenden. Reconciling catch differences from multiple fishery independent gill net surveys. Fisheries Research, 188(??):17–22, April 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S016578361630412X.

Kolding:2012:RLL

[KvZ12] Jeppe Kolding and Paul A. M. van Zwieten. Relative lake level fluctuations and their influence on productivity and resilience in tropical lakes and reservoirs. Fisheries Research, 115–116(??):99–109, March 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611003377.

Kim:2010:ASN

[KW10] Yong-Hae Kim and Dae-Sung Whang. An actively stimulating net panel and rope array inside a model codend to increase juvenile red seabream escapement. Fisheries Research, 106(1):71–75, October 2010. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610001608.

Kitson-Walters:2018:FSP

[KWCT^{+}18] Kimani A. Kitson-Walters, Adam S. Candy, Nathan K. Truelove, Marcia E. Roye, Mona K. Webber, Karl A. Aiken, and Stephen J. Box. Fine-scale population structure of Lobatus gigas in Jamaica’s exclusive economic zone considering hydrodynamic influences. Fisheries Research, 199(??):53–62, March 2018. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783617303181.

Klaer:2012:EPH

[KWF12] Neil L. Klaer, Sally E. Wayte, and Gavin Fay. An evaluation of the performance of a harvest strategy that uses an average-length-based assessment method. Fisheries Research, 134–136(??):42–51, December 2012. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612002561.

Kayaba:2015:NMI

[KWM^{+}15] Takaaki Kayaba, Toshihiro Wada, Osamu Murakami, Sayumi Sawaguchi, and Ryo Kawabe. A new method for identifying the male spawning period of fish using sperm duct volume as an index. Fisheries Research, 170(??):199–204, October 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783615001897.

Kinzey:2015:STB

[KWR15] Douglas Kinzey, George M. Watters, and Christian S. Reiss. Selectivity and two biomass measures in an agebased assessment of Antarctic krill (Euphausia superba). Fisheries Research, 168(??):72–84, August 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361500123X.

Kinzey:2019:ERV

[KWR19] Douglas Kinzey, George M. Watters, and Christian S. Reiss. Estimating recruitment variability and productivity in Antarctic krill. Fisheries Research, 217(??):98–107, September 2019. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783618302686.

King:2015:GAS

[KWS^{+}15] J. R. King, M. Wetklo, J. Supernault, M. Taguchi, K. Yokawa, O. Sosa-Nishizaki, and R. E. Withler. Genetic analysis of stock structure of blue shark (Prionace glauca) in the north Pacific Ocean. Fisheries Research, 172(??):181–189, December 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615300023.

Korman:2017:EEC

[KY17] Josh Korman and Michael D. Yard. Effects of environmental covariates and density on the catchability of fish populations and interpretation of catch per unit effort trends. Fisheries Research, 189(??):18–34, May 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361730005X.

Kai:2019:PEI

[KY19a] Mikihiko Kai and Hiroki Yokoi. Performance evaluation of information criteria for estimating a shape parameter in a Bayesian state-space biomass dynamics model. Fisheries Research, 219(??):Article 105326, November 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361930181X.

Kaifu:2019:IDC

[KY19b] Kenzo Kaifu and Kazuki Yokouchi. Increasing or decreasing? — current status of the Japanese eel stock. Fisheries Research, 220(??):Article 105348, December 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619302036.

Kurosaka:2012:TBM

[KYO^{+}12] Kohei Kurosaka, Hideyuki Yamashita, Michio Ogawa, Hiroshi Inada, and Takafumi Arimoto. Tentacle-breakage mechanism for the neon flying squid Ommastrephes bartramii during the jigging capture process. Fisheries Research, 121–122(??):9–16, June 2012. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612000112.

Kim:2011:DSI

[KZ11] Do Hoon Kim and Chang Ik Zhang. Developing socioeconomic indicators for an ecosystem-based fisheries management approach: an application to the Korean large purse seine fishery. Fisheries Research, 112(3):134–139, December 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783611000567.

Ley:2013:MMP

[LA13] J. A. Ley and M. S. Allen. Modeling marine protected area value in a catch-and-release dominated estuarine fishery. Fisheries Research, 144(??):60–73, July 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612003049.

Laptikhovsky:2010:SRS

[LAB10] Vladimir Laptikhovsky, Alexander Arkhipkin, and Paul Brickle. Squid as a resource shared by fish and humans on the Falkland Islands’ shelf. Fisheries Research, 106(2):151–155, November 2010. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610001050.

Laptikhovsky:2013:SBM

[LAB13] Vladimir Laptikhovsky, Alexander Arkhipkin, and Paul Brickle. From small bycatch to main commercial species: Explosion of stocks of rock cod Patagonotothen ramsayi (Regan) in the Southwest Atlantic. Fisheries Research, 147(??):399–403, October 2013. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613001471.

Laurrabaquio-A:2019:GER

[LAIVA^{+}19] Nadia Sandoval Laurrabaquio-A., Valentina Islas-Villanueva, Douglas H. Adams, Manuel Uribe-Alcocer, Jaime R. Alvarado-Bremer, and Píndaro Díaz-Jaimes. Genetic evidence for regional philopatry of the bull shark (Carcharhinus leucas), to nursery areas in estuaries of the Gulf of Mexico and western North Atlantic Ocean. Fisheries Research, 209(??):67–74, January 2019. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783618302480.

Lambert:2013:LTC

[Lam13] Yvan Lambert. Long-term changes in life history characteristics and reproductive potential of northern Gulf of St. Lawrence cod (Gadus morhua) and consequences for the stock productivity. Fisheries Research, 138(??):5–13, February 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783612001774.

Leland:2013:DPS

[LBB^{+}13a] Jesse C. Leland, Paul A. Butcher, Matt K. Broadhurst, Brian D. Paterson, and David G. Mayer. Damage and physiological stress to juvenile eastern rock lobster (Sagmariasus verreauxi) discarded after trapping and hand collection. Fisheries Research, 137(??):63–70, January 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612002706.

Leland:2013:RTE

[LBB^{+}13b] Jesse C. Leland, Paul A. Butcher, Matt K. Broadhurst, Brian D. Paterson, and David G. Mayer. Relative trap efficiency for recreationally caught eastern Australian blue swimmer crab (Portunus pelagicus) and associated injury and mortality of discards. Fisheries Research, 147(??):304–311, October 2013. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783613001707.

Lennox:2018:PRB

[LBCD18] Robert J. Lennox, Jacob W. Brownscombe, Steven J. Cooke, and Andy J. Danylchuk. Post-release behaviour and survival of recreationally-angled arapaima (Arapaima cf. arapaima) assessed with accelerometer biologgers. Fisheries Research, 207(??):197–203, November 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618301437.

Laptikhovsky:2016:NAE

[LBF^{+}16] V. Laptikhovsky, C. Barrett, C. Firmin, P. Hollyman, A. Lawler, R. Masefield, R. McIntyre, D. Palmer, M. Soeffker, and M. Parker-Humphreys. A novel approach for estimation of the natural mortality of the common whelk, Buccinum undatum (L.) and role of hermit crabs in its shell turnover. Fisheries Research, 183(??):146–154, November 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578361630176X.

LeBlanc:2015:BSA

[LBH15] Stéphan N. LeBlanc, Hugues P. Benoît, and Heather L. Hunt. Broad-scale abundance changes are more prevalent than acute fishing impacts in an experimental study of scallop dredging intensity. Fisheries Research, 161(??):8–20, January 2015. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783614001891.

Lallemand:2016:EEB

[LBHP16] Philippe Lallemand, Mike Bergh, Margaret Hansen, and Martin Purves. Estimating the economic benefits of MSC certification for the South African hake trawl fishery. Fisheries Research, 182(??):98–115, October 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361630025X.

Lu:2015:ASI

[LBMC15] Ching-Ping Lu, Jaime R. Alvarado Bremer, Jessica L. McKenzie, and Wei-Chuan Chiang. Analysis of sailfish (Istiophorus platypterus) population structure in the North Pacific Ocean. Fisheries Research, 166(??):33–38, June 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783614002860.

Lefebvre:2018:DNP

[LBS^{+}18] Lyndsey S. Lefebvre, Sabrina G. Beyer, David M. Stafford, Neosha S. Kashef, Edward J. Dick, Susan M. Sogard, and John C. Field. Double or nothing: Plasticity in reproductive output in the chilipepper rockfish (Sebastes goodei). Fisheries Research, 204(??):258–268, August 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618300766.

Landa:2013:AGA

[LBV13] Jorge Landa, Joaquín Barrado, and Francisco Velasco. Age and growth of anglerfish (Lophius piscatorius) on the Porcupine Bank (west of Ireland) based on illicia age estimation. Fisheries Research, 137(??):30–40, January 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612002639.

Leitao:2014:RCT

[LBZE14] F. Leitão, V. Baptista, D. Zeller, and K. Erzini. Reconstructed catches and trends for mainland Portugal fisheries between 1938 and 2009: implications for sustainability, domestic fish supply and imports. Fisheries Research, 155(??): 33–50, July 2014. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783614000526.

Li:2017:WDL

[LBZE17] Yang Li, James R. Bence, Zhen Zhang, and Mark P. Ebener. Why do lake whitefish move long distances in Lake Huron? Bayesian variable selection of factors explaining fish movement distance. Fisheries Research, 195(??):169–179, November 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783617301960.

Lyach:2018:NTC

[LČ18] Roman Lyach and Martin Čech. A new trend in Central European recreational fishing: More fishing visits but lower yield and catch. Fisheries Research, 201(??):131–137, May 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783618300262.

Lorenzen:2019:DDL

[LC19] Kai Lorenzen and Edward V. Camp. Density-dependence in the life history of fishes: When is a fish recruited? Fisheries Research, 217(??):5–10, September 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618302650.

Lee:2019:AGM

[LCA^{+}19] Brendon Lee, Kate Cockroft, Alexander I. Arkhipkin, Stephen R. Wing, and Haseeb S. Randhawa. Age, growth and mortality estimates for the ridge-scaled grenadier Macrourus carinatus (Günther, 1878) in the south-western Atlantic. Fisheries Research, 218(??):174–185, October 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783619301286.

Llompart:2017:CAR

[LCB17] F. M. Llompart, D. C. Colautti, and C. R. M. Baigún. Conciliating artisanal and recreational fisheries in anegada Bay, Argentina. Fisheries Research, 190(??):140–149, June 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783617300206.

Larocque:2012:MBF

[LCBD12] Sarah M. Larocque, Steven J. Cooke, and Gabriel Blouin-Demers. Mitigating bycatch of freshwater turtles in passively fished fyke nets through the use of exclusion and escape modifications. Fisheries Research, 125–126(??):149–155, August 2012. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783612000987.

Lundin:2015:SSW

[LCF15] Mikael Lundin, Linda Calamnius, and Arne Fjälling. Size selection of whitefish (Coregonus maraena) in a pontoon trap equipped with an encircling square mesh selection panel. Fisheries Research, 161(??):330–335, January 2015. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783614002653.

Liu:2018:UMD

[LCH^{+}18] Qi Liu, Fan Cui, Pengfei Hu, Gaotong Yi, Yanwei Ge, Wenlei Liu, Hongwei Yan, Lianshun Wang, Haiying Liu, Jun Song, Yusheng Jiang, Lei Zhang, and Zhong Tu. Using of microsatellite DNA profiling to identify hatchery-reared seed and assess potential genetic risks associated with large-scale release of swimming crab Portunus trituberculatus in Panjin, China. Fisheries Research, 207(??):187–196, November 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783618301395.

Lundin:2011:SSH

[LCHL11] Mikael Lundin, Linda Calamnius, Lars Hillström, and Sven-Gunnar Lunneryd. Size selection of herring (Clupea harengus membras) in a pontoon trap equipped with a rigid grid. Fisheries Research, 108(1):81–87, February 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610003139.

Lundin:2012:SJH

[LCL12] Mikael Lundin, Linda Calamnius, and Sven-Gunnar Lunneryd. Survival of juvenile herring (Clupea harengas membras) after passing through a selection grid in a pontoon trap. Fisheries Research, 127–128(??):83–87, September 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783612001658.

Lundin:2015:ESG

[LCLM15] Mikael Lundin, Linda Calamnius, Sven-Gunnar Lunneryd, and Carin Magnhagen. The efficiency of selection grids in perch pontoon traps. Fisheries Research, 162(??):58–63, February 2015. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783614002859.

Levsen:2018:ASC

[LCM^{+}18] Arne Levsen, Paolo Cipriani, Simonetta Mattiucci, Melanie Gay, Lee C. Hastie, Ken MacKenzie, Graham J. Pierce, Cecilie S. Svanevik, Danjal P. Højgaard, Giuseppe Nascetti, Angel F. González, and Santiago Pascual. Anisakis species composition and infection characteristics in Atlantic mackerel, Scomber scombrus, from major European fishing grounds — reflecting changing fish host distribution and migration pattern. Fisheries Research, 202(??):112–121, June 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783617302138.

Lennert-Cody:2013:DPS

[LCMdSM13] Cleridy E. Lennert-Cody, Mark N. Maunder, Alexandre Aires da Silva, and Mihoko Minami. Defining population spatial units: Simultaneous analysis of frequency distributions and time series. Fisheries Research, 139(??):85–92, March 2013. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783612002974.

Lennert-Cody:2016:PSV

[LCMF^{+}16] Cleridy E. Lennert-Cody, Mark N. Maunder, Paul C. Fiedler, Mihoko Minami, Tim Gerrodette, Jeremy Rusin, Carolina V. Minte-Vera, Michael Scott, and Stephen T. Buckland. Purse-seine vessels as platforms for monitoring the population status of dolphin species in the eastern tropical Pacific Ocean. Fisheries Research, 178(??):101–113, June 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783615301028.

Lennert-Cody:2010:EAS

[LCMTM10] Cleridy E. Lennert-Cody, Mihoko Minami, Patrick K. Tomlinson, and Mark N. Maunder. Exploratory analysis of spatial–temporal patterns in length–frequency data: an example of distributional regression trees. Fisheries Research, 102(3):323–326, March 2010. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783609003166.

Li:2019:SBE

[LCRW19] Bai Li, Yong Chen, Kathleen Reardon, and Carl Wilson. A simulation-based evaluation of sea sampling design for American lobster in the Gulf of Maine. Fisheries Research, 216(??):145–154, August 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619301092.

Lin:2010:EJE

[LCST10] Yu-Jia Lin, Yi-Jay Chang, Chi-Lu Sun, and Wann-Nian Tzeng. Evaluation of the Japanese eel fishery in the lower reaches of the Kao-Ping River, southwestern Taiwan using a per-recruit analysis. Fisheries Research, 106(3):329–336, December 2010. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783610002080.

Lin:2017:EAS

[LCWC17] Dongming Lin, Xinjun Chen, Yanran Wei, and Yong Chen. The energy accumulation of somatic tissue and reproductive organs in post-recruit female Illex argentinus and the relationship with sea surface oceanography. Fisheries Research, 185(??):102–114, January 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616303162.

Li:2016:MHS

[LCZ^{+}16] Gang Li, Jie Cao, Xiaorong Zou, Xinjun Chen, and Jocelyn Runnebaum. Modeling habitat suitability index for Chilean jack mackerel (Trachurus murphyi) in the South East Pacific. Fisheries Research, 178(??):47–60, June 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615301351.

Liu:2019:HAP

[LCZ^{+}19] Chang Liu, Geoffrey W. Cowles, Douglas R. Zemeckis, Gavin Fay, Arnault Le Bris, and Steven X. Cadrin. A hardware-accelerated particle filter for the geolocation of demersal fishes. Fisheries Research, 213(??):160–171, May 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783619300190.

Li:2015:TDT

[LDM^{+}15] Xinya Li, Zhiqun D. Deng, Jayson J. Martinez, Tao Fu, P. Scott Titzler, James S. Hughes, Mark A. Weiland, Richard S. Brown, Bradly A. Trumbo, Martin L. Ahmann, and Jon F. Renholds. Three-dimensional tracking of juvenile salmon at a mid-reach location between two dams. Fisheries Research, 167(??):216–224, July 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615000892.

Lambooij:2012:EBS

[LDR^{+}12] E. Lambooij, H. Digre, H. G. M. Reimert, I. G. Aursand, L. Grimsmo, and J. W. van de Vis. Effects of on-board storage and electrical stunning of wild cod (Gadus morhua) and haddock (Melanogrammus aeglefinus) on brain and heart activity. Fisheries Research, 127–128(??):1–8, September 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783612001324.

Lin:2019:ESF

[LDS^{+}19] Chenyu Lin, Huichao Dai, Xiaotao Shi, Zhiqun Daniel Deng, Jingqiao Mao, Shoujiang Zhao, Jia Luo, and Junjun Tan. An experimental study on fish attraction using a fish barge model. Fisheries Research, 210(??):181–188, February 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578361830273X.

Larsen:2014:AAR

[LE14] Finn Larsen and Ole R. Eigaard. Acoustic alarms reduce bycatch of harbour porpoises in Danish North Sea gillnet fisheries. Fisheries Research, 153(??):108–112, May 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783614000186.

Legault:2014:ATA

[Leg14] Christopher M. Legault. The ability of two age composition error distributions to estimate selectivity and spawning stock biomass in simulated stock assessments. Fisheries Research, 158(??):172–180, October 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613003019.

Langangen:2011:SDP

[LEO^{+}11] Øystein Langangen, Eric Edeline, Jan Ohlberger, Ian J. Winfield, Janice M. Fletcher, J. Ben James, Nils Chr. Stenseth, and L. Asbjørn Vøllestad. Six decades of pike and perch population dynamics in Windermere. Fisheries Research, 109(1):131–139, April 2011. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361100049X.

Lucey:2013:OFN

[LF13] Sean M. Lucey and Michael J. Fogarty. Operational fisheries in New England: Linking current fishing patterns to proposed ecological production units. Fisheries Research, 141(??):3–12, April 2013. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612001580.

Liu:2015:SVB

[LFCC15] Bilin Liu, Zhou Fang, Xinjun Chen, and Yong Chen. Spatial variations in beak structure to identify potentially geographic populations of Dosidicus gigas in the Eastern Pacific Ocean. Fisheries Research, 164(??):185–192, April 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783614003361.

Lowry:2011:EUS

[LFG11] Michael Lowry, Heath Folpp, and Marcus Gregson. Evaluation of an underwater solid state memory video system with application to fish abundance and diversity studies in southeast Australia. Fisheries Research, 110(1):10–17, June 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783611001123.

Lopez:2019:EPB

[LFS^{+}19] Jon Lopez, Jose Mari Ferarios, Josu Santiago, Macarena Ubis, Gala Moreno, and Hilario Murua. Evaluating potential biodegradable twines for use in the tropical tuna FAD fishery. Fisheries Research, 219(??):Article 105321, November 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783619301687.

Leite:2013:MAF

[LG13] Marta C. F. Leite and Maria A. Gasalla. A method for assessing fishers’ ecological knowledge as a practical tool for ecosystem-based fisheries management: Seeking consensus in Southeastern Brazil. Fisheries Research, 145(??):43–53, August 2013. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783613000659.

Linnane:2021:ELS

[LGH^{+}21] Adrian Linnane, Caleb Gardner, David Hobday, André Punt, Richard McGarvey, John Feenstra, Janet Matthews, and Bridget Green. Evidence of large-scale spatial declines in recruitment patterns of southern rock lobster Jasus edwardsii, across south-eastern Australia. Fisheries Research, 105(3):163–171, ???? 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610000871.

Li:2017:IAV

[LGZC17] Yunkai Li, Yi Gong, Yuying Zhang, and Xinjun Chen. Inter-annual variability in trophic patterns of jumbo squid (Dosidicus gigas) off the exclusive economic zone of Peru, implications from stable isotope values in gladius. Fisheries Research, 187(??):22–30, March 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616303708.

Leporati:2015:SWP

[LH15] Stephen. C. Leporati and Anthony. M. Hart. Stylet weight as a proxy for age in a merobenthic octopus population. Fisheries Research, 161(??):235–243, January 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614002380.

Leigh:2018:NTD

[LH18] George M. Leigh and William S. Hearn. Nonparametric tests of double-tagging assumptions. Fisheries Research, 197(??):45–49, January 2018. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617302631.

Lovgren:2016:BSS

[LHF16] Johan Lövgren, Bent Herrmann, and Jordan Feekings. Bellshaped size selection in a bottom trawl: a case study for Nephrops directed fishery with reduced catches of cod. Fisheries Research, 184(??):26–35, December 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616300881.

Leporati:2015:OLH

[LHL^{+}15] S. C. Leporati, A. M. Hart, R. Larsen, L. E. Franken, and M. De Graaf. Octopus life history relative to age, in a multigeared developmental fishery. Fisheries Research, 165(??): 28–41, May 2015. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783615000053.

Lehikoinen:2017:RCF

[LHLR17] Aleksi Lehikoinen, Outi Heikinheimo, Hannu Lehtonen, and Pekka Rusanen. The role of cormorants, fishing effort and temperature on the catches per unit effort of fisheries in Finnish coastal areas. Fisheries Research, 190(??):175–182, June 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783617300450.

Larsen:2016:SSR

[LHS^{+}16] Roger B. Larsen, Bent Herrmann, Manu Sistiaga, Eduardo Grimaldo, Ivan Tatone, and Iñigo Onandia. Size selection of redfish (Sebastes spp.) in a double grid system: Estimating escapement through individual grids and comparison to former grid trials. Fisheries Research, 183(??):385–395, November 2016. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783616302259.

Larsen:2018:CGA

[LHS^{+}18a] Roger B. Larsen, Bent Herrmann, Manu Sistiaga, Jure Brčić, Jesse Brinkhof, and Ivan Tatone. Could green artificial light reduce bycatch during Barents Sea deep-water shrimp trawling? Fisheries Research, 204(??):441–447, August 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783618300973.

Larsen:2018:BRN

[LHS^{+}18b] Roger B. Larsen, Bent Herrmann, Manu Sistiaga, Jesse Brinkhof, and Eduardo Grimaldo. Bycatch reduction in the Norwegian deep-water shrimp (Pandalus borealis) fishery with a double grid selection system. Fisheries Research, 208(??):267–273, December 2018. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618302297.

Larsen:2018:SSC

[LHS^{+}18c] Roger B. Larsen, Bent Herrmann, Manu Sistiaga, Eduardo Grimaldo, Ivan Tatone, and Jesse Brinkhof. Size selection of cod (Gadus morhua) and haddock (Melanogrammus aeglefinus) in the Northeast Atlantic bottom trawl fishery with a newly developed double steel grid system. Fisheries Research, 201(??):120–130, May 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618300286.

Leavitt:2018:SDA

[LHSS18] Jasper S. Leavitt, Carl J. Huntsberger, Ronald J. Smolowitz, and Liese A. Siemann. The seasonal distribution and abundance of barndoor skate on Georges Bank based on scallop dredge surveys. Fisheries Research, 199(??):202–211, March 2018. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783617303132.

Li:2016:MSV

[LJB16] Yan Li, Yan Jiao, and Joan A. Browder. Modeling spatially-varying ecological relationships using geographically weighted generalized linear model: a simulation study based on longline seabird bycatch. Fisheries Research, 181(??):14–24, September 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616300923.

Li:2011:DUC

[LJH11] Yan Li, Yan Jiao, and Qing He. Decreasing uncertainty in catch rate analyses using Delta-AdaBoost: an alternative approach in catch and bycatch analyses with high percentage of zeros. Fisheries Research, 107(1–3):261–271, January 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783610002936.

Lyons:2013:DRG

[LJJ^{+}13] Kady Lyons, Erica T. Jarvis, Salvador J. Jorgensen, Kevin Weng, John O’Sullivan, Chuck Winkler, and Christopher G. Lowe. The degree and result of gillnet fishery interactions with juvenile white sharks in southern California assessed by fishery-independent and -dependent methods. Fisheries Research, 147(??):370–380, October 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613001732.

Lam:2015:HCS

[LKD15] Chi Hin Lam, Dale A. Kiefer, and Michael L. Domeier. Habitat characterization for striped marlin in the Pacific Ocean. Fisheries Research, 166(??):80–91, June 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0165783615000296. See corrigendum [LKD18].

Lam:2018:CHC

[LKD18] Chi Hin Lam, Dale A. Kiefer, and Michael L. Domeier. Corrigendum to “Habitat characterization for striped marlin in the Pacific Ocean” [Fish. Res. 166 (2015) 80–91]. Fisheries Research, 199(??):271, March 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617303338. See [LKD15].

Lan:2012:RBA

[LKL^{+}12] Kuo-Wei Lan, Hiroshi Kawamura, Ming-An Lee, Hsueh-Jung Lu, Teruhisa Shimada, Kohtaro Hosoda, and Futoki Sakaida. Relationship between albacore (Thunnus alalunga) fishing grounds in the Indian Ocean and the thermal environment revealed by cloud-free microwave sea surface temperature. Fisheries Research, 113(1):1–7, January 2012. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783611002943.

Linley:2016:GSF

[LKNL16] Timothy J. Linley, Eirik J. Krogstad, Megan K. Nims, and Russell B. Langshaw. Geochemical signatures in fin rays provide a nonlethal method to distinguish the natal rearing streams of endangered juvenile Chinook salmon Oncorhynchus tshawytscha in the Wenatchee River, Washington. Fisheries Research, 181(??):234–246, September 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783616300996.

Livingston:2011:PTE

[LKR11] Patricia A. Livingston, Gordon H. Kruse, and Laura J. Richards. Progress toward ecosystem-based approaches for the assessment of fisheries under data-limited situations. Fisheries Research, 112(3):105–107, December 2011. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783611003201.

Liu:2017:FDR

[LKZ^{+}17] Hui Liu, Mandy Karnauskas, Xinsheng Zhang, Brian Linton, and Clay Porch. Forecasting dynamics of red snapper (Lutjanus campechanus) in the U.S. Gulf of Mexico. Fisheries Research, 187(??):31–40, March 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616303551.

Lew:2014:FHW

[LL14a] Daniel K. Lew and Douglas M. Larson. Is a fish in hand worth two in the sea? Evidence from a stated preference study. Fisheries Research, 157(??):124–135, September 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783614001398.

Lu:2014:CFS

[LL14b] Hsueh-Jung Lu and Hsiao-Ling Lee. Changes in the fish species composition in the coastal zones of the Kuroshio Current and China Coastal Current during periods of climate change: Observations from the set-net fishery (1993–2011). Fisheries Research, 155(??):103–113, July 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614000782.

Lima:2014:PSR

[LLH^{+}14] Françoise D. Lima, Tatiana S. Leite, Manuel Haimovici, Marcelo F. Nóbrega, and Jorge E. Lins Oliveira. Population structure and reproductive dynamics of Octopus insularis (Cephalopoda: Octopodidae) in a coastal reef environment along northeastern Brazil. Fisheries Research, 152(??):86– 92, April 2014. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783613002026.

Link:2012:ECR

[LLM12] Jason S. Link, Sean M. Lucey, and Jessica H. Melgey. Examining cannibalism in relation to recruitment of silver hake Merluccius bilinearis in the U.S. northwest Atlantic. Fisheries Research, 114(??):31–41, February 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611001718.

Logan:2019:CWA

[LLM^{+}19] James M. Logan, Michael J. Lawrence, George E. Morgan, William M. Twardek, Robert J. Lennox, and Steven J. Cooke. Consequences of winter air exposure on walleye (Sander vitreus) physiology and impairment following a simulated ice-angling event. Fisheries Research, 215(??):106– 113, July 2019. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783619300797.

Lee:2018:PST

[LLP18] Chun-Woo Lee, Jihoon Lee, and Subong Park. Prediction of shooting trajectory of tuna purse seine fishing. Fisheries Research, 208(??):189–201, December 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618302170.

Lan:2015:EVS

[LLWC15] Kuo-Wei Lan, Ming-An Lee, Sheng-Ping Wang, and Zhao-Yang Chen. Environmental variations on swordfish (Xiphias gladius) catch rates in the Indian Ocean. Fisheries Research, 166(??):67–79, June 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783614002550.

LaMesa:2011:GAS

[LMB^{+}11] Gabriele La Mesa, Andrea Molinari, Simone Bava, Maria Grazia Finoia, Riccardo Cattaneo-Vietti, and Leonardo Tunesi. Gradients of abundance of sea breams across the boundaries of a Mediterranean marine protected area. Fisheries Research, 111(1–2):24–30, September 2011. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783611002256.

Lorenzon:2013:HPP

[LMBF13] Simonetta Lorenzon, Marzia Martinis, Diego Borme, and Enrico Antonio Ferrero. Hemolymph parameters as physiological biomarkers for monitoring the effects of fishing and commercial maintenance methods in Squilla mantis (Crustacea, Stomatopoda). Fisheries Research, 137(??):9–17, January 2013. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783612002615.

Lai:2019:CEC

[LMHR19] Eva K. M. Lai, Ute Mueller, Glenn A. Hyndes, and Karina L. Ryan. Comparing estimates of catch and effort for boat-based recreational fishing from aperiodic access-point surveys. Fisheries Research, 219(??):Article 105305, November 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783619301523.

Lozano-Montes:2013:EEE

[LMLBC13] Hector Lozano-Montes, Neil R. Loneragan, Russ Babcock, and Nick Caputi. Evaluating the ecosystem effects of variation in recruitment and fishing effort in the western rock lobster fishery. Fisheries Research, 145(??):128–135, August 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578361300012X.

Lynch:2017:EPS

[LMM^{+}17] Scott D. Lynch, Benjamin J. Marcek, Heather M. Marshall, Peter G. Bushnell, Diego Bernal, and Richard W. Brill. The effects of pop-up satellite archival tags (PSATs) on the metabolic rate and swimming kinematics of juvenile sandbar shark Carcharhinus plumbeus. Fisheries Research, 186 (part 1)(??):205–215, February 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616302661.

Lourenco:2012:STR

[LMN^{+}12] Sílvia Lourenço, Ana Moreno, Luís Narciso, Ángel F. González, and João Pereira. Seasonal trends of the reproductive cycle of Octopus vulgaris in two environmentally distinct coastal areas. Fisheries Research, 127– 128(??):116–124, September 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S016578361200135X.

Lee:2011:ENM

[LMPM11] Hui-Hua Lee, Mark N. Maunder, Kevin R. Piner, and Richard D. Methot, Jr. Estimating natural mortality within a fisheries stock assessment model: an evaluation using simulation analysis based on twelve stock assessments. Fisheries Research, 109(1):89–94, April 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611000403.

Lee:2012:CSS

[LMPM12a] Hui-Hua Lee, Mark N. Maunder, Kevin R. Piner, and Richard D. Methot, Jr. Can steepness of the stock– recruitment relationship be estimated in fishery stock assessment models? Fisheries Research, 125–126(??):254–261, August 2012. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783612001099.

Lee:2012:RRE

[LMPM12b] Hui-Hua Lee, Mark N. Maunder, Kevin R. Piner, and Richard D. Methot, Jr. Reply to ‘the reliability of estimates of natural mortality from stock assessment models’. Fisheries Research, 119–120(??):154–155, May 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612000951.

Lopez:2014:ECS

[LMSM14] Jon Lopez, Gala Moreno, Igor Sancristobal, and Jefferson Murua. Evolution and current state of the technology of echo-sounder buoys used by Spanish tropical tuna purse seiners in the Atlantic, Indian and Pacific Oceans. Fisheries Research, 155(??):127–137, July 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614000794.

Langlois:2015:CLS

[LNC^{+}15a] Tim J. Langlois, Stephen J. Newman, Mike Cappo, Euan S. Harvey, Ben M. Rome, Craig L. Skepper, and Corey B. Wakefield. Corrigendum to ‘length selectivity of commercial fish traps assessed from in situ comparisons with stereo-video: Is there evidence of sampling bias?’. Fisheries Research, 162(??):47, February 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361400294X.

Langlois:2015:LSC

[LNC^{+}15b] Tim J. Langlois, Stephen J. Newman, Mike Cappo, Euan S. Harvey, Ben M. Rome, Craig L. Skepper, and Corey B. Wakefield. Length selectivity of commercial fish traps assessed from in situ comparisons with stereovideo: Is there evidence of sampling bias? Fisheries Research, 161(??):145–155, January 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783614001969.

Leroy:2015:LLI

[LNL^{+}15] Bruno Leroy, Simon Nicol, Antony Lewis, John Hampton, Dale Kolody, Sylvain Caillot, and Simon Hoyle. Lessons learned from implementing three, large-scale tuna tagging programmes in the western and central Pacific Ocean. Fisheries Research, 163(??):23–33, March 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613002117.

Lundin:2011:SEE

[LOC^{+}11] Mikael Lundin, Mikael Ovegård, Linda Calamnius, Lars Hillström, and Sven-Gunnar Lunneryd. Selection efficiency of encircling grids in a herring pontoon trap. Fisheries Research, 111(1–2):127–130, September 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611002323.

Lagares:2016:DCS

[LOdH16] Encarnación Cordón Lagares, Félix García Ordaz, and Juan José García del Hoyo. The determinants that cause small-scale vessels to exit fishing: the case of the Spanish small-scale purse seine fishery. Fisheries Research, 181(??):155–162, September 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361630128X.

Lorenzen:2016:TNP

[Lor16] Kai Lorenzen. Toward a new paradigm for growth modeling in fisheries stock assessments: Embracing plasticity and its consequences. Fisheries Research, 180(??):4–22, August 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783616300066.

Lee:2018:ETV

[LP18] Qi Lee and André E. Punt. Extracting a time-varying climate-driven growth index from otoliths for use in stock assessment models. Fisheries Research, 200(??):93–103, April 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783617303569.

Leleu:2014:MEC

[LPC^{+}14] Kevin Leleu, Dominique Pelletier, Eric Charbonnel, Yves Letourneur, Frédérique Alban, Frédéric Bachet, and Charles F. Boudouresque. Métiers, effort and catches of a Mediterranean small-scale coastal fishery: the case of the Côte Bleue Marine Park. Fisheries Research, 154(??):93– 101, June 2014. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783614000460.

Linnane:2011:AES

[LPHH11] Adrian Linnane, Shane Penny, Matthew Hoare, and Peter Hawthorne. Assessing the effectiveness of size limits and escape gaps as management tools in a commercial rock lobster (Jasus edwardsii) fishery. Fisheries Research, 111(1–2):1–7, September 2011. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783611002232.

Lehuta:2013:SVC

[LPM^{+}13] Sigrid Lehuta, Pierre Petitgas, Stéphanie Mahévas, Martin Huret, Youen Vermard, Andrés Uriarte, and Nicholas R. Record. Selection and validation of a complex fishery model using an uncertainty hierarchy. Fisheries Research, 143(??):57–66, June 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783613000118.

Lee:2014:ULP

[LPMM14] Hui-Hua Lee, Kevin R. Piner, Richard D. Methot, Jr., and Mark N. Maunder. Use of likelihood profiling over a global scaling parameter to structure the population dynamics model: an example using blue marlin in the Pacific Ocean. Fisheries Research, 158(??):138–146, October 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783613003111.

Lee:2019:UCA

[LPTK19] Huihua Lee, Kevin R. Piner, Ian G. Taylor, and Toshihide Kitakado. On the use of conditional age at length data as a likelihood component in integrated population dynamics models. Fisheries Research, 216(??):204–211, August 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619301043.

Luo:2015:NMI

[LQL^{+}15] Wei Luo, Hongyue Qu, Jingyi Li, Xin Wang, and Qiang Lin. A novel method for the identification of seahorses (genus Hippocampus) using cross-species amplifiable microsatellites. Fisheries Research, 172(??):318–324, December 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783615300266.

Levine:2019:DWT

[LR19] Mike Levine and Alex De Robertis. Don’t work too hard: Subsampling leads to efficient analysis of large acoustic datasets. Fisheries Research, 219(??):Article 105323, November 2019. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783619301705.

Llarena-Reino:2019:MGI

[LRAP19] María Llarena-Reino, Elvira Abollo, and Santiago Pascual. Morphological and genetic identification of Pennella instructa (Copepoda: Pennellidae) on Atlantic swordfish (Xiphias gladius, L. 1758). Fisheries Research, 209(??):178–185, January 2019. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783618302479.

Ligas:2010:ALT

[LRM^{+}10] A. Ligas, S. De Ranieri, D. Micheli, B. Reale, P. Sartor, M. Sbrana, and P. Belcari. Analysis of the landings and trawl survey time series from the Tyrrhenian Sea (NW Mediterranean). Fisheries Research, 105(1):46–56, June 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783610000676.

Leonarduzzi:2014:PCE

[LRM14] Ezequiel Leonarduzzi, Karina A. Rodrigues, and Gustavo J. Macchi. Proximate composition and energy density in relation to Argentine hake females (Merluccius hubbsi) morphometrics and condition indices. Fisheries Research, 160(??):33–40, December 2014. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614001532.

Landa:2017:RLC

[LRM17] Jorge Landa and Enrique Rodríguez-Marín. Response to the letter critical of a recent fisheries research publication (reference: 150706-002900). Fisheries Research, 185(??):200, January 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783616302624. See [LRML^{+}15,

Cor17].

Landa:2015:GBT

[LRML^{+}15] Jorge Landa, Enrique Rodriguez-Marin, Patricia L. Luque, Marta Ruiz, and Pablo Quelle. Growth of bluefin tuna (Thunnus thynnus) in the North-eastern Atlantic and Mediterranean based on back-calculation of dorsal fin spine annuli. Fisheries Research, 170(??):190–198, October 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783615001812. See letter [Cor17] and

response [LRM17].

Loewen:2015:DNF

[LRY^{+}15] Tracey N. Loewen, James D. Reist, Panseok Yang, Alison Koleszar, John A. Babaluk, Neil Mochnacz, and Norman M. Halden. Discrimination of northern form Dolly Varden Char (Salvelinus malma malma) stocks of the North Slope, Yukon and Northwest Territories, Canada via otolith trace elements and^{87}Sr/^{86}Sr isotopes. Fisheries Research, 170(??):116–124, October 2015. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783615001745.

Lehtonen:2010:LCG

[LS10] Esa Lehtonen and Petri Suuronen. Live-capture of grey seals in a modified salmon trap. Fisheries Research, 102(1–2):214–216, February 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783609002719.

Lamborn:2019:HPA

[LS19] Chase C. Lamborn and Jordan W. Smith. Human perceptions of, and adaptations to, shifting runoff cycles: a case-study of the Yellowstone River (Montana, USA). Fisheries Research, 216(??):96–108, August 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619300967.

Levsen:2018:SZN

[LSC^{+}18] Arne Levsen, Cecilie S. Svanevik, Paolo Cipriani, Simonetta Mattiucci, Mélanie Gay, Lee C. Hastie, Ivana Bušelić, Ivona Mladineo, Horst Karl, Ute Ostermeyer, Kurt Buchmann, Dánjal P. Højgaard, Ángel F. González, Santiago Pascual, and Graham J. Pierce. A survey of zoonotic nematodes of commercial key fish species from major European fishing grounds — introducing the FP7 PARASITE exposure assessment study. Fisheries Research, 202(??):4– 21, June 2018. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783617302576.

Liu:2013:GDP

[LSH13] Shuang Liu, Jinsheng Sun, and Luis A. Hurtado. Genetic differentiation of Portunus trituberculatus, the world’s largest crab fishery, among its three main fishing areas. Fisheries Research, 148(??):38–46, November 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613001847.

Losee:2017:MPA

[LSJ17] James P. Losee, Todd R. Seamons, and Joe Jauquet. Migration patterns of anadromous cutthroat trout in South Puget Sound: a fisheries management perspective. Fisheries Research, 187(??):218–225, March 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616304167.

Li:2011:LTV

[LSJD11] Zhonglu Li, Xiujuan Shan, Xianshi Jin, and Fangqun Dai. Long-term variations in body length and age at maturity of the small yellow croaker (Larimichthys polyactis Bleeker, 1877) in the Bohai Sea and the Yellow Sea, China. Fisheries Research, 110(1):67–74, June 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611001226.

Lynch:2012:PMU

[LSL12] Patrick D. Lynch, Kyle W. Shertzer, and Robert J. Latour. Performance of methods used to estimate indices of abundance for highly migratory species. Fisheries Research, 125–126(??):27–39, August 2012. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361200080X.

Linnane:2019:TPS

[LSM^{+}19] Adrian Linnane, Anthony D. M. Smith, Richard McGarvey, John E. Feenstra, Janet M. Matthews, Klaas Hartmann, and Caleb Gardner. Trends in productivity of southern rock lobster Jasus edwardsii, across south-eastern Australia: Evidence of a regime shift? Fisheries Research, 219(??):Article 105308, November 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619301559.

Lane:2016:PPG

[LSR16] Henry S. Lane, Jane E. Symonds, and Peter A. Ritchie. The phylogeography and population genetics of Polyprion oxygeneios based on mitochondrial DNA sequences and microsatellite DNA markers. Fisheries Research, 174(??): 19–29, February 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783615300552.

Langseth:2016:MIT

[LSS^{+}16a] Brian J. Langseth, Amy M. Schueller, Kyle W. Shertzer, J. Kevin Craig, and Joseph W. Smith. Management implications of temporally and spatially varying catchability for the Gulf of Mexico menhaden fishery. Fisheries Research, 181(??):186–197, September 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616301266.

Lappalainen:2016:LMP

[LSŠ^{+}16b] Antti Lappalainen, Lauri Saks, Mira Šuštar, Outi Heikinheimo, Kristiina Jürgens, Eevi Kokkonen, Mika Kurkilahti, Aare Verliin, and Markus Vetemaa. Length at maturity as a potential indicator of fishing pressure effects on coastal pikeperch (Sander lucioperca) stocks in the northern Baltic Sea. Fisheries Research, 174(??):47–57, February 2016. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578361530059X.

Long:2017:BTI

[LSS^{+}17] James M. Long, David R. Stewart, Jeremy Shiflet, Dane Balsman, and Daniel E. Shoup. Bait type influences on catch and bycatch in tandem hoop nets set in reservoirs. Fisheries Research, 186 (part 1)(??):102–108, February 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616302673.

Lasne:2010:NST

[LST^{+}10] E. Lasne, M.-R. Sabatié, J. Tremblay, L. Beaulaton, and J.-M. Roussel. A new sampling technique for larval lamprey population assessment in small river catchments. Fisheries Research, 106(1):22–26, October 2010. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610001505.

Landsman:2017:SOI

[LSWR17] Sean Landsman, Jeffrey A. Stein, Gregory Whitledge, and Steven R. Robillard. Stable oxygen isotope analysis confirms natural recruitment of Lake Michigan-origin lake trout (Salvelinus namaycush) to the adult life stage. Fisheries Research, 190(??):15–23, June 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361730022X.

Lyle:2016:CEF

[LT16] Jeremy M. Lyle and Sean R. Tracey. Catch, effort and fishing practices in a recreational gillnet fishery: Assessing the impacts and response to management change. Fisheries Research, 177(??):50–58, May 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616300212.

Lin:2018:MBA

[LT18] Yu-Jia Lin and Wann Nian Tzeng. Modelling the bycatch of Anguilla marmorata using a generalized depletion model with an example from the Taiwanese glass eel fisheries for Anguilla japonica. Fisheries Research, 208(??):210–218, December 2018. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783618302212.

Lin:2017:SAM

[LTHRU17] Yu-Jia Lin, Wang-Nian Tzeng, Yu-San Han, and Ruben H. Roa-Ureta. A stock assessment model for transit stock fisheries with explicit immigration and emigration dynamics: Application to upstream waves of glass eels. Fisheries Research, 195(??):130–140, November 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617301832.

Lescrauwaet:2013:ICC

[LTV^{+}13] Ann-Katrien Lescrauwaet, Els Torreele, Magda Vincx, Hans Polet, and Jan Mees. Invisible catch: a century of bycatch and unreported removals in sea fisheries, Belgium 1929– 2010. Fisheries Research, 147(??):161–174, October 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783613001483.

Libre:2015:ESF

[LvVtB^{+}15] Stella V. D. Libre, George A. K. van Voorn, Guus A. ten Broeke, Megan Bailey, Paul Berentsen, and Simon R. Bush. Effects of social factors on fishing effort: the case of the Philippine tuna purse seine fishery. Fisheries Research, 172(??):250–260, December 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615300424.

Lomeli:2012:ERC

[LW12] Mark J. M. Lomeli and W. Waldo Wakefield. Efforts to reduce Chinook salmon (Oncorhynchus tshawytscha) and rockfish (Sebastes spp.) bycatch in the U.S. west coast Pacific hake (Merluccius productus) fishery. Fisheries Research, 119–120(??):128–132, May 2012. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611003328.

Lomeli:2013:FSG

[LW13] Mark J. M. Lomeli and W. Waldo Wakefield. A flexible sorting grid to reduce Pacific halibut (Hippoglossus stenolepis) bycatch in the US west coast groundfish bottom trawl fishery. Fisheries Research, 143(??):102–108, June 2013. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783613000210.

Lomeli:2016:ESG

[LW16] Mark J. M. Lomeli and W. Waldo Wakefield. Evaluation of a sorting grid bycatch reduction device for the selective flatfish bottom trawl in the U.S. West Coast fishery. Fisheries Research, 183(??):294–303, November 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616302004.

Lomeli:2019:EAI

[LW19a] Mark J. M. Lomeli and W. Waldo Wakefield. The effect of artificial illumination on Chinook salmon behavior and their escapement out of a midwater trawl bycatch reduction device. Fisheries Research, 218(??):112–119, October 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619301109.

Lucca:2019:FIO

[LW19b] Brandyn M. Lucca and Joseph D. Warren. Fisheryindependent observations of Atlantic menhaden abundance in the coastal waters south of New York. Fisheries Research, 218(??):229–236, October 2019. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619301468.

Lennox:2015:IAH

[LWCC15] Robert Lennox, Kim Whoriskey, Glenn T. Crossin, and Steven J. Cooke. Influence of angler hook-set behaviour relative to hook type on capture success and incidences of deep hooking and injury in a teleost fish. Fisheries Research, 164(??):201–205, April 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614003415.

Lomeli:2019:EBS

[LWH19] Mark J. M. Lomeli, W. Waldo Wakefield, and Bent Herrmann. Evaluating off-bottom sweeps of a U.S. West Coast groundfish bottom trawl: Effects on catch efficiency and seafloor interactions. Fisheries Research, 213(??):204–211, May 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783619300141.

Liljestrand:2019:EMM

[LWS19a] Emily M. Liljestrand, Michael J. Wilberg, and Amy M. Schueller. Estimation of movement and mortality of Atlantic menhaden during 1966–1969 using a Bayesian multi-state mark-recovery model. Fisheries Research, 210(??):204–213, February 2019. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783618302844.

Liljestrand:2019:MSD

[LWS19b] Emily M. Liljestrand, Michael J. Wilberg, and Amy M. Schueller. Multi-state dead recovery mark-recovery model performance for estimating movement and mortality rates. Fisheries Research, 210(??):214–223, February 2019. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783618302832.

Landsman:2011:EPB

[LWSC11] Sean J. Landsman, Hedrik J. Wachelka, Cory D. Suski, and Steven J. Cooke. Evaluation of the physiology, behaviour, and survival of adult muskellunge (Esox masquinongy) captured and released by specialized anglers. Fisheries Research, 110(2):377–386, July 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783611001949.

Liu:2011:TPP

[LWY^{+}11] Zi-Ming Liu, Gui-Zhong Wang, Hai-Hui Ye, Shao-Jing Li, Yong Tao, Qiong-Wu Lin, and Emadeldeen H. Mohammed. Tag performance and physiological responses of juvenile mud crabs Scylla paramamosain tagged with visible implant elastomer. Fisheries Research, 110(1):183–188, June 2011. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783611001548.

Liu:2019:CBT

[LWZ^{+}19] Xiaoxiao Liu, Jing Wang, Yunlei Zhang, Huaming Yu, Binduo Xu, Chongliang Zhang, Yiping Ren, and Ying Xue. Comparison between two GAMs in quantifying the spatial distribution of Hexagrammos otakii in haizhou Bay, China. Fisheries Research, 218(??):209–217, October 2019. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619301493.

Lin:2019:ODF

[LXCC19] Dongming Lin, Sipeng Xuan, Zimo Chen, and Xinjun Chen. The ovarian development, fecundity and hypothesis on spawning pattern of common cuttlefish Sepia officinalis off Mauritania. Fisheries Research, 210(??):193–197, February 2019. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783618302248.

Lynch:2014:DTS

[Lyn14] Tim P. Lynch. A decadal time-series of recreational fishing effort collected during and after implementation of a multiple use marine park shows high inter-annual but low spatial variability. Fisheries Research, 151(??):85–90, March 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783613002245.

Li:2015:MSB

[LYWZ15] Zengguang Li, Zhenjiang Ye, Rong Wan, and Chi Zhang. Model selection between traditional and popular methods for standardizing catch rates of target species: a case study of Japanese Spanish mackerel in the gillnet fishery. Fisheries Research, 161(??):312–319, January 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614002665.

Liu:2017:STG

[LZGL17] Bing-Jian Liu, Bai-Dong Zhang, Tian-Xiang Gao, and Jin-Xian Liu. Signatures of temporal genetic differentiation of the small yellow croaker (Larimichthys polyactis) revealed by microsatellite markers. Fisheries Research, 194(??):50–54, October 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783617301327.

Li:2017:EAH

[LZX^{+}17] Min Li, Chongliang Zhang, Binduo Xu, Ying Xue, and Yiping Ren. Evaluating the approaches of habitat suitability modelling for whitespotted conger (Conger myriaster). Fisheries Research, 195(??):230–237, November 2017. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783617302072.

Matthias:2016:CGM

[MAA^{+}16] B. G. Matthias, R. N. M. Ahrens, M. S. Allen, L. A. Lombardi-Carlson, and G. R. Fitzhugh. Comparison of growth models for sequential hermaphrodites by considering multi-phasic growth. Fisheries Research, 179(??):67–75, July 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783616300285.

Matthias:2018:UED

[MAA^{+}18] Bryan G. Matthias, Robert N. M. Ahrens, Micheal S. Allen, Travis Tuten, Zachary A. Siders, and Kyle L. Wilson. Understanding the effects of density and environmental variability on the process of fish growth. Fisheries Research, 198(??):209–219, February 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617302321.

Marenghi:2017:DEN

[MAAW^{+}17] Frank Marenghi, Kathryn Ashton-Alcox, Richard Wong, Bellamy Reynolds, and Gulnihal Ozbay. Dredge efficiency on natural oyster grounds in Delaware Bay and its application in monitoring the Eastern oyster (Crassostrea virginica) stock in Delaware, USA. Fisheries Research, 186 (part 1)(??):292–300, February 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616303356.

Maxwell:2012:AEP

[MABA12] David L. Maxwell, Michael J. Armstrong, Steven Beggs, and John N. Aldridge. Annual egg production estimates of cod (Gadus morhua), plaice (Pleuronectes platessa) and haddock (Melanogrammus aeglefinus) in the Irish Sea: the effects of modelling choices and assumptions. Fisheries Research, 117–118(??):146–155, April 2012. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611002979.

MacCall:2013:UDM

[Mac13] Alec D. MacCall. Use of the delta method to evaluate the precision of assessments that fix parameter values. Fisheries Research, 142(??):56–60, May 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612002317.

Merten:2014:MDC

[MAH14] Wessley Merten, Richard Appeldoorn, and Donald Hammond. Movements of dolphinfish (Coryphaena hippurus) along the U.S. east coast as determined through mark and recapture data. Fisheries Research, 151(??):114–121, March 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578361300249X.

Merten:2016:MDD

[MAH16] Wessley Merten, Richard Appeldoorn, and Donald Hammond. Movement dynamics of dolphinfish (Coryphaena hippurus) in the northeastern Caribbean Sea: Evidence of seasonal re-entry into domestic and international fisheries throughout the western central Atlantic. Fisheries Research, 175(??):24–34, March 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615301181.

Mamauag:2013:FVA

[MAM^{+}13] Samuel S. Mamauag, Porfirio M. Aliño, Renmar Jun S. Martinez, Richard N. Muallil, Maria Victoria A. Doctor, Emerlinda C. Dizon, Rollan C. Geronimo, Fleurdeliz M. Panga, and Reniel B. Cabral. A framework for vulnerability assessment of coastal fisheries ecosystems to climate change — tool for understanding resilience of fisheries (VA–TURF). Fisheries Research, 147(??):381–393, October 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613001719.

March:2014:GAF

[MAP14] David March, Josep Alós, and Miquel Palmer. Geospatial assessment of fishing quality considering environmental and angler-related factors. Fisheries Research, 154(??):63–72, June 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783614000368.

Marszalec:2018:AQP

[Mar18] Daniel Marszalec. Auctions for quota: a primer and perspectives for the future. Fisheries Research, 203(??):84–92, July 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578361730200X.

McKeown:2015:IGO

[MAS15] Niall J. McKeown, Alexander I. Arkhipkin, and Paul W. Shaw. Integrating genetic and otolith microchemistry data to understand population structure in the Patagonian hoki (Macruronus magellanicus). Fisheries Research, 164(??):1– 7, April 2015. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783614002975.

McKeown:2017:RGP

[MAS17] Niall J. McKeown, Alexander I. Arkhipkin, and Paul W. Shaw. Regional genetic population structure and fine scale genetic cohesion in the southern blue whiting Micromesistius australis. Fisheries Research, 185(??):176–184, January 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783616302995.

Maunder:2011:REL

[Mau11] Mark N. Maunder. Review and evaluation of likelihood functions for composition data in stock-assessment models: Estimating the effective sample size. Fisheries Research, 109(2–3):311–319, May 2011. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611000890.

Maunder:2012:ESR

[Mau12] Mark N. Maunder. Evaluating the stock–recruitment relationship and management reference points: Application to summer flounder (Paralichthys dentatus) in the U.S. mid-Atlantic. Fisheries Research, 125–126(??):20–26, August 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783612000811.

Mayfield:2010:EFA

[May10] S. Mayfield. Enhancing fishery assessments for an Australian abalone fishery using commercial weight-grade data. Fisheries Research, 105(1):28–37, June 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610000548.

Matsushita:2012:FRC

[MAY12] Yoshiki Matsushita, Toru Azuno, and Yukiko Yamashita. Fuel reduction in coastal squid jigging boats equipped with various combinations of conventional metal halide lamps and low-energy LED panels. Fisheries Research, 125–126(??):14– 19, August 2012. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783612000793.

Montenegro:2016:BSS

[MB16] Carlos Montenegro and Márcia Branco. Bayesian statespace approach to biomass dynamic models with skewed and heavy-tailed error distributions. Fisheries Research, 181(??):48–62, September 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783616300893.

Molton:2013:HLC

[MBB13] Kyle J. Molton, Travis O. Brenden, and James R. Bence. Harvest levels that conserve spawning biomass can provide larger and more stable and sustainable yields in intermixed fisheries. Fisheries Research, 147(??):264–283, October 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783613001689.

Mazloumi:2017:DCG

[MBF^{+}17] N. Mazloumi, P. Burch, A. J. Fowler, Z. A. Doubleday, and B. M. Gillanders. Determining climate-growth relationships in a temperate fish: a sclerochronological approach. Fisheries Research, 186 (part 1)(??):319–327, February 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578361630340X.

Maxwell:2019:UAT

[MBF19a] Suzanne L. Maxwell, Greg B. Buck, and April V. Faulkner. Using acoustic telemetry to expand sonar escapement indices of Chinook salmon to in-river abundance estimates. Fisheries Research, 220(??):Article 105347, December 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619302024.

Millar:2019:FIS

[MBF^{+}19b] Russell B. Millar, Luísa Barros, Paulo Fonseca, Paulo T. Santos, and Aida Campos. Further improvements in sorting grids for the crustacean trawl fishery off the southern coast of Portugal. Fisheries Research, 219(??):Article 105329, November 2019. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783619301845.

Miller:2010:INA

[MBG^{+}10] Jessica A. Miller, M. Renee Bellinger, James T. Golden, Lorissa Fujishin, and Michael A. Banks. Integration of natural and artificial markers in a mixed stock analysis of Chinook salmon (Oncorhynchus tshawytscha). Fisheries Research, 102(1–2):152–159, February 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783609003075.

MacKenzie:2013:PHM

[MBHGN13] K. MacKenzie, P. Brickle, W. Hemmingsen, and M. George-Nascimento. Parasites of hoki, Macruronus magellanicus, in the Southwest Atlantic and Southeast Pacific Oceans, with an assessment of their potential value as biological tags. Fisheries Research, 145(??):1–5, August 2013. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783613000763.

Marsden:2016:FYE

[MBJ^{+}16] J. Ellen Marsden, Thomas R. Binder, James Johnson, Ji He, Natalie Dingledine, Janice Adams, Nicholas S. Johnson, Tyler J. Buchinger, and Charles C. Krueger. Fiveyear evaluation of habitat remediation in Thunder Bay, Lake Huron [sic: Superior]: Comparison of constructed reef characteristics that attract spawning lake trout. Fisheries Research, 183(??):275–286, November 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616302016.

Melli:2019:RSA

[MBK19] Valentina Melli, Matt K. Broadhurst, and Steven J. Kennelly. Refining a Simple Anterior Fish Excluder (SAFE) for penaeid trawls. Fisheries Research, 214(??):1–9, June 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619300232.

Morales-Bojorquez:2010:CSA

[MBNM10] Enrique Morales-Bojórquez and Manuel O. Nevárez-Martínez. Catch-at-size analysis for Dosidicus gigas in the central Gulf of California, Mexico in 1996–2002. Fisheries Research, 106(2):214–221, November 2010. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783610001815.

Marschoff:2012:SRP

[MBOAA12] Enrique R. Marschoff, Esteban R. Barrera-Oro, Nadia S. Alescio, and David G. Ainley. Slow recovery of previously depleted demersal fish at the South Shetland Islands, 1983–2010. Fisheries Research, 125–126(??):206–213, August 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783612000975.

Martinez-Banos:2018:EHM

[MBRDM18] Pedro Martínez-Baños, John Gabriel Ramírez, Montserrat Demestre, and Francesc Maynou. European hake (Merluccius merluccius) assessment based on size frequencies and basic biological parameters in the SW Mediterranean. Fisheries Research, 205(??):35–42, September 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618301024.

Maria:2014:PSA

[MBS14] Maite De María, Francisco Rafael Barboza, and Diana Szteren. Predation of South American sea lions (Otaria flavescens) on artisanal fisheries in the Rio de la Plata estuary. Fisheries Research, 149(??):69–73, January 2014. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783613002166.

Murphy:2015:MWL

[MBS^{+}15] Hannah M. Murphy, Terry Bungay, Edward W. P. Stern, Paul D. Winger, and Dominique Robert. Measurement of warp loads of bongo plankton samplers. Fisheries Research, 162(??):43–46, February 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361400280X.

McHugh:2015:CTC

[MBSM15] Matthew J. McHugh, Matt K. Broadhurst, David J. Sterling, and Russell B. Millar. Comparing three conventional penaeid-trawl otter boards and the new batwing design. Fisheries Research, 167(??):180–189, July 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615000582.

Miethe:2014:IAF

[MBvDN14] Tanja Miethe, Francois Bastardie, Christian von Dorrien, and J. Rasmus Nielsen. Impact assessment of a fisheries closure with effort and landings spatial analyses: a case study in the Western Baltic Sea. Fisheries Research, 157(??):170–179, September 2014. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614001386.

Montgomery:2012:ERM

[MBW12] S. S. Montgomery, I. M. Barchia, and C. T. Walsh. Estimating rates of mortality in stocks of Metapenaeus macleayi in estuaries of eastern Australia. Fisheries Research, 113(1):55–67, January 2012. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611002955.

Morgan:2010:CTS

[MC10] Alexia Morgan and John K. Carlson. Capture time, size and hooking mortality of bottom longline-caught sharks. Fisheries Research, 101(1–2):32–37, January 5, 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783609002471.

Murray:2016:IER

[MCB^{+}16] Fiona Murray, Philip Copland, Philip Boulcott, Mike Robertson, and Nick Bailey. Impacts of electrofishing for razor clams (Ensis spp.) on benthic fauna. Fisheries Research, 174(??):40–46, February 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615300746.

Mannheim:2018:WAR

[MCB^{+}18] Samantha L. Mannheim, Amber-Robyn Childs, Edward C. Butler, Alexander C. Winkler, Matthew C. Parkinson, Matthew W. Farthing, Tamzyn Zweig, Meaghen McCord, Natalia Drobniewska, and Warren M. Potts. Working with, not against recreational anglers: Evaluating a proenvironmental behavioural strategy for improving catchand-release behaviour. Fisheries Research, 206(??):44–56, October 2018. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783618301218.

Mancinelli:2017:ABC

[MCC^{+}17] Giorgio Mancinelli, Paula Chainho, Lucrezia Cilenti, Silvia Falco, Kostas Kapiris, George Katselis, and Filipe Ribeiro. On the Atlantic blue crab (Callinectes sapidus Rathbun 1896) in southern European coastal waters: Time to turn a threat into a resource? Fisheries Research, 194(??):1–8, October 2017. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783617301236.

Murchie:2013:MPB

[MCD^{+}13] Karen J. Murchie, Steven J. Cooke, Andy J. Danylchuk, Sascha E. Danylchuk, Tony L. Goldberg, Cory D. Suski, and David P. Philipp. Movement patterns of bonefish (Albula vulpes) in tidal creeks and coastal waters of eleuthera, the Bahamas. Fisheries Research, 147(??):404–412, October 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783613000878.

Mann:2018:CRS

[MCDP18] Bruce Q. Mann, Paul D. Cowley, Stuart W. Dunlop, and Warren M. Potts. Is catch-and-release shore angling compatible with the conservation goals of marine protected areas? A case study from the iSimangaliso Wetland Park in South Africa. Fisheries Research, 208(??):179–188, December 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783618302200.

Martin:2014:UPO

[MCEP14] Dustin R. Martin, Christopher J. Chizinski, Kent M. Eskridge, and Kevin L. Pope. Using posts to an online social network to assess fishing effort. Fisheries Research, 157(??):24–27, September 2014. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614000952.

Moffett:2011:EPM

[MCH11] Cinamon Moffett, Yong Chen, and Margaret Hunter. Evaluating port monitoring program: a case study of the northern shrimp fishery in the Gulf of Maine. Fisheries Research, 108(2–3):321–326, March 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611000063.

Mandelman:2013:STP

[MCI^{+}13] J. W. Mandelman, A. M. Cicia, G. W. Ingram, W. B. Driggers, K. M. Coutre, and J. A. Sulikowski. Short-term post-release mortality of skates (family Rajidae) discarded in a western North Atlantic commercial otter trawl fishery. Fisheries Research, 139(??):76–84, March 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612003062.

Munoz-Colmenero:2017:EMS

[MCJD^{+}17] Marta Muñoz-Colmenero, Francis Juanes, Eduardo Dopico, Jose Luis Martinez, and Eva Garcia-Vazquez. Economy matters: a study of mislabeling in salmon products from two regions, Alaska and Canada (Northwest of America) and Asturias (Northwest of Spain). Fisheries Research, 195(??):180–185, November 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617301959.

Miller:2012:MBF

[MCM12a] Dana D. Miller, Maurice Clarke, and Stefano Mariani. Mismatch between fish landings and market trends: a western European case study. Fisheries Research, 121–122(??):104– 114, June 2012. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783612000471.

Mulazzani:2012:PFG

[MCM12b] Luca Mulazzani, Luca Camanzi, and Giulio Malorgio. Price formation and geographic market integration: an empirical investigation of Adriatic small pelagic species. Fisheries Research, 119–120(??):99–107, May 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611003912.

Morris:2018:ESS

[MCMK18] C. J. Morris, D. Cote, B. Martin, and D. Kehler. Effects of 2D seismic on the snow crab fishery. Fisheries Research, 197(??):67–77, January 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617302606.

Montecinos:2016:SIT

[MCN16] Sandra Montecinos, Leonardo R. Castro, and Sergio Neira. Stable isotope (δ ^{13}C and δ ^{15}N) and trophic position of Patagonian sprat (Sprattus fuegensis) from the Northern Chilean Patagonia. Fisheries Research, 179(??):139–147, July 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783616300443.

Medina:2014:FAM

[MCP14] Gabriela Medina, Leonardo Castro, and Silvio Pantoja. Fatty acids in Merluccius australis tissues, a comparison between females from inshore and offshore spawning areas in the Chilean Patagonia. Fisheries Research, 160(??):41–49, December 2014. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578361300283X.

Maunder:2016:GTE

[MCP^{+}16] Mark N. Maunder, Paul R. Crone, André E. Punt, Juan L. Valero, and Brice X. Semmens. Growth: Theory, estimation, and application in fishery stock assessment models. Fisheries Research, 180(??):1–3, August 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616300613.

Marquet:2017:SCH

[MCP^{+}17a] Nathalie Marquet, Chantal Conand, Deborah M. Power, Adelino V. M. Canário, and Mercedes González-Wangüemert. Sea cucumbers, Holothuria arguinensis and H. mammata, from the southern Iberian Peninsula: Variation in reproductive activity between populations from different habitats. Fisheries Research, 191(??):120–130, July 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617300747.

Maunder:2017:DCW

[MCP^{+}17b] Mark N. Maunder, Paul R. Crone, André E. Punt, Juan L. Valero, and Brice X. Semmens. Data conflict and weighting, likelihood functions and process error. Fisheries Research, 192(??):1–4, August 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617300735.

Maunder:2014:STE

[MCVS14] M. N. Maunder, P. R. Crone, J. L. Valero, and B. X. Semmens. Selectivity: Theory, estimation, and application in fishery stock assessment models. Fisheries Research, 158(??):1–4, October 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783614001106.

Maunder:2010:DMC

[MD10] Mark N. Maunder and Richard B. Deriso. Dealing with missing covariate data in fishery stock assessment models. Fisheries Research, 101(1–2):80–86, January 5, 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783609002537.

Maunder:2013:SRM

[MD13] Mark N. Maunder and Richard B. Deriso. A stock– recruitment model for highly fecund species based on temporal and spatial extent of spawning. Fisheries Research, 146(??):96–101, September 2013. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613001021.

Masello:2016:DSR

[MD16] Arianna Masello and Omar Defeo. The deep-sea red crab Chaceon notialis (Geryonidae) in the southwestern Atlantic Ocean: Spatial patterns and long-term effects of fishing. Fisheries Research, 183(??):254–262, November 2016. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783616302053.

Moreno:2016:FAD

[MDC^{+}16] G. Moreno, L. Dagorn, M. Capello, J. Lopez, J. Filmalter, F. Forget, I. Sancristobal, and K. Holland. Fish aggregating devices (FADs) as scientific platforms. Fisheries Research, 178(??):122–129, June 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615300977.

Mazloumi:2017:ETS

[MDG17] N. Mazloumi, Z. A. Doubleday, and B. M. Gillanders. The effects of temperature and salinity on otolith chemistry of King George whiting. Fisheries Research, 196(??):66–74, December 2017. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783617302242.

Maunder:2015:USS

[MDH15] Mark N. Maunder, Richard B. Deriso, and Charles H. Hanson. Use of state-space population dynamics models in hypothesis testing: advantages over simple log-linear regressions for modeling survival, illustrated with application to longfin smelt (Spirinchus thaleichthys). Fisheries Research, 164(??):102–111, April 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614003105.

Mamet:2019:HGD

[MDL19] Leandro Nicolás Getino Mamet, Liza Gómez Daglio, and Francisco Javier García-De León. High genetic differentiation in the edible cannonball jellyfish (Cnidaria: Scyphozoa: Stomolophus spp.) from the Gulf of California, Mexico. Fisheries Research, 219(??):Article 105328, November 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619301833.

Mehault:2010:VTE

[MDPCSR10] Sonia Mehault, Rosario Domínguez-Petit, Santiago Cerviño, and Fran Saborido-Rey. Variability in total egg production and implications for management of the southern stock of European hake. Fisheries Research, 104(1–3):111–122, May 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783610000834.

Mormede:2017:USP

[MDPH17] Sophie Mormede, Alistair Dunn, Steve Parker, and Stuart Hanchet. Using spatial population models to investigate the potential effects of the Ross Sea region marine protected area on the Antarctic toothfish population.

Fisheries Research, 190(??):164–174, June 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617300528.

Moore:2019:DMC

[MDR^{+}19] Claire Moore, Sarah Davie, Marianne Robert, Lionel Pawlowski, Paul Dolder, and Colm Lordan. Defining métier for the Celtic Sea mixed fisheries: a multiannual international study of typology. Fisheries Research, 219(??):Article 105310, November 2019. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619301572.

Matson:2018:MMP

[MDS^{+}18] Ronan Matson, Karen Delanty, Samuel Shephard, Brian Coghlan, and Fiona Kelly. Moving from multiple pass depletion to single pass timed electrofishing for fish community assessment in wadeable streams. Fisheries Research, 198(??):99–108, February 2018. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617302849.

Murua:2015:IOT

[MEM15] Hilario Murua, J. Paige Eveson, and Francis Marsac. The Indian Ocean tuna tagging programme: Building better science for more sustainability. Fisheries Research, 163(??):1– 6, March 2015. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783614002136.

Munoz-Exposito:2017:NAO

[MEMdU^{+}17] Pedro Muñoz-Expósito, David Macías, José María Ortíz de Urbina, Salvador García-Barcelona, María José Gómez, and José C. Báez. North Atlantic oscillation affects the physical condition of migrating bullet tuna Auxis rochei (Risso, 1810) from the Western Mediterranean Sea. Fisheries Research, 194(??):84–88, October 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617301443.

Macdonald:2013:IMM

[MFC^{+}13] Jed I. Macdonald, Jessica H. Farley, Naomi P. Clear, Ashley J. Williams, Thor I. Carter, Campbell R. Davies, and Simon J. Nicol. Insights into mixing and movement of South Pacific albacore Thunnus alalunga derived from trace elements in otoliths. Fisheries Research, 148(??):56–63, November 2013. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783613001859.

Madsen:2010:DNC

[MFHK10] Niels Madsen, Rikke Petri Frandsen, René Holst, and Ludvig A. Krag. Development of new concepts for escape windows to minimise cod catches in Norway lobster fisheries. Fisheries Research, 103(1–3):25–29, April 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610000123.

Marsh:2015:VTP

[MFHK15] Jennifer M. Marsh, Robert J. Foy, Nicola Hillgruber, and Gordon H. Kruse. Variability in trophic positions of four commercially important groundfish species in the Gulf of Alaska. Fisheries Research, 165(??):100–114, May 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783615000156.

Mina:2016:EBF

[MFIG16] José Santiago Arroyo Mina, Daniel A. Revollo Fernández, Alonso Aguilar Ibarra, and Nikolaos Georgantzis. Economic behavior of fishers under climate-related uncertainty: Results from field experiments in Mexico and Colombia. Fisheries Research, 183(??):304–317, November 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616301576.

Militz:2017:CPT

[MFKS17] Thane A. Militz, Simon Foale, Jeff Kinch, and Paul C. Southgate. Consumer perspectives on theoretical certification schemes for the marine aquarium trade. Fisheries Research, 193(??):33–42, September 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617300899.

Millar:2016:MCP

[MFMM16] Colin P. Millar, Robert J. Fryer, Karen J. Millidine, and Iain A. Malcolm. Modelling capture probability of Atlantic salmon (Salmo salar) from a diverse national electrofishing dataset: Implications for the estimation of abundance. Fisheries Research, 177(??):1–12, May 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616300017.

Marquez-Farias:2013:IRP

[MFRJ13] J. Fernando Márquez-Farías and F. Jorge Rosales-Juárez. Intrinsic rebound potential of the endangered (Totoaba macdonaldi) population, endemic to the Gulf of California, México. Fisheries Research, 147(??):150–153, October 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783613001501.

Militelli:2019:EPV

[MFRM19] María I. Militelli, Carla Firpo, Karina A. Rodrigues, and Gustavo J. Macchi. Egg production and validation of clutch fullness indices scale of southern king crab, Lithodes santolla, in the Central Patagonian Sector, Argentina (44^{o}– 48^{o}S). Fisheries Research, 211(??):40–45, March 2019. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783618302972.

Moreira:2018:PSB

[MFS^{+}18] C. Moreira, E. Froufe, A. N. Sial, A. Caeiro, P. Vaz-Pires, and A. T. Correia. Population structure of the blue jack mackerel (Trachurus picturatus) in the NE Atlantic inferred from otolith microchemistry. Fisheries Research, 197(??):113–122, January 2018. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617302266.

Moreira:2019:OSA

[MFVPC19] C. Moreira, E. Froufe, P. Vaz-Pires, and A. T. Correia. Otolith shape analysis as a tool to infer the population structure of the blue jack mackerel, Trachurus picturatus, in the NE Atlantic. Fisheries Research, 209(??):40–48, January 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783618302455.

Mendoza:2010:UCT

[MGB10] Manuel Mendoza, Teresa García, and Jorge Baro. Using classification trees to study the effects of fisheries management plans on the yield of Merluccius merluccius (Linnaeus, 1758) in the Alboran Sea (Western Mediterranean). Fisheries Research, 102(1–2):191–198, February 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783609003142.

Muallil:2011:WEA

[MGC^{+}11] Richard N. Muallil, Rollan C. Geronimo, Deborah Cleland, Reniel B. Cabral, Maria Victoria Doctor, Annabelle Cruz-Trinidad, and Porfirio M. Aliño. Willingness to exit the artisanal fishery as a response to scenarios of declining catch or increasing monetary incentives. Fisheries Research, 111(1–2):74–81, September 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361100230X.

Melvin:2016:FVS

[MGLT16] Gary D. Melvin, François Gerlotto, Carolina Lang, and Pedro Trillo. Fishing vessels as scientific platforms: an introduction. Fisheries Research, 178(??):1–3, June 2016. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783616300522.

McAdam:2012:TDS

[MGM12] Bruce J. McAdam, Timothy B. Grabowski, and Gudrún Marteinsdóttir. Testing for differences in spatial distributions from individual based data. Fisheries Research, 127–128(??):148–153, September 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361200104X.

Mattiucci:2018:PGS

[MGP^{+}18] Simonetta Mattiucci, Lucilla Giulietti, Michela Paoletti, Paolo Cipriani, Melanie Gay, Arne Levsen, Regina Klapper, Horst Karl, Miguel Bao, Graham J. Pierce, and Giuseppe Nascetti. Population genetic structure of the parasite Anisakis simplex (s. s.) collected in Clupea harengus L. from North East Atlantic fishing grounds. Fisheries Research, 202(??):103–111, June 2018. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617302163.

Melvin:2013:RSB

[MGR13] Edward F. Melvin, Troy J. Guy, and Lorraine B. Read. Reducing seabird bycatch in the South African joint venture tuna fishery using bird-scaring lines, branch line weighting and nighttime setting of hooks. Fisheries Research, 147(??):72–82, October 2013. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361300132X.

Melvin:2014:BPS

[MGR14] Edward F. Melvin, Troy J. Guy, and Lorraine B. Read. Best practice seabird bycatch mitigation for pelagic longline fisheries targeting tuna and related species. Fisheries Research, 149(??):5–18, January 2014. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613001768.

Maunder:2011:UCV

[MH11] Mark N. Maunder and Shelton J. Harley. Using cross validation model selection to determine the shape of nonparametric selectivity curves in fisheries stock assessment models. Fisheries Research, 110(2):283–288, July 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611001664.

Mulcahy:2014:EMF

[MH14] Daniel M. Mulcahy and Craig A. Harms. Experimental methods fail to address the questions posed in studies of surgical techniques. Fisheries Research, 156(??):1–5, August 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783614000745.

Mason:2019:CLE

[MHB^{+}19] Julia G. Mason, Elliott L. Hazen, Steven J. Bograd, Heidi Dewar, and Larry B. Crowder. Community-level effects of spatial management in the California drift gillnet fishery. Fisheries Research, 214(??):175–182, June 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619300530.

Munroe:2018:MLD

[MHC^{+}18] Daphne M. Munroe, Dale Haidvogel, Joseph C. Caracappa, John M. Klinck, Eric N. Powell, Eileen E. Hofmann, Burton V. Shank, and Deborah R. Hart. Modeling larval dispersal and connectivity for Atlantic sea scallop (Placopecten magellanicus) in the Middle Atlantic Bight. Fisheries Research, 208(??):7–15, December 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618301899.

Madsen:2017:DTS

[MHFH17] Niels Madsen, René Holst, Rikke Frandsen, and Kurt Hansen. Development and test of selective sorting grids used in the Norway lobster (Nephrops norvegicus) fishery. Fisheries Research, 185(??):26–33, January 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616303307.

Madsen:2012:DTR

[MHFK12] Niels Madsen, Kurt E. Hansen, Rikke Petri Frandsen, and Ludvig Ahm Krag. Development and test of a remotely operated minisampler for discrete trawl sampling. Fisheries Research, 123–124(??):16–20, July 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611003456.

Mapp:2017:OSS

[MHK^{+}17] James Mapp, Ewan Hunter, Jeroen Van Der Kooij, Sally Songer, and Mark Fisher. Otolith shape and size: the importance of age when determining indices for fish-stock separation. Fisheries Research, 190(??):43–52, June 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617300267.

Maxner:2016:IFI

[MHRD16] Emily Maxner, Norman M. Halden, James D. Roth, and Gail K. Davoren. Intrinsic factors influence the timing of arrival of capelin (Mallotus villosus) to spawning grounds in coastal Newfoundland. Fisheries Research, 179(??):202– 212, July 2016. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783616300583.

Murua:2010:DEP

[MIÁ^{+}10] H. Murua, L. Ibaibarriaga, P. Álvarez, M. Santos, M. Korta, M. Santurtun, and L. Motos. The daily egg production method: a valid tool for application to European hake in the Bay of Biscay? Fisheries Research, 104(1–3):100–110, May 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783609001611.

Moura:2014:LSD

[MJC^{+}14] Teresa Moura, Emma Jones, Maurice W. Clarke, Charles F. Cotton, Paul Crozier, Ross K. Daley, Guzman Diez, Helen Dobby, Jan E. Dyb, Inge Fossen, Sarah B. Irvine, Klara Jakobsdottir, Luis J. López-Abellán, Pascal Lorance, Pedro Pascual-Alayón, Ricardo B. Severino, and Ivone Figueiredo. Large-scale distribution of three deep-water squaloid sharks: Integrating data on sex, maturity and environment. Fisheries Research, 157(??):47–61, September 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361400112X.

McKinzie:2014:FSH

[MJL14] Megan K. McKinzie, Erica T. Jarvis, and Christopher G. Lowe. Fine-scale horizontal and vertical movement of barred sand bass, Paralabrax nebulifer, during spawning and nonspawning seasons. Fisheries Research, 150(??):66–75, February 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783613002439.

Mueller:2017:REM

[MJL^{+}17] Robert P. Mueller, Jill Janak, Stephanie A. Liss, Richard S. Brown, Zhiqun Deng, and Ryan A. Harnish. Retention and effects of miniature transmitters in juvenile American eels. Fisheries Research, 195(??):52–58, November 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617301728.

McCutcheon:2013:EPL

[MK13] Sara M. McCutcheon and Stephen M. Kajiura. Electrochemical properties of lanthanide metals in relation to their application as shark repellents. Fisheries Research, 147(??):47–54, October 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783613001318.

Makiguchi:2017:STE

[MK17] Yuya Makiguchi and Takahito Kojima. Short term effects of relative tag size and surgical implantation on feeding behaviour, survival rate, plasma lactate and growth rate in juvenile to adult rainbow trout (Oncorhynchus mykiss). Fisheries Research, 185(??):54–61, January 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616303289.

Melvin:2016:AAD

[MKH16] Gary D. Melvin, Rudy Kloser, and Taina Honkalehto. The adaptation of acoustic data from commercial fishing vessels in resource assessment and ecosystem monitoring. Fisheries Research, 178(??):13–25, June 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615300862.

Melli:2019:CAB

[MKHK19] V. Melli, L. A. Krag, B. Herrmann, and J. D. Karlsen. Can active behaviour stimulators improve fish separation from Nephrops (Nephrops norvegicus) in a horizontally divided trawl codend? Fisheries Research, 211(??):282–290, March 2019. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783618303400.

Maina:2016:MAI

[MKK^{+}16] I. Maina, S. Kavadas, S. Katsanevakis, S. Somarakis, G. Tserpes, and S. Georgakarakos. A methodological approach to identify fishing grounds: a case study on Greek trawlers. Fisheries Research, 183(??):326–339, November 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783616302107.

Mirimin:2016:GAO

[MKM^{+}16] Luca Mirimin, Sven Kerwath, Brett Macey, Stephen J. Lamberth, Paul D. Cowley, Aletta Bester van der Merwe, Paulette Bloomer, and Rouvay Roodt-Wilding. Genetic analyses of overfished silver kob Argyrosomus inodorus (Scieanidae) stocks along the southern African coast. Fisheries Research, 176(??):100–106, April 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615301260.

Mason:2010:HRH

[ML10] Tom J. Mason and Christopher G. Lowe. Home range, habitat use, and site fidelity of barred sand bass within a southern California marine protected area. Fisheries Research, 106(1):93–101, October 2010. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610001773.

Meynecke:2011:CCF

[ML11] Jan-Olaf Meynecke and Shing Yip Lee. Climate-coastal fisheries relationships and their spatial variation in Queensland, Australia. Fisheries Research, 110(2):365–376, July 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783611001937.

Miller:2017:SBR

[ML17] Timothy J. Miller and Christopher M. Legault. Statistical behavior of retrospective patterns and their effects on estimation of stock and harvest status. Fisheries Research, 186 (part 1)(??):109–120, February 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361630251X.

Miralles:2016:PGS

[MLBGV16] Laura Miralles, Ling S. Lee, Yaisel Borrell, and Eva Garcia-Vazquez. Population genetic structure of the European conger (Conger conger) in North East Atlantic and West Mediterranean Sea. Fisheries Research, 174(??):245–249, February 2016. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783615301077.

Marinovic:2016:CSG

[MLBTM16] Zoran Marinović, Jelena Lujić, Violeta Bolić-Trivunović, and Goran Marković. Comparative study of growth in Carassius gibelio (Bloch, 1782) and Rutilus rutilus (L., 1758) from two Serbian reservoirs: Multi-model analysis and inferences. Fisheries Research, 173(??):11–19, January 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783615001265.

McLean:2019:IPB

[MLC^{+}19a] M. F. McLean, M. K. Litvak, S. J. Cooke, K. C. Hanson, D. A. Patterson, S. G. Hinch, and G. T. Crossin. Immediate physiological and behavioural response from catchand-release of wild white sturgeon (Acipenser transmontanus Richardson, 1836). Fisheries Research, 214(??):65– 75, June 2019. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783619300396.

Moore:2019:AVN

[MLC^{+}19b] Claire Moore, Deirdre Lynch, Maurice Clarke, Rick Officer, Jane Mills, Jean Louis-Defour, and Deirdre Brophy. Age verification of north Atlantic sprat. Fisheries Research, 213(??):144–150, May 2019. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619300189.

McGarvey:2010:IRC

[MLF^{+}10] Richard McGarvey, Adrian J. Linnane, John E. Feenstra, André E. Punt, and Janet M. Matthews. Integrating recapture-conditioned movement estimation into spatial stock assessment: a South Australian lobster fishery application. Fisheries Research, 105(2):80–90, July 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610000706.

McIntyre:2015:SMC

[MLM15] Rosslyn McIntyre, Andy Lawler, and Robin Masefield. Size of maturity of the common whelk, Buccinum undatum: Is the minimum landing size in England too low? Fisheries Research, 162(??):53–57, February 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614002951.

Moreno:2014:EHP

[MLP^{+}14] A. Moreno, S. Lourenço, J. Pereira, M. B. Gaspar, H. N. Cabral, G. J. Pierce, and A. M. P. Santos. Essential habitats for pre-recruit Octopus vulgaris along the Portuguese coast. Fisheries Research, 152(??):74–85, April 2014. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783613001860.

Moran-Lopez:2017:ITT

[MLT17] R. Morán-López and O. Uceda Tolosa. Image techniques in turbid rivers: a ten-year assessment of cyprinid stocks composition and size. Fisheries Research, 195(??):186–193, November 2017. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783617302035.

Moran-Lopez:2018:ONA

[MLT18] R. Morán-López and O. Uceda Tolosa. Obstacle negotiation attempts by leaping cyprinids indicate bank-side spawning migration routes. Fisheries Research, 197(??):84–87, January 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783617302588.

Mateo:2017:ICH

[MLTD17] Maria Mateo, Patrick Lambert, Stéphane Tétard, and Hilaire Drouineau. Impacts that cause the highest direct mortality of individuals do not necessarily have the greatest influence on temperate eel escapement. Fisheries Research, 193(??):51–59, September 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617300917.

Munoz:2013:EAF

[MLV13] Marta Muñoz, Josep Lloret, and Sílvia Vila. Effects of artisanal fisheries on the scorpaenids (Scorpaena spp.) reproduction in the marine protected area of Cap de Creus (NW Mediterranean). Fisheries Research, 138(??):146–151, February 2013. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783612002494.

McClelland:2011:UPT

[MM11] Gary McClelland and Jason Melendy. Use of parasites as tags in delineating stocks of Atlantic cod (Gadus morhua) from the southern Gulf of St. Lawrence and the Cape Breton Shelf. Fisheries Research, 107(1–3):233–238, January 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783610002894.

Mbaru:2013:EGA

[MM13] E. K. Mbaru and T. R. McClanahan. Escape gaps in African basket traps reduce bycatch while increasing body sizes and incomes in a heavily fished reef lagoon. Fisheries Research, 148(??):90–99, November 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361300204X.

McAdam:2014:BFP

[MM14] Bruce J. McAdam and C. Tara Marshall. Bayesian fitting of probabilistic maturation reaction norms to population-level data. Fisheries Research, 159(??):105–113, November 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783614001866.

Mullon:2018:CBF

[MM18] Christian Mullon and Charles Mullon. A constraint-based framework to study competition and cooperation in fishing. Fisheries Research, 203(??):74–83, July 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617302187.

Morson:2018:DDC

[MMAA^{+}18] Jason M. Morson, Daphne M. Munroe, Kathy A. Ashton-Alcox, Eric N. Powell, David Bushek, and Jennifer Gius. Density-dependent capture efficiency of a survey dredge and its influence on the stock assessment of eastern oysters (Crassostrea virginica) in Delaware Bay. Fisheries Research, 205(??):115–121, September 2018. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618301127.

Meynecke:2015:PPR

[MMAH15] Jan-Olaf Meynecke, John Mayze, and Hilke Alberts-Hubatsch. Performance and physiological responses of combined t-bar and PIT tagged giant mud crabs (Scylla serrata). Fisheries Research, 170(??):212–216, October 2015.

CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783615001939.

Maxwell:2013:ASS

[MMBD13] Kerry E. Maxwell, Thomas R. Matthews, Rodney D. Bertelsen, and Charles D. Derby. Age and size structure of Caribbean spiny lobster, Panulirus argus, in a notake marine reserve in the Florida keys, USA. Fisheries Research, 144(??):84–90, July 2013. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612003530.

Maggs:2013:CLN

[MMC13] J. Q. Maggs, B. Q. Mann, and P. D. Cowley. Contribution of a large no-take zone to the management of vulnerable reef fishes in the South-West Indian Ocean. Fisheries Research, 144(??):38–47, July 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612002998.

Moreno:2011:BDC

[MMD^{+}11] Carlos A. Moreno, Carlos Molinet, Patricio Díaz, Manuel Díaz, José Codjambassis, and Alejandra Arévalo. Bathymetric distribution of the Chilean red sea urchin (Loxechinus albus, Molina) in the inner seas of northwest Patagonia: Implications for management. Fisheries Research, 110(2):305– 311, July 2011. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578361100169X.

Martins:2018:WFL

[MMDH18] Ivan Machado Martins, Rodrigo Pereira Medeiros, Maikon Di Domenico, and Natalia Hanazaki. What fishers’ local ecological knowledge can reveal about the changes in exploited fish catches. Fisheries Research, 198(??):109–116, February 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783617302837.

Marin-Monroy:2016:RSD

[MMdlP16] Elvia Aída Marín-Monroy and Miguel Ángel Ojeda-Ruiz de la Peña. The role of socioeconomic disaggregated indicators for fisheries management decisions: the case of Magdalena–Almejas Bay, BCS, Mexico. Fisheries Research, 177(??):116–123, May 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616300091.

McGarvey:2016:UBM

[MMF^{+}16] Richard McGarvey, Janet M. Matthews, John E. Feenstra, André E. Punt, and Adrian Linnane. Using bioeconomic modeling to improve a harvest strategy for a quota-based lobster fishery. Fisheries Research, 183(??):549–558, November 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783616301424.

Malca:2017:FLA

[MMF^{+}17] Estrella Malca, Barbara Muhling, James Franks, Alberto García, Jason Tilley, Trika Gerard, Walter Ingram, and John T. Lamkin. The first larval age and growth curve for bluefin tuna (Thunnus thynnus) from the Gulf of Mexico: Comparisons to the Straits of Florida, and the Balearic Sea (Mediterranean). Fisheries Research, 190(??):24–33, June 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783617300280.

Metcalf:2010:QAR

[MMG10] S. J. Metcalf, K. Moyle, and D. J. Gaughan. Qualitative analysis of recreational fisher response and the ecosystem impacts of management strategies in a data-limited situation. Fisheries Research, 106(3):289–297, December 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783610001979.

Manjarres-Martinez:2015:EMS

[MMGEH15] Luis M. Manjarrés-Martínez, Juan C. Gutiérrez-Estrada, and José A. Hernando. Effects of mesh size and towing speed on the multispecies catch rates of historical swept area surveys. Fisheries Research, 164(??):143–152, April 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783614003312.

Manjarres-Martinez:2010:SPT

[MMGEMGS10] Luis M. Manjarrés-Martínez, Juan Carlos Gutiérrez-Estrada, Julio Mazenet-González, and Milagrosa C. Soriguer. Seasonal patterns of three fish species in a Caribbean coastal gill-net fishery: Biologically induced or climate-related aggregations? Fisheries Research, 106(3):358–367, December 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783610002146.

Macbeth:2012:ARP

[MMJ^{+}12] William G. Macbeth, Russell B. Millar, Daniel D. Johnson, Charles A. Gray, Rebecca S. Keech, and Damian Collins. Assessment of relative performance of a square-mesh codend design across multiple vessels in a demersal trawl fishery. Fisheries Research, 134–136(??):29–41, December 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783612002184.

Mbaru:2010:LWR

[MMK10] E. K. Mbaru, C. M. Mlewa, and E. N. Kimani. Length– weight relationship of 39 selected reef fishes in the Kenyan coastal artisanal fishery. Fisheries Research, 106(3):567–569, December 2010. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783610002213.

Mehault:2016:SDN

[MMK16] Sonia Méhault, Fabien Morandeau, and Dorothée Kopp. Survival of discarded Nephrops norvegicus after trawling in the Bay of Biscay. Fisheries Research, 183(??):396–400, November 2016. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783616302235.

Mosca:2016:SPN

[MML^{+}16] Frederic Mosca, Guillaume Matte, Olivier Lerda, Florent Naud, Didier Charlot, Maxence Rioblanc, and Christophe Corbières. Scientific potential of a new 3D multibeam echosounder in fisheries and ecosystem research. Fisheries Research, 178(??):130–141, June 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615301144.

Mele:2010:MGP

[MMM^{+}10] Salvatore Mele, Paolo Merella, David Macias, María J. Gómez, Giovanni Garippa, and Francisco Alemany. Metazoan gill parasites of wild albacore Thunnus alalunga (Bonaterre, 1788) from the Balearic Sea (western Mediterranean) and their use as biological tags. Fisheries Research, 102(3):305–310, March 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783610000068.

Martins:2019:EII

[MMMPP19] Eduardo G. Martins, Vanessa Minke-Martin, David A. Patterson, and Michael Power. The effects of inter- and intrapopulation variability on the use of adipose fin-muscle tissue comparisons for stable isotope studies of sockeye salmon. Fisheries Research, 211(??):148–157, March 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618303205.

Maynou:2013:SSF

[MMNCR^{+}13] Francesc Maynou, Beatriz Morales-Nin, Miguel Cabanellas-Reboredo, Miquel Palmer, Eugenio García, and Antonio María Grau. Small-scale fishery in the Balearic Islands (W Mediterranean): a socio-economic approach. Fisheries Research, 139(??):11–17, March 2013. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612003414.

Munga:2013:SCD

[MMO^{+}13] Cosmas N. Munga, Stephen Mwangi, Harrison Ong’anda, Renison Ruwa, Julius Manyala, Johan C. Groeneveld, Edward Kimani, and Ann Vanreusel. Species composition, distribution patterns and population structure of penaeid shrimps in Malindi-Ungwana Bay, Kenya, based on experimental bottom trawl surveys. Fisheries Research, 147(??):93–102, October 2013. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613001306.

Merillet:2018:SDN

[MMR^{+}18] Laurène Mérillet, Sonia Méhault, Thomas Rimaud, Corentine Piton, Fabien Morandeau, Marie Morfin, and Dorothée Kopp. Survivability of discarded Norway lobster in the bottom trawl fishery of the Bay of Biscay. Fisheries Research, 198(??):24–30, February 2018. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617302953.

Mateos-Molina:2014:MEC

[MMSUAGC14] D. Mateos-Molina, M. T. Schärer-Umpierre, R. S. Appeldoorn, and J. A. García-Charton. Measuring the effectiveness of a Caribbean oceanic island no-take zone with an asymmetrical BACI approach. Fisheries Research, 150(??):1–10, February 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783613002270.

McClenachan:2014:DCS

[MNAA^{+}14] Loren McClenachan, Benjamin P. Neal, Dalal Al-Abdulrazzak, Taylor Witkin, Kara Fisher, and John N. Kittinger. Do community supported fisheries (CSFs) improve sustainability? Fisheries Research, 157(??):62–69, September 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783614000988.

Morales-Nin:2015:HRR

[MNCPMG15] Beatriz Morales-Nin, Federico Cardona-Pons, Francesc Maynou, and Antoni Maria Grau. How relevant are recreational fisheries? Motivation and activity of resident and tourist anglers in Majorca. Fisheries Research, 164(??):45– 49, April 2015. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783614003038.

Morales-Nin:2012:SOM

[MNGPM^{+}12] B. Morales-Nin, A. J. Geffen, S. Pérez-Mayol, M. Palmer, R. González-Quirós, and A. Grau. Seasonal and ontogenic migrations of meagre (Argyrosomus regius) determined by otolith geochemical signatures. Fisheries Research, 127–128(??):154–165, September 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612000926.

Morales-Nin:2010:MOA

[MNGPMdMG10] Beatriz Morales-Nin, Amalia Grau, S. Pérez-Mayol, and María del Mar Gil. Marking of otoliths, age validation and growth of Argyrosomus regius juveniles (Sciaenidae). Fisheries Research, 106(1):76–80, October 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361000161X.

Marannu:2017:OMA

[MNT^{+}17] Savitri Marannu, Mitsuhiro Nakaya, Tetsuya Takatsu, Shin ichi Takabatake, Mikimasa Joh, and Yutaro Suzuki. Otolith microstructure of arabesque greenling Pleurogrammus azonus: a species with a long embryonic period. Fisheries Research, 194(??):129–134, October 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617301492.

Mireles:2012:CAI

[MNW12] Carlos Mireles, Royden Nakamura, and Dean E. Wendt. A collaborative approach to investigate site fidelity, home range, and homing behavior of cabezon (Scorpaenichthys marmoratus). Fisheries Research, 113(1):133–142, January 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783611003262.

Monnahan:2016:ELW

[MOA^{+}16] Cole C. Monnahan, Kotaro Ono, Sean C. Anderson, Merrill B. Rudd, Allan C. Hicks, Felipe Hurtado-Ferro, Kelli F. Johnson, Peter T. Kuriyama, Roberto R. Licandeo, Christine C. Stawitz, Ian G. Taylor, and Juan L. Valero. The effect of length bin width on growth estimation in integrated age-structured stock assessments. Fisheries Research, 180(??):103–112, August 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615301259.

Ma:2018:PSI

[MOS^{+}18] Hongguang Ma, Tom K. Ogawa, Thomas R. Sminkey, F. Jay Breidt, Virginia M. Lesser, Jean D. Opsomer, John R. Foster, and David A. Van Voorhees. Pilot surveys to improve monitoring of marine recreational fisheries in Hawai’i. Fisheries Research, 204(??):197–208, August 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618300389.

Maunder:2013:RIA

[MP13] Mark N. Maunder and André E. Punt. A review of integrated analysis in fisheries stock assessment. Fisheries Research, 142(??):61–74, May 2013. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612002627.

Mallet:2014:UVT

[MP14] Delphine Mallet and Dominique Pelletier. Underwater video techniques for observing coastal marine biodiversity: a review of sixty years of publications (1952–2012). Fisheries Research, 154(??):44–62, June 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614000356.

Maunder:2017:DDC

[MP17] Mark N. Maunder and Kevin R. Piner. Dealing with data conflicts in statistical inference of population assessment models that integrate information from multiple diverse data sets. Fisheries Research, 192(??):16–27, August 2017. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783616301394.

MacKenzie:2018:EDP

[MP18] Ken MacKenzie and Campbell Pert. Evidence for the decline and possible extinction of a marine parasite species caused by intensive fishing. Fisheries Research, 198(??):63–65, February 2018. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783617302898.

Melia:2012:BAA

[MPA^{+}12] Paco Melià, Mario Petrillo, Giancarlo Albertelli, Alberta Mandich, and Marino Gatto. A bootstrap approach to account for uncertainty in egg production methods applied to small fish stocks. Fisheries Research, 117–118(??):130– 136, April 2012. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783611002001.

Marongiu:2017:TDD

[MPB^{+}17] Martina F. Marongiu, Cristina Porcu, Andrea Bellodi, Rita Cannas, Alessandro Cau, Danila Cuccu, Antonello Mulas, and Maria C. Follesa. Temporal dynamics of demersal chondrichthyan species in the central western Mediterranean Sea: the case study in Sardinia Island. Fisheries Research, 193(??):81–94, September 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617300929.

Mauvisseau:2017:WDQ

[MPF^{+}17] Q. Mauvisseau, M. Parrondo, M. P. Fernández, L. García, J. L. Martínez, E. García-Vázquez, and Y. J. Borrell. On the way for detecting and quantifying elusive species in the sea: the Octopus vulgaris case study. Fisheries Research, 191(??):

41–48, July 2017. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783617300607.

Murphy:2018:RVD

[MPR18] Hannah M. Murphy, Pierre Pepin, and Dominique Robert. Re-visiting the drivers of capelin recruitment in Newfoundland since 1991. Fisheries Research, 200(??):1–10, April 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783617303405.

Montealegre-Quijano:2014:SDS

[MQCS^{+}14] Santiago Montealegre-Quijano, Andrei T. C. Cardoso, Renato Z. Silva, Paul G. Kinas, and Carolus M. Vooren. Sexual development, size at maturity, size at maternity and fecundity of the blue shark Prionace glauca (Linnaeus, 1758) in the Southwest Atlantic. Fisheries Research, 160(??):18–32, December 2014. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578361400085X.

Murua:2010:NEH

[MQGK10] H. Murua, I. Quincoces, D. Garcia, and M. Korta. Is the Northern European hake, Merluccius merluccius, management procedure robust to the exclusion of reproductive dynamics? Fisheries Research, 104(1–3):123–135, May 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783610000822.

Merino:2015:IPM

[MQM^{+}15] Gorka Merino, Antoni Quetglas, Francesc Maynou, Antoni Garau, Haritz Arrizabalaga, Hilario Murua, Josu Santiago, Manuel Barange, Raúl Prellezo, Dorleta García, Jordi Lleonart, Georges Tserpes, Christos Maravelias, Natacha Carvalho, Melanie Austen, Jose A. Fernandes, Pere Oliver, and Antoni Maria Grau. Improving the performance of a Mediterranean demersal fishery toward economic objectives beyond MSY. Fisheries Research, 161(??):131–144, January 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783614001982.

Montealegre-Quijano:2010:DAL

[MQV10] Santiago Montealegre-Quijano and Carolus M. Vooren. Distribution and abundance of the life stages of the blue shark Prionace glauca in the Southwest Atlantic. Fisheries Research, 101(3):168–179, January 15, 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783609002653.

Magnaye:2019:IGP

[MRD19] M. Magnaye, R. M. Rideout, and G. Davoren. Irregular growth patterns in the otoliths of a short-lived forage fish do not reliably indicate reproductive history. Fisheries Research, 218(??):120–126, October 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619301237.

Marquez:2010:UDG

[MREV10] Federico Márquez, Jorgelina Robledo, Gabriela Escati Peñaloza, and Silvina Van der Molen. Use of different geometric morphometrics tools for the discrimination of phenotypic stocks of the striped clam Ameghinomya antiqua (Veneridae) in north Patagonia, Argentina. Fisheries Research, 101(1–2):127–131, January 5, 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783609002641.

Macchi:2018:SFA

[MRLD18] Gustavo J. Macchi, Karina Rodrigues, Ezequiel Leonarduzzi, and Marina V. Diaz. Is the spawning frequency of Argentine hake, Merluccius hubbsi, affected by maternal attributes or physical variables? Fisheries Research, 204(??):147–155, August 2018. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618300390.

McKeown:2015:SSP

[MRS15] Niall J. McKeown, Jean-Paul Robin, and Paul W. Shaw. Species-specific PCR-RFLP for identification of early life history stages of squid and other applications to fisheries research. Fisheries Research, 167(??):207–209, July 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783615000442.

Murphy:2017:IDE

[MRT17] James T. Murphy, Louis J. Rugolo, and Benjamin J. Turnock. Integrating demographic and environmental variables to calculate an egg production index for the Eastern Bering Sea snow crab (Chionoecetes opilio). Fisheries Research, 193(??):143–157, September 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617301054.

Murphy:2018:EAT

[MRT18] James T. Murphy, Louis J. Rugolo, and Benjamin J. Turnock. Estimation of annual, time-varying natural mortality and survival for Eastern Bering Sea snow crab (Chionoecetes opilio) with state-space population models. Fisheries Research, 205(??):122–131, September 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618301000.

Miller:2012:NSB

[MS12] Robert J. Miller and Stephen J. Smith. Nova Scotia’s bloodworm harvest: Assessment, regulation, and governance. Fisheries Research, 113(1):84–93, January 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611002992.

Martell:2014:TDG

[MS14] S. Martell and I. Stewart. Towards defining good practices for modeling time-varying selectivity. Fisheries Research, 158(??):84–95, October 2014. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361300252X.

Maceina:2015:PBE

[MS15a] Michael J. Maceina and Steven M. Sammons. Polychlorinated biphenyl exposure and fish recruitment from 1988 to 2002 in the upper Hudson River, New York, USA. Fisheries Research, 170(??):228–236, October 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615001836.

Morgan:2015:RSD

[MS15b] Alexia C. Morgan and James A. Sulikowski. The role of spiny dogfish in the northeast United States continental shelf ecosystem: How it has changed over time and potential interspecific competition for resources. Fisheries Research, 167(??):260–277, July 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615000946.

Monnahan:2018:EHS

[MS18] Cole C. Monnahan and Ian J. Stewart. The effect of hook spacing on longline catch rates: Implications for catch rate standardization. Fisheries Research, 198(??):150–158, February 2018. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783617302722.

Martin:2017:NAR

[MSCP17] Dustin R. Martin, Daizaburo Shizuka, Christopher J. Chizinski, and Kevin L. Pope. Network analysis of a regional fishery: Implications for management of natural resources, and recruitment and retention of anglers. Fisheries Research, 194(??):31–41, October 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617301340.

Masse:2016:PBS

[MSD^{+}16] J. Massé, F. Sanchez, D. Delaunay, J. M. Robert, and P. Petitgas. A partnership between science and industry for a monitoring of anchovy & sardine in the Bay of Biscay: When fishermen are actors of science. Fisheries Research, 178(??):26–38, June 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615301417.

MacCall:2016:RCN

[MSDT16] Alec D. MacCall, William J. Sydeman, Peter C. Davison, and Julie A. Thayer. Recent collapse of northern anchovy biomass off California. Fisheries Research, 175(??):87–94, March 2016. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783615301363.

Matic-Skoko:2011:AGV

[MSFŠ^{+}11] Sanja Matić-Skoko, Josipa Ferri, Frane Škeljo, Vlasta Bartulović, Katarina Glavić, and Branko Glamuzina. Age, growth and validation of otolith morphometrics as predictors of age in the forkbeard, Phycis phycis (Gadidae). Fisheries Research, 112(1–2):52–58, December 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611002797.

Machado-Schiaffino:2011:PSL

[MSGV11] G. Machado-Schiaffino and E. Garcia-Vazquez. Population structure of long tailed hake Macruronus magellanicus in the Pacific and Atlantic oceans: Implications for fisheries management. Fisheries Research, 111(3):164–169, November 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783611002529.

Moore:2012:SVL

[MSN^{+}12] Bradley R. Moore, Colin A. Simpfendorfer, Stephen J. Newman, Jason M. Stapley, Quentin Allsop, Michelle J. Sellin, and David J. Welch. Spatial variation in life history reveals insight into connectivity and geographic population structure of a tropical estuarine teleost: king threadfin, Polydactylus macrochir. Fisheries Research, 125–126(??):214–224, August 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783612001087.

Morgan:2014:VCP

[MSR14] M. J. Morgan, P. A. Shelton, and R. M. Rideout. Varying components of productivity and their impact on fishing mortality reference points for Grand Bank Atlantic cod and American plaice. Fisheries Research, 155(??):64–73, July 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783614000599.

Marshall:2015:VPR

[MSRB15] Heather Marshall, Gregory Skomal, Paige G. Ross, and Diego Bernal. At-vessel and post-release mortality of the dusky (Carcharhinus obscurus) and sandbar (C. plumbeus) sharks after longline capture. Fisheries Research, 172(??):373–384, December 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783615300205.

Macchi:2014:AFR

[MSRM^{+}14] Gustavo J. Macchi, Fran Saborido-Rey, Hilario Murua, Gabriel Claramunt, and Paulo De Tarso Chaves. Advances in fisheries research in ibero-America. Fisheries Research, 160(??):1–7, December 2014. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614001908.

Malta:2016:LTV

[MSS^{+}16] T. Malta, P. T. Santos, A. M. P. Santos, M. Rufino, and A. Silva. Long-term variations in ibero-Atlantic sardine (Sardina pilchardus) population dynamics: Relation to environmental conditions and exploitation history. Fisheries Research, 179(??):47–56, July 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616300315.

Matsumoto:2014:BST

[MST14] Takayuki Matsumoto, Keisuke Satoh, and Mikio Toyonaga. Behavior of skipjack tuna (Katsuwonus pelamis) associated with a drifting FAD monitored with ultrasonic transmitters in the equatorial central Pacific Ocean. Fisheries Research, 157(??):78–85, September 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614001271.

Montanini:2015:ISI

[MSV^{+}15] Stefano Montanini, Marco Stagioni, Giovanni Valdrè, Stefano Tommasini, and Maria Vallisneri. Intra-specific and inter-specific variability of the sulcus acusticus of sagittal otoliths in two gurnard species (Scorpaeniformes, Triglidae). Fisheries Research, 161(??):93–101, January 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614002173.

Miyamoto:2019:PTP

[MSV^{+}19] Hiroomi Miyamoto, Satoshi Suyama, Dharmamony Vijai, Hideaki Kidokoro, Miyako Naya, Taiki Fuji, and Mitsuo Sakai. Predicting the timing of Pacific saury (Cololabis saira) immigration to Japanese fishing grounds: a new approach based on natural tags in otolith annual rings. Fisheries Research, 209(??):167–177, January 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618302510.

Moore:2017:OCP

[MSWW17] Bradley R. Moore, Jason M. Stapley, Ashley J. Williams, and David. J. Welch. Overexploitation causes profound demographic changes to the protandrous hermaphrodite king threadfin (Polydactylus macrochir) in Queensland’s Gulf of Carpentaria, Australia. Fisheries Research, 187(??):199– 208, March 2017. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578361630399X.

MacCall:2013:HSS

[MT13] Alec D. MacCall and Steven L. H. Teo. A hybrid stock synthesis — virtual population analysis model of Pacific bluefin tuna. Fisheries Research, 142(??):22–26, May 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612001579.

Maunder:2019:MTV

[MT19] Mark N. Maunder and James T. Thorson. Modeling temporal variation in recruitment in fisheries stock assessment: a review of theory and practice. Fisheries Research, 217(??):71–86, September 2019. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618303564.

Major:2017:FAB

[MTC^{+}17] Robert N. Major, David I. Taylor, Stephen Connor, Geoffrey Connor, and Andrew G. Jeffs. Factors affecting bycatch in a developing New Zealand scampi potting fishery. Fisheries Research, 186 (part 1)(??):55–64, February 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616302570.

Madsen:2010:ISB

[MTH10] Niels Madsen, Vesa Tschernij, and Renè Holst. Improving selectivity of the Baltic cod pelagic trawl fishery: Experiments to assess the next step. Fisheries Research, 103(1–3):40–47, April 2010. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610000445.

Marquez:2017:USS

[MTV17] Federico Márquez, M. Magdalena Trivellini, and Silvina Van der Molen. Use of shell shape variation as an assessment tool in the southernmost razor clam fishery. Fisheries Research, 186 (part 1)(??):216–222, February 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361630282X.

Munyandorero:2012:CRM

[Mun12a] Joseph Munyandorero. Corrigendum to “A recruitmentmortality model in the precautionary management toolkit of African tropical inland, single-species fisheries” [Fish.

Res. 127–128 (2012) 26–33]. Fisheries Research, 134–136 (??):134, December 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783612002585.

See [Mun12b].

Munyandorero:2012:RMM

[Mun12b] Joseph Munyandorero. A recruitment-mortality model in the precautionary management toolkit of African tropical inland, single-species fisheries. Fisheries Research, 127–128(??):26–33, September 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/

S0165783612001555. See corrigendum [Mun12a].

Munyandorero:2018:EUC

[Mun18] Joseph Munyandorero. Embracing uncertainty, continual spawning, estimation of the stock–recruit steepness, and size-limit designs with length–based per-recruit analyses for African tropical fisheries. Fisheries Research, 199(??):137– 157, March 2018. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783617303259.

Mortensen:2017:EFD

[MUO^{+}17] Lars O. Mortensen, Clara Ulrich, Hans Jakob Olesen, Heidrikur Bergsson, Casper W. Berg, Nikolaos Tzamouranis, and Jørgen Dalskov. Effectiveness of fully documented fisheries to estimate discards in a participatory research scheme. Fisheries Research, 187(??):150–157, March 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616303952.

Murray:2011:IBS

[Mur11] Kimberly T. Murray. Interactions between sea turtles and dredge gear in the U.S. sea scallop (Placopecten magellanicus) fishery, 2001–2008. Fisheries Research, 107(1–3):137–146, January 2011. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783610002766.

Murray:2015:ILO

[Mur15] Kimberly T. Murray. The importance of location and operational fishing factors in estimating and reducing loggerhead turtle (Caretta caretta) interactions in U.S. bottom trawl gear. Fisheries Research, 172(??):440–451, December 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783615300588.

Murphy:2019:PIS

[Mur19] James T. Murphy. Patterns and implications of skipmolting for the Eastern Bering Sea snow and Tanner crab (Chionoecetes opilio and C bairdi). Fisheries Research, 210(??):63–70, February 2019. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618302753.

Minte-Vera:2016:GFI

[MVMCC16] Carolina V. Minte-Vera, Mark N. Maunder, John M. Casselman, and Steven E. Campana. Growth functions that incorporate the cost of reproduction. Fisheries Research, 180(??):31–44, August 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615301193.

Minte-Vera:2017:GBR

[MVMdS^{+}17] Carolina V. Minte-Vera, Mark N. Maunder, Alexandre M. Aires da Silva, Keisuke Satoh, and Koji Uosaki. Get the biology right, or use size-composition data at your own risk. Fisheries Research, 192(??):114–125, August 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617300231.

Magee:2018:CTJ

[MVML18] Christopher Magee, Michelle Voyer, Alistair McIlgorm, and Owen Li. Chasing the thrill or just passing the time? Trialing a new mixed methods approach to understanding heterogeneity amongst recreational fishers based on motivations. Fisheries Research, 199(??):107–118, March 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361730334X.

Minte-Vera:2019:IMS

[MVMSdS19] Carolina V. Minte-Vera, Mark N. Maunder, Kurt M. Schaefer, and Alexandre M. Aires da Silva. The influence of metrics for spawning output on stock assessment results and evaluation of reference points: an illustration with yellowfin tuna in the eastern Pacific Ocean. Fisheries Research, 217(??):35–45, September 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618302637.

Maunder:2011:AEN

[MW11] Mark N. Maunder and Richard A. Wong. Approaches for estimating natural mortality: Application to summer flounder (Paralichthys dentatus) in the U.S. mid-Atlantic. Fisheries Research, 111(1–2):92–99, September 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611002335.

Methot:2013:SSB

[MW13] Richard D. Methot, Jr. and Chantell R. Wetzel. Stock synthesis: a biological and statistical framework for fish stock assessment and fishery management. Fisheries Research, 142(??):86–99, May 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783612003293.

Midway:2015:STV

[MWA^{+}15] Stephen R. Midway, Tyler Wagner, Stephen A. Arnott, Patrick Biondo, Fernando Martinez-Andrade, and Thomas F. Wadsworth. Spatial and temporal variability in growth of southern flounder (Paralichthys lethostigma). Fisheries Research, 167(??):323–332, July 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783615000995.

Moraga:2015:DLC

[MWC15] Andrew. D. Moraga, Alexander D. M. Wilson, and Steven J. Cooke. Does lure colour influence catch per unit effort, fish capture size and hooking injury in angled largemouth bass? Fisheries Research, 172(??):1–6, December 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615001903.

Midway:2013:IMM

[MWR^{+}13a] Stephen R. Midway, J. Wilson White, William Roumillat, Christopher Batsavage, and Frederick S. Scharf. Improving macroscopic maturity determination in a pre-spawning flatfish through predictive modeling and whole mount methods. Fisheries Research, 147(??):359–369, October 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783613001744.

Morgan:2013:EAT

[MWR13b] M. Joanne Morgan, Peter J. Wright, and Rick M. Rideout. Effect of age and temperature on spawning time in two gadoid species. Fisheries Research, 138(??):42–51, February 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783612000999.

Martins:2018:SRP

[MWR18] Camila L. Martins, Terence I. Walker, and Richard D. Reina. Stress-related physiological changes and post-release survival of elephant fish (Callorhinchus milii) after longlining, gillnetting, angling and handling in a controlled setting. Fisheries Research, 204(??):116–124, August 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618300237.

Martin:2013:RBC

[MWSN^{+}13] Marcelo A. San Martín, Rodrigo Wiff, J. C. Saavedra-Nievas, Luis A. Cubillos, and Sergio Lillo. Relationship between Chilean hake (Merluccius gayi gayi) abundance and environmental conditions in the central-southern zone of Chile. Fisheries Research, 143(??):89–97, June 2013. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783613000131.

Myers:2015:SSC

[MYJ^{+}15] Jared T. Myers, Daniel L. Yule, Michael L. Jones, Tyler D. Ahrenstorff, Thomas R. Hrabik, Randall M. Claramunt, Mark P. Ebener, and Eric K. Berglund. Spatial synchrony in cisco recruitment. Fisheries Research, 165(??):11–21, May 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783615000028.

Ma:2018:CAF

[MZC^{+}18] Tingting Ma, Wei Zhou, Like Chen, Longhua Wu, Christie Peter, and Guibin Dai. Concerns about the future of Chinese fisheries based on illegal, unreported and unregulated fishing on the Hanjiang River. Fisheries Research, 199(??):212–217, March 2018. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783617303119.

Nakatsuka:2017:BPP

[Nak17a] Shuya Nakatsuka. Best practices for providing scientific recommendations in regional fisheries management organizations: Lessons from bluefin tunas. Fisheries Research, 195(??):194–201, November 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617302023.

Nakatsuka:2017:MSE

[Nak17b] Shuya Nakatsuka. Management strategy evaluation in regional fisheries management organizations — how to promote robust fisheries management in international settings. Fisheries Research, 187(??):127–138, March 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616304039.

Nielsen:2017:FMV

[NAR^{+}17] Max Nielsen, Peder Andersen, Lars Ravensbeck, Frederik Laugesen, Dadi Már Kristófersson, and Hans Ellefsen. Fisheries management and the value chain: the Northeast Atlantic pelagic fisheries case. Fisheries Research, 186 (part 1)(??):36–47, February 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616302569.

Nielsen:2014:ETV

[NB14] Anders Nielsen and Casper W. Berg. Estimation of timevarying selectivity in stock assessments using state-space models. Fisheries Research, 158(??):96–101, October 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783614000228.

Nielsen:2019:RCI

[NB19] Anders Nielsen and Casper W. Berg. Response to: Comments on incongruous formulations in the SAM (state-space assessment model) model and consequences for fish stock assessment. Fisheries Research, 210(??):228–229, February 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783618302431.

Nielsen:2019:PPD

[NBF^{+}19] Kåre N. Nielsen, Alan R. Baudron, Niall G. Fallon, Paul G. Fernandes, Mika Rahikainen, and Michaela Aschan. Participatory planning and decision support for ecosystem based fisheries management of the west coast of Scotland. Fisheries Research, 211(??):59–68, March 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618302893.

Nowara:2017:DAS

[NBG^{+}17] G. B. Nowara, P. Burch, N. Gasco, D. C. Welsford, T. D. Lamb, C. Chazeau, G. Duhamel, P. Pruvost, S. Wotherspoon, and S. G. Candy. Distribution and abundance of skates (Bathyraja spp.) on the Kerguelen Plateau through the lens of the toothfish fisheries. Fisheries Research, 186 (part 1)(??):65–81, February 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361630234X.

Newman:2015:CMS

[NBS15] David Newman, Jim Berkson, and Lisa Suatoni. Current methods for setting catch limits for data-limited fish stocks in the United States. Fisheries Research, 164(??):86–93, April 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783614003117.

Nyboer:2013:MHR

[NC13] Elizabeth A. Nyboer and Lauren J. Chapman. Movement and home range of introduced Nile perch (Lates niloticus) in Lake Nabugabo, Uganda: Implications for ecological divergence and fisheries management. Fisheries Research, 137(??):18–29, January 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361200238X.

Natale:2015:DSS

[NCP15] Fabrizio Natale, Natacha Carvalho, and Anton Paulrud. Defining small-scale fisheries in the EU on the basis of their operational range of activity the Swedish fleet as a case study. Fisheries Research, 164(??):286–292, April 2015. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783614003695.

Needle:2015:UST

[Nee15] Coby L. Needle. Using self-testing to validate the SUR-BAR survey-based assessment model. Fisheries Research, 171(??):78–86, November 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783615000910.

Nicola:2018:LGC

[NEJ^{+}18] Graciela G. Nicola, Benigno Elvira, Bror Jonsson, Daniel Ayllón, and Ana Almodóvar. Local and global climatic drivers of Atlantic salmon decline in southern Europe. Fisheries Research, 198(??):78–85, February 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617302874.

Noack:2017:CSC

[NFK^{+}17] Thomas Noack, Rikke P. Frandsen, Ludvig A. Krag, Bernd Mieske, and Niels Madsen. Codend selectivity in a commercial Danish anchor seine. Fisheries Research, 186 (part 1)(??):283–291, February 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616303344.

Narvarte:2011:ADE

[NGMA11] Maite Narvarte, Raúl González, Alonso Medina, and María Soledad Avaca. Artisanal dredges as efficient and rationale harvesting gears in a Patagonian mussel fishery. Fisheries Research, 111(1–2):108–115, September 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611002359.

Ndjaula:2013:EBE

[NGvdL^{+}13] Hilkka O. N. Ndjaula, Kenneth G. Gerow, Carl D. van der Lingen, Coleen L. Moloney, and Astrid Jarre. Establishing a baseline for evaluating changes in body condition and population dynamics of sardine (Sardinops sagax) in the southern Benguela ecosystem. Fisheries Research, 147(??):253–263, October 2013. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783613001586.

Nijman:2015:TSF

[Nij15] Vincent Nijman. Targeted shark fisheries, DNA barcoding and regulating shark fin trade in Indonesia. Fisheries Research, 168(??):33–34, August 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361500106X.

Nichol:2019:AYS

[NKW^{+}19] Daniel G. Nichol, Stan Kotwicki, Thomas K. Wilderbuer, Robert R. Lauth, and James N. Ianelli. Availability of yellowfin sole Limanda aspera to the eastern Bering Sea trawl survey and its effect on estimates of survey biomass. Fisheries Research, 211(??):319–330, March 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618303291.

Neira:2011:DBS

[NL11] Francisco J. Neira and Jeremy M. Lyle. DEPM-based spawning biomass of Emmelichthys nitidus (Emmelichthyidae) to underpin a developing mid-water trawl fishery in south-eastern Australia. Fisheries Research, 110(2):236– 243, July 2011. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783611001585.

Norman-Lopez:2013:LPP

[NLPS^{+}13] Ana Norman-López, Éva Plagányi, Tim Skewes, Elvira Poloczanska, Darren Dennis, Mark Gibbs, and Peter Bayliss. Linking physiological, population and socio-economic assessments of climate-change impacts on fisheries. Fisheries Research, 148(??):18–26, November 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612001063.

Nam:2011:ICS

[NMH^{+}11] Nguyen Thanh Nam, Todd William Miller, Nguyen Xuan Huan, Vu Trung Tang, and Koji Omori. Integrating community structure and stable isotope analysis to assess a heavily exploited coastal marine ecosystem off Central Vietnam. Fisheries Research, 110(2):268–276, July 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611001639.

Newman:2012:ABD

[NML12] Stephen J. Newman, Michael C. Mackie, and Paul D. Lewis. Age-based demography and relative fisheries productivity of Spanish mackerel, Scomberomorus commerson (Lacepede) in Western Australia. Fisheries Research, 129–130(??):46–60, October 2012. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612001798.

Nurdin:2017:ARS

[NMLZ17] Suhartono Nurdin, Muzzneena A. Mustapha, Tukimat Lihan, and Mukti Zainuddin. Applicability of remote sensing oceanographic data in the detection of potential fishing grounds of Rastrelliger kanagurta in the archipelagic waters of Spermonde, Indonesia. Fisheries Research, 196(??):1–12, December 2017. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783617302126.

Nevarez-Martinez:2010:PDJ

[NMMBCV^{+}10] Manuel O. Nevárez-Martínez, Enrique Morales-Bojórquez, Celio Cervantes-Valle, J. Pablo Santos-Molina, and Juana López-Martínez. Population dynamics of the jumbo squid (Dosidicus gigas) in the 2002–2008 fishing seasons off Guaymas, Mexico. Fisheries Research, 106(2):132–140, November 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783610001803.

Nissling:2017:EBF

[NNP17] Anders Nissling, Sofia Nyberg, and Christoph Petereit. Egg buoyancy of flounder, Platichthys flesus, in the Baltic Sea — adaptation to salinity and implications for egg survival. Fisheries Research, 191(??):179–189, July 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617300577.

Ndjaula:2010:LTC

[NNS^{+}10] Hilkka O. N. Ndjaula, Richard D. M. Nash, Aril Slotte, Arne Johannessen, and Olav Sigurd Kjesbu. Long-term changes in the total egg production of Norwegian springspawning herring Clupea harengus (L.) — implications of variations in population structure and condition factor. Fisheries Research, 104(1–3):19–26, May 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610000895.

Nogueira:2017:DGA

[NPGT17] Adriana Nogueira, Xabier Paz, and Diana González-Troncoso. Demersal groundfish assemblages and depthrelated trends on Flemish Cap (NAFO division 3M): 2004– 2013. Fisheries Research, 186 (part 1)(??):192–204, February 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783616302715.

Nash:2010:IMA

[NPK^{+}10] Richard D. M. Nash, Graham M. Pilling, Laurence T. Kell, Pieter-Jan Schön, and Olav Sigurd Kjesbu. Investment in maturity-at-age and -length in northeast Atlantic cod stocks. Fisheries Research, 104(1–3):89–99, May 2010. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783610000652.

Nogueira:2018:CPC

[NPRGTSR18] Adriana Nogueira, Alfonso Pérez-Rodríguez, Diana González-Troncoso, and Fran Saborido-Rey. Could population and community indicators contribute to identify the driver factors and describe the dynamic in the Flemish Cap demersal assemblages? Fisheries Research, 204(??):33–40, August 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783618300274.

Niklitschek:2016:DDR

[NS16] E. J. Niklitschek and G. Skaret. Distribution, density and relative abundance of Antarctic krill estimated by maximum likelihood geostatistics on acoustic data collected during commercial fishing operations. Fisheries Research, 178(??):114–121, June 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361530093X.

Neves:2018:AEI

[NSM^{+}18] Vanessa Neves, Diogo Silva, Filipe Martinho, Carlos Antunes, Sandra Ramos, and Vânia Freitas. Assessing the effects of internal and external acoustic tagging methods on European flounder Platichthys flesus. Fisheries Research, 206(??):202–208, October 2018. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618301498.

Noack:2019:GPC

[NSM^{+}19] Thomas Noack, Daniel Stepputtis, Niels Madsen, Kai Wieland, Stefanie Haase, and Ludvig Ahm Krag. Gear performance and catch process of a commercial Danish anchor seine. Fisheries Research, 211(??):204–211, March 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783618303242.

Nyamweya:2017:PLV

[NST^{+}17] Chrispine Sangara Nyamweya, Erla Sturludottir, Tumi Tomasson, Anthony Taabu-Munyaho, Murithi Njiru, and Gunnar Stefansson. Prediction of Lake Victoria’s response to varied fishing regimes using the Atlantis ecosystem model. Fisheries Research, 194(??):76–83, October 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617301418.

Neves:2018:IRS

[NVS^{+}18] Ana Neves, Ana Rita Vieira, Vera Sequeira, Rafaela Barros Paiva, and Leonel Serrano Gordo. Insight on reproductive strategy in Portuguese waters of a commercial protogynous species, the black seabream Spondyliosoma cantharus (Sparidae). Fisheries Research, 206(??):85–95, October 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783618301401.

Neitzel:2017:RIB

[NvZH^{+}17] Sophie M. Neitzel, Paul A. M. van Zwieten, Astrid Hendriksen, Deirdre Duggan, and Simon R. Bush. Returning information back to fishers: Graphical and numerical literacy of small-scale Indonesian tuna fishers. Fisheries Research, 196(??):96–105, December 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617302199.

Nguyen:2014:UOB

[NWL^{+}14] Truong X. Nguyen, Paul D. Winger, George Legge, Earl G. Dawe, and Darrell R. Mullowney. Underwater observations of the behaviour of snow crab (Chionoecetes opilio) encountering a shrimp trawl off northeast Newfoundland. Fisheries Research, 156(??):9–13, August 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614001489.

Nguyen:2015:CSF

[NWO^{+}15] Truong X. Nguyen, Paul D. Winger, Dave Orr, George Legge, Harold Delouche, and Alex Gardner. Computer simulation and flume tank testing of scale engineering models: How well do these techniques predict full-scale at-sea performance of bottom trawls? Fisheries Research, 161(??):217–225, January 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614002525.

Noriega:2011:TAC

[NWS^{+}11] Rocio Noriega, Jonathan Mark Werry, Wayne Sumpton, David Mayer, and Shing Yip Lee. Trends in annual CPUE and evidence of sex and size segregation of Sphyrna lewini: Management implications in coastal waters of northeastern Australia. Fisheries Research, 110(3):472–477, August 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783611002220.

Ojaveer:2011:IED

[OALS11] E. Ojaveer, T. Arula, A. Lankov, and H. Shpilev. Impact of environmental deviations on the larval and year-class abundances in the spring spawning herring (Clupea harengus membras L.) of the Gulf of Riga (Baltic Sea) in 1947– 2004. Fisheries Research, 107(1–3):159–168, January 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578361000281X.

Ospina-Alvarez:2012:CEB

[OÁPP12] Andrés Ospina-Álvarez, Isabel Palomera, and Carolina Parada. Changes in egg buoyancy during development and its effects on the vertical distribution of anchovy eggs. Fisheries Research, 117–118(??):86–95, April 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611000506.

Ofstad:2013:AGA

[OAPS13] L. H. Ofstad, C. Angus, T. Pedersen, and P. Steingrund. Age and growth of anglerfish (Lophius piscatorius) in Faroese waters. Fisheries Research, 139(??):51–60, March 2013. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783612001671.

Okochi:2016:RBF

[OAT^{+}16] Yumi Okochi, Osamu Abe, Sho Tanaka, Yukio Ishihara, and Akio Shimizu. Reproductive biology of female Pacific bluefin tuna, Thunnus orientalis, in the Sea of Japan. Fisheries Research, 174(??):30–39, February 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615300667.

ODriscoll:2018:FSE

[OCL^{+}18] Richard L. O’Driscoll, Simonepietro Canese, Yoann Ladroit, Steve J. Parker, Laura Ghigliotti, Sophie Mormede, and Marino Vacchi. First in situ estimates of acoustic target strength of Antarctic toothfish (Dissostichus mawsoni). Fisheries Research, 206(??):79–84, October 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618301449.

Oyugi:2011:LHT

[OCN^{+}11] Dalmas O. Oyugi, Julien Cucherousset, Micheni J. Ntiba, Seth M. Kisia, David M. Harper, and J. Robert Britton. Life history traits of an equatorial common carp Cyprinus carpio population in relation to thermal influences on invasive populations. Fisheries Research, 110(1):92–97, June 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783611001263.

ODriscoll:2016:IAS

[ODD16] Richard L. O’Driscoll, Adam J. Dunford, and Alistair Dunn. Industry acoustic surveys of spawning southern blue whiting on the Bounty Platform, New Zealand. Fisheries Research, 178(??):61–70, June 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615001484.

Otxotorena:2010:EAG

[ODdA^{+}10] U. Otxotorena, G. Díez, E. López de Abechuco, M. Santurtún, and P. Lucio. Estimation of age and growth of juvenile hakes (Merluccius merluccius Linnaeus, 1758) of the Bay of Biscay and Great Sole by means of the analysis of macro and microstructure of the otoliths. Fisheries Research, 106(3):337–343, December 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610002092.

Oyafuso:2017:HBS

[ODMF17] Zack. S. Oyafuso, Jeffrey C. Drazen, Cordelia H. Moore, and Erik C. Franklin. Habitat-based species distribution modelling of the Hawaiian deepwater snapper-grouper complex. Fisheries Research, 195(??):19–27, November 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783617301662.

Ozbilgin:2015:SSH

[ÖEG^{+}15] Hüseyin Özbilgin, Ahmet Raif Eryaşar, Gökhan Gökçe, Yeliz Doğanyılmaz Özbilgin, Adem Sezai Bozaoğlu, Ebrucan Kalecik, and Bent Herrmann. Size selectivity of hand and machine woven codends and short term commercial loss in the Northeastern Mediterranean. Fisheries Research, 164(??):73–85, April 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614003166.

Ortega-Garcia:2015:EEF

[OGCBMR15] Sofía Ortega-García, Edgardo Camacho-Bareño, and Raúl O. Martínez-Rincón. Effects of environmental factors on the spatio-temporal distribution of striped marlin catch rates off Cabo San Lucas, Baja California Sur, Mexico. Fisheries Research, 166(??):47–58, June 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783614002744.

Ogle:2017:ABS

[Ogl17] Derek H. Ogle. An algorithm for the von Bertalanffy seasonal cessation in growth function of Pauly et al. (1992). Fisheries Research, 185(??):1–5, January 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616303137.

Onate-Gonzalez:2017:IBS

[OGSNH^{+}17] Erick C. Oñate-González, Oscar Sosa-Nishizaki, Sharon Z. Herzka, Christopher G. Lowe, Kady Lyons, Omar Santana-Morales, Chugey Sepulveda, César Guerrero-Ávila, Emiliano García-Rodríguez, and John B. O’Sullivan. Importance of Bahia Sebastian Vizcaino as a nursery area for white sharks (Carcharodon carcharias) in the Northeastern Pacific: a fishery dependent analysis. Fisheries Research, 188(??):125–137, April 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616304337.

Olale:2012:DID

[OH12] Edward Olale and Spencer Henson. Determinants of income diversification among fishing communities in Western Kenya. Fisheries Research, 125–126(??):235–242, August 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578361200121X.

Omori:2016:EUC

[OHLBL16] Kristen L. Omori, John M. Hoenig, Mark A. Luehring, and Kathy Baier-Lockhart. Effects of underestimating catch and effort on surplus production models. Fisheries Research, 183(??):138–145, November 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616301588.

Olsen:2019:EGS

[OHSG19] Leonore Olsen, Bent Herrmann, Manu Sistiaga, and Eduardo Grimaldo. Effect of gear soak time on size selection in the snow crab pot fishery. Fisheries Research, 214(??):157– 165, June 2019. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783619300426.

Olsen:2014:QCB

[OJT^{+}14] Stein Harris Olsen, Sjurdur Joensen, Torbjørn Tobiassen, Karsten Heia, Leif Akse, and Heidi Nilsen. Quality consequences of bleeding fish after capture. Fisheries Research, 153(??):103–107, May 2014. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614000198.

Okemwa:2016:CCS

[OKAKO16] G. M. Okemwa, B. Kaunda-Arara, E. N. Kimani, and B. Ogutu. Catch composition and sustainability of the marine aquarium fishery in Kenya. Fisheries Research, 183(??):19–31, November 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783616301333.

ONeill:2016:ITT

[OKB^{+}16] F. G. O’Neill, R. J. Kynoch, L. Blackadder, R. J. Fryer, A. R. Eryaşar, E. Notti, and A. Sala. The influence of twine tenacity, thickness and bending stiffness on codend selectivity. Fisheries Research, 176(??):94–99, April 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615301648.

Ovegaard:2011:SSC

[OKPL11] M. Ovegård, S. Königson, A. Persson, and S. G. Lunneryd. Size selective capture of Atlantic cod (Gadus morhua) in floating pots. Fisheries Research, 107(1–3):239–244, January 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783610002900.

ONeill:2014:STE

[OLK^{+}14] F. G. O’Neill, E. K. Lines, R. J. Kynoch, R. J. Fryer, and S. Maguire. A short-term economic assessment of incentivised selective gears. Fisheries Research, 157(??):13–23, September 2014. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783614000927.

Orue:2019:FSD

[OLM^{+}19] Blanca Orue, Jon Lopez, Gala Moreno, Josu Santiago, Guillermo Boyra, Jon Uranga, and Hilario Murua. From fisheries to scientific data: a protocol to process information from fishers’ echo-sounder buoys. Fisheries Research, 215(??):38–43, July 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783619300633.

Olaussen:2015:CVH

[OLS15] Jon Olaf Olaussen, Yajie Liu, and Anders Skonhoft. Conservation versus harvest of wild Atlantic salmon. The cost of sea lice induced mortality. Fisheries Research, 168(??):63– 71, August 2015. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783615001125.

Ozgul:2019:HRR

[ÖLTA19] Aytaç Özgül, Altan Lök, T. Tansel Tanrıkul, and Josep Alós. Home range and residency of Scorpaena porcus and Scorpaena scrofa in artificial reefs revealed by fine-scale acoustic tracking. Fisheries Research, 210(??):22–30, February 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783618302777.

Olsen:2018:IVL

[OMB18] Zachary Olsen, Dusty McDonald, and Britt Bumguardner. Intraspecific variation in life history strategies and implications for management: a case study of black drum (Pogonias cromis) in the Upper Laguna Madre, Texas USA. Fisheries Research, 207(??):55–62, November 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618301802.

Okamura:2014:ESB

[OMI^{+}14] Hiroshi Okamura, Murdoch K. McAllister, Momoko Ichinokawa, Lynne Yamanaka, and Kendra Holt. Evaluation of the sensitivity of biological reference points to the spatiotemporal distribution of fishing effort when seasonal migrations are sex-specific. Fisheries Research, 158(??):116–123, October 2014. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783613002506.

Osio:2015:AVM

[OOM15] Giacomo Chato Osio, Alessandro Orio, and Colin P. Millar. Assessing the vulnerability of Mediterranean demersal stocks and predicting exploitation status of un-assessed stocks. Fisheries Research, 171(??):110–121, November 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783615000429.

Olsen:2012:PRM

[OOTV12] Rolf Erik Olsen, Frode Oppedal, Maria Tenningen, and Aud Vold. Physiological response and mortality caused by scale loss in Atlantic herring. Fisheries Research, 129–130(??):21–27, October 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783612001804.

Ottosen:2017:MPF

[OPE^{+}17] Katharina Maj Ottosen, Martin Wæver Pedersen, Sólvá Káradottir Eliasen, Petur Steingrund, Eydfinn Magnussen, and Till Andreas Soya Rasmussen. Migration patterns of the Faroe Plateau cod (Gadus morhua, L.) revealed by data storage tags. Fisheries Research, 195(??):37–45, November 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783617301698.

Oberg:2015:CDE

[OPF15] E. W. Oberg, K. O. Perez, and L. A. Fuiman. Carbon dioxide is an effective anesthetic for multiple marine fish species. Fisheries Research, 165(??):22–27, May 2015. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783615000077.

Ono:2015:TOG

[OPH15] Kotaro Ono, André E. Punt, and Ray Hilborn. Think outside the grids: an objective approach to define spatial strata for catch and effort analysis. Fisheries Research, 170(??):89–101, October 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615001708.

OBryhim:2017:FSI

[OPL17] Jason R. O’Bryhim, E. C. M. Parsons, and Stacey L. Lance. Forensic species identification of elasmobranch products sold in Costa Rican markets. Fisheries Research, 186 (part 1)(??):144–150, February 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616302752.

Ono:2012:MPA

[OPR12] Kotaro Ono, André E. Punt, and Etienne Rivot. Model performance analysis for Bayesian biomass dynamics models using bias, precision and reliability metrics. Fisheries Research, 125–126(??):173–183, August 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612001026.

Ortiz:2011:RCR

[ORMG11] Nicolás Ortiz, María E. Ré, Federico Márquez, and Nora G. Glembocki. The reproductive cycle of the red octopus Enteroctopus megalocyathus in fishing areas of Northern Patagonian coast. Fisheries Research, 110(1):217–223, June 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783611001251.

Ormseth:2011:AVA

[OS11] Olav A. Ormseth and Paul D. Spencer. An assessment of vulnerability in Alaska groundfish. Fisheries Research, 112(3):127–133, December 2011. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611000816.

OBoyle:2012:SCI

[OS12] Robert O’Boyle and Michael Sinclair. Seal–cod interactions on the Eastern Scotian Shelf: Reconsideration of modelling assumptions. Fisheries Research, 115–116(??):1–13, March 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783611003249.

ONeill:2019:ICL

[OS19] F. G. O’Neill and K. Summerbell. The influence of continuous lines of light on the height at which fish enter demersal trawls. Fisheries Research, 215(??):131–142, July 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578361930075X.

Ottosen:2018:DTS

[OSMP18] Katharina Maj Ottosen, Petur Steingrund, Eydfinn Magnussen, and Mark R. Payne. Distribution and timing of spawning Faroe Plateau cod in relation to warming spring temperatures. Fisheries Research, 198(??):14–23, February 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783617302941.

Ohshimo:2018:LTC

[OSO^{+}18] Seiji Ohshimo, Takuya Sato, Yumi Okochi, Yukio Ishihara, Atsushi Tawa, Masanori Kawazu, Yuko Hiraoka, Hiroshi Ashida, and Nobuaki Suzuki. Long-term change in reproductive condition and evaluation of maternal effects in Pacific bluefin tuna, Thunnus orientalis, in the Sea of Japan. Fisheries Research, 204(??):390–401, August 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618300912.

Olsen:2013:CIS

[OTA^{+}13] Stein Harris Olsen, Torbjørn Tobiassen, Leif Akse, Tor H. Evensen, and Kjell Ø. Midling. Capture induced stress and live storage of Atlantic cod (Gadus morhua) caught by trawl: Consequences for the flesh quality. Fisheries Research, 147(??):446–453, October 2013. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613000775.

Oshima:2019:DRB

[OTH^{+}19] Tatsuki Oshima, Yoshimi Takao, Seizo Hasegawa, Takuto Kimura, Takayoshi Uehara, and Ippei Fusejima. Differences in reaction of bigeye tuna (Thunnus obesus) and skipjack tuna (Katsuwonus pelamis) to intermittent light. Fisheries Research, 214(??):148–156, June 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619300402.

Oesterwind:2010:BMS

[OtHH^{+}10] Daniel Oesterwind, Remment ter Hofstede, Brian Harley, Heinz Brendelberger, and Uwe Piatkowski. Biology and meso-scale distribution patterns of North Sea cephalopods. Fisheries Research, 106(2):141–150, November 2010. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783610001335.

Olin:2016:EGC

[OTK^{+}16] M. Olin, J. Tiainen, M. Kurkilahti, M. Rask, and H. Lehtonen. An evaluation of gillnet CPUE as an index of perch density in small forest lakes. Fisheries Research, 173(??):20–25, January 2016. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783615001678.

Ochavillo:2011:PSC

[OTST11] Domingo Ochavillo, Saolotoga Tofaeono, Marlowe Sabater, and Elizabeth L. Trip. Population structure of Ctenochaetus striatus (Acanthuridae) in Tutuila, American Samoa: the use of size-at-age data in multi-scale population size surveys. Fisheries Research, 107(1–3):14–21, January 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610002493.

Punt:2017:EAM

[PACF17] André E. Punt, Caitlin Allen Akselrud, and Lee Cronin-Fine. The effects of applying mis-specified age- and sizestructured models. Fisheries Research, 188(??):58–73, April 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783616304027.

Perez:2013:EUR

[PAH^{+}13] Juan Cicuendez Perez, Marlene Alvarez Alvarez, Jukka Heikkonen, Jordi Guillen, and Thomas Barbas. The efficiency of using remote sensing for fisheries enforcement: Application to the Mediterranean bluefin tuna fishery. Fisheries Research, 147(??):24–31, October 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613001070.

Pullen:2019:TRM

[PALC19] Christopher E. Pullen, Robert Arlinghaus, Robert J. Lennox, and Steven J. Cooke. Telemetry reveals the movement, fate, and lure-shedding of northern pike (Esox lucius) that break the line and escape recreational fisheries capture. Fisheries Research, 211(??):176–182, March 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618303254.

Pokki:2018:VRS

[PAM^{+}18] H. Pokki, J. Artell, J. Mikkola, P. Orell, and V. Ovaskainen. Valuing recreational salmon fishing at a remote site in Finland: a travel cost analysis. Fisheries Research, 208(??):145–156, December 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S016578361830211X.

Pierre:2012:CTW

[PAR^{+}12] Johanna P. Pierre, Edward R. Abraham, Yvan Richard, John Cleal, and David A. J. Middleton. Controlling trawler waste discharge to reduce seabird mortality. Fisheries Research, 131–133(??):30–38, November 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361200210X.

Paulino:2017:UNV

[PAX^{+}17] Carlos Paulino, Eloy Aroni, Han Xu, Edward Alburqueque, and Hervé Demarcq. Use of nighttime visible images in the study of the spatial and temporal variability of fishing areas of jumbo flying squid (Dosidicus gigas) outside Peruvian EEZ 2004–2015. Fisheries Research, 191(??):144–153, July 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783617300760.

Pegg:2011:EII

[PB11] Josephine Pegg and J. Robert Britton. Effects of inter- and intra-specific competition on the growth rates of juvenile European barbel Barbus barbus used in the stock enhancement of UK fisheries. Fisheries Research, 112(1–2):8–12, December 2011. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783611002724.

Peraltilla:2014:SMS

[PB14] Salvador Peraltilla and Sophie Bertrand. In situ measurements of the speed of Peruvian anchovy schools. Fisheries Research, 149(??):92–94, January 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613002129.

Plaganyi:2011:IUF

[PBB11] Éva Plagányi, Doug Butterworth, and Markus Burgener. Illegal and unreported fishing on abalone — quantifying the extent using a fully integrated assessment model. Fisheries Research, 107(1–3):221–232, January 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610002882.

Piddocke:2015:AVL

[PBB^{+}15a] Toby P. Piddocke, Gavin L. Butler, Paul A. Butcher, Steven W. Purcell, Daniel J. Bucher, and Les Christidis. Age validation in the Lutjanidae: a review. Fisheries Research, 167(??):48–63, July 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615000387.

Porsmoguer:2015:HEM

[PBB^{+}15b] Sebastián Biton Porsmoguer, Daniela Bănaru, Charles F. Boudouresque, Ivan Dekeyser, and Christophe Almarcha. Hooks equipped with magnets can increase catches of blue shark (Prionace glauca) by longline fishery. Fisheries Research, 172(??):345–351, December 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615300254.

Pikitch:2018:SCB

[PBB^{+}18] Ellen K. Pikitch, P. Dee Boersma, Ian L. Boyd, David O. Conover, Philippe Cury, Timothy E. Essington, Selina S. Heppell, Edward D. Houde, Marc Mangel, Daniel Pauly, Eva Plaganyi, Keith Sainsbury, and Robert S. Steneck. The strong connection between forage fish and their predators: a response to Hilborn et al. (2017). Fisheries Research, 198(??):220–223, February 2018. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617302059. See [HAB^{+}17a].

Picciulin:2013:PAM

[PBC^{+}13] M. Picciulin, M. Bolgan, A. Codarin, R. Fiorin, M. Zucchetta, and S. Malavasi. Passive acoustic monitoring of Sciaena umbra on rocky habitats in the Venetian littoral zone. Fisheries Research, 145(??):76–81, August 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613000519.

Perzia:2016:SMG

[PBC^{+}16] Patrizia Perzia, Pietro Battaglia, Pierpaolo Consoli, Franco Andaloro, and Teresa Romeo. Swordfish monitoring by a GIS-based spatial and temporal distribution analysis on harpoon fishery data: a case of study in the central Mediterranean Sea. Fisheries Research, 183(??):424–434, November 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578361630217X.

Powell:2021:EBS

[PBD21] Eric N. Powell, Eleanor A. Bochenek, and John DePersenaire. Evaluation of bag-and-size-limit options in the management of summer flounder Paralichthys dentatus. Fisheries Research, 105(3):215–227, ???? 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610001281.

Pascual:2013:TFP

[PBG^{+}13] M. Pascual, A. Borja, I. Galparsoro, J. Ruiz, E. Mugerza, I. Quincoces, A. Murillas, and L. Arregi. Total fishing pressure produced by artisanal fisheries, from a Marine Spatial Planning perspective: a case study from the Basque Country (Bay of Biscay). Fisheries Research, 147(??):240–252, October 2013. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783613001628.

Pope:2019:CSS

[PBK^{+}19] John G. Pope, Valerio Bartolino, Nataliia Kulatska, Barbara Bauer, Jan Horbowy, Joana P. C. Ribeiro, Erla Sturludottir, and Robert Thorpe. Comparing the steady state results of a range of multispecies models between and across geographical areas by the use of the Jacobian matrix of yield on fishing mortality rate. Fisheries Research, 209(??):259–270, January 2019. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783618302327.

Palmieri:2014:JBN

[PBLT14] Maria Giovanna Palmieri, Alberto Barausse, Tiziana Luisetti, and Kerry Turner. Jellyfish blooms in the Northern Adriatic Sea: Fishermen’s perceptions and economic impacts on fisheries. Fisheries Research, 155(??):51–58, July 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783614000617.

Pierce:2018:ANI

[PBM^{+}18] G. J. Pierce, M. Bao, K. MacKenzie, A. Dunser, L. Giulietti, P. Cipriani, S. Mattiucci, and L. C. Hastie. Ascaridoid nematode infection in haddock (Melanogrammus aeglefinus) and whiting (Merlangius merlangus) in Northeast Atlantic waters. Fisheries Research, 202(??):122–133, June 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783617302564.

Pedreschi:2019:IEA

[PBM^{+}19] Debbi Pedreschi, Paul Bouch, Meadhbh Moriarty, Eugene Nixon, Antony M. Knights, and David G. Reid. Integrated ecosystem analysis in Irish waters; providing the context for ecosystem-based fisheries management. Fisheries Research, 209(??):218–229, January 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618302649.

Parisi-Baradad:2010:ATI

[PBML^{+}10] V. Parisi-Baradad, A. Manjabacas, A. Lombarte, R. Olivella, Ò. Chic, J. Piera, and E. García-Ladona. Automated taxon identification of teleost fishes using an otolith online database — AFORO. Fisheries Research, 105(1):13– 20, June 2010. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783610000524.

Pilling:2016:CBE

[PBR^{+}16] Graham M. Pilling, Aaron M. Berger, Chris Reid, Shelton J. Harley, and John Hampton. Candidate biological and economic target reference points for the south Pacific albacore longline fishery. Fisheries Research, 174(??):167–178, February 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783615300941.

Punt:2019:EIS

[PC19] André E. Punt and Jason M. Cope. Extending integrated stock assessment models to use non-depensatory three-parameter stock-recruitment relationships. Fisheries Research, 217(??):46–57, September 2019. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617301819.

Prellezo:2016:BGV

[PCG16] Raúl Prellezo, Itsaso Carmona, and Dorleta García. The bad, the good and the very good of the landing obligation implementation in the Bay of Biscay: a case study of Basque trawlers. Fisheries Research, 181(??):172–185, September 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783616301291.

Pointin:2018:MAN

[PCP^{+}18] Fabien Pointin, Anne-Sophie Cornou, Romain Prod’homme, Nicolas Taupin, and Marie-Joëlle Rochet. A method to address the non-random spatial distribution of on-board observer data to map landings and discards. Fisheries Research, 199(??):242–251, March 2018. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617302989.

Pope:2016:IAS

[PCWM16] Kevin L. Pope, Christopher J. Chizinski, Christopher L. Wiley, and Dustin R. Martin. Influence of anglers’ specializations on catch, harvest, and bycatch of targeted taxa. Fisheries Research, 183(??):128–137, November 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616301722.

Punt:2014:CMA

[PD14] André E. Punt and Martin Dorn. Comparisons of metaanalytic methods for deriving a probability distribution for the steepness of the stock–recruitment relationship. Fisheries Research, 149(??):43–54, January 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613002257.

Palmer:2018:GGE

[PD18] Michael C. Palmer and Chad Demarest. Getting to good enough: Performance of a suite of methods for spatially allocating fishing effort to management areas. Fisheries Research, 204(??):26–32, August 2018. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618300316.

Politis:2012:ESS

[PDBM12] Philip J. Politis, Joseph T. DeAlteris, Russell W. Brown, and Archie T. Morrison. Effects of sea-state on the physical performance of a survey bottom trawl. Fisheries Research, 123–124(??):26–36, July 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611003468.

Punt:2018:RIA

[PDF^{+}18] André E. Punt, Jemery Day, Gavin Fay, Malcolm Haddon, Neil Klaer, L. Richard Little, Kristin Privitera-Johnson, Anthony D. M. Smith, David C. Smith, Miriana Sporcic, Robin Thomson, Geoffrey N. Tuck, Judy Upston, and Sally Wayte. Retrospective investigation of assessment uncertainty for fish stocks off southeast Australia. Fisheries Research, 198(??):117–128, February 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617302758.

Punt:2016:FSA

[PDH^{+}16] André E. Punt, Catherine M. Dichmont, Rosemary Hurst, Pamela M. Mace, and David C. Smith. Fishery systems in Australia and New Zealand: Success and challenges. Fisheries Research, 183(??):485–487, November 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616302326.

Punt:2011:COE

[PDP^{+}11] André E. Punt, Roy Deng, Sean Pascoe, Catherine M. Dichmont, Shijie Zhou, Éva E. Plagányi, Trevor Hutton, William N. Venables, Rob Kenyon, and Tonya van der Velde. Calculating optimal effort and catch trajectories for multiple species modelled using a mix of sizestructured, delay-difference and biomass dynamics models. Fisheries Research, 109(1):201–211, April 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611000749.

Punt:2017:DWT

[PDS^{+}17] André E. Punt, Roy A. Deng, M. S. M. Siddeek, Rik. C. Buckworth, and Vicki Vanek. Data weighting for tagging data in integrated size-structured models. Fisheries Research, 192(??):94–102, August 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615301636.

Plaganyi:2014:WSF

[PE14] Éva E. Plagányi and Timothy E. Essington. When the SURFs up, forage fish are key. Fisheries Research, 159(??):68–74, November 2014. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783614001647.

Plaza:2011:FRB

[PEAC11] Guido Plaza, Victor Espejo, Viviana Almanza, and Gabriel Claramunt. Female reproductive biology of the silverside Odontesthes regia. Fisheries Research, 111(1–2):31–39, September 2011. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783611002268.

Pena:2018:RCM

[Peñ18] Marian Peña. Robust clustering methodology for multifrequency acoustic data: a review of standardization, initialization and cluster geometry. Fisheries Research, 200 (??):49–60, April 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783617303557.

Pinter:2019:RHR

[PEU19] Kurt Pinter, John Epifanio, and Günther Unfer. Release of hatchery-reared brown trout (Salmo trutta) as a threat to wild populations? A case study from Austria. Fisheries Research, 219(??):Article 105296, November 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619301298.

Pappalardo:2015:DBS

[PF15a] Anna M. Pappalardo and Venera Ferrito. DNA barcoding species identification unveils mislabeling of processed flatfish products in southern Italy markets. Fisheries Research, 164(??):153–158, April 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614003294.

Parsons:2015:RBU

[PF15b] Glenn R. Parsons and Daniel G. Foster. Reducing bycatch in the United States Gulf of Mexico shrimp trawl fishery with an emphasis on red snapper bycatch reduction. Fisheries Research, 167(??):210–215, July 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615000545.

Post:2019:BWD

[PFJ19] Søren Post, Heino O. Fock, and Teunis Jansen. Blue whiting distribution and migration in Greenland waters. Fisheries Research, 212(??):123–135, April 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618303515.

Pita:2016:UTE

[PFVGGM16] Pablo Pita, Duarte Fernández-Vidal, Javier García-Galdo, and Ramón Muíño. The use of the traditional ecological knowledge of fishermen, cost-effective tools and participatory models in artisanal fisheries: Towards the comanagement of common octopus in Galicia (NW Spain). Fisheries Research, 178(??):4–12, June 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615300308.

Postuma:2014:RAT

[PG14] F. A. Postuma and M. A. Gasalla. Reproductive activity of the tropical arrow squid Doryteuthis plei around São Sebastião Island (SE Brazil) based on a 10-year fisheries monitoring. Fisheries Research, 152(??):45–54, April 2014.

CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783613003056.

Pascual:2018:FNP

[PG18] Santiago Pascual and Ángel F. González. The fish nematode problem in major European fish stocks. Fisheries Research, 202(??):1–3, June 2018. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617301157.

Pedrosa-Gerasmio:2015:GDP

[PGAS15] Ivane R. Pedrosa-Gerasmio, Altair B. Agmata, and Mudjekeewis D. Santos. Genetic diversity, population genetic structure, and demographic history of Auxis thazard (Perciformes), Selar crumenophthalmus (Perciformes), Rastrelliger kanagurta (Perciformes) and Sardinella lemuru (Clupeiformes) in Sulu–Celebes Sea inferred by mitochondrial DNA sequences. Fisheries Research, 162(??):64–74, February 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783614002999.

Pascual:2010:CDO

[PGG10] S. Pascual, A. F. González, and A. Guerra. Coccidiosis during octopus senescence: Preparing for parasite outbreak. Fisheries Research, 106(2):160–162, November 2010. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578361000130X.

Perrot:2010:CDB

[PGJ10] Y. Perrot, J. Guillard, and E. Josse. Convergence and divergence between two multibeam sonars (SIMRAD SM20 and RESON SeaBat 6012) used to extract the spatial, morphologic and energy parameters of fish schools. Fisheries Research, 106(3):378–385, December 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361000216X.

Pugh:2013:SDF

[PGL^{+}13] Tracy L. Pugh, Jason S. Goldstein, Kari L. Lavalli, Michael Clancy, and Winsor H. Watson. At-sea determination of female American lobster (Homarus americanus) mating activity: Patterns vs. expectations. Fisheries Research, 147(??):327–337, October 2013. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613001720.

Peterson:2017:NUH

[PGL17] Cassidy D. Peterson, James Gartland, and Robert J. Latour. Novel use of hook timers to quantify changing catchability over soak time in longline surveys. Fisheries Research, 194(??):99–111, October 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617301376.

Pajuelo:2011:PBS

[PGLG11] José G. Pajuelo, Serrana García, José M. Lorenzo, and José A. González. Population biology of the shark, Squalus megalops, harvested in the central-east Atlantic Ocean. Fisheries Research, 108(1):31–41, February 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610003036.

Pereyra:2010:LGD

[PGM^{+}10] Silvana Pereyra, Graciela García, Philip Miller, Sebastián Oviedo, and Andrés Domingo. Low genetic diversity and population structure of the narrownose shark (Mustelus schmitti). Fisheries Research, 106(3):468–473, December 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783610002420.

Pleizier:2015:VWL

[PGP^{+}15] Naomi Pleizier, Lee F. G. Gutowsky, Victor M. Peddemors, Steven J. Cooke, and Paul A. Butcher. Variation in whole-, landed- and trimmed-carcass and fin-weight ratios for various sharks captured on demersal set-lines off eastern Australia. Fisheries Research, 167(??):190–198, July 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783615000533. See corrigen-

dum [PGP^{+}16].

Pleizier:2016:CVW

[PGP^{+}16] Naomi Pleizier, Lee F. G. Gutowsky, Victor M. Peddemors, Steven J. Cooke, and Paul A. Butcher. Corrigendum to “Variation in whole-, landed- and trimmed- carcass and fin-weight ratios for various sharks captured on demersal set-lines off eastern Australia” [Fish. Res. 167 (2015) 190– 198]. Fisheries Research, 179(??):361, July 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616300650. See [PGP^{+}15].

Pajuelo:2010:BIC

[PGS10] José G. Pajuelo, José A. González, and José I. Santana. Bycatch and incidental catch of the black scabbardfish (Aphanopus spp.) fishery off the Canary Islands. Fisheries Research, 106(3):448–453, December 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610002286.

Polanco:2011:WAC

[PGS^{+}11] J. Polanco, U. Ganzedo, J. Sáenz, A. M. Caballero-Alfonso, and J. J. Castro-Hernández. Wavelet analysis of correlation among Canary Islands octopus captures per unit effort, sea-surface temperatures and the North Atlantic oscillation. Fisheries Research, 107(1–3):177–183, January 2011. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783610002833.

Pukk:2016:GDB

[PGVV16] Lilian Pukk, Riho Gross, Markus Vetemaa, and Anti Vasemägi. Genetic discrimination of brackish and freshwater populations of Eurasian perch (Perca fluviatilis L.) in the Baltic Sea drainage: implications for fish forensics. Fisheries Research, 183(??):155–164, November 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616301746.

Park:2017:RPG

[PGW17] Joo Myun Park, Troy F. Gaston, and Jane E. Williamson. Resource partitioning in gurnard species using trophic analyses: the importance of temporal resolution. Fisheries Research, 186 (part 1)(??):301–310, February 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616303332.

Piroddi:2015:RIM

[PGZ^{+}15] Chiara Piroddi, Michele Gristina, Kyrstn Zylich, Krista Greer, Aylin Ulman, Dirk Zeller, and Daniel Pauly. Reconstruction of Italy’s marine fisheries removals and fishing capacity, 1950–2010. Fisheries Research, 172(??):137–147, December 2015. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783615002015.

Pepin:2012:SVI

[PH12] Pierre Pepin and James A. Helbig. Sampling variability of ichthyoplankton surveys — exploring the roles of scale and resolution on uncertainty. Fisheries Research, 117–118(??):137–145, April 2012. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611002372.

Pinder:2019:TEP

[PHB19a] Adrian C. Pinder, Andrew J. Harrison, and J. Robert Britton. Temperature effects on the physiological status and reflex impairment in European grayling Thymallus thymallus from catch-and release angling. Fisheries Research, 211(??):169–175, March 2019. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618303266.

Pope:2019:SUE

[PHB^{+}19b] John G. Pope, Troels Jacob Hegland, Marta Ballesteros, Kåre Nolde Nielsen, and Mika Rahikainen. Steps to unlocking ecosystem based fisheries management: Towards displaying the N-dimensional potato. Fisheries Research, 209(??):117–128, January 2019. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618302182.

Plet-Hansen:2019:MDM

[PHBU19] Kristian S. Plet-Hansen, Heidrikur Bergsson, and Clara Ulrich. More data for the money: Improvements in design and cost efficiency of electronic monitoring in the Danish cod catch quota management trial. Fisheries Research, 215(??):114–122, July 2019. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619300748.

Pinchuk:2016:DBM

[PHE16] Alexei I. Pinchuk, H. Rodger Harvey, and Ginny L. Eckert. Development of biochemical measures of age in the Alaskan red king crab Paralithodes camtschaticus (Anomura): Validation, refinement and initial assessment. Fisheries Research, 183(??):92–98, November 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616301564.

Punt:2014:MSS

[PHFW14] André E. Punt, Felipe Hurtado-Ferro, and Athol R. Whitten. Model selection for selectivity in fisheries stock assessments. Fisheries Research, 158(??):124–134, October 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783613001550.

Pita:2018:ESE

[PHG^{+}18] Pablo Pita, Kieran Hyder, Pedro Gomes, Cristina Pita, Mafalda Rangel, Pedro Veiga, José Vingada, and Sebastián Villasante. Economic, social and ecological attributes of marine recreational fisheries in Galicia, Spain. Fisheries Research, 208(??):58–69, December 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618302121.

Plaganyi:2019:MIM

[PHG^{+}19] Éva E. Plagányi, Michael D. E. Haywood, Rebecca J. Gorton, Margaret C. Siple, and Roy Aijun Deng. Management implications of modelling fisheries recruitment. Fisheries Research, 217(??):169–184, September 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619300724.

Pitman:2013:FFI

[PHK13] Luke R. Pitman, James A. Haddy, and Rudy J. Kloser. Fishing and fecundity: the impact of exploitation on the reproductive potential of a deep-water fish, orange roughy (Hoplostethus atlanticus). Fisheries Research, 147(??):312–319, October 2013. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783613001604.

Pitman:2014:RCF

[PHK14] Luke R. Pitman, James A. Haddy, and Rudy J. Kloser. Response to comment on “fishing and fecundity: the impact of exploitation on the reproductive potential of a deep-water fish, orange roughy (Hoplostethus atlanticus)”. Fisheries Research, 155(??):196–197, July 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614000873.

Punt:2016:CSS

[PHLT16] André E. Punt, Malcolm Haddon, L. Richard Little, and Geoffrey N. Tuck. Can a spatially-structured stock assessment address uncertainty due to closed areas? A case study based on pink ling in Australia. Fisheries Research, 175(??):10– 23, March 2016. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783615301314.

Palmer:2016:DPA

[PHM^{+}16a] Michael C. Palmer, Patty Hersey, Heidi Marotta, Gina Reppucci Shield, and Sarah B. Cierpich. The design and performance of an automated observer deployment system for the Northeastern United States groundfish fishery. Fisheries Research, 179(??):33–46, July 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616300261.

Punt:2016:EGW

[PHM16b] André E. Punt, Malcolm Haddon, and Richard McGarvey. Estimating growth within size-structured fishery stock assessments: What is the state of the art and what does the future look like? Fisheries Research, 180(??):147–160, August 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783614003324.

Pol:2016:ICM

[PHRH16a] Michael V. Pol, Bent Herrmann, Chris Rillahan, and Pingguo He. Impact of codend mesh sizes on selectivity and retention of Acadian redfish Sebastes fasciatus in the Gulf of Maine trawl fishery. Fisheries Research, 184(??):54–63, December 2016. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783616302028.

Pol:2016:SRP

[PHRH16b] Michael V. Pol, Bent Herrmann, Chris Rillahan, and Pingguo He. Selectivity and retention of pollock Pollachius virens in a Gulf of Maine trawl fishery. Fisheries Research, 184(??):47–53, December 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615300382.

Punt:2015:WAC

[PHT15] André E. Punt, Malcolm Haddon, and Geoffrey N. Tuck. Which assessment configurations perform best in the face of spatial heterogeneity in fishing mortality, growth and recruitment? A case study based on pink ling in Australia.

Fisheries Research, 168(??):85–99, August 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615001253.

Pascoe:2017:IRI

[PICK17] Sean Pascoe, James Innes, Anthony Courtney, and Marco Kienzle. Impact of reducing investment disincentives on the sustainability of the Moreton Bay prawn trawl fishery. Fisheries Research, 186 (part 1)(??):121–130, February 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783616302600.

Pitcher:2014:QIE

[Pit14] C. Roland Pitcher. Quantitative indicators of environmental sustainability risk for a tropical shelf trawl fishery. Fisheries Research, 151(??):136–147, March 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613002531.

Petreman:2014:OBS

[PJM14] Ian C. Petreman, Nicholas E. Jones, and Scott W. Milne. Observer bias and subsampling efficiencies for estimating the number of migrating fish in rivers using dualfrequency IDentification SONar (DIDSON). Fisheries Research, 155(??):160–167, July 2014. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614000824.

Perez-Jimenez:2015:SSS

[PJML15] Juan Carlos Pérez-Jiménez and Iván Mendez-Loeza. The small-scale shark fisheries in the southern Gulf of Mexico: Understanding their heterogeneity to improve their management. Fisheries Research, 172(??):96–104, December 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783615300138.

Pedersen:2017:SGC

[PJR17] M. I. Pedersen, N. Jepsen, and G. Rasmussen. Survival and growth compared between wild and farmed eel stocked in freshwater ponds. Fisheries Research, 194(??):112–116, October 2017. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783617301406.

Pastor:2013:CMM

[PKAL13] Jérémy Pastor, Barbara Koeck, Patrick Astruch, and Philippe Lenfant. Coastal man-made habitats: Potential nurseries for an exploited fish species, Diplodus sargus (Linnaeus, 1758). Fisheries Research, 148(??):74–80, November 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783613002075.

Pacheco:2011:CCH

[PKH^{+}11] J. C. Pacheco, D. W. Kerstetter, F. H. Hazin, Humberto Hazin, R. S. S. L. Segundo, J. E. Graves, F. Carvalho, and P. E. Travassos. A comparison of circle hook and J hook performance in a western equatorial Atlantic Ocean pelagic longline fishery. Fisheries Research, 107(1–3):39–45, January 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783610002511.

Pascoe:2016:EUB

[PKHD16] Sean Pascoe, Viktoria Kahui, Trevor Hutton, and Catherine Dichmont. Experiences with the use of bioeconomic models in the management of Australian and New Zealand fisheries. Fisheries Research, 183(??):539–548, November 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578361630008X.

Plaza:2010:DPS

[PKO10] Guido Plaza, Satoshi Katayama, and Michio Omori. Daily patterns of settlement and individual growth rates of youngof-the-year of the rockfish Sebastes inermis in a Sargassum bed. Fisheries Research, 103(1–3):48–55, April 2010. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783610000457.

Pazhayamadom:2013:SSC

[PKRC13] Deepak George Pazhayamadom, Ciarán J. Kelly, Emer Rogan, and Edward A. Codling. Self-starting CUSUM approach for monitoring data poor fisheries. Fisheries Research, 145(??):114–127, August 2013. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613000350.

Pazhayamadom:2015:DIC

[PKRC15] Deepak George Pazhayamadom, Ciarán J. Kelly, Emer Rogan, and Edward A. Codling. Decision Interval Cumulative Sum Harvest Control Rules (DI-CUSUM-HCR) for managing fisheries with limited historical information. Fisheries Research, 171(??):154–169, November 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614002768.

Pearson:2015:ESS

[PKWP15] Kristen Nicole Pearson, William Louis Kendall, Dana Leonard Winkelman, and William Riley Persons. Evidence for skipped spawning in a potamodromous cyprinid, humpback chub (Gila cypha), with implications for demographic parameter estimates. Fisheries Research, 170(??):50–59, October 2015. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783615001563.

Pearson:2016:TBP

[PKWP16] Kristen Nicole Pearson, William Louis Kendall, Dana Leonard Winkelman, and William Riley Persons. Tradeoffs between physical captures and PIT tag antenna array detections: a case study for the Lower Colorado River Basin population of humpback chub (Gila cypha). Fisheries Research, 183(??):263–274, November 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S016578361630203X.

Perretti:2018:WDS

[PL18] Charles T. Perretti and Christopher M. Legault. When does smoothing the output of index-based stock assessments improve estimates of fish population biomass? Fisheries Research, 207(??):1–8, November 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618301607.

Pristupa:2016:PIG

[PLA16] Alexey O. Pristupa, Machiel Lamers, and Bas Amelung. Private informational governance in post-Soviet waters: Implications of the Marine Stewardship Council certification in the Russian Barents Sea region. Fisheries Research, 182(??):128–135, October 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615300151.

Plank:2018:HSF

[Pla18] Michael J. Plank. How should fishing mortality be distributed under balanced harvesting? Fisheries Research, 207(??):171–174, November 2018. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618301735.

Pintassilgo:2018:IFA

[PLGL18] Pedro Pintassilgo, Marita Laukkanen, Lone Grønbæk, and Marko Lindroos. International fisheries agreements and nonconsumptive values. Fisheries Research, 203(??):46–54, July 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783617301789.

Piner:2016:EUR

[PLM16] Kevin R. Piner, Hui-Hua Lee, and Mark N. Maunder. Evaluation of using random-at-length observations and an equilibrium approximation of the population age structure in fitting the von Bertalanffy growth function. Fisheries Research, 180(??):128–137, August 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615001733.

Pelletier:2011:CVC

[PLMT^{+}11] Dominique Pelletier, Kévin Leleu, Gérard Mou-Tham, Nicolas Guillemot, and Pascale Chabanet. Comparison of visual census and high definition video transects for monitoring coral reef fish assemblages. Fisheries Research, 107(1–3):84–93, January 2011. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610002705.

Pereira:2012:CBU

[PLPN12] Julio César Pereira, Roseli Aparecida Leandro, Miguel Petrere, and Tom Nishida. Comparison between univariate and bivariate geostatistical models for estimating catch per unit of effort (cpue): a simulation study. Fisheries Research, 121–122(??):115–125, June 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612000495.

Pita:2016:GID

[PLSM^{+}16] Alfonso Pita, Ana Leal, Angie Santafé-Muñoz, Carmen Piñeiro, and Pablo Presa. Genetic inference of demographic connectivity in the Atlantic European hake metapopulation (Merluccius merluccius) over a spatio-temporal framework. Fisheries Research, 179(??):291–301, July 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616300868.

Picquelle:2011:PGS

[PM11] Susan J. Picquelle and Kathryn L. Mier. A practical guide to statistical methods for comparing means from two-stage sampling. Fisheries Research, 107(1–3):1–13, January 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783610002183.

Punt:2013:SSA

[PM13] André E. Punt and Mark N. Maunder. Stock synthesis: Advancing stock assessment application and research through the use of a general stock assessment computer program. Fisheries Research, 142(??):1–2, May 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612003268.

Park:2011:SSK

[PMB^{+}11] Hae-Hoon Park, Russell B. Millar, Bong-Seong Bae, Heui-Chun An, Young Yull Chun, Jae Hyeong Yang, and Sang Chul Yoon. Size selectivity of Korean flounder (Glyptocephalus stelleri) by gillnets and trammel nets using an extension of SELECT for experiments with differing mesh sizes. Fisheries Research, 107(1–3):196–200, January 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783610002857.

Paoletti:2018:SSR

[PMC^{+}18] Michela Paoletti, Simonetta Mattiucci, Alessandra Colantoni, Arne Levsen, Melanie Gay, and Giuseppe Nascetti. Species-specific real Time-PCR primers/probe systems to identify fish parasites of the genera Anisakis, Pseudoterranova and Hysterothylacium (Nematoda: Ascaridoidea). Fisheries Research, 202(??):38–48, June 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617301984.

Piras:2016:PMI

[PMF^{+}16] Camilla Piras, Monica Mion, Tomaso Fortibuoni, Gianluca Franceschini, Elisa Punzo, Pierluigi Strafella, Marija Despalatović, Ivan Cvitković, and Saša Raicevich. A photographic method to identify benthic assemblages based on demersal trawler discards. Fisheries Research, 178(??):142– 151, June 2016. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783615300655.

Parkinson:2018:AER

[PMG18] Eric A. Parkinson, Murdoch K. McAllister, and Theresa I. Godin. Angler effort response to sterile rainbow trout stocking in small BC lakes. Fisheries Research, 204(??):245–257, August 2018. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783618300572.

Prchalova:2010:FAD

[PMK^{+}10] Marie Prchalová, Tomáš Mrkvička, Jan Kubečka, Jiří Peterka, Martin Čech, Milan Muška, Michal Kratochvíl, and Mojmír Vašek. Fish activity as determined by gillnet catch: a comparison of two reservoirs of different turbidity. Fisheries Research, 102(3):291–296, March 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610000032.

Punt:2012:EED

[PML^{+}12] André E. Punt, Richard McGarvey, Adrian Linnane, Justin Phillips, Lianos Triantafillos, and John Feenstra. Evaluating empirical decision rules for southern rock lobster fisheries: a South Australian example. Fisheries Research, 115–116(??):60–71, March 2012. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611003390.

Pinto:2015:SIF

[PMM^{+}15] Angela Di Pinto, Patrizia Marchetti, Anna Mottola, Giancarlo Bozzo, Elisabetta Bonerba, Edmondo Ceci, Marilisa Bottaro, and Giuseppina Tantillo. Species identification in fish fillet products using DNA barcoding. Fisheries Research, 170(??):9–13, October 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615001472.

Prchalova:2011:MGC

[PMP^{+}11] Marie Prchalová, Tomáš Mrkvička, Jiří Peterka, Martin Čech, Luděk Berec, and Jan Kubečka. A model of gillnet catch in relation to the catchable biomass, saturation, soak time and sampling period. Fisheries Research, 107(1–3):201–209, January 2011. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610002869.

Purcell:2016:TSS

[PNAL16] Steven W. Purcell, Poasi Ngaluafe, Karibanang T. Aram, and Watisoni Lalavanua. Trends in small-scale artisanal fishing of sea cucumbers in Oceania. Fisheries Research, 183(??):99–110, November 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616301473.

Punt:2018:WES

[POM^{+}18] André E. Punt, Daniel K. Okamoto, Alec D. MacCall, Andrew O. Shelton, Derek R. Armitage, Jaclyn S. Cleary, Ian P. Davies, Sherri C. Dressel, Tessa B. Francis, Phillip S. Levin, R. Russ Jones, Harvey Kitka, Lynn Chi Lee, Jim A. McIsaac, Melissa R. Poe, Steve Reifenstuhl, Jennifer J. Silver, Jörn O. Schmidt, Thomas F. Thornton, Rudiger Voss, and John Woodruff. When are estimates of spawning stock biomass for small pelagic fishes improved by taking spatial structure into account? Fisheries Research, 206(??):65–78, October 2018. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578361830122X.

Pope:2019:OST

[Pop19] John G. Pope. T-ONS a swift transportable and user friendly integrative model of the North Sea for decision support. Fisheries Research, 215(??):9–20, July 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619300554.

Prestrelo:2019:NPC

[POV19] Luana Prestrelo, Rafaela Oliveira, and Marcelo Vianna. A new proposal to classify small fishing vessels to improve tropical estuarine fishery management. Fisheries Research, 211(??):100–110, March 2019. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618303229.

Parga:2015:HLS

[PPA^{+}15] María L. Parga, Maite Pons, Sandra Andraka, Liliana Rendón, Takahisa Mituhasi, Martín Hall, Lucas Pacheco, Alvaro Segura, Michael Osmond, and Nick Vogel. Hooking locations in sea turtles incidentally captured by artisanal longline fisheries in the Eastern Pacific Ocean. Fisheries Research, 164(??):231–237, April 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614003385.

Prieto:2013:BRS

[PPBM13] Rui Prieto, Christopher Kim Pham, Cristina Brito, and Telmo Morato. Biomass removal from shore-based whaling in the Azores. Fisheries Research, 143(??):98–101, June 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783613000179.

Portela:2010:PDO

[PPdR^{+}10] J. M. Portela, G. J. Pierce, J. L. del Río, M. Sacau, T. Patrocinio, and R. Vilela. Preliminary description of the overlap between squid fisheries and VMEs on the high seas of the Patagonian shelf. Fisheries Research, 106(2):229–238, November 2010. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783610001396.

Perdichizzi:2011:DPP

[PPG^{+}11] Anna Perdichizzi, Laura Pirrera, Daniela Giordano, Francesco Perdichizzi, Barbara Busalacchi, Adriana Profeta, Teresa Bottari, and Paola Rinelli. Distribution patterns and population structure of Illex coindetii (Cephalopoda: Ommastrephidae) in the Southern Tyrrhenian Sea: Historical series of 14 years trawl survey. Fisheries Research, 109(2–3):342– 350, May 2011. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783611000932.

Pope:2017:ENR

[PPH^{+}17] Kevin L. Pope, Larkin A. Powell, Brian S. Harmon, Mark A. Pegg, and Christopher J. Chizinski. Estimating the number of recreational anglers for a given waterbody. Fisheries Research, 191(??):69–75, July 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617300644.

Park:2018:OSD

[PPLK18] Hyun Je Park, Tae Hee Park, Chung-Il Lee, and Chang-Keun Kang. Ontogenetic shifts in diet and trophic position of walleye pollock, Theragra chalcogramma, in the western East Sea (Japan Sea) revealed by stable isotope and stomach content analyses. Fisheries Research, 204(??):297–304, August 2018. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783618300808.

Pita:2017:TGE

[PPVP17] Alfonso Pita, Montse Pérez, Francisco Velasco, and Pablo Presa. Trends of the genetic effective population size in the southern stock of the European hake. Fisheries Research, 191(??):108–119, July 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617300590.

Pickering:2010:MCC

[PQ10] Tyler Pickering and Pedro A. Quijon. Missing claws in a commercial catch of the American lobster: Implications for life history and the fishery. Fisheries Research, 102(3):335– 339, March 2010. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783610000056.

Potts:2018:EGG

[PR18] Stephen E. Potts and Kenneth A. Rose. Evaluation of GLM and GAM for estimating population indices from fishery independent surveys. Fisheries Research, 208(??):167–178, December 2018. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783618302145.

Perales-Raya:2010:AEO

[PRBGS^{+}10] Catalina Perales-Raya, Aurora Bartolomé, M. Teresa García-Santamaría, Pedro Pascual-Alayón, and Eduardo Almansa. Age estimation obtained from analysis of octopus (Octopus vulgaris Cuvier, 1797) beaks: Improvements and comparisons. Fisheries Research, 106(2):171–176, November 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783610001062.

Perez-Ramirez:2016:MSC

[PRCGD16] Mónica Pérez-Ramírez, Mauricio Castrejón, Nicolás L. Gutiérrez, and Omar Defeo. The Marine Stewardship Council certification in Latin America and the Caribbean: a review of experiences, potentials and pitfalls. Fisheries Research, 182(??):50–58, October 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615301302.

Prellezo:2017:EEV

[Pre17] Raúl Prellezo. Expected economic value of the information provided by fishery research surveys. Fisheries Research, 190(??):95–102, June 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617300322.

Pontavice:2018:ISH

[PRL^{+}18] Hubert Du Pontavice, Marine Randon, Sigrid Lehuta, Youen Vermard, and Marie Savina-Rolland. Investigating spatial heterogeneity of von Bertalanffy growth parameters to inform the stock structuration of common sole, Solea solea, in the Eastern English Channel. Fisheries Research, 207(??):28–36, November 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618301450.

Perez-Rodriguez:2013:DGC

[PRMKASR13] Alfonso Pérez-Rodríguez, Joanne Morgan, Mariano Koen-Alonso, and Fran Saborido-Rey. Disentangling genetic change from phenotypic response in reproductive parameters of Flemish Cap cod Gadus morhua. Fisheries Research, 138(??):62–70, February 2013. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612002731.

Pascual:2018:AEH

[PRP^{+}18] S. Pascual, H. Rodríguez, G. J. Pierce, L. C. Hastie, and A. F. González. The NE Atlantic European hake: a neglected high exposure risk for zoonotic parasites in European fish markets. Fisheries Research, 202(??):69–78, June 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783617303430.

Puncher:2019:CIA

[PRR^{+}19] Gregory Neils Puncher, Sherrylynn Rowe, George A. Rose, Nathalie M. Leblanc, Geneviève J. Parent, Yanjun Wang, and Scott A. Pavey. Chromosomal inversions in the Atlantic cod genome: Implications for management of Canada’s northern cod stock. Fisheries Research, 216(??):29–40, August 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783619300852.

Punt:2012:HWC

[PS12] André E. Punt and Cody Szuwalski. How well can F_{MSY} and B_{MSY} be estimated using empirical measures of surplus production? Fisheries Research, 134–136(??):113–124, December 2012. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783612002603.

Phillipson:2013:SSF

[PS13] Jeremy Phillipson and David Symes. Science for sustainable fisheries management: an interdisciplinary approach. Fisheries Research, 139(??):61–64, March 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612002810.

Pulver:2019:FID

[PS19] Jeffrey R. Pulver and J. A. Stephen. Factors that influence discarding in the Gulf of Mexico commercial grouper-tilefish IFQ reef fish fishery. Fisheries Research, 218(??):218–228, October 2019. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783619301481.

Phillips:2015:HCP

[PSE15] Sophy R. McCully Phillips, Finlay Scott, and Jim R. Ellis. Having confidence in productivity susceptibility analyses: a method for underpinning scientific advice on skate stocks? Fisheries Research, 171(??):87–100, November 2015. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578361500017X.

Prichard:2019:GVU

[PSJ^{+}19] Carson G. Prichard, James J. Student, Jory L. Jonas, Nicole M. Watson, and Kevin L. Pangle. Geologic variability underlying stream catchment areas correlates with fish otolith microchemistry across disparate glacial till depths. Fisheries Research, 216(??):109–119, August 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619300979.

Paufve:2019:ASE

[PSL^{+}19] Matthew R. Paufve, Suresh A. Sethi, Brian F. Lantry, Brian C. Weidel, and Lars G. Rudstam. Assessing the spawning ecology of fish in situ using a benthic pump sampler. Fisheries Research, 214(??):19–24, June 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619300372.

Papetti:2018:GVS

[PSM^{+}18] Chiara Papetti, Luca Schiavon, Massimo Milan, Magnus Lucassen, Jilda Alicia Caccavo, Marta Paterno, Elisa Boscari, Ilaria Anna Maria Marino, Leonardo Congiu, and Lorenzo Zane. Genetic variability of the striped venus Chamelea gallina in the northern Adriatic Sea. Fisheries Research, 201(??):68–78, May 2018. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618300134.

Punt:2014:ESR

[PSS14] André E. Punt, Cody S. Szuwalski, and William Stockhausen. An evaluation of stock–recruitment proxies and environmental change points for implementing the US Sustainable Fisheries Act. Fisheries Research, 157(??):28–40, September 2014. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783614000976.

Punt:2015:ABS

[PSS15] André E. Punt, Nan-Jay Su, and Chi-Lu Sun. Assessing billfish stocks: a review of current methods and some future directions. Fisheries Research, 166(??):103–118, June 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783614002422.

Parkes:2016:ECS

[PSU^{+}16] Graeme Parkes, Jill H. Swasey, Fiona M. Underwood, Timothy P. Fitzgerald, Kent Strauss, and David J. Agnew. The effects of catch share management on MSC certification scores. Fisheries Research, 182(??):18–27, October 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783615301004.

Perretti:2019:STD

[PT19] Charles T. Perretti and James T. Thorson. Spatio-temporal dynamics of summer flounder (Paralichthys dentatus) on the Northeast US shelf. Fisheries Research, 215(??):62–68, July 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783619300712.

Phung:2015:MDL

[PTB^{+}15] Anh Tuan Phung, Ingrid Tulp, Willy Baeyens, Marc Elskens, Martine Leermakers, and Yue Gao. Migration of diadromous and landlocked smelt populations studied by otolith geochemistry. Fisheries Research, 167(??):123–131, July 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783615000417.

Pichon:2017:SIS

[PTBT17] Céline Le Pichon, Évelyne Tales, Jérôme Belliard, and Christian E. Torgersen. Spatially intensive sampling by electrofishing for assessing longitudinal discontinuities in fish distribution in a headwater stream. Fisheries Research, 185(??):90–101, January 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616303198.

Poh:2018:EPR

[PTC^{+}18] Brian Poh, James R. Tweedley, Jennifer A. Chaplin, Kerry M. Trayler, and Neil R. Loneragan. Estimating predation rates of restocked individuals: the influence of timing-of-release on metapenaeid survival. Fisheries Research, 198(??):165–179, February 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617302679.

Pang:2018:VCC

[PTF^{+}18] Yumeng Pang, Yongjun Tian, Caihong Fu, Bin Wang, Jianchao Li, Yiping Ren, and Rong Wan. Variability of coastal cephalopods in overexploited China Seas under climate change with implications on fisheries management. Fisheries Research, 208(??):22–33, December 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618301942.

Palmer:2017:CSR

[PTG^{+}17] Miquel Palmer, Borja Tolosa, Antoni Maria Grau, Maria del Mar Gil, Clara Obregón, and Beatriz Morales-Nin. Combining sale records of landings and fishers knowledge for predicting métiers in a small-scale, multi-gear, multispecies fishery. Fisheries Research, 195(??):59–70, November 2017.

CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783617301753.

Poirier:2018:NBR

[PTM^{+}18] Luke A. Poirier, Skylar Tang, Joshua Mohan, Ellen O’Connor, Emily Dennis, Mohammad Abdullah, Daohan Zhou, Henrik Stryhn, Sophie St-Hilaire, and Pedro A. Quijón. A novel bycatch reduction device (BRD) and its use in a directed fishery for non-indigenous green crabs (C. maenas) in Atlantic Canada. Fisheries Research, 204(??):165–171, August 2018. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618300547.

Pennino:2016:SMA

[PTSC^{+}16] M. Grazia Pennino, Mario J. F. Thomé-Souza, Adriana R. Carvalho, Luiz Carlos da Silveira Fontes, Cristiano Parente, and Priscila F. M. Lopes. A spatial multivariate approach to understand what controls species catch composition in small-scale fisheries. Fisheries Research, 175(??):132–141, March 2016. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578361530151X.

Peterson:2018:TEP

[PTTC18] Damon Peterson, Randi B. Trantham, Tulley G. Trantham, and Colleen A. Caldwell. Tagging effects of passive integrated transponder and visual implant elastomer on the small-bodied white sands pupfish (Cyprinodon tularosa). Fisheries Research, 198(??):203–208, February 2018. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783617302333.

Punt:2013:PMP

[PTW^{+}13] André E. Punt, Fabian Trinnie, Terence I. Walker, Richard McGarvey, John Feenstra, Adrian Linnane, and Klaas Hartmann. The performance of a management procedure for rock lobsters, Jasus edwardsii, off western Victoria, Australia in the face of non-stationary dynamics. Fisheries Research, 137(??):116–128, January 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361200286X.

Pulver:2017:SSF

[Pul17] J. R. Pulver. Sink or swim? Factors affecting immediate discard mortality for the Gulf of Mexico commercial reef fish fishery. Fisheries Research, 188(??):166–172, April 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783616304374.

Punt:2011:ICC

[Pun11] André E. Punt. The impact of climate change on the performance of rebuilding strategies for overfished groundfish species of the U.S. west coast. Fisheries Research, 109(2–3):320–329, May 2011. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611000907.

Punt:2017:SID

[Pun17] André E. Punt. Some insights into data weighting in integrated stock assessments. Fisheries Research, 192(??):52– 65, August 2017. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783615301582.

Punt:2018:SCE

[Pun18] M. J. Punt. Sunk costs equal sunk boats? The effect of entry costs in a transboundary sequential fishery. Fisheries Research, 203(??):55–62, July 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617302114.

Punt:2019:MRS

[Pun19a] André E. Punt. Modelling recruitment in a spatial context: a review of current approaches, simulation evaluation of options, and suggestions for best practices. Fisheries Research, 217(??):140–155, September 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617302424.

Punt:2019:SSA

[Pun19b] André E. Punt. Spatial stock assessment methods: a viewpoint on current issues and assumptions. Fisheries Research, 213(??):132–143, May 2019. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619300177.

Pena:2010:MLM

[PUSI10] M. Peña, A. Uriarte, M. Santos, and L. Ibaibarriaga. A maximum likelihood method for obtaining incubation temperature of eggs of synchronous spawning fishes. Fisheries Research, 103(1–3):9–17, April 2010. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361000010X.

Prince:2015:LBS

[PVKH15] Jeremy Prince, Steven Victor, Valentino Kloulchad, and Adrian Hordyk. Length based SPR assessment of eleven Indo–Pacific coral reef fish populations in Palau. Fisheries Research, 171(??):42–58, November 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0165783615001885. See corrigendum [PVKH18].

Prince:2018:CLB

[PVKH18] Jeremy Prince, Steven Victor, Valentino Kloulchad, and Adrian Hordyk. Corrigendum to “Length based SPR assessment of eleven Indo–Pacific coral reef fish populations in Palau” [Fish. Res. 171 (2015) 42–58]. Fisheries Research, 204(??):458, August 2018. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617303545. See [PVKH15].

Pereira:2019:CSO

[PVMG19] Fábio Pereira, Paulo Vasconcelos, Ana Moreno, and Miguel B. Gaspar. Catches of Sepia officinalis in the smallscale cuttlefish trap fishery off the Algarve coast (southern Portugal). Fisheries Research, 214(??):117–125, June 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619300220.

Pope:2019:PCI

[PW19] John G. Pope and Charlotte T. Weber. A parable of compliance issues and their link to EBFM outcomes. Fisheries Research, 211(??):51–58, March 2019. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618302819.

Pickering:2010:CVA

[PWA10] Tyler Pickering, Jim Williams, and Joe Apaloo. Catch variation among traps in an American lobster (Homarus americanus) trawl. Fisheries Research, 106(3):555–559, December 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783610002122.

Parker:2016:IBV

[PWB^{+}16] D. Parker, H. Winker, A. T. F. Bernard, E. R. Heyns-Veale, T. J. Langlois, E. S. Harvey, and A. Götz. Insights from baited video sampling of temperate reef fishes: How biased are angling surveys? Fisheries Research, 179(??):191–201, July 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783616300558.

Parasiewicz:2016:LVE

[PWM^{+}16] Piotr Parasiewicz, Wiesław Wiśniewolski, Marian Mokwa, Sabina Zioła, Paweł Prus, and Małgorzata Godlewska. A low-voltage electric fish guidance system — NEPTUN. Fisheries Research, 181(??):25–33, September 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616300716.

Paramo:2012:DSF

[PWSP12] Jorge Paramo, Matthias Wolff, and Ulrich Saint-Paul. Deepsea fish assemblages in the Colombian Caribbean Sea. Fisheries Research, 125–126(??):87–98, August 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612000914.

Queirolo:2019:JSB

[QCZC^{+}19] Dante Queirolo, Ana Luíza Couto-Ziezkowski, José Cusba, Pedro Apablaza, and Mauricio Ahumada. Jumbo squid behaviour in response to a rigid grid in the Chilean hake trawl fishery. Fisheries Research, 216(??):1–5, August 2019. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619300773.

Quinn:2017:CCF

[QDG17] Terrance J. Quinn, Richard B. Deriso, and Philip D. Ganz. Combining the Cohen–Fishman growth increment model with a Box–Cox transformation. Fisheries Research, 187(??):11–21, March 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361630354X.

Queirolo:2011:ACR

[QEH^{+}11a] Dante Queirolo, Karim Erzini, Carlos F. Hurtado, Mauricio Ahumada, and Milagrosa C. Soriguer. Alternative codends to reduce bycatch in Chilean crustacean trawl fisheries. Fisheries Research, 110(1):18–28, June 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611001147.

Queirolo:2011:SCB

[QEH^{+}11b] Dante Queirolo, Karim Erzini, Carlos F. Hurtado, Erick Gaete, and Milagrosa C. Soriguer. Species composition and bycatches of a new crustacean trawl in Chile. Fisheries Research, 110(1):149–159, June 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611001330.

Quynh:2018:FPM

[QHSI18] Chi Nguyen Thi Quynh, Atakelty Hailu, Steven Schilizzi, and Sayed Iftekhar. Fisher participation in monitoring:

Does it help reduce excessive investment in fishing capacity? Fisheries Research, 206(??):138–149, October 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783618301292.

Quinones:2013:JBD

[QMAM13] Javier Quiñones, Aldrin Monroy, E. Marcelo Acha, and Hermes Mianzan. Jellyfish bycatch diminishes profit in an anchovy fishery off Peru. Fisheries Research, 139(??):47–50, March 2013. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783612001439.

Queirolo:2010:DOC

[QMGP10] Dante Queirolo, Ivonne Montenegro, Erick Gaete, and Guido Plaza. Direct observation of Chilean hake (Merluccius gayi gayi) behaviour in response to trawling in a South Central Chilean fishery. Fisheries Research, 102(3):327–329, March 2010. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578360900321X.

Quintanilla:2014:SAM

[QQG14] J. M. Quintanilla, L. F. Quintanilla, and A. García. A semi-automated method for daily age estimation in larval populations by discriminant function models. Fisheries Research, 157(??):7–12, September 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614000940.

Quinn:2017:TTP

[Qui17] Brady K. Quinn. Threshold temperatures for performance and survival of American lobster larvae: a review of current knowledge and implications to modeling impacts of climate change. Fisheries Research, 186 (part 1)(??):383–396, February 2017. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783616303150.

Quiroz:2010:IUE

[QWC10] J. C. Quiroz, Rodrigo Wiff, and Bruno Caneco. Incorporating uncertainty into estimation of natural mortality for two species of Rajidae fished in Chile. Fisheries Research, 102(3):297–304, March 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610000044.

Quiroz:2011:VEY

[QWCB11] J. C. Quiroz, Rodrigo Wiff, Luis A. Cubillos, and Mauricio A. Barrientos. Vulnerability to exploitation of the yellownose skate (Dipturus chilensis) off southern Chile. Fisheries Research, 109(2–3):225–233, May 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611000105.

Queiros:2013:GVS

[QWMD13] Ana M. Queirós, Adrian Weetman, H. Anne McLay, and Helen Dobby. Geographical variation in size at the onset of maturity of male and female Norway lobster Nephrops norvegicus (L., Homarida: Decapoda) in Scottish waters. Fisheries Research, 139(??):132–144, March 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612003281.

Raghavan:2011:DMT

[RADR11] Rajeev Raghavan, Anvar Ali, Neelesh Dahanukar, and Alison Rosser. Is the Deccan Mahseer, Tor khudree (Sykes, 1839) (Pisces: Cyprinidae) fishery in the Western Ghats Hotspot sustainable? A participatory approach to stock assessment. Fisheries Research, 110(1):29–38, June 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783611001172.

Rodrigues:2019:CSF

[RAG19] Amanda Ricci Rodrigues, Patrízia Raggi Abdallah, and Maria A. Gasalla. Cost structure and financial performance of marine commercial fisheries in the South Brazil Bight.

Fisheries Research, 210(??):162–174, February 2019. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783618302868.

Rodriguez:2018:SPR

[RAGP18] Helena Rodríguez, Elvira Abollo, Ángel F. González, and Santiago Pascual. Scoring the parasite risk in highlyvaluable fish species from southern ICES areas. Fisheries Research, 202(??):134–139, June 2018. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617301741.

Raberinary:2012:RCO

[RB12] D. Raberinary and S. Benbow. The reproductive cycle of Octopus cyanea in southwest Madagascar and implications for fisheries management. Fisheries Research, 125–126(??):190–197, August 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783612001051.

Rogers:2019:ATF

[RBG^{+}19a] Matthew Rogers, Ryan Bare, Andrew Gray, Taylor Scott-Moelder, and Ron Heintz. Assessment of two feeds on survival, proximate composition, and amino acid carbon isotope discrimination in hatchery-reared Chinook salmon. Fisheries Research, 219(??):Article 105303, November 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578361930150X.

Rouyer:2019:TAB

[RBG^{+}19b] Tristan Rouyer, Sylvain Bonhommeau, Nicolas Giordano, Saviour Ellul, Giovanni Ellul, Simeon Deguara, Bertrand Wendling, Mohamed Moez Belhaj, Vincent Kerzerho, and Serge Bernard. Tagging Atlantic bluefin tuna from a farming cage: an attempt to reduce handling times for large scale deployments. Fisheries Research, 211(??):27–31, March 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783618302959.

Robotham:2010:AIS

[RBGE^{+}10] Hugo Robotham, Paul Bosch, Juan Carlos Gutiérrez-Estrada, Jorge Castillo, and Inmaculada Pulido-Calvo. Acoustic identification of small pelagic fish species in Chile using support vector machines and neural networks. Fisheries Research, 102(1–2):115–122, February 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783609002793.

Revill:2013:MAP

[RBM13] Andrew S. Revill, Matt K. Broadhurst, and Russell B. Millar. Mortality of adult plaice, Pleuronectes platessa and sole, Solea solea discarded from English Channel beam trawlers. Fisheries Research, 147(??):320–326, October 2013. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783613001690.

Riou:2016:UOC

[RBM^{+}16] A. Riou, G. Bareille, F. Morat, K. Pothin, N. Bru, and P. Chabanet. The usefulness of otolith chemistry to determine the life history of the honeycomb grouper around Réunion Island (SW Indian Ocean). Fisheries Research, 179(??):104–114, July 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616300297.

Romeo:2015:SHF

[RBR^{+}15] Teresa Romeo, Pietro Battaglia, Saša Raicevich, Patrizia Perzia, and Franco Andaloro. Swordfish harpoon fishery in the Mediterranean Sea: Recent data to implement the Marine Strategy Framework Directive and the EcAp (Ecosystem Approach) process. Fisheries Research, 161(??):191–199, January 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614002276.

Ravard:2014:EPI

[RBT14] David Ravard, Anik Brind’Amour, and Verena M. Trenkel. Evaluating the potential impact of fishing on demersal species in the Bay of Biscay using simulations and survey data. Fisheries Research, 157(??):86–95, September 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783614000897.

Reubens:2013:AWA

[RBV^{+}13] J. T. Reubens, U. Braeckman, J. Vanaverbeke, C. Van Colen, S. Degraer, and M. Vincx. Aggregation at windmill artificial reefs: CPUE of Atlantic cod (Gadus morhua) and pouting (Trisopterus luscus) at different habitats in the Belgian part of the North Sea. Fisheries Research, 139(??):28– 34, March 2013. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578361200327X.

Rogers:2014:TOE

[RBW14] Mark W. Rogers, Andrew B. Barbour, and Kyle L. Wilson. Trade-offs in experimental designs for estimating post-release mortality in containment studies. Fisheries Research, 151(??):130–135, March 2014. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613002488.

Ragonese:2017:MSR

[RC17] Sergio Ragonese and Leonardo Cannizzaro. The missed sides in the reconstruction of Italy’s marine fisheries total removals by Piroddi et al. (2015). Fisheries Research, 186 (part 1)(??):23–24, February 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616302508.

Robotham:2011:CMC

[RCBPK11] H. Robotham, J. Castillo, P. Bosch, and J. Perez-Kallens. A comparison of multi-class support vector machine and classification tree methods for hydroacoustic classification of fish-schools in Chile. Fisheries Research, 111(3):170–176, November 2011. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783611002694.

Rossi:2019:IFS

[RCBS19] Steven P. Rossi, Sean P. Cox, Hugues P. Benoît, and Douglas P. Swain. Inferring fisheries stock status from competing hypotheses. Fisheries Research, 216(??):155–166, August 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783619301080.

Rincon:2019:GCA

[RCE^{+}19] Margarita María Rincón, Rachele Corti, Bjarki Thor Elvarsson, Fernando Ramos, and Javier Ruiz. Granger-causality analysis of integrated-model outputs, a tool to assess external drivers in fishery. Fisheries Research, 213(??):42–55, May 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783619300013.

Russo:2016:MLP

[RCF^{+}16] Tommaso Russo, Paolo Carpentieri, Fabio Fiorentino, Enrico Arneri, Michele Scardi, Alfonso Cioffi, and Stefano Cataudella. Modeling landings profiles of fishing vessels: an application of self-organizing maps to VMS and logbook data. Fisheries Research, 181(??):34–47, September 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578361630100X.

Riccioni:2017:ECL

[RCF^{+}17] Giulia Riccioni, Alessia Cariani, Giorgia Ferrara, Rita Cannas, Riccardo Melis, Marco Stagioni, Piero Addis, and Fausto Tinti. Evolutionary constraints limiting the variation of expressed sequence tag-linked microsatellite loci, prevent the detection of local adaptation in Mediterranean bluefin tuna. Fisheries Research, 190(??):157–163, June 2017. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783617300486.

Romo-Curiel:2015:OBG

[RCHSN^{+}15] Alfonsina E. Romo-Curiel, Sharon Z. Herzka, Oscar Sosa-Nishizaki, Chugey A. Sepulveda, and Scott A. Aalbers.

Otolith-based growth estimates and insights into population structure of White Seabass, Atractoscion nobilis, off the Pacific coast of North America. Fisheries Research, 161(??):374–383, January 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614002719.

Reynolds:2018:MER

[RCLS18] Emily M. Reynolds, James H. Cowan, Kristy A. Lewis, and Kirsten A. Simonsen. Method for estimating relative abundance and species composition around oil and gas platforms in the northern Gulf of Mexico, U.S.A. Fisheries Research, 201(??):44–55, May 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361830002X.

Richards:2015:EMO

[RDM^{+}15] Russell G. Richards, Andrew T. Davidson, Jan-Olaf Meynecke, Kerrod Beattie, Vanessa Hernaman, Tim Lynam, and Ingrid E. van Putten. Effects and mitigations of ocean acidification on wild and aquaculture scallop and prawn fisheries in Queensland, Australia. Fisheries Research, 161(??):42–56, January 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614002100.

Rios:2017:OII

[RDP^{+}17] Noelia Ríos, Matija Drakulic, Iosu Paradinas, Anastasia Milliou, and Ruth Cox. Occurrence and impact of interactions between small-scale fisheries and predators, with focus on Mediterranean monk seals (Monachus monachus Hermann 1779), around Lipsi Island complex, Aegean Sea, Greece. Fisheries Research, 187(??):1–10, March 2017. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783616303411.

Ralston:2018:PMS

[RDR^{+}18] Stephen Ralston, Emmanis Dorval, Laura Ryley, Keith M. Sakuma, and John C. Field. Predicting market squid (Doryteuthis opalescens) landings from pre-recruit abundance. Fisheries Research, 199(??):12–18, March 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361730317X.

Reubens:2011:AFB

[RDV11] J. T. Reubens, S. Degraer, and M. Vincx. Aggregation and feeding behaviour of pouting (Trisopterus luscus) at wind turbines in the Belgian part of the North Sea. Fisheries Research, 108(1):223–227, February 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610003103.

Riepe:2017:WDB

[RFC^{+}17] Carsten Riepe, Marie Fujitani, Julien Cucherousset, Thilo Pagel, Mathieu Buoro, Frédéric Santoul, Rémy Lassus, and Robert Arlinghaus. What determines the behavioral intention of local-level fisheries managers to alter fish stocking practices in freshwater recreational fisheries of two European countries? Fisheries Research, 194(??):173–187, October 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783617301480.

Rey:2012:DOM

[RFPE^{+}12] Javier Rey, Lourdes Fernández-Peralta, Alba Esteban, Ramón García-Cancela, Francisca Salmerón, Miguel Ángel Puerto, and Carmen Piñeiro. Does otolith macrostructure record environmental or biological events? The case of black hake (Merluccius polli and Merluccius senegalensis). Fisheries Research, 113(1):159–172, January 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611003304.

Rey:2016:OMA

[RFPG^{+}16] Javier Rey, Lourdes Fernández-Peralta, Alberto García, Enrique Nava, María Carmen Clemente, Pablo Otero, Elisa Isabel Villar, and Carmen Gloria Piñeiro. Otolith microstructure analysis reveals differentiated growth histories in sympatric black hakes (Merluccius polli and Merluccius senegalensis). Fisheries Research, 179(??):280–290, July 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783616300844.

Roeger:2016:WFF

[RFS16] Jessica Roeger, Simon Foale, and Marcus Sheaves. When ‘fishing down the food chain’ results in improved food security: Evidence from a small pelagic fishery in Solomon Islands. Fisheries Research, 174(??):250–259, February 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783615301132.

Raoult:2018:RBS

[RG18] V. Raoult and T. F. Gaston. Rapid biomass and sizefrequency estimates of edible jellyfish populations using drones. Fisheries Research, 207(??):160–164, November 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783618301796.

Rodriguez:2018:REA

[RGAP18] Helena Rodríguez, Ángel F. González, Elvira Abollo, and Santiago Pascual. Re-evaluation of anchovies (Engraulis encrasicolus) as an important risk factor for sensitization to zoonotic nematodes in Spain. Fisheries Research, 202(??):49–58, June 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783617303211.

Regueira:2014:HSP

[RGG14] M. Regueira, A. F. González, and A. Guerra. Habitat selection and population spreading of the horned octopus Eledone cirrhosa (Lamarck, 1798) in Galician waters (NW Atlantic). Fisheries Research, 152(??):66–73, April 2014. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783613002130.

Robinson:2013:JSD

[RGGdL13] Carlos J. Robinson, Jaime Gómez-Gutiérrez, and David Alberto Salas de León. Jumbo squid (Dosidicus gigas) landings in the Gulf of California related to remotely sensed SST and concentrations of chlorophyll a (1998–2012). Fisheries Research, 137(??):97–103, January 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612002755.

Ravago-Gotanco:2019:RGS

[RGK19] Rachel Ravago-Gotanco and Kenneth M. Kim. Regional genetic structure of sandfish Holothuria (Metriatyla) scabra populations across the Philippine archipelago. Fisheries Research, 209(??):143–155, January 2019. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618302625.

Rodriguez-Garcia:2019:RTV

[RGSAMV19] Oswaldo U. Rodríguez-García, Juan Carlos Seijo, Juan A. De Anda-Montañez, and Raul Villanueva. Recovery timelines of vulnerable high value species under moratoria: Dealing with uncertain levels of illegal fishing. Fisheries Research, 220(??):Article 105345, December 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619302000.

Rosen:2013:DTI

[RH13] Shale Rosen and Jens Christian Holst. DeepVision in-trawl imaging: Sampling the water column in four dimensions. Fisheries Research, 148(??):64–73, November 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613001835.

Rudershausen:2016:COT

[RHB16] P. J. Rudershausen, J. E. Hightower, and J. A. Buckel. Can optimal trap mesh size be predicted from body depth in a laterally-compressed fish species? Fisheries Research, 179(??):259–270, July 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616300637.

Rankin:2017:DEC

[RHB^{+}17] Polly S. Rankin, Robert W. Hannah, Matthew T. O. Blume, Timothy J. Miller-Morgan, and Jerry R. Heidel. Delayed effects of capture-induced barotrauma on physical condition and behavioral competency of recompressed yelloweye rockfish, Sebastes ruberrimus. Fisheries Research, 186 (part 1)(??):258–268, February 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616302971.

Rapp:2012:PBC

[RHC^{+}12] Tobias Rapp, Jan Hallermann, Steven J. Cooke, Stefan K. Hetz, Sven Wuertz, and Robert Arlinghaus. Physiological and behavioural consequences of capture and retention in carp sacks on common carp (Cyprinus carpio L.), with implications for catch-and-release recreational fishing. Fisheries Research, 125–126(??):57–68, August 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612000847.

Rhodes:2018:YCP

[RHOCB^{+}18] Kevin L. Rhodes, Dalia X. Hernandez-Ortiz, Javier Cuetos-Bueno, McKye Ioanis, Welbert Washington, and Ryan Ladore. A 10-year comparison of the Pohnpei, Micronesia, commercial inshore fishery reveals an increasingly unsustainable fishery. Fisheries Research, 204(??):156–164, August 2018. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783618300535.

Runcie:2016:FIS

[RHW^{+}16] Rosa Runcie, David Holts, James Wraith, Yi Xu, Darlene Ramon, Rand Rasmussen, and Suzanne Kohin. A fishery-independent survey of juvenile shortfin mako (Isurus oxyrinchus) and blue (Prionace glauca) sharks in the Southern California Bight, 1994–2013. Fisheries Research, 183(??):233–243, November 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616301990.

Riley:2010:PCE

[RIB^{+}10] W. D. Riley, A. T. Ibbotson, W. R. C. Beaumont, P. Rycroft, and A. C. Cook. A portable, cost effective, pass-through system to detect downstream migrating salmonids marked with 12 mm passive integrated transponder tags. Fisheries Research, 101(3):203–206, January 15, 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783609002677.

Richards:2016:SPP

[RJ16] R. Anne Richards and Larry D. Jacobson. A simple predation pressure index for modeling changes in natural mortality: Application to Gulf of Maine northern shrimp stock assessment. Fisheries Research, 179(??):224–236, July 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783616300595.

Rotherham:2012:SEF

[RJKG12] Douglas Rotherham, Daniel D. Johnson, Caitlin L. Kesby, and Charles A. Gray. Sampling estuarine fish and invertebrates with a beam trawl provides a different picture of populations and assemblages than multi-mesh gillnets. Fisheries Research, 123–124(??):49–55, July 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611003481.

Riha:2012:SSM

[ŘJP^{+}12] Milan Říha, Tomáš Jůza, Marie Prchalová, Tomáš Mrkvička, Martin Čech, Vladislav Draštík, Milan Muška, Michal Kratochvíl, Jiří Peterka, Michal Tušer, Mojmír Vašek, and Jan Kubečka. The size selectivity of the main body of a sampling pelagic pair trawl in freshwater reservoirs during the night. Fisheries Research, 127–128(??):56–60, September 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783612001415.

Rogers:2017:PRS

[RKH^{+}17] P. J. Rogers, I. Knuckey, R. J. Hudson, A. D. Lowther, and L. Guida. Post-release survival, movement, and habitat use of school shark Galeorhinus galeus in the Great Australian Bight, Southern Australia. Fisheries Research, 187(??):188– 198, March 2017. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783616303964.

Reid:2012:CGS

[RKP^{+}12] D. G. Reid, R. J. Kynoch, I. Penny, K. Summerbell, A. Edridge, and F. G. O’Neill. A comparison of the GOV survey trawl with a commercial whitefish trawl. Fisheries Research, 121–122(??):136–143, June 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612000525.

Ramos:2015:PIO

[RLC^{+}15] Jorge Ramos, Pedro G. Lino, Miguel Caetano, Fábio Pereira, Miguel Gaspar, and Miguel Neves dos Santos. Perceived impact of offshore aquaculture area on smallscale fisheries: a fuzzy logic model approach. Fisheries Research, 170(??):217–227, October 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615001824.

Ramirez:2017:ISA

[RLC^{+}17] John Gabriel Ramírez, Jordi Lleonart, Marta Coll, Francisco Reyes, and Gina Marcela Puentes. Improving stock assessment and management advice for data-poor smallscale fisheries through participatory monitoring. Fisheries Research, 190(??):71–83, June 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617300243.

Rogers:2011:RVP

[RLFJ11] Bonnie L. Rogers, Christopher G. Lowe, and Esteban Fernández-Juricic. Recovery of visual performance in rosy rockfish (Sebastes rosaceus) following exophthalmia resulting from barotrauma. Fisheries Research, 112(1–2):1–7, December 2011. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783611002700.

Radhakrishnan:2012:AOS

[RLJ^{+}12] K. V. Radhakrishnan, Yuxuan Li, K. V. Jayalakshmy, Ming Liu, Brian R. Murphy, and Songguang Xie. Application of otolith shape analysis in identifying different ecotypes of Coilia ectenes in the Yangtze Basin, China. Fisheries Research, 125–126(??):156–160, August 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612001002.

Ramon:2010:RCS

[RLM10] Montserrat Ramón, Jordi Lleonart, and Enric Massutí. Royal cucumber (Stichopus regalis) in the northwestern Mediterranean: Distribution pattern and fishery. Fisheries Research, 105(1):21–27, June 2010. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610000536.

Rabaoui:2017:NHL

[RLM^{+}17] Lotfi Rabaoui, Yu-Jia Lin, Rommel H. Maneja, Mohammad A. Qurban, Pulikkodan Abdurahiman, Panickan Premlal, Khaled Al-Abdulkader, and Rubén H. Roa-Ureta. Nursery habitats and life history traits of the green tiger shrimp Penaeus semisulcatus (De Haan, 1844) in the Saudi waters of the Arabian Gulf. Fisheries Research, 195(??):1–11, November 2017. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783617301686.

Rodrigues:2018:MCF

[RLMM18] Karina A. Rodrigues, Ezequiel Leonarduzzi, Gustavo J. Macchi, and Maria I. Militelli. Maternal condition, fecundity and oocyte quality of Argentine hake (Merluccius hubbsi) from the northern stock. Fisheries Research, 197(??):105–112, January 2018. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617302539.

Rosas-Luis:2014:FHT

[RLSPdR14] Rigoberto Rosas-Luis, Pilar Sánchez, Julio M. Portela, and José Luis del Rio. Feeding habits and trophic interactions of Doryteuthis gahi, Illex argentinus and Onykia ingens in the marine ecosystem off the Patagonian shelf. Fisheries Research, 152(??):37–44, April 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613002828.

Rosas-Luis:2014:THO

[RLVS14] Rigoberto Rosas-Luis, Roger Villanueva, and Pilar Sánchez. Trophic habits of the ommastrephid squid Illex coindetii and Todarodes sagittatus in the northwestern Mediterranean Sea. Fisheries Research, 152(??):21–28, April 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613002361.

Rasheed:2010:SSM

[RM10a] Shazia Rasheed and Javed Mustaquim. Size at sexual maturity, breeding season and fecundity of three-spot swimming crab Portunus sanguinolentus (Herbst, 1783) (Decapoda, Brachyura, Portunidae) occurring in the coastal waters of Karachi, Pakistan. Fisheries Research, 103(1–3):56– 62, April 2010. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783610000469.

Rodrigues:2010:SRP

[RM10b] Karina A. Rodrigues and Gustavo G. Macchi. Spawning and reproductive potential of the northern stock of Argentine hake (Merluccius hubbsi). Fisheries Research, 106(3):560–566, December 2010. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610002201.

Romeo:2019:EEP

[RM19] Giuseppa Romeo and Claudio Marcianò. Evaluating the economic performance of fishing systems using fuzzy multicriteria analysis in a fishery local action group in South Italy. Fisheries Research, 218(??):259–268, October 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619301456.

Rudershausen:2021:DTS

[RMB^{+}21] P. J. Rudershausen, W. A. Mitchell, J. A. Buckel, E. H. Williams, and E. Hazen. Developing a two-step fisheryindependent design to estimate the relative abundance of deepwater reef fish: Application to a marine protected area off the southeastern United States coast. Fisheries Research, 105(3):254–260, ???? 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610001086.

Rumisha:2018:GDG

[RMGK18] Cyrus Rumisha, Robinson H. Mdegela, Paul S. Gwakisa, and Marc Kochzius. Genetic diversity and gene flow among the giant mud crabs (Scylla serrata) in anthropogenic-polluted mangroves of mainland Tanzania: Implications for conservation. Fisheries Research, 205(??):96–104, September 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783618301140.

Rodriguez-Mendoza:2019:SCB

[RMMSR19] Rebeca Rodríguez-Mendoza, Marta Muñoz, and Fran Saborido-Rey. Structure and connectivity of bluemouth, Helicolenus dactylopterus, populations in the NE Atlantic and Mediterranean. Fisheries Research, 213(??):56–66, May 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783619300062.

Russo:2018:MCL

[RMP^{+}18] T. Russo, E. B. Morello, A. Parisi, G. Scarcella, S. Angelini, L. Labanchi, M. Martinelli, L. D’Andrea, A. Santojanni, E. Arneri, and S. Cataudella. A model combining landings and VMS data to estimate landings by fishing ground and harbor. Fisheries Research, 199(??):218–230, March 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783617303107.

Raby:2018:SIE

[RMT^{+}18] Graham D. Raby, Vanessa Messmer, Andrew J. Tobin, Andrew S. Hoey, Fredrik Jutfelt, Josefin Sundin, Steven J. Cooke, and Timothy D. Clark. Swim for it: Effects of simulated fisheries capture on the post-release behaviour of four Great Barrier Reef fishes. Fisheries Research, 206(??):129–137, October 2018. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618301486.

Robertson:2015:CRR

[Rob15] Bruce C. Robertson. Comment on “Review of research and assessments on the efficacy of sea lion exclusion devices in reducing the incidental mortality of New Zealand sea lions Phocarctos hookeri in the Auckland Islands squid trawl fishery”. Fisheries Research, 165(??):127–129, May 2015. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783614003580.

Richard:2013:EPI

[ROCC13] Alexandre Richard, Jane O’Rourke, Arnaud Caudron, and Franck Cattanéo. Effects of passive integrated transponder tagging methods on survival, tag retention and growth of age-0 brown trout. Fisheries Research, 145(??):37–42, August 2013. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783613000672.

Rodgveller:2019:ULA

[Rod19] Cara J. Rodgveller. The utility of length, age, liver condition, and body condition for predicting maturity and fecundity of female sablefish. Fisheries Research, 216(??):18–28, August 2019. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783619300785.

Russo:2011:NII

[RPC11] Tommaso Russo, Antonio Parisi, and Stefano Cataudella. New insights in interpolating fishing tracks from VMS data for different métiers. Fisheries Research, 108(1):184–194, February 2011. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783610003450.

Reid:2018:CEG

[RPH^{+}18] Kerry Reid, Eric P. Palkovacs, Daniel J. Hasselman, Diana Baetscher, Jared Kibele, Ben Gahagan, Paul Bentzen, Meghan C. McBride, and John Carlos Garza. Comprehensive evaluation of genetic population structure for anadromous river herring with single nucleotide polymorphism data. Fisheries Research, 206(??):247–258, October 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783618301139.

Robbins:2011:APM

[RPK11] W. D. Robbins, V. M. Peddemors, and S. J. Kennelly. Assessment of permanent magnets and electropositive metals to reduce the line-based capture of Galapagos sharks, Carcharhinus galapagensis. Fisheries Research, 109(1):100–106, April 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783611000427.

Russo:2011:WBR

[RPP^{+}11] Tommaso Russo, Antonio Parisi, Marina Prorgi, Fabrizio Boccoli, Innocenzo Cignini, Maurizio Tordoni, and Stefano Cataudella. When behaviour reveals activity: Assigning fishing effort to métiers based on VMS data using artificial neural networks. Fisheries Research, 111(1–2):53–64, September 2011. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783611002281.

Roth:2012:SAC

[RR12] Bjorn Roth and Bjørn Tore Rotabakk. Stress associated with commercial longlining and recreational fishing of saithe (Pollachius virens) and the subsequent effect on blood gases and chemistry. Fisheries Research, 115–116(??):110–114, March 2012. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783611001925.

Rogers:2016:LLH

[RRM16] Robert Rogers, Sherrylynn Rowe, and M. Joanne Morgan. Lack of life history change in two over-exploited haddock (Melanogrammus aeglefinus) stocks. Fisheries Research, 181(??):77–83, September 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616300984.

Rezende:2019:MST

[RRO^{+}19] Gabriela A. Rezende, Marie-Christine Rufener, Ileana Ortega, Vinicius Mendes Ruas, and Luiz Felipe C. Dumont. Modelling the spatio-temporal bycatch dynamics in an estuarine small-scale shrimp trawl fishery. Fisheries Research, 219(??):Article 105336, November 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619301912.

Richards:2018:PMR

[RRPG18] R. J. Richards, V. Raoult, D. M. Powter, and T. F. Gaston. Permanent magnets reduce bycatch of benthic sharks in an ocean trap fishery. Fisheries Research, 208(??):16–21, December 2018. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783618301966.

Rogers:2019:FHM

[RRRM19] Robert Rogers, Sherrylynn Rowe, Rick M. Rideout, and M. Joanne Morgan. Fecundity of haddock (Melanogrammus aeglefinus) off southern Newfoundland. Fisheries Research, 220(??):Article 105339, December 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619301948.

Runde:2019:LDS

[RRS^{+}19] Brendan J. Runde, Paul J. Rudershausen, Beverly Sauls, Chloe S. Mikles, and Jeffrey A. Buckel. Low discard survival of gray triggerfish in the southeastern US hook-and-line fishery. Fisheries Research, 219(??):Article 105313, November 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783619301602.

Rees:2013:IIO

[RRSB13] Siân E. Rees, Lynda D. Rodwell, Spike Searle, and Andrew Bell. Identifying the issues and options for managing the social impacts of Marine Protected Areas on a small fishing community. Fisheries Research, 146(??):51–58, September 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783613000908.

Romero:2013:UDE

[RRW^{+}13] María Alejandra Romero, Matías Ocampo Reinaldo, Gabriela Williams, Maite Narvarte, Domingo Antonio Gagliardini, and Raúl González. Understanding the dynamics of an enclosed trawl demersal fishery in Patagonia (Argentina): a holistic approach combining multiple data sources. Fisheries Research, 140(??):73–82, February 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783612003463.

Rotabakk:2011:QAA

[RSAB11] Bjørn Tore Rotabakk, Dagbjørn Skipnes, Leif Akse, and Sveinung Birkeland. Quality assessment of Atlantic cod (Gadus morhua) caught by longlining and trawling at the same time and location. Fisheries Research, 112(1–2):44–51, December 2011. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783611002785.

Richmond:2014:QBR

[RSE^{+}14] Marshall C. Richmond, John A. Serkowski, Laurie L. Ebner, Mirjam Sick, Richard S. Brown, and Thomas J. Carlson. Quantifying barotrauma risk to juvenile fish during hydroturbine passage. Fisheries Research, 154(??):152–164, June 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783614000095.

Rodriguez-Sanchez:2016:HTS

[RSEERR^{+}16] Victoria Rodríguez-Sánchez, Lourdes Encina-Encina, Amadora Rodríguez-Ruiz, Agustín Monteoliva, and Ramona Sánchez-Carmona. Horizontal target strength of Cyprinus carpio using 200 kHz and 430 kHz split-beam systems. Fisheries Research, 174(??):136–142, February 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783615300874.

Rodriguez-Sanchez:2015:HTS

[RSEERRSC15] Victoria Rodríguez-Sánchez, Lourdes Encina-Encina, Amadora Rodríguez-Ruiz, and Ramona Sánchez-Carmona. Horizontal target strength of Luciobarbus sp. in ex situ experiments: Testing differences by aspect angle, pulse length and beam position. Fisheries Research, 164(??):214–222, April 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783614003464.

Rodriguez-Sanchez:2016:DCR

[RSEERRSC16] Victoria Rodríguez-Sánchez, Lourdes Encina-Encina, Amadora Rodríguez-Ruiz, and Ramona Sánchez-Carmona. Do close range measurements affect the target strength (TS) of fish in horizontal beaming hydroacoustics? Fisheries Research, 173(??):4–10, January 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783615001101.

Robinson:2011:ITS

[RSG^{+}11] Jan Robinson, Melita A. Samoilys, Edwin Grandcourt, Danny Julie, Maria Cedras, and Calvin Gerry. The importance of targeted spawning aggregation fishing to the management of Seychelles’ trap fishery. Fisheries Research, 112(1–2):96–103, December 2011. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611002840.

Ryan:2019:EST

[RSGR19] Diarmuid Ryan, Samuel Shephard, Patrick Gargan, and William Roche. Estimating sea trout (Salmo trutta L.) growth from scale chemistry profiles: an objective approach using LA-ICPMS. Fisheries Research, 211(??):69–80, March 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783618302984.

Roth:2018:FAE

[RSQ18] Curtis J. Roth, Daniel J. Schill, and Michael C. Quist. Fight and air exposure times of caught and released salmonids from the South Fork Snake River. Fisheries Research, 201(??):38–43, May 2018. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618300146.

Rohtla:2017:DNM

[RSV^{+}17] Mehis Rohtla, Roland Svirgsden, Aare Verliin, Kateriina Rumvolt, Lagle Matetski, Kristiina Hommik, Lauri Saks, and Markus Vetemaa. Developing novel means for unravelling population structure, provenance and migration patterns of European whitefish Coregonus lavaretus s.l. in the Baltic Sea. Fisheries Research, 187(??):47–57, March 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783616303575.

Robertis:2014:STS

[RT14] Alex De Robertis and Kevin Taylor. In situ target strength measurements of the scyphomedusa Chrysaora melanaster. Fisheries Research, 153(??):18–23, May 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614000046.

Rodgveller:2017:FUP

[RTML17] Cara J. Rodgveller, Cindy A. Tribuzio, Patrick W. Malecha, and Chris R. Lunsford. Feasibility of using pop-up satellite archival tags (PSATs) to monitor vertical movement of a Sebastes: a case study. Fisheries Research, 187(??):96–102, March 2017. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783616303976.

Rakowitz:2012:UHF

[RTŘ^{+}12] Georg Rakowitz, Michal Tušer, Milan Říha, Tomáš Jůza, Helge Balk, and Jan Kubečka. Use of high-frequency imaging sonar (DIDSON) to observe fish behaviour towards a surface trawl. Fisheries Research, 123–124(??):37–48, July 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578361100347X.

Rohtla:2015:OTB

[RTSV15] Mehis Rohtla, Imre Taal, Roland Svirgsden, and Markus Vetemaa. Old timers from the Baltic Sea: Revisiting the population structure and maximum recorded age of ide Leuciscus idus. Fisheries Research, 165(??):74–78, May 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783615000132.

Roa-Ureta:2015:SAS

[RU15] Rubén H. Roa-Ureta. Stock assessment of the Spanish mackerel (Scomberomorus commerson) in Saudi waters of the Arabian Gulf with generalized depletion models under data-limited conditions. Fisheries Research, 171(??):68–77, November 2015. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783614002598.

Roa-Ureta:2015:HSF

[RUMBA15] Ruben H. Roa-Ureta, Carlos Molinet, Nancy Barahona, and Pablo Araya. Hierarchical statistical framework to combine generalized depletion models and biomass dynamic models in the stock assessment of the Chilean sea urchin (Loxechinus albus) fishery. Fisheries Research, 171(??):59–67, November 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783614003622.

Roch:2018:FHI

[RvAGB18] S. Roch, L. von Ammon, J. Geist, and A. Brinker. Foraging habits of invasive three-spined sticklebacks (Gasterosteus aculeatus) — impacts on fisheries yield in Upper Lake Constance. Fisheries Research, 204(??):172–180, August 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578361830050X.

Ruesink:2014:STR

[RvRW^{+}14] J. L. Ruesink, K. van Raay, A. Witt, S. Herrold, N. Freshley, A. Sarich, and A. C. Trimble. Spatio-temporal recruitment variability of naturalized Manila clams (Ruditapes philippinarum) in willapa Bay, Washington, USA. Fisheries Research, 151(??):199–204, March 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613002890.

Rose:2019:SCI

[RW19] George A. Rose and Carl J. Walters. The state of Canada’s iconic northern cod: a second opinion. Fisheries Research, 219(??):Article 105314, November 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619301614.

Rooper:2015:EUL

[RWRT15] Christopher N. Rooper, Kresimir Williams, Alex De Robertis, and Vanessa Tuttle. Effect of underwater lighting on observations of density and behavior of rockfish during camera surveys. Fisheries Research, 172(??):157–167, December 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783615300217.

Salmi:2016:CCO

[SA16] Juhani A. Salmi and Heikki Auvinen. Comments on the criticism in ‘Overestimated effect of cormorant predation on fisheries catches’ presented by Heikinheimo and Lehtonen, 2015. Fisheries Research, 179(??):358–360, July 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783616300674. See [HL16, SAR^{+}15].

Smith:2011:MSD

[SAB^{+}11] Steven G. Smith, Jerald S. Ault, James A. Bohnsack, Douglas E. Harper, Jiangang Luo, and David B. McClellan. Multispecies survey design for assessing reef-fish stocks, spatially explicit management performance, and ecosystem condition. Fisheries Research, 109(1):25–41, April 2011. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783611000312.

Stobart:2012:EHP

[SABG12] Ben Stobart, Diego Alvarez-Barastegui, and Raquel Goñi. Effect of habitat patchiness on the catch rates of a Mediterranean coastal bottom long-line fishery. Fisheries Research, 129–130(??):110–118, October 2012. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612001981.

Stratoudakis:2015:BDL

[SAF^{+}15] Yorgos Stratoudakis, Manuela Azevedo, Inês Farias, Carlos Macedo, Teresa Moura, Manuel J. Pólvora, Cristina Rosa, and Ivone Figueiredo. Benchmarking for data-limited fishery systems to support collaborative focus on solutions. Fisheries Research, 171(??):122–129, November 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614002872.

Schmidt:2015:PRA

[SAH15] Beatriz F. Schmidt, Alberto F. Amorim, and Alexandre W. S. Hilsdorf. PCR–RFLP analysis to identify four ray species of the genus Dasyatis (Elasmobranchii, Dasyatidae) fished along the southeastern and southern coast of Brazil. Fisheries Research, 167(??):71–74, July 2015. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783615000314.

Schuchert:2010:TAC

[SAK10] Pia C. Schuchert, Alexander I. Arkhipkin, and Alan E. Koenig. Traveling around Cape Horn: Otolith chemistry reveals a mixed stock of Patagonian hoki with separate Atlantic and Pacific spawning grounds. Fisheries Research, 102(1–2):80–86, February 2010. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783609002768.

Sharawy:2017:DAD

[SAK^{+}17a] Zaki Zaki Sharawy, Eman Mamdouh Abbas, Amal Ragae Khafage, Asmaa Galal-Khallaf, Rania Fahmy Ismail, Hamdy Omar Ahmed, Khaled Mohammed-Geba, and Mikio Kato. Descriptive analysis, DNA barcoding and condition index of penaeids (Crustacea: Decapoda) from the Egyptian Mediterranean coast. Fisheries Research, 188(??):6–16, April 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783616304179.

Smichi:2017:ESQ

[SAK^{+}17b] Nabil Smichi, Baha Eddine Abdelmalek, Nadia Kharrat, Assaad Sila, Ali bougatef, Youssef Gargouri, and Ahmed Fendri. The effects of storage on quality and nutritional aspects of farmed and wild sea bass (Dicentrachus labrax) muscle: In vitro oils digestibility evaluation. Fisheries Research, 188(??):74–83, April 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616304155.

Salenius:2018:IMN

[Sal18] Fredrik Salenius. International management of North Atlantic pelagic fisheries — the role of competing species and exploiters. Fisheries Research, 203(??):12–21, July 2018.

CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783617302151.

Sampson:2011:ASR

[Sam11] David B. Sampson. The accuracy of self-reported fisheries data: Oregon trawl logbook fishing locations and retained catches. Fisheries Research, 112(1–2):59–76, December 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783611002815.

Sampson:2014:FSR

[Sam14] David B. Sampson. Fishery selection and its relevance to stock assessment and fishery management. Fisheries Research, 158(??):5–14, October 2014. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613002312.

Salmi:2015:PPF

[SAR^{+}15] Juhani A. Salmi, Heikki Auvinen, Jari Raitaniemi, Mika Kurkilahti, Juha Lilja, and Riikka Maikola. Perch (Perca fluviatilis) and pikeperch (Sander lucioperca) in the diet of the great cormorant (Phalacrocorax carbo) and effects on catches in the Archipelago Sea, Southwest coast of Finland. Fisheries Research, 164(??):26–34, April 2015. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578361400304X. See letter [HL16] and

response [SA16].

Sauls:2014:RSG

[Sau14] Beverly Sauls. Relative survival of gags Mycteroperca microlepis released within a recreational hook-and-line fishery: Application of the Cox regression model to control for heterogeneity in a large-scale mark-recapture study. Fisheries Research, 150(??):18–27, February 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361300235X.

Streich:2018:HSP

[SAWS18] Matthew K. Streich, Matthew J. Ajemian, Jennifer J. Wetz, and Gregory W. Stunz. Habitat-specific performance of vertical line gear in the western Gulf of Mexico: a comparison between artificial and natural habitats using a paired video approach. Fisheries Research, 204(??):16–25, August 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783618300250.

Smallwood:2012:SDZ

[SB12] Claire B. Smallwood and Lynnath E. Beckley. Spatial distribution and zoning compliance of recreational fishing in Ningaloo Marine Park, north-western Australia. Fisheries Research, 125–126(??):40–50, August 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612000501.

Sumaila:2014:FEJ

[SB14] U. Rashid Sumaila and Mahamudu Bawumia. Fisheries, ecosystem justice and piracy: a case study of Somalia. Fisheries Research, 157(??):154–163, September 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361400143X.

Spaet:2015:FMS

[SB15] Julia L. Y. Spaet and Michael L. Berumen. Fish market surveys indicate unsustainable elasmobranch fisheries in the Saudi Arabian Red Sea. Fisheries Research, 161(??):356–364, January 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783614002689.

Steer:2016:LAS

[SB16] Mike Steer and Michelle Besley. The licence amalgamation scheme: Taming South Australia’s complex multispecies, multi-gear marine scalefish fishery. Fisheries Research, 183(??):625–633, November 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361630145X.

Sterling:2017:ECP

[SB17] David Sterling and Cheslav Balash. Engineering and catching performance of five netting materials in commercial prawn-trawl systems. Fisheries Research, 193(??):223–231, September 2017. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783617301194.

Sypitkowski:2010:EFE

[SBA10] Eben Sypitkowski, Curtis Bohlen, and William G. Ambrose. Estimating the frequency and extent of bloodworm digging in Maine from aerial photography. Fisheries Research, 101(1–2):87–93, January 5, 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783609002562.

Schwinn:2017:CSM

[SBAK17] Michael Schwinn, Henrik Baktoft, Kim Aarestrup, and Anders Koed. A comparison of the survival and migration of wild and F1-hatchery-reared brown trout (Salmo trutta) smolts traversing an artificial lake. Fisheries Research, 196(??):47–55, December 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617302254.

Smejkal:2019:SSP

[ŠBB^{+}19] Marek Šmejkal, Petr Blabolil, Daniel Bartoň, Jindřich Duras, Lukáš Vejřík, Zuzana Sajdlová, Luboš Kočvara, and Jan Kubečka. Sex-specific probability of PIT tag retention in a cyprinid fish. Fisheries Research, 219(??):Article 105325, November 2019. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783619301729.

Sulikowski:2018:ECD

[SBC^{+}18] James A. Sulikowski, Hugues P. Benoît, Connor W. Capizzano, Ryan J. Knotek, John W. Mandelman, Ted Platz, and David B. Rudders. Evaluating the condition and discard mortality of winter skate, Leucoraja ocellata, following capture and handling in the Atlantic monkfish (Lophius americanus) sink gillnet fishery. Fisheries Research, 198(??):159–164, February 2018. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617302692.

Silva:2012:ENG

[SBD^{+}12] Catarina N. S. Silva, Matt K. Broadhurst, José H. Dias, André P. Cattani, and Henry L. Spach. The effects of Nordmøre-grid bar spacings on catches in a Brazilian artisanal shrimp fishery. Fisheries Research, 127– 128(??):188–193, September 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783612000355.

Sistiaga:2016:SSP

[SBH^{+}16] Manu Sistiaga, Jesse Brinkhof, Bent Herrmann, Eduardo Grimaldo, Lise Langård, and Dagfinn Lilleng. Size selective performance of two flexible sorting grid designs in the Northeast Arctic cod (Gadus morhua) and haddock (Melanogrammus aeglefinus) fishery. Fisheries Research, 183(??):340–351, November 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616302119.

Stewart:2019:EDM

[SBJ^{+}19] David R. Stewart, Matthew J. Butler, Lacrecia A. Johnson, Aaron Cajero, Amber N. Young, and Grant M. Harris. Efficacy of depletion models for estimating abundance of endangered fishes in streams. Fisheries Research, 209(??):208–217, January 2019. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783618302674.

Szuwalski:2019:GFF

[SBL^{+}19] Cody S. Szuwalski, Gregory L. Britten, Roberto Licandeo, Ricardo O. Amoroso, Ray Hilborn, and Carl Walters. Global forage fish recruitment dynamics: a comparison of methods, time-variation, and reverse causality.

Fisheries Research, 214(??):56–64, June 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619300074.

Shahrestani:2017:DNF

[SBLB17] Suzan Shahrestani, Hongsheng Bi, Viacheslav Lyubchich, and Kevin M. Boswell. Detecting a nearshore fish parade using the adaptive resolution imaging sonar (ARIS): an automated procedure for data analysis. Fisheries Research, 191(??):190–199, July 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617300802.

Spencer:2017:POP

[SBLL17] D. M. Spencer, I. W. Brown, S. Y. Lee, and C. J. Lemckert. Physical oceanographic processes affecting catchability of spanner crab (Ranina ranina) — a review. Fisheries Research, 186 (part 1)(??):248–257, February 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616302983.

Stallings:2014:CCB

[SBLM14] Christopher D. Stallings, Joshua P. Brower, Jennifer M. Heinlein Loch, and Alejandra Mickle. Catch comparison between otter and rollerframe trawls: Implications for sampling in seagrass beds. Fisheries Research, 155(??):177– 184, July 2014. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783614000848.

Segvic-Bubic:2016:TST

[ŠBMG^{+}16] Tanja Šegvić-Bubić, Federico Marrone, Leon Grubišić, David Izquierdo-Gomez, Ivan Katavić, Marco Arculeo, and Sabrina Lo Brutto. Two seas, two lineages: How genetic diversity is structured in Atlantic and Mediterranean greater amberjack Seriola dumerili Risso, 1810 (Perciformes, Carangidae). Fisheries Research, 179(??):271–279, July 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578361630087X.

Silva:2011:RNG

[SBS^{+}11] Catarina N. S. Silva, Matt K. Broadhurst, Andréia Schwingel, José H. Dias, André P. Cattani, and Henry L. Spach. Refining a Nordmøre-grid for a Brazilian artisanal penaeid-trawl fishery. Fisheries Research, 109(1):168–178, April 2011. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783611000555.

Schramm:2017:EHT

[SBS17] Michael P. Schramm, Mark Bevelhimer, and Constantin Scherelis. Effects of hydrokinetic turbine sound on the behavior of four species of fish within an experimental mesocosm. Fisheries Research, 190(??):1–14, June 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617300218.

Stevens:2018:IVC

[SBT^{+}18] Philip W. Stevens, Ross E. Boucek, Alexis A. Trotter, Jared L. Ritch, Eric R. Johnson, Colin P. Shea, David A. Blewett, and Jennifer S. Rehage. Illustrating the value of cross-site comparisons: Habitat use by a large, euryhaline fish differs along a latitudinal gradient. Fisheries Research, 208(??):42–48, December 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618301954.

Santos:2019:DPI

[SC19] Catarina C. Santos and Rui Coelho. Distribution patterns and indicators of the smooth hammerhead shark (Sphyrna zygaena) in the Atlantic Ocean. Fisheries Research, 212(??):107–113, April 2019. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361830359X.

Satterthwaite:2015:SCO

[SCC^{+}15] William H. Satterthwaite, Javier Ciancio, Eric Crandall, Melodie L. Palmer-Zwahlen, Allen M. Grover, Michael R. O’Farrell, Eric C. Anderson, Michael S. Mohr, and John Carlos Garza. Stock composition and ocean spatial distribution inference from California recreational Chinook salmon fisheries using genetic stock identification. Fisheries Research, 170(??):166–178, October 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783615001800.

Staton:2017:DEM

[SCF^{+}17a] Benjamin A. Staton, Matthew J. Catalano, Troy M. Farmer, Asheber Abebe, and F. Stephen Dobson. Development and evaluation of a migration timing forecast model for Kuskokwim River Chinook salmon. Fisheries Research, 194(??):9–21, October 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617301248.

Staton:2017:SIB

[SCF17b] Benjamin A. Staton, Matthew J. Catalano, and Steven J. Fleischman. From sequential to integrated Bayesian analyses: Exploring the continuum with a Pacific salmon spawnerrecruit model. Fisheries Research, 186 (part 1)(??):237–247, February 2017. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783616302946.

Suuronen:2012:LIF

[SCG^{+}12] Petri Suuronen, Francis Chopin, Christopher Glass, Svein Løkkeborg, Yoshiki Matsushita, Dante Queirolo, and Dominic Rihan. Low impact and fuel efficient fishing — looking beyond the horizon. Fisheries Research, 119–120(??):135– 146, May 2012. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783611003870.

Sajina:2011:SSA

[SCJ^{+}11] A. M. Sajina, S. K. Chakraborty, A. K. Jaiswar, D. G. Pazhayamadam, and Deepa Sudheesan. Stock structure analysis of Megalaspis cordyla (Linnaeus, 1758) along the Indian coast based on truss network analysis. Fisheries Research, 108(1):100–105, February 2011. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610003188.

Sun:2015:RBB

[SCL^{+}15] Chi-Lu Sun, Hsiao-Yun Chang, Tsung-Yun Liu, Su-Zan Yeh, and Yi-Jay Chang. Reproductive biology of the black marlin, Istiompax indica, off southwestern and eastern Taiwan. Fisheries Research, 166(??):12–20, June 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614002732.

Shideler:2015:LGG

[SCLS15] Geoffrey S. Shideler, David W. Carter, Christopher Liese, and Joseph E. Serafy. Lifting the goliath grouper harvest ban: Angler perspectives and willingness to pay. Fisheries Research, 161(??):156–165, January 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614002239.

Szuwalski:2016:ISA

[SCOC16] Cody S. Szuwalski, Mauricio Castrejon, Dan Ovando, and Brandon Chasco. An integrated stock assessment for red spiny lobster (Panulirus penicillatus) from the Galapagos Marine Reserve. Fisheries Research, 177(??):82–94, May 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783616300029.

Smart:2017:SDA

[SCT^{+}17] Jonathan J. Smart, Andrew Chin, Andrew J. Tobin, William T. White, Brian Kumasi, and Colin A. Simpfendorfer. Stochastic demographic analyses of the silvertip shark (Carcharhinus albimarginatus) and the common blacktip shark (Carcharhinus limbatus) from the Indo–Pacific. Fisheries Research, 191(??):95–107, July 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617300620.

Sinopoli:2012:AFA

[SCV^{+}12] Mauro Sinopoli, Luca Castriota, Pietro Vivona, Michele Gristina, and Franco Andaloro. Assessing the fish assemblage associated with FADs (Fish Aggregating Devices) in the southern Tyrrhenian Sea using two different professional fishing gears. Fisheries Research, 123–124(??):56–61, July 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783611003493.

Sanchez:2010:RBS

[SD10] P. Sánchez and M. Demestre. Relationships between somatic tissues, reproductive organs and fishing season in the commercially exploited European squid Loligo vulgaris from the western Mediterranean Sea. Fisheries Research, 106(2):125–131, November 2010. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610001797.

Spencer:2013:IWS

[SD13] Paul D. Spencer and Martin W. Dorn. Incorporation of weight-specific relative fecundity and maternal effects in larval survival into stock assessments. Fisheries Research, 138(??):159–167, February 2013. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612001592.

Sardenne:2015:DAT

[SDC^{+}15] Fany Sardenne, Emmanuelle Dortel, Gaël Le Croizier, Julien Million, Maylis Labonne, Bruno Leroy, Nathalie Bodin, and Emmanuel Chassot. Determining the age of tropical tunas in the Indian Ocean from otolith microstructures. Fisheries Research, 163(??):44–57, March 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614000903.

Sturludottir:2018:EEM

[SDE^{+}18] Erla Sturludottir, Christopher Desjardins, Bjarki Elvarsson, Elizabeth A. Fulton, Rebecca Gorton, Kai Logemann, and Gunnar Stefansson. End-to-end model of Icelandic waters using the Atlantis framework: Exploring system dynamics and model reliability. Fisheries Research, 207(??):9–24, November 2018. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783618301620.

Sajdlova:2015:FBR

[SDJ^{+}15] Z. Sajdlová, V. Draštík, T. Jůza, M. Říha, J. Frouzová, M. Čech, M. Vašek, M. Muška, P. Blabolil, M. Tušer, M. Kratochvíl, J. Peterka, T. Mrkvička, H. Balk, and J. Kubečka. Fish behaviour in response to a midwater trawl footrope in temperate reservoirs. Fisheries Research, 172(??):105–113, December 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615300084.

Soetaert:2016:SEE

[SDV^{+}16] Maarten Soetaert, Annemie Decostere, Bart Verschueren, Jimmy Saunders, Annemie Van Caelenberge, Velmurugu Puvanendran, Atle Mortensen, Luc Duchateau, Hans Polet, and Koen Chiers. Side-effects of electrotrawling: Exploring the safe operating space for Dover sole (Solea solea L.) and Atlantic cod (Gadus morhua L.). Fisheries Research, 177(??):95–103, May 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616300200.

Shephard:2018:LBI

[SDWG18] Samuel Shephard, Ian C. Davidson, Alan M. Walker, and Patrick G. Gargan. Length-based indicators and reference points for assessing data-poor stocks of diadromous trout Salmo trutta. Fisheries Research, 199(??):36–43, March 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783617303326.

Sippel:2015:UMD

[SEG^{+}15] Tim Sippel, J. Paige Eveson, Benjamin Galuardi, Chi Lam, Simon Hoyle, Mark Maunder, Pierre Kleiber, Felipe Carvalho, Vardis Tsontos, Steven L. H. Teo, Alexandre Aires da Silva, and Simon Nicol. Using movement data from electronic tags in fisheries stock assessment: a review of models, technology and experimental design. Fisheries Research, 163(??):152–160, March 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614001404.

Seung:2021:EEI

[Seu21] Chang K. Seung. Estimating economic information for fisheries using unequal probability sampling. Fisheries Research, 105(3):134–140, ???? 2021. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610000809.

Stamoulis:2013:SAI

[SF13] Kostantinos A. Stamoulis and Alan M. Friedlander. A seascape approach to investigating fish spillover across a marine protected area boundary in Hawai’i. Fisheries Research, 144(??):2–14, July 2013. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612002858.

Schaefer:2019:SVR

[SF19] Kurt M. Schaefer and Daniel W. Fuller. Spatiotemporal variability in the reproductive dynamics of Skipjack Tuna (Katsuwonus pelamis) in the eastern Pacific Ocean. Fisheries Research, 209(??):1–13, January 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618302376.

Schaefer:2011:MBH

[SFB11] Kurt M. Schaefer, Daniel W. Fuller, and Barbara A. Block. Movements, behavior, and habitat utilization of yellowfin tuna (Thunnus albacares) in the Pacific Ocean off Baja California, Mexico, determined from archival tag data analyses, including unscented Kalman filtering. Fisheries Research, 112(1–2):22–37, December 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361100275X.

Stocks:2017:CCC

[SFBV17] Allison P. Stocks, Sarah J. Foster, Nguyen K. Bat, and Amanda C. J. Vincent. Catch as catch can: Targeted and indiscriminate small-scale fishing of seahorses in Vietnam. Fisheries Research, 196(??):27–33, December 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617302047.

Schaefer:2015:MDM

[SFH^{+}15] Kurt Schaefer, Daniel Fuller, John Hampton, Sylvain Caillot, Bruno Leroy, and David Itano. Movements, dispersion, and mixing of bigeye tuna (Thunnus obesus) tagged and released in the equatorial Central Pacific Ocean, with conventional and archival tags. Fisheries Research, 161(??):336–355, January 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361400263X.

Staalhammar:2014:ELT

[SFL^{+}14] M. Stålhammar, T. Fränstam, J. Lindström, J. Höjesjö, R. Arlinghaus, and P. A. Nilsson. Effects of lure type, fish size and water temperature on hooking location and bleeding in northern pike (Esox lucius) angled in the Baltic Sea. Fisheries Research, 157(??):164–169, September 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614001283.

Shertzer:2017:IGM

[SFPB17] Kyle W. Shertzer, John Fieberg, Jennifer C. Potts, and Michael L. Burton. Identifying growth morphs from mixtures of size-at-age data. Fisheries Research, 185(??):83–89, January 2017. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783616303253.

Shephard:2013:MRC

[SFR^{+}13] Samuel Shephard, Tak Fung, Axel G. Rossberg, Keith D. Farnsworth, David G. Reid, Simon P. R. Greenstreet, and Steve Warnes. Modelling recovery of Celtic Sea demersal fish community size-structure. Fisheries Research, 140(??):91–95, February 2013. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612003542.

Santos:2015:ATI

[SG15] Jorge Santos and Johan C. Groeneveld. Accounting for tag-induced growth retardation in spiny lobsters using censoring and modelling approaches. Fisheries Research, 161(??):166–173, January 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614002227.

Sethi:2017:AAJ

[SGA17] Suresh Andrew Sethi, Jonathon Gerken, and Joshua Ashline. Accurate aging of juvenile salmonids using fork lengths. Fisheries Research, 185(??):161–168, January 2017. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783616303058.

Sass:2018:ECR

[SGA^{+}18] Greg G. Sass, Jereme W. Gaeta, Micheal S. Allen, Cory D. Suski, and Stephanie L. Shaw. Effects of catch-and-release angling on a largemouth bass (Micropterus salmoides) population in a north temperate lake, 2001–2005. Fisheries Research, 204(??):95–102, August 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361830047X.

Stephenson:2010:ABN

[SGB^{+}10] John R. Stephenson, Andrew J. Gingerich, Richard S. Brown, Brett D. Pflugrath, Zhiqun Deng, Thomas J. Carlson, Mike J. Langeslay, Martin L. Ahmann, Robert L. Johnson, and Adam G. Seaburg. Assessing barotrauma in neutrally and negatively buoyant juvenile salmonids exposed to simulated hydro-turbine passage using a mobile aquatic barotrauma laboratory. Fisheries Research, 106(3):271–278, December 2010. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783610001839.

Steele:2011:CSE

[SGC11] John H. Steele, Dian J. Gifford, and Jeremy S. Collie. Comparing species and ecosystem-based estimates of fisheries yields. Fisheries Research, 111(3):139–144, November 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783611002682.

Speirs:2010:LSP

[SGGH10] D. C. Speirs, E. J. Guirey, W. S. C. Gurney, and M. R. Heath. A length-structured partial ecosystem model for cod in the North Sea. Fisheries Research, 106(3):474–494, December 2010. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783610002432.

Sequeira:2010:MBT

[SGN^{+}10] Vera Sequeira, Leonel Serrano Gordo, Ana Neves, Rafaela B. Paiva, Henrique N. Cabral, and Joana F. Marques. Macroparasites as biological tags for stock identification of the bluemouth, Helicolenus dactylopterus (Delaroche, 1809) in Portuguese waters. Fisheries Research, 106(3):321–328, December 2010. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783610002079.

Sydeman:2018:FHB

[SGRS^{+}18] William J. Sydeman, Marisol García-Reyes, Amber I. Szoboszlai, Sarah Ann Thompson, and Julie A. Thayer. Forecasting herring biomass using environmental and population parameters. Fisheries Research, 205(??):141–148, September 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783618301267.

Salvatore:2019:EFP

[SGSFL19] Pablo Di Salvatore, María Gowland-Sainz, Olga Florentin, and Gustavo A. Lovrich. Effects of fishery practices on fecundity of two lithodid crab species of commercial interest in Southern South America. Fisheries Research, 211(??):275– 281, March 2019. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783618303382.

Szaloky:2014:AEB

[SGT^{+}14] Z. Szalóky, Á.I. György, B. Tóth, A. Sevcsik, A. Specziár, B. Csányi, J. Szekeres, and T. Erős. Application of an electrified benthic frame trawl for sampling fish in a very large European river (the Danube River) — is offshore monitoring necessary? Fisheries Research, 151(??):12–19, March 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783613002981.

Simonovic:2018:JRC

[SGT^{+}18] Predrag Simonović, Goran Grubić, Ana Tošić, Jelena Čanak Atlagić, Dubravka Škraba Jurlina, and Vera Nikolić. Justification for retention of the catch-and-release in the wild brown trout Salmo cf. trutta fishery. Fisheries Research, 200(??):17–24, April 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783617303521.

Serena:2016:RBM

[SGW16] M. Serena, T. R. Grant, and G. A. Williams. Reducing bycatch mortality in crustacean traps: Effect of trap design on platypus and yabby retention rates. Fisheries Research, 175(??):43–50, March 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615301338.

Stephen:2010:CCC

[SH10] Jessica A. Stephen and Patrick J. Harris. Commercial catch composition with discard and immediate release mortality proportions off the southeastern coast of the United States.

Fisheries Research, 103(1–3):18–24, April 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610000111.

Siceloff:2013:FST

[SH13a] Laughlin Siceloff and W. Huntting Howell. Fine-scale temporal and spatial distributions of Atlantic cod (Gadus morhua) on a western Gulf of Maine spawning ground. Fisheries Research, 141(??):31–43, April 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612001294.

Su:2013:ALH

[SH13b] Zhenming Su and Ji X. He. Analysis of Lake Huron recreational fisheries data using models dealing with excessive zeros. Fisheries Research, 148(??):81–89, November 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783613002051.

Sepulveda:2015:PRS

[SHA^{+}15] C. A. Sepulveda, C. Heberer, S. A. Aalbers, N. Spear, M. Kinney, D. Bernal, and S. Kohin. Post-release survivorship studies on common thresher sharks (Alopias vulpinus) captured in the southern California recreational fishery. Fisheries Research, 161(??):102–108, January 2015. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783614002112.

Shih:2014:FAA

[SHC14] Chia-Lung Shih, Chien-Chung Hsu, and Chiee-Young Chen. First attempt to age yellowfin tuna, Thunnus albacares, in the Indian Ocean, based on sectioned otoliths. Fisheries Research, 149(??):19–23, January 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613002191.

Sistiaga:2015:ELS

[SHG^{+}15] Manu Sistiaga, Bent Herrmann, Eduardo Grimaldo, Roger B. Larsen, and Ivan Tatone. Effect of lifting the sweeps on bottom trawling catch efficiency: a study based on the Northeast Arctic cod (Gadus morhua) trawl fishery. Fisheries Research, 167(??):164–173, July 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615000363.

Sistiaga:2016:ESB

[SHG^{+}16] Manu Sistiaga, Bent Herrmann, Eduardo Grimaldo, Roger B. Larsen, and Ivan Tatone. The effect of sweep bottom contact on the catch efficiency of haddock (Melanogrammus aeglefinus). Fisheries Research, 179(??):302–307, July 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783616300856.

Sistiaga:2021:ADS

[SHGL21] Manu Sistiaga, Bent Herrmann, Eduardo Grimaldo, and Roger B. Larsen. Assessment of dual selection in grid based selectivity systems. Fisheries Research, 105(3):187– 199, ???? 2021. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783610001098.

Stouten:2011:PIM

[SHGP11a] Hendrik Stouten, Aimé Heene, Xavier Gellynck, and Hans Polet. Policy instruments to meet fisheries management objectives in Belgian fisheries. Fisheries Research, 111(1–2):8–23, September 2011. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611002244.

Stouten:2011:SGB

[SHGP11b] Hendrik Stouten, Aimé Heene, Xavier Gellynck, and Hans Polet. Strategic groups in the Belgian fishing fleet. Fisheries Research, 108(1):121–132, February 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361000322X.

Stawitz:2019:HDG

[SHJ19] Christine C. Stawitz, Melissa A. Haltuch, and Kelli F. Johnson. How does growth misspecification affect management advice derived from an integrated fisheries stock assessment model? Fisheries Research, 213(??):12–21, May 2019. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619300049.

Spencer:2012:SMT

[SHM12] Paul D. Spencer, Dana H. Hanselman, and Denise R. McKelvey. Simulation modeling of a trawl-acoustic survey design for patchily distributed species. Fisheries Research, 125–126(??):289–299, August 2012. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361200118X.

Santos:2016:RFB

[SHM^{+}16] Juan Santos, Bent Herrmann, Bernd Mieske, Daniel Stepputtis, Uwe Krumme, and Hans Nilsson. Reducing flatfish bycatch in roundfish fisheries. Fisheries Research, 184(??):64–73, December 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615300710.

Sato:2015:ACC

[SHPSK15] Mei Sato, John K. Horne, Sandra L. Parker-Stetter, and Julie E. Keister. Acoustic classification of coexisting taxa in a coastal ecosystem. Fisheries Research, 172(??):130–136, December 2015. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783615001964.

Southwick:2018:EMR

[SHR^{+}18a] Rob Southwick, John C. Holdsworth, Trish Rea, Lisa Bragg, and Thomas Allen. Estimating marine recreational fishing’s economic contributions in New Zealand. Fisheries Research, 208(??):116–123, December 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618301863.

Sweetman:2018:MDV

[SHR18b] P. C. Sweetman, J. A. Haddy, and S. G. Robertson. Multidecadal variation in cohort specific sex ratios and otolith increment growth characteristics of juvenile blue grenadier (Macruronus novaezelandiae). Fisheries Research, 201(??): 79–87, May 2018. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578361830016X.

Stewart:2013:CSA

[SHT^{+}13] Ian J. Stewart, Allan C. Hicks, Ian G. Taylor, James T. Thorson, Chantell Wetzel, and Sven Kupschus. A comparison of stock assessment uncertainty estimates using maximum likelihood and Bayesian methods implemented with the same model framework. Fisheries Research, 142(??):37– 46, May 2013. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783612002081.

Spirkovski:2019:GNR

[SIBR^{+}19] Z. Spirkovski, D. Ilik-Boeva, D. Ritterbusch, R. Peveling, and M. Pietrock. Ghost net removal in ancient Lake Ohrid: a pilot study. Fisheries Research, 211(??):46–50, March 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783618302923.

Simon:2015:AGE

[Sim15] Janek Simon. Age and growth of European eels (Anguilla anguilla) in the Elbe River system in Germany. Fisheries Research, 164(??):278–285, April 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614003610.

Skov:2017:YPD

[SJA17] Christian Skov, Teunis Jansen, and Robert Arlinghaus. 62 years of population dynamics of European perch (Perca fluviatilis) in a mesotrophic lake tracked using angler diaries: the role of commercial fishing, predation and temperature. Fisheries Research, 195(??):71–79, November 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617301716.

Sen:2011:SSA

[SJJ^{+}11] Swatipriyanka Sen, Shrinivas Jahageerdar, A. K. Jaiswar, S. K. Chakraborty, A. M. Sajina, and G. R. Dash. Stock structure analysis of Decapterus russelli (Ruppell, 1830) from east and west coast of India using truss network analysis. Fisheries Research, 112(1–2):38–43, December 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783611002773.

Sobrino:2011:DCP

[SJR^{+}11] I. Sobrino, A. Juarez, J. Rey, Z. Romero, and J. Baro. Description of the clay pot fishery in the Gulf of Cadiz (SW Spain) for Octopus vulgaris: Selectivity and exploitation pattern. Fisheries Research, 108(2–3):283–290, March 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783610003474.

Siikavuopio:2017:EPH

[SJS^{+}17] Sten Ivar Siikavuopio, Philip James, Even Stenberg, Tor Evensen, and Bjørn-Steinar Sæther. Evaluation of protein hydrolysate of by-product from the fish industry for inclusion in bait for longline and pot fisheries of Atlantic cod. Fisheries Research, 188(??):121–124, April 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361630409X.

Schuhbauer:2013:ARF

[SK13] Anna Schuhbauer and Volker Koch. Assessment of recreational fishery in the Galapagos Marine Reserve: Failures and opportunities. Fisheries Research, 144(??):103–110, July 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783613000155.

Stottrup:2018:HGS

[SKB^{+}18] Josianne G. Støttrup, Alexandros Kokkalis, Elliot John Brown, Jeppe Olsen, Stine Kærulf Andersen, and Eva Maria Pedersen. Harvesting geo-spatial data on coastal fish assemblages through coordinated citizen science. Fisheries Research, 208(??):86–96, December 2018. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0165783618302133. See corrigendum [SKB^{+}19].

Stottrup:2019:CHG

[SKB^{+}19] Josianne G. Støttrup, Alexandros Kokkalis, Elliot John Brown, Jeppe Olsen, Stine Kærulf Andersen, and Eva Maria Pedersen. Corrigendum to “Harvesting geo-spatial data on coastal fish assemblages through coordinated citizen science” [Fish. Res. 208 (2018) 86–96]. Fisheries Research, 212(??):208, April 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783618303813. See [SKB^{+}18].

Svalheim:2017:EES

[SKDO^{+}17] Ragnhild Aven Svalheim, Anders Karlsson-Drangsholt, Stein Harris Olsen, Helge Kreutzer Johnsen, and Øyvind Aas-Hansen. Effects of exhaustive swimming and subsequent recuperation on flesh quality in unstressed Atlantic cod (Gadus morhua). Fisheries Research, 193(??):158–163, September 2017. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783617301133.

Stewart:2010:TTW

[SKF^{+}10] Ian J. Stewart, Aimee A. Keller, Erica L. Fruh, Victor H. Simon, and Beth H. Horness. Throwing in the towel: When do adverse conditions dictate a weather day during a bottom trawl survey? Fisheries Research, 102(1–2):130–140, February 2010. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783609003026.

Samy-Kamal:2015:ESC

[SKFL15] Mohamed Samy-Kamal, Aitor Forcada, and José Luis Sánchez Lizaso. Effects of seasonal closures in a multi-specific fishery. Fisheries Research, 172(??):303–317, December 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783615300369.

Savina:2017:EFS

[SKFM17] Esther Savina, Ludvig Ahm Krag, Rikke Petri Frandsen, and Niels Madsen. Effect of fisher’s soak tactic on catch pattern in the Danish gillnet plaice fishery. Fisheries Research, 196(??):56–65, December 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617302230.

Somerville:2014:MME

[SKH14] G. J. Somerville, M. Krkosek, and C. D. Hepburn. A matrix model and elasticity analysis for New Zealand’s blackfoot pāua Haliotis iris. Fisheries Research, 151(??):158–168, March 2014. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783613002865.

Skjaeraasen:2013:SES

[SKK^{+}13] Jon Egil Skjæraasen, Knut Korsbrekke, Olav Sigurd Kjesbu, Merete Fonn, Trygve Nilsen, and Richard D. M. Nash. Size-, energy- and stage-dependent fecundity and the occurrence of atresia in the Northeast Arctic haddock Melanogrammus aeglefinus. Fisheries Research, 138(??):120–127, February 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783612001312.

Sullivan:2012:SLS

[SKR^{+}12] B. J. Sullivan, P. Kibel, G. Robertson, B. Kibel, M. Goren, S. G. Candy, and B. Wienecke. Safe leads for safe heads: safer line weights for pelagic longline fisheries. Fisheries Research, 134–136(??):125–132, December 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612002524.

Shimose:2018:IFC

[SKT18] Tamaki Shimose, Minoru Kanaiwa, and Toshiyuki Tanabe. Influence of the fat content on the auction price of Pacific bluefin tuna Thunnus orientalis at Japanese fish markets.

Fisheries Research, 204(??):224–230, August 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618300754.

Sakamoto:2017:TDO

[SKY^{+}17] Tatsuya Sakamoto, Kosei Komatsu, Michio Yoneda, Toyoho Ishimura, Tomihiko Higuchi, Kotaro Shirai, Yasuhiro Kamimura, Chikako Watanabe, and Atsushi Kawabata. Temperature dependence of δ ^{18}O in otolith of juvenile Japanese sardine: Laboratory rearing experiment with micro-scale analysis. Fisheries Research, 194(??):55–59, October 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578361730125X.

Sala:2010:EMC

[SL10] Antonello Sala and Alessandro Lucchetti. The effect of mesh configuration and codend circumference on selectivity in the Mediterranean trawl Nephrops fishery. Fisheries Research, 103(1–3):63–72, April 2010. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610000500.

Sala:2011:EMS

[SL11a] Antonello Sala and Alessandro Lucchetti. Effect of mesh size and codend circumference on selectivity in the Mediterranean demersal trawl fisheries. Fisheries Research, 110(2):252–258, July 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783611001615.

Shainee:2011:CPF

[SL11b] M. Shainee and B. J. Leira. On the cause of premature FAD loss in the Maldives. Fisheries Research, 109(1):42– 53, April 2011. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783611000348.

Suuronen:2012:RSD

[SL12] Petri Suuronen and Esa Lehtonen. The role of salmonids in the diet of grey and ringed seals in the Bothnian Bay, northern Baltic Sea. Fisheries Research, 125–126(??):283–288, August 2012. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783612001269.

Staalhammar:2012:ICR

[SLB^{+}12] Martin Stålhammar, Rasmus Linderfalk, Christer Brönmark, Robert Arlinghaus, and P. Anders Nilsson. The impact of catch-and-release on the foraging behaviour of pike (Esox lucius) when released alone or into groups. Fisheries Research, 125–126(??):51–56, August 2012. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783612000483.

Surette:2015:UIF

[SLCL15] Tobie Surette, Claude H. LeBlanc, Ross R. Claytor, and Christophe Loots. Using inshore fishery acoustic data on Atlantic herring (Clupea harengus) spawning aggregations to derive annual stock abundance indices. Fisheries Research, 164(??):266–277, April 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361400366X.

Sharma:2014:IIL

[SLH^{+}14] Rishi Sharma, Adam Langley, Miguel Herrera, James Geehan, and Saang-Yoon Hyun. Investigating the influence of length–frequency data on the stock assessment of Indian Ocean bigeye tuna. Fisheries Research, 158(??):50–62, October 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783614000204.

Siskey:2016:PSC

[SLL^{+}16] M. R. Siskey, V. Lyubchich, D. Liang, P. M. Piccoli, and D. H. Secor. Periodicity of strontium: Calcium across annuli further validates otolith-ageing for Atlantic bluefin tuna (Thunnus thynnus). Fisheries Research, 177(??):13–17, May 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783616300042.

Sala:2015:SMB

[SLP^{+}15] Antonello Sala, Alessandro Lucchetti, Anna Perdichizzi, Bent Herrmann, and Paola Rinelli. Is square-mesh better selective than larger mesh? A perspective on the management for Mediterranean trawl fisheries. Fisheries Research, 161(??):182–190, January 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614002252.

Sippel:2017:STM

[SLPT17] Tim Sippel, Hui Hua Lee, Kevin Piner, and Steven L. H. Teo. Searching for M: Is there more information about natural mortality in stock assessments than we realize? Fisheries Research, 192(??):135–140, August 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616304192.

Schvezov:2019:ETA

[SLT^{+}19] Natasha Schvezov, Gustavo A. Lovrich, Federico Tapella, Maria Gowland-Sainz, and M. Carolina Romero. Effect of the temperature of air exposure on the oxidative stress status of commercial male southern king crab Lithodes santolla. Fisheries Research, 212(??):188–195, April 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618303709.

Sarmiento-Lezcano:2018:ABL

[SLTPC^{+}18] Airam-Nauzet Sarmiento-Lezcano, Raül Triay-Portella, José Juan Castro, Uriel Rubio-Rodríguez, and José G. Pajuelo. Age-based life-history parameters of the mesopelagic fish Notoscopelus resplendens (Richardson, 1845) in the Central Eastern Atlantic. Fisheries Research, 204(??):412–423, August 2018. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783618300900.

Song:2015:DSP

[SLXZ15] Liming Song, Jie Li, Weiyun Xu, and Xinfeng Zhang. The dynamic simulation of the pelagic longline deployment. Fisheries Research, 167(??):280–292, July 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615000958.

Song:2018:CPP

[SLZG18] Na Song, Pengfei Li, Xiumei Zhang, and Tianxiang Gao. Changing phylogeographic pattern of Fenneropenaeus chinensis in the Yellow Sea and Bohai Sea inferred from microsatellite DNA: Implications for genetic management. Fisheries Research, 200(??):11–16, April 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617303387.

Stewart:2014:HRS

[SM14] Ian J. Stewart and Steven J. D. Martell. A historical review of selectivity approaches and retrospective patterns in the Pacific halibut stock assessment. Fisheries Research, 158(??):40–49, October 2014. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613002221.

Sahashi:2015:NRS

[SM15] Genki Sahashi and Kentaro Morita. Non-random sampling for reproductive status induces bias in probabilistic maturation reaction norm midpoints. Fisheries Research, 170(??):24–29, October 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783615001460.

Stewart:2017:IPE

[SM17] Ian J. Stewart and Cole C. Monnahan. Implications of process error in selectivity for approaches to weighting compositional data in fisheries stock assessments. Fisheries Research, 192(??):126–134, August 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616302077.

Sacco:2017:MSH

[SMB^{+}17] Francesco Sacco, Federico Marrone, Sabrina Lo Brutto, Amina Besbes, Ahmed Nfati, Mark Gatt, Samar Saber, Fabio Fiorentino, and Marco Arculeo. The Mediterranean Sea hosts endemic haplotypes and a distinct population of the dolphinfish Coryphaena hippurus Linnaeus, 1758 (Perciformes, Coryphaenidae). Fisheries Research, 186 (part 1)(??):151–158, February 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616302764.

Stoner:2012:MAQ

[SMBP^{+}12] Allan W. Stoner, Karl W. Mueller, Nancy J. Brown-Peterson, Martha H. Davis, and Catherine J. Booker. Maturation and age in queen conch (Strombus gigas): Urgent need for changes in harvest criteria. Fisheries Research, 131–133(??):76–84, November 2012. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612002238.

Santos:2013:ODR

[SMCR13] Ana Ribeiro Santos, Cóilín Minto, Paul Connolly, and Emer Rogan. Oocyte dynamics and reproductive strategy of Aphanopus carbo in the NE Atlantic — implications for fisheries management. Fisheries Research, 143(??):161–173, June 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783613000076.

Stratoudakis:2016:FCD

[SMD^{+}16] Yorgos Stratoudakis, Patrick McConney, John Duncan, Abdul Ghofar, Nancy Gitonga, Kolliyil S. Mohamed, Melita Samoilys, Keith Symington, and Luis Bourillon. Fisheries certification in the developing world: Locks and keys or square pegs in round holes? Fisheries Research, 182(??):39–49, October 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615300679.

Saber:2016:CBF

[SMdUK16] Sámar Saber, David Macías, Josetxu Ortiz de Urbina, and Olav Sigurd Kjesbu. Contrasting batch fecundity estimates of albacore (Thunnus alalunga), an indeterminate spawner, by different laboratory techniques. Fisheries Research, 176(??):76–85, April 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361530165X.

Shiffman:2017:FPR

[SMGH17] David S. Shiffman, Catherine Macdonald, Harry Y. Ganz, and Neil Hammerschlag. Fishing practices and representations of shark conservation issues among users of a land-based shark angling online forum. Fisheries Research, 196(??):13–26, December 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361730214X.

Smith:2014:AVL

[Smi14] Joanne Smith. Age validation of lemon sole (Microstomus kitt), using marginal increment analysis. Fisheries Research, 157(??):41–46, September 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614000939.

Somerton:2017:EUA

[SMI17] David A. Somerton, Robert A. McConnaughey, and Steven S. Intelmann. Evaluating the use of acoustic bottom typing to inform models of bottom trawl sampling efficiency. Fisheries Research, 185(??):14–16, January 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783616303228.

Schmalenbach:2011:MRS

[SMJB11] Isabel Schmalenbach, Folke Mehrtens, Michael Janke, and Friedrich Buchholz. A mark-recapture study of hatchery-reared juvenile European lobsters, Homarus gammarus, released at the rocky island of Helgoland (German Bight, North Sea) from 2000 to 2009. Fisheries Research, 108(1):22–30, February 2011. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610003012.

Sato:2013:CEP

[SMK^{+}13] Noriyosi Sato, Hiroshi Minami, Nobuhiro Katsumata, Daisuke Ochi, and Kotaro Yokawa. Comparison of the effectiveness of paired and single tori lines for preventing bait attacks by seabirds and their bycatch in pelagic longline fisheries. Fisheries Research, 140(??):14–19, February 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783612003426.

Soares:2019:SSR

[SMNdC^{+}19] Júlia Benevenuti Soares, Cassiano Monteiro-Neto, Marcus Rodrigues da Costa, Raquel Rennó M. Martins, Francyne Carolina dos Santos Vieira, Magda Fernandes de Andrade-Tubino, Ana Luiza Bastos, and Rafael de Almeida Tubino. Size structure, reproduction, and growth of skipjack tuna (Katsuwonus pelamis) caught by the pole-and-line fleet in the southwest Atlantic. Fisheries Research, 212(??):136–145, April 2019. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783618303552.

Sandlund:2016:MGP

[SMØ16] Odd Terje Sandlund, Jon Museth, and Sondre Øistad. Migration, growth patterns, and diet of pike (Esox lucius) in a river reservoir and its inflowing river. Fisheries Research, 173(??):53–60, January 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615300564.

Sepulveda:2018:FAO

[SMO^{+}18a] Maritza Sepúlveda, Tamara Martínez, Doris Oliva, Pablo Couve, Guido Pavez, Claudia Navarro, Milan Stehlik, L. René Durán, and Guillermo Luna-Jorquera. Factors affecting the operational interaction between the South American sea lions and the artisan gillnet fishery in Chile. Fisheries Research, 201(??):147–152, May 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618300213.

Silva:2018:GDP

[SMO^{+}18b] Danillo Silva, Kely Martins, Joiciane Oliveira, Raimundo da Silva, Iracilda Sampaio, Horacio Schneider, and Grazielle Gomes. Genetic differentiation in populations of lane snapper (Lutjanus synagris — Lutjanidae) from Western Atlantic as revealed by multilocus analysis. Fisheries Research, 198(??):138–149, February 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617302734.

Smith:2018:ESS

[SMP18] Geoffrey H. Smith, Debra J. Murie, and Daryl C. Parkyn. Effects of sex-specific fishing mortality on sex ratio and population dynamics of Gulf of Mexico greater amberjack. Fisheries Research, 208(??):219–228, December 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618302108.

Sys:2017:TRF

[SMPB17] Klaas Sys, Jef Van Meensel, Hans Polet, and Jeroen Buysse. A temporal race-for-fish: the interplay between local hotspots of flatfish and exploitation competition between beam trawlers after a seasonal spawning closure. Fisheries Research, 193(??):21–32, September 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617300851.

Sasikumar:2015:CAF

[SMRS15] Geetha Sasikumar, K. S. Mohamed, Prathibha Rohit, and G. Sampathkumar. Can an aggregation-fishery be responsible for recruitment overfishing? A case study on cuttlefish stock associated with moored fish aggregation devices (FADs). Fisheries Research, 172(??):148–156, December 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783615300096.

Smith:2017:EDT

[SMS17] Nicholas J. Smith, Patricia L. McCall, and Trent M. Sutton. Effects of different tagging protocols on survival, growth, and tag retention in juvenile least cisco Coregonus sardinella. Fisheries Research, 187(??):68–72, March 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783616303733.

Shao:2019:TSC

[SMS^{+}19] Huamei Shao, Kenji Minami, Hokuto Shirakawa, Yohei Kawauchi, Ryuichi Matsukura, Makoto Tomiyasu, and Kazushi Miyashita. Target strength of a common kelp species, Saccharina japonica, measured using a quantitative echosounder in an indoor seawater tank. Fisheries Research, 214(??):110–116, June 2019. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619300098.

Savina:2019:EST

[SNK19] Esther Savina, Thomas Noack, and Junita D. Karlsen. Effects of season, target species and codend design on the survival of undersized plaice (Pleuronectes platessa) discarded in the bottom otter trawl mixed fisheries in Skagerrak. Fisheries Research, 219(??):Article 105311, November 2019. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619301584.

Sossoukpe:2013:GMP

[SNOD^{+}13] E. Sossoukpe, F. K. E. Nunoo, P. K. Ofori-Danson, E. D. Fiogbe, and H. R. Dankwa. Growth and mortality parameters of P. senegalensis and P. typus (Sciaenidae) in nearshore waters of Benin (West Africa) and their implications for management and conservation. Fisheries Research, 137(??):71–80, January 2013. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612002688.

Sequeira:2012:LHP

[SNP^{+}12] Vera Sequeira, Ana Neves, Rafaela Barros Paiva, João Pereira de Lima, Ana Rita Vieira, and Leonel Serrano Gordo. Life history parameters as possible discriminators of bluemouth Helicolenus dactylopterus (Delaroche, 1809) populations in Portuguese waters. Fisheries Research, 125–126(??):69–76, August 2012. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783612000896.

Satterthwaite:2018:IOD

[SO18] William H. Satterthwaite and Michael R. O’Farrell. Inferred ocean distributions of genetically similar Chinook salmon stocks compared across run timing and river/hatchery of origin. Fisheries Research, 199(??):171–176, March 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783617303144.

Samoilys:2017:AFK

[SOMO17] Melita A. Samoilys, Kennedy Osuka, George W. Maina, and David O. Obura. Artisanal fisheries on Kenya’s coral reefs: Decadal trends reveal management needs. Fisheries Research, 186 (part 1)(??):177–191, February 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616302375.

Song:2018:BCK

[Son18] Kyung-Jun Song. Bycatch of cetaceans in Korea fisheries in the East Sea. Fisheries Research, 197(??):7–9, January 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783617302709.

Shabangu:2014:MAA

[SOY14] Fannie W. Shabangu, Egil Ona, and Dawit Yemane. Measurements of acoustic attenuation at 38 kHz by windinduced air bubbles with suggested correction factors for hull-mounted transducers. Fisheries Research, 151(??):47– 56, March 2014. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783613003020.

Schroeder:2010:SIS

[SP10a] R. Schroeder and J. A. A. Perez. The study of intraspecific growth variability of Illex argentinus (Cephalopoda: Teuthida) in Brazilian waters as reconstructed from the gladius microstructure. Fisheries Research, 106(2):163–170, November 2010. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783610001116.

Schwarz:2010:GMI

[SP10b] Richard Schwarz and José Angel Alvarez Perez. Growth model identification of short-finned squid Illex argentinus (Cephalopoda: Ommastrephidae) off southern Brazil using statoliths. Fisheries Research, 106(2):177–184, November 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783610001384.

Szuwalski:2012:IRP

[SP12] Cody Szuwalski and André E. Punt. Identifying research priorities for management under uncertainty: the estimation ability of the stock assessment method used for eastern Bering Sea snow crab (Chionoecetes opilio). Fisheries Research, 134–136(??):82–94, December 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612002536.

Sinclair:2015:RLR

[SP15a] Mike Sinclair and Mike Power. The role of “ larval retention” in life-cycle closure of Atlantic herring (Clupea harengus) populations. Fisheries Research, 172(??):401–414, December 2015. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783615300357.

Szuwalski:2015:CAA

[SP15b] Cody S. Szuwalski and André E. Punt. Can an aggregate assessment reflect the dynamics of a spatially structured stock? Snow crab in the eastern Bering Sea as a case study. Fisheries Research, 164(??):135–142, April 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614003142.

Sharma:2019:RTE

[SPB^{+}19] Rishi Sharma, Clay E. Porch, Elizabeth A. Babcock, Mark N. Maunder, and André E. Punt. Recruitment: Theory, estimation, and application in fishery stock assessment models. Fisheries Research, 217(??):1–4, September 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619300827.

Sulikowski:2013:ONS

[SPC^{+}13] James A. Sulikowski, Bianca K. Prohaska, Amy E. Carlson, Angela M. Cicia, Christopher T. Brown, and Alexia C. Morgan. Observations of neonate spiny dogfish, Squalus acanthias, in Southern New England: a first account of a potential pupping ground in the Northwestern Atlantic. Fisheries Research, 137(??):59–62, January 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612002664.

Spurgeon:2018:MOC

[SPH18] Jonathan J. Spurgeon, Mark A. Pegg, and Norman M. Halden. Mixed-origins of channel catfish in a large-river tributary. Fisheries Research, 198(??):195–202, February 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783617302436.

Stari:2010:SAL

[SPM^{+}10] Traiani Stari, Katharine F. Preedy, Eddie McKenzie, William S. C. Gurney, Michael R. Heath, Philip A. Kunzlik, and Douglas C. Speirs. Smooth age length keys: Observations and implications for data collection on North Sea haddock. Fisheries Research, 105(1):2–12, June 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610000512.

Sembiring:2015:DBR

[SPM^{+}15a] Andrianus Sembiring, Ni Putu Dian Pertiwi, Angka Mahardini, Rizki Wulandari, Eka Maya Kurniasih, Andri Wahyu Kuncoro, N. K. Dita Cahyani, Aji Wahyu Anggoro, Maria Ulfa, Hawis Madduppa, Kent E. Carpenter, Paul H. Barber, and Gusti Ngurah Mahardika. DNA barcoding reveals targeted fisheries for endangered sharks in Indonesia. Fisheries Research, 164(??):130–134, April 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614003282.

Shrives:2015:WBU

[SPM15b] Jonathan Peter Shrives, Sophie Emma Pickup, and Greg Michael Morel. Whelk (Buccinum undatum L.) stocks around the Island of Jersey, Channel Islands: Reassessment and implications for sustainable management. Fisheries Research, 167(??):236–242, July 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783615000922.

Sande:2019:TAE

[SPM^{+}19] Hege Anett Sande, Anders Persson, Sigbjørn Mehl, Jennifer A. Devine, Thassya C. dos Santos Schmidt, Ørjan Karlsen, Jane Aanestad Godiksen, and Olav Sigurd Kjesbu. Temperature and age effects on latitudinal growth dynamics of the commercially valuable gadoid Northeast Arctic saithe (Pollachius virens). Fisheries Research, 213(??):94– 104, May 2019. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783619300050.

Scordino:2017:ECD

[SPMS17] Jonathan Joseph Scordino, Joseph Robert Petersen, Joshua Levi Monette, and Joe Scordino. Evaluation of the čibu· d, traditional halibut hook of the Makah Tribe, for reducing catch of non-target species in recreational Pacific halibut fisheries. Fisheries Research, 185(??):17–25, January 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361630323X.

Stefansson:2019:IEA

[SPR^{+}19] Gunnar Stefansson, André E. Punt, Javier Ruiz, Ingrid van Putten, Sveinn Agnarsson, and Anna Kristín Daníelsdóttir. Implementing the ecosystem approach to fisheries management. Fisheries Research, 216(??):174–176, August 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619301110.

Shikon:2019:SVN

[SPS^{+}19] Violaine Shikon, Pierre Pepin, David C. Schneider, Martin Castonguay, and Dominique Robert. Spatiotemporal variability in Newfoundland capelin (Mallotus villosus) larval abundance and growth: Implications for recruitment. Fisheries Research, 218(??):237–245, October 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619301122.

Santos:2019:GMT

[SPV19] Sérgio Ricardo Santos, Leila Maria Pessôa, and Marcelo Vianna. Geometric morphometrics as a tool to identify species in multispecific flatfish landings in the tropical Southwestern Atlantic. Fisheries Research, 213(??):190–195, May 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783619300153.

Stepanuk:2018:SPO

[SRB^{+}18] Julia E. F. Stepanuk, Andrew J. Read, Robin W. Baird, Daniel L. Webster, and Lesley H. Thorne. Spatiotemporal patterns of overlap between short-finned pilot whales and the U.S. pelagic longline fishery in the Mid-Atlantic Bight: an assessment to inform the management of fisheries bycatch. Fisheries Research, 208(??):309–320, December 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578361830198X.

Santis:2019:IAS

[SRB19] Vanessa De Santis, Catherine Gutmann Roberts, and J. Robert Britton. Influences of angler subsidies on the trophic ecology of European barbel Barbus barbus. Fisheries Research, 214(??):35–44, June 2019. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619300360.

Sushchik:2017:EST

[SRG17] Nadezhda N. Sushchik, Anastasia E. Rudchenko, and Michail I. Gladyshev. Effect of season and trophic level on fatty acid composition and content of four commercial fish species from Krasnoyarsk Reservoir (Siberia, Russia). Fisheries Research, 187(??):178–187, March 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616304015.

Siccha-Ramirez:2018:SIV

[SRMP^{+}18] Zoila Raquel Siccha-Ramirez, Francesco Maroso, Belén G. Pardo, Carlos Fernández, Paulino Martínez, and Claudio Oliveira. SNP identification and validation on genomic DNA for studying genetic diversity in Thunnus albacares and Scomberomorus brasiliensis by combining RADseq and long read high throughput sequencing. Fisheries Research, 198(??):189–194, February 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617302448.

Saldana-Ruiz:2017:HRG

[SRSNC17] Luz Erandi Saldaña-Ruiz, Oscar Sosa-Nishizaki, and Daniel Cartamil. Historical reconstruction of Gulf of California shark fishery landings and species composition, 1939– 2014, in a data-poor fishery context. Fisheries Research, 195(??):116–129, November 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617301947.

Saborido-Rey:2013:FRF

[SRT13] Fran Saborido-Rey and Edward A. Trippel. Fish reproduction and fisheries. Fisheries Research, 138(??):1–4, February 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783612003360.

Souto:2018:BFW

[SSB^{+}18] Marina Do Souto, Mariela L. Spinelli, Daniel R. Brown, Marcelo Pájaro, Marina V. Diaz, and Fabiana L. Capitanio. Benefits of frontal waters for the growth of Engraulis anchoita larvae: the influence of food availability. Fisheries Research, 204(??):181–188, August 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618300559.

Stein:2012:IHS

[SSC^{+}12] Jeffrey A. Stein, Aaron D. Shultz, Steven J. Cooke, Andy J. Danylchuk, Kit Hayward, and Cory D. Suski. The influence of hook size, type, and location on hook retention and survival of angled bonefish (Albula vulpes). Fisheries Research, 113(1):147–152, January 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611003286.

Soeth:2019:SSA

[SSD^{+}19] Marcelo Soeth, Henry Louis Spach, Felippe Alexandre Daros, Johnatas Adelir-Alves, Ana Claudia Oliveria de Almeida, and Alberto Teodorico Correia. Stock structure of Atlantic spadefish Chaetodipterus faber from Southwest Atlantic Ocean inferred from otolith elemental and shape signatures. Fisheries Research, 211(??):81–90, March 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783618303035.

Sacristan:2019:ESS

[SSFGL19] Hernán Javier Sacristán, Pablo Di Salvatore, Analía Verónica Fernández-Gimenez, and Gustavo Alejandro Lovrich. Effects of starvation and stocking density on the physiology of the male of the southern king crab Lithodes santolla. Fisheries Research, 218(??):83–93, October 2019. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783619301225.

Sukumaran:2017:GPS

[SSG17] Sandhya Sukumaran, Wilson Sebastian, and A. Gopalakrishnan. Genetic population structure and historic demography of Indian mackerel, Rastrelliger kanagurta from Indian peninsular waters. Fisheries Research, 191(??):1–9, July 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783617300553.

Silva:2014:DSI

[SSH14] Luis Silva, Ignacio Sobrino, and Eva Hernández. Deep-sea incirrate octopus Velodona togata (Chun, 1915) in waters off Mozambique (SW Indian Ocean): Reproductive biology, distribution, and abundance. Fisheries Research, 152(??):98– 105, April 2014. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578361300177X.

Stepputtis:2016:BHS

[SSHM16] Daniel Stepputtis, Juan Santos, Bent Herrmann, and Bernd Mieske. Broadening the horizon of size selectivity in trawl gears. Fisheries Research, 184(??):18–25, December 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578361530076X.

Silva:2013:SPB

[SSJ^{+}13] Filipa F. G. Silva, Aril Slotte, Arne Johannessen, James Kennedy, and Olav Sigurd Kjesbu. Strategies for partition between body growth and reproductive investment in migratory and stationary populations of spring-spawning Atlantic herring (Clupea harengus L.). Fisheries Research, 138(??):71–79, February 2013. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612002196.

Su:2015:EIS

[SSP^{+}15] Nan-Jay Su, Chi-Lu Sun, André E. Punt, Su-Zan Yeh, and Gerard DiNardo. Environmental influences on seasonal movement patterns and regional fidelity of striped marlin Kajikia audax in the Pacific Ocean. Fisheries Research, 166(??):59–66, June 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783614002434.

Santiago:2018:AMA

[SSR18] Jose L. Santiago and Juan C. Surís-Regueiro. An applied method for assessing socioeconomic impacts of European fisheries quota-based management. Fisheries Research, 206(??):150–162, October 2018. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618301462.

Somarakis:2012:HVD

[SSS^{+}12] Stylianos Somarakis, Eudoxia Schismenou, Apostolos Siapatis, Marianna Giannoulaki, Argyris Kallianiotis, and Athanassios Machias. High variability in the daily egg production method parameters of an eastern Mediterranean anchovy stock: Influence of environmental factors, fish condition and population density. Fisheries Research, 117–118(??):12–21, April 2012. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610003127.

Sethi:2013:BIT

[ST13] Suresh Andrew Sethi and Theresa L. Tanner. Bayesian implementation of a time stratified Lincoln–Petersen estimator for salmon abundance in the upper Matanuska River, Alaska, USA. Fisheries Research, 145(??):90–99, August 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783613000374.

Stamoulis:2016:RNS

[ST16] Antonios Stamoulis and Els Torreele. The response of the North Sea demersal fish community to changing fishing pressure as seen through the prism of the large fish indicator. Fisheries Research, 181(??):222–233, September 2016. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783616301369.

Stange:2016:BKB

[Sta16] Kari Stange. Building a knowledge base for management of a new fishery: Boarfish (Capros aper) in the Northeast Atlantic. Fisheries Research, 174(??):94–102, February 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783615300692.

Sydeman:2017:BPA

[STAN^{+}17] William J. Sydeman, Sarah Ann Thompson, Tycho Anker-Nilssen, Mayumi Arimitsu, Ashley Bennison, Sophie Bertrand, Philipp Boersch-Supan, Charlotte Boyd, Nicole C. Bransome, Robert J. M. Crawford, Francis Daunt, Robert W. Furness, Dimas Gianuca, Amanda Gladics, Laura Koehn, Jennifer W. Lang, Elizabeth Logerwell, Taryn L. Morris, Elizabeth M. Phillips, Jennifer Provencher, André E. Punt, Claire Saraux, Lynne Shannon, Richard B. Sherley, Alejandro Simeone, Ross M. Wanless, Sarah Wanless, and Stephani Zador. Best practices for assessing forage fish fisheries-seabird resource competition. Fisheries Research, 194(??):209–221, October 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783617301455.

Specziar:2012:REE

[STCE12] A. Specziár, P. Takács, I. Czeglédi, and T. Erős. The role of the electrofishing equipment type and the operator in assessing fish assemblages in a non-wadeable lowland river. Fisheries Research, 125–126(??):99–107, August 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361200094X.

Stewart:2011:EAC

[Ste11] John Stewart. Evidence of age-class truncation in some exploited marine fish populations in New South Wales, Australia. Fisheries Research, 108(1):209–213, February 2011.

CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783610003024.

Stevenson:2018:DRS

[Ste18] Duane E. Stevenson. Documenting the reliability of species identifications in the North Pacific Observer Program. Fisheries Research, 201(??):26–31, May 2018. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618300043.

Skuvadal:2011:EBW

[STJ11] Fródi B. Skúvadal, Bjarti Thomen, and Jan Arge Jacobsen. Escape of blue whiting (Micromesistius poutassou) and herring (Clupea harengus) from a pelagic survey trawl. Fisheries Research, 111(1–2):65–73, September 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611002293.

Stoner:2012:EVP

[Sto12] Allan W. Stoner. Evaluating vitality and predicting mortality in spot prawn, Pandalus platyceros, using reflex behaviors. Fisheries Research, 119–120(??):108–114, May 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783611003924.

Sanchez:2016:PGJ

[STY^{+}16] Gustavo Sanchez, Satoshi Tomano, Carmen Yamashiro, Ricardo Fujita, Toshie Wakabayashi, Mitsuo Sakai, and Tetsuya Umino. Population genetics of the jumbo squid Dosidicus gigas (Cephalopoda: Ommastrephidae) in the northern Humboldt Current system based on mitochondrial and microsatellite DNA markers. Fisheries Research, 175(??):1–9, March 2016. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783615301284.

Stemle:2016:DPI

[SUR16] Adam Stemle, Hirotsugu Uchida, and Cathy A. Roheim. Have dockside prices improved after MSC certification? Analysis of multiple fisheries. Fisheries Research, 182(??):116–123, October 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361530031X.

Svedang:2014:CFA

[Sve14] Henrik Svedäng. Comments to Froese et al. (2008): Size matters: Ne quid nimis. Fisheries Research, 149(??):74–75, January 2014. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783613002038. See [FSPWG08,

FWG15].

Svedang:2015:SSH

[Sve15] Henrik Svedäng. On size selectivity and L_{opt} as a harvest strategy: Reply to Froese et al. (2014). Fisheries Research, 164(??):331–332, April 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614003646. See [FSPWG08, Sve14].

Sorell:2017:DCR

[SVG^{+}17] Joan Miquel Sorell, José Luis Varela, Nicolas Goñi, David Macías, Haritz Arrizabalaga, and Antonio Medina. Diet and consumption rate of Atlantic bluefin tuna (Thunnus thynnus) in the Strait of Gibraltar. Fisheries Research, 188(??):112–120, April 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783616304313.

Sequeira:2013:WSO

[SVN^{+}13] Vera Sequeira, Ana Rita Vieira, Ana Neves, Rafaela Barros Paiva, Adelino Canário, and Leonel Serrano Gordo. Whole or sectioned otoliths? Choosing the best method for aging bluemouth, Helicolenus dactyloterus (Delaroche, 1809). Fisheries Research, 147(??):235–239, October 2013. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783613001641.

Seijo:2016:BOA

[SVPC16] Juan Carlos Seijo, Raúl Villanueva-Poot, and Anthony Charles. Bioeconomics of ocean acidification effects on fisheries targeting calcifier species: a decision theory approach. Fisheries Research, 176(??):1–14, April 2016. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783615301521.

Sutela:2016:LFI

[SVR16] Tapio Sutela, Teppo Vehanen, and Martti Rask. A littoral fish index that responds to eutrophication in boreal lakes. Fisheries Research, 173(??):88–92, January 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615001861.

Sadighzadeh:2014:UOS

[SVV^{+}14] Z. Sadighzadeh, T. Valinassab, G. Vosugi, A. A. Motallebi, M. R. Fatemi, A. Lombarte, and V. M. Tuset. Use of otolith shape for stock identification of John’s snapper, Lutjanus johnii (Pisces: Lutjanidae), from the Persian Gulf and the Oman Sea. Fisheries Research, 155(??):59–63, July 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783614000708.

Sylvia:2019:UMR

[SW19] Andrea Sylvia and Michael J. Weber. Use of a markrecapture model to evaluate largemouth bass delayed tournament mortality. Fisheries Research, 219(??):Article 105335, November 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783619301900.

Sepulveda:2019:PRS

[SWA19] C. A. Sepulveda, M. Wang, and S. A. Aalbers. Post-release survivorship and movements of bigeye thresher sharks, Alopias superciliosus, following capture on deep-set buoy gear. Fisheries Research, 219(??):Article 105312, November 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619301596.

Seiler:2012:ASA

[SWB12] Jan Seiler, Alan Williams, and Neville Barrett. Assessing size, abundance and habitat preferences of the ocean perch Helicolenus percoides using a AUV-borne stereo camera system. Fisheries Research, 129–130(??):64–72, October 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783612001956.

Schueller:2014:PTA

[SWC14] Amy M. Schueller, Erik H. Williams, and Robin T. Cheshire. A proposed, tested, and applied adjustment to account for bias in growth parameter estimates due to selectivity. Fisheries Research, 158(??):26–39, October 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613002518.

Saul:2013:MSD

[SWD^{+}13] S. E. Saul, J. F. Walter, D. J. Die, D. F. Naar, and B. T. Donahue. Modeling the spatial distribution of commercially important reef fishes on the West Florida shelf. Fisheries Research, 143(??):12–20, June 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613000052.

Selden:2018:OSP

[SWG18] Rebecca L. Selden, Robert R. Warner, and Steven D. Gaines. Ontogenetic shifts in predator diet drive tradeoffs between fisheries yield and strength of predator–prey interactions. Fisheries Research, 205(??):11–20, September 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578361830095X.

Sexton:2017:CSP

[SWH17] Stuart C. Sexton, Tim M. Ward, and Charlie Huveneers. Characterising the spawning patterns of Jack mackerel (Trachurus declivis) off eastern Australia to optimise future survey design. Fisheries Research, 186 (part 1)(??):223–236, February 2017. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783616302843.

Samedy:2015:HRT

[SWL^{+}15] V. Samedy, M. Wach, J. Lobry, J. Selleslagh, M. Pierre, E. Josse, and P. Boët. Hydroacoustics as a relevant tool to monitor fish dynamics in large estuaries. Fisheries Research, 172(??):225–233, December 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615300345.

Sun:2011:MSB

[SYJ^{+}11] Xiaofeng Sun, Yong Yin, Yicheng Jin, Xinyu Zhang, and Xiufeng Zhang. The modeling of single-boat, mid-water trawl systems for fishing simulation. Fisheries Research, 109(1): 7–15, April 2011. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783610003528.

Sun:2015:AGB

[SYL^{+}15] Chi-Lu Sun, Su-Zan Yeh, Chien-Shan Liu, Nan-Jay Su, and Wei-Chuan Chiang. Age and growth of black marlin (Istiompax indica) off eastern Taiwan. Fisheries Research, 166(??): 4–11, June 2015. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783614002720.

Song:2010:DIH

[SZ10] Liming Song and Yingqi Zhou. Developing an integrated habitat index for bigeye tuna (Thunnus obesus) in the Indian Ocean based on longline fisheries data. Fisheries Research, 105(2):63–74, July 2010. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610000688.

Seung:2011:DSI

[SZ11] Chang Seung and Chang Ik Zhang. Developing socioeconomic indicators for fisheries off Alaska: a multiattribute utility function approach. Fisheries Research, 112(3):117–126, December 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783611001366.

Siddeek:2017:EDW

[SZPP17] M. S. M. Siddeek, J. Zheng, A. E. Punt, and D. Pengilly. Effect of data weighting on the mature male biomass estimate for Alaskan golden king crab. Fisheries Research, 192(??):103–113, August 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617300292.

Siddeek:2016:EST

[SZPV16] M. S. M. Siddeek, J. Zheng, A. E. Punt, and Vicki Vanek. Estimation of size–transition matrices with and without molt probability for Alaska golden king crab using tag– recapture data. Fisheries Research, 180(??):161–168, August 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783615001782.

Storai:2011:BLE

[SZR^{+}11] Tiziano Storai, Luca Zinzula, Simone Repetto, Marco Zuffa, Alexia Morgan, and John Mandelman. Bycatch of large elasmobranchs in the traditional tuna traps (tonnare) of Sardinia from 1990 to 2009. Fisheries Research, 109(1):74–79, April 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783611000373.

Szuwalski:2016:BBE

[Szu16] C. S. Szuwalski. Biases in biomass estimates: the effect of bin width in size-structured stock assessment methods. Fisheries Research, 180(??):169–176, August 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615002003.

Traver:2012:PBD

[TAS12] Michele L. Traver, Larry Alade, and Katherine A. Sosebee. Population biology of a data poor species, offshore hake (Merluccius albidus) in the northwest Atlantic, United States. Fisheries Research, 114(??):42–51, February 2012.

CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783611002736.

Thorson:2010:ESM

[TB10] James T. Thorson and Jim Berkson. Evaluating singleand multi-species procedures to estimate time-varying catchability functional parameters. Fisheries Research, 101(1–2):38–49, January 5, 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783609002483.

Teh:2019:EIG

[TBC^{+}19] Louise S. L. Teh, Natachia Bond, Krishna K. C., Evan Fraser, Ratha Seng, and U. Rashid Sumaila. The economic impact of global change on fishing and non-fishing households in the tonle sap ecosystem, pursat, Cambodia. Fisheries Research, 210(??):71–80, February 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618302741.

Thompson:2016:EEE

[TBF^{+}16] Benjamin S. Thompson, Annabelle J. Bladon, Zubair H. Fahad, Samiul Mohsanin, and Heather J. Koldewey. Evaluation of the ecological effectiveness and social appropriateness of fishing regulations in the Bangladesh sundarbans using a new multi-disciplinary assessment framework. Fisheries Research, 183(??):410–423, November 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616302223.

Taylor:2013:TCS

[TBS13] Matthew D. Taylor, Jessica Baker, and Iain M. Suthers. Tidal currents, sampling effort and baited remote underwater video (BRUV) surveys: Are we drawing the right conclusions? Fisheries Research, 140(??):96–104, February 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783612003578.

Truesdell:2017:EME

[TBSE17] Samuel B. Truesdell, James R. Bence, John M. Syslo, and Mark P. Ebener. Estimating multinomial effective sample size in catch-at-age and catch-at-size models. Fisheries Research, 192(??):66–83, August 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616303563.

Tzeng:2012:DBB

[TC12] Chih-Hsiang Tzeng and Tai-Sheng Chiu. DNA barcodebased identification of commercially caught cutlassfishes (family: Trichiuridae) with a phylogenetic assessment. Fisheries Research, 127–128(??):176–181, September 2012. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783612000537.

Thorson:2015:CCS

[TC15] James T. Thorson and Jason M. Cope. Catch curve stockreduction analysis: an alternative solution to the catch equations. Fisheries Research, 171(??):33–41, November 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783614001507.

Tanaka:2016:CMS

[TC16a] Kisei Tanaka and Yong Chen. Corrigendum to “Modeling spatiotemporal variability of the bioclimate envelope of Homarus americanus in the coastal waters of Maine and New Hampshire” [Fish. Res. 177 (2016) 137–152]. Fisheries Research, 181(??):247, September 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616301114. See [TC16b].

Tanaka:2016:MSV

[TC16b] Kisei Tanaka and Yong Chen. Modeling spatiotemporal variability of the bioclimate envelope of Homarus americanus in the coastal waters of Maine and New Hampshire. Fisheries Research, 177(??):137–152, May 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0165783616300108. See corrigendum [TC16a].

Thorson:2017:UUM

[TC17] James T. Thorson and Jason M. Cope. Uniform, uninformed or misinformed?: the lingering challenge of minimally informative priors in data-limited Bayesian stock assessments. Fisheries Research, 194(??):164–172, October 2017. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783617301546.

Tetzlaff:2011:ETM

[TCAP11] Jakob C. Tetzlaff, Matthew J. Catalano, Micheal S. Allen, and William E. Pine. Evaluation of two methods for indexing fish year-class strength: Catch-curve residuals and cohort method. Fisheries Research, 109(2–3):303–310, May 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783611000889.

Tulk:2017:SSR

[TCBR17] Frederick J. Tulk, Noel G. Cadigan, John Brattey, and Dominique Robert. Spatial synchronicity in recruitment of Atlantic cod (Gadus morhua) stocks off Newfoundland and Labrador and the Flemish Cap. Fisheries Research, 191(??): 49–59, July 2017. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783617300516.

Tobin:2013:IVS

[TCS13] Andrew Tobin, Leanne Currey, and Colin Simpfendorfer. Informing the vulnerability of species to spawning aggregation fishing using commercial catch data. Fisheries Research, 143(??):47–56, June 2013. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613000143.

Tsai:2015:SCS

[TCS^{+}15] Chung-Nan Tsai, Wei-Chuan Chiang, Chi-Lu Sun, Kwang-Tsao Shao, Shu-Ying Chen, and Su-Zan Yeh. Stomach content and stable isotope analysis of sailfish (Istiophorus platypterus) diet in eastern Taiwan waters. Fisheries Research, 166(??):39–46, June 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614003154.

Thoya:2019:EAW

[TD19] Pascal Thoya and Tim M. Daw. Effects of assets and weather on small-scale coastal fishers’ access to space, catches and profits. Fisheries Research, 212(??):146–153, April 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783618303680.

Thorson:2019:SWC

[TDH19] James T. Thorson, Martin W. Dorn, and Owen S. Hamel. Steepness for West Coast rockfishes: Results from a twelveyear experiment in iterative regional meta-analysis. Fisheries Research, 217(??):11–20, September 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618300882.

Tuuli:2016:MIC

[TdMWC16] Cynthia Darta Tuuli, Yvonne Sadovy de Mitcheson, and N. G. Wai-Chuen. Molecular identification of croaker dried swim bladders (maw) on sale in Hong Kong using 16S rRNA nucleotide sequences and implications for conservation. Fisheries Research, 174(??):260–269, February 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783615301156.

Thiel:2012:ABN

[TEFAR12] Martin Thiel, Nuxia L. Espinoza-Fuenzalida, Enzo Acuña, and Marcelo M. Rivadeneira. Annual brood number and breeding periodicity of squat lobsters (Decapoda: Anomura: Galatheidae) from the continental shelf of the SE Pacific — implications for fisheries management. Fisheries Research, 129–130(??):28–37, October 2012. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612001762.

Tuset:2011:EDF

[TEGM^{+}11] Victor M. Tuset, Domingo I. Espinosa, Antonio García-Mederos, José I. Santana, and José A. González. Egg development and fecundity estimation in deep-sea red crab, Chaceon affinis (Geryonidae), off the Canary Islands (NE Atlantic). Fisheries Research, 109(2–3):373–378, May 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783611000920.

Tillett:2012:ASI

[TFB^{+}12] Bree J. Tillett, Iain C. Field, Corey J. A. Bradshaw, Grant Johnson, Rik C. Buckworth, Mark G. Meekan, and Jennifer R. Ovenden. Accuracy of species identification by fisheries observers in a north Australian shark fishery. Fisheries Research, 127–128(??):109–115, September 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612001361.

Tuser:2014:EPB

[TFB^{+}14] Michal Tušer, Jaroslava Frouzová, Helge Balk, Milan Muška, Tomáš Mrkvička, and Jan Kubečka. Evaluation of potential bias in observing fish with a DIDSON acoustic camera. Fisheries Research, 155(??):114–121, July 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614000770.

Takahashi:2015:DTO

[TFH^{+}15] Yuki Takahashi, Yasuzumi Fujimori, Fuxiang Hu, Xiaoli Shen, and Nobuo Kimura. Design of trawl otter boards using computational fluid dynamics. Fisheries Research, 161(??):400–407, January 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614002562.

Trygonis:2016:SDF

[TGDB16] Vasilis Trygonis, Stratis Georgakarakos, Laurent Dagorn, and Patrice Brehmer. Spatiotemporal distribution of fish schools around drifting fish aggregating devices. Fisheries Research, 177(??):39–49, May 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616300133.

Twardek:2018:CCR

[TGE^{+}18] W. M. Twardek, T. O. Gagne, L. K. Elmer, S. J. Cooke, M. C. Beere, and A. J. Danylchuk. Consequences of catch-and-release angling on the physiology, behaviour and survival of wild steelhead Oncorhynchus mykiss in the Bulkley River, British Columbia. Fisheries Research, 206(??):235–246, October 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618301553.

Taylor:2013:SBA

[TGM13] Ian G. Taylor, Vladlena Gertseva, and Sean E. Matson. Spine-based ageing methods in the spiny dogfish shark, Squalus suckleyi: How they measure up. Fisheries Research, 147(??):83–92, October 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613001288.

Taylor:2013:SRR

[TGMM13] Ian G. Taylor, Vladlena Gertseva, Richard D. Methot, Jr., and Mark N. Maunder. A stock–recruitment relationship based on pre-recruit survival, illustrated with application to spiny dogfish shark. Fisheries Research, 142(??):15–21, May 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783612001567.

Tokac:2014:PMC

[THA^{+}14] Adnan Tokaç, Bent Herrmann, Celalettin Aydın, Hakan Kaykaç, Arcan Ünlüler, and Gökhan Gökçe. Predictive models and comparison of the selectivity of standard (T0) and turned mesh (T90) codends for three species in the Eastern Mediterranean. Fisheries Research, 150(??):76–88, February 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783613002427.

Thogersen:2012:LEF

[THF12] Thomas Thøgersen, Ayoe Hoff, and Hans Frost. Linking effort and fishing mortality in a mixed fisheries model: Comparing linear versus non-linear assumptions. Fisheries Research, 127–128(??):9–17, September 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612001440.

Tokac:2016:USS

[THG^{+}16] Adnan Tokaç, Bent Herrmann, Gökhan Gökçe, Ludvig A. Krag, Davod S. Nezhad, Altan Lök, Hakan Kaykaç, Celalettin Aydın, and Ali Ulaş. Understanding the size selectivity of red mullet (Mullus barbatus) in Mediterranean trawl codends: a study based on fish morphology. Fisheries Research, 174(??):81–93, February 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615300783.

Tharwat:2018:BBM

[THHG18] Alaa Tharwat, Ahmed Abdelmonem Hemedan, Aboul Ella Hassanien, and Thomas Gabel. A biometric-based model for fish species classification. Fisheries Research, 204(??):324–336, August 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783618300821.

Tidd:2012:DPE

[THKB12] Alex N. Tidd, Trevor Hutton, Laurence T. Kell, and Julia L. Blanchard. Dynamic prediction of effort reallocation in mixed fisheries. Fisheries Research, 125–126(??):243–253, August 2012. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783612001208.

Thorson:2019:CGD

[Tho19a] James T. Thorson. Corrigendum to “Guidance for decisions using the Vector Autoregressive Spatio-Temporal (VAST) package in stock, ecosystem, habitat and climate assessments” [Fish. Res. 210 (February) 2019, 143–161]. Fisheries Research, 215(??):143–150, July 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619300591. See [Tho19b].

Thorson:2019:GDU

[Tho19b] James T. Thorson. Guidance for decisions using the Vector Autoregressive Spatio-Temporal (VAST) package in stock, ecosystem, habitat and climate assessments. Fisheries Research, 210(??):143–161, February 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0165783618302820. See corrigendum [Tho19a].

Thorson:2019:PLS

[Tho19c] James T. Thorson. Perspective: Let’s simplify stock assessment by replacing tuning algorithms with statistics. Fisheries Research, 217(??):133–139, September 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S016578361830033X. See corrigendum [Tho21].

Thorson:2021:CPL

[Tho21] James T. Thorson. Corrigendum to “Perspective: Let’s simplify stock assessment by replacing tuning algorithms with statistics” [Fish. Res. 219 (2019) 133–139]. Fisheries Research, 235(??):Article 105852, March 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620303696. See [Tho19c].

Torres-Irineo:2014:CFP

[TIGCDL14] Edgar Torres-Irineo, Daniel Gaertner, Emmanuel Chassot, and Michel Dreyfus-León. Changes in fishing power and fishing strategies driven by new technologies: the case of tropical tuna purse seiners in the eastern Atlantic Ocean. Fisheries Research, 155(??):10–19, July 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614000575.

Tang:2014:HBA

[TJJ14] Man Tang, Yan Jiao, and Jess W. Jones. A hierarchical Bayesian approach for estimating freshwater mussel growth based on tag-recapture data. Fisheries Research, 149(??):24–32, January 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783613002154.

Thorson:2017:MBE

[TJMT17] James T. Thorson, Kelli F. Johnson, Richard D. Methot, Jr., and Ian G. Taylor. Model-based estimates of effective sample size in stock assessment models using the Dirichletmultinomial distribution. Fisheries Research, 192(??):84–93, August 2017. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783616301941.

Tuset:2019:OPV

[TJROFS19] Víctor Manuel Tuset, Alba Jurado-Ruzafa, José Luis Otero-Ferrer, and María Teresa G. Santamaría. Otolith phenotypic variability of the blue jack mackerel, Trachurus picturatus, from the Canary Islands (NE Atlantic): Implications in its population dynamic. Fisheries Research, 218(??):48–58, October 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783619301134.

Thorson:2016:IGM

[TK16] James T. Thorson and Kasper Kristensen. Implementing a generic method for bias correction in statistical models using random effects, with spatial and population dynamics examples. Fisheries Research, 175(??):66–74, March 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783615301399.

Tian:2011:IDV

[TKF11] Yongjun Tian, Hideaki Kidokoro, and Tadanori Fujino. Interannual-decadal variability of demersal fish assemblages in the Tsushima warm current region of the Japan Sea: Impacts of climate regime shifts and trawl fisheries with implications for ecosystem-based management. Fisheries Research, 112(3):140–153, December 2011. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611000543.

Thorson:2015:I

[TKO^{+}15] James T. Thorson, Laurence T. Kell, José A. A. De Oliveira, David B. Sampson, and André E. Punt. Introduction. Fisheries Research, 171(??):1–3, November 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615002106.

Truelove:2015:GAR

[TLCSG^{+}15] Nathan K. Truelove, Kim Ley-Cooper, Iris Segura-García, Patricia Briones-Fourzán, Enrique Lozano-Álvarez, Bruce F. Phillips, Stephen J. Box, and Richard F. Preziosi. Genetic analysis reveals temporal population structure in Caribbean spiny lobster (Panulirus argus) within marine protected areas in Mexico. Fisheries Research, 172(??):44–49, December 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783615001794.

Tang:2019:IAS

[TLLZ19] Yanli Tang, Yue Liu, Changdong Liu, and Wuhao Zhang. Improving the accordion-shaped trap selectivity for black rockfish by mounting escape vents: a case study from the small-scale fishery in Shandong, China. Fisheries Research, 219(??):Article 105317, November 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361930164X.

Tyrrell:2011:IIP

[TLM11] M. C. Tyrrell, J. S. Link, and H. Moustahfid. The importance of including predation in fish population models: Implications for biological reference points. Fisheries Research, 108(1):1–8, February 2011. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610003504.

Taylor:2013:HDC

[TM13] Ian G. Taylor and Richard D. Methot, Jr. Hiding or dead? A computationally efficient model of selective fisheries mortality. Fisheries Research, 142(??):75–85, May 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361200269X.

Thorson:2015:PIT

[TMC15] James T. Thorson, Cole C. Monnahan, and Jason M. Cope. The potential impact of time-variation in vital rates on fisheries management targets for marine fishes. Fisheries Research, 169(??):8–17, September 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615001307.

Truesdale:2019:GMC

[TMC19] Corinne L. Truesdale, M. Conor McManus, and Jeremy S. Collie. Growth and molting characteristics of Jonah crab (Cancer borealis) in Rhode Island Sound. Fisheries Research, 211(??):13–20, March 2019. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361830300X.

Tholke:2019:SRM

[TMM19] Emilee K. Tholke, W. David McElroy, and Richard S. McBride. Spermatogenesis, reproductive maturation, and spawning seasonality of male winter flounder, Pseudopleuronectes americanus: Comparisons among fishery stocks. Fisheries Research, 214(??):85–97, June 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619300438.

Thorson:2016:RMC

[TMV16] James T. Thorson and Carolina V. Minte-Vera. Relative magnitude of cohort, age, and year effects on size at age of exploited marine fishes. Fisheries Research, 180(??):45–53, August 2016. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783614003427.

Takasuka:2019:VSA

[TNKO19] Akinori Takasuka, Mitsuo Nyuji, Hiroshi Kuroda, and Yoshioki Oozeki. Variability of swept area by sea-surface tows of a neuston net: Balance of resistance, clogging, and over-inflow effects. Fisheries Research, 210(??):175–180, February 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578361830290X.

Tavares:2012:PSD

[TOA12] Rafael Tavares, Mauricio Ortiz, and Freddy Arocha. Population structure, distribution and relative abundance of the blue shark (Prionace glauca) in the Caribbean Sea and adjacent waters of the North Atlantic. Fisheries Research, 129–130(??):137–152, October 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612002020.

Tanaka:2010:IGO

[TOTI10] Hiroshige Tanaka, Seiji Ohshimo, Nobuo Takagi, and Toshio Ichimaru. Investigation of the geographical origin and migration of anchovy Engraulis japonicus in Tachibana Bay, Japan: a stable isotope approach. Fisheries Research, 102(1–2):217–220, February 2010. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783609003038.

Tenningen:2015:ENV

[TPM15] Maria Tenningen, Héctor Peña, and Gavin J. Macaulay. Estimates of net volume available for fish shoals during commercial mackerel (Scomber scombrus) purse seining. Fisheries Research, 161(??):244–251, January 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614002409.

Tillotson:2017:CCD

[TQ17] Michael D. Tillotson and Thomas P. Quinn. Climate and conspecific density trigger pre-spawning mortality in sockeye salmon (Oncorhynchus nerka). Fisheries Research, 188(??):138–148, April 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783616304325.

Taddese:2019:DRB

[TRC19] Fasil Taddese, Malcolm R. Reid, and Gerard P. Closs. Direct relationship between water and otolith chemistry in juvenile estuarine triplefin Forsterygion nigripenne. Fisheries Research, 211(??):32–39, March 2019. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618303023.

Twist:2016:MPA

[TRH16] Brenton A. Twist, William J. Rayment, and Christopher D. Hepburn. Movement patterns of adult scallops (Pecten novaezealandiae) within a customary fisheries reserve: Implications for fine scale spatial management. Fisheries Research, 174(??):160–166, February 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615300965.

Torres:2012:LWR

[TRS12] M. A. Torres, F. Ramos, and I. Sobrino. Length–weight relationships of 76 fish species from the Gulf of Cadiz (SW Spain). Fisheries Research, 127–128(??):171–175, September 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783612000562.

Thorson:2019:CER

[TRW19] James T. Thorson, Merrill B. Rudd, and Henning Winker. The case for estimating recruitment variation in datamoderate and data-poor age-structured models. Fisheries Research, 217(??):87–97, September 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618301978.

Takahara:2010:IJC

[TS10] H. Takahara and Y. Sakurai. Infection of the Japanese common squid, Todarodes pacificus (Cephalopoda: Ommastrephidae) by larval anisakid nematodes. Fisheries Research, 106(2):156–159, November 2010. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610001128.

Topping:2011:HRM

[TS11] Darin T. Topping and Stephen T. Szedlmayer. Home range and movement patterns of red snapper (Lutjanus campechanus) on artificial reefs. Fisheries Research, 112(1–2):77–84, December 2011. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783611002827.

Tveit:2019:EDT

[TSHB19] Guro Møen Tveit, Manu Sistiaga, Bent Herrmann, and Jesse Brinkhof. External damage to trawl-caught northeast Arctic cod (Gadus morhua): Effect of codend design. Fisheries Research, 214(??):136–147, June 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619300517.

Tascheri:2021:SMS

[TSNRU21] R. Tascheri, J. C. Saavedra-Nievas, and R. Roa-Ureta. Statistical models to standardize catch rates in the multi-species trawl fishery for Patagonian grenadier (Macruronus magellanicus) off Southern Chile. Fisheries Research, 105(3):200– 214, ???? 2021. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578361000127X.

Tenningen:2011:AEN

[TSS11] Maria Tenningen, Aril Slotte, and Dankert Skagen. Abundance estimation of Northeast Atlantic mackerel based on tag recapture data — a useful tool for stock assessment?

Fisheries Research, 107(1–3):68–74, January 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610002572.

Thorson:2014:CPS

[TT14] James T. Thorson and Ian G. Taylor. A comparison of parametric, semi-parametric, and non-parametric approaches to selectivity in age-structured assessment models. Fisheries Research, 158(??):74–83, October 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613002294.

Tolotti:2013:SDC

[TTF^{+}13] Mariana Travassos Tolotti, Paulo Travassos, Flávia Lucena Frédou, Catarina Wor, Humber Agrelli Andrade, and Fábio Hazin. Size, distribution and catch rates of the oceanic whitetip shark caught by the Brazilian tuna longline fleet. Fisheries Research, 143(??):136–142, June 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613000180.

Teodoro:2016:DIJ

[TTMC16] S. S. A. Teodoro, M. Terossi, F. L. Mantelatto, and R. C. Costa. Discordance in the identification of juvenile pink shrimp (Farfantepenaeus brasiliensis and F. paulensis: Family penaeidae): an integrative approach using morphology, morphometry and barcoding. Fisheries Research, 183(??):244–253, November 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616301989.

Teh:2011:QOS

[TTS11] Louise S. L. Teh, Lydia C. L. Teh, and U. Rashid Sumaila. Quantifying the overlooked socio-economic contribution of small-scale fisheries in Sabah, Malaysia. Fisheries Research, 110(3):450–458, August 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611002189.

Tlusty:2016:STA

[TTT16] Michael F. Tlusty, Matthew Thompson, and Heather Tausig. Statistical tools to assess the breadth and depth of shrimp aquaculture certification schemes. Fisheries Research, 182(??):172–176, October 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615301053.

Tunca:2016:BAM

[TÜM^{+}16] Sezgin Tunca, Vahdet Ünal, Bülent Miran, Harun Güçlüsoy, and Ana Gordoa. Biosocioeconomic analysis of marine recreational fisheries: a comparative case study from the Eastern Mediterranean, Turkey. Fisheries Research, 174(??):270–279, February 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783615301211.

Temperoni:2013:FFA

[TV13] B. Temperoni and M. D. Viñas. Food and feeding of Argentine hake (Merluccius hubbsi) larvae in the Patagonian nursery ground. Fisheries Research, 148(??):47–55, November 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783613001896. See corrigen-

dum [TV14].

Temperoni:2014:CFF

[TV14] B. Temperoni and M. D. Viñas. Corrigendum to “Food and feeding of Argentine hake (Merluccius hubbsi) larvae in the Patagonian nursery ground” [Fish. Res. 148 (2013) 47– 55]. Fisheries Research, 151(??):205, March 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613002944. See [TV13].

Tanner:2012:TOG

[TVCT12] Susanne E. Tanner, Rita P. Vasconcelos, Henrique N. Cabral, and Simon R. Thorrold. Testing an otolith geochemistry approach to determine population structure and movements of European hake in the northeast Atlantic Ocean and Mediterranean Sea. Fisheries Research, 125–126(??):198–205, August 2012. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612000938.

Tripp-Valdez:2021:PGS

[TVdLOG^{+}21] Miguel A. Tripp-Valdez, Francisco J. García de León, Sofía Ortega-García, Daniel Lluch-Cota, Juana López-Martínez, and Pedro Cruz. Population genetic structure of dolphinfish (Coryphaena hippurus) in the Gulf of California, using microsatellite loci. Fisheries Research, 105(3):172–177, ???? 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783610000883.

Towler:2010:IMA

[TW10] Richard Towler and Kresimir Williams. An inexpensive millimeter-accuracy electronic length measuring board. Fisheries Research, 106(1):107–111, October 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610001517.

Thorson:2013:AST

[TW13] James T. Thorson and Eric J. Ward. Accounting for space– time interactions in index standardization models. Fisheries Research, 147(??):426–433, October 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613000805.

Thorson:2014:AVE

[TW14] James T. Thorson and Eric J. Ward. Accounting for vessel effects when standardizing catch rates from cooperative surveys. Fisheries Research, 155(??):168–176, July 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783614000836.

Taylor:2015:DPL

[TWCA15] Geraldine C. Taylor, Olaf L. F. Weyl, Paul D. Cowley, and Micheal S. Allen. Dispersal and population-level mortality of Micropterus salmoides associated with catch and release tournament angling in a South African reservoir. Fisheries Research, 162(??):37–42, February 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614002823.

Thorsen:2010:FGA

[TWM^{+}10] Anders Thorsen, Peter R. Witthames, Gudrun Marteinsdóttir, Richard D. M. Nash, and Olav Sigurd Kjesbu. Fecundity and growth of Atlantic cod (Gadus morhua L.) along a latitudinal gradient. Fisheries Research, 104(1–3):45–55, May 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783610000846.

Taylor:2011:IPR

[TWM11] Stephen M. Taylor, James A. C. Webley, and David G. Mayer. Improving the precision of recreational fishing harvest estimates using two-part conditional general linear models. Fisheries Research, 110(3):408–414, August 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783611001901.

Tarnecki:2016:PGM

[TWSA16] Joseph H. Tarnecki, Amy A. Wallace, James D. Simons, and Cameron H. Ainsworth. Progression of a Gulf of Mexico food web supporting Atlantis ecosystem model development. Fisheries Research, 179(??):237–250, July 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616300534.

Tang:2019:EMS

[TXH^{+}19] Hao Tang, Liuxiong Xu, Fuxiang Hu, Taisei Kumazawa, Mamoru Hirayama, Cheng Zhou, Xuefang Wang, and Wei Liu. Effect of mesh size modifications on the sinking performance, geometry and forces acting on model purse seine nets. Fisheries Research, 211(??):158–168, March 2019. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783618303308.

Turra:2016:ARH

[TXP^{+}16] Alexander Turra, Luciana Yokoyama Xavier, Maíra Pombo, Camila de Carvalho Paschoal, and Márcia Regina Denadai. Assessment of recreational harvesting of the trigonal clam Tivela mactroides: Socioeconomic aspects and environmental perception. Fisheries Research, 174(??):58–67, February 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783615300722.

Tang:2017:EEV

[TXZ^{+}17] Hao Tang, Liuxiong Xu, Cheng Zhou, Xuefang Wang, Guoping Zhu, and Fuxiang Hu. The effect of environmental variables, gear design and operational parameters on sinking performance of tuna purse seine setting on free-swimming schools. Fisheries Research, 196(??):151–159, December 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783617302205.

Teixeira:2016:IBR

[TZW16] Daniella Teixeira, Mitchell T. Zischke, and James A. C. Webley. Investigating bias in recreational fishing surveys: Fishers listed in public telephone directories fish similarly to their unlisted counterparts. Fisheries Research, 181(??):127–136, September 2016. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783616301254.

Urquhart:2011:SAS

[UARC11] Julie Urquhart, Tim Acott, Matt Reed, and Paul Courtney. Setting an agenda for social science research in fisheries policy in Northern Europe. Fisheries Research, 108(2–3):240– 247, March 2011. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783610003516.

Uriarte:2012:RES

[UASM12] Andrés Uriarte, Ana Alday, María Santos, and Lorenzo Motos. A re-evaluation of the spawning fraction estimation procedures for Bay of Biscay anchovy, a species with short interspawning intervals. Fisheries Research, 117–118(??):96– 111, April 2012. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783611001111.

Uehara:2019:EDM

[UEOA19] Masato Uehara, Akihiko Ebisawa, Itaru Ohta, and Yoshimasa Aonuma. Effectiveness of deepwater marine protected areas: Implication for Okinawan demersal fisheries management. Fisheries Research, 215(??):123–130, July 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619300839.

Urra:2017:DAD

[UGGR^{+}17] Javier Urra, Teresa García, Helena Gallardo-Roldán, Estefanía León, Matías Lozano, Jorge Baro, and José L. Rueda. Discard analysis and damage assessment in the wedge clam mechanized dredging fisheries of the northern Alboran Sea (W Mediterranean Sea). Fisheries Research, 187(??):58–67, March 2017. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783616303459.

Uematsu:2018:NOI

[UIH^{+}18] Yuki Uematsu, Taiki Ishihara, Yuko Hiraoka, Tamaki Shimose, and Seiji Ohshimo. Natal origin identification of Pacific bluefin tuna (Thunnus orientalis) by vertebral first annulus. Fisheries Research, 199(??):26–31, March 2018. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783617303247.

Uriarte:2019:EDD

[UJLC^{+}19] Amaya Uriarte, Carolina Johnstone, Raul Laiz-Carrión, Alberto García, Joel K. Llopiz, Akihiro Shiroza, Jose M. Quintanilla, Diego Lozano-Peral, Patricia Reglero, and Francisco Alemany. Evidence of density-dependent cannibalism in the diet of wild Atlantic bluefin tuna larvae (Thunnus thynnus) of the Balearic Sea (NW-Mediterranean). Fisheries Research, 212(??):63–71, April 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618303588.

Uchikawa:2014:FHJ

[UK14] Kazuhisa Uchikawa and Hideaki Kidokoro. Feeding habits of juvenile Japanese common squid Todarodes pacificus: Relationship between dietary shift and allometric growth. Fisheries Research, 152(??):29–36, April 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613001653.

Urbach:2012:GMK

[UKK^{+}12] Davnah Urbach, Minho Kang, Sukyung Kang, Ki Baek Seong, Suam Kim, Ulf Dieckmann, and Mikko Heino. Growth and maturation of Korean chum salmon under changing environmental conditions. Fisheries Research, 134–136(??):104–112, December 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612002548.

Ulman:2016:MHC

[UP16] Aylin Ulman and Daniel Pauly. Making history count: the shifting baselines of Turkish fisheries. Fisheries Research, 183(??):74–79, November 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616301503.

Underwood:2016:BBB

[UPS16] Fiona M. Underwood, Graeme Parkes, and Jill H. Swasey. Building Bayesian belief networks to investigate how fishery performance responds to management interventions. Fisheries Research, 182(??):28–38, October 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615301570.

Uphoff:2013:DDA

[USH^{+}13] Christopher S. Uphoff, Casey W. Schoenebeck, W. Wyatt Hoback, Keith D. Koupal, and Kevin L. Pope. Degree-day accumulation influences annual variability in growth of age-0 walleye. Fisheries Research, 147(??):394–398, October 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783613001513.

Volstad:2014:PBS

[VAB^{+}14] Jon Helge Vølstad, Paula Santana Afonso, Ana Paula Baloi, Narci de Premegi, Jørgen Meisfjord, and Massimiliano Cardinale. Probability-based survey to monitor catch and effort in coastal small-scale fisheries. Fisheries Research, 151(??):39–46, March 2014. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613002968.

Vandergoot:2015:APF

[VB15] Christopher S. Vandergoot and Travis O. Brenden. Accuracy and precision of fishery and demographic estimates from a spatial tag-recovery model when inter-regional movements are treated as fixed. Fisheries Research, 164(??):8– 25, April 2015. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783614002987.

Veldhuizen:2018:FWC

[VBdB^{+}18] L. J. L. Veldhuizen, P. B. M. Berentsen, I. J. M. de Boer, J. W. van de Vis, and E. A. M. Bokkers. Fish welfare in capture fisheries: a review of injuries and mortality. Fisheries Research, 204(??):41–48, August 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618300298.

Verhelst:2018:ESE

[VBR^{+}18a] Pieterjan Verhelst, Raf Baeyens, Jan Reubens, Jean-Phillippe Benitez, Johan Coeck, Peter Goethals, Michaël Ovidio, Jenna Vergeynst, Tom Moens, and Ans Mouton. European silver eel (Anguilla anguilla L.) migration behaviour in a highly regulated shipping canal. Fisheries Research, 206(??):176–184, October 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618301504.

Verhelst:2018:DME

[VBR^{+}18b] Pieterjan Verhelst, David Buysse, Jan Reubens, Ine Pauwels, Bart Aelterman, Stijn Van Hoey, Peter Goethals, Johan Coeck, Tom Moens, and Ans Mouton. Downstream migration of European eel (Anguilla anguilla L.) in an anthropogenically regulated freshwater system: Implications for management. Fisheries Research, 199(??):252–262, March 2018. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578361730293X.

Varisco:2019:FRV

[VCS^{+}19] Martín Varisco, Julia Colombo, Pablo Di Salvatore, Pamela Balzi, Nelson Bovcon, Gustavo Lovrich, and Julio Vinuesa. Fisheries-related variations in the fecundity of the southern king crab in Patagonia. Fisheries Research, 218(??):105–111, October 2019. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783619301171.

Vandeperre:2014:DEB

[VdSS^{+}14] Frederic Vandeperre, Alexandre Aires da Silva, Marco Santos, Rogerio Ferreira, Alan B. Bolten, Ricardo Serrao Santos, and Pedro Afonso. Demography and ecology of blue shark (Prionace glauca) in the central North Atlantic. Fisheries Research, 153(??):89–102, May 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614000083.

Volpedo:2010:EPL

[VF10] Alejandra V. Volpedo and Daniela V. Fuchs. Ecomorphological patterns of the lapilli of Paranoplatense Siluriforms (South America). Fisheries Research, 102(1–2):160–165, February 2010. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783609003099.

Valley:2013:RSS

[VF13] Kenneth J. La Valley and Rachel Gallant Feeney. Reconciling spatial scales and stock structures for fisheries science and management. Fisheries Research, 141(??):1–2, April 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783613000660.

Vidal:2010:DOV

[VFdS10] Érica A. G. Vidal, Lidia Fuentes, and Lisiane B. da Silva. Defining Octopus vulgaris populations: a comparative study of the morphology and chromatophore pattern of paralarvae from Northeastern and Southwestern Atlantic. Fisheries Research, 106(2):199–208, November 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361000202X.

Vasconcelos:2017:FRS

[VFFG17] Joana Vasconcelos, Graça Faria, Rita Freitas, and Leonel Serrano Gordo. Fecundity regulation strategy of the blue jack mackerel, Trachurus picturatus (Bowdich, 1825), off Madeira Island (NE Atlantic). Fisheries Research, 190 (??):150–156, June 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783617300462.

Vendeville:2016:REC

[VFMS16] Philippe Vendeville, Hosni Fadhel, Amira Magraoui, and Jacques Sacchi. Restoring the ecosystem creates wealth. The case of the Northern coast of Tunisia’s deep-water rose shrimp trawl fishery. Fisheries Research, 183(??):55–73, November 2016. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783616301321.

Veiga:2011:STH

[VGE11] Pedro Veiga, Jorge M. S. Gonçalves, and Karim Erzini. Short-term hooking mortality of three marine fish species (Sparidae) caught by recreational angling in the south Portugal. Fisheries Research, 108(1):58–64, February 2011. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783610003061.

Vainikka:2012:EES

[VH12] Anssi Vainikka and Pekka Hyvärinen. Ecologically and evolutionarily sustainable fishing of the pikeperch Sander lucioperca: Lake Oulujärvi as an example. Fisheries Research, 113(1):8–20, January 2012. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611002967.

Vasconcelos:2017:UPB

[VHS^{+}17] Joana Vasconcelos, Margarida Hermida, Aurélia Saraiva, José Antonio González, and Leonel Serrano Gordo. The use of parasites as biological tags for stock identification of blue jack mackerel, Trachurus picturatus, in the North-eastern Atlantic. Fisheries Research, 193(??):1–6, September 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783617300826.

Viana:2019:MMS

[Via19] Daniel F. Viana. Management of mobile species with spatial property rights: Challenges and opportunities. Fisheries Research, 218(??):29–34, October 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619301195.

Varela:2017:FHJ

[VIFLP17] José Luis Varela, Ketty María Intriago, Julio César Flores, and Cristhian Ronald Lucas-Pilozo. Feeding habits of juvenile yellowfin tuna (Thunnus albacares) in Ecuadorian waters assessed from stomach content and stable isotope analysis. Fisheries Research, 194(??):89–98, October 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783617301431.

Venerus:2013:ABE

[VIP13] Leonardo A. Venerus, Alejo J. Irigoyen, and Ana M. Parma. Assessment of biases in the estimation of tag shedding rates through a mark-resight experiment. Fisheries Research, 140(??):133–148, February 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613000039.

Vainikka:2017:SDT

[VJH17] Anssi Vainikka, Eglė Jakubavičiūtė, and Pekka Hyvärinen. Synchronous decline of three morphologically distinct whitefish (Coregonus lavaretus) stocks in Lake Oulujärvi with concurrent changes in the fish community. Fisheries Research, 196(??):34–46, December 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617302278.

Vogel:2017:IGS

[VKM^{+}17] Camille Vogel, Dorothée Kopp, Fabien Morandeau, Marie Morfin, and Sonia Méhault. Improving gear selectivity of whiting (Merlangius merlangus) on board French demersal trawlers in the English Channel and North Sea. Fisheries Research, 193(??):207–216, September 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617301182.

Vainikka:2012:CEP

[VKNK12] Anssi Vainikka, Jani Koskimäki, Petri T. Niemelä, and Raine Kortet. Composition of the Eurasian perch (Perca fluviatilis) catches in ice fishing: Does capture order predict body size? Fisheries Research, 115–116(??):24–30, March 2012. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578361100333X.

Verschueren:2019:RCR

[VLSP19] Bart Verschueren, Heleen Lenoir, Maarten Soetaert, and Hans Polet. Revealing the by-catch reducing potential of pulse trawls in the brown shrimp (Crangon crangon) fishery. Fisheries Research, 211(??):191–203, March 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618303230.

Vigier:2018:TSI

[VMB18] Audric Vigier, Stéphanie Mahévas, and Michel Bertignac. Towards a spatial integrated stock assessment model for European hake northern stock. Fisheries Research, 199 (??):158–170, March 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783617303363.

vanMarlen:2014:CCF

[vMWvOKvB14] B. van Marlen, J. A. M. Wiegerinck, E. van Os-Koomen, and E. van Barneveld. Catch comparison of flatfish pulse trawls and a tickler chain beam trawl. Fisheries Research, 151(??):57–69, March 2014. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613002853.

Vasilakopoulos:2012:DIE

[VOM12] P. Vasilakopoulos, F. G. O’Neill, and C. T. Marshall. Differential impacts of exploitation rate and juvenile exploitation on NE Atlantic fish stock dynamics over the past half century. Fisheries Research, 134–136(??):21–28, December 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783612002512.

vanPoorten:2018:EFE

[vPB18] Brett T. van Poorten and Scott Brydle. Estimating fishing effort from remote traffic counters: Opportunities and challenges. Fisheries Research, 204(??):231–238, August 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783618300742.

vanPoorten:2017:RBA

[vPBWA17] Brett T. van Poorten, Boyd Barrett, Carl J. Walters, and Robert N. M. Ahrens. Are removal-based abundance models robust to fish behavior? Fisheries Research, 196(??):160–169, December 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617301650.

Valentin:2014:CMG

[VPC^{+}14] Alexandra E. Valentin, Xavier Penin, Jean-Pierre Chanut, Don Power, and Jean-Marie Sévigny. Combining microsatellites and geometric morphometrics for the study of redfish (Sebastes spp.) population structure in the Northwest Atlantic. Fisheries Research, 154(??):102–119, June 2014. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783614000484.

vanPoorten:2015:IRA

[vPCWV15] Brett T. van Poorten, Thomas R. Carruthers, Hillary G. M. Ward, and Divya A. Varkey. Imputing recreational angling effort from time-lapse cameras using an hierarchical Bayesian model. Fisheries Research, 172(??):265–273, December 2015. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783615300412.

vanPutten:2011:NAR

[vPHG11] Ingrid van Putten, Katell G. Hamon, and Caleb Gardner. Network analysis of a rock lobster quota lease market. Fisheries Research, 107(1–3):122–130, January 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610002742.

vanPutten:2019:FEO

[vPKF^{+}19] Ingrid van Putten, Matt Koopman, Aysha Fleming, Alistair J. Hobday, Ian Knuckey, and Shijie Zhou. Fresh eyes on an old issue: Demand-side barriers to a discard problem. Fisheries Research, 209(??):14–23, January 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361830242X.

Villanueva-Poot:2017:DPT

[VPSH^{+}17] Raúl Villanueva-Poot, Juan Carlos Seijo, Maren Headley, Ana Minerva Arce, Eloy Sosa-Cordero, and Daniel Bernardo Lluch-Cota. Distributional performance of a territorial use rights and co-managed small-scale fishery. Fisheries Research, 194(??):135–145, October 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617301522.

vanPoorten:2016:HCB

[vPW16] Brett T. van Poorten and Carl J. Walters. How can bioenergetics help us predict changes in fish growth patterns? Fisheries Research, 180(??):23–30, August 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615300400.

vanPoorten:2016:PCC

[vPWW16] Brett T. van Poorten, Carl J. Walters, and Hillary G. M. Ward. Predicting changes in the catchability coefficient through effort sorting as less skilled fishers exit the fishery during stock declines. Fisheries Research, 183(??):379–384, November 2016. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783616302120.

Vazquez-Rowe:2010:LCA

[VRMF10] Ian Vázquez-Rowe, Ma. Teresa Moreira, and Gumersindo Feijoo. Life cycle assessment of horse mackerel fisheries in Galicia (NW Spain): Comparative analysis of two major fishing methods. Fisheries Research, 106(3):517–527, December 2010. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578361000247X.

Vazquez-Rowe:2011:LCA

[VRMF11] Ian Vázquez-Rowe, Ma. Teresa Moreira, and Gumersindo Feijoo. Life cycle assessment of fresh hake fillets captured by the Galician fleet in the northern stock. Fisheries Research, 110(1):128–135, June 2011. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0165783611001317. See corrigendum [BMF^{+}10].

Vazquez-Rowe:2012:CLC

[VRMF12] Ian Vázquez-Rowe, Ma. Teresa Moreira, and Gumersindo Feijoo. Corrigendum to: “Life Cycle Assessment of fresh Hake fillets captured by the Galician fleet in the Northern Stock” [Fish. Res. 110 (2011) 128–135]. Fisheries Research, 127–128(??):194, September 2012. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612000513. See [VRMF11].

Vazquez-Rowe:2013:CFA

[VRMF13] Ian Vázquez-Rowe, Ma. Teresa Moreira, and Gumersindo Feijoo. Carbon footprint analysis of goose barnacle (Pollicipes pollicipes) collection on the Galician coast (NW Spain). Fisheries Research, 143(??):191–200, June 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783613000520.

Varela:2013:GGP

[VRS13] Andrea I. Varela, Peter A. Ritchie, and Peter J. Smith. Global genetic population structure in the commercially exploited deep-sea teleost orange roughy (Hoplostethus atlanticus) based on microsatellite DNA analyses. Fisheries Research, 140(??):83–90, February 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612003554.

Villalobos-Rojas:2011:GDC

[VRW11] Fresia Villalobos-Rojas and Ingo S. Wehrtmann. Gonad development in the commercially exploited deepwater shrimp Solenocera agassizii (Decapoda: Solenoceridae) from Pacific Costa Rica, Central America. Fisheries Research, 109(1):150–156, April 2011. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361100052X.

Virbickas:2016:CFC

[VS16] Tomas Virbickas and Saulius Stakėnas. Composition of fish communities and fish-based method for assessment of ecological status of lakes in Lithuania. Fisheries Research, 173(??):70–79, January 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615300618.

Varela:2019:SCS

[VSLC^{+}19] José Luis Varela, Joan Miquel Sorell, Raúl Laiz-Carrión, Ignacio Baro, Amaya Uriarte, David Macías, and Antonio Medina. Stomach content and stable isotope analyses reveal resource partitioning between juvenile bluefin tuna and Atlantic bonito in Alboran (SW Mediterranean). Fisheries Research, 215(??):97–105, July 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619300815.

Varela:2018:NIT

[VSM^{+}18] José Luis Varela, Joan Miquel Sorell, David Macías, Nicolas Goñi, Haritz Arrizabalaga, and Antonio Medina. New insight into the trophic biology of age-0 Atlantic bluefin tuna in the western Mediterranean using stomach content and stable isotope analyses. Fisheries Research, 208(??):274–285, December 2018. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361830225X.

Vila:2010:FDP

[VSTS10] Y. Vila, L. Silva, M. A. Torres, and I. Sobrino. Fishery, distribution pattern and biological aspects of the common European squid Loligo vulgaris in the Gulf of Cadiz. Fisheries Research, 106(2):222–228, November 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610001372.

Veilleux:2011:SGS

[VvHE^{+}11] Heather D. Veilleux, Lynne van Herwerden, Richard D. Evans, Michael J. Travers, and Stephen J. Newman. Strong genetic subdivision generates high genetic variability among eastern and western Australian populations of Lutjanus carponotatus (Richardson). Fisheries Research, 108(1):74–80, February 2011. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783610003115.

Virgili:2018:CUI

[VVL18] Massimo Virgili, Claudio Vasapollo, and Alessandro Lucchetti. Can ultraviolet illumination reduce sea turtle bycatch in Mediterranean set net fisheries? Fisheries Research, 199(??):1–7, March 2018. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361730320X.

Varkey:2018:DGB

[VvPM^{+}18] Divya A. Varkey, Brett van Poorten, Murdoch K. McAllister, Theresa Godin, Eric Parkinson, and Rajeev Kumar. Describing growth based on landscape characteristics and stocking strategies for rainbow trout. Fisheries Research, 208(??):229–238, December 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618301851.

VanDeHey:2014:EGS

[VWHB14] Justin A. VanDeHey, David W. Willis, Jason M. Harris, and Brian G. Blackwell. Effects of gizzard shad introductions on walleye and yellow perch populations in prairie glacial lakes. Fisheries Research, 150(??):49–59, February 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613002397.

Velez-Zuazo:2015:WBS

[VZASM^{+}15] Ximena Velez-Zuazo, Joanna Alfaro-Shigueto, Jeffrey Mangel, Riccardo Papa, and Ingi Agnarsson. What barcode sequencing reveals about the shark fishery in Peru. Fisheries Research, 161(??):34–41, January 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783614001933.

Winter:2015:EIR

[WA15] Andreas Winter and Alexander Arkhipkin. Environmental impacts on recruitment migrations of Patagonian longfin squid (Doryteuthis gahi) in the Falkland Islands with reference to stock assessment. Fisheries Research, 172(??):85–95, December 2015. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783615300163.

Walker:2016:ENA

[WAB^{+}16] Ricardo W. Walker, Neil K. Ashton, Richard S. Brown, Stephanie A. Liss, Alison H. Colotelo, Bernardo V. Beirão, Richard L. Townsend, Z. Daniel Deng, and M. Brad Eppard. Effects of a novel acoustic transmitter on swimming performance and predator avoidance of juvenile Chinook salmon: Determination of a size threshold. Fisheries Research, 176(??):48–54, April 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615301594.

Walden:2013:EHN

[Wal13] John B. Walden. Economic health of the northeast (U.S.) multispecies trawl fleet 1996–2010. Fisheries Research, 139(??):98–104, March 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783612002986.

Ward:2012:BCT

[WAP^{+}12] Hillary G. M. Ward, Paul J. Askey, John R. Post, Divya A. Varkey, and Murdoch K. McAllister. Basin characteristics and temperature improve abundance estimates from standard index netting of rainbow trout (Oncorhynchus mykiss) in small lakes. Fisheries Research, 131–133(??):52–59, November 2012. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783612002172.

Witthames:2013:CDD

[WAT^{+}13] Peter Robin Witthames, Michael Armstrong, Anders Thorsen, Per Solemdal, and Olav Sigurd Kjesbu. Contrasting development and delivery of realised fecundity in Atlantic cod (Gadus morhua) stocks from cold and warm waters. Fisheries Research, 138(??):128–138, February 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783612002640.

Wayte:2013:MII

[Way13] Sally E. Wayte. Management implications of including a climate-induced recruitment shift in the stock assessment for jackass morwong (Nemadactylus macropterus) in southeastern Australia. Fisheries Research, 142(??):47–55, May 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783612002159.

Walsh:2015:BCS

[WB15] William A. Walsh and Jon Brodziak. Billfish CPUE standardization in the Hawaii longline fishery: Model selection and multimodel inference. Fisheries Research, 166(??):151– 162, June 2015. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783614002410.

Was:2016:LTS

[WB16] Anna Was and Rafał Bernaś. Long-term and seasonal genetic differentiation in wild and enhanced stocks of sea trout (Salmo trutta m. trutta L.) from the Vistula River, in the southern Baltic — management implications. Fisheries Research, 175(??):57–65, March 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615301296.

Wurtzell:2016:IBA

[WBB^{+}16] Katharine V. Wurtzell, Adam Baukus, Curtis J. Brown, J. Michael Jech, Andrew J. Pershing, and Graham D. Sherwood. Industry-based acoustic survey of Atlantic herring distribution and spawning dynamics in coastal Maine waters. Fisheries Research, 178(??):71–81, June 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361530134X.

Walker:2013:DUD

[WBD^{+}13] Ricardo W. Walker, Richard S. Brown, Katherine A. Deters, M. Brad Eppard, and Steven J. Cooke. Does UV disinfection compromise sutures? An evaluation of tissue response and suture retention in salmon surgically implanted with transmitters. Fisheries Research, 147(??):32–35, October 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783613001276.

Ward:2016:SBJ

[WBI^{+}16] Tim M. Ward, Owen W. Burnell, Alex Ivey, Stuart C. Sexton, Jonathan Carroll, John Keane, and Jeremy M. Lyle. Spawning biomass of jack mackerel (Trachurus declivis) off eastern Australia: Critical knowledge for managing a controversial fishery. Fisheries Research, 179(??):10–22, July 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783616300273.

Wanke:2017:HSD

[WBM17] Thomas Wanke, Uwe Brämick, and Thomas Mehner. High stock density impairs growth, female condition and fecundity, but not quality of early reproductive stages in vendace (Coregonus albula). Fisheries Research, 186 (part 1)(??):159–167, February 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783616302831.

Wu:2021:PYT

[WCC^{+}21] Georgiana Cho-Chen Wu, Hsin-Chieh Chiang, Yao-Wen Chou, Zong-Ruie Wong, Chien-Chung Hsu, Chiee-Young Chen, and Hsi-Yuan Yang. Phylogeography of yellowfin tuna (Thunnus albacares) in the Western Pacific and the Western Indian Oceans inferred from mitochondrial DNA. Fisheries Research, 105(3):248–253, ???? 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610000792.

Winker:2018:JJA

[WCK18] Henning Winker, Felipe Carvalho, and Maia Kapur. JABBA: Just another Bayesian biomass assessment. Fisheries Research, 204(??):275–288, August 2018. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783618300845.

Webster:2017:DFT

[WCM^{+}17] Fiona J. Webster, Philippa J. Cohen, Siola’a Malimali, Mele Tauati, Karen Vidler, Sione Mailau, Lavinia Vaipuna, and Viliami Fatongiatau. Detecting fisheries trends in a comanaged area in the Kingdom of Tonga. Fisheries Research, 186 (part 1)(??):168–176, February 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616302818.

Warner:2012:RBM

[WCS^{+}12] David M. Warner, Randall M. Claramunt, Jeffrey S. Schaeffer, Daniel L. Yule, Tom R. Hrabik, Bernie Pientka, Lars G. Rudstam, Jeffrey D. Holuszko, and Timothy P. O’Brien. Relationship between mid-water trawling effort and catch composition uncertainty in two large lakes (Huron and Michigan) dominated by alosines, osmerids, and coregonines. Fisheries Research, 123–124(??):62–69, July 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361100350X.

Willems:2016:RBT

[WDBH16] Tomas Willems, Jochen Depestele, Annelies De Backer, and Kris Hostens. Ray bycatch in a tropical shrimp fishery: Do bycatch reduction devices and turtle excluder devices effectively exclude rays? Fisheries Research, 175(??):35–42, March 2016. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783615301326.

Wang:2017:DBF

[WDL17] Hui-Yu Wang, Chieh A. Dong, and Hsiu-Chin Lin. DNA barcoding of fisheries catch to reveal composition and distribution of cutlassfishes along the Taiwan coast. Fisheries Research, 187(??):103–109, March 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616304003.

Witkin:2015:OBF

[WDM15] Taylor Witkin, Sahan T. M. Dissanayake, and Loren Mc-Clenachan. Opportunities and barriers for fisheries diversification: Consumer choice in New England. Fisheries Research, 168(??):56–62, August 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615001095.

Ward:2019:ALT

[WDNB^{+}19] Eric J. Ward, Helena Delgado-Nordmann, Richard E. Brenner, Anne H. Beaudreau, Steven D. Moffitt, and Andrew O. Shelton. Assessing long-term changes in sex ratios of Pacific herring in Prince William Sound, Alaska. Fisheries Research, 211(??):300–308, March 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618302996.

Williams:2011:EIF

[WDS^{+}11] A. Williams, J. Dowdney, A. D. M. Smith, A. J. Hobday, and M. Fuller. Evaluating impacts of fishing on benthic habitats: a risk assessment framework applied to Australian fisheries. Fisheries Research, 112(3):154–167, December 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783611000488.

Waddy:2017:RCM

[WFHG^{+}17] Susan L. Waddy, Nathaniel Feindel, Natalie Hamilton-Gibson, David E. Aiken, Vicky Merritt, and Nicole Leavitt. Reproductive cycles and mating capacity in male American lobsters (Homarus americanus). Fisheries Research, 186 (part 1)(??):358–366, February 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616303277.

Wiedenmann:2019:EPD

[WFJ19] John Wiedenmann, Christopher M. Free, and Olaf P. Jensen. Evaluating the performance of data-limited methods for setting catch targets through application to data-rich stocks: a case study using Northeast U.S. fish stocks. Fisheries Research, 209(??):129–142, January 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618302534.

Wakefield:2011:SNH

[WFLP11] Corey B. Wakefield, David V. Fairclough, Rod C. J. Lenanton, and Ian C. Potter. Spawning and nursery habitat partitioning and movement patterns of Pagrus auratus (Sparidae) on the lower west coast of Australia. Fisheries Research, 109(2–3):243–251, May 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611000798.

Wiff:2018:ESS

[WFNC18] Rodrigo Wiff, Andrés Flores, Sergio Neira, and Bruno Caneco. Estimating steepness of the stock-recruitment relationship in Chilean fish stocks using meta-analysis. Fisheries Research, 200(??):61–67, April 2018. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617303399.

Watson:2019:AEM

[WGC^{+}19] Jabin R. Watson, Harriet R. Goodrich, Rebecca L. Cramp, Matthew A. Gordos, and Craig E. Franklin. Assessment of the effects of microPIT tags on the swimming performance of small-bodied and juvenile fish. Fisheries Research, 218(??):22–28, October 2019. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S016578361930116X. See corrigendum [WGC^{+}21].

Watson:2021:CAE

[WGC^{+}21] Jabin R. Watson, Harriet R. Goodrich, Rebecca L. Cramp, Matthew A. Gordos, and Craig E. Franklin. Corrigendum to “Assessment of the effects of microPIT tags on the swimming performance of small-bodied and juvenile fish” [Fish. Res. 218 October (2019) 22–28]. Fisheries Research, 241(??):Article 106005, September 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621001338. See [WGC^{+}19].

Wenzel:2013:NPR

[WGDB13] Lauren Wenzel, Jordan Gass, Mimi D’Iorio, and Jason Blackburn. A national perspective on the role of Marine Protected Areas in sustaining fisheries. Fisheries Research, 144(??):23–27, July 2013. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612002718.

Wattage:2011:EVC

[WGM^{+}11] P. Wattage, H. Glenn, S. Mardle, T. Van Rensburg, A. Grehan, and N. Foley. Economic value of conserving deep-sea corals in Irish waters: a choice experiment study on marine protected areas. Fisheries Research, 107(1–3):59–67, January 2011. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783610002559.

Williams-Grove:2017:DPT

[WGS17] Laura Jay Williams-Grove and Stephen T. Szedlmayer. Depth preferences and three-dimensional movements of red snapper, Lutjanus campechanus, on an artificial reef in the northern Gulf of Mexico. Fisheries Research, 190(??):61– 70, June 2017. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783617300036.

Weber:2016:CRI

[WHB^{+}16] Michael J. Weber, Matthew J. Hennen, Michael L. Brown, David O. Lucchesi, and Todd R. St. Sauver. Compensatory response of invasive common carp Cyprinus carpio to harvest. Fisheries Research, 179(??):168–178, July 2016. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783616300546.

Watson:2019:RMD

[WHC^{+}19] Fiona M. Watson, Leanne J. Hepburn, Tom Cameron, Will J. F. Le Quesne, and Edward A. Codling. Relative mobility determines the efficacy of MPAs in a two species mixed fishery with conflicting management objectives. Fisheries Research, 219(??):Article 105334, November 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619301894.

Wienbeck:2014:CAL

[WHF^{+}14] Harald Wienbeck, Bent Herrmann, Jordan P. Feekings, Daniel Stepputtis, and Waldemar Moderhak. A comparative analysis of legislated and modified Baltic Sea trawl codends for simultaneously improving the size selection of cod (Gadus morhua) and plaice (Pleuronectes platessa). Fisheries Research, 150(??):28–37, February 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613002348.

White:2010:ASP

[Whi10] J. Wilson White. Adapting the steepness parameter from stock–recruit curves for use in spatially explicit models. Fisheries Research, 102(3):330–334, March 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783609003245.

Wilson:2013:GDM

[WHK13] James Wilson, Anne Hayden, and Michael Kersula. The governance of diverse, multi-scale fisheries in which there is a lot to learn. Fisheries Research, 141(??):24–30, April 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783612001890.

Wienbeck:2011:END

[WHMS11] Harald Wienbeck, Bent Herrmann, Waldemar Moderhak, and Daniel Stepputtis. Effect of netting direction and number of meshes around on size selection in the codend for Baltic cod (Gadus morhua). Fisheries Research, 109(1):80– 88, April 2011. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783611000385.

Windsland:2013:EBG

[WHNS13] Kristin Windsland, Carsten Hvingel, Einar M. Nilssen, and Jan H. Sundet. Evaluation of von Bertalanffy growth curves for the introduced red king crab (Paralithodes camtschaticus) in Norwegian waters. Fisheries Research, 145(??):15– 21, August 2013. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783613000714.

Williams:2015:EIE

[WHP15] Kresimir Williams, John K. Horne, and André E. Punt. Examining influences of environmental, trawl gear, and fish population factors on midwater trawl performance using acoustic methods. Fisheries Research, 164(??):94–101, April 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783614003130.

Williams:2015:CRM

[WHS15] Laura Jay Williams, Jennifer L. Herbig, and Stephen T. Szedlmayer. A cage release method to improve fish tagging studies. Fisheries Research, 172(??):125–129, December 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783615300060.

Watson:2018:VMS

[WHS^{+}18] Jordan T. Watson, Alan C. Haynie, Patrick J. Sullivan, Larry Perruso, Shay O’Farrell, James N. Sanchirico, and Franz J. Mueter. Vessel monitoring systems (VMS) reveal an increase in fishing efficiency following regulatory changes in a demersal longline fishery. Fisheries Research, 207(??):85–94, November 2018. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783618301760.

Woods:2017:BHS

[WJ17] Pamela Woods and Jónas Páll Jonasson. Bayesian hierarchical surplus production model of the common whelk Buccinum undatum in Icelandic waters. Fisheries Research, 194(??):117–128, October 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617301388.

Wysiecki:2017:IED

[WJC17] Agustín María De Wysiecki, Andrés Javier Jaureguizar, and Federico Cortés. The importance of environmental drivers on the narrownose smoothhound shark (Mustelus schmitti) yield in a small-scale gillnet fishery along the southern boundary of the Río de la Plata estuarine area. Fisheries Research, 186 (part 1)(??):345–355, February 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616303393.

Wayte:2010:EHS

[WK10] Sally E. Wayte and Neil L. Klaer. An effective harvest strategy using improved catch-curves. Fisheries Research, 106(3):310–320, December 2010. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610002055.

Winker:2013:CTA

[WKA13] Henning Winker, Sven E. Kerwath, and Colin G. Attwood. Comparison of two approaches to standardize catch-per-unit-effort for targeting behaviour in a multispecies handline fishery. Fisheries Research, 139(??):118–131, March 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783612003311.

Winker:2014:PCN

[WKA14] Henning Winker, Sven E. Kerwath, and Colin G. Attwood. Proof of concept for a novel procedure to standardize multispecies catch and effort data. Fisheries Research, 155(??):149–159, July 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783614000563.

Won:2011:CAH

[WKT^{+}11] Nam-Il Won, Tomohiko Kawamura, Hideki Takami, Hiroshi Hoshikawa, and Yoshiro Watanabe. Comparison of abalone (Haliotis discus hannai) catches in natural habitats affected by different current systems: Implication of climate effects on abalone fishery. Fisheries Research, 110(1):84–91, June 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578361100124X.

Wells:2013:AGN

[WKT^{+}13] R. J. David Wells, Suzanne Kohin, Steven L. H. Teo, Owyn E. Snodgrass, and Koji Uosaki. Age and growth of North Pacific albacore (Thunnus alalunga): Implications for stock assessment. Fisheries Research, 147(??):55–62, October 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783613001331.

Whitten:2013:ACS

[WKTD13] Athol R. Whitten, Neil L. Klaer, Geoffrey N. Tuck, and Robert W. Day. Accounting for cohort-specific variable growth in fisheries stock assessments: a case study from south-eastern Australia. Fisheries Research, 142(??):27–36, May 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783612002068.

Winger:2015:EPB

[WLBB15] Paul D. Winger, George Legge, Christopher Batten, and Georgina Bishop. Evaluating potential biodegradable twines for use in the snow crab fishery off Newfoundland and Labrador. Fisheries Research, 161(??):21–23, January 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783614001957.

Weltersbach:2019:ELB

[WLGS19] Marc Simon Weltersbach, Wolf-Christian Lewin, Joachim Paul Gröger, and Harry Vincent Strehlow. Effect of lure and bait type on catch, size, hooking location, injury and bycatch in the western Baltic Sea recreational cod fishery. Fisheries Research, 210(??):121–130, February 2019. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783618302716.

Wang:2015:MSA

[WLZ^{+}15] Le Wang, Shufang Liu, Zhimeng Zhuang, Haoran Lin, and Zining Meng. Mixed-stock analysis of small yellow croaker Larimichthys polyactis providing implications for stock conservation and management. Fisheries Research, 161(??):86–92, January 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783614001945.

Warden:2011:RPS

[WM11] Melissa L. Warden and Kimberly T. Murray. Reframing protected species interactions with commercial fishing gear: Moving toward estimating the unobservable. Fisheries Research, 110(3):387–390, August 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611002049.

Watson:2013:FDA

[WM13] Reg A. Watson and Telmo Morato. Fishing down the deep: Accounting for within-species changes in depth of fishing. Fisheries Research, 140(??):63–65, February 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612003487.

Wang:2017:WCD

[WM17] Sheng-Ping Wang and Mark N. Maunder. Is down-weighting composition data adequate for dealing with model misspecification, or do we need to fix the model? Fisheries Research, 192(??):41–51, August 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616304143.

Whitlock:2012:EFN

[WMB12] R. E. Whitlock, M. K. McAllister, and B. A. Block. Estimating fishing and natural mortality rates for Pacific bluefin tuna (Thunnus orientalis) using electronic tagging data. Fisheries Research, 119–120(??):115–127, May 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783612000021. See corrigen-

dum [WMB16].

Whitlock:2016:CEF

[WMB16] R. E. Whitlock, M. K. McAllister, and B. A. Block. Corrigendum to “Estimating fishing and natural mortality rates for Pacific bluefin tuna (Thunnus orientalis) using electronic tagging data” [Fish. Res. 119–120C (2012) 115–127]. Fisheries Research, 181(??):248–253, September 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616300662. See [WMB12].

Watson:2016:PGS

[WMC^{+}16] Hayley Victoria Watson, Niall Joseph McKeown, Ilaria Coscia, Emma Wootton, and Joseph Edward Ironside. Population genetic structure of the European lobster (Homarus gammarus) in the Irish Sea and implications for the effectiveness of the first British marine protected area. Fisheries Research, 183(??):287–293, November 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616302041.

Wang:2014:SDP

[WMdS14] Sheng-Ping Wang, Mark N. Maunder, and Alexandre Aires da Silva. Selectivity’s distortion of the production function and its influence on management advice from surplus production models. Fisheries Research, 158(??):181–193, October 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783614000253.

Wuenschel:2013:RBT

[WMF13] M. J. Wuenschel, R. S. McBride, and G. R. Fitzhugh. Relations between total gonad energy and physiological measures of condition in the period leading up to spawning: Results of a laboratory experiment on black sea bass (Centropristis striata). Fisheries Research, 138(??):110–119, February 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783612001683.

Warren-Myers:2018:ESI

[WMIDS18] Fletcher Warren-Myers, Brett A. Ingram, Tim Dempster, and Stephen E. Swearer. Enriched stable isotope marking of hatchery trout via immersion: a method to monitor restocking success. Fisheries Research, 197(??):78–83, January 2018. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578361730259X.

Weigel:2015:AEE

[WMM^{+}15] Jean-Yves Weigel, Pierre Morand, Thuntada Mawongwai, Jean-François Noël, and Ruangrai Tokrishna. Assessing economic effects of a marine protected area on fishing households. a Thai case study. Fisheries Research, 161(??):64–76, January 2015. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783614002094.

Wang:2015:IMM

[WMNC15] Sheng-Ping Wang, Mark N. Maunder, Tom Nishida, and Ying-Ru Chen. Influence of model misspecification, temporal changes, and data weighting in stock assessment models: Application to swordfish (Xiphias gladius) in the Indian Ocean. Fisheries Research, 166(??):119–128, June 2015. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783614002483.

Wang:2014:EVR

[WMP^{+}14] Sheng-Ping Wang, Mark N. Maunder, Kevin R. Piner, Alexandre Aires da Silva, and Hui-Hua Lee. Evaluation of virgin recruitment profiling as a diagnostic for selectivity curve structure in integrated stock assessment models. Fisheries Research, 158(??):158–164, October 2014. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783613003032.

Wilhelm:2017:FGI

[WMPR17] M. R. Wilhelm, C. L. Moloney, S. Paulus, and J-P. Roux. Fast growth inferred for northern Benguela shallowwater hake Merluccius capensis using annual survey- and monthly commercial length-frequency distributions. Fisheries Research, 193(??):7–14, September 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617300619.

Webley:2011:ICI

[WMT11] James A. C. Webley, David G. Mayer, and Stephen M. Taylor. Investigating confidence intervals generated by zeroinflated count models: Implications for fisheries management. Fisheries Research, 110(1):177–182, June 2011. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783611001378.

Wessels:2010:EFM

[WMvdL10] Grea Wessels, Coleen L. Moloney, and Carl D. van der Lingen. The effects of freezing on the morphometrics of sardine Sardinops sagax (Jenyns, 1842). Fisheries Research, 106(3):528–534, December 2010. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610002523.

Wieland:2011:PBE

[WOPB11] Kai Wieland, Hans J. Olesen, Eva Maria Fenger Pedersen, and Jan E. Beyer. Potential bias in estimates of abundance and distribution of North Sea cod (Gadus morhua) due to strong winds prevailing prior or during a survey. Fisheries Research, 110(2):325–330, July 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361100172X.

Wetzel:2011:MPD

[WP11] Chantell R. Wetzel and André E. Punt. Model performance for the determination of appropriate harvest levels in the case of data-poor stocks. Fisheries Research, 110(2):342– 355, July 2011. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783611001731.

Wetzel:2015:EPD

[WP15] Chantel R. Wetzel and André E. Punt. Evaluating the performance of data-moderate and catch-only assessment methods for U.S. west coast groundfish. Fisheries Research, 171(??):170–187, November 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361500185X.

Wetzel:2017:PTO

[WP17] Chantel R. Wetzel and André Punt. The performance and trade-offs of alternative harvest control rules to meet management goals for U.S. west coast flatfish stocks. Fisheries Research, 187(??):139–149, March 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616304040.

Williams:2015:ICP

[WPCO15] Samuel M. Williams, Julian G. Pepperell, Sean W. Corley, and Jennifer R. Ovenden. Isolation and characterisation of 18 polymorphic microsatellite loci for black marlin (Istiompax indica) and their utility for Pacific billfish species. Fisheries Research, 166(??):29–32, June 2015. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783614002197.

Wilber:2019:CSC

[WPS19] Dara H. Wilber, Mark S. Peterson, and William T. Slack. Cross-site comparisons of Gulf sturgeon prey assemblages throughout the northern Gulf of Mexico reveal regional differences. Fisheries Research, 211(??):121–130, March 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783618303059.

Wright:2018:ISS

[WRG^{+}18] Peter J. Wright, Thomas Régnier, Fiona M. Gibb, Julian Augley, and Sandhya Devalla. Identifying stock structuring in the sandeel, Ammodytes marinus, from otolith microchemistry. Fisheries Research, 199(??):19–25, March 2018.

CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783617303235.

Weston:2015:SDS

[WRH^{+}15] Laura F. Weston, Cecile C. Reed, Marc Hendricks, Henning Winker, and Carl D. van der Lingen. Stock discrimination of South African sardine (Sardinops sagax) using a digenean parasite biological tag. Fisheries Research, 164(??):120–129, April 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578361400318X.

Wright:2013:MCE

[Wri13] Peter J. Wright. Methodological challenges to examining the causes of variation in stock reproductive potential. Fisheries Research, 138(??):14–22, February 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612001749.

Wirgin:2018:DPO

[WRMM18] Isaac Wirgin, Nirmal K. Roy, Lorraine Maceda, and Mark T. Mattson. DPS and population origin of subadult Atlantic sturgeon in the Hudson River. Fisheries Research, 207(??):165–170, November 2018. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618301747.

Wells:2010:RVO

[WRP10] R. J. David Wells, Jay R. Rooker, and Eric D. Prince. Regional variation in the otolith chemistry of blue marlin (Makaira nigricans) and white marlin (Tetrapturus albidus) from the western North Atlantic Ocean. Fisheries Research, 106(3):430–435, December 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783610002262.

Wilberg:2013:ELE

[WRR^{+}13] Michael J. Wilberg, Jason M. Robinson, Sarah A. M. Rains, Jennifer L. Humphrey, and Romuald N. Lipcius. Effects of location errors on estimates of dredge catchability from depletion based methods. Fisheries Research, 148(??):1–8, November 2013. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578361300180X.

Watkins:2016:UFV

[WRR^{+}16] J. L. Watkins, K. Reid, D. Ramm, X. Y. Zhao, M. Cox, G. Skaret, S. Fielding, X. L. Wang, and E. Niklitschek. The use of fishing vessels to provide acoustic data on the distribution and abundance of Antarctic krill and other pelagic species. Fisheries Research, 178(??):93–100, June 2016. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783615300229.

Weltersbach:2018:EMP

[WSF^{+}18] Marc Simon Weltersbach, Harry Vincent Strehlow, Keno Ferter, Thomas Klefoth, Martin de Graaf, and Malte Dorow. Estimating and mitigating post-release mortality of European eel by combining citizen science with a catch-and-release angling experiment. Fisheries Research, 201(??):98– 108, May 2018. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783618300171.

Werner:2016:TSP

[WSG16] K.-M. Werner, A. Staby, and A. J. Geffen. Temporal and spatial patterns of reproductive indices of European hake (Merluccius merluccius) in the northern North Sea and Norwegian coastal areas. Fisheries Research, 183(??):200–209, November 2016. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783616301977.

Waterhouse:2014:UAF

[WSMS14] Lynn Waterhouse, David B. Sampson, Mark Maunder, and Brice X. Semmens. Using areas-as-fleets selectivity to model spatial fishing: Asymptotic curves are unlikely under equilibrium conditions. Fisheries Research, 158(??):15–25, October 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783614000174.

Webber:2016:VGA

[WT16] D’Arcy N. Webber and James T. Thorson. Variation in growth among individuals and over time: a case study and simulation experiment involving tagged Antarctic toothfish. Fisheries Research, 180(??):67–76, August 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361530062X.

Weber:2016:EPD

[WTF16] Michael J. Weber, Megan D. Thul, and Mark Flammang. Effectiveness of pulsed direct current at reducing walleye escapement from a simulated reservoir. Fisheries Research, 176(??):15–21, April 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615301442.

Witthames:2010:FVF

[WTK10] Peter Robin Witthames, Anders Thorsen, and Olav Sigurd Kjesbu. The fate of vitellogenic follicles in experimentally monitored Atlantic cod Gadus morhua (L.): Application to stock assessment. Fisheries Research, 104(1–3):27–37, May 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783609003105.

Wild:2017:FTA

[WTS^{+}17] Lauren Wild, Aaron Thode, Janice Straley, Stephen Rhoads, Dan Falvey, and Joseph Liddle. Field trials of an acoustic decoy to attract sperm whales away from commercial longline fishing vessels in western Gulf of Alaska. Fisheries Research, 196(??):141–150, December 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361730231X.

Williams:2016:SCL

[WUGG16] Alan Williams, Judy Upston, Mark Green, and Ken Graham. Selective commercial line fishing and biodiversity conservation co-exist on seamounts in a deepwater marine reserve. Fisheries Research, 183(??):617–624, November 2016.

CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783616300455.

Wor:2018:SRA

[WvPLW18] Catarina Wor, Brett van Poorten, Roberto Licandeo, and Carl J. Walters. Stock reduction analysis using catch-at-length data: Length-SRA. Fisheries Research, 208(??):124–132, December 2018. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783618302091.

Westerberg:2011:POP

[WW11a] Håkan Westerberg and Karin Westerberg. Properties of odour plumes from natural baits. Fisheries Research, 110(3):459–464, August 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783611002190.

Winger:2011:SEU

[WW11b] Paul D. Winger and Philip J. Walsh. Selectivity, efficiency, and underwater observations of modified trap designs for the snow crab (Chionoecetes opilio) fishery in Newfoundland and Labrador. Fisheries Research, 109(1):107–113, April 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783611000440.

Woodcock:2014:TES

[WW14] Skye H. Woodcock and Benjamin D. Walther. Trace elements and stable isotopes in Atlantic tarpon scales reveal movements across estuarine gradients. Fisheries Research, 153(??):9–17, May 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783614000058.

Whitney:2016:NMD

[WWG^{+}16] Nicholas M. Whitney, Connor F. White, Adrian C. Gleiss, Gail D. Schwieterman, Paul Anderson, Robert E. Hueter, and Gregory B. Skomal. A novel method for determining post-release mortality, behavior, and recovery period using acceleration data loggers. Fisheries Research, 183(??):210–221, November 2016. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783616301916.

Williams:2013:WPT

[WWH13] Kresimir Williams, Chris D. Wilson, and John K. Horne. Walleye pollock (Theragra chalcogramma) behavior in midwater trawls. Fisheries Research, 143(??):109–118, June 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783613000209.

Wang:2015:GHL

[WWLY15] Le Wang, Zi Yi Wan, Huan Sein Lim, and Gen Hua Yue. Genetic heterogeneity and local adaptation of Asian seabass across Indonesian Archipelago revealed with gene-associated SNP markers. Fisheries Research, 170(??):205–211, October 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783615001927.

Wakefield:2014:RRM

[WWN^{+}14] Corey B. Wakefield, Ashley J. Williams, Stephen J. Newman, Melanie Bunel, Christopher E. Dowling, Chella A. Armstrong, and Tim J. Langlois. Rapid and reliable multivariate discrimination for two cryptic eteline snappers using otolith morphometry. Fisheries Research, 151(??):100–106, March 2014. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783613002385.

Wakefield:2015:VGL

[WWN^{+}15] Corey B. Wakefield, Ashley J. Williams, Stephen J. Newman, Melanie Bunel, Dion K. Boddington, Elodie Vourey, and David V. Fairclough. Variations in growth, longevity and natural mortality for the protogynous hermaphroditic eightbar grouper Hyporthodus octofasciatus between the Indian and Pacific Oceans. Fisheries Research, 172(??):26–33, December 2015. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783615001988.

Westlake:2018:BRD

[WWR18] Emma L. Westlake, Mike Williams, and Nick Rawlinson. Behavioural responses of draughtboard sharks (Cephaloscyllium laticeps) to rare earth magnets: Implications for shark bycatch management within the Tasmanian southern rock lobster fishery. Fisheries Research, 200(??):84–92, April 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783618300018.

Wiedenmann:2015:AES

[WWSM15] John Wiedenmann, Michael J. Wilberg, Andrea Sylvia, and Thomas J. Miller. Autocorrelated error in stock assessment estimates: Implications for management strategy evaluation. Fisheries Research, 172(??):325–334, December 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783615300461.

Wang:2018:EAS

[WXZ^{+}18] Jing Wang, Binduo Xu, Chongliang Zhang, Ying Xue, Yong Chen, and Yiping Ren. Evaluation of alternative stratifications for a stratified random fishery-independent survey. Fisheries Research, 207(??):150–159, November 2018. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783618301905.

Xavier:2019:EPP

[XdA19] Érika Fialho Morais Xavier and Humber Agrelli de Andrade. Evaluation of the precision of parameters of the Schaefer production model applied to South Atlantic swordfish (Xiphias gladius). Fisheries Research, 211(??):21–26, March 2019. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783618303011.

Xu:2019:MAI

[XDL^{+}19] Lei Xu, Kay Van Damme, Hong Li, Yingying Ji, Xuehui Wang, and Feiyan Du. A molecular approach to the identification of marine fish of the Dongsha Islands (South China Sea). Fisheries Research, 213(??):105–112, May 2019. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619300116.

Xue:2017:EER

[XGT^{+}17] Ying Xue, Lisha Guan, Kisei Tanaka, Zengguang Li, Yong Chen, and Yiping Ren. Evaluating effects of rescaling and weighting data on habitat suitability modeling. Fisheries Research, 188(??):84–94, April 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616304118.

Xu:2019:CSO

[XLCC19] Luoliang Xu, Bai Li, Xinjun Chen, and Yong Chen. A comparative study of observation-error estimators and statespace production models in fisheries assessment and management. Fisheries Research, 219(??):Article 105322, November 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783619301699.

Xu:2019:SDD

[XLCMMV19] Haikun Xu, Cleridy E. Lennert-Cody, Mark N. Maunder, and Carolina V. Minte-Vera. Spatiotemporal dynamics of the dolphin-associated purse-seine fishery for yellowfin tuna (Thunnus albacares) in the eastern Pacific Ocean. Fisheries Research, 213(??):121–131, May 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361930013X.

Xu:2016:UAL

[XTP^{+}16] Yi Xu, Steven L. H. Teo, Kevin R. Piner, Kuo-Shu Chen, and R. J. David Wells. Using an approximate lengthconditional approach to estimate von Bertalanffy growth parameters of North Pacific albacore (Thunnus alalunga). Fisheries Research, 180(??):138–146, August 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615300631.

Xue:2019:CAP

[XYL^{+}19] Dong-Xiu Xue, Qiao-Li Yang, Yu-Long Li, Shao-Bing Zong, Tian-Xiang Gao, and Jin-Xian Liu. Comprehensive assessment of population genetic structure of the overexploited Japanese grenadier anchovy (Coilia nasus): Implications for fisheries management and conservation. Fisheries Research, 213(??):113–120, May 2019. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619300128.

Yandi:2018:ISU

[YA18] Ilhan Yandi and Ilhan Altinok. Irreversible starvation using RNA/DNA on lab-grown larval anchovy, Engraulis encrasicolus, and evaluating starvation in the field-caught larval cohort. Fisheries Research, 201(??):32–37, May 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783618300122.

Yule:2013:UCT

[YAH^{+}13] Daniel L. Yule, Jean V. Adams, Thomas R. Hrabik, Mark R. Vinson, Zebadiah Woiak, and Tyler D. Ahrenstorff. Use of classification trees to apportion single echo detections to species: Application to the pelagic fish community of Lake Superior. Fisheries Research, 140(??):123–132, February 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783612003566.

Yamaguchi:2019:MRD

[YAT^{+}19] Tadanori Yamaguchi, Takako Aketagawa, Katsumi Takayama, Naoki Hirose, and Michiya Matsuyama. Migratory routes of different sized swordtip squid (Uroteuthis edulis) caught in the Tsushima Strait. Fisheries Research, 209(??):24–31, January 2019. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783618302285.

Yu:2018:OWI

[YC18] Wei Yu and Xinjun Chen. Ocean warming-induced rangeshifting of potential habitat for jumbo flying squid Dosidicus gigas in the Southeast Pacific Ocean off Peru. Fisheries Research, 204(??):137–146, August 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618300523.

Ye:2011:UEI

[YCQ11] Yimin Ye, Kevern Cochrane, and Yongsong Qiu. Using ecological indicators in the context of an ecosystem approach to fisheries for data-limited fisheries. Fisheries Research, 112(3):108–116, December 2011. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611002219.

Yu:2016:STD

[YCYC16] Wei Yu, Xinjun Chen, Qian Yi, and Yong Chen. Spatiotemporal distributions and habitat hotspots of the winter– spring cohort of neon flying squid Ommastrephes bartramii in relation to oceanographic conditions in the Northwest Pacific Ocean. Fisheries Research, 175(??):103–115, March 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783615301491.

Yee:2010:VVO

[Yee10] Thomas W. Yee. VGLMs and VGAMs: an overview for applications in fisheries research. Fisheries Research, 101(1–2):116–126, January 5, 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783609002616.

Yergey:2012:EDM

[YGA^{+}12] Matthew E. Yergey, Thomas M. Grothues, Kenneth W. Able, Callie Crawford, and Kevin DeCristofer. Evaluating discard mortality of summer flounder (Paralichthys dentatus) in the commercial trawl fishery: Developing acoustic telemetry techniques. Fisheries Research, 115–116(??):72– 81, March 2012. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783611003389.

Yu:2018:CIH

[YGZ^{+}18] Wei Yu, Ai Guo, Yang Zhang, Xinjun Chen, Weiguo Qian, and Yuesong Li. Climate-induced habitat suitability variations of chub mackerel Scomber japonicus in the East China Sea. Fisheries Research, 207(??):63–73, November 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783618301772.

Yu:2013:PCB

[YJC13] Hao Yu, Yan Jiao, and Laurence W. Carstensen. Performance comparison between spatial interpolation and GLM/ GAM in estimating relative abundance indices through a simulation study. Fisheries Research, 147(??):186–195, October 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783613001549.

Yu:2012:PCT

[YJSR12] Hao Yu, Yan Jiao, Zhenming Su, and Kevin Reid. Performance comparison of traditional sampling designs and adaptive sampling designs for fishery-independent surveys: a simulation study. Fisheries Research, 113(1):173–181, January 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783611003274.

Youcef:2015:VLG

[YLA15] Wahiba Ait Youcef, Yvan Lambert, and Céline Audet. Variations in length and growth of Greenland halibut juveniles in relation to environmental conditions. Fisheries Research, 167(??):38–47, July 2015. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615000260.

Yamashita:2012:CPC

[YMA12] Yukiko Yamashita, Yoshiki Matsushita, and Toru Azuno. Catch performance of coastal squid jigging boats using LED panels in combination with metal halide lamps. Fisheries Research, 113(1):182–189, January 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611003316.

Yochum:2015:EFR

[YRH15] Noëlle Yochum, Craig S. Rose, and Carwyn F. Hammond. Evaluating the flexibility of a reflex action mortality predictor to determine bycatch mortality rates: a case study of Tanner crab (Chionoecetes bairdi) bycaught in Alaska bottom trawls. Fisheries Research, 161(??):226–234, January 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783614002264.

Young:2019:WGD

[YRP^{+}19] H. J. Young, V. Raoult, M. E. Platell, J. E. Williamson, and T. F. Gaston. Within-genus differences in catchability of elasmobranchs during trawling. Fisheries Research, 211(??):141–147, March 2019. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618303278.

Yeon:2011:EBA

[YSS^{+}11] Inja Yeon, Mi Young Song, Myoung Ho Sohn, Hak Jin Hwang, Yang Jae Im, Do Hoon Kim, Hee Won Park, and Chang Ik Zhang. An ecosystem-based assessment of the Korean blue crab trammel net fishery in the Yellow Sea and management implications. Fisheries Research, 112(3):179–188, December 2011. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611002712.

Ye:2012:EIF

[YVJ12] Yimin Ye and John Valbo-Jørgensen. Effects of IUU fishing and stock enhancement on and restoration strategies for the stellate sturgeon fishery in the Caspian Sea. Fisheries Research, 131–133(??):21–29, November 2012. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783612002111.

Yuniarta:2017:UCE

[YvZG^{+}17] Shinta Yuniarta, Paul A. M. van Zwieten, Rolf A. Groeneveld, Sugeng Hari Wisudo, and E. C. van Ierland. Uncertainty in catch and effort data of small- and medium-scale tuna fisheries in Indonesia: Sources, operational causes and magnitude. Fisheries Research, 193(??):173–183, September 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783617301145.

Yates:2019:DAT

[YZW^{+}19] Peter Yates, Philippe Ziegler, Dirk Welsford, Simon Wotherspoon, Paul Burch, and Dale Maschette. Distribution of Antarctic toothfish Dissostichus mawsoni along East Antarctica: Environmental drivers and management implications. Fisheries Research, 219(??):Article 105338, November 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783619301936.

Zenone:2017:CDV

[ZBB17] Aubree M. Zenone, Deron E. Burkepile, and Kevin M. Boswell. A comparison of diver vs. acoustic methodologies for surveying fishes in a shallow water coral reef ecosystem. Fisheries Research, 189(??):62–66, May 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617300073.

Zisserson:2017:IBW

[ZC17] Ben Zisserson and Adam Cook. Impact of bottom water temperature change on the southernmost snow crab fishery in the Atlantic Ocean. Fisheries Research, 195(??):12–18, November 2017. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783617301649.

Zhang:2011:EBR

[ZCC11] Yuying Zhang, Yong Chen, and Yi-Jay Chang. Estimating biological reference points using individual-based per-recruit models for the Gulf of Maine American lobster, Homarus americanus, fishery. Fisheries Research, 108(2–3):385–392, March 2011. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783611000336.

Zhu:2012:IUS

[ZCD^{+}12] Jiangfeng Zhu, Yong Chen, Xiaojie Dai, Shelton J. Harley, Simon D. Hoyle, Mark N. Maunder, and Alexandre M. Aires da Silva. Implications of uncertainty in the spawner– recruitment relationship for fisheries management: an illustration using bigeye tuna (Thunnus obesus) in the eastern Pacific Ocean. Fisheries Research, 119–120(??):89–93, May 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783611003869.

Zhang:2011:RPJ

[ZCLB11] Zane Zhang, Alan Campbell, Dan Leus, and Dominique Bureau. Recruitment patterns and juvenile–adult associations of red sea urchins in three areas of British Columbia. Fisheries Research, 109(2–3):276–284, May 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611000853.

Zhang:2016:EIL

[ZCR16] Chongliang Zhang, Yong Chen, and Yiping Ren. An evaluation of implementing long-term MSY in ecosystembased fisheries management: Incorporating trophic interaction, bycatch and uncertainty. Fisheries Research, 174(??):179–189, February 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783615301041.

Zukowski:2011:UFL

[ZCW11] Sylvia Zukowski, Allan Curtis, and Robyn J. Watts. Using fisher local ecological knowledge to improve management: the Murray crayfish in Australia. Fisheries Research, 110(1):120–127, June 2011. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611001299.

Zhang:2018:EFE

[ZCX^{+}18] Chongliang Zhang, Yong Chen, Binduo Xu, Ying Xue, and Yiping Ren. Evaluating fishing effects on the stability of fish communities using a size-spectrum model. Fisheries Research, 197(??):123–130, January 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617302527.

Zhang:2013:EHC

[ZCZ^{+}13] Yuying Zhang, Yong Chen, Jiangfeng Zhu, Siquan Tian, and Xinjun Chen. Evaluating harvest control rules for bigeye tuna (Thunnus obesus) and yellowfin tuna (Thunnus albacares) fisheries in the Indian Ocean. Fisheries Research, 137(??):1–8, January 2013. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612002597.

Zhang:2013:CBR

[ZD13] Zane Zhang and Jason S. Dunham. Construction of biological reference points for management of the Dungeness crab, Cancer magister, fishery in the Fraser River Delta, British Columbia, Canada. Fisheries Research, 139(??):18– 27, March 2013. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783612003402.

Zwolinski:2019:REE

[ZD19] Juan P. Zwolinski and David A. Demer. Re-evaluation of the environmental dependence of Pacific sardine recruitment. Fisheries Research, 216(??):120–125, August 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619300876.

Zhu:2019:OTD

[ZDZY19] Guoping Zhu, Bo Deng, Haiting Zhang, and Qingyuan Yang. Ontogenetic and temporal diet variation in adult Antarctic krill Euphausia superba at South Georgia during austral winter revealed by stable isotope analysis. Fisheries Research, 215(??):1–8, July 2019. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619300542.

Zischke:2014:TLS

[ZG14] Mitchell T. Zischke and Shane P. Griffiths. Time-location sampling with capture-recapture to assess specialised recreational fisheries. Fisheries Research, 157(??):136–146, September 2014. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783614001490.

Zischke:2012:CES

[ZGT12] Mitchell T. Zischke, Shane P. Griffiths, and Ian R. Tibbetts. Catch and effort from a specialised recreational pelagic sport fishery off eastern Australia. Fisheries Research, 127–128(??):61–72, September 2012. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612001403.

Zhang:2010:UGL

[ZH10] Zane Zhang and John Holmes. Use of generalized linear Bayesian model to mitigate the impact of spatial contraction in fishing pattern on the estimation of relative abundance. Fisheries Research, 106(3):413–419, December 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783610002237.

Zhang:2013:ELS

[Zha13] Zane Zhang. Evaluation of logistic surplus production model through simulations. Fisheries Research, 140(??):36–45, February 2013. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783612003475.

Zhou:2016:ERA

[ZHDS16] Shijie Zhou, Alistair J. Hobday, Cathy M. Dichmont, and Anthony D. M. Smith. Ecological risk assessments for the effects of fishing: a comparison and validation of PSA and SAFE. Fisheries Research, 183(??):518–529, November 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783616302272.

Ziegler:2012:FTF

[Zie12] Philippe E. Ziegler. Fishing tactics and fleet structure of the small-scale coastal scalefish fishery in Tasmania, Australia. Fisheries Research, 134–136(??):52–63, December 2012. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783612002573.

Zischke:2012:RBS

[Zis12] Mitchell T. Zischke. A review of the biology, stock structure, fisheries and status of wahoo (Acanthocybium solandri), with reference to the Pacific Ocean. Fisheries Research, 119–120(??):13–22, May 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783611003699.

Zare:2017:SAT

[ZKSF17] Parviz Zare, Svetlana M. Kasatkina, Sergey V. Shibaev, and Hassan Fazli. In situ acoustic target strength of anchovy kilka (Clupeonella engrauliformis) in the Caspian Sea (Iran). Fisheries Research, 186 (part 1)(??):311–318, February 2017. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783616303381.

Zischke:2016:OMF

[ZLT^{+}16] Mitchell T. Zischke, Lenore Litherland, Benjamin R. Tilyard, Nicholas J. Stratford, Ebony L. Jones, and You-Gan Wang. Otolith morphology of four mackerel species (Scomberomorus spp.) in Australia: Species differentiation and prediction for fisheries monitoring and assessment. Fisheries Research, 176(??):39–47, April 2016. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783615301557.

Zaidman:2015:GVM

[ZM15] Paula C. Zaidman and Enrique Morsan. Growth variability in a metapopulation: the case of the southern geoduck (Panopea abbreviata). Fisheries Research, 172(??):423–431, December 2015. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783615300576.

Zhu:2016:EGW

[ZMdSC16] Jiangfeng Zhu, Mark N. Maunder, Alexandre M. Aires da Silva, and Yong Chen. Estimation of growth within stock synthesis models: Management implications when using length-composition data. Fisheries Research, 180(??):87– 91, August 2016. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783615300953.

Zudaire:2013:FRS

[ZMG^{+}13] Iker Zudaire, Hilario Murua, Maitane Grande, Maria Korta, Haritz Arrizabalaga, Juan Jose Areso, and Alicia Delgado-Molina. Fecundity regulation strategy of the yellowfin tuna (Thunnus albacares) in the Western Indian Ocean. Fisheries Research, 138(??):80–88, February 2013. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783612002366.

Zudaire:2014:AML

[ZMG^{+}14] Iker Zudaire, Hilario Murua, Maitane Grande, Fabrice Pernet, and Nathalie Bodin. Accumulation and mobilization of lipids in relation to reproduction of yellowfin tuna (Thunnus albacares) in the Western Indian Ocean. Fisheries Research, 160(??):50–59, December 2014. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613003044.

Zhang:2017:DBR

[ZQZ^{+}17] Yan-Hong Zhang, Geng Qin, Hui-Xian Zhang, Xin Wang, and Qiang Lin. DNA barcoding reflects the diversity and variety of brooding traits of fish species in the family Syngnathidae along China’s coast. Fisheries Research, 185(??):137–144, January 2017. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783616303083.

Zeller:2011:BSE

[ZRH^{+}11] D. Zeller, P. Rossing, S. Harper, L. Persson, S. Booth, and D. Pauly. The Baltic Sea: Estimates of total fisheries removals 1950–2007. Fisheries Research, 108(2–3):356–363, March 2011. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783611000294.

Zhou:2011:QER

[ZSF11] Shijie Zhou, Anthony D. M. Smith, and Mike Fuller. Quantitative ecological risk assessment for fishing effects on diverse data-poor non-target species in a multi-sector and multigear fishery. Fisheries Research, 112(3):168–178, December 2011. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783610002481.

Zeng:2018:GDE

[ZTC^{+}18] Xu Zeng, Kisei R. Tanaka, Yong Chen, Kai Wang, and Shouyu Zhang. Gillnet data enhance performance of rockfishes habitat suitability index model derived from bottomtrawl survey data: a case study with Sebasticus marmoratus. Fisheries Research, 204(??):189–196, August 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783618300377.

Zajicek:2018:GAS

[ZW18] Petr Zajicek and Christian Wolter. The gain of additional sampling methods for the fish-based assessment of large rivers. Fisheries Research, 197(??):15–24, January 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783617302667.

Zhang:2012:TLC

[ZXH^{+}12] Min Zhang, Congxin Xie, Lars-Anders Hansson, Wanming Hu, and Jiapu Che. Trophic level changes of fishery catches in Lake Chaohu, Anhui Province, China: Trends and causes. Fisheries Research, 131–133(??):15–20, November 2012. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783612001993.

Zhou:2015:KDT

[ZXHK15] Cheng Zhou, Liuxiong Xu, Fuxiang Hu, and Taisei Kumazawa. The kinetic deformation of tuna purse seine: a model experiment on different shooting patterns at uniform current. Fisheries Research, 169(??):18–25, September 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783615001319.

Zhu:2018:PGB

[ZYSZ18] Guoping Zhu, Yang Yang, Qi Song, and Haiting Zhang. Precision of growth band determination from eyestalk sections of Antarctic krill (Euphausia superba) preserved in formalin. Fisheries Research, 197(??):1–6, January 2018. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783617302680.

Zhang:2014:IPS

[ZYW^{+}14] Chi Zhang, Zhenjiang Ye, Rong Wan, Qiuyun Ma, and Zengguang Li. Investigating the population structure of small yellow croaker (Larimichthys polyactis) using internal and external features of otoliths. Fisheries Research, 153(??):41– 47, May 2014. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783613003068.

Zhu:2014:DLH

[ZZA^{+}14] Guoping Zhu, Min Zhang, Julian Ashford, Xiaorong Zou, Xinjun Chen, and Yingqi Zhou. Does life history connectivity explain distributions of Chilean jack mackerel Trachurus murphyi caught in international waters prior to decline of the southeastern Pacific fishery? Fisheries Research, 151(??):20–25, March 2014. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783613002919.

Zhang:2015:HSC

[ZzCbC^{+}15] Jun Zhang, Zuo zhi Chen, Guo bao Chen, Peng Zhang, Yong song Qiu, and Zhuang Yao. Hydroacoustic studies on the commercially important squid Sthenoteuthis oualaniensis in the South China Sea. Fisheries Research, 169(??):45–51, September 2015. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783615001447.